A Complete Bibliography of Publications in the Journal of Theoretical Biology: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

(F M) [LIK^{+}05]. (M, R) [LMM03, LSAA^{+}06]. 16 [Han04]. 2 [KK09, Kin04]. 24 [KG06b]. 3 [Pal08, wTA09, YY07]. ^{+} [GB01]. ^{−} [MS01a]. ^{2+} [BGP00, DBR^{+}05, FB08, FK01a, FWLN04, FKAC06, GBD00, HS07, KBT08, Kar03, MGP^{+}08, MTS05, PAH06, Saf09, WHH07]. _{0} [NM06]. _{1} [NM06]. _{2} [DNS00]. _{Na} [CPM^{+}09]. α [Daw09, Kun05, LKM^{+}08, SHF02, WMP03]. β [BNT^{+}00, Cib08, CPG09, CHN08, DS00a, HHP05, KR05, NS09b, Ped05, Ped07, TPD^{+}00, TSC04, WKG03, vLBJK07]. · [HK05]. d [Ait08]. D_{0} [BB03]. dN/dS [RSH^{+}06]. γ [HHP05, JSA^{+}07, JHJ^{+}09b, RWCK08, SBPC05]. I_{to} [CPM^{+}09]. k [DCL09, KTH09]. κ [HJ06, LPB^{+}04]. λ [BM04, TB04]. μ [SPC02]. N [CGH01, SPS09, BHC06, SA08b]. π [AS07, MSGS08]. pK_{a} [WP01]. Ψ [MPOBD^{+}09]. R^{∗} [JD09]. R^{2} [WL09]. R_{0} [BB03, HS02b, Kao06, KG00, TK09a, TK09b]. σ54 [VdL03]. σ70 [LL06]. t [LLCM01]. × [ANT09, LFFT06, WM00]. → [Leh02].

-activated [HS07]. -Adic [KK09]. -Adrenoceptors [BNT^{+}00]. -Amyloid [WKG03, CPG09]. -Arrestin [CHN08]. -ATP [NM06]. -ATPase [GG07]. -base [YY07]. -binding [Kar03]. -catenin [Daw09, vLBJK07]. -Cell [TPD^{+}00]. -Cells [DS00a, KR05, Ped05, Ped07]. -Core [KTH09]. -D [Kin04, Pal08]. -dependent [VdL03]. -dimensional [wTA09]. -h [KG06b]. -lowering [MBD08]. -nearest [DCL09]. -Opioid [SPC02]. -person [SA08b, SPS09]. -player [BHC06]. -proteobacteria [Kun05]. -regulated [MTS05]. -RNA [MPOBD^{+}09]. -segment [Han04]. -SelectinE [VS08]. -species [CGH01]. -spectrum [WMP03]. -tubulin [Cib08]. -Turn [JSA^{+}07, JHJ^{+}09b].

/Apaf [NS06]. /Apaf-1 [NS06].

015 [ZW07].

1 [BRND09, CMB^{+}01, GSRC^{+}06, KS08a, LGK^{+}09, MW01, MPOBD^{+}09, MGAD09a, MV02a, NS06, PPD09, RG06, RGFP07, Sne03]. 1-Chains [SHF02]. 12 [Kli06, LL06]. 126 [ZZW07]. 15 [WCLL08, WCH^{+}09]. 16mer [RPB03]. 16S [AIK00]. 1D [Cyt04, DECEK06].

2 [FWCN05, GFWT04, MGAD09b, MV02b]. 209 [SI04c]. 212 [FK03a]. 214 [LW04, TT03]. 218 [BSRH03]. 221 [BKE04, TRM03a, TRM03b]. 223 [BBM04]. 224 [Gie06]. 226 [HTN04a, NI04a, UI04b]. 228 [Ano05h]. 231 [HH05b, Ker06, OI05]. 233 [AMD06, BZ05a]. 234 [SHI06b]. 235 [PCZL06]. 236 [GWM06, SRAL12]. 240 [MT06a, NI07a]. 241 [RMAI09]. 242 [AGW^{+}08]. 244 [SML08]. 245 [BLZ07a, ML07a, VGMM^{+}07a, VGMM^{+}08]. 246 [JCRJ07b]. 248 [TQUN08]. 249 [KCP09, Paw09b]. 24h [Cal06]. 251 [LH09, ML08b]. 252 [IvDHI08a]. 253 [MI08a]. 254 [CSP^{+}09, LVL08a]. 256 [Lei10]. 257 [EPJ^{+}11]. 258 [PA09b, RH09a, TK09a]. 259 [ML12]. 260 [DBBW11]. 261 [LGK^{+}12]. 29 [SMPM09]. 2D [GR05, GDPMDS^{+}09, LDXW06]. 2D-visualization [GDPMDS^{+}09]. 2DD [ZC06]. 2DD-curves [ZC06]. 2nd [MMTS02].

3 [MV02c, YCS04]. 3/4 [DRW01]. 3/4-law [DRW01]. 3CL [NLM^{+}08]. 3D [HNTHA07, MLA04, MGDM08, QWQ07, YSW09].

4-law [DRW01]. 4th [MMTS02].

5 [AB07b]. 5’-nucleotides [AB07b]. 5-trisphosphate [LGB03a, TLZ05, WMS08b]. 538 [BLZ07a].

6th [Ano00-30].

A-tracts [AAL08]. ABC [Hop02]. ABC-Type [Hop02]. Abcg2 [IZGG05]. Ability [BB07b, Ger09, HFY07, LdGH09, MH05]. abortion [SK09a]. abrupt [Tas05]. Absence [DPA03, RR08, YD09]. Absorption [FI00, Pic06]. abstained [MM09]. Abstract [OI00, WH07]. Abundance [ML08a, ML08b, AJ05, ANT09, BH08, Chi07, Eng07, PM08, WOLS07]. abutting [BFG08b, BFG08a]. Accelerated [HLW00]. Accelerated-mortality [HLW00]. Accelerating [TMBD09]. accessibility [GKTN07]. accidental [SJ03]. accompanying [Bat06]. account [Ped07]. Accounting [CPR00, Smi08d]. Accretive [KS01a]. Accumulation [AH00, Arc05, Cha01b, Jac03, Kaw01, KK00b, IHV^{+}06]. accuracy [AKLS05, JHJ09a]. accurately [SLH05]. acetylcholine [CRR08]. achieve [Jam08b]. Achieving [Ste00]. Acid [HCMF01, HCB^{+}02, LMT02, MCWF01, AC04, BKD^{+}06, BS05b, BGG^{+}09, CZC05, CC06a, CTZ^{+}06, CL07a, CPG09, Dal06, DCC^{+}08, DCL09, For07a, GKNT09, GKTN07, HCR^{+}04, Hul05, JK04, Kal07, KS04, KGB04, Kur07, LLS^{+}09, Lin08, LF08, MBBR06, RA08, SYC06, SW08c, WBSY06, WYC06, XWC08, YTL08, YZHZ09, hZzGqX^{+}09, ZF08, ZDC08, ZC09, ZCLZ07]. acidic [Di 05]. acidity [SGGM05, SGM08]. Acids [CDDW02, DSY01a, DSY01b, For07b, GKNT09, MPOBD^{+}09, TLZ^{+}08, TGN07, YDL06b, vdGMG^{+}09]. Acinar [FY00, GCB^{+}07]. acorn [NAV04]. Acoustical [SPVN06]. Acoustically [HHR01]. Acquisition [BCV00, KNN01, HML04, YKG^{+}05]. across [AAL08, GINT09, GS08b, Mag04, Sav04, Sel06, TBCD06]. Act [CSM02, Jam01c, Dor03a, VPGA07]. Actin [DECEK06, Wan00, AGW^{+}06, AGW^{+}08, FPBM08, GP08b, MO07, NMH07]. actinomycetes [GT03]. Action [Di 00c, Jac03, Mur07, SV05, BS04, BOCF08, Hop06, LRT04, Mit04, WCLL08, ZKH^{+}05]. Action-response [Di 00c]. activated [HS07, RR09, TSRB08, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, VLFN00]. activating [VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08]. Activation [BSRH02, BSRH03, CMPL^{+}00, GBD00, GF02, SPC02, WKG03, BCL08, CZM09, KSK08, KH09c, LGB03a, LGB03b, MGAR07, PC09b, RK06, RWCK08, Saf05, SF04, SS05a, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, VGMVR^{+}06, vdBR04a]. activation-deactivation [MGAR07]. Activations [CC01c]. Active [Hop02, KRB05a, PTN06, PGH^{+}04, TGR^{+}00, BFGD07, Dim05, HHH06, MS05a, MS06a, MS07b, MS08c, PTON08]. activities [LTG^{+}04, OCA08]. Activity [ABP^{+}00, ARW00, CS00, MM02a, MW01, RWR00, SRMW00, BFG08b, BFG08a, IvDHI08a, IvDHI08b, KNC^{+}04, KL09, LBSP06, LP07b, OR05, Ort06, PRT04, RKRW08, WR07b, dVG04, dVG06]. activity-dependent [RKRW08]. actomyosin [LN05a]. acupuncture [Sil09]. Acute [AM01, CRNP07, DRV^{+}08, DRV^{+}06, KCVC04, LGK^{+}09, LGK^{+}12, RRC^{+}06]. adaptability [ALM09]. Adaptation [DCC^{+}08, FK01b, Gre00, HB02a, HB02b, JWH08, LVH04, LL00, MBP07, SW08a, WT09a, BTA08, DMW04, FK03a, GZL^{+}03, JAA^{+}07, JAJA07, JBJ^{+}08, KGD09, Kon06, Kri09, Mas08, MP09a, Orr06, RAHO06, Wax06]. adaptation-Beyond [Wax06]. adapted [Laz03]. Adaptive [Abr09, Amz04, DHRM08, For01, Gre00, HCC00, HS02a, Man01, Orr03, PN01, AH03b, AA04, Bon06, BHWB05, DMW04, DHP06, GDS07, GI09, GSNH08, Jam06, LAG09, MAC06, RBJ06, Ros05, SSK06, TKN07, TOB08, UD07, WSM05, ZF06]. Addendum [BLZ07a, Gie06]. Adding [CPMG^{+}08]. Addition [SKK^{+}07, HW09]. Additional [DFC^{+}02]. Additive [HL06, LFFT06, NP08, Pec06]. adductors [VAAH05]. Adenine [NTK02]. Adenoma [Pin00]. adenovirus [ZW07]. adherence [KW07]. adherens [Daw09]. Adhesion [Hog00, TMI03a, TM00, TS02, APS06, GC08, KD03, KJD08, Pal08, PDB08, SS06c, WSM05, ZJG03]. adhesions [CSON^{+}05, SMCT08]. Adhesive [LCHK09, vLBJK07]. Adic [KK09]. adipose [KSK08]. Adjacent [RMBM00]. adjust [Cal06]. Adjustment [HSB01, BdOP06]. adjustments [Cib08, DCC^{+}08]. administered [NKL06]. administration [DRV^{+}06, FM09]. Adrenoceptors [BNT^{+}00]. adsorption [Kat03]. adult [WNT08]. adults [SGD04]. Advanced [OS03]. Advantage [Eic01, SI09]. Advantaged [HLW00]. Advantageous [AH00, RS04, BH04]. advantages [CW08b, HMGK06]. advection [FM07a]. advection-diffusion [FM07a]. adversity [WVA05a]. AE1 [RB02a]. Aedes [NBMS06]. Aerodynamic [Sac05, AO05, Iim07]. aerodynamics [USTG09]. aeroelasticity [USTG09]. aeruginosa [RDSB^{+}03]. Affect [CR01, CK08, HDZ^{+}07, MFI09, Pal08, PDM04]. Affected [DLGC02, FPM^{+}06]. affecting [AMC^{+}09, Yam03]. Affects [For00a, BdABA09, RWCK08]. afferent [Sil09]. Affinity [IM02, MH02, Raf02, HBS07, KD03, MPOBD^{+}09, TSK09, VRB01]. Africa [MBH03]. African [EL09]. after [AMP06, BFGS07, Di 06b, FM09, IZGG05, Jam09b, Luc05, MGP00, ØKRG04, SSB^{+}07, TP05]. afterpotentials [Dim05]. Against [ABP^{+}00, CSD09, FMP01, Fra00a, Gar09, HM05, IS09, ITI^{+}09, KSEK09, MFM^{+}03, NGN^{+}04, PBH04, SR08, SI01, SI04c, Wod07]. Agamosperms [NSM02]. Age [Cha01b, TTR00, Tyr01, BSJ04, ELSFB07, Jam08c, LYZ08, PL09]. Age-dependent [Tyr01]. age-infection-structured [LYZ08]. Age-specific [Cha01b]. age-structured [ELSFB07]. Ageing [Cha01b, KK00b, GPB^{+}04, Lan03]. Agent [BGB08, CSR^{+}05, MKLD02, SRAL12, DLM08, GST08, KJ08, Lan03, LK08a, MD03, MDD06, RRH08, Sch08a, SJGK04, WWS^{+}06, ZAD07]. Agent-based [BGB08, CSR^{+}05, MKLD02, SRAL12, DLM08, GST08, KJ08, LK08a, MD03, MDD06, RRH08, Sch08a, SJGK04, ZAD07]. agent-mathematical [WWS^{+}06]. agents [RSC^{+}06]. Agglomeration [MKLD02, CCV08]. Aggregate [Bas01, HTN04a, HTN04b]. Aggregation [GF02, PPL^{+}00, HI05, HMGK06, LdGH09, NWT09, RPNH03, SPVN06, WM04]. aggressiveness [Hur06, NBMS06]. Aging [ABTR07, BTA08, EKIL01, GG01a, Gol08, Jaz01, JW01, Man01, OSCK01, AAK^{+}09, BL08, Hul05, Wod07]. Agonist [KBT08, BCL08]. agonists [HW09]. agroecosystem [CIV09]. Agrotis [RRS^{+}07]. aid [HCC00]. AIDS [GS00, Tho00]. airway [FB08, LWFP08, WTC09]. Ait. [DFSD05]. al [Ril02]. al. [Sch08b]. Alanine [Raf02]. Alanine-scan [Raf02]. Alarm [CMB^{+}01]. albedo [WAL06]. albicans [SSF09]. Algae [TPWG01, BG08]. Algal [Han02, KUK07]. algebra [LS04b, WL04]. Algebraic [AFD^{+}06]. Algorithm [QA01, RPB03, WPE03, XWD^{+}01, CSW^{+}08, DCL09, FWE06, GLSW07, ZLXY08]. Algorithm-based [QA01]. Algorithms [mLLS^{+}06, AHCN07, PLH05, RB08]. alien [YHI07]. aligned [XWD^{+}01]. Alignment [PBMU^{+}09, PAD00, XWD^{+}01, CSON^{+}05, LZ09b, MNL^{+}07]. Alignment-free [PBMU^{+}09]. Alignments [FSS06b, LMT02]. alkylation [Wal07b]. all-alpha [MSGS08]. Allee [BB02, FM02, FR02, MPL06b, vKdRP05]. Alleles [WM00, Wax09]. Allocation [DLGC02, LD00, PG01a, Wah02, EW09, RHF07, SS03, Sat04, YOYT07]. Allometric [Bio08, FI00, LST00, San03, WMLC02, BDMR06, Cud05, Dem06, FDR04, HJ07, Kai04, MGL05, Pac09]. Allometry [Nik02, Cle07, KE09, KSG03, LBCL09, PTD09]. allosteric [RGW^{+}05]. allow [CM09, FM07b]. allowance [LFP^{+}05]. allowing [CSS08, Lit07]. Allows [DBG01]. Almost [MC06]. along [CXM^{+}09, FWLN04, KG02, MCK07, Mur07, QXDW06]. Alpha [RPB03, BGG^{+}09, MSGS08, MRJR09, SLH05, YDL06a]. alpha-deficient [BGG^{+}09]. alpha-helix [YDL06a]. alphabet [LF08]. alphoid [RGPB08]. already [BKKR08]. altering [BL08]. Alternating [Nei01, BC04a]. Alternative [KLK06, SKR06b, SAC^{+}04b, YN02, AWW06, BFA08, CSS06, GMFS06, LGCL07, SH08a]. Altitudes [Kan01]. Altruism [Di 00c, EFW07, FLS01, Gin03, MW00, Pec06, SC05, vVH07, FZ04, FZ07, GA08b, MM03b, NI06, NI07a, Row06, ST05, vV09]. Altruistic [OMT03, JB08, LT09]. Alveolus [Bra01]. Alzheimer [EKS02]. amantadine [BMD^{+}08]. ambiguity [GF09]. ambiguous [PBR01]. Ameboid [CLM07]. Amended [VP01]. Amendments [Luc05]. American [OBPH^{+}08]. Amino [CDDW02, HCMF01, HCB^{+}02, KS04, Kur07, LMT02, CZC05, CC06a, CTZ^{+}06, CL07a, DCC^{+}08, DCL09, GKNT09, GKTN07, HCR^{+}04, KGB04, LLS^{+}09, Lin08, LF08, MBBR06, RA08, SYC06, SW08c, TLZ^{+}08, TGN07, WYC06, XWC08, YDL06b, YTL08, YZHZ09, hZzGqX^{+}09, ZF08, ZDC08, ZCLZ07, vdGMG^{+}09]. Aminoacyl [CDF00, dFG08, AB07b, KS04]. Aminoacyl-tRNA [CDF00, dFG08, KS04]. aminoacylation [JS07]. ammonium [CTS^{+}08]. amoebic [GSNH08, NS07]. among [AAEW09, AH03a, BEK^{+}03, EFW07, Fis03, HL06, Kun05, LL01b, MT06a, MT06b, Orr06, PDC04, SI01, SI04c]. Amount [AGT^{+}01, KMC^{+}07, SH04]. AMPA [Saf05]. amperometric [SN05b]. amphibian [BFGD07]. amphiphilic [ZF08, ZCLZ07]. Amplification [CLB05, For02, Leh02, MB00, KA03a, RCS05]. amplified [LS04a]. amplifiers [GCC08]. Amplifying [Har06]. Amplitude [KG06b, IB06]. amplitudes [PPD09]. Amylase [FY00]. Amyloid [FW00, WKG03, CPG09]. anabolism [Bat09]. Anaerobic [PGLG01, AD07b]. anaesthetics [ZTKH09]. analgesia [Sil09]. analogous [Cud05, Tas05]. analogue [DMW04]. Analyse [CLO^{+}02]. analyser [Ano05h, Hof04]. Analyses [Di 01b, OF01, Bla04, GXG03, OI07, OY03, Pac09, ST05, YAL04]. analysing [YD09]. Analysis [Ace00, AøP03, AGT^{+}01, Ano01b, BZ04, BG07, BWHK02, CDDW02, CT02, CCP^{+}00, CPC^{+}00, CBH02, DFC^{+}02, DVL^{+}00, FTEG02, FGH01, GG01c, GBCC01, HRSV06, Her00, HMN09, HBB08, JMB00, LM01a, LSAA^{+}06, LM01b, LST00, MPD^{+}07, ND04, OR05, PC02, PG00, PKW^{+}00, PGF^{+}08, PV00, ST01, SP00, SML02, SHRR06, SHHD02, SNT03, TML02, TWE04, UL06, VLFN00, VGCGM^{+}02, WKG03, YTAK01, YL00, öOW01, ACD04, AK09, AKdlPP06, BB06a, Bon04b, CSA07, COdAM05, CGS05, CD07, DGZ07, DHW^{+}09, EBId09, FLWB07, GMSW04, GKS06, GZL^{+}03, GPMW06, GBK^{+}07b, GNLEK07, HHP05, HAC^{+}09, HTCS07, Hop06, HGH08, IS03, Iro09, JB06, JWWS08, KR07, KRGH07, KI05a, LN03, LG04, LH06, LTLM09, MD04, ML09a, ML12, MLA04, MB09, MHRK08, MO07, MZK08, MvdO06, MR07b, MRJR09, MNMH^{+}08]. analysis [MGAD09b, MSSS09, NAP04, NBT07, Nar06a, Nie06, NAS07, NLM^{+}08, OST09, Pán08, PP04, PDC04, Paw07b, Paw09b, Pep04a, PWG09, PA09a, PSPF07, PRP^{+}03, QWQ09, RCA09, RSSM06, RRC^{+}06, RAK^{+}08, RDC09, RH08, RLCIB05, RGB06, RAHO06, RBP^{+}09, ST03, SRS09, SPAH06, SF04, Sch08a, SV05, SWRH03, SCS04a, SB07, SW08c, SS09b, ST05, TKKA04, Toy09, TOB08, TBB^{+}06, VGMVR^{+}06, Voi03, WBSY06, Wes03, WC07b, YTL08, YPY^{+}09, YY07, ZSZ^{+}06, ZLN07, ZZLT07, ZLXY08, ZC09, dPCP^{+}08]. Analytic [CT02, DALP03, HC07a]. Analytical [MBB02, WCH^{+}09, FTG07, Gie06, KSB07, MZK08, Moc05, Nis06, PBvdG09, ZAdlPLM07]. analyze [IGHW07]. analyzed [VGS^{+}05]. Analyzing [BdBH09, GCH^{+}07, CEP05]. Anastomoses [GP00, RMBM00]. anatomical [BG06, TPP^{+}04]. Anatomy [AO02, GCYH01]. Ancestor [Di 00b, Di 01a, Di 03b, Di 03a, Di 06b, Lac09]. Ancestral [Möh00]. Ancestry [SS02, Bel06]. Anchorage [FMP01]. ancient [WLZ^{+}06]. Anelosimus [SMD00]. anemia [CSD09]. aneurysm [KH07a]. aneurysms [KH09b]. Angiogenesis [GGV^{+}08, HS00c, ASMM08, BRS^{+}09, JTGP06, LS08a, MAC06, PSJ04, SS05b]. angiogenic [BGB08]. angiopoietins [PSJ04]. angiotypes [Kok04]. angle [Mag04, WS09]. angles [BCV^{+}08, WZ08]. angular [RP09a]. Anhydrase [Tho02, Gam06]. Animal [CKJ^{+}02, LJW02, TT02, TT03, TT01, YTAK01, BCV^{+}08, Ben04b, BWMS07, FHL^{+}06, GAK^{+}06, JBF^{+}03, LNRR06, LAG09, MNL^{+}07, Niw03, Pen03, YM05a, YM05b]. Animals [MGL03, Tay00, BH04, CQLV^{+}03, HNTHA07, MH05, NS03a, OS03, SSD09]. anisogamy [Bon06, Dus06]. Anisotropic [SIMK02, FH07, IHV^{+}06, ZJG03, Zhu09]. anisotropy [LDT09, PPGS03]. ankle [BL08, Mag04]. annealing [FPBM08]. Annihilation [ACK00]. annotated [CB08c]. Announcement [Ano04b, Ano04a, Ano05a, Ano00a, Ano01a]. Annual [Ano00e, CEP05, YFH^{+}07]. Annuals [ARKL02]. Anomalocaris [Usa06]. Ant [CMW02, CB00, DS00b, ORM03b, SFV02, TBS^{+}02, Bur09, JSCN04, JR06, NBT07, RRH08, VTG^{+}06]. Antagonism [SGCK^{+}08, TN01]. antagonist [BCL08, GSG^{+}07]. antagonistic [HK09, UH09]. Antagonists [CD02]. Antarctic [GMMR07]. anterior [MKL09]. Antheraxanthin [LBS00, DVC^{+}04]. anthocyanins [HG08]. Anthozoa [Ber07]. Anti [WJ01, LS08a, RK06, RRC^{+}06, RSC^{+}06, RB08, TMH04, ZJ06]. anti-angiogenesis [LS08a]. anti-herbivore [ZJ06]. anti-inflammation [RRC^{+}06]. anti-inflammatory [RK06]. anti-invasive [RSC^{+}06]. anti-neoplastic [TMH04]. anti-predator [RB08]. Anti-viral [WJ01]. Antiangiogenic [HFH03]. Antibacterial [ABP^{+}00]. Antibiotic [DPA05, DMO^{+}07, HML09, MWD08]. antibiotic-resistance [MWD08]. Antibodies [ISC01, TGR^{+}00, AMP06, NKL06]. Antibody [GG03a, PD01, AB06, BSS^{+}07, WM04]. antibody-dependent [AB06, BSS^{+}07]. anticancer [BFGS07]. anticipated [ISW04]. anticipation [SW08a]. anticodon [Tak06]. antifolate [AIKP06]. Antigen [ABP^{+}03, FFD^{+}02, SBZ^{+}08, vdBR03, AMP06, ACSY04, NA04, vdBWLS07]. Antigen-driven [FFD^{+}02, SBZ^{+}08]. Antigen-independent [ABP^{+}03, ACSY04]. antigenic [JG08, RG05]. antigens [CTB^{+}05]. Antimicrobial [Ano01e, Mac01, AAEW09]. Antioxidants [Sta00c]. antiparallel [Kur08b]. Antipredator [PBR03, BBD06, PBR01]. Antipredatory [MKN02]. antiretroviral [DP04, GSRC^{+}06, PPD09, Smi08d]. Antisense [FC01, Dor03a]. antithrombin [ARR08]. Antiviral [AN00, HLA09, ABvdD^{+}08]. Ants [BZ00, IST01, BGD^{+}06, Roe08]. any [Di 01b]. AP [MW01]. AP-1 [MW01]. Apaf-1 [NS06]. ape [GR05]. Aperiodic [CR02]. apex [SP07b]. aphid [AK04]. aphids [Roe08]. Apical [PKL02]. Apis [GHHR03]. Apomixis [RKNK01]. Apoptosis [PN02b, Ark05, CL07a, CL07b, ML09a, ML12, OAKC08, ZLLZ09]. apoptosome [NS06]. Aposematic [LY01, MR07a, LY03a]. Apparent [Kin04, KMGDG04, WS04, BBB^{+}07, SH03, SW08a, SB07]. appearance [KCV05]. appendix [BBB^{+}07]. Appetitive [RRS^{+}07]. applicability [WBH04]. Applicable [KK00a, Bat09]. Application [BCRG04, BGG^{+}09, CRJC04, FLWB07, GLE^{+}09, GG01c, HNTHA07, LF01, LMF03, LL01a, LSMLB^{+}02, Mag04, MCWF01, MSI01, NM06, NLS08, OSCK01, RSD^{+}01, RC00, SZLK^{+}05, SP07b, TH03a, TRM03c, WP01, YDL06a, AB06, AAK^{+}09, AB03, BC09a, BKM09, BB06b, CLMMP06, CBB07, CLM07, DFCL08, DFSD05, GMK06, GSB05, HC07b, JDMZ^{+}07, LMH04, LDXW06, MRA06, OS03, PBB03, Rej07, RRK06, RGRB04, RBP^{+}09, RB08, TLZ^{+}08, TRM03a, TRM03b, XSD^{+}05, XFBC07, Mei05]. Applications [FSS06b, KSM02, Ano05h, BZ04, Hof04, LS03, LS04a, LP07c, MH09a, YSW09, dPGR06]. Applied [Ano02a, VLFN00, HTCS07, LJ09b, Lit07, TBR08, Zam03]. Applying [NGTB06, CKE06, HK07]. appressorial [TGT06]. Approach [AMW00, CS02, GS02b, HS02b, HA00, Kal00, PH03, TLA00, XD02, ZSZ01, AJOK09, AM08, APS06, AB03, BBKGP08, BZM05, BMR08, BWMS07, BCRG04, BRC^{+}03, BTA08, CL07a, CCC08, DT07a, DCP^{+}08, DFSD05, FTG07, FCP03, FSS06a, GB07, Gro04, GST08, HI05, HF03a, HQP^{+}09, IGHW07, JB04, JB06, JWB^{+}09, KSB07, KYZ^{+}08, Kau04, KMT06, KOT07, LM09, LRB^{+}06, LS04b, LJTD05, LKW06, Lio09, LP07b, LSD^{+}07, MH06a, MCK07, MBBR06, MNL^{+}07, MI08a, MI08b, MSSS09, MV07, Ots08, PT09, RKF06, RR09, RRS06, SMH07, STK08, SJD^{+}09, TPP^{+}04, Twa04, Ver04, VN08, WT05, XBCF05, XWC08, YLCW06, YZGW06, ZSRB07, hZzGqX^{+}09]. Approached [QLHL09]. Approaches [Jaz01, HK07, HLP06, JAHKH09, LJ09b, PHdB09, ZW07]. appropriate [DBBC08]. Approximate [ZDC08, GSB05, KSB07]. approximating [BGG06]. Approximation [Ell01, LS04a, TE04, TE05]. approximations [LS03, LD09a, PD04, Sav04, XFBC07]. April [Ano00q, Ano00u, Ano01v, Ano01-27, Ano02p, Ano02t, Ano03-36, Ano03-40, Ano04-46, Ano04-48, Ano05-46, Ano05-52, Ano06-43, Ano06-45, Ano07-44, Ano07-53, Ano08-43, Ano08-47, Ano09-40, Ano09-47]. aqueous [VS08]. aquifer [RDL07]. Arabidopsis [Kor09, LMT05, MAB00, Mei05]. arbitrary [ANT09]. arbors [RKH^{+}06]. Arbuscular [PBT02]. ARC [PCB07]. archaeal [FM03]. Archaeological [OLS^{+}02]. archaic [MB09]. Architectural [Whi05]. Architecture [ORM03a, SMG^{+}03, WM00, Bon04a, Bon04b, DFSD05, HML04, PRT04, PPRI08]. Area [Ant02, CC01a, Eng07, LJ06, PC09a, PPL^{+}00]. areas [AKS07]. arginines [Gre05]. arid [GJ07]. Arino [Kim07]. arise [MGDDH08, SC05]. arising [CKS07]. Aristotelian [LCL03]. Arithmetic [Gin00a]. Arms [WBB02, Cib08]. ARO [SMPM09]. arousal [PR08]. arrangement [Kun06]. Arrangements [Kun01, FSS06b, Kun03]. array [YYA09]. Arrestin [CHN08]. Arrhythmia [öWO01, ZTKH09]. Arrhythmias [CBBH01]. arrival [GBB08, LCL07a]. Arrivals [BCV00]. Art [Ost04, Rot04, Win04]. artefacts [HH04b]. Arterial [BE03, GBLV03, SKN^{+}03, DWBB04, Hol06]. arterio [WBR09]. arterio-venous [WBR09]. arteriolar [BFG08b, BFG08a]. Artery [RMBM00]. Artery/Graft [RMBM00]. arthropod [CGKC07]. arthropods [MRA06]. Arthur [TG04]. article [Ort06]. articular [BM06]. artifacts [CdOWS04]. Artificial [Bok06, HAG03, JST^{+}02, HDW^{+}09, LWRK07, SDBH04, VTG^{+}06, VN08]. Artist [Win04]. ascending [Ait08, PR08]. Ascertained [Jam00]. Asexual [CMB02, DEM^{+}00b, PDC02, Fuk04, LT09]. aspartate [LZS^{+}08]. aspartic [VGMVR^{+}06]. Aspects [WRF01, PW09b, SML07, SML08, dSKL09]. aspen [PBB03]. aspiration [BBSLN08]. assay [CLH07, Hua03]. assays [AC04]. Assemblages [Bon03]. Assembling [MCWF01, MPD^{+}07]. Assembly [GAA02, Vei03, AJSL07, BLMV05b, KMT06, KKR^{+}07, NS06, Paw09a, Twa04]. assess [RHF07]. Assessing [CAHH06, FMF^{+}00, KLL08, LPJB^{+}08, LGS^{+}09, NLS08]. Assessment [Hay00, SP00, Arc07a, BS04, BFA08, Gos06, QSAL08, YHI03]. Assessments [Pel00]. assignment [WT09b]. Assimilate [TH03a]. Assimilation [KRNKH00, EBId09, RH09a, RH09b]. Associated [Che00, FW00, Lan00a, LY01, ASMM08, BL08, GKM^{+}00, Gie03, HSK^{+}08, KLK06, OR05, SFC^{+}09, TMH04, WR07a, YKG^{+}05]. association [GFWT04, Hui09]. Associations [Jam01c, SCH05b, SCS07]. Associative [GAT07, WS02a]. Assortative [CT06, RC09]. assuming [Tak06]. assumption [Bea06, EAM07, HE08a, MHKS03]. Assumptions [HS01, CZM09, Dus06, MGDDH08, SG07, Win06]. Assurance [FSL03]. Asthma [Sch02a, HSK^{+}08]. astrocytes [BFG08b, BFG08a, BFG09, LH08, LH09]. astrocytic [SM06b]. Asymmetric [BS01, FK03b, HG02, HGC03, LL08, OM08, PNG03, Yan05, YM05b, ZWT^{+}08]. asymmetries [ZXWF08]. Asymmetry [ES01a, Laz02, MC01, PKL01, FKS07, HF03a, PCH^{+}05]. Asymptotic [GL09, MGAR07, Zah00]. asymptotics [Sim08]. asynchronous [BCRG04, TC09]. asynchronously [LFC04]. AT2 [Plu06]. AT3 [Plu06]. AT3-profiling [Plu06]. Atherosclerosis [Gam06]. atherosclerotic [ZHB04]. athletics [GMMGDA05]. Atlantic [GNLEK07, LS08b]. Atmospheric [SM01, Wal00, KRNG08]. atom [HK05]. ATP [BPLS06, MS06a, NM06, NHTM09, SM06b, WHH07]. ATP-dependent [MS06a]. ATP-induced [WHH07]. ATP-mediated [SM06b]. ATPase [AGT^{+}01, FKAC06, GG07]. ATPases [NM08]. atrial [CE08, RK05]. atrioventricular [RK05]. attachment [MD04, SPG06]. Attack [Fra00a, Fra02, PKW^{+}00, VMW01, KYS06, LPC06, MW09]. Attacks [Sch02a, Lin04]. Attempt [Eic01]. attention [LAG09]. attenuation [CPG09, MKA05, SZ04]. attract [KYS06]. attraction [Bye09, SRS09]. attractive [CMG06, RRH08]. attractiveness [CTB09]. Attractor [MD01, WS02a, Hav04]. attractors [RK07]. Attrition [HG02, ELL04, HGC03]. ATW [Ano04c]. auditory [SP05a, SGTF07]. Augmented [NSS^{+}08a, hZzGqX^{+}09]. August [Ano00w, Ano00-27, Ano01t, Ano01y, Ano02q, Ano02w, Ano03-32, Ano03-37, Ano04-43, Ano04-47, Ano05-48, Ano05-51, Ano06-46, Ano06-51, Ano07-46, Ano07-52, Ano08-45, Ano08-50, Ano09-42, Ano09-48]. Aunt [GK00]. aureus [ABP^{+}00, MPM07, PWVvR09, MWD08]. Author [Ano03a, Ano03b, Ano03c, Ano04h, Ano04i, Ano04d, Ano04e, Ano04f, Ano04g, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d]. autism [KV05]. auto [Tao04a, XT06, hZzGqX^{+}09]. auto-regulatory [Tao04a, XT06]. autocatalysis [MS05b]. autocatalytic [HS04]. autoimmune [ITM^{+}07, ITI^{+}09]. Autoimmunity [NGN^{+}04]. Automata [KCT02, MD06, STK08]. Automatic [XWD^{+}01]. Automaton [CMB02, HBLG02, KTH^{+}00, PGLG01, ABM03, BMH07, BST05, GA07, GA08a, GFWT04, PA09b, PA09c, XWC08]. automimicry [LY03a]. autonomous [Cin03, MC06]. autonomy [BFK08]. Autoparasitoids [SMG01]. Autopoietic [LMM03]. Autoregressive [CPR00]. autoregulation [TIM06]. Autotoxicity [CEP05, Sin07]. autotroph [MRC09]. autotroph-based [MRC09]. Autumn [Arc00, Arc07a, Arc07b, Arc09a]. Auxin [Kra01c, FMI05, FPM^{+}06, FM06]. Auxin-mediated [Kra01c]. Availability [TTR00, AKS07, EW09, MFI09]. Avalanches [NL03b, RKYH06, SVGK07]. avascular [RSC^{+}06]. average [DCC^{+}08, PL09, WN05]. Avian [ITL09, FS06, ITKL08, ITLN09, JITJ09, MBB^{+}06, PB07]. avidity [KSEK09]. avoid [KY03, PBC09b]. avoidance [Yam03]. away [Akt04]. Axially [KRNKH00]. Axillary [GBCC01, GHC03]. Axis [Yos03]. Axisymmetric [GH02]. axon [LRT04]. axonal [ML09b, OLBM08]. Axoneme [DF00, Cib08]. axons [CBF05].

B [TL00, LIK^{+}05, AMP06, CRNP07, DFP01, HJ06, Jam07, JAFW05, KD03, LPB^{+}04, MH07, NA04, TBR08]. B-DNA [TL00]. Baby [GCYH01, KS03]. Bacillariophyta [SWM09]. Bacillus [KiHM^{+}03, MBP07, AIK00]. Back [MD01]. backbone [WZ08]. Background [HH05a, MSDM06]. backness [Nar06b]. Backward [RMP08, CMW08]. Bacteria [Di 03a, AAEW09, Amz04, AR08, BKP09, EHG03, HFS06, JG08, Joh08, KNT^{+}09, Kun03, MRC09, RAZ03, ROR05]. bacteria-derived [JG08]. Bacterial [AMW00, Ats01, BP01, GLV02, GEK04, NP07, RAZ03, TGP^{+}00, CRL06, Cog06, DRV^{+}08, GASA09, Kun06, Luc05, MX08, Mig06, PLG^{+}06, PGF^{+}08, RZF03, RSH^{+}06, SCS07, YMLK04]. Bacteriophage [AC07a, PJ01, RFH^{+}02, AC07b, BM04, KT07, PBHS05, TB04]. Bacteriophages [Wil01, AR08]. bacterium [SRN07]. bad [BS06, Yan09]. Balance [CD02, CSP01, FLBB01, LGB02, The00, AF09, BL08, BGG^{+}09, Fuk04, HH07, HTCS07, Kel07, MGM07, MS08b, MSSS09, SS06c, SHRR06, SS09b]. Balanced [KW00a, PP04]. ballistic [RB09a]. Banavar [MGL05, MGL06a]. Banding [BPZ^{+}01]. bands [GBZ06, GG09]. Bank [ARKL02]. Banker [ARKL02]. barbata [See00]. bare [For09]. bargaining [dSKL09]. Barker [Jam01b]. barking [SG07]. barley [OF05]. barley-malting [OF05]. barnacles [YOYT07, YYY^{+}08]. Baron [KV05]. Barr [CMB^{+}01, SDL^{+}08]. barriers [BS05b, MS06a]. basal [vAR09, vAGDR09]. Base [OKS01, For07b, JS07, LRD04, Mei05, OS03, YY07, ZXWF08]. base-pattern [LRD04]. base-specific [Mei05]. Based [AøP03, Ano09a, CSFH^{+}01, Cum00, GBG01, GS02b, GRG02, HGV01, IST01, ISWT02, PDA^{+}00, PBR03, SK01, Wal00, AM09, AWO^{+}09, AV05, BFH^{+}01, BKMH07, Ben03, BGB08, BBM03, BBM04, BRS^{+}09, Bon04a, BSJ04, CSR^{+}05, CC06b, COdAM05, CCZC08, CP03, CSM05, DGM05, DLM08, Di 01b, Dic08, DFCL08, DFSD05, DBF07, FM07a, FmW08, FVP^{+}07, FBM06, FCD^{+}05, FP04b, Fra03, FH07, GA09, GLV02, GKK06, GG09, Gro02, GST08, GNLEK07, Han04, HK07, HP04, JAFW05, JAV07, KYZ^{+}08, KG00, KR07, KD03, KRB05a, KJ08, KSS07, KLL08, KSR07, LK08a, LH06, Lio09, LW08b, LZ09b, ML09a, ML12, MKLD02, MD03, MDD06, MW01, MSS05, MS06a, MRC09, MB07, MLWL06, MD08, MM00b, MGDM08, MGDBdM08, NN07, NBT07, Nar06b, NA02, PWZ09, PBMU^{+}09, PWG09, PR08, PLG^{+}06, PRP^{+}03, QWQ07, QA01]. based [RRH08, RWK08, RLCIB05, SK09a, SWN07, Sch08a, SI06, SCH05b, SMPM09, SJGK04, STK08, SKK^{+}07, Sta09, SRAL12, Tak06, TSZ^{+}07, TGV^{+}07, UM08, Usa06, VGS^{+}05, VRB01, VCGV07, VN08, Wal08, WYC06, WZ08, WM04, WC07b, YDL06b, YPY^{+}09, YAL04, YSW09, hZzGqX^{+}09, ZC06, ZAD07, ZF08, ZLXY08, ZYD^{+}05, ZT06]. Baseline [BB03]. bases [MA07b, Wal07b]. Basic [Dus01, MH09a, Now00, CHCC^{+}04, IMN04a, NI07b, RR08, Ste09, Yan08]. Basis [ADMZ02, HFH03, PGF^{+}08, PD00, RB02c, TGR^{+}00, øPVO02, AC04, JB04, TK05]. Basket [RB02a]. bass [CH07]. bat [DHR^{+}07, RWID^{+}08]. batch [LTLM09]. bathtub [BLZ07a, BLZ07b]. bats [DHRM08]. Bayesian [BMR08, BE08, Ded09, GRBR04, HTCS07, HLP06, LD09a, RGRB04, VTL05, Van06a]. BCG [Ger09]. Bcrp1 [IZGG05]. Bcrp1/Abcg2 [IZGG05]. Bdellovibrios [Wil01]. BDI [JST^{+}08]. BDI-modelling [JST^{+}08]. be [BS01, BH03b, Buc04, Di 01b, EFW07, Kom07, KV04, MH06b, MS03b, Par04, RH04, SR08, THKU07, WFGP04, vV06, vV09]. bean [FHL^{+}06]. Bearing [GB02]. Beating [DF00]. Beautiful [Kan07]. because [Di 01b, For00a]. becomes [BM09]. Beddington [LWC09]. Beddington-DeAnglis [LWC09]. Bee [SB00, AO03b, CM03, GHHR03, HD08]. been [JHJ09a]. Bees [BHOR01, HD08, RP09a]. Beetles [PKW^{+}00]. before [Hau07, MHMGM08, TP05]. Begging [DSK02]. Behaves [SM02]. Behavior [AMW00, ACK00, BKCR01, CR02, KG02, MKN02, Mur00, OMT03, TM00, ZKH^{+}05, BBD06, BR09, BLRR08, BOvD08, BEK^{+}03, HTN04a, HTN04b, LPC06, MGS08, RV05, RBS05, Tul03, UD07, Won06]. Behavioral [Cha03, FSGB^{+}02, FHL^{+}06, NDD08]. Behaviors [Cha01a, BCH03, BFA08, LSH06, YKiP^{+}09]. Behaviour [TGR^{+}00, YTAK01, BG08, CM03, CW08b, Har07b, HWCF07, HBWB08, LPJB^{+}08, LYZ08, MSH05, SSR04, TOB08, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, vV06]. behaviour-pharmacokinetic [LPJB^{+}08]. Behavioural [Jam01c, PCA^{+}09, RB08]. Beijing [WR04]. being [Hau07]. bellied [DPV00]. Belousov [SRR08]. below [ES00]. bending [AAL08]. Beneficial [BR02, JCK09, Bar05, BR09, DPA03, DPA05, Orr06]. Benefit [GG01c, Ste00, Gol08, RSC^{+}06]. Benefits [BKM09, KCHP08, LAG09, MH09b, NP08]. Benthic [Mcn00, MN01]. Bertalanffy [LP07c, RBP^{+}09]. Bertrand [BZ00]. bet [YCS04]. beta [GBZ06, OHM05, YDL06a]. beta-strand [YDL06a]. better [DqLmW07]. Between [ARKL02, CD02, DPA03, FDG02, Had01, ISC01, PL01, SCH02c, TMI03b, AB06, ALM09, AWO^{+}09, AF09, ARR08, Bat09, BGE06a, BdSFFDMC09, Bla04, BB03, BJ02, Cha03, CXM^{+}09, CCF06, CTB09, CGKC07, DLGC02, DBBW09, DBBW11, Ded08, DI09, DFP^{+}08, Dru03, FLBB01, FL03, FS04b, FHD09, FM06, Fur02, Ger09, Gol07, GG02, Gra02, HH07, Hog00, HVPN02, Jam01c, Jav00, JCK09, KF06, KNT^{+}09, KW00b, Ken07, KRN01, KMC^{+}07, KRB05b, LTG^{+}04, LRT04, MZ02, MX08, Mar09a, MB06, MR07a, MLWL06, MLPJ09, MKE^{+}09, Nak01, Nak03, NM08, NSS08b, NA03, NdGG06, OPN07, OKS01, PCH^{+}05, PPR01, PDIS00, PG01a, PL09, PDC02, Raf02, RR09, RDC09, RGSFM07, RGFP07, Row06, SSLB07, SHI03, SB09c, Tak06, TW04, TNTJ08, UD07, UCSZ07, WT09a, WKB07, WT05, Won05, WB06]. between [XFBC07, Xie09, YCC02, Yan08, YB05, ZSF^{+}07, ZZLT07, vLBJK07, vdGMG^{+}09]. Between-generation [SCH02c]. Between-group [DPA03]. Beyond [BLS06, CXM^{+}09, Wax06, PD04, SSJ09]. bgl [RGF^{+}08]. bi [JSCN04, HB02a, HB02b]. bi-directional [JSCN04]. Bi-trophic [HB02a, HB02b]. Bias [SMNL07, DZB^{+}04, GW06, HP04, MX08, SIHH04, SWI07]. Biased [GJE02, YMC01, Blu07, BR09, CH05b, KGL09, QXDW06, Wak05]. Biases [WKL01, Ded08, YDL06b]. biasing [Ded09]. bibliography [CB08c]. biclustering [ZLXY08]. bicoid [LVABV05]. Bicoordinate [Ben03]. Bidirectional [TYW05, JCRJ07b, JCRJ07a]. bidomain [Rot04, TPP^{+}04]. Bifurcation [AKdlPP06, PG00, QB00, CD07, CMW08, GKS06, GK09, HdGH07, NK08]. Bifurcations [SS06a, TYI^{+}06, BHC06, BGF07, RMP08]. Big [Den08, Kan05, LD09a]. bilateral [Ber07]. Bilayer [CMPL^{+}00]. Biliary [EM00]. bilinear [MPOBD^{+}09]. bimolecular [TE05]. Binary [DALP03, HH08, PE04, SG07, Yam03]. Binding [FFN00, KMP03, Mac00, MJW00, PD01, SRN^{+}00, TML02, WQ00, AS07, BM04, CZM09, DHW^{+}09, FWCN05, GSG^{+}07, GFWT04, HBB08, Hua03, Jus08, Kar03, KS08a, KmMK04, LVABV05, MD04, Mei05, PBvdG09, PE04, Sar04, STW^{+}09, TR08, VSP06, WOLS07, YCC^{+}06]. bio [MPOBD^{+}09]. bio-macromolecular [MPOBD^{+}09]. Bioactivity [BS00, HDW^{+}09, Kli06]. Biochemical [Alb02, DALP03, MWB^{+}02, PD01, Voi03, BBCQ04, BB07a, DL08, GAT07, KA03a, Lie05, MLA04, MA04, MS05b, NSS08b, Péc05, SHI05, SHI06b, SSB06, SML07, SML08, UW09]. Biochemistry [JST^{+}02, PTFF05]. Bioconvection [GH02]. Biodiversity [BDBR07, BLMV05a, BLMV05b, DLRL08, AK09, Est07, EAM07, HE08a, MWL^{+}07, MBRRI08, SWN07, WR03]. biofilm [Joh08, ZFVH05]. Biofilms [BBB^{+}07, ES08, MRC09]. biogenesis [KS04]. Bioinformatic [KMP03, RB02c]. bioinformatics [JHJ09a, MPOBD^{+}09, Ano01-33]. Biokinetic [Has01a]. Biol [AGW^{+}08, AMD06, BZ05a, BBM04, BKE04, CSP^{+}09, DBBW11, EPJ^{+}11, FK03a, Gie06, GWM06, HTN04a, IvDHI08a, JCRJ07b, KCP09, Ker06, LH09, LGK^{+}12, Lei10, LVL08a, LW04, ML07a, ML12, MT06a, ML08b, MI08a, NI07a, NI04a, Paw09b, PCZL06, PA09b, RH09a, RMAI09, SHI06b, SI04c, SRAL12, SML08, TK09a, TRM03a, TRM03b, TQUN08, UI04b, VGMM^{+}07a, VGMM^{+}08]. Biological [BP00, CSM02, CA00, CD00, Dai02, DMW04, De 02a, FBUL05, GJE02, Gin00a, HRBL02, Jaz01, KS01c, Lac01, LFLM00, MZ02, Nak01, Nak03, OCP^{+}00, PPR01, PSM06, RC00, RKL02, Tor01, WBB02, BCRG04, BLS06, BGO08, CC08, CMG06, CIV09, CPMG^{+}08, DqLmW07, Dal06, DW05, DGS09, FP04a, GXG03, GDPMDS^{+}09, GF09, HVY^{+}07, Hor08, JJEX09, LME06, MD04, PCZL05, PCZL06, PKS^{+}08, RG03, San03, SPG06, Vin05, Wag03, Wil09, Zam03, dPGR06]. Biologically [Rae02]. Biology [Agi04, Ano00e, Ano01h, Ano02a, Ano05h, BLZ07a, BSRH03, HH05b, Laz02, LHD^{+}01, OI05, TT03, VW00, WMLC02, BC09a, BNRW04, BIB07, CLMMP06, Dem06, KM08, MHRK08, SBI07]. bioluminescence [DS05]. biomarker [APL08]. biomarkers [VGDSU09]. Biomass [GS02a, HM01, Nik02, BR04, Niw05]. Biomass-Diversity [GS02a]. Biomechanical [ACD04, AF09, GT03, TGT06, BG06, Wes03]. Biomechanics [CC01c, LDT09]. biomembranes [MS05a, MS06a]. biomimetic [RRR^{+}08]. Biomolecular [WPE03, BMR08]. Biomolecules [Kal00]. biophysical [EH08, RRR04]. bioreactor [WBD^{+}09, ZC09]. bioregulatory [GGV^{+}08]. biosynthesis [RGW^{+}05]. biotic [BTCD07, FSS06a]. biotin [dQW07]. BiP [GAA02]. BiP- [GAA02]. Biparental [DMZ00]. biped [KH07b]. bipedalism [Syl06]. Biphasic [QASL08, QSAL08]. birch [Che07]. Bird [Ale01, ELB02, MM00b, SPH03, Ale09, AKS07, Lin04, Sac05]. Birds [GA02, LZGL03, Sac07]. Birth [Jam01b, Wel02, Bog04, Hel08, ING04, Jam06, Jam08b, Jam09b, Sta09]. birth-death [Sta09]. Bisexual [GMM09]. Bistability [MKL09, TB04, vKdRP05, LHFH08, NPN09, PHL08, SGCK^{+}08]. Bit [FMD01]. Bits [PFdP^{+}07]. Black [Ano01e, Mac01]. blindness [Wal07a]. blips [PPD09, RP09b]. Block [Che03a, CRB05, LAG07]. Blocks [öOW01]. Blood [DEM^{+}00b, SP02, WGH01, AM06, APL08, BGE06a, BGE06b, BFGB04, BFG05, BFG08a, GRW03, JCK09, KCV05, Kin07, LT08, PRP^{+}03, RGB06, WBR08, WBR09]. blood-borne [APL08]. Blooms [CSM02, CCV08, HBOS05, SC03]. bluefin [GNLEK07]. Blueprints [KTE08]. bm86 [RAHO06]. BMP [ZLN07]. board [Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano04-29, Ano04z, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y]. Board [Ano06g, Ano06h, Ano06i, Ano06j, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. Bodies [FMF^{+}00]. Body [KMGV00, LST00, MGL03, Can07, CM09, CSS08, CTB09, Hel08, HJ07, HNTHA07, NM06, Usa06, VPGA07, WM04]. Boiss [FLS^{+}04]. BOLD [BFG08b]. boldness [JZST09]. Bolus [NB02, MF05]. bond [RN07]. bonded [SWB06]. Bonds [Tho02, PSSY09]. Bone [CACC02, FK01b, KG02, MGAD09a, MGAD09b, SNT03, ZCC01, Cur04, DMW04, FK03a, FH07, FHD09, GGV^{+}08, GBGAKD05, IvDHI08a, IvDHI08b, KSG03, LTG^{+}04, MRGADD08, NP05, THL03, TT04]. bone-healing [IvDHI08a, IvDHI08b]. Bones [Tay00]. Boolean [ML12, AHCN07, CAS05, DTG09, DB08, GBK^{+}07b, IiKY07, iKIY09, MWS09, ML09a, QD09, Rae02, TC09, VFC^{+}09, WW07]. Boophilus [RAHO06]. boost [CW07]. Boosting [GG03a]. border [HMB^{+}08]. born [JHJ09a]. borne [APL08, CBC^{+}09, HRvdD08, RPNH03]. Botanical [SMG^{+}03]. Both [ZCC01, DL08, HC07a, Lei09, Lei10, OW07, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08]. bottlenecks [AB04]. bottom [IPY07]. bottom-heavy [IPY07]. Botts [VGCGM^{+}02]. bouncing [BD08b]. Boundaries [BB02, DFC^{+}02, RN07]. Boundary [Har01, MZ02, SBH01, HKP07, Rej07]. bounded [JBF^{+}03, LP08]. bounds [FS09]. Bout [YTAK01]. Boutons [KFG^{+}02]. bovine [EL09, GLE^{+}09]. bowel [Agi04, BBB^{+}07]. bownockeri [SKY09]. Boyden [CC09a]. brachiating [GR05]. Brachiopoda [SKY09]. Brain [KTH^{+}00, NKC^{+}08, Phi02, SP02, AMD05, AMD06, GMK06, GK06a, GT06a, HKP07, KV05, KJJB07, LSMZ08, MDD06, Mat06, NL03a, RRR04, SCNP^{+}06, TN04, TTN05, ZAD07]. brains [LD09a]. Branch [BM03b, SW03, LBS06, Xie07]. Branching [GL01, GBCC01, HGV01, Kun05, OSCK01, Zam01, ZMQX01, Cam03, CLMMP06, DECEK06, FFHK09, GASA09, GMM09, GvO04, GHC03, HM06, HIM09, II06, Kun03, MAL03, MHKS03, SZC^{+}03, SK09b]. brant [PB07]. breakdown [KCS^{+}06]. Breaking [Di 00c, OPN07]. breaks [Wal07b]. Breast [SA01, EAC^{+}06, ECAV07, FBUL05, Lin07, MMUGD09, SA05b, TCP05]. breathing [BTS08]. breeder [STMH04]. breeding [BWMS07, FS06]. bridge [Dru03, STSD09, Zah00]. bridges [GZ04]. Bridging [GS00, HP00a, LSH06]. bright [NGN^{+}04]. Brood [ORM03b, HB09]. brothers [Bla04]. Brown [BDMR06, MGL05, Cle07]. Brownian [Kut03, LDW04, LDW05, YMC01]. brucei [NO04]. Brusch [Rot08]. bryophyte [BC09b]. Bt [MMV^{+}04, MGL^{+}06b, MGL^{+}06b]. Bt-plant [MGL^{+}06b]. Bt-resistant [MMV^{+}04]. Bubble [HHR01, HVN07, SWRH03]. bubbles [KRN03]. Buckling [FMP01]. bud [HIM09]. Budburst [Chu00]. budding [Ber03, Iro09, iTYU06]. budgets [MKE^{+}09]. buffering [GKS06]. Build [GMdA^{+}07]. Build-up [GMdA^{+}07]. builders [Gol07]. Building [HW01, VCBV^{+}06]. bumpiness [Kin07]. Burden [HFH03, RRS06]. Burst [PB02, Bul08, Ver04]. burster [LFSG^{+}05]. Bursting [DS00a, FDA^{+}09, Ped05, Ped07, RWK08, WB04, WKRD09]. Businesses [CMW02]. Butler [For09]. butterfly [OST09]. bystander [LFP^{+}05, ØØS06].

C [CS09, DqLmW07, KRN01, NS06, ARR08, Bat09, DLRP07, DL08, KS08a, KOK06]. C-mol [Bat09]. C-to-U [DL08]. C. [MY01, MGS08, NM09, RW08]. Ca [AGT^{+}01, BGP00, DBR^{+}05, FB08, FK01a, FWLN04, FKAC06, GBD00, HS07, KBT08, Kar03, MGP^{+}08, MTS05, PAH06, Saf09, WHH07]. Ca-ATPase [AGT^{+}01]. Cabanes [Ril02]. cache [PM03]. cadmium [MCMS06]. Caenorhabditis [KCS^{+}06, RDSB^{+}03]. Calcium [ESGK02, GTG01, GG03b, GS08a, PH03, SRMW00, VLFN00, BFGD07, FL03, Gam06, GKS06, HGP^{+}07, KGG08, LH09, LK08b, LGB03a, MS03a, MRGADD08, WTC09, WMS08b, LH08]. Calcium-activated [VLFN00]. Calcium-dependent [GS08a]. calcium-induced [LK08b, WTC09]. Calculate [HS02b]. Calculating [Kon09, WW08]. Calculation [FSG00, For07a, Mei05, Tur02, HP04, LMF03, SHI05, SHI06b]. Calculations [SP01]. calculus [Bra07, NTU06]. calorie [Bar05]. calsequestrin [LK08b]. Cambium [Kra02]. Can [BS01, BR02, BLT03, BTL08, Dor03a, EKIL01, HHR01, ITL09, NS03a, Par04, PL01, PG01b, SGTF07, VPGA07, Arc09a, BN04, Blu07, Bok06, CS00, Di 00c, EFW07, FI06, FZ04, Fow09, GOP09, HB01, Kom07, MGDDH08, MS03b, MGL^{+}06b, MPA^{+}08, OI06, PN01, PDM04, PE04, PGH^{+}04, RZF03, RHF07, RR09, RB09a, RMRG09, SR08, SH08b, THKU07, Uit09, WFGP04, WN05, WG03, WVA05b, vV09]. Canada [TP05]. canalization [FI06]. Cancer [CP07, DGD^{+}09, LY02, QHF^{+}07, SHHD02, SRCDS08, TS02, WW00, ABM04, AS09, BCJ^{+}08, BRS^{+}09, DCP^{+}08, EAC^{+}06, ECAV07, Fra03, GC08, GA08a, HPZ09, IZGG05, JWB^{+}09, Kam03, KSO06, KSN03, Kom06, LKW06, LVL08a, LVL08b, MCM^{+}09, MFM^{+}03, MMUGD09, Ort06, ØØS06, SD04, SBGA04, TSC04, VHF08, VGDSU09, Wod07, ZAD07, dVG04, dVG06]. cancer-related [MMUGD09]. cancerous [vLBJK07]. cancers [SSB^{+}07]. Candida [SSF09]. candidate [Yan06]. cannibalism [NI04b, NI04a]. cannibalistic [Wak04]. cannot [Di 01b, Sch05a]. canonical [HR08, Voi03]. canopies [BH04, BLL08, HHH06]. canopy [MDCC06, VN07]. Cap [BG00b, NKC^{+}08]. Capabilities [SSJ09, SP00]. capability [MAHD06]. Capable [SCGFS00]. Capacitance [GB01]. Capacity [ES00, BdABA09, SF08a, Whi07]. Capillaries [MKB03, Kin07]. Capillary [GP00, GP01]. capped [MJ07a]. capped-rate [MJ07a]. capsid [AKR09, PT09, Twa04]. capsids [KMT06, KTE08, TH06]. Capture [FSGB^{+}02, CB07b, RHM^{+}08]. Capturing [MSSS09, DRV^{+}06]. carbohydrate [NM08]. Carbon [DLGC02, TPWG01, BS05b, BSJ04, EBId09, Gam06, Gie03]. Carbonic [Tho02, Gam06]. carboxylase [dQW07]. carcinogenesis [LVL08a, LVL08b, SGG^{+}07]. carcinogens [KDKS06]. carcinoma [BMH07, FBUL05]. Cardiac [CBBH01, Cyt04, JK06, KC03, KMGDG04, MS08c, Rot04, SWC^{+}08]. cardio [CKE06]. cardio-respiratory [CKE06]. cardiomyocyte [TOB08]. Care [MSWH00, WH03, MPM07, NWT09, SWI07, WTSN06]. careful [SWRH03]. cargo [FWE06, ML09b]. cargo-exerted [ML09b]. Caribbean [SZC^{+}03]. Carlo [ACK08, GFW^{+}09, HW05, HW06, SPAH06, SIHH04, WFGP04]. carnivory [Bal04]. Carnot [Smi08b]. carotid [HdGH07, VPGA07]. carrier [APL08, FMI05, FM06, NPN09]. carriers [CWDM06, Jam07, TGV^{+}07]. Carrying [ES00, BdABA09]. cartilage [BM06]. Carving [KF03]. cascaded [RAK^{+}08]. Cascades [Har02, SSB^{+}02, BH03a, FMHW08, LP07a, SV07]. cascading [AB04]. Case [AP01, AM01, ACCC02, BR00, Kor07, RB02c, AHT^{+}07, BG08, CM09, DB08, Dor03a, DSA^{+}06, FPL03, FHL^{+}06, GXG03, HJ07, KMW03, Lew03, MS03a, SSD09, TSS06, TGN07, VCBV^{+}06, YM05a, YM05b]. cases [CC08, CHCC^{+}04]. caspase [SS05a]. caspase-3 [SS05a]. caste [Tof06]. Castelvecchio [Ano02a]. Cat [BKCR01]. catabolism [Bat06]. catalysed [BS05b]. Catalysis [For00b, MKT^{+}00, CY08, WZW^{+}07]. Catalyst [ALSM06]. Catalyst-induced [ALSM06]. catalysts [RN07]. Catalytic [KY02, MS01a, Tho02, TGR^{+}00, dCZJ^{+}04]. catalyzed [Alb08]. catastrophe [SD04, SSDM06]. Catastrophic [ACL03, wTA09, YN02]. Catches [BZ00]. categorization [KJ08]. catenin [Daw09, vLBJK07]. caterpillar [LDT09]. catfishes [VAAH05]. Catheter [ABR02]. cation [AS07, MSGS08]. cation- [MSGS08]. cattle [IGHW07]. caudatum [UKY^{+}09]. Cauliflower [LN02]. Causal [BMR08, Kru02, Jäc07]. causality [BB07a]. causation [MLS09]. cause [GHHR03, GT06b, MS08b, WG03]. Caused [RBBH02, TMI03a, KJD05, KC07, SW08b]. causes [GKB03, GS08b, PLG^{+}06]. cautionary [Van06a]. cautions [BDMP^{+}08]. Cave [OCP^{+}00]. cavities [TTN05]. Cavity [COS01]. CBPP [EL09]. CD25 [LGCL07]. CD28 [CRB05, SF04]. CD4 [LGCL07, AWJ02, CRB05, PD01]. CD8 [ABP^{+}03, HW06, LVV^{+}01, vdBWLS07]. CD8-driven [vdBWLS07]. celestial [BH04]. Cell [AMW00, BG00a, BFGS07, CMTU01, CMS^{+}00, Coo01, DFP01, EP00, FMF^{+}00, FK01c, GS00, HEH^{+}09, HSB01, Hog00, JST^{+}02, JL01, LCL07a, LVV^{+}01, Moc02, OI00, PKL02, PD01, RBN^{+}01, SHHD02, SLH^{+}09, SK01, SRN^{+}00, TMI03a, TTN05, TSZ^{+}07, TMI03b, TPD^{+}00, Tyr01, TN01, UI04a, UI04b, VRB01, VP01, WJ01, YBV^{+}00, Agi04, ABM04, ACSY04, ABTR07, AGZ^{+}06, APS06, AB00, AB03, BLS^{+}09, BGK09, Bel06, BGB08, BFK08, BB07a, BI09, BC06, BdOP06, CLHW07, CLH07, CSR^{+}05, CDFP04, CE08, CC09a, Cin03, CL05, CS07, CLM07, CSD09, CWJ07, Cyt04, DHYHR09, DBBW09, DBBW11, DECEK06, DSVBW07, EZM05, FWLN04, Fli05, FFHK09, FFD^{+}02, FKS07, GC08, GA09, GCP04, GCB^{+}07, GCS07, GSNH08, GZ04, HHBY06, Har07b, HJ06, HK07, HTN04a, HTN04b, HPZ09, HC07a, HD08, Iro09, IvDHI08a]. cell [IvDHI08b, JAFW05, KBT08, KP06, Kin04, Koh07, KJD08, Lan00a, LWRK07, LPLC00, LLC03, LHFH08, LFP^{+}05, Lit07, Lo07, LCHB06, Man06, MS08c, MH07, MR07b, MRJR09, ML04, MSL04, MS06b, MCN08, NK08, NS07, NT04, OOHI06, ØKRG04, Ots08, OHM05, PS03, Pal08, PW09a, Paw07b, Paw09b, PGMK^{+}03, PTFF05, PAVS03, PRP^{+}03, QO08, RZF03, Ros03, SF04, SS06c, SWM09, SMPM09, SBPC05, SLB07, SG06, SMCT08, Ste04, SMPvdB08, SRAL12, SML07, SML08, iTYU06, TCP05, VHF08, WWS^{+}06, WSM05, WO04, WL03, WM04, Wod07, WH07, YLCW06, YHM^{+}06, YYA09, ZAD07, Zhu09, ZAB^{+}09, dT07b, vdBR04a, vdBWLS07]. cell- [JAFW05]. cell-autonomous [Cin03]. Cell-based [TSZ^{+}07, JAFW05, SMPM09]. cell-cell [APS06]. cell-cycle [ABM04, BB07a, VHF08]. Cell-fate [LCL07a]. cell-getting [Cyt04]. cell-heating [HD08]. cell-phenotype [IvDHI08a, IvDHI08b]. Cell-Rearrangement [TMI03b]. cell-to-cell [HJ06]. Cells [AWJ02, Ano01h, Bar01, BKCR01, BGP00, BPZ^{+}01, DS00a, ERM00, FMF^{+}00, FY00, GB02, KK00b, LSMLB^{+}02, MLMW01, MW01, TM00, TLA00, UI02, WS02b, WW00, ABM04, AMP06, BFG08a, CSON^{+}05, CCT^{+}09, DT07a, Daw09, DBB09, DdB03, Fin06, FB06, HW06, Jäc07, JCK09, KR05, KF04, KD03, KSEK09, KLN^{+}09, Koh07, KM05, KI04, KI05a, LFC04, LGCL07, MS05a, MH07, MRJR09, NMH07, NA04, Noe00, OAKC08, Ped05, Ped07, Rej07, RR09, RWP^{+}08, RG06, Ros03, SI09, Saf09, SB05, SMPM09, SLHN06, SMNL07, SRCDS08, TN04, Thi04, TMH04, Wal08, WCLL08, WCH^{+}09, vLBJK07, vLHH06]. Cellular [BMH07, BWHK02, CLO^{+}02, CMB02, HW01, HS00c, HQP^{+}09, JG01, KTH^{+}00, KCT02, PGLG01, ABM03, Ark05, ACLH05, AMD05, AMD06, BST05, CKS04, CD05, CL05, DO04, DLF^{+}07, GA07, GA08a, GFWT04, GZ04, IvDH^{+}08, KD03, LK08a, MWS09, MD06, MHMGM08, NOT04, PBHS05, PWK03, PWG09, PA09b, PA09c, RM04, STK08, SM06b, SML07, SML08, TK05, THL03, XWC08, Yag09, YCS04]. centenary [Jam09a]. Center [GBG01, KS01c, MHDOG01, MH02, Hot03, KD03, RZF03, TKM06]. Centers [PDIS00, WS03b, SBZ^{+}08]. Central [JR08, RHG00, LLCM01]. Centre [IM02]. centres [GSM06]. centromere [RGPB08]. cerebellar [Saf05, Saf09]. Cerebral [Car02b, MBD08, KH07a, KH09b, RGB06]. cerevisiae [GKM^{+}00, Thi04]. Certain [JE01, AH08, Wil08b]. Cessation [HSB01]. CG [KTP09]. Chagas [ISC01]. Chain [HS02b, JBP02, RWF01, WRF01, AC04, GFW^{+}09, IGHW07, LS03, Lal06, MSGS08, PB06]. Chains [Koc00, SHF02]. chalk [CH07]. Challenge [ZSS02]. chamber [CC09a, Luc05]. chambers [LB05]. Chance [HHR01]. Change [FW00, JW01, Lac01, SSWF01, BGE06a, BGE06b, BN04, BW06, BFG08a, BTCD07, Hav04, LW07, MS06a, MM03a, MPL06a, NS07, OBPH^{+}08, SS06b, SW08a, SZ09, TY09]. Changes [FMF^{+}00, HVPN02, Kar03, MFB01, OKS01, Tas02, YFKP03, Abr09, Amz04, AGZ^{+}06, ACLH05, BFG08b, CSON^{+}05, FCP03, GBZ06, LZ09a, OCA08, OS03, OD03, PCA^{+}09, RLCIB05, SdAC^{+}08, SGG^{+}07, TK08, Tas05, ZTKH09, dM09]. changing [BTCD07, SP07b]. Changizi [How09]. Channel [CPC^{+}00, DS00a, FSG00, GJE02, GB01, KSM02, SS00, TTKZ01, VLFN00, WT01, DHW^{+}09, GFW^{+}09, LSS06, MTS05, Tlu07]. channeling [HGH08]. Channels [CMPL^{+}00, Lan00b, Bye05, Dim05, GKXS07, Gre05, HS07, OP05, OUPG09]. Chaos [FK01c, SI00, SA08b, YAL04, MS06b, YPY^{+}09]. Chaotic [Hok00, SR08, SH07, WS02a, BM08, KNS05]. Chara [BPZ^{+}01]. Character [Fra02, Mad00, Tur02, KC07]. characterisation [UW09]. characterising [RH08]. characteristic [AHT^{+}07]. characteristics [AO05, Che06b, IvDH^{+}08, LR07, LM08, Mag04, Sac05, SM06a, SCH05b, SBZ^{+}08]. Characterization [FmW08, NAV04, SW03, TSRB08, DqLmW07, Her09, LSS06, LTW06]. Characterizing [IB06, MGT^{+}06]. Characters [WS03a]. Chargaff [ZXWF08]. charge [GKTN07, SFC^{+}09]. charges [Kur08a]. cheaters [EFW07]. cheating [Arc09a]. Chemical [Bye05, Smi08b, Bak04, FR07, Fer09b, HS04, Leh00, LF08, MGAR07, MHMGM08, Sel06, SRR08, SCS04a]. Chemically [SCH02c, WD03, CEP05, OQGC07]. chemicals [SB06]. chemistry [HDW^{+}09, MHMGM08, Smi08c]. chemistry-driven [HDW^{+}09]. chemoattractant [JA05]. chemoautotrophic [MB09]. chemomechanics [GG07]. Chemoreception [Mau02]. chemostat [GC09, LWC09, RDH09, Thi04, GB02]. chemostat-cultivated [Thi04]. chemotactic [SMNL07, SNA^{+}08]. Chemotaxis [JA05, NSI08, RWP^{+}08]. chemotaxis-driven [RWP^{+}08]. Chemotherapeutic [HFH03]. Chemotherapy [AM01, Gar02, CA09, FM09, KSO06, MG09, ØKRG04, SHRR06, dPGR06]. chi [PKS^{+}08]. chi-square [PKS^{+}08]. Chiasma [GG01c]. Chick [PMMS01, BS05c]. Childhood [GG03a, TP05]. chimpanzee [WSM05]. chirality [LDW05, SW04a]. Chlamydia [WM04]. chloride [CTS^{+}08, Luc05]. chlorine [Fer09b]. Chlorophyll [Laz03, GZL^{+}03, LIK^{+}05, LJ09b]. Chloroplasts [TPWG01]. Choice [Bjö02, NS03b, SI01, HH04c, HH05b, LKK07, NP08, NWP07, PB07, Ros03, SWI07, SHI03, SI04c, WA07]. choices [RDYH09]. cholinergic [SM09a]. chondrocyte [CC09a]. choose [ITKL08, Paw07a]. Chou [GKNT09, Lin08, hZzGqX^{+}09, ZF08, ZCLZ07]. chromatids [JM07]. Chromatin [Ish00, DRMLS09]. Chromosomal [For04, KRN01, KSN03]. Chromosome [HAC^{+}09, IST01, ISWT02, RPB03, Amz04, GT06b, HK05, Ken07]. Chromosome-specific [HAC^{+}09]. chromosomes [CW07]. chronic [CM05a, DKLL05, KLL08, ML04, PTD09, WL07a]. CI [BM04]. Ciliary [DF00]. ciliates [EPR07, PER03]. Ciocco [Ano02a]. Circadian [Dai02, GRG02, KMI02, LGE00, RGG00, RVMR01, UHK01, UHI02, BMT04, Cal06, IB06, KA03b, KI05b, KG06b, LM07, LG04, LMT05, RSSM06, Ref04, RCS05, SRR08, SG04, SKK^{+}07, THKU07, TIM06, TYI^{+}06, WLHB07, XK07]. Circle [öOW01]. circles [DdB03]. Circuit [SM02, BS05a, HE08b]. Circuits [CMW02, MWS09]. circulans [KiHM^{+}03]. Circular [FM03, LM01a, DM07]. Cirrhosis [EM00]. Cities [CMW02]. Citric [MCWF01, Kal07]. city [AKS07]. Cl [FTEG02]. clamp [LSS06]. Clarifying [Yos03]. clariid [VAAH05]. Class [CLXC03, Kun01, LL01a, PD01, QB00, QA01, APS08, AS09, CTZ^{+}06, CCZC08, LS08a, LLS^{+}09, MBBR06]. class-wise [APS08, AS09]. Classes [CDF00, LN02, CZC05, CFLC06, JAJH07, LGCL07, XWC08, YPY^{+}09, ZDC08, ZCLZ07, dFG08]. Classical [ACK00, DKD02, HBWB08, PSM06]. Classification [LY02, QWQ09, Rei02a, Tak06, TU00, AS09, DKS04, KGB04, MRJR09, MBBR06, MK03, MGDM08, MGDBdM08, MMUGD09, TK08]. classifier [BE08, CFLC06]. classify [GKNT09, Pán08]. Classifying [SMG^{+}03]. claws [PM07]. Cleavage [Yos03, ZSRB07]. click [BEBV03]. climate [BTCD07, MPL06a, OBPH^{+}08, Whi07]. climbing [LS07]. Clinical [MAC06]. cliquishness [HH06b]. Clock [Dai02, HRBL02, KW00c, KMI02, MM00b, Cal06, Cin03, FP04b, GP08a, KI05b, LM07, LG04, LMT05, RGSFM07, RCS05, TYI^{+}06, WLHB07]. Clock-and-Compass [MM00b]. Clock-Controlled [Dai02]. Clocks [Ano01c, GG01b, Ost04, RSSM06]. Clonal [Kor07, PKL01, GA08a, NAP04, NMI06, RLCIB05, UKY^{+}09]. clonally [JWH08]. Clones [PDC02]. cloning [CLS08]. Closed [CC01a]. closely [RSH^{+}06]. closing [VAAH05]. Clostridium [AIK00]. closure [LSAA^{+}06, MLS09]. closures [MDL04]. clots [ARR08]. clouds [BH04]. clues [BDMR06, MGL05]. Cluster [CKM05, PD04, CDDW02]. Clustering [NA02, GKNT09, KTP09, KK09, MLPJ09, TBC^{+}08]. Clusters [MKLD02, SB00, HSF09]. CML [ML04]. Co [Ste02b, BCH03, CXZ^{+}09, SW08b, MB09]. Co-Evolution [Ste02b, BCH03]. co-evolutionary [SW08b]. co-transcriptional [CXZ^{+}09]. CO2 [KRNG08]. Coalescence [FB01, SLB07]. Coalescent [Möh00, DBF07, Sta09]. coast [LS08b, DFCL08]. cockchafers [HHH06]. cockroach [WLHB07]. Code [AGCLMM03, CDF00, Di 00a, Di 01a, Di 01b, GS02c, KS01b, LM01a, Ste02b, AD07b, Che03a, Che06a, CL09, Dav09, Di 05, Di 09, Fol08, GNH^{+}05, JS07, KS04, Pat05, Poh08, Rak04, Tlu07, ZF06]. Codes [De 02a, FM03, MWCS04, SH08a]. codescript [Wil09]. Coding [HGB^{+}00, Leh00, AFZ08, DBHS00, KMC^{+}07, MX08, NA03, SP05a, TFYY03, TGN07, YY07, ZW03, ZYD^{+}05]. codling [TNTJ08]. Codon [AS00, LN02, CSW^{+}08, DZB^{+}04, GW06, NGT05, Plu06, RAHO06, SWB06, Tak06, ZW03]. Coefficient [DFC^{+}02, WL02]. Coefficient-Orchestrated [WL02]. coefficients [HP04]. Coevolution [Arc00, GS02b, Gin03, LBF01, LLF02, GNH^{+}05, HK09, NI06, NI07a]. Coevolutionary [PSWK09]. coexist [DW08b]. Coexistence [AF01, AN00, CMB02, CVH03, FM02, FSS06a, HI04, HLH01, LL01b, PDC02, SA07a, TH05, WZL00, Ama06, BLMV05a, BIB07, Dam04, FS04a, GEF09, LS09, MN06, MN07a, MI08a, MI08b, MBJ09, PNG03, PM03, RAZ03, RDH09, SB05, TYW05, vdBvdB09]. Cofactor [MS01a]. cofactors [ZF08]. Cognitive [KF03, Khr04]. Cognizer [Nak03]. COGs [SRS09]. Cohen [KV05]. coherence [BD08a, HE08b, KGG07, Rob03]. Coherent [UI02]. cohesion [LN03]. Coincidence [GV03]. Coinfection [MPA^{+}08]. coital [Jam08b]. cold [DCC^{+}08, JBJ^{+}08, NKC^{+}08]. cold-adaptation [JBJ^{+}08]. coli [AHT^{+}07, AS00, BKP09, CB07a, GGK05, Koc00, LR07, LSD^{+}00, LL06, LLO08, MBP07, NGT05, NPN09, PCSL^{+}06, RFH^{+}02, RGF^{+}08, RGW^{+}05, SWB06, TBB^{+}06, TBC^{+}08, WDH^{+}09]. Colicin [CMPL^{+}00]. colitis [Agi04]. Collagen [SHF02, DWBB04, FHD09, GHA03, GMAH07, HdGH07, KP06, KH07a, QXDW06, PSSY09]. Collagenase [MW01, QXDW06]. Collagens [SHF02]. collapse [AM04b, AM06]. collapses [MN06, MN07a]. Collective [CKJ^{+}02, JSCN04, LAG09, NDD08, YKiP^{+}09, CGF^{+}08, DO04, MAHD06, ZKH^{+}05]. Collie [HDRM00]. Collision [BTL08]. collisional [RBS05]. Collisions [ACK00]. colon [GT08, LVL08a, LVL08b, MMUGD09]. Colonies [CMW02, DS00b, SFV02, AO03b, BC06, Mig06, Tof06]. colonization [CCF06, LdGH09, YNO09]. Colony [GTDA02, NC02, AK04, CM03, KiHM^{+}03, NBT07, VGS^{+}05, WRG^{+}04]. colony-based [NBT07]. colony-stimulating [VGS^{+}05]. color [KJ08]. Coloration [LY01]. colorectal [VGDSU09]. colorful [LY03a]. Colour [Arc07b, SZ01]. colouration [SG07]. Colours [Arc00, Arc07a, Arc07b, Arc09a, HH04b]. Column [YN02]. comb [HD08]. combination [ZW07]. combinations [Jam09a, Kur08b]. Combinatorial [Cha01a, SH06, TSK09]. Combined [KDKS06, PSSY09]. combines [BNRW04]. Combining [YCS04, SSD09]. come [Hau07]. comedo [FBUL05]. Comment [BDMR06, For02, Sol01, Tor00, BM03a, How09, Jam08c, Ped05, Sch08b]. Commentary [KRN01]. comments [Jam08a]. commitment [MSS05]. committed [CPMG^{+}08]. Common [SS00, SS02, AR05, Bol06, Di 06b, Lac09, Wak07, Zhu09]. Communication [BJ02, GG03b, PBR03, DPV00, Bye05, Fei08, HH08, RG03]. Communication-Based [PBR03]. communities [BH08, CE05, CK08, CCF06, Eng07, Fow09, MK09, ML08a, ML08b, PDC04, wTA09, TY09]. Community [BH01, KDK02, Paw09a, RRKF09, SSA00, DLF^{+}07, Fow09, JI05, MGC04, MI08a, MI08b, SAC^{+}04b]. compact [Cur04, FFIS07, MH06a]. compactification [MHKS03]. Comparable [MCM^{+}09]. Comparative [KMI02, LR07, LD00, Nar06a, PC02, SRS09, ADHM09, AS00, BRC07, OP05, Pán08, WW08, XSD^{+}05, dPCP^{+}08]. compare [DW08a]. Comparing [ACSY04, CC09b, EN02, HIN00, PHdB09, RGG00, SMG^{+}03, DGS09]. Comparison [BP00, BGE06a, DOT02, GHC03, HLP06, KW00b, LW08b, SHF02, Sne03, WZ08, ARR08, FS04b, GNH^{+}05, KS08a, MBP07, Mei05, RWP^{+}08, SZLK^{+}05, SA07b, Tas05, TPN07]. Comparisons [GKK06, Mad00, RSH^{+}06, Kun06, LKC07, Ste09]. compartment [Lit07]. Compartmental [BKE04, Mur00, SP02, Gie03, LK06, Lit07, TWE04, VTC08a, VTC08b, BKE03]. compartmentalization [BTL08, CHdV06]. compartmentation [FKAC06, KSK08]. Compartments [Kut03, MKB03, RSS04]. Compass [MM00b]. compensated [THKU07]. Compensation [ADMZ02, Dam04, KI05b, RCS05, SRR08]. Competence [Mac00]. Competing [SCS07, BLMV05a, BH03b, SB05, TH05, TY09]. Competition [ARKL02, BP00, BW01, BIB07, BJ02, DFP01, GG02, Gro02, KY00, Rev00, Rev02, WZL00, ALSM06, ACK09, BP03, BdABA09, Blu07, BB03, BCL08, CW08a, CXM^{+}09, CB07b, CCF06, CSM05, Dam04, Eam06, EFW07, FT09a, FR06, Gol07, Gro04, HH06a, Hua03, INSR08, KNS05, KCHP08, MS08a, NSS08b, PDM04, PB08, PNG03, RMF08, RS04, SA05a, VN07, Wak05, WS04, WLHB07, YOYT07]. competitions [ZWT^{+}08]. Competitive [AN00, CVH03, DM00, OKTS02, Wil06a, CKS07, Fow09, Kom07, LS09, MT06a, MT06b, MN06, MN07a, RSS04, ZJ05]. competitors [MBJ09, YB05]. complementarities [NN07]. Complementarity [FPC01, RB02c]. Complementary [KNT^{+}09, NA03, ZXWF08]. complete [WLZ^{+}06, ZYD^{+}05]. Complex [BNT^{+}00, BSRH02, BSRH03, DP04, HDRM00, HS03, Kru02, LSMLB^{+}02, Man01, PDPL05, SWC^{+}08, WB04, BPC08, BBPSV05, Bie06, GMMGD^{+}08, GK04, HK09, JST^{+}08, KYZ^{+}08, iKIY09, Kut05, LP07a, MK06, NS09b, NL04, NO04, PSSY09, PWVvR09, RD07, Sac04a, SKI^{+}06, Tak06, VGS^{+}05, VS08, WS03b]. complexes [Kut05, She06]. complexing [Gre05]. Complexities [ZYW07]. Complexity [Cha01a, CK08, GDS07, GMMGDA05, GD08, KGD09, Ker04, Ker06, Kon06, Kon09, LW08b, Loc08, MPL06b, MGDM08, PKS^{+}08, RWID^{+}08, RKH^{+}06, RHH08, SSM09, UD07, Yag09]. complexity-stability [KGD09, Kon06]. compliant [BGE05a]. Component [Sil02, MvdO06, SPAH06, SKK^{+}07, VGDSU09, YKiP^{+}09]. components [GCH^{+}07, IHV^{+}06, LFFT06, MB09]. Composite [KCV05]. Composition [HM01, LN02, CZC05, CC06a, CTZ^{+}06, CL07a, DCC^{+}08, DCL09, For07b, GKNT09, Hul05, Kur07, Lin08, MBBR06, MGDM08, SYC06, WYC06, XWC08, hZzGqX^{+}09, ZF08, ZDC08, ZCLZ07]. Compositional [SSKL01]. compound [GMMR07, TK08]. comprehensive [HHBY06]. Compression [ZM03, MBB08]. Comprising [FY00]. computable [MLS09]. Computation [SV05, Smi08c]. Computational [Ano09a, CTB^{+}05, COS01, CBBH01, GMMR07, GP00, GP01, GG09, GZG07, HGV01, KTCD00, Len01, Saf09, SKY09, SS05b, YMLK04, ZLN07, BWMS07, Bon04a, DWBB04, FDA^{+}09, GBGAKD05, HL05, HGH08, JA05, JTGP06, KP06, KSK08, KOT07, LS04b, MGT^{+}06, MWD08, OAKC08, Pán08, PDB08, PD06, RLCIB05, SMCT08, SKS09, TGV^{+}07, Zac09, ZSRB07]. Computer [AIKP06, NJVA04, BS05c, GT08, HED06, UKY^{+}09]. computer-simulated [UKY^{+}09]. Concentrating [TPWG01]. Concentration [GB01, SCC^{+}00, WGH00, YFKP03, FL03, HW09, JK03, Kur08a, Tho05]. concentrations [APL08, GMFS06, TE07]. concentric [Ber03]. Concept [GBG01, PDA^{+}00, GF09, NI07b, Pen03]. Conception [DML02, Jam08b]. Concepts [CGH01, LHD^{+}01]. Conceptual [For07a, Pep04a, Ver04]. concerning [Koh07]. concerted [Gam06]. Conciousness [Ano01f]. Concord [ES01a]. Concurrent [BSWM00]. Condition [MGP00, KST07]. Conditional [HS00b, Fra03, Men07]. conditioned [Ger08]. Conditions [IS09, Jam01c, KYS06, SM01, TGR^{+}00, AA04, BBKGP08, BLMV05a, Bat09, BP08, BBK04, CSP^{+}08, CSP^{+}09, Fei08, GMM09, Jäg08, MWS09, MS03a, Mig06, NOT04, SA07a, YY06, YFK05, OSK^{+}05]. conductance [RWK08]. conductance-based [RWK08]. conductances [FB06]. Conduction [EMS02, HD08]. conduit [HVN07]. conduits [KRN03]. Cone [Moc02, TMI03b]. cones [GMMR07]. Conference [Ano00e, Ano00-30, Ano01-33, Ano09a, Len01]. conferring [PCB07]. confidence [WL09]. Conflict [ES01a, Had01, Mce03, KI09, WT05, YM04, Yan05]. conflicts [WNT08, YM05a, YM05b]. confocal [TR08]. Conformation [LFLM00, YDL06b, AH08]. Conformational [FW00, Hav04, LZ09a]. Conformations [SV03]. Conformist [HB01]. Confounders [Jam01c]. Congo [CHCC^{+}04]. Conical [Yos03]. conjectures [FP04b]. Conjugates [LPLC00]. connectance [Est07, Fow09]. Connected [Leh00, FPL03]. connecting [GSNH08]. Connection [KRB05b]. Connections [VW00, Sta09]. connective [GMAH07]. connectivity [AMS09, IKD04]. conotoxin [MBBR06]. Consciousness [Tan01, Bod08, OSJ08, Sev06]. Consecutive [FY00]. consensus [DT06]. Consequence [GTDA02, Puj02, SKN^{+}03, Amz04, FK03c]. Consequences [BJD^{+}02, BB06b, FR02, Has01a, KVMV01, LLF02, MJW00, PPRI08, BCL08, KV04, SF04, Wil08a, vKdRP05, vKPdR07]. Conservation [AS00, KCS^{+}06, YHI03, YYI03, SBI07, TH03b, YHI04, YHI07]. Conservative [BKKR08]. conserved [YMLK04]. consideration [GNH^{+}05, JEHK06, SS06b]. Considerations [Dor03b, Ped07, Pro03, Fer09b]. considered [Buc04, KV04]. Considering [Laz03, DSU^{+}04]. consistency [CS06, SHP09]. consistent [Bel06, KSPA^{+}08]. Conspecifics [CMB02]. Conspicuity [Rei02a]. conspicuousness [MR07a]. Constant [DNS00, CLHW07, MPOBD^{+}09, MLWL06, SW08c, TE07]. Constituent [HCMF01, HCB^{+}02]. constitutive [IS09, LDT09]. Constrained [BR01, WD03, AC04]. constraint [AC07a, ACK09, CWJ07, LVH04]. Constraints [CP03, KMGV00, KVMV01, Leh00, LL00, LH01, BBCQ04, Bye05, CW08a, Jam06, KSG03, LB05, LK08c, MSS07, Paw09a, Ace00]. Constraints-based [CP03]. construct [SJD^{+}09]. Constructal [Mig06, BLL08]. Construction [BSRH02, BSRH03, HCMF01, KB02, OLS^{+}02, PV00, WS02a, WCA06, BFK08, Bur09, LB05]. Consumer [VW00, KUK07, PM03, vKPdR07]. consumer-driven [KUK07]. Consumer-Resource [VW00, vKPdR07]. consuming [RMAI06a, RMAI09]. consumption [SX06]. Contact [HED06, LP00, RGZ09, UI02, KSY^{+}08, Mas09, MNP06, PSWK09, RA08, TY06, dM09]. contacts [NdGG06, RWP^{+}08]. contagion [DW05]. contagious [EL09]. Contain [JW01]. containing [HW09, KTP09]. containment [Mar09b]. contamination [GVB^{+}08]. contemporary [Hel08]. Content [Che03b, KMP03, KRN01, TGR^{+}00, WJMH00, ZSZ01, BSJ04, OR04]. Context [VO00, CSR^{+}05, KYS06, SRAL12]. contexts [Sch08a]. contextual [ST05]. contingent [Akt04]. Continuation [CD07]. Continuous [BM03b, IKD04, AD07a, BGE05b, BR04, BÅ03, CHR06, DCC^{+}08, DS04, EHG03, Ezo09, GXG03, GPN05, KAI08, LB07, LLW09, MDL04, Sch05a, SRG^{+}03, VA04a, YM04, Yan05, Yan06]. continuous-quality [BÅ03]. continuous-time [GXG03]. Continuum [KS09, KA02, WM00, APS06, SLHN06, TWO^{+}09]. Contractile [Yos03, TOB08]. Contraction [DSCD02, BFG05, GD06, GG09, LN05a, OD03]. contractions [RSB09, TSRB08]. Contrast [Möl02]. Contrasting [JAFW05]. contrasts [LD09b, WW08]. contribute [GOP09]. contributes [Bar05]. contributing [HL05, JAJH07]. Contribution [WCLL08, AGW^{+}06, AGW^{+}08, ES08, Toy09]. Contributions [Her00]. Control [AHCN07, CSC03, Car02a, CSM02, DFC^{+}02, HS01, Hav04, HP00b, HW01, Jam00, KY02, MW01, MCMS06, MCB07, OB04, PWK03, SKW00, VW00, WSC02, Wil01, BZ05a, BZ05b, BTS08, BB06a, BGMM08, BRND09, BdOP06, CP07, CCC08, CFCGCC03, CB07a, CCT^{+}09, DSS08, DMW04, FB08, FYX^{+}09, FLWB07, Fra08, GMK06, GB07, GGH^{+}05, GRH^{+}07, GN07, GC09, GZK06, HLA09, HRvdD08, IS03, JMvdB09, JJEX09, JITJ09, Kee05, LWC09, LMS^{+}08, MS08b, MWB05, NS09b, NTU06, NSI08, NAS07, NT04, OS06, OCA08, PCZL05, PCZL06, PNG03, RBW09, RGZ09, RB09b, SC03, SMH07, SJGK04, SHP09, SD06a, SD06b, SHI03, TBR08, Wes08, WS09, WLHB07, ZT06]. Controlled [Dai02, Rae02, KH09b, RvMK^{+}05]. controller [VPGA07]. controllers [CIV09]. Controlling [UM08, VP01, MSS05]. controversial [DHW^{+}09]. controversies [FP04b, Gol08]. Convection [KVMV01, LSMZ08]. convection-enhanced [LSMZ08]. convective [ES08]. Convergence [Rou03, DT06, Sta08]. Convergent [BLS^{+}09, ZJG03]. Conversion [HCC00, LBS00]. convex [TN04]. Cooling [See00, NKC^{+}08]. cooperate [NA04]. Cooperation [EL05, Gar09, GSB01, Har01, HNO03, INSR08, NSS08b, SL00, SR02, Ste00, Tor00, VP01, Akt04, BHC06, CH05a, GGPFM08, HMND06, HH06b, JG06, JB08, JLCS08, KNT^{+}09, KT03, LL08, LNH08, Men07, MS09a, ND09, OI06, OI07, OB08, SN09, SFVA09, SPS09, SSS08, Tan07a, Uit09, VSA07, Wak07, vV09]. cooperation-threshold [BHC06]. Cooperative [Har01, HB01, SD02, KAI08, KA03a, LVABV05, Sch09, STMH04, Toy09]. Cooperativity [JNWJWB04]. cooption [BR06]. coordinate [PDC04]. coordinated [BC04a, SZLM09]. cope [SHI03]. copies [OW07, RPB03]. copying [BFA08]. coral [AROS07, ILDP04, MHKS03, MHKS04, SBI07, wTA09]. Corals [MSI01, Mig06, SZC^{+}03]. Cord [BG00a]. cords [AP09, BFGS07]. Core [KTH09, IB06, KLN^{+}09]. core-temperature [IB06]. Coreceptor [RB02b, vdBWLS07]. cornfield [KT03]. corollary [SVN^{+}05]. correction [SIHH04, HH04b]. Corrector [ZSS02]. correlate [Bod08, OSJ08]. Correlated [AS00, OSBH02, WH00, CH05b, Fen03, Hie05, Jam04, PB06]. Correlation [Mad00, PCH^{+}05, BK05, GPG07, HK09, JWWS08, MLPJ09, NAV04, RRKF09, SP05a, YAL04]. correlations [Ded08, Ken07, PD04, Rob03]. correspondence [Nos06]. Corrigendum [AGW^{+}08, AMD06, CSP^{+}09, DBBW11, EPJ^{+}11, IvDHI08a, JCRJ07b, LH09, LGK^{+}12, Lei10, ML12, MT06a, ML08b, NI07a, Paw09b, PCZL06, PA09b, RH09a, RMAI09, SHI06b, SRAL12, SML08, TK09a, TRM03a, VGMM^{+}07a, VGMM^{+}08]. corrugation [KKPB09]. Cortex [Car02b, SGTF07]. Cortical [RWR00, FH07]. Cortically [FWLN04]. corticothalamic [RR08]. Corticotrophs [SRMW00]. cosexual [SS03]. Cost [BJ02, FBM06, GG01c, Loc08, Sel07, ELL04, FLS^{+}04, GR05, GNH^{+}05, JWH08, Nis06, RRS06, RBS05]. Cost-Benefit [GG01c]. Cost-effective [FBM06]. Costly [De 02a, GSB01, HB01, Jäg08, RON09, SN09, SS09a]. Costs [PAD00, BG06, KCHP08, NI04b, NI04a, Roe07]. Cotransporter [FTEG02]. Could [Gra07, Hau07, Möl02, Gol09]. count [CY09]. Counterpart [EN02]. Counting [Cha00]. Country [Tho00]. Coupled [DALP03, OCP^{+}00, PPR01, SLIL07, WKL01, BH08, BGO08, BTL08, Cin03, CSON^{+}05, DBB09, FB06, GP08a, HMN09, JEDH08, KA03b, Lew05, OF01, PM03, RB09b, Smi09, WW05b, WL04]. coupled-oscillator [Cin03]. Coupling [Bah07, BTS08, DLGC02, PL01, SI00, UHI02, öWO01, BTL08, CV08, GFWT04, KMC^{+}07, MSdlPS09, RSG09, SMCT08]. course [RKF06]. courtship [SS09a]. CoV [NLM^{+}08]. covariance [LFFT06, ST05, hZzGqX^{+}09]. CovR [MCB07]. CovR/S [MCB07]. Cow [HCMF01, HCB^{+}02, HCR^{+}04]. Cowan [MBB02]. cows [BKD^{+}06]. crab [SS06b]. Crabtree [Thi04]. cranial [MBB^{+}06]. Crassa [RVMR01, SG04]. crawling [ACD04]. create [Bul08, Kon06, PAA07]. created [KUK07]. creatine [FKAC06]. CRH [SRMW00]. CRH-induced [SRMW00]. cricket [MDCC06]. crisis [SF08b]. Criteria [BH01, BD08b, YTAK01, BB03, SJD^{+}09, Ten08]. criterion [KGB04]. Critical [APL08, CC01a, CF09, GHA03, NOT04, DLRP07, DRLB05]. Criticality [ORM03b, ORM03a, LBCL09, RKYH06, SJ03]. Critique [MGL06a, Zam03, Bak07]. Cro [BM04]. Crohn [Agi04]. Cronk [PKA02]. Crop [HB02a, EL09]. crop-livestock [EL09]. Cross [DSY01a, DSY01b, PGM00, Zah00, AB07a, CKM05, HSL04, LSMB^{+}03, RG05, RDC09, SP05a, ST03, STSD09]. Cross-bridge [Zah00, STSD09]. cross-correlation [SP05a]. cross-immunity [CKM05]. Cross-linking [DSY01a, DSY01b]. cross-pollination [LSMB^{+}03]. cross-protection [RG05, RDC09]. cross-reaction [AB07a]. cross-regulation [HSL04]. cross-regulatory [ST03]. Cross-resistance [PGM00]. Crosslinking [TH06]. crosslinks [FHD09]. Crossovers [HLW00]. crosstalk [BTL08, PP04]. Crowding [ORM03a]. crown [LBCL09]. crown/root [LBCL09]. CRP [Mac00]. CRP-Binding [Mac00]. crumpled [CMG06]. Crustacea [Hok00]. cruzi [DLM08, ISC01]. crypt [dT07b]. crystalline [GMMR07]. crystallization [SPN06, SN06]. crystals [FHD09, Lap03]. CSF [FBM06, FM09, ØKRG04]. CT [RGB06]. CTL [Wod01, WJ01]. Cucumbers [Kal00]. culling [LMS^{+}08]. cultivated [Thi04]. Cultural [MZ02, Ded09, SGS^{+}05, Sch09, Whi07]. Culture [GB02, Wal07a, EHG03, EG07, Gin03, HHBY06]. cultured [TOB08]. Cultures [GLV02, GCYH01, DBBW09, DBBW11, FVP^{+}07, GPN05, Har07b, PLG^{+}06, RNP04, SRG^{+}03]. Cumulative [Ste00, EG07, EBId09]. cupboard [For09]. curb [Mar04]. cure [MG09]. Current [GJE02, KFG^{+}02, CRL06, SH05, WCLL08]. Currents [TTKZ01, GKXS07, WCH^{+}09]. curse [KB04]. Curvature [IHV^{+}06, PM07, TL00]. Curvature-induced [IHV^{+}06]. curve [LDXW06, YSW09]. Curved [Lor06]. Curves [JL01, PPL^{+}00, SP01, Eng07, SLH05, YS07, ZC06]. Cutaneous [BSWM00, BKF00, BGE06a, BGE06b, Sil09]. cuticle [LDT09]. Cutoff [YZHZ09, Kat03]. CV [VL09]. Cyanidioschyzon [Di 08]. Cyanobacteria [GRG02, BdSFFDMC09]. Cyanobacterial [RGG00, TIM06]. Cyanobacterium [WAMO00]. Cybernetic [Kor01]. Cycle [CMS^{+}00, GRG02, KKDV01, LGB02, LBS00, MCWF01, MFB01, SM02, Tyr01, TN01, ABM04, Bel06, BFGS07, BB07a, Cal06, CS07, DBBW09, DBBW11, Iro09, Jam08b, Kal07, MB09, MCN08, NM08, NHTM09, NT04, PBC09a, PFRR08, RNSU04, RD07, SHP09, SG06, Ste04, iTYU06, VHF08, Voi03, YHM^{+}06, ZAB^{+}09]. cycle-specific [BFGS07]. Cycles [Ano01c, GG01b, HBLG02, HRBL02, JBP02, ZCC01, BPC08, CMF08, KJD05, KG06b, Smi08b, SH07, THKU07, XBT09]. Cycles- [ZCC01]. cyclic [AB07b, KNS05, KG06a, NMH07, RMF08, SSL08]. cyclic-3 [AB07b]. Cyclical [RKNK01, SMBM00, AM08, BBM03, BBM04, FBM06, CM05b]. Cyclically [Pál01]. Cycling [TGR^{+}00, FFIS07, PSWK09]. Cydia [TNTJ08]. Cylinder [GH02]. cylindrical [IHV^{+}06]. CYP2D6 [DFP^{+}08]. cysteines [JNWJWB04, SWB06]. Cysts [Fer09b]. Cytochrome [SPN06, LIK^{+}05, NS06, PCSL^{+}06, SN05b]. Cytochromes [FW01]. cytogenetic [COdAM05]. cytokeratin [PAVS03]. Cytokine [YBV^{+}00, YCS04]. Cytokine-modulated [YBV^{+}00]. cytometric [Mei05]. Cytoplasm [Mal02, Pic06]. cytoplasmic [CCT^{+}09, Dru03, TYW05]. Cytosine [NTK02]. Cytoskeletal [Lan00b, FKS07, KS09]. Cytoskeleton [CLO^{+}02, TM00, Wan00, LCL07a, LCL07b, LL09]. cytoskeleton-induced [LCL07a, LCL07b, LL09]. Cytosol [BGP00]. Cytotoxic [AB00, CDFP04].

D [Kin04, Pal08]. dab [KRGH07]. daily [Ale09]. Dairy [HCMF01, HCB^{+}02, HCR^{+}04, IGHW07, LGS^{+}09, TBB^{+}06, TBC^{+}08, XBCF05]. daisies [SA05a]. Daisyworld [ACL03, LL00, Puj02, SA05a, WAL06, WC07b, Ack04, ACLW03]. Damage [BL00, KKR^{+}07, Mar03, Bon06, FCD^{+}05, GVB^{+}08, MRGADD08, SB09c, THL03, YKG^{+}05]. damage-stimulated [THL03]. Damaged [KK00b]. Damuth [MGL06a]. Dance [Nir02]. danger [JBK04, SHI03]. Dangerous [MW09, WBB02]. Daphnia [DBGM08, FSGB^{+}02]. Dark [GRG02, Laz03, DVC^{+}04, GZL^{+}03, KG06b]. Dark-adapted [Laz03]. Darwin [Tor01]. Darwinian [DGZ07, DML02, LL00, LL01b, MHMGM08, Sta02, Sug02, TU00]. Data [CSFH^{+}01, CMTU01, LY02, MBB08, MJW00, PC02, SZ01, SHHD02, TGP^{+}00, AFD^{+}06, AAK^{+}09, BKM09, BGE06a, CLM07, DKS04, DSA^{+}06, GFW^{+}09, KLL08, LVL08a, LVL08b, LMT05, Luc05, Mei05, MKB03, RH09a, RH09b, RYAI06, SIHH04, SHP09, SVS04, SVGK07, SH08b, WR04, YD09, Zam03, ZLXY08]. database [YDL06b]. databases [Sch08b]. datasets [Sta08]. date [Sch08b]. daughters [KV05, Kan07]. day [Ale09, Cal06]. dC [HK05]. de-differentiation [HC07a]. De-epoxidase [LBS00]. deactivation [MGAR07]. Dead [KKDV01, SPVN06]. dealing [LP07b]. DeAnglis [LWC09]. Death [CPC^{+}00, RBN^{+}01, CL05, DdB03, HLW00, LCL07a, LCL07b, LL09, PBHS05, Sta09]. debate [Chi07]. debris [AR08, RAZ03]. DEBtox [BDMP^{+}08, BdSFFDMC09]. Decay [SS01, ZSS02, DP04]. decaying [LL05]. December [Ano00s, Ano00x, Ano01z, Ano01-31, Ano02r, Ano02x, Ano03-35, Ano03-39, Ano04-53, Ano04-55, Ano05-43, Ano05-49, Ano06-42, Ano06-47, Ano07-48, Ano07-54, Ano08-41, Ano08-42, Ano09-41, Ano09-46]. deception [BK01]. decide [RR09]. Deciphering [VdL03]. decision [APS08, AMD05, AMD06, CC06b, LP08, MAHD06, NDD08, RRH08, WA07, YCS04]. decision-making [LP08, MAHD06, NDD08, YCS04]. decision-proces [AMD05, AMD06]. Decisions [WH00, BFK08, HB09, LD09a]. Decline [DdB03, LCL07a, LCL07b, LL09, YS07]. declining [CCM07]. decoding [Tak06]. decomposition [FP03, PSPF07]. Deconstructing [Vin05]. decorrelation [SP05a]. Decoupling [Arc09a, GMK06, Syl06]. decoy [GSG^{+}07]. Decoys [Wil01]. decrease [VSA07]. Decreased [Her00, WG03]. deep [FS09]. Deeply [LBCL09]. defeats [KT03]. defection [Fis06, KT03]. Defective [Fra00b, KK00b, MB00]. Defectors [HB01, PB03]. Defects [LY00, MKD^{+}05]. defence [BBK04, BSR06, Gar09, PBH04]. Defense [FSL03, Fra00a, SI01, SI02, IS09, NGN^{+}04, SI04b, SI04c, ZJ06]. defenses [Jes06]. defibrillation [KC03, Kee04]. deficiency [ARR08]. deficient [BGG^{+}09]. define [OI04, OI05]. Defining [LBJE03, WMP03]. Definition [Kor01, MH09a, SLP00, TRM03c, PGF^{+}08, RvMK^{+}05, TRM03a, TRM03b]. deforestation [RKOS09]. deformable [MF05]. Deformation [Bar01, DD02, Wan00]. degenerate [DB05]. degeneration [Rǎd09]. degenerative [NGN^{+}04]. Degradation [KK00b, LGB02, KNT^{+}09, SS05a]. degree [Est07, GGPFM08, RMAI06b, TY06, vdBR04b]. degrees [Eti09]. Dehydroascorbate [RBBH02]. dehydrogenase [DSA^{+}06]. Deinococcus [SB09c]. Delay [EMS02, LGE00, AM04a, DP04, JE09, SG04, SG06, WWS08, XM09]. Delayed [KK00b, ZRSK07, AWW06, EHG03, GZL^{+}03, MS09b, RDYH09, RSG09, WCLL08, XBT09]. delays [ACR06, LKM06, MC06]. Delbrück [Fra03]. Deleterious [Kaw01, Pál01, Arc05, Yam03]. deleteriousness [KMC^{+}07]. Deletion [MB00]. Deletions [HR00]. delicatissima [SWM09]. delivery [BFGS07, LSMZ08, OBL04, TGV^{+}07]. delta [MRJR09]. demand [RH08]. Demands [AøP03]. deme [LL08]. Demographic [AK09, AROS07, CLMMP06, DDJ06, MM09]. demography [DHR^{+}07, MSdlPS09, NI07b]. Demon [Hop02]. demonstrates [HHP05]. DEMSIM [DGM05]. denaturation [OQGC07]. Dendrite [HGV01, Nie06]. dendritic [GvO04, MRJR09, NA04, RWP^{+}08, RKH^{+}06]. dengue [AB06]. Dense [Wil01, BCKE^{+}08]. Density [CLZZ02, HI00a, JJ01, Jes06, PJ01, dV06, Ama04, AD06b, CPM^{+}09, DMW04, DFSD05, ELSFB07, EH08, FPL03, GEF09, KRNG08, MJ07b, MBJ09, NAV04, PRT04, RYAI06, Sin07, SSL08, Sta00a, Sta00b, Sta00c, iTYU06, TMBD09]. density-based [DFSD05]. Density-dependence [dV06]. Density-dependent [HI00a, Jes06, PJ01, Ama04, FPL03, GEF09, Sin07, SSL08, TMBD09]. density-independent [ELSFB07]. Dentition [Lin01]. depend [KW07]. Dependence [DFC^{+}02, HGB^{+}00, LJW02, MGL03, MFB01, APL08, Gos06, JEHK06, Pep04a, Pep04b, PC09a, SRR08, SFC^{+}09, WR03, dV06]. Dependencies [GS02a]. Dependency [BG00b, AFD^{+}06, CSR^{+}05, SRAL12, Uit09]. Dependent [EMS02, LPLC00, RFH^{+}02, YK03, YLL00, AB06, Ama04, BPC08, BFR05, BSS^{+}07, CGH01, CHN08, DBR^{+}05, DBB09, FMI05, FI00, FPL03, GS08a, GEF09, HI00a, JE09, Jes06, KGL09, KGG08, MS06a, NWP07, O’K05, PJ01, PB08, RSH^{+}06, RKRW08, SS03, Sat04, SDS04, SV05, SSR04, She06, Sin07, SSL08, SX06, iTYU06, TTT09, TL02, TMBD09, Tyr01, VFC^{+}09, VdL03, WDH^{+}09, Wod01, WVA05b, YM05a, YM05b, YL00, ZXWF08]. depending [RG05, SF04]. depends [GABK08]. dephosphorylation [Ano01c, GG01b]. Depletion [RGG00, CDB09, HS07]. depolarizations [DSS08]. depolarizing [Dim05]. deposition [CTB09]. depression [FDA^{+}09, MZK08, RSB09]. deprivation [PR08]. Derivation [Rou03, RRC^{+}06]. Derived [KMH00, DBBW09, DBBW11, JG08, RWK08]. Describe [Has01a, LMT02, PN01, BGE05b]. described [GMMGD^{+}08, MY09]. describes [SVGK07, WH07]. Describing [DO01, LBS00, Pie09a, GVB^{+}08, GT06b, KI04, KI05a]. description [BFGB04, BFG05, KRB05a, NM06, PBvdG09, PD04, SP07b]. descriptions [Mat06]. descriptor [TLZ^{+}08]. descriptors [BRCB04, MPOBD^{+}09]. desensitization [LGB03a, LGB03b, TLZ05]. desert [ACL03, MRA06, PM03]. desertification [LBCL09]. desertion [SE07]. deserved [SS06d]. Design [Ace00, KSM02, LFW02, Pel00, SJD^{+}09, TO06, AF09, BLOL07, CA09, DMPS05, FMHW08, GLSW07, HDW^{+}09, ISW04, IvDH^{+}08, ITI^{+}09, JAV07, LWFP08, LH06, RSSM06, TGT06, Vin05, WRKK09, WZW^{+}07]. designed [HM05]. Designing [Noe00]. despite [Sch06]. Destruction [Her00, SFS^{+}01, ML08a, ML08b]. destructive [RB09a]. Desynchronization [CMTU01]. Desynchronized [PC09b]. detailed [YAL04]. details [HLA09]. detect [BH04, BIS^{+}07, Ger09, HL05, HHH06]. detectable [HRZ06]. Detecting [HS04, YDFQ05, MGT^{+}06]. Detection [CWDM06, CACC02, JLS01, Noe00, LRB^{+}06, MSDM06, RP09a, YTL08]. determinant [Kli06]. Determinants [KD06, NSH^{+}03]. Determination [Alb08, Hol06, Rei02b, BGF03, CWJ07, Fli05, FFHK09, GMMR07, JAJH07, LVH04, MSP03, TSRB08]. Determine [CSFH^{+}01, YTAK01, WFGP04]. determined [Kur08a, VL09]. Determines [DFP01, Wan00, DO04, Lin04]. Determining [IvDH^{+}08, YLCW06, GMdA^{+}07, Mit04]. Deterministic [EN02, Lor06, PGM00, SYSY02, BLF^{+}09, GKK06, JG08, KRB05b, Man06, SD06a, Sim08, Wit03, WT09b]. Detours [Ale01]. detrital [MB07]. detrital-based [MB07]. Deuterium [PBL06]. Deutsch [Rot08]. Develop [Yos03]. developing [KF04]. Development [IV02, Lin01, LH01, Mac00, MAB00, Pin00, RFH^{+}02, SMG^{+}03, VW00, ZAD07, BKMH07, BMS05, CLHW07, CGS05, Dru03, ECAV07, FM04, FK03c, GT08, GZ04, HMB^{+}08, HM06, HIM09, HKP07, JMvdB09, LPJB^{+}08, LPT05, MSMKM06, Osb08, PBB03, Rej07, RLCIB05, SK05, Sch08a, WRG^{+}04, Has01a]. Developmental [BFK08, PKL01, PBT02, CWJ07, IS08, Ker04, Ker06, RKH08, Sel06, Shi06a, VTL05]. Deviation [NA02]. deviations [GKXS07, Gra07]. Devonian [KRNKH00, SKY09]. dG [HK05]. DGYW [KS08b]. DGYW/WRCH [KS08b]. Diabetes [TPD^{+}00, KSEK09, MKD^{+}05]. Diabetic [CCP^{+}00, RBBH02]. diagnoses [Izs05]. diagnosis [CFCGCC03]. diagnostics [Ano05h, Hof04]. diagram [RNP04]. dialects [SGS^{+}05]. Diameter [HVPN02]. diamondback [NLS08]. diapause [MM08a]. diaphragm [LKM^{+}08]. diarrhoea [GLE^{+}09]. Diastolic [BBSLN08]. Diatom [LH01]. diauxic [NP07]. Dichotomy [Kun06]. Dictated [DSCD02]. Dictyostelium [KI04, KI05a, UI02, DO04, UI04a, UI04b, VW03]. did [Gab06]. Diel [HS01]. diet [RYAI06]. Difference [TMI03b, DFP^{+}08, WW05b]. Differences [KRN01, Wel00b, Lap03, MWCS04, NA03]. Different [LJ09b, PN02b, RGG00, Tay00, CC09b, DP08, ETH04, Eti09, FDA^{+}09, GSRC^{+}06, GGH^{+}05, GEF04, MBP07, Mig06, PHdB09, RNS04, RRKF09, SF04, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, WT07, YZHZ09, vVH07]. Differential [Hog00, RGG00, RDSB^{+}03, DB08, FPM^{+}06, HF03a, KSK08, MGHF06, PBB03, SH03, SBGA04, UI04a, UI04b, ZJG03]. differentially [YDFQ05]. differentiated [WCH^{+}09]. Differentiating [FMF^{+}00]. Differentiation [CMW02, Coo01, For00a, FK01c, Hog00, PD01, VP01, CD05, DHYHR09, HC07a, HH05a, LWRK07, NK08, Ots08, QO08, RDYH09, Ros03, YCS04, dT07b]. diffusely [WMS08b]. diffusing [Kut05]. Diffusion [BSWM00, BCL08, CLZZ02, CC01a, CC08, DNS00, LL09, Mat06, SCGFS00, SIMK02, Sta00b, TLA00, TGP^{+}00, AC04, BFGB04, Cin06, DALP03, FM07a, Gie03, Gie06, KL09, MCK07, MW07, MS06a, MSMKM06, PD06, QXDW06, Rel04, RB06, Saf05, TK05, TN04, VGBA06, WWS08, WO04, WDH^{+}09, WB06, XM09]. diffusion-constrained [AC04]. Diffusion-limited [BCL08]. diffusion-reaction [Gie03, VGBA06]. Diffusional [LK08c]. diffusive [NPN09, TKM06]. diffusivity [Kin04]. Digestion [LJW02, VTC08a, VTC08b]. Digit [MMTS02]. Digital [Ano00a]. dihedral [WZ08]. Dihydric [CCGC02, Ril00]. Dilemma [EL05, Ezo09, HS02a, HNO03, IKD04, KK00c, LBSS02, Nei01, VMW02, AD07a, Arc09b, CJLS08, JLCS08, LB07, LW04, MS09a, Sch05a, WR07a, WVA05b, WL07b]. Dilemmas [HB01, HMND06, Hau06, MP09b]. dilute [IPY07]. dimension [Ben04b, Niw05, QWQ09, UCSZ07]. dimension-to-biomass [Niw05]. Dimensional [GBG01, KTH^{+}00, ACLW03, AO05, BZ04, CD05, CHR06, DSU^{+}04, HTN04a, HTN04b, HQP^{+}09, Iim07, LMT02, MBB08, RA08, SIK03, SI05, wTA09, TGV^{+}07, UCSZ07, WDH^{+}09, WLFC08, ZAD07]. Dimensions [Dus01, MV02a, Zam01, Bye09, DBGM08, TQUN07, TQUN08]. Dimer [MS01a, CY08, Kli06]. Dimeric [DBHS00, HQP08]. Dimerization [BHJ03, GS02c, Daw09, SYYI07, WSM05, WL04]. dinoflagellates [DS05]. dinucleotides [Mic07]. dioecious [PM09]. dioxide [Gam06]. Diploid [GS00, YL00, HA09, SP05b, YY06]. dipole [LN05a]. dipteran [BEBV03]. Direct [ON07, RON09, WH01, BS06, PTON08, SA07b]. Directed [Arm01, Fox05, LTW06, MM00a, ANL01, MNP06, MLPJ09, RAHO06, WRKK09, WS02b]. Direction [KG02, SI06, YMC01, LDW05]. Directional [JP00, JSCN04, Wit03]. Directionality [Mas09, SIMK02, MBRRI08]. directly [LPJB^{+}08]. Disadvantaged [HLW00]. disappearance [BM09, KCV05]. disc [Won05]. disc-sphere [Won05]. discoideum [DO04, VW03]. discontinuities [KMGDG04]. Discount [Ste00]. discounting [HMND06]. Discovery [XWD^{+}01, JPJ06, YZHZ09]. discrepancy [LRT04]. Discrete [BM03b, DS04, GLV02, KKK00, Tas02, TS02, YCC02, AM04a, Dam04, DGM05, GK04, HC07a, MSdlPS09, NLS08, SN07]. Discrete-time [YCC02, GK04]. Discriminant [Lin08, JHJ^{+}09b]. discriminate [SPE08]. Discrimination [For01, MAHD06, Wil06a]. discriminator [BS06]. discriminator-Errors [BS06]. Discussion [NI07b]. Disease [ACCC02, BW01, EKS02, FPL03, HS08, MLPJ09, SMBM00, Agi04, AMC^{+}09, BSGT08, BB06a, BSS^{+}07, BH03b, Bul06, BMS05, CTS06, DMH08, DHR^{+}07, GBB08, HRvdD08, ITM^{+}07, ITI^{+}09, JMvdB09, JEDH08, Lan03, LT06, LZ09a, LMS^{+}08, MP09a, MGT^{+}06, OS06, PCB07, PRT04, RPNH03, SJ03, TK09a, TK09b, Yan08]. disease-control [OS06]. Disease-induced [HS08]. Diseases [GG03a, CBC^{+}09, DHRM08, FYX^{+}09, FMLP06, GWPW04, GOP09, HDF04, JJEX09, KYZ^{+}08, Mat06, Sch03, TP05, dM09]. Disentangling [Her09]. disinfection [ES08]. dismutase [GMFS06, KLK06]. disorder [BSGT08]. disordered [BE08]. disorders [RKF06, dBCP^{+}07]. Disparate [GK00, BTL08]. Dispersal [CR01, KY00, YCC02, YALT00, Ama04, Ama06, BBKGP08, BNRW04, BTCD07, BH03b, CG06, Eti09, FM04, HWCF07, LSMB^{+}03, MCM^{+}09, MC06, MA03b, MBRRI08, NBT07, SSL08, SSS08, TMBD09, WW05b, XBGN05]. dispersal-driven [WW05b]. Dispersing [HLH01, Mcn00, MN01, Hie05, HM07]. Dispersion [QB00, KNS05]. displacement [GAK^{+}06, GNLEK07, KC07, YLCW06]. displacements [CGF^{+}08]. Displaying [Ano01c, GG01b, Leh02, LRD04]. Displays [Szá03]. disrupt [HHP05]. disruption [LDW04]. disruptive [BdABA09]. Dissecting [MM08b]. dissimilarities [PDC04]. dissipation [CQLV^{+}03, LPT05]. Distance [OKS01, PE00, AB06, BTCD07, CF07, GMMGDA05, GMY09, MN01, Mic07, Mur07, TH03a, TLC07]. distances [BGG06]. Distinct [KP06, MWB05]. Distinctive [Dus02, RHM^{+}08]. distinctiveness [MR07a]. Distinguish [DOT02, NA03, QLHL09, ZYD^{+}05]. distinguishing [ZW03]. Distorted [YMC01]. Distortion [Che00]. distributed [CMF08, FLG^{+}07, JI05, OT09, PPD09, WMS08b]. Distribution [AH03a, Ano01g, Bea00, DFC^{+}02, FPS08, IKS00, Ken01, Ken02, Koc00, LLCM01, Mcn00, QB00, RH02, SRN^{+}00, WGH00, YTAK01, Ait08, BH08, Bal04, BGK09, BM06, Eng07, Est07, FR06, Fra03, GMMR07, GP08b, HF03a, HL06, Hui09, Kok04, LS04a, LBS06, LS08b, MS08a, Mil08, Orr06, PM08, RDYH09, Saf09, SH03, SVGK07, SNA^{+}08, TBCD06, Wal07b, YB05, ZF08, Zhu09]. Distributions [DBHS00, EMS02, MZ02, YN02, BLZ07a, BLZ07b, BSJ04, Bye09, Chi07, GR08a, GNLEK07, HEH^{+}09, HSF09, Jam09a, LP07a, LP07b, Niw03, RMAI06b, SPAH06, TY06]. Distributive [MGL06a, BDMR06, BES06, MGL05]. Disturbance [OKTS02, BR00, CBR04, CCF06, NBT07]. disturbances [CCC08, Hie05, HM07, wTA09]. Disturbed [GS02a]. disulfide [FW00, SWB06]. disulfide-bonded [SWB06]. Disulphide [WP01]. divergence [FS09]. Diverse [GVK00, PFRR08, UB01, Wil01]. diversifies [KOK06]. Diversity [BLMV01, GS02a, LJK06, LL01a, OKTS02, Sch00, WVA05a, Ari05, COdAM05, CCF06, FPS08, Gre09, HFS06, KUK07, LV08, MPN^{+}05, Moc08, MSS07, MS06b, NdGG06, RHM^{+}08, SMPvdB08, TK05, vdBR03]. diversity-stability [LV08]. Dividing [KK00b, LFC04]. Diving [HHR01, HF03b]. Diving-induced [HHR01]. Division [Wah02, DW08b, HEH^{+}09, Paw07b, Paw09b, SZLM09, Tan07c, YHM^{+}06]. DNA [AAL08, Aba09, AFZ08, AJSL07, AF01, BL00, BZ04, Bie06, BM03b, CSC03, CRL06, DB01, DVL^{+}00, DBHS00, Dru03, EN03, FCD^{+}05, GS06, GVB^{+}08, Gru00, HK05, Has01a, KmMK04, LCBTP03, LDW05, LTW06, LDXW06, MS01a, Mei05, MM00a, MM02b, MK03, Mur07, Nar07, NA02, PCH^{+}05, PBMU^{+}09, PSSY09, PN02b, QWQ07, RZF03, RPB03, RAHO06, STW^{+}09, SB09c, SLL06, SK01, TML02, TL00, VO00, VA04a, Wal07b, Xie09, Yag09, YY07, YCC^{+}06, YSW09, ZSZ^{+}06, ZC06, ZXWF08, ZAB^{+}09]. DNA- [STW^{+}09]. DNA-binding [YCC^{+}06]. DNA-membrane [RZF03]. DNA-Proofreading [Has01a]. DNA/collagen [PSSY09]. dnaQ [Dor03a]. DNAs [MB06]. Do [Ded09, FB00, FM07b, Kin07, Möl01, PD00, WSC02, YOYT07, AFZ08, BIS^{+}07, BKF00, Den08, HHH06, LY03b, PBC09b]. Does [Dru03, GDD^{+}03, Kon06, NBT07, Sug02, Tan07a, VSA07, Yos03, BBD06, BB05, BEBV03, CK08, EG07, GMFS06, MV06, MJ07b, SS00, Tan07c, VA06, WKB07]. Dollar [Yos03]. dolphins [BLT03]. Domain [CLXC03, DSY01a, DSY01b, SHF02, Ari05, BL09, CC06a, Di 03a, Di 08, JBK04, KS08a, LGB03b, MBB08, MSMKM06, PCB07, RN07, SC09, SCS04a]. Domains [XWD^{+}01]. Dominance [BC02, Bro02, HR00, Agu08, LKK07, PB08]. dominant [GV03]. dominant-recessive [GV03]. dominated [MGC04]. dominates [AB04]. Dominoes [Ano01c, GG01b]. Don [LHDvdM04]. donation [Mar09a]. Donnan [Kur08a]. Donut [LZ09a]. Donut-shaped [LZ09a]. Dopamine [PB02, SM09a]. Dormancy [KY00, ITM^{+}07]. dorsal [ZLN07]. dorsal-ventral [ZLN07]. Dosage [HR00, Vei03, RCS05]. Dose [BNT^{+}00, BFGS07, Hua03, MV07, ØKRG04]. Dose-Response [BNT^{+}00, MV07]. Dosing [HFH03, Cog06]. Double [CMB^{+}01, AO03b, HK05, KS08b, Lan03, PDC04, Wal07b, WBSY06]. double-agent [Lan03]. double-motif [KS08b]. double-strand [Wal07b]. Double-stranded [CMB^{+}01]. Doublefoot [MSMKM06]. doublet [Pat05]. Douglas [AF02b]. dove [AKdlPP06, NS03a]. Doves [Cro00]. down [BTA08, CF09, DFW^{+}07, PT09, SB05]. down-modulation [SB05]. downwelling [HHH06]. Doxorubicin [Jac03, EPJ^{+}09, EPJ^{+}11]. DQN [mLLS^{+}06]. Dr. [Gel07]. dragonfly [KKPB09]. Drastic [Wak04]. dreaming [Muz05]. Drift [GA02, LACL03, LFFT06]. Drive [De 02a, Blu07, Mar09b, WKRD09, ZXWF08]. Driven [HHR01, HRBL02, CABB09, EL05, FFD^{+}02, FWE06, HDW^{+}09, KUK07, LCL03, NOT04, Ped07, Rel04, RWP^{+}08, SBZ^{+}08, VL09, WW05b, ZW07, vdBWLS07]. drives [LWFP08]. Driving [Pro03]. Drosophila [AWAB05, AD06a, Ano01b, AGZ^{+}06, HMB^{+}08, KHHS09, LM07, LVABV05, MSP03, OY03, RCS05, ST01, UHK01, XK07, AO03a, LdGH09]. Drug [GABK08, RB02b, Wod01, ADHM09, BFGS07, DLRP07, DF09, DHW^{+}09, Gar02, JEHK06, Kom06, KW07, LPJB^{+}08, LSMZ08, LYZ08, OBL04, Pep04a, Pep04b, RGFP07, STK08, Smi08d, VHF08, XAP07]. Drug-efficacy [GABK08]. drug-treatment [LYZ08]. drug/metabolite [DF09]. dT [HK05]. Dual [AWJ02, Ezo09, CXM^{+}09, FB06, QWQ07, QWQ09, TBCD06, WE06]. dual-lattice [CXM^{+}09, TBCD06]. dual-recorded [FB06]. Duchenne [DC09]. duct [MV02a, MV02b, MV02c, MV02a, MV02b, MV02c]. ductal [BMH07, FBUL05]. Ductile [FHD09]. Ducts [MV02c]. Due [KY00, BGE06a, BGE06b, BFG08b, BFG08a, HD08, Lac01, NGT05, Shi06a, TTKZ01, Wal07b]. duplicates [OW07]. duplication [BLS06, PSSS03]. duplications [KK06]. Duration [SS09a, CY07, DMO^{+}07, Dus06, Hop06, Jam08b, JEDH08]. During [ISC01, JW01, Nir02, AGZ^{+}06, ACLH05, BFGD07, BGG^{+}09, Buc04, BRND09, CXZ^{+}09, CLB05, CPC^{+}00, CS07, CSD09, DVC^{+}04, DBG06, Dru03, FL03, GGH^{+}05, GZL^{+}03, GK09, GRBR04, HMB^{+}08, HM06, HHH06, Jam09b, KP06, LS03, LW08a, LHDvdM04, MSMKM06, MPL06a, NB02, OLBM08, PL09, RWCK08, RWP^{+}08, RGRB04, RRH08, RLCIB05, Ros03, SJGK04, SPC02, SGG^{+}07, SCC^{+}00, TT04, The00, UL06, WBR09, YFKP03]. dusk [HHH06]. dusk-active [HHH06]. DVM [LSH06]. dwarf [YOYT07]. dyadic [BC04a]. dyes [Mei05]. Dynamic [BC04b, CXZ^{+}09, CMPL^{+}00, CCC08, FR02, GZK06, HIM09, JG06, MSWH00, MKE^{+}09, NSM02, Ort06, SZ04, SL00, SI04b, VCBV^{+}06, WSS07, XWD^{+}01, YLM03, ZC09, dVG04, dVG06, BKM09, BD08b, GDC^{+}06, HM07, IB06, KNWCB07, KH07b, LR07, LAG09, MAC06, MKB03, NLM^{+}08, PB07, SGTF07, SK05, SHI05, SHI06b, STMH04, TSZ^{+}07, VAAH05]. Dynamic-persistence [JG06]. Dynamical [ACLH05, AR08, BBPSV05, BLMV01, ETTV08, FC01, HLH01, ITM^{+}07, IKS00, MHDOG01, PWG09, TK05, AVSHV04, CBF05, HSC07, Hor08, LGCL07, MBB^{+}06, MP09b, Paw09a, QHF^{+}07, vKPdR07]. dynamically [DBB09]. Dynamics [AWJ02, AP01, Ano00-30, Ano01c, Ano09a, ACCC00, Bas01, BKE03, BH01, CLZZ02, Car02a, CCP^{+}00, COS01, Dai02, Dam04, DTG09, DSCD02, DF00, DHM01, EN03, EMS02, FMD01, Fra00b, Fur02, GG03a, GG01b, HM01, HDHS02, HDRM00, JJ01, JSV02, KMH00, KTH^{+}00, KRR09, KS01c, KNN01, KN03, KPS02, KJ02, LPC06, LLF02, LVV^{+}01, MKLD02, MI06, MA07a, MSI01, PN01, PN02a, PKL02, PG00, QA01, Rae02, RSBY03, RRR^{+}08, RSG09, SSA00, SB09b, SCH02c, SMG01, SE02b, SVS^{+}02, SHHD02, SLL06, SSF09, SRWL02, SB00, TT02, TT03, TT01, TBB^{+}06, VS08, WWS08, YCC02, YALT00, vAGDR09, vdBR04a, AAEW09, AJOK09, AM04a, Ama04, AKS07, AM04b, AM06, Ark05, AB03, AMC^{+}09, BH08, Bea06, BTS08, BKE04, BB07a, CMF08, CSD04, CKS07, CAHH06, CRR08]. dynamics [Cog07, CSM05, DLRP07, DHYHR09, DGM05, DGD^{+}09, DKLL05, DHP06, DBB09, Eam06, EZM05, ETH04, FPBM08, FMI05, FP07, FDA^{+}09, Fra08, FPL03, FJBK05, GK06a, GDS07, GSRC^{+}06, GASA09, GK04, GPMW06, GS08a, GT08, GPN05, HCM^{+}07, HK09, HDZ^{+}07, Has06, HA09, HTN04a, HTN04b, HPZ09, HBB08, HS07, IYGA08, IiKY07, IB06, ING04, IGHW07, IMN04b, JM07, JB04, JEDH08, JZST09, JR06, JST^{+}08, KBT08, KBD06, KI09, iKIY09, KHHS09, Koh07, Kom04, KiHM^{+}03, KRN03, KC05, KV04, KCVC04, LL08, LCL03, LCBTP03, Lap03, LB07, LCHK09, LGB03a, LGB03b, Lew05, LL07, LDW04, LDW05, LTLM09, MOBN09, Man06, MCM^{+}09, Mas08, Mas09, MS08b, MB05, MGL^{+}06b, MNI06, MP05, Mil05, MK09, MM08b, MNL^{+}07, MRF07, MS09b, MDL04, NWT09, NO04, NBMS06, NLS08, OLBM08, OI04]. dynamics [OI05, OI07, OBN07, PAA07, PW09a, PCS^{+}06, Péc05, PDPL05, PHG04, PFRR08, PRT04, PB09, RB06, RDL07, RR08, RWK08, RRS06, RGFP07, RRR04, SSB^{+}07, SBI07, SLIL07, SI04a, SI06, SW08b, SWRH03, STK08, SKY09, SRG^{+}03, Sim08, SSL08, SV03, SML07, SML08, SM09a, TIN06, TC09, TSN05, TPN07, TSC04, VWR07, VCGV07, Wak07, WZL^{+}08, WG06, WT07, WLHB07, XBCF05, Xie09, XFAS06, ZAdlPLM07, ZZW07, ZCD05, vLHH06, MV02a]. dynein [Cib08, GE05]. Dynnik [NM08]. dysregulation [Agi04]. dystrophy [DC09].

E. [AHT^{+}07, CB07a, LR07, LSD^{+}00, LLO08, SWB06, TBC^{+}08, WDH^{+}09]. each [BZ04]. earliest [NDE06]. Early [KRNKH00, PGLG01, Wel00a, Wel00b, Zhi02, AWAB05, BKKR08, Dru03, EAC^{+}06, Gab06, GT06a, GT08, HIM09, HKP07, Kal07, MCK07, MCM^{+}09, MSDM06, Rej07, vdGMG^{+}09]. early-stage [BKKR08]. easier [SSM09]. Ebola [CHCC^{+}04]. eccentric [TSRB08]. Eccentricity [FK01b, FK03a]. ECG [ACLH05, JDMZ^{+}07]. Echinodermata [Kal00]. echocardiogram [Han04]. echocardiogram-based [Han04]. ECMs [FS04b]. ECoGs [OR04]. Ecological [BLMV01, Di 00c, Dus02, FPC01, KC07, MS08b, BFP07, CB08a, CE05, DP08, FCP03, GDS07, GN07, HBWB08, LL07, MCM^{+}09, NWP07, PCS^{+}06, RAZ03, SLIL07, SH07, VWR07]. ecological-social [SLIL07]. Ecology [CDHJ02, Kee00, AJ05, BR08, RE04, RB08]. Economical [DBG01]. Ecosystem [Tsc00, AB04, FFIS07, GvHP^{+}07, JD09, LPT05, SWN07, SD06a, SD06b]. Ecosystems [CMW02, Gro02, LLF02, Wil01, BLMV05a, BLMV05b, GJ07, MGC04, MB07, Mic05, Sav04]. Ecotypes [SKI^{+}06]. ectotherms [DF08]. edge [LRB^{+}06]. edge-detection [LRB^{+}06]. editing [DL08]. editor [Ret04]. Editorial [Ano00b, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano04-29, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30]. Editorial [Mil09, Ano04j, Ano06g, Ano06h, Ano06i, Ano06j, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. editors [Gel07]. Edmund [EK08]. EEG [GBZ06, RRR04, UL06]. EEGs [OR04]. Effect [BL08, BB02, CLZZ02, EW09, FM02, GB07, GP00, GP01, HBLG02, KY00, KT01, LL05, MCF04, MLMW01, NWT09, Nar07, OMT03, RVMR01, SM06a, See00, SN09, TZS07, WGH00, Wod07, ADHM09, AP09, BFGS07, Bog04, BEBV03, BFA08, BC06, CC09a, CCF06, DDJ06, GR08a, GASA09, GC08, GT06b, HRSV06, HM07, ING04, JMvdB09, KNS05, Kat03, KDKS06, KC03, LIK^{+}05, LFFT06, ML07a, ML07b, MGP^{+}08, Mar09b, MTS05, MGHF06, MPL06b, NPN09, ØKRG04, ØØS06, PDB08, PLSGG05, SES08, Sin07, SSS08, iTYU06, Thi04, TBC^{+}08, VGS^{+}05, Wak04, WWS^{+}06, XSD^{+}05, XBT09, ZAB^{+}09, vKdRP05]. Effective [Bye09, TH03b, BC09b, FBM06, HP04, KMW03]. effectively [Tak06, Tan07a]. Effectiveness [TY06]. effector [HW06]. Effects [ABR02, AMC^{+}09, BPV01, BZ00, CC01a, Che06b, Cog06, CPR00, DMH08, DFC^{+}02, Fen03, FM04, FR02, GMAH07, GTG01, GPG07, GF02, HS00c, IKD04, KUK07, KKPB09, KA02, LFLM00, LL07, LST00, Pál01, PGLG01, PN02b, Rev02, SCH02c, SZLK^{+}05, SWI07, Ste04, ST05, SFS^{+}01, Tay00, Vei03, WH00, WH01, WLHB07, YALT00, AAL08, ABM04, ARR08, AD06b, BRCB04, BMH07, Bar05, Bea06, BR04, BGMM08, CMF08, CHCC^{+}04, CAHH06, CH05b, CBC^{+}09, CM03, GVB^{+}08, GT06a, Gol07, GvO04, HL05, Hau06, HL06, JK03, Jes06, JCK09, JSCN04, Ker04, Ker06, KHHS09, KLK06, Kur08a, LKK07, LMF08, LKC07, LHFH08, LFP^{+}05, LW07, MKS^{+}09, MBD08, MK09, ND04, OBPH^{+}08, Orr06, PLB^{+}05, PB09, ROR05, RSS04, RCS05, RA06, SSB^{+}07, Sac07, SS03, Sar04, Sat04, SWN07, SSL08, SX06]. effects [USTG09, WZL^{+}08, YHBW04, ZJ06]. efficacy [BRS^{+}09, DLRP07, GABK08, vdBR03]. efficiencies [GG09]. Efficiency [Puj02, Lal06, LW08a, LMVPM07, MT06a, MT06b, VTG^{+}06, VSP06]. Efficient [Arm01, DSS08, DBG01, CB07a, DMPS05, KD03, MP05, MLS09, SHI05, SHI06b, ZKH^{+}05]. efflux [FM06, MGP^{+}08, NPN09]. Effort [YYI03, HDZ^{+}07, SS09a, YHI03, YHI04]. EGF [AD06b]. EGF-receptor [AD06b]. Egg [Lan00a, Nir02, CH07, FWLN04, LHDvdM04, ZAB^{+}09]. Eggs [Dus02, Yos03, WB06]. Egocentric [MRA06]. eight [OI06]. Einstein [AH03b]. ejaculates [HH06a]. Ejection [KKDV01, AKR09, Han04]. Elastic [FS04b, AGW^{+}06, AGW^{+}08, GG09, KSG03, RBS05, TOB08]. elastic-force [KSG03]. Elasticity [Nie02, Tho05, FH07]. elbow [GMSW04]. Electric [LN05a]. Electrical [ARW00, RWR00, SRMW00]. Electrically [BGP00]. electrocorticographic [Rob03]. Electrodiffusion [GJE02]. electroencephalographic [OR05, Rob03]. Electrolytes [SS00]. Electromagnetic [LFLM00, TK04]. electromechanical [Mas03]. Electron [FW01, MS01a, PDIS00, Ban06, Bat09]. Electronic [CMW02]. electrophysiological [ACLH05]. electrophysiology [GKXS07]. electroporation [PGN08]. electrostatic [AAL08, SPN06]. elegans [MY01, MGS08, NM09, RW08, RDSB^{+}03]. Element [JMB00, SNT03, BBSLN06, DGS09, LME06]. elementary [LP07b, WM04]. Elements [Mac00, PKW^{+}00, QA01, KSS07, RSG09, YMLK04]. elevational [Sav04]. eliminated [EPR07]. elimination [PER03]. Elongation [HGV01, Hau07, OLBM08, ZZW07]. Elucidating [JBJ^{+}08, vLBJK07]. elusive [Mit04]. Embolism [HVPN02, HVN07, KRN03]. Embryo [PMMS01, AWAB05, ST03, ZLN07]. Embryonic [ZSRB07, HMB^{+}08, Ros03, SNCM09]. Embryos [KMGV00, KVMV01]. Emergence [AP04, GD08, JLCS08, KTCD00, MGC04, NI07b, Pro03, RWR00, RC09, Dal06, GA08a, GSM06, HH06b, LL08, MS09a, SH08a, TM06, Tlu07]. Emergent [CCM07, DS00a, GTDA02, ORM03a, WBR08, CSP^{+}08, CSP^{+}09, vKdRP05]. Emerging [MKLD02, FYX^{+}09, GDS07, Tan07a]. emission [BKM09, KY03, MKB03]. emitting [GSM06]. Emory [Ano01-33]. empathy [Fis06]. Emphasis [HCMF01, HCB^{+}02, LM07, SB07]. Empirical [Gin00a, KK00a, QSAL08, EHG03, HLP06]. empty [Lio09, SN09]. encapsulated [GCS07]. Encephalization [Cha03, JCW^{+}03]. Encephalopathies [CPC^{+}00]. encephalopathy [BKM09]. encoded [BMT04]. Encoding [LMT02, LLS^{+}09, Wal08]. Encounter [Dus00, Dus06, LBS06, Lew03]. Encounters [BB02, Nak01, Nak03, UCSZ07]. endemic [CMW08, dM09]. endemically [LGS^{+}09]. Endochondral [Cub00]. endocytosis [TWE04]. Endogenous [EM00]. Endolymph [MV02b]. endoplasmic [DBR^{+}05]. endosseous [MGAD09a, MGAD09b]. Endosymbiosis [FFIS07, dBCP^{+}07]. Endothelial [Bar01, CSON^{+}05, KP06, KM05, LMF08, PSJ04]. endotoxin [DRV^{+}06]. enemies [MPL06a, PBH04]. enemy [Bon03]. Energetic [DHW^{+}09, BG06, NHTM09, RBS05]. energetics [Kal07]. energies [For07a, RA08, ZXWF08]. Energy [AH03a, RRS06, Sco00, SE02b, Smi08a, WBSY06, AH08, Alp05, Bak04, Bal04, Bat06, CQLV^{+}03, CGS05, DBR^{+}05, EH08, Fin06, FS04b, GR05, HD08, JD09, LK08c, MKE^{+}09, NM08, Nis06, Pen03, SdAC^{+}08, SSD09, VHF06, VHF08, WDH^{+}09, WG03, AB04]. energy-dependent [DBR^{+}05]. Enforcement [HB01]. enfuvirtide [MD08]. engagement [SF04]. engine [SRCDS08]. Engineered [DBG01]. Engineering [AH00, Ano02a, Ano04b, LS04b, SJD^{+}09, WBD^{+}09]. Engineers [KV05, GvHP^{+}07]. enhance [BKMH07, HM05, PC09b, SFVA09]. enhanced [LSMZ08, Ped05, WBD^{+}09]. Enhancement [DS00a, GRG02, AB06, BSS^{+}07, DLRL08, RSB09, YHI07]. enhancer [Mur07]. Enhances [AGT^{+}01, FSGB^{+}02, JJ01, LdGH09, ZF06, vdBR03]. enhancing [JAJA07, JBJ^{+}08]. enlarged [VAAH05]. enough [vV06]. Enquist [BDMR06, MGL05, Cle07]. enriched [MGM07]. Enrichment [GEF04, MPN07, MN07b]. Ensemble [AH03a, MKT^{+}00, DRV^{+}08, Kau04]. ensure [Dim05]. Ensuring [BB07a]. enterotoxin [Luc05]. entities [Agi04]. entrained [KG06b, THKU07]. entrant [CBBH01]. Entropic [Nie02, AW06]. Entropy [DSY01a, DSY01b, DPV00, COdAM05, GRBR04, KYZ^{+}08, KmMK04, LKM^{+}08, MB07, RGRB04, VA04a, YDFQ05, ZDC08]. entropy-based [COdAM05, KYZ^{+}08]. entry [LT06]. Environment [Di 00b, PG01a, PL01, TTT09, Wel02, YYI03, ABM03, AD07b, BL06b, Bra07, CGKC07, GASA09, MB06, RKOS09, RW08, SW09]. Environment-dependent [TTT09]. Environmental [CPR00, GN07, Lac01, MGDDH08, PKL01, RTK02, Wel00a, BDBR07, Bat09, BW06, BGO08, BDR08, CCC08, CXM^{+}09, DLRL08, DI09, Gol07, HCEK08, HMP04, LVH04, LRHB09, PTFF05, PBB03, RRKF09, SC03, SW08a, SA07a, TK08, TY09, TBCD06, VdL03]. Environmentally [AN09]. Environments [CC01a, DO00, Gag00, Gre00, KF03, Rec02, SML02, CQLV^{+}03, CGF^{+}08, DCC^{+}08, MBP07, MSH05, YHI03, YHI07]. Enzymatic [SM02, ASMD06, ANL01, RA06]. Enzyme [ADMZ02, LBS00, MKT^{+}00, Alb08, BS05b, CZC05, HGH08, KI04, KI05a, LP07a, LH06, OCA08, RGFP07, SZ04, TE04, TE07, WZW^{+}07, ZCLZ07]. enzyme-catalysed [BS05b]. enzyme-catalyzed [Alb08]. Enzyme-Substrate [ADMZ02]. Enzymes [TGR^{+}00, MGDM08, QLHL09, SB09a, SRG^{+}03, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08]. Enzymes/non [MGDM08]. Enzymes/non-enzymes [MGDM08]. enzymic [MH06a]. epicardial [CPM^{+}09]. Epidemic [CV08, HS02b, MLPJ09, Tho00, Vaz07, BBPSV05, GR08a, ITKL08, JMvdB09, JEDH08, KR07, LTI08, MK06, MM09, PD04, Rel04, RMP08, RH09a, RH09b, RDC09, RGZ09, TK09a, TK09b, WC07a, WG06, XAP07, ZT08]. epidemics [CCM07, FM04, HDZ^{+}07, HRvdD08, ITLN09, MNP06, PL09, PCA^{+}09, SK05, TP05, VdGN^{+}05]. Epidemiological [Jam01c, GLE^{+}09, GWM04, GN07, Wil08a]. Epidemiology [OS06, CB05, SJ03]. epidermal [SMH07, WWS^{+}06]. epigene [TG09a]. epigenetic [DRMLS09, SSF09]. Epilepsy [Jam00, PD06]. epileptic [KRR09, UL06]. epimutation [SP05b]. Epistasis [LK03, TH00, WM00, KYZ^{+}08, MGT^{+}06]. epistatic [HL05, LFFT06, SW08b]. epithelial [Agi04, DT07a, Daw09, PAVS03]. epithelium [DSZ09, LK08a, WTC09]. Epitope [Lan00a, SB05]. epizootiology [DHR^{+}07]. epoxidase [LBS00]. Epstein [CMB^{+}01, SDL^{+}08]. Epstein-Barr [SDL^{+}08]. equality [BP08]. Equation [Gra02, AWW06, BÅ03, CHR06, FM07a, KMT06, Kom04, LMF03, MW07, MNL^{+}07, ON06, Pie09b, PBB03, SBGA04, TW04, vV05]. equation-free [MNL^{+}07]. Equations [CT02, PC02, DB08, HDNC04, Lio09, MH06a, RWK08, TSS06, WW05b]. Equilibria [Gag00, YN02, CMW08, EN02, FP07, RMP08]. Equilibrium [BLMV01, CLZZ02, DSZ09, Hua03, Mie05, MJW00, TML02, Tsc00, ATO^{+}09, HH04a, MP05, TAN09, WN05]. equitans [Di 06a]. Equivalence [LBJE03, Bat09]. era [KM08]. ergodic [RK07]. ERPs [OR04]. Erratum [Ano05h, BZ05a, BBM04, BKE04, FK03a, GWM06, HTN04a, HH05b, KCP09, Ker06, LVL08a, LW04, ML07a, MI08a, NI04a, OI05, RF04a, SI04c, TRM03b, TQUN08, UI04b, Ano00c, Ano00d, Ano01e, Ano01b, Ano01c, Ano01f, Ano01g, Ano01d]. Error [CDF00, DW08b, De 03, FLG^{+}07, GVK00, Sel06, FSS06a, PCK^{+}05, SD04, TGN07, VTL05]. Error-prone [DW08b, PCK^{+}05]. Error-proneness [De 03]. Errors [BS06, De 02a, FGH01, GW06, HH04b, RCA09, TM06]. Erythrocyte [CTS^{+}08, FVP^{+}07, PBZ06, PBZ08, Won06, dPCP^{+}08]. erythrocytes [Won05]. erythroid [CPMG^{+}08]. erythropoiesis [CPMG^{+}08]. ESAD [GV03]. Escape [Fur02, MGP00, IMN04b, WL07b]. escaping [LPC06]. Escherichia [AS00, BKP09, GGK05, Koc00, LL06, MBP07, NGT05, NPN09, PCSL^{+}06, RFH^{+}02, RGF^{+}08, RGW^{+}05, TBB^{+}06]. ESMTB [Ano01h]. Esophageal [NB02]. ESS [BHC06, GV03, HG02, LM01b, SS03, SP06, Yan06]. essay [PTFF05]. essence [MSSS09]. Essential [Fei08, KLL07b, NM09]. established [GT06b]. establishment [AWAB05]. estimable [PBB03]. Estimate [FGH01, TGP^{+}00, Ben04b, Bok06, ELSFB07, VTC08a, VTC08b].

Estimated [HI00a, OKS01, Par04]. Estimates [MC01, AS09, CM09, HW05, HW06, SLH05]. Estimating [BSJ04, CPR00, FB06, GB02, GLK^{+}02, GA02, NB02, PC02, PKL01, RYAI06, SIHH04, BDMP^{+}08, HNTHA07, VTC08a]. Estimation [BKD^{+}06, DNS00, ISWT02, KKK00, KMW03, Lal06, RRR04, SZ01, BGK09, CSS08, Eti09, GBB08, GAK^{+}06, HJ07, KHE06, LSS06, LH06, Nis06, RDF03]. estimators [WTL08]. estradiol [SVG^{+}08]. ET-KNN [DCL09]. Ethylene [WAMO00]. etiology [Rǎd09]. Euclidean [CM09]. Euglena [HP00b]. Eukarya [Di 08]. Eukaryotes [ISWT02, AFZ08, GS06, PLB^{+}05, SM03]. Eukaryotic [BG00b, TPWG01, TN01, Bel06, BOvD08, Nar06b, SG06, SN04]. Euphausia [GMMR07]. European [NAV04]. eusocial [BIS^{+}07]. Evade [CMB^{+}01]. Evaluating [HK07, PKA02]. Evaluation [CRL06, HCB^{+}02, MKA05, SA05b, VDV00, ADHM09, GEK04, JAJH07, TNP07]. evasion [ZKH^{+}05]. Even [GRW03]. evenness [Gos06, Ric03]. event [DGM05, GBZ06]. event-related [GBZ06]. Events [FB01, KKK00, RRK06]. Evidence [Di 03b, Jam00, JLS01, KSK08, Arc07b, BBM03, BBM04, CH05a, DCL09, Kan06, Smi09]. evidence-theoretic [DCL09]. Evidenced [DSY01a, DSY01b]. Evolution [Ano01e, Ano01h, AF01, ACCC02, BES06, BH01, BC04a, De 02a, Dem00a, DML02, Di 00c, DSK02, DHM01, DPA03, EG07, FDG02, For00b, FLG^{+}09, Fra02, Gag00, GA09, HI00b, HS00a, Har01, Hau07, HR00, IST01, JB08, Kal00, Kao06, KW00a, Kel01, KRR09, KY00, KY03, KYS06, KT03, KT01, Lac01, LP07a, Mac01, MZ02, Mce03, MW09, Mic07, MW00, MS01b, MM03b, MM00a, MN06, MN07a, NL03b, NI04b, NI04a, NP05, OMT03, PB08, Pep00, PN00, PD00, PBT02, RB02c, Sch09, SCH05b, SE02b, SPS09, Ste02b, Sug02, SA07b, The00, TU00, VCGV07, WM00, Wak05, Wak07, WD03, YHBW04, Zhi02, AC07a, Agu08, AH03b, AHT^{+}07, Akt04, AJ05, AG06b, Arc07b, AD07b, BHC06, BM08, BPC08, Ben04a, Bon06, BB03, BHWB05, BCH03, BR09]. evolution [BB09, Bul06, Bye05, CSP06, CS08, CL09, CRL06, Chu08, CC09b, Ded09, DLB07, DMQ04, DPA05, Dus06, EK08, Ezo09, Fei08, FR07, FZ07, FABdC04, Fox05, FS07, GK06b, GA08a, GT06a, GINT09, GR08b, GDC^{+}06, HI07, HLA09, Här07a, How09, IMN05, JL06, JI05, Kal07, Kam03, KAI08, KB04, KSB07, KGL09, KWGE04, KS04, KSG03, KOK06, KD06, LJ06, LJ09a, LGK^{+}09, LGK^{+}12, LKK07, LVH04, MR03, MHMGM08, Men07, MVS^{+}06, MWB05, MK09, O’K05, OAC03, Osb08, Ots08, PB03, PBL06, PGH^{+}04, PBH04, QB06, RV05, RGFP07, Rot09, Row06, SP07a, SGS^{+}05, SG07, Sch05a, SSLB07, SPF08, SWI07, SN09, SH08a, SC09, SF08b, Smi04, SA08a, SPB06, ST05, SA08b, TFYY03, TCP05, TMBD09, Wag03, WOLS07, WSM05, WRKK09, Whi07, Wil09, Yam03]. evolution [Yun05, ZJ05, ZJ06, ZF06, ZWT^{+}08]. Evolutionarily [Ant02, BSR06, LBSS02, RHM^{+}08, YLL00, AM04a, AD07a, Kom07, LW04, MH05, YYY^{+}08]. Evolutionary [AH00, Apa09, BK01, CDHJ02, CGH01, CG03, CH07, Dem02, FP07, FB00, Fra08, Had01, IMN04b, Jäg08, JZST09, KMC^{+}07, KNN01, LL08, LB07, MDD06, Nei04, NJVA04, OPN07, ON08, OKS01, PN01, PN02a, PDA^{+}00, QA01, SSWF01, SK09b, VO00, WZL^{+}08, WL07b, ZSF^{+}07, ZCD05, AJ05, ACK08, BC09a, BN04, BP08, BR08, BL06b, Cor05, Di 06b, DL08, Fis06, FMH08, GA07, Has06, HA09, HBWB08, II06, KV05, KF06, Ken07, LJTD05, LM08, MSP03, Mas08, Mas09, MWCS04, Mie05, Net09, NI04c, OI07, OBN07, PTN06, PHP03, RSSM06, SW08b, Sim08, TSN05, TIN07, TPN07, TLC07, Wax06, WT07, WN05, Win06, dSKL09, vdGMG^{+}09, MS09b]. evolvability [ABKR07, Gol08]. Evolve [PD00, Dic08, FZ04, Gab06, Gol09, NS03a, Uit09]. Evolving [Hog00, MS09a, PSSS03, Wah02, GD08]. evoultionary [TIN06]. EX1 [Ano03-42, Ano03-41, Ano03-43, Ano04-54, Ano04-50, Ano04-56, Ano04-49, Ano04-51, Ano04-52, Ano04-53, Ano04-55, Ano05-55, Ano05-56]. EX2 [Ano03-42, Ano03-41, Ano03-43, Ano04-54, Ano04-50, Ano04-49, Ano04-51, Ano04-52, Ano04-53, Ano04-55, Ano05-55, Ano05-56]. EX584 [Ano04-56]. Exact [BÅ03]. Examination [CC01c, DRW01, ML09b, MM03a, Sta08]. examine [CKE06]. examining [BMH07, HW09]. Example [CBH02, IGHW07, Kal00, FMLP06, FLG^{+}09, LRHB09, YM04]. examples [CWJ07]. exception [BM09]. excess [TE07]. Exchange [McN06, Mcn00, MN01, SBH01, WGH01, Bat06, BT06, BTS08, PM03]. exchanging [San03, SBH01]. excision [DdB03, KKR^{+}07]. Excitable [ACK00, ERM00, BGP00, CD07, GSM06, RSG09]. Excitation [MV02b]. Excitatory [TTKZ01]. exciton [Sin06]. Excluded [AW06]. Exclusion [BL09, RBBH02, FT09a]. Exercise [Sco00, GGH^{+}05]. exerted [ML09b]. exhaustive [OI07]. exhibit [Kin07]. Exhibiting [TH00, DF08, RA06]. eximius [SMD00]. exist [BB05, Dru03]. Existence [IMNT09, LGCL07]. exists [BEBV03]. Exit [öOW01, TQUN07, TQUN08]. Exocytosis [FY00]. exonuclease [Xie09]. exotic [GT06b]. expanded [NGTB06]. expanding [HH06c]. expansibility [Est07]. Expansion [FB01, JE09, NMI06]. expansions [MPL06a, VA06]. Expectation [MM00b]. Expected [Dic08, FB00, RHM^{+}08]. Expenditure [Sco00]. experience [PTFF05]. Experiment [SE02a, LSS06]. Experimental [CSFH^{+}01, CMTU01, KS01c, LH06, MS01b, Nas01b, PH03, Wan00, WCLL08, BGE06a, CTS^{+}08, CLS08, FS04b, JR06, LMT05, MCC^{+}09, SC03, WRKK09, KSM02]. Experimentally [FHD09, MW01]. Experiments [ABR02, Kru02, MM00a, AKLS05, DMPS05, ISW04, IvDH^{+}08, SVS04, Tan07b]. Experts [RMAI06a, RMAI09]. Expiration [LWFP08]. Explain [BSRH02, BSRH03, MFB01, BTL08, Buc04, CS00, EG07, Osb08, PE04, RMRG09, SGTF07, Sch05a, SCNP^{+}06, Wak04, vV06, vdBvdB09]. Explained [öOW01]. Explaining [Dus02, SSF09, Twa04]. Explains [Dus00, KK00a, CM09, Dus06, MGP^{+}08, RBS05]. Explanation [GK00, FHD09, GMFS06]. explanations [Jam09b]. Explicit [Sch02b, CG06, CSM05, HH04a, KB04, LKM06, LSMB^{+}03, RHH08]. Explicitly [Smi08d]. Exploit [Möl02]. exploitation [De 02b, DBGM08, HH06a]. exploited [SBI07]. Exploiters [LBF01, Gol07]. Exploration [AOH03, FF02, CB08a, MCC^{+}09]. Exploratory [SMD00]. explore [Pal08]. Exploring [AS07, BFH^{+}01, BOvD08, CPG09, EKS02, ELJ06, HML09, MDD06, PL09, SWB06]. Explosions [DMW08]. exponent [GDC^{+}06, HJ07]. Exponential [SS01, CL05, MAL03, Niw03]. exponentially [HI07, LL05]. exposed [ML08a, ML08b, wTA09]. Exposure [ES08, BGMM08, GRH^{+}07, LMF03, Lit07, Luc05]. expressed [YDFQ05]. Expression [CSP01, CP03, FC01, HR00, Ish00, LSD^{+}00, AO03a, BZ05a, BZ05b, DGM05, DKS04, DSA^{+}06, FM07b, Gie06, HAC^{+}09, IS08, IZGG05, KLL08, Lei09, Lei10, LPMC^{+}06, mLLS^{+}06, Moc08, MA04, MKA05, MLS09, NS09a, NTU06, PLB^{+}05, PHG04, PE04, SVS04, SVGK07, SP06, SHI03, SVN^{+}05, TZS07, TSZ^{+}07, Wil06b]. Extended [RSBY03, TS02, BGF07, CXZ^{+}09, DFCL08, KS08b, LHDvdM04, MS07a, RPNH03, SS06a]. extending [BCRG04]. extension [BLS^{+}09, Hol06, IS03, KV05, MCN08, TSRB08, ZJG03]. extensions [LD09a]. extensive [DFP^{+}08]. extent [Gie06, RG05]. extermination [YHI07]. External [DF00, GRG02, XT06, Ben04a, GVB^{+}08, LN05b, LHDvdM04, MS03a, MJ07a, OB04, Tao04a]. Extinction [BB02, HIN00, HI00a, HI05, JE01, JSV02, Nås01a, OSBH02, PCZL05, PCZL06, Bok06, ES00, FPS08, GMM09, GT06b, KJJ07, KBD06, Par04]. extinctions [AB04]. Extracellular [HSMM07, LGB02, VP01, FDA^{+}09, FS04b, FH07, GHA03, Kur08a, TN04, TTN05]. Extracting [AHT^{+}07, Tan07b, Wan06]. extracts [ZAB^{+}09]. extraordinary [Wak05]. Extrapair [FSL03]. extravascular [FH07]. Extreme [PPEøP02, GRW03, KV05, WMP03]. extremes [TSK09]. extrinsic [GS08b]. eye [EH08, GMMR07]. eyes [CS08, How09].

F [NMH07, NM06, NM06]. F-actin [NMH07]. F1 [GG07]. Fabaceae [PCB07]. face [Unn09]. faces [KSB08]. facilitates [PM07, SA07c]. Facilitation [MS08a, BFGD07, CXM^{+}09]. facing [CS08, How09]. factitious [RKF06]. Factor [BPV01, FFN00, KMP03, Wil01, BM06, CZM09, PE04, PSJ04, SMPM09, VGS^{+}05, WOLS07, WWS^{+}06, YYA09]. factorial [DMW04]. Factors [CD00, FMP01, HIN00, Jam01c, PB07, VW00, Yam03, BL08, CHdV06, FS04a, RG06]. Facultative [NSM02, YK03, Ezo09]. FADH [MS01a]. failed [MG09]. Failure [ZCC01, BLZ07a, BLZ07b, Smi08d, TSC04]. Fairness [Här07a]. FAK [CLB05]. FAK/Src [CLB05]. Falciparum [GLK^{+}02, DEM^{+}00b, FVP^{+}07, MB05]. false [HH04b]. Families [Jam00, RMAI06a, RMAI09]. Family [TL02, Bog04, BKKR08, CZC05, KTE08, LZ09b, PCB07, SLH05, YMLK04]. Family-structured [TL02]. far [ES00]. farm [XFBC07]. Fas [OAKC08]. fascicle [LW08a]. Fast [Cin06, PAD00, KH09c, MSdlPS09]. Fast-tracking [Cin06]. fastest [TIN07]. fastidiosa [VRAF06]. fasting [BGG^{+}09]. fat [Alp05, CTB09]. fate [AMP06, BFK08, FFHK09, LCL07a, PWZ09]. Fatigue [Mar03, Tay00, ZCC01]. fatty [BKD^{+}06, BGG^{+}09, Hul05]. favor [KOK06, NBT07, Ped07]. Favour [Di 00c]. favoured [SPF08]. fd [AHT^{+}07]. fearfulness [JZST09]. Feasibility [BGP00, HH08]. feasible [CK08]. feathers [DFW^{+}07]. feature [LZ09a]. Features [BS00, APS08, CCZC08, LZ09b, PWZ09, UKY^{+}09]. February [Ano00p, Ano00r, Ano01u, Ano01-28, Ano02s, Ano02y, Ano03-34, Ano03-38, Ano04-54, Ano04-42, Ano05-47, Ano05-53, Ano06-48, Ano06-52, Ano07-43, Ano07-49, Ano08-44, Ano08-48, Ano09-38, Ano09-44]. fecal [IGHW07]. fecundity [AC07b, SS03]. Feedback [CA09, CSS08, CD02, ESGK02, KCA03, TMS00, UHK01, Zhi02, AN09, FM06, GC09, KL09, LWC09, LJL08, RGF^{+}08, RK06, RCS05, SZ04, SA05a, SHI03, SSF09, TZS07, WCA06, XK07, XT06]. feedbacks [BDBR07]. Feedforward [KH09c, HJ05, SHI03]. Feeding [COS01, LPJB^{+}08, RPNH03, SKY09]. feet [BR09]. Feline [FPL03]. Female [BFA08, Dus00, Cor05, FHL^{+}06, PDM04, SWI07, Wak05]. female-biased [Wak05]. Females [NS03b, BP03, CMCH08, GOP09]. femur [SB07]. fermentation [ZC09]. Fertility [FSL03, GRW03, MI08a, MI08b]. Fertilization [Nir02, FWLN04, LHDvdM04, MA03a, SK09a]. Fertilizations [FSL03]. Fetal [De 02b]. fetus [Buc04]. FGF [FWCN05, GP08a, GFWT04]. FGF-2 [FWCN05, GFWT04]. Fiber [DSCD02, SHF02, HdGH07, KH09a, PC09b]. Fiber-Forming [SHF02]. fibers [CSON^{+}05, Tas05]. fibre [DWBB04, KH07a, VTC08a, VTC08b, WBD^{+}09]. fibre-enhanced [WBD^{+}09]. Fibril [FW00, GMAH07, ZZW07]. Fibrillar [SHF02]. fibrillation [JB04, RK05]. fibrils [GHA03, GMAH07, QXDW06]. fibroblast [KH09b]. fibroblast-controlled [KH09b]. fibroblasts [AGW^{+}06, AGW^{+}08]. fidelity [SA08a, Xie09]. Field [CSM02, CRR08, GKK06, LS03, LS04a, LB06, OHM05, SP07b, Wal08, WKB07, WKRD09, vAR09, vAGDR09]. Fields [HFGB02, KFG^{+}02, LFLM00, Lin01, Ben03, ES08, JA05, Lew03, NOT04, Pie09a, YLCW06]. fighting [MH05]. FIH [DMPP09]. filament [BEF^{+}06, BLF^{+}09, DECEK06, GP08b, LCL03]. filiform [CGKC07]. filling [BBSLN06, BBSLN08, HTN04a, HTN04b]. filter [SSK06, vdBR03]. filtering [HQP^{+}09]. filters [MWCS04]. fim [CB07a]. final [TK09a, TK09b]. Find [Kru02]. Finding [Hot03, PLH05, SSM09, Ten08, TKN07]. fine [CLS08]. Finger [Jam01a]. Finger-length [Jam01a]. Fingerprint [KN05, LZ09a]. Finite [JMB00, KN03, Paw07a, SNT03, BBSLN06, BP08, DGS09, FP07, FAW06, Fuk05, HA09, LL08, LT09, LME06, WT07]. Finite-Element [JMB00]. Finnish [Hel08]. Fir [AF02b]. Fire [FLBB01, LL05, UL06, VL09]. Firing [PB02, WKRD09, FL03, LL05, WCLL08, vAR09]. First [BGE05a, For00a, FY00, Gra02, PBB03, Tho00, TGR^{+}00, WL02, BGE05b, Maz08, Tak06, vdGMG^{+}09]. First-Order [TGR^{+}00, FY00, PBB03]. First-World [Tho00]. fischeri [KH08]. Fish [Arm01, IK02, Kov02, CKE06, FM07a, GBK^{+}07a, GT06b, HMP04, ILDP04, LP07c, Niw04, Niw05, QASL08, QSAL08, RBP^{+}09, Ver04, YM04, ZKH^{+}05]. Fisher [LH06, Orr06, SLH05, Wax06, Wax09]. Fisher-curves [SLH05]. Fisherian [Win06]. Fishes [COS01, Wes03]. fishing [BR04]. fission [BS05a, ZSZ^{+}06]. fit [SH08b]. Fitness [CSFH^{+}01, GRG02, RGG00, SL00, AH03b, AHT^{+}07, AH08, Ait08, DGZ07, FLG^{+}07, FP07, FZ07, FLS^{+}04, GCP04, GINT09, Gra06, GPG07, OBPH^{+}08, Orr06, TDW07, WR07a, WN05, WC07b, vV09]. fitness-associated [WR07a]. fitted [LVL08a, LVL08b]. Fitting [AH08, BEF^{+}06, GFW^{+}09]. FitzHugh [ACK00]. five [GR05]. five-link [GR05]. Fixation [HA09, Wax09, KK06, MB09, PW09a, PM07, TIN06, TDW07]. fixed [BLS^{+}09, Fuk04, SH04]. fixing [PBT02]. flagella [Dru03]. Flagellar [Ats01, Kee05]. Flapping [TT02, TT03]. flare [ITM^{+}07]. flare-up [ITM^{+}07]. flattest [SES08]. Flavonoids [ABP^{+}00]. Flexibilities [SHF02]. Flexibility [IK02, TL00, LCBTP03, RSSM06]. flexible [CM03, KLL07b, MS03a, MGT^{+}06]. fliers [RH04]. Flies [FDG02]. Flight [Ale09, SP01, TT02, TT03, TT01, BEBV03, CF07, LZGL03, Lor06, OST09, PBR01, RRS^{+}07]. Flights [PE00]. flock [LPC06]. flocks [BZM05]. Floral [MT06a, MT06b]. Flow [AF02a, AF02b, Bar01, QB00, SNCM09, SNT03, AB04, BGE06a, BGE06b, DS05, ES08, FFTS09, FM06, HC07b, HH06c, KNS05, LMF08, Mei05, PDB08, RGB06, SLIL07, SZLK^{+}05, Smi08a, SS05b, WBR08]. flow-induced [DS05]. flow-tissue [SS05b]. flower [EW09]. Flowering [GBCC01, GHC03]. flowers [MT06a, MT06b]. Flows [FGH01, SKY09]. flu [ITL09]. Fluctuating [PKL01, HF03a, MB06, YHI03, YHI04, YHI07]. Fluctuation [BL06b, YMC01, KF06, Kut05, Lei09, Lei10, MI08a, MI08b]. Fluctuation-induced [YMC01]. Fluctuations [PE04, Lie05, MAL03, PTFF05, SC03, dSKL09]. Fluid [COS01, DF00, SNT03, CGKC07, GCB^{+}07, LCL03, Lap03, SKY09, WTC09]. Fluorescence [Laz03, GZL^{+}03, LIK^{+}05, LJ09b]. fluorescens [FFME08]. Flux [BKF00, CSP01, SS09b, Agu08, Che06b, FMI05, Gie03, HTCS07, JD09, LP07b, MSSS09, SPAH06, WMP03, WFGP04]. flux-summation [Agu08]. Fluxes [DLGC02, Lan00b, OCA08]. fly [ST03]. flying [Bye09, Iim07, MA03b]. focal [SMCT08]. focus [DHRM08, KS08b]. folate [AIKP06]. fold [SC09]. Folding [Dem02, GAA02, LRD04, NJVA04, AH08, BKKR08, CXZ^{+}09, For07a, HQP08, MLWL06, NSH^{+}03, NSG08, PBMU^{+}09, RGPB08, ZXWF08]. folding/unfolding [HQP08]. Folds [DML02, BM03a, Shi06a]. Foliar [HG08, JJEX09]. Follicular [HBLG02]. follow [GDD^{+}03]. followed [FHD09]. Following [UI02, BFGB04, BFG05, Dus01, LGB03a, Lit07, MPL06a, NKC^{+}08]. Food [Bar05, CC01b, DHM01, Est07, Går00, GDS07, Har02, II06, JSV02, KJJ07, LBS06, LBJE03, MS02, RMAI06a, RMAI09, TTR00, VDV00, AP04, BDBR07, Bal04, BRCB04, BLMV05b, CSA07, DI09, DMQ04, FV09, GD08, KGD09, KJD05, KRGH07, Kon06, LM08, MB07, Paw09a, PM08, PB09, RMAI06b, UD07, Van06b]. Food-web [GDS07, II06, BRCB04, Kon06, RMAI06b]. foodchain [GEF04]. Foot [TK09a, TK09b, KH07b]. Foot-and-mouth [TK09a, TK09b]. footprinting [YMLK04]. forager [AO03b, HC07b]. Foragers [Bea00, RHG00, DBGM08, HMGK06, PBR01]. Foraging [PE00, SCH02c, YHG^{+}02, CM03, JR06, KGD09, KHHS09, Kon06, KV04, LNRR06, MSH05, PB06, RRH08, SSR04, VTG^{+}06]. Force [Nie02, Pro03, BB06b, DO04, HQP^{+}09, KSG03, LN05b, RSB09, Wes03]. Forced [KVMV01, CD07, NAS07]. forces [CABB09, KS09, ML09b, NSG08, SP07a, SS06c]. forcing [BGO08, GN07, MS03a]. Forecasting [XBGN05]. foreign [Buc04]. Foreignness [vdBR04b]. Forest [SI00, SI04a, BK05, SIHH04, SI06, XBT09]. forested [SLIL07]. Forests [PPL^{+}00, BLL08, PAA05]. forever [SS06d]. form [BGD^{+}06, HI04, KWGE04, Sac04a, Usa06, Whi05]. Formal [Gra02, SSWF01, BCRG04]. Formation [BS05c, Bro02, Car02b, Cum00, DL00, EKS02, FW00, JG01, KW00c, Kra02, Moc02, SCGFS00, SWS02, WB02, ACL03, ACLW03, ARR08, AGZ^{+}06, BLF^{+}09, Ber03, BI09, BC02, BSJ04, CKM05, Cog06, Cog07, DVC^{+}04, FI06, FM06, GBK^{+}07a, HKP07, IHV^{+}06, II06, Joh08, KLN^{+}09, KN05, LDW04, LF08, MNI06, Mig06, NS09b, NOT04, PAA05, RN07, RM04, RRR^{+}08, SJGK04, She06, SP07b, Uit09, UI04a, UI04b, UMI09, vVH07]. formations [SPH03]. formats [RNS04]. Formed [LSMLB^{+}02, SIMK02, FZ04]. Forming [SI00, SHF02]. Forms [DML02, BM03a, DSS08, GEF04, SP05b]. formula [AH03b]. Formulae [AH03a]. formulas [HC07a]. Formulation [CMS^{+}00, Kor01, AWW06, LJ09b, NTU06, OS03]. Forsdyke [KRN01]. forward [BB06b, CS08, How09]. forward-facing [CS08, How09]. fossil [Di 06a]. fossils [Par04]. Found [SHF02, BKKR08]. foundations [QASL08]. Founder [PNG03]. Four [FLBB01, DSU^{+}04]. four-dimensional [DSU^{+}04]. Fourier [DVL^{+}00, HQP^{+}09]. Fractal [Gar04, QB00, Zam01, Ben04b, PPRI08, UCSZ07, ZYD^{+}05, CML08]. Fractal-multifractal [CML08]. Fractals [PPR01]. fraction [Han04]. fractional [HCM^{+}07]. fractionated [SSB^{+}07]. fractionation [DSU^{+}04]. Fractions [KKDV01, BSJ04]. Fracture [BPV01, GGV^{+}08, GBGAKD05, IvDH^{+}08, THL03]. fragility [CAS05]. fragment [ZZW07]. fragmentation [CSM05, FPBM08, Kon03]. fragments [SV03]. frame [PV08]. frameshifts [Sel07]. Framework [BPV01, PN01, CB08a, CSD04, Cui07, Hor08, KI04, KI05a, LFC04, MH09b, MGT^{+}06, MSL04, PP04, Rej07, SWN07, SN06, SML07, SML08]. Frameworks [LK03]. France [SGS^{+}05]. FRAP [TR08]. fraternal [Bog04]. fratricide [AB03]. Free [Bak04, Bea00, CW08a, RGG00, AH08, CL05, DO00, Fer09b, FR06, FM07b, IiKY07, KB04, KBD06, KHHS09, MS08a, MA03a, MNL^{+}07, PBMU^{+}09]. Free-living [CW08a, KB04]. Free-running [RGG00]. free-spawning [MA03a]. freely [Iim07]. Frequencies [HGB^{+}00, Cib08]. Frequency [AJOK09, CGH01, Ken02, LAD03, SVS^{+}02, SCS04a, TL02, YK03, YLL00, YL00, BR01, BFR05, Fli05, GBZ06, Izs05, KGL09, O’K05, PB08, RHF07, She06, Sta08, SFC^{+}09, WVA05b]. Frequency-Dependent [YK03, YLL00, CGH01, TL02, YL00, BFR05, KGL09, O’K05, PB08, She06, WVA05b]. Friendship [HH06b]. Frog [AN09, ZAB^{+}09]. frontness [Nar06b]. Fronts [Sch02b]. frq [BMT04]. FTTP [YDL06a]. fuel [Lin04, dBWBP06]. Fuji [AH03b]. Fuji-type [AH03b]. Fulfilment [DS00b]. Fully [DALP03, JEDH08]. Function [CACC02, SLP00, BBB^{+}07, DB05, DHP06, FCD^{+}05, GZG07, Här07a, HMN09, HPZ09, LWC09, LJL08, Men07, NS09a, NAV04, NAS07, PCZL05, PCZL06, PWG09, RGFP07, Sar04, WH07, ZAdlPLM07]. function-valued [DHP06, Men07]. Functional [BFP07, BJD^{+}02, DSY01a, DSY01b, FKAC06, Kal00, KA02, Kun00, Laz02, LDW09, Noe00, Phi02, WRF01, AIKP06, Bon04a, Bon04b, CC06a, DMQ04, Kam03, KSS07, KV04, LLS^{+}09, LLW09, SC09]. functionalities [DHYHR09]. Functionality [NA02]. Functionally [RKH^{+}06, DLB07]. Functioning [LLF02, SWN07]. Functions [FPC01, Rae02, BRC07, Ben04a, CZM09, GMY09, GZ04, HWCF07, PBB03, RGB06, vLBJK07]. Fundamental [YL00, PHP03]. Fungal [GBG01, TKM06]. Fungi [BJD^{+}02, DO00, GG02]. furrow [AGZ^{+}06]. Further [Di 03b, Kan05, Kan06, Kan07]. Fusco [PKA02]. fusion [BI09, CCZC08, MD08, ZSZ^{+}06]. fusion/fission [ZSZ^{+}06]. Futile [JBP02, BH03b]. Future [Ste00, KY03]. Fuzzy [SYC06, BZM05, FLWB07, GKNT09, LM09]. fuzzy-logic [FLWB07].

G [CHN08, FBM06, FM09, KRN01, ØKRG04, BTL08, BCL08, BNT^{+}00, KOK06, OF01, WKG03, WKL01, WL04]. G-CSF [FBM06, FM09, ØKRG04]. G-Protein [WKL01, BTL08, BCL08, WL04]. G-protein-coupled [OF01]. G1 [Coo01, CLW^{+}08]. G1/S [CLW^{+}08]. Gaia [Sta02, Sug02]. gain [CSS08]. gains [SHI05, SHI06b]. gait [BSGT08, GSB05, KSPA^{+}08]. gaits [BD08b]. galactose [SSK06]. galloping [RBS05]. galvanotaxis [OOHI06]. Game [Bea00, BK01, Di 00c, FZ07, MSWH00, PN01, YL00, Arc09a, BC02, BLRR08, CCC08, DT06, Ezo09, Här07a, HB09, Has06, HA09, Hur06, KAI08, KNS05, LL08, LJTD05, MDD06, Mar09a, MT06a, MT06b, MH06b, MGS09, NS03a, OT09, Rǎd08, RKF06, RKOS09, Sch05a, SA07c, TPN07, Uit09, WZL^{+}08, WVA05b, YPY^{+}09, YAL04, dSKL09]. Game-theoretic [MSWH00, BLRR08, HB09, MGS09, RKF06]. Games [BP00, BS01, Cro00, HDHS02, HNO03, Roy00, Ste00, WB02, ATO^{+}09, ANT09, Apa09, BHC06, BP03, BC04a, CS06, FP07, Has06, HH08, Jäg08, JG06, Mas08, Mie05, PTN06, PTON08, SN09, SPS09, SA08b, SSS08, TAN09, Wak07, WN05, WVA05a]. Gamete [Dus00, Dus06]. Gametocytes [DEM^{+}00b]. Gamma [FCD^{+}05, RWR00, GBZ06]. Ganglia [KFG^{+}02, vAR09, vAGDR09]. ganglia-thalamocortical [vAR09, vAGDR09]. Ganglion [BKCR01, MRJR09]. Gap [Ano01b, GS00, HP00a, MY01, PAD00, ST01, AD06a, DBB09, GBGAKD05, PAA05, PAA07, SI04a]. Gap-gene [Ano01b, ST01]. gapped [LBCL09]. Gas [SBH01, WGH01, BT06, BTS08, KRN03, SWRH03]. Gaseous [CB00]. Gases [Wal00]. Gastric [KKDV01, JK04]. Gastrin [RB02c]. gastrocnemius [LW08a]. GATA [RG06, YCS04]. GATA-1 [RG06]. GATA-3 [YCS04]. gated [Gre05]. Gating [KSM02, WT01, HS07, MTS05].

Gaussian [FLG^{+}07]. Gaussian-distributed [FLG^{+}07]. GB [NTK02]. GC [MX08]. Gel [Tas02]. gels [Tas05]. gelsolin [MO07]. geminiviral [PCB07]. GENE [BZ05a, AIK00, BKMH07, CdOWS04, CS02, CSP01, CP03, DMPP09, Eld00, FC01, FB01, Gin03, HCC00, HR00, Ish00, Kun01, Kun03, LSD^{+}00, MS01b, SCGFS00, TG09a, Aba09, AKLS05, ACK08, Ano01b, BM08, BZ05b, BLS06, Bra05, BOvD08, CZC05, CSW^{+}08, DGM05, EPR07, HAC^{+}09, HE08b, HC07b, IYGA08, IS08, IZGG05, KMC^{+}07, KK06, KZ05, KLL08, Kun05, Kun06, LS04b, Lei09, Lei10, LPMC^{+}06, MSP03, Mar09b, MI06, Moc05, Moc08, MA04, MKA05, NS09a, NTU06, Nar06a, NP07, OW07, OS06, PSSS03, PHG04, PE04, PHdB09, RKYH06, RRK06, RK07, RAHO06, RCS05, ST01, SBPC05, SVS04, SVGK07, SR08, SP06, SBZ^{+}08, SHI03, SM03, SCS04a, SVN^{+}05, Tan08, Tao04b, TZS07, WLZ^{+}06, WSS07, Wil06b, WW07, XSD^{+}05, ZAD07, vLHH06, BZ05b]. Gene-based [BKMH07]. Gene-culture [Gin03]. gene-duplication [BLS06]. gene-for-gene [OS06]. gene-modified [vLHH06]. gene-orientation [SM03]. gene-protein [ZAD07]. genealogical [MCF04]. Genealogy [DMZ00, FB01, RSBY03]. Genera [RH02]. General [LRHB09, OSCK01, Pie09b, SKW00, TGR^{+}00, Tsc00, VGCGM^{+}02, WC02, CB08a, GBK^{+}07a, HH04a, LFC04, LSH06, LHDvdM04, MLWL06, PCK^{+}05, SWN07, VGMVR^{+}06, MWS09]. generalise [Ded09]. generalised [RH08]. generalist [vdBvdB09]. Generalists [Wah02]. Generalization [Ghi02, WYC06]. generalizations [Kan05]. Generalized [ACCC00, DLGC02, DFC^{+}02, GDPMDS^{+}09, HG02, HW09, PN01, PC02, SV05, CY09, DTG09, DW05, HGC03, Kan06, Kan07, Kri09, VTC08a, VTC08b]. Generalizing [RHH08]. generate [PFRR08]. Generated [HFGB02, AP04, DO04, FK03c, LFFT06, SI05]. generates [SH03]. Generating [SKY09, CFM04, ZKH^{+}05]. Generation [Ber07, Nie02, UI02, Wil09, HFS06, INSR08, JDMZ^{+}07, Koh07, KA03b, LFC04, NSI08, SSJ09, SCH02c]. generic [Ban06, JJEX09]. Generous [RON09, Sch09]. Genes [De 02b, LN02, LY02, AO03a, AD06a, AS09, Ded08, DZB^{+}04, Di 06b, Di 08, Fol08, GMM09, IMKN05, Ken07, KSN03, MX08, Mit04, PWZ09, PER03, SP06, TYI^{+}06, YDFQ05]. Genetic [AGCLMM03, AAK^{+}09, BMS06, BHJ03, CDF00, Cha01b, Che06a, Ded09, Di 00a, Di 01a, Di 01b, DM00, GS02c, JP00, KS01b, KCA03, Leh00, MWCS04, MAB00, NMI06, QA01, SVS04, Ste02b, WM00, WMLC02, ABKR07, AD07b, BW06, CAS05, Che03a, CL09, Dav09, Ded08, Di 05, Di 09, DSU^{+}04, Fol08, GNH^{+}05, HK09, HL06, JS07, KF06, KYZ^{+}08, Kau04, KOT07, KS04, LRD04, LKK07, LMT05, LFFT06, MWS09, MM08b, MGT^{+}06, NAP04, NP08, OM08, OB04, PLH05, Pat05, Poh08, QB06, Rak04, RHF07, RB08, SVGK07, SH08a, Tao04a, TB04, Tlu07, TKKA04, TC09, VGDSU09, WCA06, Wil09, XT06, ZF06, ZRSK07]. genetic-component [VGDSU09]. genetically [FLG^{+}09, MGL^{+}06b, RBW09]. Genetics [FPS01, MK01, Möh00, OSCK01, PV00, PBT02, TH00, Cam03, IMKN05, MW07, UH09, WW05a, Win06]. Genetics-the [Möh00]. Geneva [CAHH06]. Genome [Ano04b, Che03b, GKM^{+}00, LN02, PPEøP02, SP00, ZSZ^{+}06, Cui07, FP03, FSS06b, Fuk04, Fuk05, GS06, Gar04, GGK05, HK05, QHF^{+}07, SWB06, WLZ^{+}06, ZYD^{+}05]. Genome-Scale [PPEøP02, SP00, FP03]. Genome-wide [ZSZ^{+}06, Cui07, GGK05, QHF^{+}07]. genomes [Ben04a, FM03, MX08, MBB08, OAC03, RSH^{+}06, TSS06, Wil08b]. Genomic [GK00, Kro08, WH03, LVL08a, LVL08b, NA03, SRS09, Voi03, Wil06a]. Genotype [MDD06, Pec06]. Genotype-Phenotype [MDD06, Pec06]. genotyping [LS04a]. Geobacteraceae [YMLK04]. Geographic [LNRR06]. geographical [ITLN09]. Geometric [Gin00a, BHWB05, KSG03, Orr06, SS09b, XWC08, ZLXY08]. geometric-shape [KSG03]. Geometrical [LH01, Nos06, RMBM00, TN04, Wax06]. geometries [Tak06]. Geometry [Cum01, HL00, JBK04, LF01, CM09, KC05, PAA05, SML07, SML08, Wax06]. Georgia [Ano01-33]. germ [CWJ07, GZ04]. Germinal [IM02, KS01c, MHDOG01, MH02, KD03, SBZ^{+}08]. germinating [OF05]. germination [NAV04]. Germline [Zhi02]. gestation [Jam08b]. getting [Cyt04]. giant [LRT04, Rap08]. Giardia [Fer09b]. giraffe [MBD08]. Girdling [FGH01]. given [AC07a, BE08, MH06b]. gives [FSG00]. Gland [KKDV01, SVS^{+}02]. Glandular [KKDV01]. glass [Won05]. Gliding [TT01, KKPB09, OST09]. glioma [KLN^{+}09]. Global [CC01b, FMP01, GBB08, LYZ08, OI07, JB04, LLS^{+}09, MHRK08, PAA07, Wal07a, WMS08b]. Globeflowers [FDG02]. Globular [AH03a, ZSZ01, BC04b, JNWJWB04, ROR05]. glomerulus [Saf05]. Glucagon [RB02c]. Gluconeogenesis [JBP02]. Glucose [AMV^{+}02, RBBH02, TMS00, TPD^{+}00, JCK09, LKM06, LT08]. Glucose-induced [AMV^{+}02]. glucose-insulin [LKM06]. glutamate [BFG09, Saf05]. glutamatergic [BFG08a]. glutathione [MCMS06]. Glycogen [GBD00, FKAC06]. Glycolysis [JBP02, KSB08, NO04, NM08, SRR08, dPCP^{+}08]. glycolytic [GA08a, KR05, RvMK^{+}05, WB04]. glycoproteins [AS07]. Glycosides [Kal00]. GNC [LRD04]. GnRH [KPS02]. Goal [FPC01]. goals [RSSM06]. Gompertz [Gol09, Lo07]. Gompertz-Makeham [Gol09]. Gonadal [Jam01c]. gondii [KCP07, KCP09]. Good [HDHS02, BS06, HH04c, HH05b, JG06, TK09a, TK09b]. goodness [OI04, OI05]. goods [KAI08, Wak07]. Goodwin [RVMR01]. gorgonian [SZC^{+}03]. Gorshkov [BDMR06]. govern [Gra07]. governed [BR04]. gp120 [Lan00a]. gracilis [HP00b]. grade [KJJB07, LAG07]. graded [DMPP09, PE04]. Gradient [BGP00, Ben03, HKP07, KI04, KI05a, MGS08, MM03a, NM09, NSI08, Nar06b, SMNL07, SN04, VN07]. gradient-sensing [SMNL07]. Gradients [Dus01, Har01, Wal00, CXM^{+}09, Gie03, Gie06, Sav04, TBCD06, ZLN07]. Graft [RMBM00]. Grain [LY00]. gram [Kun03]. gram-positive [Kun03]. Grammar [KNN01, JPJ06]. Grant [Jam08a]. granule [Saf09]. granulocyte [ØKRG04, VGS^{+}05]. granuloma [SJGK04]. granulopoiesis [VGS^{+}05]. Graph [RSD^{+}01, FABdC04, GBD06, MMUGD09, OPN07, PBMU^{+}09, VN08]. graph-based [VN08]. graphical [DqLmW07, LDXW06, QWQ07, YSW09]. graphs [GDPMDS^{+}09, KST07, LTW06, ON06, ON07, ON08]. gravitropic [AF09]. Gravity [SBvS06, FPM^{+}06]. grazing [CQLV^{+}03, HFY07, YFH^{+}07]. great [CAHH06]. Green [Möl02, HHH06, LY03b, NAS07, PWG09]. gregariousness [vKPdR07]. Grey [HDRM00, DBG06]. grid [SI06]. grid-based [SI06]. grip [SBvS06]. grisea [TGT06]. Grober [EK08]. groove [AAL08]. grounds [KRB05a]. Group [BR02, DPA03, OMT03, PT09, Pep00, VMW01, WB02, vV09, Arc09b, BR09, BFA08, Chu08, DPA05, DBF07, Fer09a, FSS06a, JG06, MNL^{+}07, NDD08, Niw03, PBR01, Sch08a, Sch09, ST05, Uit09, YYY^{+}08, vV06, vVH07]. Group-beneficial [DPA03, DPA05]. group-level [Chu08, MNL^{+}07]. Groupings [BW01]. Groups [BZ00, CKJ^{+}02, BBK04, CLMMP06, FZ04, LAG09, MM09, Rot09, SA07b, Toy09, VN08, dFG08]. grow [PBC09a, YOYT07]. Growing [BLS06, For02, KMH00, BL09, HI07, Har07b, JCW^{+}03, KS09, KY03, MSMKM06, SZLK^{+}05, SSD09, WBR08]. Growth [AMW00, BPV01, BP01, CAF03, CABB09, FFN00, HHR01, IKS00, KS01a, KTCD00, KTH^{+}00, KiHM^{+}03, Kra01c, MLMW01, MSI01, PGLG01, RFH^{+}02, RNP04, SS01, VG01, Wel00a, ALSM06, ABM03, AG06a, ARR08, AM04b, AM06, AD06b, BS05a, BKMH07, BM06, BR06, BCKE^{+}08, BC06, BGD^{+}06, BdOP06, CKS04, CW07, DCP^{+}08, DF08, ES08, EHG03, EH08, FPM^{+}06, GA07, GA09, GT06a, GT03, GCH^{+}07, GDD^{+}03, GDC^{+}06, GC09, HEH^{+}09, HVN07, KGG08, KH07a, KH09b, KRB05b, Lap03, LKW06, LDW09, Lo07, ML07a, ML07b, MAL03, MFI09, MHKS03, MHKS04, MJ07a, Nar06a, NP07, Nie06, NP05, Osb08, Paw07b, Paw09b, Pie09b, PA09b, PA09c, PSJ04, QASL08, QSAL08, ROR05, Rej07, RSC^{+}06, RLCIB05, RBP^{+}09, SMPM09, SGGM05, SMCT08, SSD09, SZLM09, SKS09, SP07b, TTT09, TKM06, VN07]. growth [VHF06, Wak04, WWS^{+}06, WBH04, WBR09, WLFC08, WAL06, Yan08, YHM^{+}06, YS07, Zac09, ZAdlPLM07, ZHB04]. GTP [BNT^{+}00]. GTP-Synthase [BNT^{+}00]. gTWH [Kan06, Kan07]. Guanine [NTK02]. Guide [Gin03]. guided [AJSL07, PER03]. Guo [Sch08b]. gusts [USTG09]. gut [PBC09b].

H [NM08, RGF^{+}08, KG06b, GB01]. H-NS-mediated [RGF^{+}08]. H-transporting [NM08]. H. [CS09, DT07a]. Habitat [FR06, Kaw01, Kon03, OKTS02, PPL^{+}00, Sch00, SFS^{+}01, AWO^{+}09, FL09, ML08a, ML08b, ZJ06]. habitat-based [AWO^{+}09]. Habitats [WZL00]. Haemophilus [LKC07, Mac00, PPEøP02, SP00]. Hair [HBLG02, MDCC06, MAB00]. hairpins [JPJ06]. hairs [CGKC07]. Haldane [For00a]. half [APL08, dQW07]. half-lives [APL08]. half-sites [dQW07]. Halobacterium [CR02]. Hamilton [Row06, vV07]. Handedness [Jam01b, Kel01, Laz02, BFR05]. handicap [De 03]. Handicaps [NS03b, Szá03]. handled [KV04]. handling [Saf09]. hantavirus [AWO^{+}09]. haplo [YY06]. haplo-diploid [YY06]. haploid [Fuk05, JWH08, SP05b, TSS06]. Happens [SSA00]. hard [NGMS08]. Hardness [AHCN07]. Harmonic [Raf02, Rak04]. harsh [CQLV^{+}03]. Harvesting [Bra07, Nik02, MS08b]. having [CMG06, SI09]. hawk [AKdlPP06, NS03a]. hawk-dove [NS03a]. Hawks [Cro00]. Hayflick [GKB03]. Hazard [HPZ09]. HBV [WWS08, XSD^{+}05, XM09]. Head [ACK00, MV02c, MBD08]. Head-on [ACK00]. head-raising [MBD08]. heal [BLT03]. Healing [BPV01, CS00, CLH07, GGV^{+}08, GBGAKD05, IvDHI08a, IvDH^{+}08, IvDHI08b]. health [AAK^{+}09, CHCC^{+}04]. healthy [vAR09]. Heart [DO01, See00, GG09, OD03, SNCM09]. Hearts [Hok00]. Heat [BKF00, Kan01, See00, DFW^{+}07, LMF03, PGN08, SZ09]. Heating [See00, HD08]. Heavy [Sco00, GVB^{+}08, IPY07]. Hebbian [RCA09]. Height [Hel08, MS06a, Pen03]. Heinrich [NM08, CB08b, CB08c, Maz08, MH08, Rap08]. helices [HK05]. Helicobacter [JK04]. Helix [ZSZ01, Kur07, Kur08b, YDL06a]. helix-helix [Kur07, Kur08b]. Helix/Strand [ZSZ01]. Helle [Jam08c]. helminth [BIB07, PBC09a]. helminths [PBC09b]. Help [EKIL01, WJ01]. Helper [AWJ02, Lan00a, Wod01, YBV^{+}00]. Helper-dependent [Wod01]. Helper-independent [Wod01]. helpful [BdSFFDMC09]. Hematagenous [Ano01g]. Hematogenous [Ken01, Ken02]. hematopoiesis [CM05a, CM05b]. hematopoietic [RG06]. Hemimelia [Hay01]. hemodynamic [GMK06, KHE06]. hemoglobin [TGV^{+}07]. hemoglobin-based [TGV^{+}07]. hemoglobins [Sar04]. hens [PMB06]. heparan [FWCN05]. Heparin [FFN00, MJW00]. hepatic [BKM09, BGG^{+}09, TWO^{+}09]. Hepatitis [TBR08, AM09, CRNP07, DLRP07, Jam07]. hepatocyte [SMPM09]. Hepatocytes [GTG01, MGP^{+}08]. hepatopulmonary [CBB07]. Herbivore [LLF02, KYS06, ZJ06]. herbivores [Kon03]. Herbivory [Gro02, Bal04, HFY07, IS09]. Herd [MKN02, VMW01, VMW02, GLE^{+}09, IGHW07, LGS^{+}09, NDE06, RV05, TBB^{+}06, TBC^{+}08]. herds [JBK04, Wil08a, XBCF05]. hereditary [GV03, TG09a, dBCP^{+}07]. Heredity [KY02]. hermaphrodite [PM09]. hermaphrodites [CH07]. hermaphroditic [Sat04]. Hermaphroditism [EK07, EK08, GR08b]. hermit [SS06b]. herpes [HHP05, NS09a, PBHS05]. Hes1 [MM08b]. Heterochromatin [Gru00]. Heterochronic [CACC02]. Heterochronies [Cub00]. Heterochrony [Kov02]. Heterogametic [For00a]. heterogeneities [CPM^{+}09, KUK07]. Heterogeneity [BR00, CC01a, FLS01, Laz03, RMRG09, TTKZ01, CL05, CCM07, DI09, DDJ06, GKB03, HDF04, KJ08, MI08a, MI08b, TBC^{+}08, Wil08a]. Heterogeneous [CAF03, Eng07, Gre00, KF03, MNL^{+}07, PCS^{+}06, SKN^{+}03, SA01, BBPSV05, BGO08, CGF^{+}08, CV08, GGPFM08, HSMM07, Kao06, Lin07, Man06, MSH05, Mas08, MSK03, PWK03, Pic06, SA05b]. heteronomous [HK05]. heterotrophic [MRC09, MKE^{+}09]. hibernation [HA07]. Hidden [JW01, WT01]. Hierarchical [Ait08, Cha01a, FFHK09, MSS05]. Hierarchy [BC02, Bro02, BMH07, CML08, DMH08, GGPFM08, SW04a]. HIF [DMPP09, KSB08]. High [CD05, Dus00, FTG07, Laz03, NS03b, TSK09, VTL05, WKL01, AC04, Dus06, HHH06, KJJB07, LJ09a, LAG07, ØKRG04, RMRG09, TBR08, Xie09]. High-dimensional [CD05]. high-dose [ØKRG04]. high-fidelity [Xie09]. high-grade [KJJB07, LAG07]. High-pressure [FTG07]. high-throughput [AC04]. Higher [GOP09, LY01, Nei01, RPB03, SL00, BRC^{+}03, GS06, RGPB08]. Higher-Order [RPB03]. Highest [ISWT02]. highlands [Ait08]. highly [MS03a, Ped07, ZCD05]. Hill [TW04, TSRB08]. Hill-type [TSRB08]. hindrance [TN04]. hip [CTB09]. histological [Hol06, Izs05]. histone [Aba09, LDW05]. historical [SB07]. Histories [JE01, RTK02]. History [KKK00, Bul06, CW08a, MVS^{+}06, MN07a, SCH05b, VA04b, Wit03, YKG^{+}05]. hit [LCL07a, LCL07b, LL09]. Hitchhiker [Gin03]. Hitting [Mcn00, MN01]. Hitting-distance [MN01]. Hitting-time [Mcn00]. HIV [HM05, LGK^{+}12, AWJ02, BRND09, CRB05, GSRC^{+}06, HW05, HW06, HLS01, IMNT09, KW07, Lan00a, LGK^{+}09, LYZ08, MBH03, MPOBD^{+}09, MD08, MGP00, PPD09, PD01, RB02b, RGFP07, RP09b, STK08, Smi08d, Sne03, SCC^{+}00, SRWL02, WSC02, WS02b, Wod01, ZT06, vLHH06]. HIV-1 [LGK^{+}12, BRND09, GSRC^{+}06, LGK^{+}09, MPOBD^{+}09, PPD09, RGFP07, Sne03]. HIV-directed [WS02b]. hive [AO03b]. HLH [DFCL08]. Hodgkin [LSS06, OUPG09]. Hoffmann [Sch02b]. holding [Hur06]. holds [AH03b]. Holistic [Kal00]. Holliday [Xie07]. Holling [PCZL06, PCZL05]. Holothurioidea [Kal00]. home [GAK^{+}06]. homeoprotein [HKP07]. Homeostasis [ESGK02, MRGADD08, Paw07b, Paw09b, PAH06, PC04]. Homeostatic [SML02, DdB03, MKS^{+}09, YLM03]. Homing [Rei02b, Wal00, GRBR04, RGRB04]. hominin [Syl06]. Homo [SKR06a]. homoeostasis [SMH07]. Homogeneous [KF03, LP00, GA08b, JAHKH09, Lew03, Saf09, TDW07]. homogenisation [HGP^{+}07]. homogenization [SJD^{+}09]. Homologous [FK03b, OKS01, HFS06, LZ09a]. Homology [SN05b, WS03a, For07b, LGB03b, YTL08]. Homosexual [MKN02]. homosexuality [Bla04, Bog04]. Homotopic [JM07]. Honey [BHOR01, SB00, CM03, GHHR03, HD08]. Honeybee [AO02, AO03b]. Hong [CFCGCC03]. hooks [Kee05]. Hopf [BGF07, GK09]. Hopfield [PD00]. Hopping [ITT09]. horizon [VCBV^{+}06]. horizontal [KZ05, KH09c, WS09]. hormesis [MV07]. Hormonal [Jam00, SHP09]. hormone [Leo07]. Hormones [Jam01c]. HORs [RGPB08]. hospital [UM08]. hospitals [DMO^{+}07]. Host [BH01, FMD01, GS02b, KT01, LAD03, PGM00, RSBY03, SCH02c, SMG01, WKB07, WC02, ADHM09, AG06b, AM06, CVH03, CBR04, CW08b, DT07a, DMH08, DLM08, ETH04, FS04a, FLG^{+}09, Fra00b, FPL03, GCP04, HI07, ING04, KSB07, KBD06, MCC^{+}09, MWB05, MRF07, PCB07, PW09a, PCS^{+}06, PSWK09, RDF03, RRS06, RGFP07, SH03, SCH05b, SI04b, SN07, TYW05, VWR07, Wil08a, WM04, XBT09, ZSF^{+}07]. host-cell [PW09a]. Host-limited [SMG01]. Host-parasite [WKB07, FS04a]. host-parasite-hyperparasite [MRF07]. Host-Parasitoid [KT01, CBR04, KBD06, PCS^{+}06, SN07]. host-parasitoid-microbe [SCH05b]. host-pathogen [RRS06, XBT09]. Host-plant [PGM00]. host-sanction [MCC^{+}09]. host-symbiont [SH03]. host-tumour [AM06]. hosted [Di 09]. hosts [BB03, JE09, MB06, MKE^{+}09, NdGG06, OSK^{+}05, PBC09a, PBC09b]. Hot [MS03b, DMW08, DCC^{+}08]. Hough [ZLXY08]. house [PMB06]. Households [ACCC00]. Howland [CS09]. HP [YAL04]. HRV [JDMZ^{+}07]. HSV [NS09a]. HSV-1 [NS09a]. HTLV [CMB^{+}01]. HTLV-1 [CMB^{+}01]. Hubbell [AK09]. Human [Ano01g, AB00, Gin00b, GS00, Kan01, Kel01, Ken01, Ken07, Lin01, LVV^{+}01, OB08, RPB03, Tho02, VG01, Wel00a, Wel02, Alp05, BM06, BLZ07a, BLZ07b, BT06, BBSLN06, BBSLN08, BBB^{+}07, BOCF08, CE08, CBC^{+}09, DBBW09, DBBW11, DZB^{+}04, DLF^{+}07, EH08, EG07, For09, Gar04, GT08, HSC07, HK05, HdGH07, HM05, Hop06, IB06, Jam08b, Jam09b, JK04, KKR^{+}07, LWFP08, LWRK07, LDW09, LW08a, LSMZ08, Mag04, MK04, MGT^{+}06, MMUGD09, PRP^{+}03, RD07, RKOS09, Rot09, SP07a, SdAC^{+}08, SB07, SKK^{+}07, VGS^{+}05, VGDSU09, Won06, YZGW06, Yan09, Yun05, ZYW07, dBCP^{+}07, vV06]. human-environment [RKOS09]. humanity [Lac09]. humans [BFR05, BB07b, Jam06, MS03b]. humoral [FABdC04, NKL06]. Hunchback [HMB^{+}08]. Huntington [BSGT08]. Hurst [VLFN00]. Hutchinson [MBJ09]. Huxley [LSS06, OUPG09]. Hyaluronan [Bra01]. Hybrid [For00a, AIKP06, CL07a, GA07, GA08a, GST08, JAJH07, JHJ09a, JHJ^{+}09b, PBB03]. hybridize [For07b]. hybridizing [CZC05]. hybridogenesis [YY06]. hybridogenetic [Arc05]. hydra [Ber03]. hydraulic [DFCL08, LMVPM07, MV06]. hydraulic-photosynthetic [DFCL08]. hydrocephalus [FTG07]. Hydrodynamic [KCT02, BBKGP08, Usa06]. Hydrodynamics [KS01a, Ver04, Mig06]. Hydrogen [Tho02, GMFS06, PSSY09]. hydrolases [dCZJ^{+}04]. hydrophobic [SPN06]. Hydrophobicity [Sil02, KSR07, LFF06, ZDC08]. hydrostats [LME06]. hydrothermal [BBKGP08, MB09]. Hymenoptera [Kro08, YY06]. Hypercycle [JL00]. Hypercycles [ZSS02, SS06a, SF08a]. Hyperexponential [VG01]. Hyperglycemic [RBBH02]. hypermutation [KS08b]. hyperparasite [MRF07]. hypertension [SVG^{+}08]. Hyperthermophile [Di 03b, Di 03a]. Hyperthermophilic [Di 00b]. Hypertrophic [CS00]. hyphae [TKM06]. hyphal [GT03]. Hypocalcemia [ESGK02]. hypophagia [Alp05]. hypothermia [GMK06, GK06a]. Hypotheses [GZ04, Jam04, RGG00, Cor05, FBM06, PHP03]. Hypothesis [CPC^{+}00, FK01c, Gru00, Hay01, Jam01b, Jam01c, Leh02, Nas01b, SS02, Sug02, VP01, Wel03, AO03b, Arc07a, Cog07, CSS06, HH04a, Jam07, Kan05, Kan06, Kan07, Kor09, MCC^{+}09, RKH08, Won05, Won06, ZAdlPLM07]. Hypothesized [Jam00]. Hypothetical [SVG^{+}08, CAHH06, TMH04, USTG09]. hypoxemia [CBB07]. Hypoxia [YFKP03, ABM04, DMPP09]. hypoxic [OBL04]. Hysteresis [TN01].

I/type [OAKC08]. Ia [CMPL^{+}00]. IbM [PGF^{+}08]. ICAMs [WSM05]. icosahedral [PT09]. ictal [FDA^{+}09]. Ideal [Bea00, FR06]. Identification [BS00, JWWS08, LKW06, BGE06b, Cud05, FYX^{+}09, HQP08]. identified [Nos06]. Identifier [Ano00a]. Identify [dCZJ^{+}04, DRLB05, LTLM09, VHF08]. Identifying [RH08, SJGK04]. Idiopathic [Her00]. idiotypic [HBS07]. IES [PER03]. IESs [EPR07]. If [MH05, BEBV03]. IFC [Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano04-29, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30]. IFC- [Ano04-29, Ano04z, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30]. IFC-Editorial [Ano03o, Ano03p, Ano03q, Ano03r, Ano03s]. IFN [RWCK08]. IFN- [RWCK08]. II [BKE04, LH08, LH09, PCZL06, TT03, BLMV05b, BGE06b, BKE03, BCV00, CM05b, DSY01b, DRV^{+}06, DLM08, DPA05, FSG00, GRBR04, Han02, KI05a, Laz03, LIK^{+}05, LJ09b, LGB03b, LLW09, LCL07b, OF05, OAKC08, PCZL05, Pep04b, PD01, QA01, QSAL08, RNS04, SKW00, Smi08b, Sta00b, TT02, Tho02, vAGDR09]. III [ARR08, LL09, RB02c, SHF02, Smi08c, Sta00c]. IL-12 [Kli06]. Illumination [Laz03]. illustrate [DBBW09, DBBW11]. illustrated [RB08]. illustrating [NGN^{+}04]. image [SA07c, XSD^{+}05, XWC08]. images [MR07b]. imaging [CLM07]. Imitation [Kra01a]. Immediate [ADMZ02]. immersed [Rej07]. immigration [ZKFW08]. immortality [ABTR07]. Immune [Fra02, HI00b, IMNT09, AB07a, BRND09, BdOP06, CTB^{+}05, CCC08, DKLL05, FJBK05, GWM04, HSC07, HML09, HM05, IMN04a, KLL07a, MD06, MK04, NKL06, NGN^{+}04, SI09, SI04b, VWR07, WL07a, vdBR03]. immune-mediated [VWR07]. Immunity [AN00, GG03a, CKM05, FABdC04, FMLP06, FGMP08, GRH^{+}07, LLC03, NDE06, RMP08]. immunization [HRSV06]. immunobiology [LGCL07]. immunocontraception [DBBC08]. immunodeficiency [HM05, IMNT09].

Immunodominance [Lan00a]. Immunogenicity [Cha00, vdBR04b]. immunogens [HM05]. immunoglobulin [SBZ^{+}08]. Immunological [UB01, AB06, Ort06, dVG04, dVG06]. immunology [AB03, DHR^{+}07]. Immunopathology [BKE03, BKE04]. immunoregulatory [Agi04]. Immunosenescence [LVV^{+}01]. immunotherapeutic [CEA07]. immunotherapies [JWB^{+}09]. immunotherapy [CP07, CA09, dPGR06]. Impact [AAEW09, FGMP08, GG03a, HKC^{+}07, PBvdG09, vLHH06, AMD05, AMD06, DMO^{+}07, DBBC08, DMQ04, EBId09, FMLP06, GKNT09, GEF04, HFS06, LGS^{+}09, MD03, MPM07, MMV^{+}04, MPN07, PW09a, RB06, SGG^{+}07, SML07, SML08]. impacts [BTCD07, HH07]. Impaired [CCP^{+}00]. impairment [IMNT09, WL07a]. impedance [BZ05b]. impedence [BZ05a]. imperfect [Fis03, Won06]. impinge [LGCL07]. implants [MGAD09a, MGAD09b]. Implementation [BL06a]. Implication [CPC^{+}00, Kan07, WZW^{+}07]. Implications [Aba09, BHHS01, BP00, Bea00, CTB09, GS02c, GRH^{+}07, GT08, GJ07, Hay00, Her00, LVV^{+}01, Pic06, RR08, SHF02, SPN06, Sne03, WRKK09, Wod01, dQW07, BNRW04, Di 04, ETTV08, FLS^{+}04, Gol08, GKB03, KRN03, LIK^{+}05, LRD04, MAC06, PCB07, PCS^{+}06, San03, SKR06a, WL04, Yag09, ZSF^{+}07, dBCP^{+}07, KS08b]. Implicit [BMR08, ELL04, LSD^{+}07, WMS08a]. implies [ZW03]. imply [MV06]. Importance [AWJ02, LY02, Bal04, For07b, Hul05, ING04, Lin04, LMS^{+}08, NLS08, PB03, SW08c, TM06]. Important [LL00, ZCC01, IvDH^{+}08, Kli06]. imposed [LB06, SSR04]. imposition [Win06]. imprinted [Cui07, SP06, Wil06b]. Imprinting [GK00, WH03, Kro08, Wil06a]. imprints [SLL06]. Improved [Sav04, WTL08, CL07a, CLS08, DFW^{+}07, LSS06, VHF08]. improvement [Cud05, Han04, LZ09b]. improves [CPMG^{+}08, HFY07]. Improving [WS09]. Impulse [Tas02, RGB06]. impulsive [GC09, LWC09, LYZ08, PCZL05, PCZL06, SX06, WZL^{+}08]. In-phase [ZW03]. Inaccurate [RE04]. inactivating [WCLL08, WCH^{+}09]. Inactivation [BP01, JL01, CPM^{+}09, Fer09b, GS08a, MI06]. inattentional [Wal07a]. Inbreeding [HP04, KY00, Lac09, WH03, Pec06, UH09]. incidence [PL09, ZKFW08]. inclines [SB09b]. include [HW06]. including [BDR08, GW06, ØØS06, PP04, RPNH03]. inclusion [XFBC07]. inclusive [FZ07, Gra06, TDW07, vV09]. incoherence [Rot07]. income [SH05]. incompatibilities [GPG07]. incompatibility [Sat04, TYW05]. Incompatible [Cur04]. Incomplete [DW08b, AAK^{+}09, RYAI06, Sta09]. Incorporating [CXM^{+}09, Cog07, DSA^{+}06, HS00c, JJ01, VHF06, IvDHI08a, IvDHI08b, LVL08a, LVL08b, MX08, QLHL09, SA05a, XK07]. Increase [HHR01, AF09, Fow09, GMFS06, Ker04, Ker06, TTN05, WN05]. Increased [GTDA02, Her00, LKM^{+}08, PGLG01, PL01, Tan07c]. increases [Fer09a, IMN05, TNP07]. Increasing [Fow09, WL09, MM08a, ZKFW08]. incurred [Roe07]. Independence [WS03a]. Independent [LD09b, ACSY04, ABP^{+}03, ELSFB07, WW08, Wod01, ZW03, dV06]. indeterminate [DF08]. Index [Ano03a, Ano03b, Ano03c, Ano03-44, Ano03-45, Ano03-46, Ano04i, Ano04d, Ano04e, Ano04f, Ano04g, Ano04-62, Ano04-57, Ano04-58, Ano04-59, Ano04-60, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05-57, Ano05-58, Ano05-59, Ano05-60, Ano05-61, Ano05-62, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06-55, Ano06-56, Ano06-57, Ano06-58, Ano06-59, Ano06-60, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07-55, Ano07-56, Ano07-57, Ano07-58, Ano07-59, Ano07-60, Ano08a, Ano08b, Ano08c, Ano08d, Ano08-52, Ano08-53, Ano08-54, Ano08-55, Ano04h, Ano04-61, Can07, CTB09, GMMR07, Hel08, KDKS06, LW07, RAHO06]. Indicator [PKL01, CF09, LPT05, NAV04, RKH08]. Indicators [PH03, KP08]. Indices [DOT02, FV09, Gos06, KSR07, MPOBD^{+}09, MGDM08, MGDBdM08, MMUGD09, PBMU^{+}09]. Indirect [Fis03, WH01, BS04, BS06, MM03b, OI04, Oht04, OI05, OI06, PB03, RS04, SA07b, SA07c, SA08b, TM06, XFBC07]. INDISIM [GLV02]. Individual [AMW00, FVP^{+}07, FSGB^{+}02, GLV02, HK07, KFG^{+}02, KG00, MAHD06, PLG^{+}06, SMPM09, SL00, DPV00, AM09, AV05, BG08, BCH03, BEK^{+}03, CGF^{+}08, CM03, CSM05, FM07a, FHL^{+}06, GA09, Gol08, GKK06, HWCF07, JCW^{+}03, LSH06, Rej07, SC05, Sch08a, SCH05b, TW04]. Individual-based [FVP^{+}07, GLV02, HK07, KG00, PLG^{+}06, AM09, FM07a, GA09, GKK06, SCH05b]. individuality [Unn09]. Individuals [DMZ00, Gar02, Arc09a, FT09a, Fis03, ZKH^{+}05]. Induced [CCP^{+}00, DO00, Han02, Laz03, MD01, WKG03, AN09, ALSM06, AMV^{+}02, Bar01, BW06, BKP09, DLRL08, DT07a, DS05, FCD^{+}05, GFWT04, Has06, HH04b, HS08, HSMM07, HHR01, IHV^{+}06, IS09, KH09a, KNWCB07, KDK02, LK08b, LPJB^{+}08, LFP^{+}05, LCL07a, LCL07b, LL09, MAC06, OAKC08, OQGC07, RMBM00, RDL07, RSB09, SHPDL03, SdAC^{+}08, SRMW00, SB09c, SNT03, Ste04, SN04, SM09a, TWE04, WHH07, WTC09, YMC01, ZTKH09]. Induces [SL00, NS07, YZHZ09]. inducibility [PBH04]. Inducible [SI01, SI04c, PE04]. Induction [GVB^{+}08, SCGFS00, GZL^{+}03, HJ05, Nar07, NKC^{+}08]. industrial [PMB06]. inelastic [GA08b]. inequity [BDMR06, MGL05]. Inert [WGH01]. inevitable [RHF07]. infant [BBK04]. Infanticide [BBK04]. infants [SGD04]. infected [FVP^{+}07, GCP04, LGS^{+}09, MSDM06, PPD09]. infecting [FLG^{+}09]. Infection [AWJ02, AB00, BKE03, ETH04, GWM04, ISC01, MGP00, PG00, PD01, PV00, Sne03, SCC^{+}00, WSC02, Wod01, BRC07, Bea06, BKE04, BIB07, BRND09, CRNP07, CSD09, CTS06, EZM05, GB07, GZK06, HSC07, HW05, HW06, IMNT09, LGK^{+}09, LGK^{+}12, LT06, LYZ08, LTLM09, MSDM06, OSK^{+}05, PL09, PMB06, RWCK08, RMRG09, RRS06, SJGK04, SDL^{+}08, SI04b, SCS04b, Smi09, SK09b, TBR08, TBC^{+}08, UM08, WC07a, WL07a, WS04, Zhd03, vLHH06]. infections [BST05, CKM05, CMF08, RPNH03, VCGV07, XBCF05]. infectious [FYX^{+}09, HDF04, JEDH08, KB04, Sch03, TP05, Yan08, dM09]. infectivity [AHT^{+}07]. Inference [AKLS05, FB01, BMR08, Fra03, NS03a]. inferences [FBUL05]. inferred [Kun05, Kun06]. Inferring [Bra05, CWJ07, GNLEK07, PHG04, RLCIB05, Ste09]. infestation [LSD^{+}07]. infested [KY03]. infinite [Fuk04, Fuk05, GL09, VCBV^{+}06]. infinite-time [VCBV^{+}06]. inflammation [KCVC04, RRC^{+}06]. inflammatory [Agi04, DRV^{+}08, DRV^{+}06, RK06, RRC^{+}06]. inflated [CY09]. Influence [ASD08, CLMMP06, DHM01, GBGAKD05, KS01a, KNC^{+}04, Kaw01, LP01, OKS01, PD01, SVS^{+}02, Tof06, WH00, AB07a, BBKGP08, BCV^{+}08, BBD06, CB07b, DP04, Fox05, HF03a, HBWB08, KHL07, MJ07b, Ort06, SZ04, VN08, dVG04, dVG06]. Influences [BPV01, SBH01, TTR00, CW08b, HK09, HBS07, LDW05]. influencing [LTLM09, PB07]. Influenza [LF01, ABvdD^{+}08, CAHH06, HLA09, HDZ^{+}07, ITKL08, ITLN09, JITJ09, MSDM06, PLH05, RDC09, XAP07, BST05, BMD^{+}08, CY07, DHW^{+}09, HSC07, LACL03]. influenzae [LKC07, Mac00, PPEøP02, SP00]. influx [MM03a, NPN09]. Information [AMW00, BJ02, CDDW02, Car02a, FPS01, Gag00, JW01, KMP03, LFSG^{+}05, PN00, RC00, SSDM06, WJMH00, ZSS02, dM09, AVSHV04, Bon04b, DL08, FCP03, FK03c, GDPMDS^{+}09, JL06, Khr04, LH06, MM03b, PSM06, RHF07, RB09b, SPAH06, SLIL07, SC09, SF08a, SF08b, SSM09, ST05, Tlu07, VN08]. information-processing [Bon04b]. Information-related [dM09]. Information-Theoretic [CDDW02]. Informed [Wil09]. Infrared [JLS01, Bak07]. infusion [NKC^{+}08, SW08c]. ingredients [GKTN07]. ingrowth [MGAD09a, MGAD09b]. inhabited [SA05a]. inherent [MGL^{+}06b]. Inheritance [DRMLS09, HS00b, SSKL01, SLL06]. inhibit [MR03, PGH^{+}04]. Inhibition [BHOR01, Has01a, RGG00, AIKP06, CKS07, Hua03, JPJ06, Nar06b, SZ04, TKP07]. Inhibition/Stimulation [Has01a]. inhibitor [GABK08, MD08, WZW^{+}07]. Inhibitors [SHHD02, NLM^{+}08, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08]. Inhibitory [EMS02, Wil01, CRB05]. inhomogeneities [LMF08, WBR08]. inhomogeneity [KC03]. Inhomogeneous [WGH00, ABM03, VGBA06]. Initial [Nir02, The00, LIK^{+}05, WG06, YS07, YFK05]. initialisation [BGB08]. Initialization [HLS01]. initiated [USTG09]. Initiation [BG00b, FKS07, PMMS01, CF07, KSN03, ØØS06, SWC^{+}08, VSP06]. Initio [NTK02]. innate [ZKH^{+}05]. inoculum [WG06]. inositol [GFW^{+}09, LGB03a, TLZ05, WMS08b]. Input [DFC^{+}02, FLBB01, KI05a, VL09]. Input-output [KI05a]. inputs [Fen03, FL03]. Insect [GTDA02, OCP^{+}00, WB06, BNRW04, CC09b, EZM05, KH09c, LBF01, LSD^{+}07, PM07, PGH^{+}04, Tof06, WRNB04]. insect-resistant [PGH^{+}04]. Insects [Möl01, Möl02, NC02, WC02, Bye09, Iim07, MMV^{+}04, PRT04, PGH^{+}04, XBT09]. inseparable [GD06]. insight [Leo07]. Insights [GZK06, KP06, KS01c, OHM05, ACK08, AM04b, LS07, RG03, SP05a, WZW^{+}07, CSD09]. Inspired [WGH00]. Instabilities [BGF07, BSS^{+}07, Mal00, WW05b]. Instability [GB02, PKL01, WAMO00, BMS06, KSN03, LVL08a, LVL08b, NMI06, VTL05]. Instars [HYA02]. institutions [BCH03]. Insulin [RB02c, TMS00, TPD^{+}00, WQ00, BM06, KR05, LKM06, OHM05]. insulin-like [BM06]. insurance [GRW03]. intake [KRGH07, LPJB^{+}08, SZLK^{+}05, SSD09]. Integral [ESGK02, SKW00]. Integrate [FLBB01, LL05, UL06, VL09]. Integrate-and-Fire [FLBB01, LL05, UL06]. integrated [Ben04a, GMK06, MB05, WWS^{+}06]. Integrating [HWCF07, VHF08, VTC08b]. Integration [AVSHV04, SS06c, BFP07, MRA06, PCB07, SV07, VdL03]. Integrative [DF00, DHR^{+}07, IST01, Jaz01, HCR^{+}04, JB04, Muz05]. integrator [Fen03]. integrin [CLB05]. integro [WW05b]. integro-difference [WW05b]. integument [LDT09]. intellectual [Rap08]. Intelligence [JST^{+}02]. Intensity [Ghi02, DGZ07, HHH06, TT04, Wal08]. intensive [MPM07]. Intentions [JST^{+}02, PTFF05]. Inter [DSCD02, FMI05, LJK06, Lit07, RA08]. inter-compartmental [Lit07]. Inter-dependent [FMI05]. inter-residue [RA08]. Inter-Sarcomere [DSCD02]. inter-specific [LJK06]. interact [FS04a, HHP05, HDZ^{+}07]. interacting [NK08, PS03, RB06, Sch06, SV07, SLB07]. Interaction [Cum01, CGKC07, Ell01, EN03, IST01, ISWT02, ISC01, Jav00, MJW00, NM08, Nir02, RGSFM07, Sta00a, Sta00b, Sta00c, AR08, BdSFFDMC09, CdOWS04, DF09, GEF04, HSK^{+}08, IS08, JK04, KJJ07, KST07, LN03, Leo07, MB06, ML04, MA04, MRF07, NL04, Nos06, OS06, OPN07, PSSS03, Péc05, PDPL05, RG06, RSS04, SGCK^{+}08, SIHH04, SB09a, SB06, SVN^{+}05, WT09b, Yan09, ZAD07, vKPdR07]. Interactions [BHHS01, CGH01, FLS01, LPLC00, LAD03, AAL08, Aba09, ACLW03, AO03a, ADHM09, AW06, AJ08, AS07, AM06, AB07b, BTCD07, CKE06, CdOWS04, CEP05, CMG06, Fox05, Her09, HGH08, JCK09, KSY^{+}08, KOT07, LJK06, LTG^{+}04, LFF06, LW07, MD06, MSGS08, MGM07, MP09b, PAA07, PCS^{+}06, RZF03, RDL07, RS04, RRS06, RDSB^{+}03, SLIL07, SB09c, SS05b, TC09, TBCD06, VWR07, WKB07, YHI07, ZSF^{+}07, ZZLT07, vLBJK07]. Interactive [Gol07, ZJ06, HNTHA07]. interactors [CE05]. Intercalation [HA00, BLS^{+}09]. Intercalation-Mediated [HA00]. Intercellular [GTG01, GG03b, TS02, UHI02, öWO01, GZ04, HKP07, JCK09, WWS^{+}06, WSM05]. Interchange [FW00]. interest [CC08, OB08]. interfaces [Kur07, Kur08b]. Interference [GG01c, JWB07, BZ05a, BZ05b, KC07]. Interfering [Fra00b]. interferons [HHP05]. interferons- [HHP05]. intergradation [KS06]. interleukin [CEA07, GSG^{+}07]. interleukin-1 [GSG^{+}07]. interleukin-21 [CEA07]. interleukins [CRB05]. Interlocked [UHK01, SSF09, WCA06, XK07]. interlocking [RCS05]. interlocking-feedback [RCS05]. Intermediate [FK03b, LBS00, CG06, LHDvdM04, PBC09b, RHH08, PBC09a]. intermittency [KvHMP09]. Internal [DF00, Ben04a, EPR07, GVB^{+}08, OB04, WE06]. Internalization [Gin03]. internally [OSK^{+}05]. International [Ano00e, Ano00-30, Ano01-33, Len01]. interneurons [SM09a]. Internodal [Dim05]. Interplay [Går00, Hog00, KH08, YCC02, CXM^{+}09, DI09, Zhd03]. Interpretation [RA06, SPH03, AH03b, FCD^{+}05, GRBR04, MKB03]. interpretations [dPGR06]. interpreters [Wil09]. Interpreting [Che00, GXG03, SLP00, MGT^{+}06, OR04]. interspecific [HB09, Sch05a, WLHB07]. intersubunit [NN07]. Interval [Sol01, LP07b]. intervality [RMAI06b]. Intervals [Hay00]. intervention [MGS09, TBB^{+}06]. intervention/prevention [TBB^{+}06]. interventions [CAHH06, GRH^{+}07, MPM07]. Intestinal [MF05, Agi04, MA07a, dT07b]. Intestine [ARW00, LBSP06]. Intra [WO04, PDM04]. Intra-membrane [WO04]. intra-sexual [PDM04]. intracarotid [NKC^{+}08]. Intracellular [AF01, FK01a, For01, GG03b, Jac03, Kar03, RWF01, SYSY02, VP01, WRF01, ASMD06, BH03a, DP04, HSL04, JST^{+}08, KSK08, Kur08a, MGP^{+}08, NS05, NS09a, WHH07, WMS08b, Yam03]. intracranial [GMK06, GK06a]. Intragenic [DZB^{+}04]. Intraguild [Har02, Rev02, Ama06, KHL07]. Intramolecular [AB07b]. intramuscular [Kok04]. intransitivity [LS09]. Intrasex [Tor00]. intraspecies [PDPL05]. intraspecific [MJ07b, PB08]. Intrinsic [GS08b, iKIY09, Tao04a, Tao04b, BS05b, Lei09, Lei10, Sil09, Yan08]. intrinsically [BE08]. Introduced [FGH01, GT06b, KC05]. introducing [SA07a]. Introduction [CW07, LHDvdM04]. Introgression [DBG01]. Introns [Eld00]. intussusceptive [SS05b, SKS09]. Invader [MD01, HH07]. Invader-Induced [MD01]. invading [DT07a]. invariance [LSAA^{+}06, PAA05]. invariant [Ken07, MV06]. Invariants [HL00, ZC06, SCJJ08, SJ09]. Invasion [BH01, MGL^{+}06b, PGLG01, TS02, BHWB05, GC08, GN07, IMN04b, KCP07, KCP09, KC05, LNH08, MM08a, MCM^{+}09, MMV^{+}04, MPL06b, SLHN06, SGGM05, TMBD09, VCGV07, WK05]. Invasive [Sch02b, RSC^{+}06, ZJ06]. invasiveness [SGG^{+}07]. inverse [CLM07, HB07]. Invertebrate [Har02]. invertebrates [CM09, MA03a]. Inverting [RB02a]. Invest [WNT08]. investigate [RSC^{+}06, RRH08]. investigated [VCGV07]. Investigating [GBK^{+}07a, UW09, BRS^{+}09, FWE06, LRB^{+}06]. Investigation [JBJ^{+}08, LK08a, TGR^{+}00, TR08, WH01, EHG03, GKB03, KSEK09, LJL08, LF08, THL03, WR07b]. Investment [SR02, YM04, Yan05]. invisible [GEF09]. Invoking [GK00]. Involuntary [Fis06]. involved [AF09, Cib08, HGH08]. Involvement [PBHS05]. involves [DCC^{+}08]. Involving [SR02, VGCGM^{+}02, MCM^{+}09]. Ion [FSG00, Lan00b, NTK02, WT01, DBG06, GKXS07, LSS06, Luc05, OP05]. Ionic [The00, YFKP03, LAG07, PAH06]. Ionizing [BL00, JL01]. ions [MS05a, MS06a, MS08c]. IP3 [FWLN04, HSF09]. iron [FFME08, LK06]. Irradiance [Han01]. irradiated [HHBY06]. irradiation [BKP09, GVB^{+}08]. Irregular [BNRW04]. irrelevant [Niw04]. Irreversible [FK01c, ML09a, ML12, TN01, BZ04]. irritation [BGE05a, BGE05b, BGE06a, BGE06b]. ischaemia [ACLH05]. Ischemia [YFKP03, LZS^{+}08]. ischemic [JK06, LAG07]. Ising [SI04a]. islet [JCK09]. Islets [AMV^{+}02, WB04]. isoforms [SNA^{+}08]. isolation [CFCGCC03]. Isomorphism [VSCG00]. isotonic [CTS^{+}08]. isotope [RA06]. isotopically [RA06]. Isotopomer [KP08]. Isotropic [LP00, Lew03]. isotropy [SJ09]. isovelocity [TSRB08]. Issue [Ano06-54]. Issues [LFW02, OLS^{+}02, For07a, Her09]. Itô [Bra07]. Italy [Ano02a]. Iterated [KK00c, LBSS02, AD07a, LB07, LW04, Sch05a]. Iteroparity [RTK02]. Itself [Di 01b]. IV [Ano02a]. Ixodes [OBPH^{+}08].

J [AGW^{+}08, AMD06, BZ05a, BBM04, BKE04, CSP^{+}09, DBBW11, EPJ^{+}11, FK03a, Gie06, GWM06, HTN04a, IvDHI08a, JCRJ07b, KCP09, Ker06, LH09, LGK^{+}12, Lei10, LVL08a, LW04, ML07a, ML12, MT06a, ML08b, MI08a, NI07a, NI04a, Paw09b, PCZL06, PA09b, RH09a, RMAI09, SHI06b, SI04c, SRAL12, SML08, TK09a, TRM03a, TRM03b, TQUN08, VGMM^{+}07a, VGMM^{+}08, DqLmW07]. J.Theor [UI04b]. JAK [SYYI07]. JAK/STAT [SYYI07]. January [Ano00k, Ano00-29, Ano01j, Ano01p, Ano02d, Ano02j, Ano03u, Ano03-28, Ano04-34, Ano04-38, Ano05-55, Ano05-56, Ano06-35, Ano06-40, Ano07-31, Ano07-37, Ano08-33, Ano08-39, Ano09z, Ano09-32]. jaw [VAAH05, Wes03]. jitter [OUPG09]. Joel [Ano01d, Has01b]. John [Ano06-54, SS06d]. Johne [LMS^{+}08]. Joint [CS02, LD00, Mag04, PBH04, WRF01]. Jones [Bak07]. Josef [BZ00]. Journal [Ano05h, BLZ07a, BSRH03, HH05b, OI05, TT03]. July [Ano00i, Ano00o, Ano01n, Ano01-32, Ano02e, Ano02k, Ano03t, Ano03z, Ano04-49, Ano04-51, Ano05-35, Ano05-42, Ano06-34, Ano06-39, Ano07-35, Ano07-41, Ano08-34, Ano08-40, Ano09-31, Ano09-37]. jump [Mur07]. jumpers [SBvS06]. jumping [SBvS06]. Junction [MY01, BFGD07, DBB09]. junctions [Daw09, Xie07]. June [Ano00-30, Ano00t, Ano00z, Ano01w, Ano01-29, Ano02n, Ano02u, Ano03-42, Ano03-33, Ano04-56, Ano04-45, Ano05-45, Ano05-54, Ano06-44, Ano06-49, Ano07-45, Ano07-51, Ano08-49, Ano09-45]. Juvenile [FSGB^{+}02]. juxtacrine [WO04].

K-12 [LL06]. KaiC [TIM06]. Kanazawa [Gel07]. karyotypic [COdAM05]. Kauffman [IiKY07]. kDa [MW01]. Keiser [Ano01d]. Keizer [Has01b]. Keloid [CS00]. keratin [BEF^{+}06, BLF^{+}09]. keratinocytes [LN05b]. Kernel [CR01, STW^{+}09]. kernels [BdABA09]. Key [FW00, RPB03, Yos03, ELSFB07, LTLM09, SLHN06]. Key-String [RPB03]. keystone [FV09, CE05]. kidney [HIM09]. Killer [WS02b]. Killers [WSC02]. Kin [MR03, WT05, vV06, vV09]. kin-selection [WT05]. Kinase [SSB^{+}02, CLB05, FKAC06, FWCN05, KL09, RR09, SV07]. kinase/phosphatase [FB08]. kinases [HSL04]. Kinds [Fur02]. Kinematic [LP00, KH09a]. kinematics [RWID^{+}08, Sch08a]. Kinesin [MFB01]. kinesis [MBB^{+}06]. Kinetic [AGT^{+}01, ADMZ02, BG00b, CT02, CCT^{+}09, EP00, FTEG02, GAA02, KCP07, KCP09, KY02, MD04, OCA08, PJ01, Pro03, Sne03, TML02, TGR^{+}00, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, VGCGM^{+}02, VGMVR^{+}06, WKG03, WL02, Alb08, AVSHV04, BS05b, CHN08, KKR^{+}07, KRB05a, KZ05, LJ09b, Lei09, Lei10, LH06, LVABV05, MH06a, PCSL^{+}06, SB09a, SGD04, TKP07, WHH07, dPCP^{+}08]. Kinetics [Ano01h, BG00a, BSRH02, BSRH03, CCGC02, Dem02, DNS00, FC01, IM02, LBS00, Ril00, SYSY02, SRN^{+}00, TPD^{+}00, VLFN00, WQ00, BRND09, CPM^{+}09, CL05, DVC^{+}04, FWCN05, Jus08, MA03a, Nar07, OP05, SRR08, STSD09, TE04, TE07, dQW07]. kinship [Kro08]. klepsydra [WE06]. kleptoparasitic [BLRR08]. kleptoparasitism [YB07]. KNN [DCL09, SYC06]. knock [SVS04]. knock-out [SVS04]. knockout [LKM^{+}08]. knockouts [BWvK^{+}08]. knot [AKR09]. knot-limited [AKR09]. Know [Akt04, MH05]. Knowledge [BFH^{+}01, HNO03]. Knowledge-based [BFH^{+}01]. Knuckle [Kel01]. Knuckle-walking [Kel01]. Kong [CFCGCC03]. Krebs [NM08, MCWF01]. krill [GMMR07]. Ku [TML02].

L [FLS^{+}04, LCHK09]. L-selectin [LCHK09]. L1 [RRK06]. labelled [SW08c]. labor [DW08b, Tan07c]. Labour [Wah02]. Labyrinthine [Car02b, SI05]. lac [Nar07, NPN09]. lactate [DSA^{+}06]. lactating [BKD^{+}06, HCR^{+}04]. lactic [ZC09]. laevis [MKL09]. lag [BGK09, PLG^{+}06, PGF^{+}08, YS07]. lagging [MX08]. Lake [Har02]. lambda [LM09, PBHS05]. Land [KRNKH00]. Landauer [Smi08c]. Landmark [Möl01, Möl02]. Landscape [KCT02, AH03b, AHT^{+}07, AH08, Ait08, RRS^{+}07, SLIL07, WDH^{+}09]. Landscapes [OSBH02, FLG^{+}07, GPG07, RBJ06, SN03, Tan07b]. Langerhans [AMV^{+}02]. Langevin [SCS04a]. Language [KN03, PN00, CS06, Ded09, LJTD05, Loc08, Paw07a]. Languages [Cha01a, Ded08]. Laplacian [MHKS03]. Large [Ace00, BBB^{+}07, Kon09, LRHB09, Nei04, NS07, OCA08, Ots08, Péc05, RH04, SBZ^{+}08, Sim08, ZLXY08, ZW03]. large-population [Sim08]. large-scale [Ots08, SBZ^{+}08, ZLXY08]. Larvae [Arm01, Bur09, ILDP04, Ver04]. larval [BBKGP08, PBC09b]. laser [TR08]. last [Di 06b]. latencies [RR08]. latency [OUPG09]. latent [Ger09, GRBR04, JWWS08, KNWCB07, PBHS05, RGRB04, RP09b, Yan08]. later [MS09b]. lateral [PD06, SB09b]. latitudinal [Sav04]. Lattice [Ell01, PAA05, CXM^{+}09, Ezo09, GDPMDS^{+}09, KS06, NBT07, PBMU^{+}09, PM09, TBCD06]. lattices [BES06, WW05b]. Laurent [PG00]. Law [DML02, LY02, Can07, DRW01, Gol09, GDD^{+}03, Niw03, Niw04, Niw05, OBN07, PWK03, PPD09, San03]. Laws [Cha01a, KK00a, Tor01, WMLC02, Bio08, CY07, Dem06, Kai04, San03]. layer [BGE05a, BGE05b, LMF08, Lew05]. Layers [SBH01, EPJ^{+}09, EPJ^{+}11]. LCMV [BKE04, BKE03]. Lead [PL01, LJ09b, MGL^{+}06b, NP05, Tan07c, WVA05b]. leading [HVN07, LR07, MX08, OI06]. Leads [Sta02, CMW08, FWLN04, Uit09]. Leaf [Ant02, MC01, SBH01, CABB09, FI06, FM06, Pie09a, SG07, Whi05]. Leafless [Nik02]. leaky [Fen03]. Learn [PD00, Bok06]. Learning [Bea00, BFGS07, KF03, Roy00, Whi07, CW08b, DRLB05, EG07, Fox05, GAT07, KGL09, Kom04, MS09a, RCA09, RKH08]. least [MC01]. leaves [FMI05, GZL^{+}03, Wil08b]. leaving [HGC03]. led [Jam04]. left [Han04]. leg [RBS05, SH08b, WS09]. legged [Nis06]. Legos [CMW02]. legume [MCC^{+}09]. legume-rhizobia [MCC^{+}09]. lek [NP08]. Length [Koc00, KS08b, OF01, OY03, FS09, Fuk05, GHA03, GKB03, GP08b, GNLEK07, Jam01a, Kee05, KSPA^{+}08, LBS06, LW08a, Mei05]. lengths [CM09]. lengthy [OW07]. Lepidoptera [Cor05, CW08b]. Leslie [GL09]. Lessons [DL00]. Lethal [BR08, SW09, KJD05]. Lethality [GGK05, Kam03]. Letter [Gel07, Ret04, Ril02, Rot08]. leukaemia [Lit07]. Leukemia [AM01, FPL03, CM05a, DKLL05, ML04, PTD09, SHRR06]. leukocyte [PDB08]. leukopoiesis [ACR06]. Level [Han01, KRNKH00, AF02b, Chu08, HW01, LIK^{+}05, MH05, MNL^{+}07, PDPL05, PD04, RDC09, ZRSK07]. Levels [FMF^{+}00, Had01, Sta00a, AA04, Gra07, KOK06, LT08, Sel06, SI04b, TG09a, TSN05, WG03]. Levenshtein [BGG06]. Levins [EN02, RHH08]. Lévy [RB09a, RBP^{+}09]. Lewis [VS08]. Lexicon [Now00]. Li [BDMR06]. libraries [Kon09, TSK09]. library [BdBH09]. Life [CW08a, HA00, JE01, KKK00, Kor01, MVS^{+}06, RTK02, Wel00a, Wel00b, BPC08, Bat09, Bul06, Byw09, CMF08, CBC08, DM07, Gab06, GR05, KTH09, MBB08, MN07a, PBC09a, SCH05b, SP05b, UKY^{+}09, VA04b, Wit03, YKG^{+}05, Pro03]. Life-history [CW08a, MVS^{+}06, MN07a, Wit03]. life-like [GR05]. Lifespan [AO02, HLW00, ALSM06, GCP04]. Lifetime [YHG^{+}02]. Ligand [BNT^{+}00, GFWT04, KCT02, SRN^{+}00, TM00, CSD04, GSG^{+}07, HBB08, Jus08, TWE04, WO04]. ligand-binding [Jus08]. Ligand-Cytoskeleton [TM00]. Ligand-induced [GFWT04, TWE04]. Ligand-Receptor-G-Protein [BNT^{+}00]. ligands [vdBR04b]. ligation [AMP06]. Light [BPZ^{+}01, CR02, FI00, GRG02, KDK02, Laz03, MFI09, Nik02, RVMR01, Cal06, CB07b, DVC^{+}04, EBId09, GZL^{+}03, HHH06, KS08a, KG06b, PNG03, SKK^{+}07, VN07]. light-based [SKK^{+}07]. light-dark [KG06b]. Light-induced [KDK02]. Light-triggered [BPZ^{+}01]. like [AH03b, BM06, EK07, GR05, JB04, SM02, SH08b]. likelihood [BdABA09, Eti09, ZW03]. likely [LP08]. Limb [IV02, MSMKM06]. Limit [Har07b, Zah00, Alp05, GKB03, JR08, MP09a]. Limitation [TGP^{+}00, Eti09, FFME08, Ste09, WB06]. limitations [LB06]. Limited [JLS01, ALSM06, AKR09, BCL08, GvO04, JBK04, Lew03, MSI01, SMG01, WR04]. Limiting [SM09b, BM09, NL03a]. limits [ALM09, KS08b, Smi08a, WT07]. line [Chi07]. Lineage [PKL02, RG06, SBZ^{+}08, SLH^{+}09]. lineages [Ros03]. Linear [DFC^{+}02, MJW00, PAD00, DW08a, GBK^{+}07a, JAHKH09, MCK07, MGS08, MM09]. lines [DBBW09, DBBW11, DSVBW07]. linguistic [DT06]. Link [Gra02, PPR01, GR05, MDD06, RYAI06, vV07]. linkage [FWE06]. linked [GMM09, NL04, OW07, Sar04, SP06]. Linker [MFB01]. Linking [HMN09, YHM^{+}06, DSY01a, DSY01b, PTN06, PTON08]. links [MLPJ09]. lion [Bur09]. lipases [KSK08]. Lipid [CMPL^{+}00, RN07, AS07, PPGS03]. lipid-binding [AS07]. lipids [KI04, KI05a]. lipopolysaccharide [DRV^{+}08]. Lipoprotein [Sta00b, TWO^{+}09]. Lipoproteins [Sta00a, Sta00b, Sta00c]. liposome [PP08]. Listeria [IGHW07]. literature [Pán08]. litters [Jam09a]. Live [RH02, CLM07, SS06d]. Lived [Di 00b]. Liver [DLRP07, GBD00, CLW^{+}08, HCR^{+}04, KLL08]. lives [APL08]. livestock [EL09]. Living [AKS07, ZSS02, BDMR06, CW08a, Di 06a, KB04, MGL05, NDD08, WH07, YYA09]. Load [GLK^{+}02, MFB01, SNT03, CMB^{+}01, DP04, GMSW04, Kam03, Lin04, NGT05, NM08, STK08, SP05b]. load-based [STK08]. Load-induced [SNT03]. Loaded [BP00]. Loading [YNL01]. loads [GHHR03, LAG09]. Lobster [FTEG02]. Local [CL09, CC01b, GC08, JE01, LLCM01, McN06, Mcn00, MN01, YCC02, Ait08, AP04, BBKGP08, BdBH09, BdOP06, CSA07, CBR04, Cib08, ECAV07, ELJ06, FGMP08, JR08, LHD^{+}01, PAA07, SIHH04, Wak05, WKB07, WMS08b]. localization [Kut05, ZLLZ09]. localize [RZF03]. Localized [Kar03, NdGG06, SI09]. locally [BH08, Hie05, HM07, KA03b]. Location [Kut03, LJW02, CL07a, CL07b]. locations [DCL09, hZzGqX^{+}09]. loci [Cui07, CY09]. Locking [HRBL02]. Locomotion [Kel01, AN09, KCS^{+}06, KH09c, Nis06, Pen03, SB09b, SH08b, WS09]. Locomotor [CR02, Syl06, Sch08a]. Locus [YK03, KW00a, Wax09, Wil06b]. log [MR07b]. log-log [MR07b]. logarithmic [KE09, Pac09, SHI05, SHI06b]. Logic [HFH03, BCRG04, BS04, FLWB07, LM09, Yun05]. logic-mathematics [Yun05]. Logical [Ano01b, Iro09, ST01, BS05a, BCRG04, ST03, SF04, AJOK09]. logics [TK08]. Logistic [Nås01a, AWW06, Cam03, LB05, MM09]. LogitBoost [CFLC06]. Lognormal [QB00, Eng07]. Lolium [CEP05]. Long [CBBH01, DHM01, HFH03, Ish00, KTP09, KSG03, Ova02, RDH09, TH03a, UB01, AF09, BRC^{+}03, DI09, DMQ04, For09, GMMGDA05, GBGAKD05, Pie09a, QD09, Smi08d, vdBvdB09]. Long-bone [KSG03]. Long-distance [TH03a]. Long-range [Ish00, Pie09a]. long-run [QD09]. Long-Term [HFH03, Ova02, DHM01, UB01, AF09, BRC^{+}03, DI09, DMQ04, Smi08d, vdBvdB09]. Longevity [GG01a, JW01, LVV^{+}01, Man01, Pel00, AAK^{+}09, Tof06]. Longitudinal [KG02, TT03, AAK^{+}09, TT02]. look [Plu06, SDL^{+}08]. Looking [SLHN06, Hok00]. loop [BFG09, HJ05, RCS05, SWC^{+}08]. looping [Nar07]. Loops [KCA03, BNRW04, FI06, JM07, RGF^{+}08, SSF09, XK07]. loosely [OW07]. lopinavir [NLM^{+}08]. Lorentzian [Che07]. lose [IFN07]. lose-shift [IFN07]. loses [SBvS06]. Loss [WL03, Ark05, Est07, MKS^{+}09, RC09]. Losses [FB00, BZ04]. Lotka [BM09, CG03, MC06, PDC02]. Lotz [NDE06]. love [vVH07]. Low [CSD09, KKDV01, MC01, Sta00a, Sta00b, Sta00c, VRB01, CMCH08, EZM05, Hav04, JAA^{+}07, JAJA07, Kao06, KSEK09, Net09]. Low-affinity [VRB01]. Low-density [Sta00a, Sta00b, Sta00c]. low-quality [CMCH08]. Lower [HYA02, SL00, CSS06, Wes03]. lowering [MBD08]. LUCA [Di 06b]. luminal [HS07]. lung [HM06, MK04]. Lungs [WGH00, WGH01, BT06]. Luria [Fra03]. lymph [MK04]. Lymphocyte [Noe00, PW09b]. Lymphocytes [AB00]. lymphoma [AMP06]. lymphomas [MH07].

Lymphotropic [AB00]. Lynn [Bak07]. lysis [ARR08, CTS^{+}08, TB04]. lysis/lysogeny [TB04]. lysogens [EBM07]. lysogeny [TB04]. Lysozyme [Lan00a]. lytic [Kom07, PW09a]. LZ [LW08b].

M [How09]. M. [SJGK04]. M2 [DHW^{+}09]. Machine [GCYH01, CTZ^{+}06, Fox05, ZLLZ09, ZCLZ07]. machinery [Hau07, PBL06]. Machines [AGCLMM03, CLXC03, AS09, dCZJ^{+}04, DB05, DRLB05, JSA^{+}07, MBBR06, QLHL09, YCC^{+}06]. Macrohistorical [VG01]. Macromolecular [HM01, Vei03, MD04, MPOBD^{+}09]. macroparasite [RDF03]. macrophage [LCHB06, MKD^{+}05, RWCK08]. macrophages [ND04, OBL04, RK06]. macroscopic [GKXS07]. Magnaporthe [TGT06]. Magnetic [Rei02b, Wal08]. magnetite [Wal08]. magnetite-based [Wal08]. magnetoreceptor [Wal08]. Magnitude [MC01]. Mahalanobis [Lin08]. maintain [OI06, OI07]. maintained [EFW07]. maintaining [MNMH^{+}08]. Maintenance [BFR05, Bro02, Hay01, RKNK01, ASD08, BDBR07, FMH08, GEK04, LKK07, NKC^{+}08, OI00, SMPvdB08, TKP07]. maize [LSMB^{+}03]. Major [Wit03, JL06, Kun06]. Makarieva [BDMR06]. make [CA00, FI06, PG01b]. Makeham [Gol09]. Making [Di 01b, LP08, MAHD06, NDD08, RRH08, YCS04]. Malaria [GLK^{+}02, CSD09, GZK06, RBW09]. malate [LZS^{+}08]. malate-aspartate [LZS^{+}08]. Male [Dus00, FSL03, GMMGD^{+}08, GK00, NS03b, Blu07, Bog04, BBK04, GZ04, KV05, MT06a, MT06b, MS03b, RRS^{+}07, SWI07]. male-biased [Blu07]. males [BBK04, CMCH08, GOP09, HH04c, HH05b, YOYT07]. males-modelling [BBK04]. Malthusian [Mar04]. malting [OF05]. mammal [CLMMP06, MGC04]. mammal-dominated [MGC04]. Mammalian [Coo01, MW01, OF01, FP04b, Gra07, GI09, HA07, JM07, Jam09a, KF04, KOK06, LG04, NT04, Ros03, SYYI07, ZSZ^{+}06]. Mammals [BKF00, Cha03, HHR01, IST01, BTS08, BGF03, BG06, DBF07, KG03, KSG03, Osb08, SI04b]. Mammary [HCMF01, HCB^{+}02, SVS^{+}02, BMH07]. management [SD06a, SD06b, TH03b, Wag03]. managing [CC09b]. mangrove [LS08b]. manifesting [Dim05]. manner [HHP05]. manuscripts [SDBH04]. many [AK04, CY09, HMGK06, KSB08]. many-wrongs [HMGK06]. Map [öOW01, MBB08, Pec06, RA08, WW05b, FWCN05, RR09, SV07, SSB^{+}02]. MAP2 [HGV01]. MAPK [ZPP08]. maple [Che07]. Mapping [CS02, RC00, SPN06, CLS08, CY09, Dav09, FHL^{+}06, LDW09]. maps [Cyt04, JQR08]. March [Ano00j, Ano00-28, Ano01l, Ano01r, Ano02c, Ano02i, Ano03y, Ano03-31, Ano04-50, Ano04-41, Ano05-36, Ano05-41, Ano06-31, Ano07-33, Ano07-39, Ano08-31, Ano08-37, Ano09-30, Ano09-36]. marginal [LR07]. marginata [BIS^{+}07]. Marine [BKF00, Hay00, HHR01, MSI01, Sol01, MA03a]. Maritan [MGL05, MGL06a]. Maritime [DFSD05]. mark [Jam09a]. Markers [ACCC02]. market [SP07a]. Markov [BLF^{+}09, BL06a, GFW^{+}09, HS02b, IGHW07, JBP02, PB06, SCJJ08, SJ09, WT01]. Markovian [GKXS07].

Marks [LM01b]. markup [KM08]. Marrakech [Ano00-30]. Mass [KDK02, SV05, TPD^{+}00, BB07b, Can07, CM09, CTB09, GSB05, Hel08, HJ07, NAV04, PP04, SH08b, TP05, VTC08a, VTC08b]. mass-balanced [PP04]. Massaging [DSK02]. mast [BEK^{+}03]. Master [KMT06]. masting [LRHB09]. Matabolic [Ano04b]. matching [CCC08, Mie05]. Mate [Bjö02, BB02, HH04c, HH05b, LKK07, NP08, WA07, BFA08, CMCH08, SWI07, Wak05]. material [FFIS07, Hol06, Vin05]. materials [FH07]. Maternal [AD06a, GK00, WH03, Buc04, FMLP06, Gra07]. mates [WT05]. Mathematical [AC04, AMP06, AMW00, Ano00e, Ano00-30, Ano09a, AP09, BPV01, BB06a, BM03b, BRND09, CSFH^{+}01, CSM02, CMTU01, CS00, CMS^{+}00, EAC^{+}06, FW01, FMF^{+}00, FBM06, FM09, GC08, HHP05, Her00, HB02a, HS00c, IM02, KHE06, KSO06, Kra01c, Kra02, Kut03, LM09, LBS00, LWRK07, LKC07, LPB^{+}04, LK03, Mau02, MAC06, MGL^{+}06b, MH02, MA07b, MS01b, MM00b, NKL06, NB02, ØKRG04, OBL04, PAH06, PKW^{+}00, RSSM06, RKF06, SWS02, SCH02c, SMG^{+}03, SB07, SGM08, SVN^{+}05, SS05a, SRN^{+}00, SZLM09, SZ09, TK08, TGP^{+}00, TLC07, WRG^{+}04, WBB02, Wan00, WBD^{+}09, YFKP03, ABM04, AM08, Ark05, AB03, BBM03, BBM04, BRS^{+}09, CP07, Cog07, CM05a, CM05b, CPG09, CPMG^{+}08, DT07a, Daw09, DRV^{+}06, DC09, DBG06, ECAV07, EPJ^{+}09, EPJ^{+}11, FBUL05, GK06a, GB07, GvHP^{+}07]. mathematical [GPB^{+}04, GCB^{+}07, GPMW06, HA07, Hop06, ITI^{+}09, JR06, JWB^{+}09, KBT08, KNC^{+}04, KKR^{+}07, KLN^{+}09, KJJB07, LN05b, LH08, LH09, LS08a, LM07, LLC03, LFC04, Leo07, MPM07, MKS^{+}09, ML04, MSL04, NS06, NS09b, Nar06b, NM08, NLS08, NGN^{+}04, NL03a, NI07b, Pep04b, PSJ04, RNSU04, RD07, RRC^{+}06, RSC^{+}06, RG06, RF04b, SHPDL03, SC03, SS06b, SE07, SJD^{+}09, SKK^{+}07, SSD09, TKN07, TBR08, TYI^{+}06, VGS^{+}05, VdL03, VN07, VA06, WWS^{+}06, WTC09, ZFVH05, dT07b, MGAD09a, RWF01]. Mathematically [Koh07]. Mathematics [Ano00e, Ano01h, EKIL01, Len01, Yun05]. Mating [Dus00, RKNK01, TNTJ08, YYY^{+}08, AR05, CC06b, CT06, HK09, KW00b, Kel07, PDM04, RHF07, RC09, SWI07]. matrices [FP03, FS04b, GKNT09, GNH^{+}05, GL09, PGLL07]. Matrix [Di 01b, LGB02, YL00, Apa09, BdSFFDMC09, DqLmW07, GHA03, GKTN07, KP06, KK09, LH06, PWG09, VCGV07]. Matter [WS02b, BÅ03, DBG06, WKB07, vdBR04b]. Matthew [EK08]. Maturation [IM02, MH02, KD03, MN06]. Maturity [TTR00, LPT05]. maximal [LIK^{+}05]. maximising [LW08a]. Maximization [Ack04, SPF08]. Maximizing [Pel00]. Maximum [Eti09, FT09b, HLW00, Now00, Puj02, TN04, BGF03, BDR08]. Maxwell [Hop02]. May [Ano00g, Ano00m, Ano01m, Ano01s, Ano02b, Ano02h, Ano03v, Ano03-27, Ano04-36, Ano04-40, Ano05-37, Ano05-40, Ano06-32, Ano06-37, Ano07-34, Ano07-40, Ano08-30, Ano08-36, Ano09-27, Ano09-33, CSM02, BH03b, Buc04, CSFH^{+}01, CRB05, CSD09, HCC00, KS03, MR03, MS08b, NA04, SC05, Wak04, ZXWF08, dM09, vV06]. Maynard [Ano06-54, Gav06, Har06, SS06d]. Mdm2 [AJOK09]. mdr1 [IZGG05]. Mean [CRR08, Orr03, Raf02, SS00, vAR09, vAGDR09, BK05, GKK06, LS03, LS04a, OHM05, Ros05, Tan08, WKB07, WKRD09]. Mean-field [vAR09, vAGDR09, GKK06, OHM05, WKB07, WKRD09]. mean-variance [BK05]. meaning [DLB07]. Meaningful [Rae02]. Means [Cha01b, Laz03, GBK^{+}07a, HM05, NA02, RAK^{+}08]. measles [BCJ^{+}08, GXG03]. measurable [BSJ04]. Measure [LL01a, MWB^{+}02, AMS09, GNH^{+}05, KmMK04, LW07]. Measurement [HS03, RR08]. Measurements [Di 01b, Alb08, CH05b, FS04b, GAK^{+}06, RGB06, SW08c]. Measures [Bas01, Kee00, Rou03, CHCC^{+}04, FMHW08, PKS^{+}08, PPRI08, Ric03, RKH^{+}06, TLC07]. Measuring [BEK^{+}03, JCRJ07b, JCRJ07a, LBJE03, PSM06, DPV00, CC09a]. Mechanical [BGF03, BS00, Ish00, JMB00, KG02, MFB01, MV02a, MV02b, MV02c, Nir02, SHF02, TL00, AN09, AJ08, AM06, BGE05a, BGE06a, BGE06b, BB07b, CABB09, Cur04, DMW04, KNC^{+}04, KJD08, LCHK09, NSS08b]. Mechanics [DF00, ZM03, BB06b, DSZ09, DP08, LBQ^{+}05, LZ09b, NM06, Pen03, THL03]. Mechanism [CCGC02, HLH01, Jac03, MS01a, MD08, NN07, PDIS00, Ril00, TPWG01, Tho02, UHI02, VGCGM^{+}02, VP01, WKG03, Yos03, BBSLN08, BTA08, BL06b, BEBV03, BI09, Buc04, CT06, CFCGCC03, CB07a, CPG09, Dal06, Dru03, HKP07, Hua03, JK06, JLCS08, KTH09, KLK06, LK08b, LVABV05, MP09b, Nie06, OQGC07, RAZ03, RKH08, RSB09, SSJ09, SC03, SB05, SA05a, SDL^{+}08, SKY09, SVG^{+}08, TKP07, TMH04, USTG09, WB04, Xie09, Zhu09, dBCP^{+}07, dV06]. Mechanism-based [MD08]. mechanism-mathematical [SC03]. Mechanisms [BFGD07, CY08, FFN00, HS01, Hog00, LFLM00, RWR00, RMAI06b, SCNP^{+}06, SML02, The00, ACSY04, ACLH05, CPM^{+}09, CRNP07, CWJ07, Dic08, FM04, GGH^{+}05, GMdA^{+}07, GZ04, GZK06, INSR08, JB04, Ken07, KLL07a, LM07, LG04, MKA05, MWD08, NLS08, PCB07, RR08, SJGK04, TR08, WRG^{+}04, YLM03, FL09]. Mechanistic [GS02b, Han01, Han02, LM01b, dBWBP06, ASMM08, DGM05, GK04, JEDH08, SN04]. mechano [IvDHI08a, IvDHI08b]. mechano-regulatory [IvDHI08a, IvDHI08b]. Mechanochemical [SNT03, LCL07a, LCL07b, LL09]. Mechanokinetic [Paw07b, Paw09b]. mechanoregulation [NP05]. Medaka [TMI03b]. Media [KMH00, BGE05b, BFGB04]. Mediated [HA00, LPLC00, SCH02c, BGB08, CEP05, CRR08, GAA02, HC07b, Kra01c, KCHP08, LK08b, LLC03, LdGH09, RGF^{+}08, RCS05, SM06b, VWR07, WMS08b, Win06, Xie07]. medical [Gol08]. mediolaterally [BLS^{+}09]. Mediterranean [FLS^{+}04, SBI07]. Medium [AF02a, GSM06]. Meeting [Ano00e]. Meiosis [RNS04]. Meiosis-II [RNS04]. Melanin [Ano01e, Mac01, MBH03]. Melanocytes [Ano01e, Mac01]. melanogaster [AO03a, LdGH09]. Melanogenic [øPVO02]. Melanoma [RSD^{+}01]. melanophores [DTM07]. Melanosomes [Ano01e, Mac01]. mellifera [GHHR03]. Melolontha [HHH06]. member [SS06a]. Membrane [Byw09, CMPL^{+}00, DB05, FSG00, GB01, Hop02, TGP^{+}00, AW06, CRJC04, CC06a, IHV^{+}06, KP06, Kin04, KI04, KI05a, Lin08, MGP^{+}08, OUPG09, Paw07b, Paw09b, PGLL07, RZF03, RM04, RAK^{+}08, SYC06, WYXC05, WYC06, WO04, ZC09]. Membrane-spanning [Byw09]. Membranes [Wil02, CML08, Hul05]. Memoriam [Ano01d, Has01b]. Memories [MH08]. Memory [Ano06-54, CKJ^{+}02, DB01, KHHS09, KK00c, Nei01, UB01, WS02a, CC08, CE08, For09, HVY^{+}07, Kim07, Koh07, SI04b]. Memory-one [KK00c]. men [Bla04, Kan06]. Meningitis [SJ03]. meningococcal [SJ03]. menstrual [RD07, SHP09]. Menten [CT02, LFW02, TE07]. Meristems [LSMLB^{+}02, PKL02]. merolae [Di 08]. mesenchymal [LWRK07]. mesendoderm [MKL09]. mesenteric [KBT08]. mesoderm [GK09, MKL09, TSZ^{+}07]. mesoscale [CLW^{+}08]. mesoscopic [BRC^{+}03]. Meta [Nak01, HI05]. Meta-observer [Nak01]. meta-population [HI05]. Metabolic [Ace00, AøP03, BDMR06, BKF00, DFC^{+}02, GGH^{+}05, MGL03, MGL05, MVM^{+}00, MCWF01, NGMS08, PPEøP02, SP00, SLP00, SGG^{+}07, TRM03c, dBCP^{+}07, BWvK^{+}08, CCT^{+}09, DSA^{+}06, ELJ06, FP03, FLWB07, GABK08, GGK05, HTCS07, HJ07, HGH08, IS03, JK06, KOT07, KP08, LSAA^{+}06, LV08, LZS^{+}08, MvdO06, MSSS09, OCA08, PFRR08, PSPF07, PRP^{+}03, SPAH06, SPF08, Thi04, TRM03a, TRM03b, WMP03, Ano02a]. metabolic-pathway [KP08]. metabolically [Ped07]. Metabolism [BSWM00, CBH02, CSP01, DRW01, EP00, HCMF01, HCB^{+}02, PGLG01, AVSHV04, AIKP06, Cle07, DGD^{+}09, HCEK08, HCR^{+}04, Kal07, KSK08, LK08c, MS07a, MHMGM08, NM08, NHTM09, SdAC^{+}08, Smi09, TWO^{+}09, VHF06, VHF08]. metabolism-vesicle [MS07a]. metabolisms [MNMH^{+}08]. metabolite [DF09]. metabolites [RH08]. metabolizers [DFP^{+}08]. Metabotropic [LGB03a, LGB03b]. metacommunities [MGM07]. metacommunity [GS08b]. metal [YMLK04]. metal-reducing [YMLK04]. metalloproteinase [KP06]. metals [GVB^{+}08]. metamorphosis [TK05]. metaphoric [Pie09a]. Metapopulation [AP01, Bas01, EH00, GG02, OSBH02, Sch00, SFS^{+}01, YCC02, CV08, DDJ06, DPA05, EL09, JEDH08, KCHP08, MWL^{+}07, MJ07b, RHH08, XFAS06]. Metapopulations [EN02, CG06, CBR04, Fer09a]. metastability [LCL07b]. Metastases [Ano01g, Ken01, Ken02, HRZ06]. metastasis [DGD^{+}09, MI06, MNI06]. Metastatic [IKS00]. metazoan [FSS06b]. Methanediazonium [NTK02]. methanogen [Di 09]. Methicillin [ABP^{+}00, MPM07, PWVvR09, MWD08]. Methicillin-resistant [ABP^{+}00, MWD08]. Methionine [MVM^{+}00, PFRR08, RNSU04]. Method [MCWF01, NA02, PKL01, PKA02, SMG^{+}03, TGR^{+}00, Tur02, BGE06a, BGE06b, BMR08, CKE06, Che06b, HL05, HW09, HNTHA07, JAJH07, JCW^{+}03, KGB04, LMH04, LSS06, MR07b, NAS07, PGMK^{+}03, RLCIB05, SHI05, SHI06b, TSRB08, Wan06, WL09, WLFC08, ZYD^{+}05, ZT06]. methodology [MHRK08]. Methods [BS00, DOT02, KKK00, BCRG04, DMW08, DKS04, GHC03, IvDH^{+}08, STW^{+}09]. Methylation [NTK02, SLL06, Wil06a]. metric [RA08]. Metronomic [HFH03]. MHC [PD01, vdBR03, vdBR04b]. Mice [BKE03, BGG^{+}09, BKE04, LKM^{+}08, NKL06, RGSFM07]. Michaelis [CT02, LFW02, TE07]. Microarray [LY02, GLSW07, ZLXY08]. Microarrays [CS07, mLLS^{+}06]. microbe [SCH05b]. Microbial [VDV00, Bat06, Bat09, KRB05b, MRC09, Nar06a, PTFF05, SRS09, WLZ^{+}06, WVA05a, WVA05b, YS07]. microbiological [GST08]. microchimerism [BI09]. Microcirculation [SP02]. microcirculatory [BGE05a]. Microcolony [Joh08]. microcracks [TT04]. microdialysis [Che06b]. microdomains [HGP^{+}07]. microenvironment [ML07a, ML07b, YDL06b]. microglia [BFG09]. micromechanical [FHD09]. Micromechanics [FH07]. Micromechanics-based [FH07]. Microorganism [HP00b, Mar04]. microorganisms [Fer09b, KNS05]. Microparasites [BH01, AG06b, FLG^{+}09]. microplus [RAHO06]. Microprobe [BPZ^{+}01]. MicroRNA [WL09, PWZ09]. microRNAs [CRB05]. Microsatellite [LS03, VA06]. Microsatellites [ZXWF08, LS04a]. microscopic [PLG^{+}06]. microscopy [TR08]. Microspheres [TM00]. microstructural [FH07]. microtine [KJD05]. Microvascular [See00, MKB03, SKS09, TGV^{+}07]. microvessels [PDB08]. Microvillar [Lan00b]. Microwave [LFLM00, LMF03]. mid [Ari05]. mid-domain [Ari05]. Middle [GMMGDA05, NA02]. Middle-scale [NA02]. might [SPE08, vdBvdB09]. migrants [PB07]. Migrating [GA02, TM00]. Migration [Ale01, BM03b, ELB02, FF02, HS01, LAD03, Ale09, BR09, CLH07, CC09a, DDJ06, JWH08, KP06, Lin04, Lit07, MSdlPS09, NBT07, PB07, SMCT08, VW03, Xie07]. migrations [MPN07]. migratory [AMS09]. Milk [DSK02, HCMF01, HCB^{+}02]. Milking [SVS^{+}02]. mimic [JI05]. Mimicry [PDA^{+}00, FS07, KS06, LY03a, MH09b, RE04, She06]. mind [Yun05]. mineral [FHD09]. mineralized [FS04b]. Minimal [BOCF08, GSNH08, Jus08, Rou03, AA04, BGG^{+}09, BG07, MGDDH08, MS06b, OHM05, SZLM09]. Minimization [CDF00, Mar03, NGT05, Sel07]. Minimize [FB00]. Minimizing [HFH03, HH07, DMO^{+}07, TGN07]. Minimum [IST01, ISWT02, Orr03, Alb08, Ros05]. mining [Pán08]. minor [AAL08]. Minority [KY02]. miRNAs [CRB05]. missense [XSD^{+}05]. missing [ISW04, vV07]. Mistakes [HG02, LP08, RHF07]. mites [CM09]. Mitochondria [KK00b, Ban06, DL08, Kun05, Mit04, SKR06b, TK04]. Mitochondrial [CSC03, FK01a, FW01, NHTM09, SKR06a, SK01, BPLS06, FSS06b, OLBM08, OS03, PCH^{+}05, PAH06, UH09, dBCP^{+}07]. mitosis [RNS04]. Mitotic [TQUN07, TQUN08, KK00b]. Mixed [BS01, CA09, Cro00, KW00b, MSMKM06, RKNK01, VDV00, WGH01, YTAK01, dPGR06, Bea06, EL09, GZK06, Kel07, Lew05, Nar06a, RNP04, TAN09, Yan06]. Mixed-mating [KW00b]. Mixed-mode [MSMKM06]. mixed-species [GZK06]. mixed-substrate [Nar06a]. Mixed-symmetry [Cro00]. Mixed-Venous [WGH01]. mixing [KR07]. Mixotrophs [KDK02].

Mixture [Lin07, CY09]. Mixtures [Jav00, BLZ07a, BLZ07b, HW09, RNP04]. MLC [FB08]. MLC-kinase [FB08]. MLC-kinase/phosphatase [FB08]. MLP [SMPM09]. MLP-29 [SMPM09]. MMP [KP06, KP06]. mobile [MB06, RMF08]. mobility [SFVA09, VSA07]. Mode [MVM^{+}00, MSMKM06]. Model [AWAB05, ARW00, AGT^{+}01, Ano01c, AMV^{+}02, Ats01, AF02b, BHHS01, Bar01, BKP09, BE03, BCV00, Bro02, BSRH02, BSRH03, CCGC02, CLO^{+}02, CSC03, CMB02, CCP^{+}00, Cha00, CSFH^{+}01, CC01b, CC01c, CDHJ02, Chu00, CSON^{+}05, CMS^{+}00, Cum00, DF00, DALP03, EN03, EN02, FW01, FSG00, FLBB01, FC01, Fin06, FMF^{+}00, FP04b, FY00, Går00, GJE02, GBG01, GLV02, GAA02, GG01b, GRG02, GG02, GLK^{+}02, GCYH01, HBLG02, Han01, Han02, HCMF01, HCB^{+}02, Has01a, HGV01, HB02a, HB02b, HS00c, IM02, IV02, ISC01, Ish00, IKS00, Jac03, JG01, JMB00, JBP02, KTCD00, KF04, KKK00, KJD05, KW00c, KCT02, Kor07, KA02, Kra01c, Kra02, KGG08, KKDV01, Kut03, LM01a, LBS00, Laz03, LGE00, LFW02, LVV^{+}01, Mal02, MKLD02, Mau02, McN06, Mcn00, MN01, MSWH00, MS06a]. Model [MS08c, MAB00, MH02, MS01b, Moc02, MBB02, MM00b, Nås01a, NSM02, NB02, Nie02, NL03b, OI00, PMMS01, PGLG01, Pel00, Pin00, PG00, PDC02, PBR03, PV00, QA01, RMBM00, RB02a, RWF01, Ril00, RKL02, RVMR01, SKW00, Sch02a, SP02, Sil02, SMG^{+}03, SK01, Sta00b, SFS^{+}01, Tas02, TH00, TH03a, TPWG01, TM00, TMI03b, TPD^{+}00, TL00, TS02, UHK01, WM00, WT01, WBB02, WAMO00, WC02, WL02, WMLC02, Xie07, YLL00, YL00, YNL01, Zah00, Zhi02, ZM03, ZSS02, ZW07, AB07a, ASMM08, AN09, AM09, ABM03, ABM04, AG06a, AWO^{+}09, ARR08, AV05, AJ05, ABTR07, AKR09, AAK^{+}09, ABvdD^{+}08, Ark05, AA04, AKdlPP06, BM08, Bak04, BK05, Ban06, BM09, BLOL07, BLMV05a, BLMV05b, BGE05a, BGE05b, BGE06a, BGE06b, BST05, BTS08, BR04, Ber03, Ber07, BLS06]. model [BPLS06, BFR05, BRS^{+}09, BGF03, BG06, BGG^{+}09, BL06a, Bol06, Bon04a, BG07, BDR08, BTL08, BLRR08, BMS06, BOCF08, BB07b, CMG06, CC06b, CBB07, CDFP04, CA09, CLW^{+}08, CKS04, CCC08, Che07, CK08, CZM09, Cin03, CRR08, Cle07, Cog07, CM05a, CM05b, CLM07, CM03, CBF05, CPMG^{+}08, CSM05, CHN08, Cud05, CY09, CW08b, DW08a, Dam04, DBBW09, DBBW11, DB08, DMW04, DECEK06, Daw09, DRV^{+}06, DCC^{+}08, DS05, DC09, DBR^{+}05, DMPS05, DLM08, DSU^{+}04, DW05, DWBB04, DBG06, DFW^{+}07, DFCL08, DPA05, DBF07, EPR07, EZM05, ECAV07, Eti09, EPJ^{+}09, EPJ^{+}11, EL09, Ezo09, FWLN04, Fen03, FVP^{+}07, FI06, FDR04, FABdC04, FM09, FFHK09, FBUL05, FMH08, Fuk04, GS06, GMK06, GK06a, GVB^{+}08, GLE^{+}09, Gav06, GASA09, GGV^{+}08, GA07, GA08a, GA09, GvHP^{+}07, GW06]. model [GPB^{+}04, GCB^{+}07, GE05, GR05, GMM09, GBZ06, GvO04, GI09, GHHR03, GT08, GD08, GDC^{+}06, GSNH08, GC09, GT06b, HA07, Han04, HSC07, HLA09, HCR^{+}04, HHBY06, HRZ06, HB09, HH04a, HTCS07, HK07, Hie05, HE08b, HVN07, HMN09, HTN04a, HTN04b, HC07b, HBOS05, HMP04, IYGA08, Iim07, IB06, IvDHI08a, IvDHI08b, JDMZ^{+}07, JHJ^{+}09b, JG08, JBF^{+}03, JJEX09, JEDH08, JTGP06, JK04, KBT08, KP06, KNC^{+}04, KH09a, KSY^{+}08, KS06, KCV05, Kee05, KM08, KNWCB07, KD03, KKR^{+}07, KSK08, KLN^{+}09, KJJB07, KLL07b, KHE06, KT03, KZ05, KG06a, KH07a, KH09b, KH09c, LK06, LK08a, LN05b, LH08, LH09, LJ09b, LM07, LLC03, LFC04, Leo07, Lew05, LPJB^{+}08, LWC09, LACL03, LKC07, LDT09, LPB^{+}04, LFP^{+}05, Lit07, LVL08a, LVL08b, LYZ08, LJL08, Lo07, LCL07a, LCL07b, LL09, LSD^{+}07, LAG07]. model [LLO08, LM08, LCHB06, LRHB09, Mag04, MGL06a, MV06, MD06, MD03, MDD06, MLA04, Mas03, MH06b, MGS08, MZK08, MPM07, MKS^{+}09, MB05, MBB^{+}06, MGS09, MLWL06, MKL09, MA07b, MSMKM06, MD08, MSDM06, ML04, MGAD09a, MRF07, MWL^{+}07, MS06b, MGDM08, MMUGD09, MCN08, MWD08, Mur07, NMH07, NGTB06, NS06, NS09b, NBT07, Nar06b, NM08, NHTM09, NKC^{+}08, Net09, NGN^{+}04, NL03a, NT04, OOHI06, Orr06, Osb08, ØØS06, OHM05, Pal08, PBZ06, Paw07b, PBZ08, Paw09b, PCSL^{+}06, Pep04b, PNG03, PA09a, PR08, PFRR08, PV08, PWVvR09, PA09b, PA09c, PSJ04, PGN08, PAH06, PM08, PC09b, PMB06, PM03, PCK^{+}05, PB07, PHL08, QXDW06, RCA09, Rǎd09, RG05, RNSU04, RNP04, RD07, RRK06, RR09, RRC^{+}06, RP09a, RSC^{+}06, RRH08, RKOS09, RG06, RF04a]. model [RF04b, RF04c, RMAI06b, Rot04, RRR04, RHH08, RKRW08, RBS05, RDSB^{+}03, RCS05, RB08, Sac04a, SHPDL03, Sar04, SPG06, SI04a, SW08b, SWM09, SJGK04, SHP09, SB06, SE07, SVGK07, SB09c, SGD04, Sil09, Sim08, STMH04, SBGA04, SG04, SG06, SKK^{+}07, SMCT08, STSD09, SJ03, SSD09, SN04, ST05, SK09b, SS05b, SKS09, Tak06, TTT09, THKU07, wTA09, Tay06, TKN07, TBR08, TSRB08, TWO^{+}09, TKM06, Tlu07, TGT06, TBCD06, TGV^{+}07, TYI^{+}06, TBB^{+}06, TBC^{+}08, UI04a, UI04b, VGS^{+}05, VdL03, VAAH05, VN07, VGMVR^{+}06, VHF06, Ver04, VA06, VTC08a, VTC08b, VGDSU09, Voi03, VSP06, WE06, WOLS07, WWS^{+}06, Wal08, WC07a, WL07a, WHH07, WWS08, WTC09, Wax06, Wax09, WL03, WW05a, Wes03, WM04, Win06, WH07, XFAS06, XAP07, XM09, YHM^{+}06, YCS04, YB05, YKG^{+}05, YAL04]. model [ZFVH05, Zac09, ZAdlPLM07, ZAD07, ZZLT07, ZT08, ZW03, ZHB04, ZT06, ZAB^{+}09, dT07b, dBWBP06, WLFC08]. Model-driven [ZW07]. model-mediated [Win06]. Modeling [AD06a, AF02a, BCJ^{+}08, BSGT08, BMD^{+}08, BLZ07a, BLZ07b, BKCR01, BSWM00, BMS05, Bye09, Can07, CXZ^{+}09, CAF03, CH05a, CMTU01, CRNP07, CS02, DMO^{+}07, DLRP07, DTM07, FM07a, FI00, Gag00, Gar02, GT06a, GS02b, GP08a, GCS07, GF02, HAG03, HSG05, HMB^{+}08, HDRM00, Hen04, HKP07, JK06, JLS01, JEHK06, KLL07a, KS01c, KCA03, KG02, LGB02, LGK^{+}09, LGK^{+}12, LG04, LTG^{+}04, Len01, LKM06, LT08, MM02a, MMV^{+}04, MRC09, MCB07, MM00a, NAP04, OLBM08, OD03, PDM04, PGM00, RGF^{+}08, RR08, RGW^{+}05, RGFP07, RP09b, RW08, RGG00, SGTF07, Saf05, SRN07, SBPC05, SYSY02, SCC^{+}00, SM06b, Sta00a, Sta00b, Sta00c, SRN^{+}00, Thi04, TLA00, TMS00, VWR07, VDV00, WM00, WQ00, Wan00, WDH^{+}09, WMS08b, WR07b, WLHB07, ZMQX01, ZPP08, dQW07, AM08, BBKGP08, Bak07]. modeling [BBM03, BBM04, BB06a, BRND09, CLH07, CSR^{+}05, CTB^{+}05, CP07, CB05, CXM^{+}09, CTS^{+}08, CV08, CSD09, DT07a, DCP^{+}08, DHRM08, DSA^{+}06, FTG07, Fer09b, FFME08, FS04b, GKXS07, GST08, HK07, HIM09, HML04, HKC^{+}07, JHJ09a, KCP07, KCP09, KSEK09, KS09, KSO06, Koh07, Kom06, KJD08, KI04, LM09, Lin07, MLA04, MCK07, MGL^{+}06b, MCMS06, MSL04, NKL06, ØKRG04, RK05, RKF06, Sch08a, STK08, SN05b, SS05a, SRAL12, SML07, SML08, SZLM09, SZ09, WRG^{+}04, WKRD09, Yan09, ZSRB07, dPGR06, vAR09, vAGDR09]. Modelled [SBH01, KS08a]. Modelling [AB06, AJ08, BCKE^{+}08, Ben04a, BGMM08, BKE03, BKE04, BTCD07, BC02, CHdV06, CSM02, CB07b, CS00, DLGC02, DFP^{+}08, DEM^{+}00b, DF08, EKIL01, EBM07, FF02, FPM^{+}06, Ghi02, HMGK06, HCMF01, HCB^{+}02, HCM^{+}07, HM06, HGP^{+}07, Kal00, KRNG08, LPLC00, LBSP06, LMT05, LSMLB^{+}02, MBD08, MG09, PS03, PCS^{+}06, PB02, PBR01, QB06, RKH08, SK05, SWS02, SG07, SCH02c, Sin07, VW03, Wal07b, WGH01, WRNB04, XK07, ZTKH09, ADHM09, APS06, AB03, AIKP06, AP09, BKMH07, BBK04, CD05, CDB09, CCT^{+}09, DFSD05, EAC^{+}06, FBM06, GC08, HF03a, JST^{+}08, JWB^{+}09, KI05a, KRB05b, LGB03a, LGB03b, LWRK07, LSMB^{+}03, MCC^{+}09, MGP^{+}08, MSS05, MAC06, MHKS04, OBL04, PHdB09, PLG^{+}06, Rej07, RGRB04, SMH07, SGM08, SVN^{+}05, WRKK09, WBD^{+}09, XBCF05, YZGW06]. modellings [Cud05]. Models [ABP^{+}03, BP00, BLMV01, BG00b, BKF00, BB02, BM03b, CZM09, CBBH01, CPR00, CSP01, Dai02, DKD02, DM00, Dus02, EP00, Ell01, GS02b, Gro02, HS02b, HLW00, HR00, HLS01, JJ01, KS01a, KW00b, KMI02, LHD^{+}01, LM01b, MS05a, MS07b, MHKS03, MWB^{+}02, Mur00, OW07, ORM03a, OSBH02, Pep00, QB00, RSD^{+}01, RSBY03, RHH08, Rua01, Sch02b, SKN^{+}03, SFV02, SIMK02, Sne03, Sol01, Tyr01, VSCG00, VDV00, WT01, Wal00, WBH04, YCC02, YTAK01, YD09, Zah00, AC04, ACR06, AMP06, AFD^{+}06, AVSHV04, Ari05, AROS07, BDMR06, BG08, BEF^{+}06, BT06, BSS^{+}07, BdSFFDMC09, Bul06, BC04b, CSP^{+}08, CSP^{+}09, CAS05, CXM^{+}09, CE08, CF09, CSS08, CRL06, CTB09, CP03, CPG09, DRV^{+}08, Ded09, DBBC08, DHR^{+}07, DS04, DGS09, DSVBW07, FR06, FSS06a, FLWB07, FJBK05, GC08]. models [GK04, Gie03, GXG03, GT03, GKK06, GMY09, GG09, GEK04, GRBR04, GBK^{+}07a, HED06, Hol06, HC07a, Hur06, IiKY07, ITM^{+}07, Jus08, KGD09, KM08, KR07, KI09, KS08b, KCHP08, KI05b, Lap03, LS08a, LFSG^{+}05, LTI08, LLW09, Lor06, LP07c, MGL05, MX08, MSdlPS09, Mat06, MM09, MTS05, MRA06, MNL^{+}07, MJ07a, MGDBdM08, Nar06a, NLS08, OAKC08, PHG04, PSM06, PBB03, PB06, RMP08, RH09a, RH09b, RK07, RDF03, RPNH03, RBP^{+}09, SNCM09, SES08, SC03, SB09a, SDS04, SI06, SCH05b, SZLK^{+}05, SMPM09, SVS04, SHRR06, SF08b, SLHN06, SN07, SLH^{+}09, Smi08d, SX06, SH08b, Sta09, Ste04, TWE04, VCGV07, WMS08a, WTL08, WW07, YB07, ZRSK07, dPCP^{+}08, vV06]. models-measles [GXG03]. modes [BS05a, Fer09a, JTGP06, KP06, LP07b, MA03b, PAH06]. Modification [OF05, TG09b]. Modifications [PKA02]. Modified [MBB02, NA02, JDMZ^{+}07, Lin08, MS03b, MGL^{+}06b, RBW09, WVA05b, vLHH06]. Modifier [VGCGM^{+}02]. modifiers [SB09a]. Modifying [Gre00]. Modular [BWHK02, PSPF07, SPAH06]. modulate [WO04]. modulated [HdGH07, YBV^{+}00]. modulates [GP08b]. modulating [TLZ05]. Modulation [FK01a, Lan00b, MW01, NS07, SB05, vdBWLS07]. modulators [HG08]. Module [KPS02, Kri09, LPB^{+}04, ST03]. Modules [Phi02, FLS^{+}04]. mol [Bat09]. molar [SPF08]. mold [TTT09, TKN07]. Molecular [AH00, AGCLMM03, For07b, For00b, HB07, IV02, JB04, KW00a, Lin01, NLM^{+}08, PDA^{+}00, RB02c, RKL02, SSKL01, SHF02, Tur02, TN01, WZW^{+}07, Yan09, ZZW07, AHT^{+}07, AJ08, AMD05, AMD06, Bok06, Cal06, Fox05, FS04b, FWE06, GG07, HA07, HBB08, ITT09, KOT07, LN05a, LBQ^{+}05, LT08, LLO08, Mas03, OAC03, RNS04, SV03, Tay06, XWD^{+}01, ZRSK07, Zhu09]. Molecule [PD01, Di 04, Di 06b, Di 08]. Molecules [HS03, KY02, GE05, PPGS03, WSM05, vdBR03]. moment [Lio09, Mag04, MDL04, Sac05, SM06a, UW09]. moment-angle [Mag04]. Moments [Kee00, XWC08]. monitoring [YHI04]. Mono [KDK02]. Mono-species [KDK02]. monocytogenes [IGHW07]. Monod [DMPS05, GC09]. monoecious [MT06a, MT06b]. monogyny [FMH08]. Monohydric [CCGC02, Ril00]. monolayered [DSZ09]. monomer [Kli06]. monomeric [Sar04]. monomers [Cib08]. monophyletic [Di 06b]. monospecies [CCV08]. monospermic [MA03a]. Monotone [DBB09]. Monotonicity [GSG^{+}07]. Monte [ACK08, GFW^{+}09, HW05, HW06, SPAH06, SIHH04, WFGP04]. mood [Net09]. Morales [VGCGM^{+}02]. Morbidity [Wel00b]. Morocco [Ano00-30]. morphogen [Cin06]. Morphogenesis [GBG01, Hog00, RB09b, HM06, HIM09, Shi06a]. Morphogenetic [Lin01]. Morphogens [Cum01]. Morphological [AF09, KS01a, LST00, TTKZ01, NS07]. Morphology [CR01, MH02, CB07b, MRJR09, TTT09]. Morphs [ARKL02]. Mortality [Cha01b, Har01, HLW00, Jam01b, Wel00b, YHG^{+}02, Abr09, ACK09, BPC08, BLZ07a, BLZ07b, Bon06, GEF09, MM09, MI08a, MI08b, MJ07a, ZAdlPLM07]. Mortality-rate [HLW00]. Mosaic [Kor07, LN02, Moc02, TMI03b, She06]. mosaicism [Fra03]. mosquito [NBMS06]. mosquitoes [Mar09b, RBW09]. most [CMB^{+}01, IvDH^{+}08, Lac09]. moth [NLS08]. Mother [Wel03]. moths [Bye05, RRS^{+}07, TNTJ08]. motif [HJ05, KS08b]. motifs [BKKR08, HVY^{+}07, IS08, PCB07, YDL06a]. motile [DECEK06]. motility [CLM07, GA09, JA05, MGS08, UI04a, UI04b]. Motion [Arm01, IK02, UI02, YMC01, BLS^{+}09, GE05, GSNH08, LAG09, Wes03]. Motional [Mal00, NS07]. Motions [Fur02, GR05, RB09a]. Motor [Ats01, Wan06, Bak04, FWE06, GG07, LN05a, MFB01]. motor-cargo [FWE06]. motor-driven [FWE06]. Motors [Ats01, ITT09, LBQ^{+}05, Mas03]. Mountain [PKW^{+}00]. mouse [DZB^{+}04, GK09, LK06, MSMKM06]. mouth [TK09a, TK09b, VAAH05]. Movement [JEDH08, Kut03, Mal02, VMW02, Akt04, AV05, BZ04, BNRW04, CBC^{+}09, DO04, FL09, FKS07, HCM^{+}07, HWCF07, HC07b, LK08a, PS03, Pal08, PDM04, WRNB04]. Movements [MV02b, DBG06, FM07a, GNLEK07, JBF^{+}03, VAAH05]. moving [Pic06]. MR [RGB06]. mRNA [CS07, HS03, WAMO00]. mRNP [Kut05]. mRPE65 [GZG07]. MRSA [MWD08]. Mt. [AH03b]. much [AFZ08, PDM04]. mucosa [GT08]. mucosal [Agi04]. Müllerian [JI05, KS06, She06]. Multi [Bol06, BC02, CLH07, Gar02, HW01, LMT02, MK06, MMUGD09, Pin00, XWD^{+}01, Zah00, AB07a, BBK04, CKM05, CCZC08, CHR06, DMW04, EL05, FMHW08, GASA09, Jus08, Lit07, MBBR06, ML08a, ML08b, PCS^{+}06, PCZL05, PCZL06, Rǎd09, RA08, SWN07, SHRR06, SH08b, SH07]. Multi-aligned [XWD^{+}01]. multi-class [MBBR06]. multi-compartment [Lit07]. Multi-dimensional [LMT02, CHR06, RA08]. Multi-drug [Gar02]. multi-etiology [Rǎd09]. multi-factorial [DMW04]. multi-features [CCZC08]. Multi-level [HW01]. multi-male [BBK04]. multi-pedal [SH08b]. multi-person [EL05]. Multi-player [BC02]. multi-predators [PCZL05, PCZL06]. Multi-scale [CLH07]. multi-site [Jus08]. Multi-species [Bol06, ML08a, ML08b, PCS^{+}06]. Multi-stage [Pin00, FMHW08]. multi-staged [SHRR06]. Multi-state [MK06, Zah00]. multi-strain [AB07a, CKM05]. Multi-target [MMUGD09]. multi-trait-based [SWN07]. multi-trophic [SH07]. multi-type [GASA09]. Multicellular [LPLC00, MKLD02, AMD05, AMD06, EPJ^{+}09, EPJ^{+}11, Pal08, PA09b, PA09c, VHF06, YFK05, ZAD07]. multicellularity [MVS^{+}06, Ots08]. Multiclass [AS09]. multicomponent [FS07]. Multifractal [GBLV03, CML08, YAL04]. Multifractality [ASMD06, Zam03]. Multifractility [Zam01]. Multigame [Gag00]. Multigroup [HLS01]. multilevel [DCP^{+}08, FZ07]. multilocus [LFFT06]. multinomial [BE08]. Multiperson [Uit09]. Multiple [BR01, CS02, EP00, Ell01, ES01a, IKS00, OKTS02, PDC02, SI01, TH00, XWD^{+}01, YHI04, ATO^{+}09, BWvK^{+}08, CWDM06, CMW08, Eti09, GHHR03, Gro04, KDKS06, LVL08a, LVL08b, LZ09b, MPA^{+}08, PLB^{+}05, RDH09, RK06, RMP08, RA06, SWI07, SKR06b, SI04c, VCGV07, Wax09, ZSF^{+}07]. Multiple-Competitive [OKTS02]. multiple-strain [CWDM06]. Multiple-year [YHI04]. Multiplicative [Kee00, KE09, BDR08, HHP05]. multiplicity [EZM05, LG04]. multipotent [RDYH09]. Multiscale [CBB07, AD06b, BRS^{+}09, RSC^{+}06, ZAD07]. multisegmental [STSD09]. multiserotype [BSS^{+}07]. Multisite [CT02, CKS07, TG09b]. multispecies [BIB07, WLFC08]. multistable [MWS09]. Multistage [MLMW01]. multistationarity [KST07]. multistep [SBGA04]. multitaskers [DW08b]. multivalent [CPG09]. Multivariate [TKKA04]. Münch [DLGC02]. Muscle [CC01c, DSCD02, HFGB02, NB02, Nie02, Zah00, BFG08b, BFG08a, FKAC06, GD06, HTCS07, JTGP06, KBT08, Kok04, LN05a, LKM^{+}08, LW08a, LK08c, LAG07, Mag04, NMH07, Nos06, PC09b, RSB09, SdAC^{+}08, TW04, TSRB08, Wes03, dBWBP06]. Muscles [Cha03, JMB00, FB08, KH09c, LBQ^{+}05, Mag04]. muscular [BL08, DC09, Kok04, LME06]. must [RH04]. mutA [Dor03a]. Mutability [LT09]. mutagen [Tan07b]. mutagen-response [Tan07b]. mutagenesis [BKP09, GWPW04, SW09]. Mutant [MB00, MGP00, HM05, MSMKM06, WK05]. mutants [BMT04, Raf02, UH09]. Mutate [MS09b]. Mutation [ATO^{+}09, Cha01b, Kam03, Kel07, KSN03, MM00a, NL03b, SN03, SP05b, TAN09, WN05, WW00, WAL06, ANT09, Blu07, Bul08, Fuk04, Fuk05, JWH08, JCRJ07b, JCRJ07a, KT07, LJ09a, RBJ06, Whi05, XSD^{+}05]. Mutation-Accumulation [Cha01b]. Mutation-selection [ATO^{+}09, Kel07, KSN03, TAN09, Fuk04]. Mutational [GKM^{+}00, FAW06, GK06b]. Mutations [AH00, AS00, Kaw01, Pál01, SK01, Arc05, BM08, EL05, KMC^{+}07, LS03, Orr06, SR08, SN05b]. Mutator [AF01, Kom04]. Mutual [IS08, RGG00, Nar06b]. Mutualism [FDG02, FFIS07, MCC^{+}09, Sch05a, YHBW04]. mutualistic [BCP^{+}07, GMdA^{+}07, JBK04, MPD^{+}07, YNO09]. Mutualists [LBF01]. Mutually [KY02, NA03]. Mycelial [BJD^{+}02]. mycobacterial [PBMU^{+}09]. Mycobacteriophage [Kun00]. Mycobacterium [MK04, SKI^{+}06, RWCK08]. Mycorrhiza [PBT02, Smi09]. mycorrhizal [YNO09]. myelin [Dim05]. Myelogenous [AM01, CM05a, DKLL05, ML04]. myeloid [PTD09]. myocardial [LZS^{+}08]. Myocardium [CBBH01, KNC^{+}04]. myocyte [Hop06]. myocytes [JK06]. myofibril [STSD09]. myosin [LBQ^{+}05, Mas03]. Myrmeleon [Bur09]. Myxobacteria [Ost04, AV05]. myxomatosis [FMLP06, FGMP08].

Nagumo [ACK00]. naive [DdB03]. naming [DT06]. nano [HDNC04]. nano-signal [HDNC04]. Nanoarchaeum [Di 06a]. nanobiology [HB07]. Native [PGLG01, JAV07, MS08b, OQGC07, YHI07]. Natural [DML02, Di 00a, FR07, GGPFM08, GA08b, HW06, KB02, KK00c, MK01, MGDBdM08, PB02, SSR04, WS06, Wel00b, YL00, BZ05a, BZ05b, Cor05, DGZ07, FMLP06, Gab06, HW05, MHMGM08, Mil08, MPL06a, PBH04, Smi08a, Smi08b, Smi08c, Wit03]. Natural/random [MGDBdM08]. Nature [CDHJ02, KS01b, Pel00, Her09, KE09, Leo07, SZC^{+}03]. Navigation [Möl01, Möl02, MM00b, Wal00, Ben03, LRB^{+}06]. navigational [GRBR04, HMGK06]. NB [PCB07]. NB-ARC [PCB07]. Near [BW06, KW00a, KY03, SJ03]. Near-periodic [BW06]. nearest [DCL09, NGTB06]. nearest-neighbor [NGTB06]. nearly [Ait08, LJ09a, Ten08]. Necessary [AA04, CD02, KST07, KE09]. Necessity [MKT^{+}00]. Neck [MFB01]. nectar [GHHR03]. Needed [BS00, RHF07]. Negative [CD02, PGM00, KLK06, SA05a, TBCD06]. Neglected [Ghi02]. Neighbor [MGS09, DCL09, NGTB06]. Neighborhood [Kaw01, BGF07]. neighbourhood [IKD04, SSS08]. nematode [HCM^{+}07, RW08]. neocortex [Bon04a, Bon04b]. neonatal [Sch08a]. neoplasm [Izs05]. neoplastic [TMH04]. neovascularization [KM05]. Nephropathy [RBBH02]. Nerve [Tas02, VW00, Tas05]. Nervous [CMW02, Rec02, Sil09, SPE08]. Nest [KB02, PRT04]. Nested [GS02b, WT01]. nestedness [BCP^{+}07]. Nests [CB00]. Net [MSS05, MCN08, NS03b, Che06b, GNLEK07]. nets [AIKP06].

Netscan [CFM04]. Network [AOH03, AJ05, Ano01c, Bjö02, BS00, CS06, GL01, GP00, GG01b, HI00b, HDZ^{+}07, KR07, KPS02, MAB00, MPN^{+}05, MBB02, PPEøP02, PRP^{+}03, PD00, See00, AJOK09, ACK08, AGZ^{+}06, BMT04, BR06, BMR08, BTA08, DB08, DLF^{+}07, FB06, GABK08, GGK05, GP08a, GBK^{+}07b, HVY^{+}07, HBS07, HSMM07, HDW^{+}09, HSK^{+}08, IiKY07, IS08, INSR08, Ker04, Ker06, KSS07, LMVPM07, MSP03, ML09a, ML12, MGL06a, MGAR07, MSS05, Mat06, MPD^{+}07, MBRRI08, MGDBdM08, MSSS09, NGTB06, NTU06, PP04, PBMU^{+}09, PAVS03, RKRW08, San03, SVS04, TK05, TTT09, TO06, Tao04a, Tao04b, TKN07, TB04, TBC^{+}08, UL06, VGDSU09, VTG^{+}06, WBR08, WBR09, XT06, YYA09, YZHZ09]. Network-based [KR07, PRP^{+}03, ML09a, ML12]. network-QSAR [VGDSU09]. Networks [BWHK02, BHJ03, CMW02, GP01, HAG03, HW01, MY01, MCWF01, MWB^{+}02, Rae02, SCGFS00, Ait08, AHCN07, ABKR07, AFD^{+}06, AGW^{+}06, AGW^{+}08, AKLS05, BDMR06, BBPSV05, BCKE^{+}08, BBCQ04, BWvK^{+}08, BEF^{+}06, BLF^{+}09, BCRG04, BLS06, Bok06, BFP07, BB09, Bra05, BOvD08, BGD^{+}06, BCP^{+}07, CAS05, CFM04, CGK^{+}05, CABB09, Cud05, DTG09, DB08, ELJ06, FM07b, FFTS09, GAT07, GDS07, GMdA^{+}07, GSNH08, HGH08, IYGA08, IS03, JQR08, Kau04, KS09, iKIY09, KSN03, LS04b, LR07, LL07, Lie05, LMT05, MGL05, MH09a, MK06, Mas09, MNP06, Moc05, Moc08, MLPJ09, MS05b, MWL^{+}07, NAP04, NL04, NSS^{+}08a, NAS07, NGMS08, OBDJ06, OB04, PP04, PSSS03, Péc05, PWK03, PWG09, PHdB09, PSWK09, PB06, QD09, QB06, RKYH06, RAK^{+}08, RK07, RB09b, SGCK^{+}08, SZC^{+}03]. networks [SK05, SPAH06, SPF08, SVGK07, SR08, SSB06, SM06b, TY06, TG09a, TKKA04, TC09, TGV^{+}07, UM08, UW09, VFC^{+}09, VdGN^{+}05, WCA06, Wil08b, WS03b, ZRSK07, ZPP08]. Neural [Bjö02, BS00, Car02a, HAG03, RWK08, SP05a, BL08, BTS08, Bod08, Bok06, HDW^{+}09, JHJ^{+}09b, NGTB06, OSJ08, RAK^{+}08, Rue07, UL06, WR07b]. Neurites [FMF^{+}00]. neurobehavioral [MKS^{+}09]. neurobiological [LAG09]. Neuroblastoma [FMF^{+}00]. Neurochemical [MA07b]. neurodegeneration [CL05]. Neurofilament [FMF^{+}00, CBF05]. neurogenesis [SLH^{+}09]. neurohormone [KCV05]. neuromuscular [BFGD07]. neuron [LL05, LCL07a, LCL07b, LL09, Sev06]. Neuronal [CPC^{+}00, MM02a, MBB02, SM02, Fen03, FL03, MR07b, OUPG09, RKH^{+}06, WCLL08, WCH^{+}09]. Neurone [FTEG02]. Neurons [PB02, FB06, LFSG^{+}05, VL09, WKRD09]. Neuropathy [RBBH02]. Neurophysical [Rob03]. neurophysiological [RRR04]. Neurospora [BMT04, RVMR01, SG04]. neurotransmitter [GBK^{+}07b, Pán08]. neurotransmitters [Che06b, MS07b]. Neurotrophic [VW00]. Neutral [AOH03, JQR08, Ait08, AK09, CGK^{+}05, EAM07, Eti09, HE08a, LJ09a, MWL^{+}07, PC04]. neutrality [ALM09]. neutralization [WM04]. neutropenia [BBM03, BBM04, CM05b, FBM06]. Neutrophil [HDRM00, LCHK09]. Newton [Can07]. NF [HJ06, LPB^{+}04]. NF- [HJ06, LPB^{+}04]. NG108 [WCLL08, WCH^{+}09]. NG108-15 [WCLL08, WCH^{+}09]. Niche [KF03, BFK08, CB08a, GEF09, GD08, SMPvdB08, SM09b]. niche-model [GD08]. night [Ale09, Cal06]. Nigral [PB02]. Nile [MY09]. Nitric [CS00, SWS02, ND04]. Nitrogen [HHR01, PBT02]. Nitrogen-fixing [PBT02]. nitzschia [SWM09]. nk} [WW05a, AH08, Ait08]. No [LBSS02, IS09, LW04, KBT08]. no-defense [IS09]. nociception [MA07b]. Node [öWO01, MK04, RK05]. nodes [MLPJ09]. Nodules [PBT02]. Noise [BHJ03, DS00a, FSGB^{+}02, MA04, Nei01, TTKZ01, UHI02, BDR08, Cal06, DLRL08, HSMM07, Lei09, Lei10, LRHB09, MKA05, OM08, OB04, Ped05, Ped07, SHPDL03, SR08, Ste04, Tan08, Tao04a, Tao04b, TKKA04, VL09, XT06]. noise-induced [HSMM07, SHPDL03, Ste04]. Noise-reduction [MA04]. Noise-Resistant [UHI02]. Noisy [RK07, DTG09, DS04, LFSG^{+}05, LMT05, Tlu07]. nomadic [HMGK06]. Non [ARKL02, ANL01, BGP00, DVC^{+}04, Di 01a, DBHS00, EN02, Fra00a, LLCM01, LHD^{+}01, MJW00, PC04, Ref04, SBH01, SS00, TH03a, WT01, AFZ08, AP04, BCV^{+}08, Ben03, BR08, DMPP09, DBBC08, Di 06b, FHL^{+}06, GC08, GZK06, IS03, JAHKH09, KJD05, KMC^{+}07, KY03, LB06, LKC07, MM09, MGL^{+}06b, MC06, MF05, NP07, NP08, NA03, NP05, PD06, RPNH03, Saf09, Sar04, SWB06, SKK^{+}07, TIM06, VTL05, ZW03, ZYD^{+}05, SH03]. non-additive [NP08]. non-autonomous [MC06]. Non-banker [ARKL02]. Non-central [LLCM01]. Non-coding [DBHS00, AFZ08, KMC^{+}07, NA03, ZW03, ZYD^{+}05]. non-deformable [MF05]. non-diauxic [NP07]. non-disulfide [SWB06]. Non-electrolytes [SS00]. Non-enzymatic [ANL01]. non-enzymes [MGDM08]. Non-equilibria [EN02]. Non-exchanging [SBH01]. Non-excitable [BGP00]. non-homogeneous [Saf09]. non-infested [KY03]. non-lethal [BR08]. non-linear [JAHKH09, MM09]. Non-local [LHD^{+}01, GC08]. non-monophyletic [Di 06b]. Non-nested [WT01]. Non-neutral [PC04]. non-normal [VTL05]. non-optimal [NP05]. non-oriented [BCV^{+}08]. non-orthogonal [Ben03]. non-phosphorylated [TIM06]. non-photic [SKK^{+}07]. Non-photochemical [DVC^{+}04]. non-preferred [KJD05]. non-random [AP04, SH03]. non-redundant [DMPP09]. non-sexual [DBBC08]. Non-specific [Fra00a, MJW00, GZK06]. non-spherical [LB06]. non-stationarity [FHL^{+}06]. Non-stationary [Ref04]. non-steady [IS03]. Non-steady-state [TH03a]. non-synaptic [PD06]. non-typeable [LKC07]. non-uniqueness [MGL^{+}06b]. Non-universality [Di 01a]. non-vertebrate [Sar04]. non-viraemic [RPNH03]. Nonequilibrium [KSM02]. noninactivating [WCH^{+}09]. Nonlinear [ARW00, KRGH07, KJD08, ML07a, ML07b, MBB^{+}06, RK05, Vei03, YZGW06, Amz04, CB05, DMQ04, DSVBW07, Fox05, Mil05, QHF^{+}07, TSC04, WLFC08, dT07b]. nonsense [GW06]. Nonspecific [BM04]. noradrenaline [BFGB04, BFG05]. norm [vVH07]. Normal [AF01, SMBM00, ABM04, GT08, IZGG05, VTL05, vLBJK07]. Normality [Gin00a]. Normotensive [CCP^{+}00]. Norms [BR02, Gin03, HB01, CSP06, NL04, OI06, OI07, Sch09]. North [OBPH^{+}08]. nosocomial [UM08]. Not-self [CMB^{+}01, For01]. Notch [GP08a, RGSFM07, BGB08]. notch-mediated [BGB08]. Note [MGL03, LRT04, Roe08]. Notes [Jam09a, NDE06]. notion [MV07]. Notions [Rei02a]. Novel [JAJH07, KSR07, MPOBD^{+}09, MKN02, UKY^{+}09, BSGT08, FTG07, HQP08, HGH08, JHJ09a, TLZ^{+}08, YSW09, ZLLZ09]. November [Ano00f, Ano00l, Ano01i, Ano01o, Ano02f, Ano02l, Ano03x, Ano03-30, Ano04-52, Ano04-37, Ano05-34, Ano05-39, Ano06-33, Ano06-38, Ano07-32, Ano07-38, Ano08-29, Ano08-35, Ano09-28, Ano09-34]. novo [RGW^{+}05]. NP [NGMS08]. NP-hard [NGMS08]. NS [RGF^{+}08]. Nuclear [CSC03, SK01, Kut05, TR08]. Nucleation [Sil02]. Nucleic [DSY01a, DSY01b, AC04, Dal06, For07a, For07b, MPOBD^{+}09, WBSY06]. nucleosomal [Aba09]. nucleosome [LDW04, LDW05]. Nucleotide [HGB^{+}00, MPOBD^{+}09, NA02, OKS01, KKR^{+}07, MWCS04, Wal07b, WTL08]. nucleotides [AB07b, KOK06, LKW06, QWQ07, QWQ09]. nucleus [Bel06, KA03b]. null [BMT04, PSPF07]. null-space [PSPF07]. Number [AIK00, Cha03, EH00, HS03, ISWT02, Ken02, Kru02, Orr03, TTKZ01, YALT00, Alb08, CHCC^{+}04, GOP09, Moc05, NS05, NDE06, NI07b, Ros05, SH04, SH05, SA05a, Wal07b, Yan08, ZYD^{+}05]. numbers [HP04]. numeric [SHPDL03]. Numerical [CMS^{+}00, EBId09, KKK00, LJL08, WPE03, BGE06a, DqLmW07, FWE06, GBK^{+}07a, LTW06, MGAD09b, SM09a, WLFC08, dQW07]. nurses [KV05]. Nutrient [BJD^{+}02, DO00, Rua01, Arc07a, Gro04, JD09, KUK07, MFI09, RF04b]. Nutrient-free [DO00]. Nutrients [Gro02, Gro04, MW09]. nutritional [SZ04].

O [DNS00]. O157 [TBB^{+}06, TBC^{+}08]. oak [NAV04]. obesity [CGS05, WG03]. Object [Ano00a, wTA09]. object-oriented [wTA09]. Objects [Nak01, Nak03, Cud05]. Obligate [RH04]. Observations [SPH03, ISW04, Ort06, SC03, Voi03, Won05]. Observed [OKS01, Wal00, BDMR06, LRT04, MGL05, PE04]. observer [Nak01]. obtain [NM09, PP04]. obtained [FP03]. Occupancy [GF02, LMH04, PBvdG09]. Occurrence [SC03]. ocean [SZC^{+}03]. octamer [Aba09, LDW05]. October [Ano00v, Ano00y, Ano01x, Ano01-30, Ano02a, Ano02o, Ano02v, Ano03-41, Ano03-43, Ano04-44, Ano05-44, Ano05-50, Ano06-50, Ano06-53, Ano07-47, Ano07-50, Ano08-46, Ano08-51, Ano09-39, Ano09-43]. octopus [SN06]. ODE [PHdB09]. Odor [LdGH09, Tur02]. Odor-mediated [LdGH09]. odorant [OY03]. odour [Rue07]. Odours [Mau02]. off [ALM09, BHWB05, Ger09, HBWB08, KSB07, PG01a, SH04, SH05, SA05a]. offenses [Jes06]. offs [BH01, RGFP07]. Offspring [Jam00, Jam01a, MMTS02, Wel03, GI09, Hel08, Jam04, Jam07, Jam08c, SH04, SH05, SE07]. often [GMFS06, RHM^{+}08, SFVA09]. Oja [RCA09]. Old [HYA02, NM08]. older [Bla04]. olfactory [Bye05]. Oligochaeta [ACD04]. Oligomeric [CPC^{+}00]. Oligonucleotides [ANL01, FC01, KTP09]. Omnivory [Van06b]. Oncogenesis [RSD^{+}01]. oncolytic [ZW07]. onconetworks [QHF^{+}07]. One [ACLW03, BP08, LCL07a, LCL07b, LL09, WDH^{+}09, BZ04, ETH04, KK00c, OBN07, PCZL05, PCZL06, SX06, WTL08]. one- [WTL08]. One-dimensional [ACLW03, WDH^{+}09, BZ04]. One-hit [LCL07a, LCL07b, LL09]. one-predator [SX06]. one-prey [PCZL05, PCZL06]. One-third [BP08, OBN07]. ones [FM07b, LD09a]. only [AD07a, Di 06b, Smi08d]. Onsager [OP05]. onset [JEHK06, QHF^{+}07, SRN07]. Ontario [CFCGCC03]. Ontogenetic [CACC02]. Ontogeny [Kov02, DF08]. ONYX [ZW07]. ONYX-015 [ZW07]. Oogamy [Dus02]. operating [RNP04]. operational [TC09]. Operon [Che03b, LMH04, Nar07, NPN09, RGF^{+}08, SZ04]. operons [YMLK04]. ophthalmology [EH08]. Opioid [SPC02]. opportunities [Roe07]. opportunity [BKMH07]. optima [Ait08]. Optimal [Bjö02, CC01c, CF07, CL00, EH00, HF03b, ISW04, IS09, JITJ09, KF03, KW07, KH07b, KT07, LD09a, LS08a, LW08a, NS05, Noe00, Pel00, PAD00, PE00, RBW09, SI01, SI02, SHI03, WH00, YFH^{+}07, YYI03, YHI07, AK04, Arc09b, Bul08, Bur09, CP07, FYX^{+}09, FL09, GMY09, HQP^{+}09, HH04c, HH05b, Kin07, KV04, LH06, LP08, LHDvdM04, NS06, NP05, PLH05, Paw07a, PB06, SH04, SH05, SS06b, SE07, SD06a, SD06b, SI04c, SN05b, Ten08, Wak05, YOYT07, Ale09]. Optimality [Bul06, NGT05, Nei01, GNH^{+}05, KCS^{+}06]. optimisation [JAHKH09]. Optimising [RB09a]. Optimization [BR01, BS00, CEA07, Gra02, Gra06, HML04, HGB^{+}00, KSM02, MV02a, MV02b, MV02c, PSSY09, Sne03, ALM09, Apa09, HFY07, Hor08, LZ09b, RAHO06, SI04b, STMH04, VCBV^{+}06, YHI04]. Optimized [SKN^{+}03, APS08, AS09, ZCD05]. Optimizing [AC07b, CSW^{+}08, GCP04, MSS07, FT09a, ILDP04]. optimum [FP04a]. Optional [HDHS02]. Options [SI01, SI04c]. Oral [COS01]. orb [VCBV^{+}06]. orb-weaving [VCBV^{+}06]. Orbital [CS09, How09]. Orbits [WS02a]. Orchestrated [WL02]. Order [IK02, MC01, RPB03, TGR^{+}00, Ait08, AW06, Bog04, BP08, FY00, Hav04, KKR^{+}07, PBB03, RGPB08, SK09a, WLZ^{+}06, ZXWF08]. order-dependent [ZXWF08]. ordered [NS09a]. ordering [FmW08]. orders [Kun03, Kun05]. ordinary [SBGA04]. Organ [ABR02, Ken02]. organelle [Pic06]. Organelles [KK00b, Dru03]. organic [BÅ03, BSJ04, Lap03]. organism [JWH08]. Organisms [Cha01a, CMW02, Mcn00, MN01, MSI01, Pál01, PKL01, Arc05, NDD08, YNO09, YFK05]. Organization [BPZ^{+}01, KB02, NC02, PAVS03, CBC08, ELJ06, FMI05, Gab06, GS06, GJ07, KG06a, LRD04, MB09, Oud05, QO08, RMF08, Sac04a, Sel06, Smi08a, WH07]. Organizational [LSAA^{+}06]. Organized [ORM03a, ORM03b]. organizing [AC04, MCM^{+}09, TG09a]. oribatid [CM09]. Orientation [HP00b, Jam01a, CS09, How09, SM03]. Orientational [CB07a, IPY07]. Oriented [SLP00, BCV^{+}08, MHKS04, wTA09]. Origin [Arc00, DSY01a, DSY01b, Di 00a, Di 01b, Eld00, GS02c, GS00, GVK00, HA00, KY02, MZ02, NA02, Ste02a, WL02, WMLC02, Arc03, BZ05a, BZ05b, Bat09, Byw09, Cor05, CBC08, Dem06, DM07, Di 04, Di 06b, Di 08, Di 09, EPR07, EG07, LD09b, LZGL03, MH07, MVS^{+}06, Syl06, Tay06, Wil09]. Originator [MOBN09]. Origins [BFG08b, BFG08a, Laz02, TCP05, Fis06, SKR06a, SKR06b]. orphan [Pán08]. orthogonal [Ben03, LJTD05]. Orthogonalization [Krá01b]. orthogonalize [SPE08]. Oryza [Mei05]. Oscillation [KMI02, SA08b, TH05]. Oscillations [BBM03, BBM04, GBD00, GG03b, ORM03a, PHL08, Rua01, UHK01, ACR06, BG07, CLB05, DBR^{+}05, Fin06, GP08a, HJ06, Hen04, HS08, JK06, KR05, KG06b, KGG07, LN05b, LH08, LH09, LKM06, MS03a, MGP^{+}08, Ped07, PM03, RvMK^{+}05, SRR08, Ste04, SH07, TYI^{+}06, WHH07, WAL06, dM09, vAGDR09]. Oscillator [GRG02, JG01, Cin03, MM08b]. Oscillators [OCP^{+}00, PPR01, RGG00, UHI02, BGO08, FP04a, KA03b, WCA06]. Oscillatory [CKS07, KMH00, Mas08, GE05, HE08b, KGG08, LG04, SMCT08]. oscillophore [RvMK^{+}05]. osmolality [Kur08a]. osmometer [Won06]. osmotic [GP08b, MM03a]. Ossification [Cub00]. osteoblast [LTG^{+}04]. osteoclast [LTG^{+}04]. other [Cha01a, Dru03, Gab06, Jam06, vV06]. other-regarding [vV06]. Ottavio [LHDvdM04]. Out-of-Sequence [MM02b]. outbreak [GR08a, HKC^{+}07, MPN^{+}05, WR04]. Outbreaks [PL01, BBPSV05, CFCGCC03, MSDM06, RGZ09, Vaz07, Yan08]. outcome [Gol07]. outcomes [Bol06]. Outer [TGP^{+}00, Lin08]. Outgroup [OLS^{+}02]. Outgrowth [DO00]. Outlaws [De 02b]. outperform [RB09a]. Output [DFC^{+}02, BK05, CLHW07, KI05a, Tan07c, Wes03]. ovarian [KSO06]. over-compensation [Dam04]. Over-exploitation [De 02b]. over-exploited [SBI07]. Overarm [CC01c]. overexpression [GMFS06, KLK06]. overgrazing [Kon03]. Overgrowth [CSM05]. overlap [Kam03, RG05]. Overlapping [SSB^{+}02]. Overproduction [SK09a]. Ovide [Kim07]. oviposition [CW08b]. ovulation [TLZ05]. ovules [SK09a]. Ovum [Nir02]. own [MH05]. Oxidation [CCGC02, Ril00]. Oxidative [SK01, SB09c]. Oxide [CS00, SWS02, ND04]. oxidization [JNWJWB04]. oxidoreductases [ZF08]. Oxygen [BL00, GP00, GP01, SP02, WGH00, CBB07, DMPP09, GT08, JTGP06, ND04, PCSL^{+}06, San03, Sar04, TGV^{+}07, WB06]. oxygenation [KHE06, PC09b].

P2X7 [BFG09]. P40 [Kli06]. P450scc [SN05b, SPN06]. p53 [BG07, PHL08, AJOK09]. p53-Mdm2 [AJOK09]. pace [GINT09]. Pacemaker [INSR08, LGE00, SKK^{+}07]. pacemakers [KMGDG04]. Pacific [PB07]. Package [SF08b]. packaging [MPOBD^{+}09]. packing [ZSRB07]. Paddlefish [FSGB^{+}02]. pads [SPG06]. Pages [Ano00q, Ano00u, Ano00w, Ano00-27, Ano00s, Ano00x, Ano00p, Ano00r, Ano00k, Ano00-29, Ano00t, Ano00z, Ano00i, Ano00o, Ano00j, Ano00-28, Ano00g, Ano00m, Ano00f, Ano00l, Ano00v, Ano00y, Ano00h, Ano00n, Ano01v, Ano01-27, Ano01t, Ano01y, Ano01z, Ano01-31, Ano01u, Ano01-28, Ano01j, Ano01p, Ano01w, Ano01-29, Ano01n, Ano01-32, Ano01l, Ano01r, Ano01m, Ano01s, Ano01i, Ano01o, Ano01x, Ano01-30, Ano01k, Ano01q, Ano02p, Ano02t, Ano02q, Ano02w, Ano02r, Ano02x, Ano02s, Ano02y, Ano02d, Ano02j, Ano02n, Ano02u, Ano02e, Ano02k, Ano02c, Ano02i, Ano02b, Ano02h, Ano02f, Ano02l, Ano02o, Ano02v, Ano02g, Ano02m, Ano03-36, Ano03-40, Ano03-32, Ano03-37, Ano03-35, Ano03-39, Ano03-42, Ano03-41, Ano03-43, Ano03-34, Ano03-38, Ano03u, Ano03-28, Ano03-33, Ano03t]. Pages [Ano03z, Ano03y, Ano03-31, Ano03v, Ano03-27, Ano03x, Ano03-30, Ano03w, Ano03-29, Ano04-46, Ano04-48, Ano04-43, Ano04-47, Ano04-54, Ano04-50, Ano04-56, Ano04-49, Ano04-51, Ano04-52, Ano04-53, Ano04-55, Ano04-42, Ano04-34, Ano04-38, Ano04-45, Ano04-41, Ano04-36, Ano04-40, Ano04-37, Ano04-44, Ano04-35, Ano04-39, Ano05-46, Ano05-52, Ano05-48, Ano05-51, Ano05-43, Ano05-49, Ano05-55, Ano05-56, Ano05-47, Ano05-53, Ano05-45, Ano05-54, Ano05-35, Ano05-42, Ano05-36, Ano05-41, Ano05-37, Ano05-40, Ano05-34, Ano05-39, Ano05-44, Ano05-50, Ano05-33, Ano05-38, Ano06-43, Ano06-45, Ano06-46, Ano06-51, Ano06-42, Ano06-47, Ano06-48, Ano06-52, Ano06-35, Ano06-40, Ano06-44, Ano06-49, Ano06-34, Ano06-39, Ano06-31, Ano06-32, Ano06-37, Ano06-33, Ano06-38, Ano06-50, Ano06-53, Ano06-36, Ano06-41, Ano07-44, Ano07-53, Ano07-46, Ano07-52, Ano07-48, Ano07-54, Ano07-43, Ano07-49]. Pages [Ano07-31, Ano07-37, Ano07-45, Ano07-51, Ano07-35, Ano07-41, Ano07-33, Ano07-39, Ano07-34, Ano07-40, Ano07-32, Ano07-38, Ano07-47, Ano07-50, Ano07-36, Ano07-42, Ano08-43, Ano08-47, Ano08-45, Ano08-50, Ano08-41, Ano08-42, Ano08-44, Ano08-48, Ano08-33, Ano08-39, Ano08-49, Ano08-34, Ano08-40, Ano08-31, Ano08-37, Ano08-30, Ano08-36, Ano08-29, Ano08-35, Ano08-46, Ano08-51, Ano08-32, Ano08-38, Ano09-40, Ano09-47, Ano09-42, Ano09-48, Ano09-41, Ano09-46, Ano09-38, Ano09-44, Ano09z, Ano09-32, Ano09-45, Ano09-31, Ano09-37, Ano09-30, Ano09-36, Ano09-27, Ano09-33, Ano09-28, Ano09-34, Ano09-39, Ano09-43, Ano09-29, Ano09-35]. Pain [HAG03]. Pair [Ell01, XFBC07, OS03, PD04, ST03]. pair-rule [ST03]. pairing [Tak06]. pairs [HGH08, Tak06]. Pairwise [Mad00, TPN07, WM00, Ste09]. palliation [MG09]. palmate [Whi05]. PAM [Di 01b, KK09]. Panama [LS08b]. Pancreatic [AMV^{+}02, DS00a, JCK09, Ped05, WB04]. pandemic [CAHH06, HLA09, ITL09, JITJ09, MPA^{+}08, PLH05, RDC09]. Papain [WP01]. Paper [LL01b, Jam08a, Jam08c, KNS05, NM08]. papers [Gel07]. Papilio [OST09]. Parabolic [SS01]. paradigms [SB09a]. Paradox [Eic01, MS08b, MPN07, MN07b, NP08]. paradoxical [OCA08]. parageneses [MB09]. Parallel [BP01, JP00, VO00, Yos03, Kur07, Kur08b]. Parallels [VW00, Bul06]. Paramecium [OOHI06, UKY^{+}09]. Parameter [BGK09, BGE06b, KKK00, DRV^{+}08, LSS06, Lie05, LH06, Rǎd08, RDF03, TSRB08, WTL08]. parameterisation [TG09b]. Parameterization [GKK06]. Parameterizing [AF02b]. Parameters [GB02, Möl01, SMG^{+}03, Alb08, BDMP^{+}08, BB07b, CLMMP06, DCC^{+}08, ELSFB07, Eti09, HNTHA07, JAJH07, JAJA07, JBJ^{+}08, Lei09, Lei10, LTLM09, MLWL06, NGTB06, NAV04, PFRR08, PBB03, RLCIB05, RRR04, VTC08a, VL09]. parametric [DSU^{+}04, KmMK04, Ric03, VGBA06]. parapatric [TYW05]. Parasite [FMD01, GS02b, GLK^{+}02, HS00a, BPC08, FS04a, KCHP08, LBF01, MRF07, NO04, WKB07, Wil08a, ZSF^{+}07]. parasite-mediated [KCHP08]. Parasites [SI01, BB03, GRW03, HI04, JWB07, KSB07, KA03a, SI04c, Yam03]. Parasitic [Fra00a, Fra02]. Parasitism [FFIS07, HB09, KCHP08]. parasitize [STMH04]. Parasitoid [CVH03, KT01, LAD03, SCH02c, CBR04, KBD06, PCS^{+}06, SCH05b, SN07]. Parasitoid-host [CVH03]. Parasitoids [CSFH^{+}01]. Paraspirifer [SKY09]. paratenic [PBC09a]. parceling [CH07]. Parental [MSWH00, NWT09, SWI07, WTSN06, YM04, Yan05]. parents [Den08, Kan05, Kan07]. parity [Jam08c, ZXWF08]. Parkinson [HSG05]. Parkinsonian [vAR09, vAGDR09]. paromomycin [MPOBD^{+}09]. Parotid [FY00, GCB^{+}07]. parsimonious [HKP07]. parsimony [FT09b]. Part [BGE06a, BGE06b, SD06a, SD06b, Wal00, Dru03, MGAD09a, MGAD09b, MV02a, MV02b, MV02c]. parthenogenesis [SDS04]. Parthenogenetic [Pál01]. parthenogens [JE09]. Partial [GHHR03, FmW08, GRH^{+}07, HW09]. partially [CML08]. particle [MD04, McN06, RW08]. Particles [BL00, Fra00b, JL01, Mcn00, MN01, Sta00b]. Particularly [JE01]. Partitioning [LM01b, Nik02, TBS^{+}02, VN07]. partner [GD06]. partners [CdOWS04, GOP09, WT09b]. Partnership [Eam06, ES01b]. Parturient [ESGK02]. Pascoli [Ano02a]. Passage [KCT02, VTC08b]. Passerine [ELB02]. passive [KH07b, SKY09]. passively [NKL06]. Past [FB01]. Patch [AP01, CC01a, DBGM08, HGC03, YALT00, HC07b, OT09, PB06, XFAS06]. patch-to-patch [HC07b]. patches [HH04c, HH05b, OUPG09, WT09a, YB05]. patchiness [MS08a]. Patchy [CCV08, MPL06b]. Paternity [FSL03]. Path [For00a, MS01a, VTG^{+}06, Ben04b, MRA06, TKN07, Toy09]. Pathogen [NC02, RSBY03, Wil01, IGHW07, NdGG06, OS06, PLSGG05, PSWK09, RG05, RRS06, SI04b, VWR07, WK05, XBT09]. pathogenesis [Agi04, RDSB^{+}03]. Pathogenic [SKR06a, Chu08, Hav04, LZ09a]. Pathogenicity [AIK00, IMN05, SJ03]. Pathogens [SI01, BR08, CW08a, Kao06, O’K05, OBPH^{+}08, SI04c, SJ03, vdBvdB09]. pathologic [VA06]. Pathological [Jam01c]. Pathology [PDA^{+}00]. Pathophysiology [EM00]. paths [Lor06]. Pathway [DFC^{+}02, MVM^{+}00, PPEøP02, SP00, SLP00, AVSHV04, BFP07, BSJ04, DMPP09, GBK^{+}07b, Khr04, KP08, NS05, NS09b, RDYH09, SYYI07, SSK06, MCMS06]. Pathway-Oriented [SLP00]. Pathways [Cha00, Cum00, JP00, SLP00, TPD^{+}00, BMR08, GP08a, KLK06, Lin07, MS03a, Nar06b, RR09, RGSFM07, SA05b, SPB06, Thi04, WMP03]. Patients [GLK^{+}02, HPZ09, IMN05, KJJB07, PPD09]. Pattern [Car02b, Cum00, DL00, FM06, GBCC01, JG01, KA03a, Kra02, Leh02, Moc02, RWF01, SCGFS00, TMI03b, ACLW03, Agi04, AO03a, Ber03, CV08, DTM07, FI06, GBGAKD05, Hot03, ISW04, IK06, KLN^{+}09, KSPA^{+}08, LBS06, LRD04, LJL08, Mig06, MSMKM06, NTU06, NOT04, PDB08, PRT04, PC09b, RM04, RRR^{+}08, SSJ09, SC09, UI04a, UI04b, Usa06, WKRD09, YYY^{+}08, ZDC08, ZKH^{+}05]. Patterned [KvHMP09]. Patterning [Mal02, AD06a, WO04, ZLN07]. Patterns [BLMV01, Cha01b, CTB09, DVL^{+}00, MKLD02, MLMW01, MS02, MBRRI08, MBJ09, SSWF01, YALT00, AMD05, AMD06, AD06b, BBPSV05, BGF07, DHYHR09, DP08, DP04, DBF07, ELJ06, FH07, HQP08, Hui09, JB06, LBCL09, LP07b, MS08a, MWCS04, Moc08, MGT^{+}06, ML08a, ML08b, OF05, Ost04, PCS^{+}06, Pie09a, RRS^{+}07, SWC^{+}08, SI05, TSZ^{+}07, UL06, VGBA06, WBR08, WSS07, ZAD07, Zhu09, dM09]. pave [KJJ07]. pay [Arc09a]. Payoff [BR09, KGL09, TNP07, WT09b]. Payoff-biased [KGL09]. PCR [JK03, LS04a, PCK^{+}05]. PDB [YDL06b]. PDI [GAA02]. PDI-mediated [GAA02]. PDP1 [XK07]. Peak [HCC00, FLG^{+}07]. pectinolytic [SRN07]. pedal [SH08b]. Pedicel [FGH01]. pedigree [Lac09]. pedigrees [SH06, TS08]. Peg [SWS02]. penetration [EPJ^{+}09, EPJ^{+}11, VHF08]. penguin [DFW^{+}07, KSPA^{+}08]. Pentadactyly [Hay01]. People [HB01]. Peptide [HM05, RB02c, AW06, RN07, TLZ^{+}08, vdBR04b, RB02c]. peptide/MHC [vdBR04b]. Peptides [ANL01, KA02, Byw09, HDW^{+}09, JG08, vdGMG^{+}09]. Peptidoglycan [Koc00]. PER/TIM [RCS05, XK07]. PER/TIM-mediated [RCS05]. percent [Smi09]. Perception [HG02, HLP06, KI04, KI05a, Lew03, LB06]. perceptual [TM06]. Percolation [GPG07, VSCG00, KTH09]. perennial [FLS^{+}04]. Perfect [AD07a]. Performance [Dus01, MV02a, MV02b, MV02c, AN09, DGS09, GMMGD^{+}08, HNTHA07, KKPB09, MKS^{+}09, MKA05]. performances [PKS^{+}08]. performing [MHRK08]. Perfusion [ABR02, SKN^{+}03, WGH00, MBD08, VPGA07, WBD^{+}09]. Perimeter [VMW01]. Period [HRBL02, BMT04, FGMP08, GZL^{+}03]. Period-2 [HRBL02]. Periodic [ACR06, AM01, CR02, CBR04, Cyt04, Eld00, GC09, Hok00, JG01, LWC09, SI00, WS02a, BW06, CLHW07, LN05b, MC06, CM05a]. periodically [CD07, NAS07]. periodicities [HAC^{+}09, JWWS08]. periodicity [YY07]. Periods [RGG00, Yan08]. Peripheral [Her00, BTS08]. Peristaltic [DD02]. perivascular [SCNP^{+}06]. Permanence [CC01b, PCZL05, PCZL06]. Permuted [Di 08]. peroxide [GMFS06]. Peroxynitrite [Sta00a, Sta00b, Sta00c]. Persistence [Bon03, EH00, Kee00, LAD03, Ova02, Sch06, SL00, YY06, BBD06, CBR04, CBC^{+}09, DMH08, EL09, FGMP08, HDF04, HM05, JG06, JWB07, MJ07b, OW07, RP09b, RPNH03]. Persistent [SPC02, OSK^{+}05]. persister [Cog06, Cog07, LLO08]. person [EL05, MP09b, SPS09, SA07c, SA08b]. personal [Kim07]. Perspective [Dem02, ESGK02, FPS01, JST^{+}02, SLP00, YHG^{+}02, GBD06, GMMGDA05, HCEK08, LAG09, MGL06a, MB07, QHF^{+}07, SS09b, SH06, TS08, VTL05, Wal07a]. Perspectives [EKIL01, KG00, GWM04, Win06]. Perturbation [FC01, RKYH06, SMBM00, AKLS05]. perturbations [Bra05, DT07a, FP04a, GLE^{+}09, Sch06, YLM03]. Perturbed [SML02]. pertussis [KMW03]. Pest [BB06a, DBG01, HB02a, HB02b, MS08b, WLHB07]. Pesticide [PG01b, KNT^{+}09]. pests [LSD^{+}07, MGL^{+}06b]. Peter [YM04]. petri [AIKP06, MSS05, MCN08]. pH [BPZ^{+}01, Di 05, LCHB06, Sar04]. phage [AHT^{+}07, Bul06, KT07, WBH04]. Phantom [Ped07]. pharaoh [KB04]. pharmacokinetic [GMK06, LPJB^{+}08]. pharmacokinetics [DFP^{+}08, DP04, MD08]. pharmacologically [BRS^{+}09]. Phase [Coo01, HRBL02, KMH00, AB07a, BGE05a, BGE05b, BBSLN06, BBSLN08, CLW^{+}08, CB07a, DC09, KMGDG04, PLG^{+}06, PGF^{+}08, Rel04, SS06a, WVA05b, YS07, ZW03]. phase-resetting [KMGDG04]. PHD [DMPP09]. Phenolic [CCGC02, Ril00]. Phenological [VA04b]. phenology [OBPH^{+}08, YFH^{+}07]. Phenomena [BPZ^{+}01, Thi04]. Phenomenological [TRM03a, TRM03b, TRM03c, MGS08, ZHB04]. Phenotype [MDD06, Wel03, AMD05, AMD06, Bul08, GA08a, GPG07, IvDHI08a, IvDHI08b, KF04, Pec06]. phenotypes [NP05, TCP05, ZAD07]. Phenotypic [BN04, MS06b, KF06, NI04c, TY09]. Phenotypical [LK03]. pheromone [Bye09, PDM04]. pheromones [RRH08]. Phloem [FGH01, TH03a, HMN09, LN03, Tho05]. phosphatase [FB08]. phosphatases [HSL04]. phosphate [CCT^{+}09, Gre05, PTFF05]. phosphoinositide [SN04]. phosphorus [HK05, HML04, SZLK^{+}05]. Phosphorylase [GBD00, FKAC06]. phosphorylated [TIM06]. Phosphorylation [Ano01c, CT02, GG01b, HGV01, CKS07, FMHW08, NS05]. Phosphorylation-dephosphorylation [Ano01c, GG01b]. photic [SKK^{+}07]. photoautotrophic [MKE^{+}09]. photochemical [DVC^{+}04]. Photodamage [Han02]. Photoinhibition [Han02, WAMO00, LIK^{+}05]. Photolyase [MS01a]. photoresponses [EBId09]. Photosynthesis [FI00, Han01]. Photosynthesis-Irradiance [Han01]. Photosynthetic [PDIS00, DFCL08, Gie03, HFY07]. Photosystem [Han02, Laz03, LIK^{+}05, LJ09b]. photosystems [BRC^{+}03]. Phototropism [RWF01, WRF01]. phreatophyte [RDL07]. phreatophyte-aquifer [RDL07]. Phycomyces [RWF01, WRF01]. phyla [Kun06]. phyllotactic [Hot03, JB06]. Phyllotaxis [NSS08b, Rei02a, HED06, SSJ09, SN05a]. Phylogenetic [CPR00, HL00, PKA02, SW03, SSM09, BL06a, DKS04, FPS08, FS09, Mic07, MSS07, RHM^{+}08, Wil08b, YMLK04]. Phylogenetics [OLS^{+}02, SDBH04, SCJJ08]. phylogenies [Bok06, CdOWS04, Ste09]. Phylogeny [Kun01, Mad00, FSS06b, OS03, PCH^{+}05]. Physarum [MSS05, TTT09]. Physical [DSY01a, DSY01b, KMGV00, LZGL03, Wil09, BL06b, CTB09, NAV04, OOHI06, SWRH03, TPP^{+}04, wTA09]. Physico [Leh00]. Physico-chemical [Leh00]. physicochemical [YTL08]. Physics [HDW^{+}09, RC00, BC09a]. Physiological [Han01, SM01, GGH^{+}05, IZGG05]. physiologically [LLW09, PR08, VGS^{+}05]. physiologically-based [VGS^{+}05]. Physiology [Ano01h, Ghi02, SKW00, GT08, Pen03, PA09a, RK05]. phytochelatin [MCMS06]. Phytophagous [WC02]. Phytoplankton [FF02, YN02, EBId09, HK07, HBOS05, MFI09, MW09, SCS04b]. phytotoxicity [Sin07]. piecewise [BLF^{+}09]. piecewise-deterministic [BLF^{+}09]. pig [LTLM09, SSD09]. Pigeon [Wal00, GRBR04, LRB^{+}06, RGRB04]. Pigeons [Rei02b]. Piglets [Jam01d]. Pigs [DSK02, AAEW09, SZLK^{+}05]. pilot [Izs05]. pinaster [DFSD05]. Pine [PKW^{+}00, DFSD05]. Pinus [DFSD05]. Pinwheel [SE02a]. pistachio [LRHB09]. pit [Bur09]. Pitfalls [Sch08b]. Pituitary [SRMW00, Leo07]. pituitary-thyroid [Leo07]. PKAc [Yan09]. Place [RHG00, AD07b, Di 05]. placement [Kor09]. placenta [KS03]. Placing [VO00]. Planar [CMPL^{+}00, BB07b, Zhu09]. Plane [Yos03, KH09c, SB09b, SH08b, WS09]. Planet [DL00]. Planetary [LL00]. planforms [SN05a]. Plankton [CSM02, LP01, LB06, Rua01, CCV08, Lew05, MK09, MPN07, WMS08a]. Planktonic [CSM02, Gro02, LP00, Lew03, SC03]. Plant [Ant02, CSFH^{+}01, CR01, FT09a, JL00, KK00a, LBF01, Laz03, LLF02, LSMLB^{+}02, PA09a, Smi09, VN07, WC02, ZM03, Bio08, BRC^{+}03, CEP05, CXM^{+}09, CB07b, CCT^{+}09, CCF06, CW08b, DL08, FMI05, FLS^{+}04, HML04, JMvdB09, JJEX09, JD09, LMVPM07, LSD^{+}07, MGL^{+}06b, Mig06, OS06, Pie09b, PGM00, RF04a, RF04b, RF04c, USTG09, WL03, ZJ06, vdBvdB09]. Plant-insect [LBF01]. Plant-mycorrhiza [Smi09]. plant-nutrient [JD09]. plant-pathogen [OS06]. Plants [KRNKH00, Kor07, LY01, MGL03, Nik02, ZMQX01, BH03b, GvHP^{+}07, IS09, KJD05, KY03, KYS06, KRN03, LY03b, LY03a, MV06, MMV^{+}04, MCMS06, Mil05, PA09a, PGH^{+}04, SS03, Sat04, SN05a, YFH^{+}07, YNO09]. Plaque [For02, AC07b, ZHB04]. Plaques [EKS02]. Plasma [SCC^{+}00, ARR08, MGP^{+}08, PBvdG09, Sta00a]. plasma-protein [PBvdG09]. plasmalemma [Paw07b, Paw09b]. Plasmid [GB02]. Plasmid-Bearing [GB02]. plasmodium [TTT09, DEM^{+}00b, FVP^{+}07, MB05]. Plasticity [SMD00, BN04, SGTF07, TK05]. plateau [Hop06]. Platelet [GF02, SMBM00]. platform [DGM05]. Platonic [DML02, BM03a]. Plausibility [MWB^{+}02]. Plausible [PDIS00, Dus06, HM05]. played [Mar09a]. player [BHC06, BC02]. Players [Ste00, Mas08, Mie05]. playing [NS03a]. pleckstrin [LGB03b]. Pleiotropy [TH00]. Plethodontidae [DOT02]. plethysms [SCJJ08]. pleuropneumonia [EL09]. plots [BHWB05]. PNN [LDXW06]. PNN-curve [LDXW06]. Pogona [See00]. point [BB06b, Che06a, NT04]. points [YD09]. Poisson [CY09]. Polarity [LFF06, AO03a]. Polarization [AAL08, KW00c, Aba09, BH04, FKS07, HH04b, HHH06, Nar06b, SN04, Zhu09]. polarization-induced [HH04b]. polarization-related [HH04b]. polarization-sensitive [HH04b]. polarized [Daw09, HHH06, Zhu09]. Polarotactic [HP00b]. Policy [Ano00b, ITKL08, TBR08]. polio [BMS05]. Pollen [SI00, HH06c, KGG08, LSMB^{+}03, USTG09]. pollinated [MT06a, MT06b, SS03]. Pollinating [FDG02, LBF01]. Pollination [OBDJ06, LSMB^{+}03, MT06a, MT06b, USTG09]. pollinator [HC07b]. pollinator-mediated [HC07b]. pollinators [EW09]. Poly [BG00b, HK05, HK05]. Polyanionic [Tas02, Tas05]. polycephalum [MSS05, TTT09]. Polyethism [BHOR01, GTDA02, Tof06]. Polygonal [SN05a]. polygyny [FS06]. polyhedra [HTN04a, HTN04b]. polyhedral [BC04b]. Polymer [MJW00, AJ08, AKR09]. polymerase [AC04, LS03, Lal06, WDH^{+}09, Xie09]. Polymerases [AF01]. polymerization [MO07]. polymerized [CML08]. polymers [CPG09]. polymorphic [BLRR08, FP07, MvdO06]. Polymorphism [For01, KW00a, RKNK01, BFR05, KOK06, NDD08, Wak04]. polymorphisms [TH05]. Polyomaviridae [KTE08]. Polyp [MHKS04]. polypeptide [LZ09a]. polysomic [TSS06]. polyspermic [MA03a]. pomonella [TNTJ08]. Ponerine [TBS^{+}02]. pool [Sat04]. pools [SBM08]. Poor [öWO01, DFP^{+}08]. Population [AMW00, Ano00-30, Ano09a, BIB07, BR02, Dai02, DMZ00, DHM01, FR02, FPS01, FB01, GVK00, HIN00, HI00a, HS00a, Hay00, IMKN05, JJ01, KKK00, KJ08, KCHP08, MP05, MW00, Möh00, NSM02, PL01, PDC02, PV00, SBI07, SGS^{+}05, SDS04, SE02b, TH00, UH09, VW00, VG01, YALT00, YYI03, vKPdR07, AKS07, AV05, BG08, BdSFFDMC09, BDR08, BEK^{+}03, Cam03, CC09a, CW07, DS04, ETH04, FM07a, Fuk04, Fuk05, GK04, GT06b, HI05, Her09, Hie05, HM07, HH05a, HP04, ING04, Kom04, LL08, LFC04, Lew05, LSH06, MAL03, Man06, MS08b, MW07, MK09, MDL04, NK08, NWT09, NLS08, PB09, RG05, RBP^{+}09, SSB^{+}07, SHRR06, Sim08, TBR08, VN07, WMS08a, WKB07, WW05a, WLHB07, XBT09, YHI03, YHI04, ZZLT07, dSKL09]. Populations [Ant02, CMTU01, CL00, DBG01, ES00, Har01, HLW00, KW00b, KK00a, KN03, LSMLB^{+}02, Rev00, RKNK01, TL02, VDV00, YBV^{+}00, Abr09, AROS07, BC09b, BP08, BLRR08, CEP05, CLS08, DBBW09, DBBW11, DLM08, ELSFB07, FP07, FS06, FAW06, FMLP06, FGMP08, FLG^{+}09, FPL03, GAK^{+}06, GGPFM08, GA08b, HHBY06, HA09, HC07a, KB04, KSB07, Kao06, Kel07, LT09, Lio09, Mar04, MSK03, Nei04, PS03, Paw07a, PGMK^{+}03, PM09, RMF08, RB06, Sav04, Sch05a, Sch06, SLB07, SSL08, SM09b, TOA^{+}09, TDW07, TYW05, TH03b, Vaz07, WT07, XBGN05, ZJ06]. pore [Kut05]. Porosity [ZFVH05]. Porous [AF02a, BGE05b, LWRK07]. posed [WR07a]. Position [BZ00, LN02, Rei02b, GABK08, LFP^{+}05, PGLL07]. position-specific [PGLL07]. Positional [GRBR04, HGB^{+}00, RGRB04, CLS08, FK03c, RB09b]. Positioning [MV02c, Yos03, Dru03]. Positions [JSV02, DRLB05, Tak06, Wan06].

Positive [CD02, KGD09, KCA03, FM06, Kon06, Kun03, SSF09, TBCD06]. positron [BKM09, MKB03]. Possibility [Hop02, MKT^{+}00, TMI03a, PTFF05]. Possible [EM00, Jam06, KK00b, SKR06b, SSWF01, The00, TU00, UI02, CdOWS04, DM07, Gre05, JB08, San03, SC03, Tak06, Xie09]. possibly [Dru03, LD09a]. Post [DKLL05, KK00b, BGMM08, GRH^{+}07, ØKRG04, SI04b, TG09b]. post-exposure [BGMM08, GRH^{+}07]. post-infection [SI04b]. Post-mitotic [KK00b]. post-translational [TG09b]. post-transplant [ØKRG04]. Post-transplantation [DKLL05]. postantibiotic [GASA09]. Postnatal [BKCR01]. posttranslational [LM07]. Potassium [VLFN00, FDA^{+}09, PD06, WCLL08]. Potential [FSG00, FW00, FFN00, HFGB02, HSB01, MPL06a, SGG^{+}07, AAK^{+}09, FLS^{+}04, Hop06, Hur06, Jäc07, Jam09b, Ken07, LKC07, LRT04, LGS^{+}09, MSP03, SV07, SPN06, TLZ05, VRAF06, WSM05, Wan06, WB04]. potentially [HGH08]. Potentials [BFH^{+}01, HGB^{+}00, BOCF08, WCLL08]. potentiate [BN04]. potentiates [Tak06]. potentiation [BFG09]. Potts [TS02]. Power [Niw03, Niw05, SP01, Tor01, CKE06, GDC^{+}06, MV06, PPD09, Wes03]. Power-law [Niw03, Niw05, PPD09]. powerlifting [GMMGD^{+}08]. PP1c [Yan09]. PPAR [BGG^{+}09, LKM^{+}08]. PQN [mLLS^{+}06]. practice [PA09a]. Prairie [BR00]. Prairies [GS02a]. Pre [DML02, BGMM08, FSS06a, GFWT04]. pre-association [GFWT04]. pre-biotic [FSS06a]. Pre-Darwinian [DML02]. pre-exposure [BGMM08]. prebiotic [MHMGM08, RN07, SF08b, vdGMG^{+}09]. Preceding [LL01b]. precise [AWAB05, CM09]. precision [HMB^{+}08, MD03]. Predation [CLZZ02, Gro04, Har02, LP01, Rev02, WH01, Wil01, AKS07, BR04, HF03b, KHL07, LB06, ZKH^{+}05, vKdRP05]. Predator [Bon03, CLZZ02, DKD02, DHM01, Fur02, Mal00, SFS^{+}01, Abr09, AKdlPP06, AMC^{+}09, BRC07, BGF07, DI09, Gar09, GEF04, HS08, Jes06, KG03, KV04, LPC06, LB06, Lin04, MDCC06, MC06, MPL06b, RE04, RB08, SX06, WTSN06, Wil08a, Zhd03]. predator-host-parasite [Wil08a]. Predator-prey [Bon03, AKdlPP06, AMC^{+}09, BRC07, BGF07, GEF04, HS08, KV04, MC06, MPL06b]. Predators [Går00, Har02, Ama06, AKdlPP06, BBD06, FL09, KYS06, Lew03, PCZL05, PCZL06, WTSN06, WNT08, WT09a]. predatory [Bur09, PM07]. predecessor [Pat05]. Predicate [MK01]. Predict [BKF00, HCMF01, HCB^{+}02, ZSZ01, BE08, BB07b, CRJC04, CFLC06, CTZ^{+}06, JS07, KS03, NGTB06, WYC06]. Predicted [BR01, Cha01b, LRT04]. Predicting [APS08, CZC05, CC06a, CCZC08, DCL09, ES00, KW00b, MPM07, MNP06, MM02b, OM08, PWZ09, PM08, STW^{+}09, SC09, TY09, TM00, TL00, XWC08, YCC^{+}06, ZF08, DBBC08, DLM08, Han04, HDW^{+}09, Lin08, MPN^{+}05, SYC06, VSP06, WYXC05, XSD^{+}05, hZzGqX^{+}09]. Prediction [CLXC03, CL07a, CL07b, DL08, DBF07, FSG00, Jam01d, LL01a, LSMZ08, PGLL07, RAK^{+}08, SNT03, VRAF06, YPY^{+}09, YY07, ZDC08, FH07, GZG06, JSA^{+}07, JHJ^{+}09b, LMH04, LL06, LLS^{+}09, PBMU^{+}09, VS08, VGDSU09, WL09, YMLK04, ZLLZ09, ZCLZ07, MPOBD^{+}09]. Predictions [Koh07, LP00, NJVA04, Kro08, PB06]. Predictive [FFME08, VDV00, Mag04, ZT06]. predictor [TK09a, TK09b]. Predicts [AMW00, AO03a, Smi08d]. Preemptive [ACK09]. prefer [CMCH08]. preferable [Bat09]. preference [Arc07b, DI09]. preferences [AR05, SWB06, WS06]. preferred [KJD05]. pregnancy [Buc04]. Preparation [HS03]. Presence [SP02, SHHD02, Wil01, BIS^{+}07, BBD06, BB07a, GSG^{+}07, MO07, SN09, VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08, YD09, vdBvdB09]. presence/absence [YD09]. present [BKKR08]. presentation [JG08, vdBR03]. preservation [IK06]. presomitic [GK09, TSZ^{+}07]. Pressure [DO01, GKM^{+}00, BC06, FTG07, GMK06, GK06a, GP08b, HED06, Mas09, MBD08, Shi06a, Tho05, YFH^{+}07, ZJ06]. pressure-concentration [Tho05]. pressures [OAC03]. prestin [SFC^{+}09]. prestin-associated [SFC^{+}09]. prevails [RON09]. prevalence [Bog04, DLM08, GOP09, TBR08, TBC^{+}08, dM09]. prevent [DSS08, ITL09]. preventing [HML09, HM05]. Prevention [ITKL08, JITJ09, TBB^{+}06, YKG^{+}05]. Prey [CLZZ02, DKD02, DHM01, FSGB^{+}02, Fur02, Mal00, RHG00, SFS^{+}01, WBB02, WTSN06, Abr09, Ama06, AKdlPP06, AMC^{+}09, BRC07, BGF07, Bon03, CDB09, DI09, FL09, GEF04, HS08, Jes06, KG03, KV04, LPC06, MC06, MPL06b, PCZL05, PCZL06, PM07, SX06, WT09a, Zhd03]. prey-dependent [SX06]. prey-flock [LPC06]. Prey-predator [WTSN06, KG03, Zhd03]. Price [Gra02, vV05]. Primary [EM00, MGP00, SCC^{+}00, WSC02, YNO09, Agi04, BRND09, DBBW09, DBBW11, FmW08, NGN^{+}04, SGTF07, SVG^{+}08, YCC^{+}06]. primate [Bon04a, Bon04b, BBK04, SKR06b, VN08]. primates [CH05a, Jam06]. primer [VSP06]. primer-binding [VSP06]. Primitive [AGCLMM03, GS02c, Nas01b, PMMS01, BS05c]. primitively [BIS^{+}07]. Primordial [CWJ07]. principal [PDC04, SPAH06]. Principle [Tor01, CG06, HMGK06, Smi08c]. Principles [GL01, KMP03, Ack04, FMHW08, RSSM06]. printed [SJD^{+}09]. Prion [PG00, GWPW04, GPMW06, Hav04, Mat06, ZZW07]. Prior [Van06a, KYS06]. Prisoner [LBSS02, AD07a, JLCS08, LW04, MS09a, Sch05a, WVA05b, WL07b, EL05, Ezo09, HS02a, IKD04, KK00c, LB07, Nei01, WR07a]. Private [Kra01a]. pro [NLM^{+}08]. Probabilistic [GF02, Khr04, McN06, TM00, BLOL07, MH06a, QD09, SWM09, VFC^{+}09, WVA05a]. Probabilities [Nak01, Nak03, APS08, HA09, Mic07]. Probability [HI07, MHDOG01, Sch02a, AS09, LP07a, Lin04, MS05b, PW09a, TDW07, VSP06, Wal07b]. Probable [PCB07, Bie06]. probe [GLSW07]. Probes [WP01]. Probing [Bea06]. Problem [GA02, MN01, Ova02, Wil02, XD02, CLM07, HB07, Jam09a, LMH04, ZSRB07]. procedure [BGE06b, CFM04, HQP08]. procedures [JHJ09a]. proces [AMD05, AMD06]. Process [Cub00, JW01, MM02a, OSCK01, PJ01, SRN^{+}00, BZ05a, BZ05b, BLF^{+}09, Cam03, CLMMP06, CC06b, Ger08, LFF06, LH06, SF04, TT04, WA07]. process-based [CC06b]. Processes [FB00, Kan01, KMI02, Möh00, WPE03, BL09, CDB09, CCF06, Dim05, FHL^{+}06, GMM09, IS08, MAL03, MCM^{+}09, MK06, MA04, NGN^{+}04, NSS08b, PD04, RBP^{+}09, SSB06, Tas05, VWR07]. Processing [Gag00, RHG00, Bon04b, JG08, Kri09, LFSG^{+}05, SV07]. processive [ITT09]. processors [PM03]. produce [CRB05]. produced [SH04]. Producer [Bea00, OT09]. producer-scrounger [OT09]. produces [WAL06]. producing [CSM02, JDMZ^{+}07]. product [Ano04b, Bat09, CZC05, LJK06]. Production [Ant02, HCMF01, HCB^{+}02, Her00, HYA02, SH05, SMBM00, SI02, WAMO00, BKD^{+}06, BPLS06, CRB05, CSD09, CSS06, FWLN04, FFME08, GCP04, LKM^{+}08, MB07, ND04, NHTM09, PCSL^{+}06, YFK05]. Productivity [Ama06, GJ07, ZJ06]. Products [VDV00, MRC09]. proenzymes [VGMM^{+}07a, VGMM^{+}07b, VGMM^{+}08]. Professors [Rot08]. Profile [BPZ^{+}01, DSU^{+}04, Wan06]. Profiles [JW01, PLB^{+}05, ZAD07]. Profiling [COdAM05, LNRR06, Plu06]. progenitor [BMH07, KM05]. progenitors [CPMG^{+}08]. Progenote [Di 01a]. progeny [DBBW09, DBBW11]. prognosis [RKF06]. Program [ABP^{+}03, Gra02]. Programme [RBN^{+}01]. Programmed [RBN^{+}01]. Programming [NJVA04, XWD^{+}01, PB07]. Progression [SA01, Lin07, QHF^{+}07, SRCDS08, SA05b, VHF08, ZAB^{+}09]. Progressive [CLS08]. Projected [OBPH^{+}08]. Projection [ZKFW08]. Prokaryotes [AøP03, dFG08]. prolactin [SVN^{+}05]. prolactin-receptor [SVN^{+}05]. Proliferating [LSMLB^{+}02, DSVBW07]. Proliferation [ABP^{+}03, Eic01, HSB01, JL00, SLHN06, VP01, ABTR07, BFP07, GPMW06, LN05b, LWRK07, WWS^{+}06]. proliferative [Shi06a]. Prolonged [MM08a]. prominent [GNH^{+}05]. promote [MI08a, MI08b]. promoter [WDH^{+}09, YZGW06, ZYW07]. promoters [LL06, PBMU^{+}09, VdL03]. promotes [BFK08, HH06a, PM03]. Promoting [KMI02, CSS08, FS04a]. Promotion [AF01]. prompt [GZL^{+}03]. Prone [JE01, DW08b, PCK^{+}05]. proneness [De 03]. proof [DT06, Pie09b]. Proofreading [BG00b, Has01a, KKR^{+}07]. propagated [LRT04]. propagating [PGN08]. Propagation [AMV^{+}02, PAA07, SRWL02, Tas02, TK04, Zhd03, CE08, FWLN04, FSS06a, FPL03, Sel06, SM06b, VN08, ZTKH09, Zhu09]. Propensity [Cal06]. Properties [AGCLMM03, ADMZ02, Ano01e, CE08, Leh00, Mac01, MHDOG01, RBJ06, Yan06, AHT^{+}07, AJ05, Cur04, Gar04, GL09, ITM^{+}07, IMN04a, JAA^{+}07, iKIY09, MBP07, OCA08, PP04, RMAI06b, SVS04, TW04, TP05, TLC07, YTL08, YZHZ09]. property [Kom04]. proportion [IK06, RB09b]. proportionate [KR07]. proportions [AWAB05]. Proposal [Kel01, IZGG05, Kau04]. Proposed [Mac00, Buc04, LZGL03]. proposing [BSGT08]. propositional [NTU06]. propulsion [MF05, MA07a]. prospects [GRH^{+}07]. prostate [LK08a]. protamine [Bie06]. protamine-DNA [Bie06]. protease [Smi08d]. protease-only [Smi08d]. protect [CSD09]. protection [GWM04, KSEK09, RG05, RDC09, Wod07, YKG^{+}05]. protective [JAFW05]. Protein [AOH03, AH03a, AøP03, BFH^{+}01, BHHS01, BNT^{+}00, CDDW02, CLXC03, Dal06, Dem02, DML02, EN03, FMF^{+}00, GB01, GAA02, GP08b, HCMF01, HCB^{+}02, JAA^{+}07, JAV07, KCT02, KJ02, LFLM00, LL01a, LLS^{+}09, LMT02, MJW00, NL03b, NJVA04, NSG08, PD00, SSLB07, SML02, Sil02, TML02, WKG03, WJMH00, WKL01, AH08, AFD^{+}06, ARR08, APS08, BM03a, BZ04, BTL08, BCL08, BD08a, BE08, CRJC04, CC06a, CFLC06, CdOWS04, CTZ^{+}06, CL07a, CCZC08, CS07, CHN08, DHYHR09, Di 04, DCL09, DRLB05, FmW08, FT07, FFTS09, GW06, GZG06, HSL04, Hav04, HGH08, HSK^{+}08, JAJH07, JBJ^{+}08, JHJ09a, JWWS08, JEHK06, Kur08a, KSR07, Lap03, LFF06, MX08, MGDBdM08, NSS^{+}08a, NS07, NSH^{+}03, OF01, Pán08, PSSS03, PBvdG09, PGLL07, RAK^{+}08, Rot07, SGCK^{+}08, SYC06, SC09, SB09c, SPN06, SW08c, SZ09, Tay06]. protein [WYXC05, WYC06, WZ08, WL04, XWC08, YLM03, YTL08, YPY^{+}09, YY07, YAL04, YZHZ09, hZzGqX^{+}09, ZAD07, ZZW07, ZDC08, ZLLZ09, Nas01b, OQGC07]. protein-coding [MX08]. protein-folding [AH08]. protein-protein [CdOWS04]. Proteinase [LGB02, VGMVR^{+}06]. Proteinase/Transglutaminase [LGB02]. Proteins [Kar03, Nas01b, PD00, ZSZ01, AS07, APL08, BC04b, CL07b, DT07a, DCC^{+}08, FMI05, HQP08, HBB08, HDW^{+}09, JAA^{+}07, JAJA07, JSA^{+}07, JHJ^{+}09b, JNWJWB04, Lin08, MBP07, MSGS08, MLWL06, MMUGD09, PCB07, ROR05, STW^{+}09, Sin06, TR08, VRAF06, YCC^{+}06, ZF06, vdGMG^{+}09]. Proteobacteria [Kun01, Kun05]. proteoglycans [FWCN05]. proteolysis [KP06]. proteome [MBP07]. Proteomic [Ano05h, Hof04]. Proto [YNL01]. Proto-tRNA [YNL01]. Protocell [MS07a, CSP^{+}08, CSP^{+}09, SZLM09]. protocols [JWB^{+}09, LS08a]. Proton [BNT^{+}00, Wil02, BS05b, DHW^{+}09]. prototypes [BdBH09]. protozoan [GRW03, NO04]. protrusion [DECEK06]. protrusions [IHV^{+}06, SMCT08]. provide [Leo07]. provides [BDMR06, MGL05]. proximal [SB07]. PrP [CPC^{+}00]. PrP106 [ZZW07]. PrP106-126 [ZZW07]. Pruning [KTH09]. Pseudo [MBBR06, SWM09, CZC05, CC06a, CTZ^{+}06, CL07a, DCL09, GKNT09, Lin08, PBMU^{+}09, RBS05, SYC06, WYC06, XWC08, hZzGqX^{+}09, ZF08, ZDC08, ZCLZ07]. pseudo-amino [CZC05, CC06a, CTZ^{+}06, CL07a, DCL09, SYC06, WYC06, ZF08, ZDC08, ZCLZ07]. pseudo-elastic [RBS05]. pseudo-folding [PBMU^{+}09]. Pseudo-nitzschia [SWM09]. pseudoknot [HR08]. Pseudomonas [FFME08, RDSB^{+}03]. Psoriasis [SWS02]. psoriatic [LN05b]. psychiatric [PFdP^{+}07]. psychological [KV05]. psychrophilicity [JAA^{+}07, JBJ^{+}08]. PU.1 [RG06]. Public [HDHS02, CHCC^{+}04, JG06, KAI08, Wak07]. puffs [GS08a]. pulled [Kal07]. Pulmonary [Bra01, CBB07]. Pulsatile [TMS00]. pulsatory [PP08]. pulse [KCV05, LL09, ZT08]. Pulsed [Fra02, LFLM00, LMF03]. Pulses [GJE02, HRBL02, RVMR01]. pump [SVG^{+}08]. pumped [Kur08a]. Pumping [BNT^{+}00, DO01, GG09]. Punctuated [Gag00]. Punish [HB01]. punisher [NI06, NI07a]. punishing [CH05a]. punishment [EFW07, JB08, NI06, NI07a, ND09, RON09, SN09]. purification [NA04]. Purine [CMB^{+}01]. Purine-load [CMB^{+}01]. Purpura [Her00]. Putting [FR06, JST^{+}02, PTFF05]. Puzzle [HW01, Twa04]. pylori [JK04, DT07a]. pyogenes [MCB07]. Pyrimidine [CMB^{+}01, RGW^{+}05]. Pyrimidine-load [CMB^{+}01]. pyrite [Kal07]. pyrite-pulled [Kal07].

q [PFdP^{+}07]. q-bits [PFdP^{+}07]. QPDR [MMUGD09]. QPSO [CSW^{+}08]. QSAR [TLZ^{+}08, VGDSU09]. QT [CBBH01]. QTN [HL06]. Qualitative [CS02, RDC09, BRCB04]. qualitatively [DDJ06, RBS05]. quality [Arc09a, BÅ03, CMCH08, Gol07]. Quantal [KFG^{+}02]. quantification [YTL08, YPY^{+}09]. quantify [CBB07]. Quantifying [AGW^{+}06, AGW^{+}08, CQLV^{+}03, MKD^{+}05, MB07, PG00, RWID^{+}08, JCW^{+}03, PGMK^{+}03]. Quantitative [Bla04, CSA07, Cha00, CGS05, CS02, DM00, KW00b, KS01c, MRJR09, TH00, AMS09, BRCB04, BFGB04, BFG05, CCGC02, Cui07, CY09, DS05, FMHW08, HL05, KZ05, LKW06, LF08, MH09b, ML09b, NKC^{+}08, PR08, PA09b, PA09c, Ril00, RYAI06, WHH07, ZAB^{+}09]. quantized [Ste04]. Quantum [BS00, GBD06, PSM06, VA04b]. quarantine [HKC^{+}07]. quarter [MV06]. quarter-power [MV06]. quartet [SJ09]. Quasi [ES00, LSMLB^{+}02, Nås01a, TE07, WS03a, TE04, TE05]. Quasi-extinction [ES00]. Quasi-Independence [WS03a]. Quasi-quiescent [LSMLB^{+}02]. Quasi-stationarity [Nås01a]. quasi-steady [TE05]. Quasi-steady-state [TE07, TE04]. Quasispecies [VSCG00, ALM09, BMS06, FLG^{+}07, SES08, TSS06]. queen [RHF07, KJ02]. queens [BIS^{+}07]. quenching [DVC^{+}04]. Question [RBN^{+}01, Yos03]. queueing [MGHF06]. Quiescent [LSMLB^{+}02, ARR08]. Quinones [PDIS00]. quorum [BdOP06, HSMM07, KH08]. quota [HK07].

R [NM08, PCB07]. rabbit [FMLP06, FGMP08]. rabbits [FLG^{+}09]. rabies [DHR^{+}07, DHRM08, HRSV06]. Rac [CSON^{+}05]. Race [WBB02]. races [GMMGDA05]. Radial [Kra01c, FPM^{+}06]. Radiate [KS01a]. Radiation [LFLM00, EH08, FCD^{+}05, LFP^{+}05, Lit07, SB09c, TK04]. radiation-induced [FCD^{+}05, LFP^{+}05, SB09c]. radiocarbon [BSJ04]. radiodurans [SB09c]. radiosensitivity [DSU^{+}04]. radiotherapy [DSU^{+}04, EAC^{+}06, SSB^{+}07]. radius [Bye09]. Raffle [BP00]. raft [IHV^{+}06]. rainbow [KRGH07]. Rainfall [NBMS06, GJ07, KvHMP09]. rainforest [PAA07]. raising [MBD08]. Ram [COS01]. Random [BZ00, FK03b, JK03, JW01, MS05b, RKNK01, SFV02, SFVA09, AP04, BEF^{+}06, BZ04, Bra07, CH05b, FPS08, FM07b, JA05, KR07, KKR^{+}07, MLWL06, Mie05, MI08a, MI08b, MGDBdM08, Mur07, PB06, RB09a, RWP^{+}08, SH03, SBM08, WVA05b]. random-community [MI08a, MI08b]. Randomized [GLSW07, Kon09]. randomly [AP04, FZ04]. Range [Ari05, JLS01, AA04, BTCD07, GAK^{+}06, HH06c, Ish00, KTP09, MPL06a, Pie09a]. Ranking [KSS07, HLP06]. Rapid [JP00, BN04, BBSLN06, BBSLN08, CHN08]. Rapidly [BR02, WCH^{+}09]. raptor [Lor06]. rat [KBT08, MRJR09]. Ratchets [YMC01]. Rate [AGT^{+}01, CMTU01, HS00a, Kan01, MGL03, MM02a, RFH^{+}02, See00, Alp05, Blu07, BTCD07, CH05b, FL03, FCD^{+}05, GVB^{+}08, GCP04, GC09, HLW00, HJ07, Jam08b, KMW03, LWFP08, MLWL06, MN06, MJ07a, NS07, NSH^{+}03, OD03, OP05, PC09a, ROR05, RWK08, SRS09, VFC^{+}09, VL09, Yan08]. Rates [BKF00, Cha01b, Dus00, LP00, ANT09, Bok06, Dus06, HEH^{+}09, ING04, KOK06, KT07, LJ09a, LBS06, Lew03, LB06, Mas08, Par04, RMRG09, SP07b, TLZ05, TMBD09, TNTJ08, WW08, Wod07, vAR09]. rather [For07b]. Ratio [GK00, Jam00, MMTS02, Now00, Tor00, CTB09, CSM05, Gra07, Jam06, Jam08b, Jam08c, Jam09b, Kli06, MT06a, MT06b, NS06, Wak05, WT05, XSD^{+}05]. rational [TG09b]. Rationality [Tul03, NWP07]. Ratios [BB02, CL00, Jam01a, Jam01d, GRW03, GI09, Jam04, Jam07, KAI08, Wak05]. Rats [CCP^{+}00, DRV^{+}08, Izs05, Sch08a]. ray [CS08, How09]. Rd [SP00]. Re [CBBH01, DRW01, SE02a, CWDM06]. Re-entrant [CBBH01]. Re-examination [DRW01]. Re-revisited [SE02a]. re-sampling [CWDM06]. reach [DT06]. Reaction [FY00, FK03b, HW01, MCK07, NTK02, PDIS00, SCGFS00, SIMK02, Sta00b, WP01, AB07a, AF09, BGE05a, CFM04, CHN08, DF09, Gie03, HS04, LS03, Lal06, MLA04, MSMKM06, MA04, NAS07, NS07, PWK03, SRR08, TK05, VGBA06, WB06, ZRSK07]. Reaction-diffusion [MCK07]. Reaction/Diffusion [Sta00b]. Reactions [Alb02, FY00, MHDOG01, SM02, TRM03c, AC04, Alb08, ASMD06, BCL08, DFC^{+}02, FP03, KD03, MH06a, RA06, SI09, SS06c, TRM03a, TRM03b, TE05]. reactivation [KNWCB07, PBHS05, Wil06b]. Reactive [DALP03, Oht04]. Reactive-diffusion [DALP03]. Reactivity [WP01, dQW07]. Real [DKD02]. Realistic [VMW02, Kon06, TY06]. Reality [MKT^{+}00]. realizing [CGK^{+}05]. Really [LBSS02, Cin03, LW04]. Reappraisal [SB09a, WB02]. Rearrangement [Moc02, TMI03b, EPR07, SBPC05]. Reasoning [KM08]. reasons [Har06]. Reassembly [MM02b]. Reception [JLS01, Bak07]. Receptor [BNT^{+}00, BSRH02, BSRH03, FTEG02, GF02, RB02c, SRN^{+}00, TTKZ01, WQ00, AMP06, AA04, AD06b, BTL08, CSD04, GSG^{+}07, GFWT04, HSL04, KS08a, LGB03a, LGB03b, LHFH08, PBvdG09, Saf05, SBPC05, SVN^{+}05, TLZ05, TWO^{+}09, WWS^{+}06, WL04, vdBWLS07]. receptor-ligand [CSD04, GSG^{+}07]. Receptors [OF01, Raf02, VRB01, WKL01, GFW^{+}09, HSF09, OY03, WMS08b]. recessive [GV03, MMV^{+}04]. Reciprocal [FLS01, Uit09]. reciprocation [BS04, CH05a, MR03]. Reciprocity [Gin00b, AD07a, BS06, BC04a, Fis03, MM03b, OI04, Oht04, OI05, OI06, ON07, OP05, PTON08, PB03, RON09, Rot09, SA07b, SA07c, SA08b, TM06]. Recognition [RKL02, CSR^{+}05, CC06b, KKR^{+}07, KS08b, LL06, PKS^{+}08, SRAL12]. recognized [Unn09]. Recombinant [WAMO00, BCJ^{+}08]. Recombination [FK03b, GK06b, GG01c, MM00a, Arc03, HFS06, PER03, WR07a]. Reconciliation [MJW00]. Reconciling [CC06b]. reconstitution [ØKRG04]. reconstruct [NGMS08]. reconstructed [Ger08]. Reconstructing [MSS05, SH06, TS08, WMP03]. Reconstruction [GKXS07, IYGA08, WLZ^{+}06, Wil08b, NSS^{+}08a, YYA09]. recorded [FB06]. Recovery [Sco00, BL08, CMW08, HQP^{+}09, TLZ05, WS04]. Recruitment [MSI01, SMD00, ILDP04, LP07c, ZAdlPLM07]. Rectangular [GJE02]. recurrence [ECAV07, YTL08, YPY^{+}09]. Recurrent [EMS02, OKTS02, VdGN^{+}05, CCV08, JWWS08]. recursive [II06, YFK05]. Recycling [MHDOG01, Rua01, GW06, Gro04, KUK07, SW08c]. Red [EP00, KJ02, CSD09, PRP^{+}03, SBI07]. Redefining [BDR08]. Redox [HSB01, HDNC04, LIK^{+}05]. reduce [GLE^{+}09, NdGG06, RHF07]. reduced [DRV^{+}06, MGP^{+}08, MH05, RRC^{+}06]. Reduces [CP03]. reducing [YMLK04]. Reduction [BS05b, BHJ03, KKDV01, MSdlPS09, DRV^{+}08, MA04, QWQ09, SWM09]. reductive [Kal07]. Redundancy [PPEøP02]. redundant [DMPP09, Fra08]. redwood [DFCL08]. Reef [Arm01, ILDP04, wTA09]. reentries [TPP^{+}04]. reentry [CE08, JB04, SWC^{+}08]. Reexcitation [CPM^{+}09]. referential [DLB07]. referring [NI04b, NI04a]. reflected [SBZ^{+}08]. Reflects [Di 01b]. reflex [SHPDL03]. refractive [GMMR07]. refuge [CC09b]. refuges [PGH^{+}04]. regard [SHP09]. regarding [Gel07, vV06]. regeneration [CLW^{+}08, DLRP07, IZGG05, SI06]. regenerative [BFG09]. regenerative-loop [BFG09]. Regime [FCP03, wTA09]. regimens [KW07]. region [MPOBD^{+}09]. regional [AP04, SAC^{+}04b]. regionally [LAG07]. regions [OS03, YY07]. regression [BR06, CY09, DW08a, LD09b]. Regressive [Lac01]. Regular [DVL^{+}00, LBJE03]. regulate [LKW06]. Regulated [FY00, SSJ09, MTS05, PSJ04, RDYH09]. regulating [DHYHR09, NLS08, NS07, TLZ05]. Regulation [BHOR01, CSP01, CP03, EH08, FFN00, Ish00, KB02, KL09, LL00, LSD^{+}00, Mac00, MVM^{+}00, MAB00, SMBM00, SWS02, TN01, YBV^{+}00, Aba09, ACSY04, BS05a, BRC^{+}03, Cib08, DGM05, DMPP09, DTM07, FM06, GZK06, HSL04, HJ05, JCK09, KLL07a, KI04, KI05a, KH08, LK06, LCHB06, MGDDH08, MJ07b, MBJ09, Nar06a, NP07, PCSL^{+}06, PRP^{+}03, PHL08, RGF^{+}08, RGW^{+}05, RB09b, RKRW08, SA07a, SCS04a, SS05a, SVG^{+}08, Tao04b, TZS07, Voi03, WSS07, XBT09, XT06, YCS04]. regulations [KOT07]. Regulative [NK08]. regulators [PLB^{+}05]. Regulatory [AO02, AGZ^{+}06, BWHK02, BdOP06, JP00, KA02, Mac00, SML02, øPVO02, AO03a, ABKR07, AKLS05, BDBR07, BCRG04, BB09, CAS05, DB08, FM07b, Fra08, IvDHI08a, IvDHI08b, Kau04, LS04b, LR07, LGCL07, LKM06, LPB^{+}04, LT08, MSS05, Moc05, Moc08, OB04, PHdB09, QB06, RKYH06, RK07, RB09b, RH08, SI09, ST03, SVGK07, SR08, TK08, Tao04a, TB04, TKKA04, WRG^{+}04, WW07, XT06, YMLK04, YLM03, ZRSK07, dV06]. Rein [SKW00]. reinfection [BB05, GWM04, GWM05, GWM06, RMRG09, Sch03]. reinforcing [BL06b, GMAH07]. Reinhart [CB08b, CB08c, Maz08, MH08, Rap08]. Related [AIK00, WAMO00, CTS06, DCC^{+}08, GBZ06, HH04b, LT06, LZ09a, MK09, MMUGD09, RSH^{+}06, WC07a, Won05, dM09]. Relatedness [Lio09, Pep00, FFTS09]. Relating [WW05b, BBSLN06]. Relation [BPZ^{+}01, UD07, AH03b, Bla04, CCF06, DVC^{+}04, FGMP08, Kok04, NDE06, TW04, Yan08]. relation-like [AH03b]. Relational [Gre09]. Relations [BR01, KGD09, KG03, ZFVH05]. Relationship [Cha03, Coo01, Raf02, AB06, BdBH09, CS07, CTB09, DBBW09, DBBW11, FL03, KF06, Kon06, LJ06, Leo07, LV08, MR07a, Niw05, PL09, UCSZ07]. Relationships [BNT^{+}00, HVPN02, Kal00, MLWL06, BEK^{+}03, Dus06, LZ09a, MKE^{+}09, PC09a, RKOS09, Sav04, WR03]. Relative [FMP01, Her00, Bal04, Cib08, GMAH07, KLL08, Lin04, NAV04, WL09, YDFQ05, vdBR04b]. Relatives [BJ02, Mar09a]. relaxation [IPY07, WL03]. Relaxing [HE08a]. Release [HFGB02, PV00, BFGB04, BFG05, CWJ07, HS07, LK08b]. Relevance [AGCLMM03, CPMG^{+}08, LMVPM07]. Relevant [DKD02, RKH^{+}06]. Reliability [BB09, GG01a, Har06]. Reliable [MC01, RB09b, VRB01]. Reliably [BKF00, Ben04b]. remembrance [Ost04]. Remigration [Hay00, Sol01]. Remodeling [Mar03, RMBM00, BR06, HdGH07, LTG^{+}04, NMH07, SKS09]. remodelling [DWBB04, KH07a, THL03, WBR09]. Remote [YTL08]. removal [HH07]. remove [TPP^{+}04]. Removing [GF09, LST00]. Renewable [PG01b]. renewal [CPMG^{+}08, DdB03]. Rényi [KmMK04, VA04a]. Reorganization [Wan00, CABB09]. reorientation [KH09a]. Repair [Has01a, MS01a, PN02b, Bon06, FCD^{+}05, GVB^{+}08, KKR^{+}07, KRN03, TT04]. reparation [VA06]. Repeat [RPB03]. Repeated [PTON08, BC04a, DRV^{+}06]. Repeats [DBHS00, RGPB08]. repellent [RRH08]. Repertoire [Cha03, VRB01, RWP^{+}08, SBM08, SMPvdB08]. Repertoires [DFP01, RKL02, RG05, TSK09]. repetition [KWGE04]. Repetitive [Tas05]. replacement [OPN07]. Replicating [KY02, ZW07]. Replication [CSC03, GKM^{+}00, HA00, Leh02, Ste02b, HHP05, NS09a, SKR06a, SKR06b, SSDM06, XSD^{+}05]. Replication-associated [GKM^{+}00]. Replicative [SK01, KF04, PC04]. Replicator [HDHS02, KI09, Kom04, AM04a, CHR06, ON06]. Replicator-dynamics [KI09]. Replicator-mutator [Kom04]. Replicators [LL01b, SS01]. Reply [CCGC02, EK08, HP00a, Rot08, For04]. repopulation [Lit07]. Representation [KA02, FM07a, HK05, Khr04, LDXW06, QWQ07, WE06, YPY^{+}09, YAL04, YSW09, ZLLZ09, ZYD^{+}05]. representations [Bat09, DqLmW07]. representative [LZ09b]. Represented [LLCM01]. repression [HJ05]. repressor [AO03b]. repressors [BM04]. reprobation [BS04]. Reproducers [Ker00]. reproducing [JWH08]. Reproduction [SI00, YK03, ACK09, ASD08, Fuk05, MS07a, NDE06, NI07b, OI00, WE06, Yan08]. reproduction-mortality [ACK09]. Reproductive [BK05, LD00, Now00, BFA08, CHCC^{+}04, CC09b, ELSFB07, FGMP08, Kom07, KC07, MGHF06, STMH04]. reprogramming [Wil06a]. Reptiles [See00]. Reputation [Mce03, MM03b, OI04, OI05, OI07]. Require [KKDV01, AFZ08]. Required [Kru02, OCA08]. Requirement [RK06, DB05, Gol08, JEHK06]. Requirements [DGS09, CGS05]. requires [DdB03, SI09]. Rerun [SSA00]. rescaling [PGMK^{+}03]. rescue [Mit04]. resemble [LY03b]. resensitization [BFGS07]. reserve [BLOL07, TMH04]. Reserves [EH00, Pel00, SH04, SH05]. reservoir [AWO^{+}09, RP09b]. resetting [FP04a, KMGDG04]. Resident [MD01]. residual [BMT04, JAA^{+}07, MTS05]. Residue [LMT02, GKTN07, RA08]. resilience [JLCS08]. Resistance [BH01, BH03b, HB02b, JL00, PG01b, Wod01, ADHM09, BR08, CC09b, DMO^{+}07, DPA05, HI07, HLA09, HML09, ITLN09, JWH08, Kom06, MV06, MWB05, MG09, MWD08, PCB07, PGM00, PGH^{+}04, RGFP07, XAP07]. Resistant [HB02a, UHI02, AAEW09, ABP^{+}00, MPM07, MMV^{+}04, MGL^{+}06b, PGH^{+}04, MWD08]. resistive [KC03]. resolution [MN07b, NP08, WT05]. Resolutions [ES01a]. Resolve [Eic01]. Resolving [Chi07, FS09, MPN07, OS03]. Resonance [AM01, FSG00, FSGB^{+}02, GRG02, ING04, KGG07, Mil05, SHRR06, VA04b]. resonant [AB07a]. Resource [AKS07, DFP01, Hur06, LM01b, PG01b, Rev00, Rev02, SFS^{+}01, VW00, YHI07, EW09, FL09, GNLEK07, KG06a, MW09, PM03, Sat04, TY09, Wak07, WR03, YOYT07, YB05, YKG^{+}05, vKPdR07]. resource-biodiversity [WR03]. resource-competing [TY09]. Resource-enhancement [YHI07]. resource-species [KG06a]. resources [BB03, KLL07b]. respect [BWvK^{+}08, GSG^{+}07, Yam03]. Respiratory [FW01, Kan01, CKE06, Hen04, MSDM06, VPGA07]. responding [CSON^{+}05]. Response [BNT^{+}00, BWHK02, CS09, FSG00, Fra02, Han01, KSM02, KW00b, Lie05, LL01b, PG01a, Ril02, SMBM00, TRM03c, VMW01, Abr09, Agi04, AA04, BRC07, Bra05, CTB^{+}05, Cog06, DRV^{+}08, DRV^{+}06, DMPP09, DKLL05, Di 00c, DSU^{+}04, EHG03, HSC07, HML09, Hua03, KRNG08, KV04, KMGDG04, LPC06, LLW09, LZS^{+}08, MS03a, MK04, MBB^{+}06, MM03a, MV07, NKL06, OM08, Ort06, PCZL05, PCZL06, PWVvR09, PCA^{+}09, PBR01, RRC^{+}06, RGB06, Rue07, SZ04, SMPM09, SMNL07, SSB06, SZ09, TY09, Tan07b, TRM03a, TRM03b, WAL06, dVG04, dVG06]. Responses [Ace00, ABP^{+}03, GS02a, Kar03, Wod01, WJ01, BdOP06, CDFP04, DMQ04, HM05, Jam04, Koh07, PE04, RRKF09, WMS08b, YLM03]. rest [GGH^{+}05]. resting [Jäc07]. Restitution [CE08]. restoration [HH07]. restraints [Osb08]. restricted [FWLN04, GG02]. restriction [Bar05, NT04]. Result [öWO01, MWCS04, TMBD09]. Resulting [MLMW01, CdOWS04, Kon03, WBR08]. Results [DRV^{+}08, MBB02, AHCN07, BTL08, KD03, wTA09]. retaliation [JB08]. Retama [FLS^{+}04]. Rete [SWS02]. reticulated [FI06]. Reticulitermes [MKN02]. reticulocyte [PBZ06]. Reticulum [AGT^{+}01, DBR^{+}05]. Retina [Moc02, EH08]. retinae [HH04b]. Retinal [BKCR01, Che00, TMI03b, MRJR09]. Retinoid [EM00]. Retinopathy [RBBH02]. retranslocation [Arc07a]. retrospective [HKC^{+}07]. retrotransposition [RRK06]. Retrotransposons [XD02]. Retroviral [PDA^{+}00, PD01, VSP06]. Return [WGH01]. reveal [HEH^{+}09, Ken07]. Revealed [BPZ^{+}01, HAC^{+}09, ZFVH05]. reveals [FPBM08, HGH08]. Reversal [YMC01, HMP04, SPC02]. Reverse [Nas01b, LS04b, SCNP^{+}06]. reversed [GT06b, SIK03]. reversible [TE04]. Review [WRF01, Voi03]. Revising [BDMR06, MGL05]. Revisited [DS00a, HS00b, JBP02, CH07, MBJ09, NM08, Poh08, Row06, SK09a, SE02a, ZAdlPLM07]. Revisiting [Coo01]. reward [CJLS08, Här07a, JLCS08, Toy09]. rex [HNTHA07]. RHDV [FLG^{+}09]. Rhizobia [PV00, MCC^{+}09]. Rho [CSON^{+}05]. rhodopsin [SN06]. Rhythm [RWR00, RVMR01, UHK01, IB06, INSR08, KA03b, SG04]. Rhythms [Dai02, LAG07, Ref04, SRR08, THKU07, TIM06, XK07]. ribosomal [NN07, OS03, Sel06, Sel07]. Ribosome [LSD^{+}00, GW06]. ribosomes [GPN05]. Ribozyme [Ste02b]. rice [BKMH07]. richness [Gos06]. right [HH07, vV09]. rigid [NM06]. rigidum [CEP05]. Rinaldo [MGL05, MGL06a]. Ring [Yos03]. rings [Ber03]. rippling [AV05]. Rise [Laz03, CY07]. rises [LMF03]. Risk [Bon03, CLZZ02, HIN00, HI00a, Har01, Jam01c, PKW^{+}00, Wag03, WH01, HI05, HF03b, KY03]. risks [BKM09]. risky [IMNT09]. ritronavir [NLM^{+}08]. rival [HH06a]. river [Cud05, MWL^{+}07, MBRRI08]. RNA [BZ05a, AS07, AJSL07, BZ05b, CXZ^{+}09, CGK^{+}05, CMB^{+}01, DW08a, DM07, Dor03a, DL08, DLM03, GWPW04, HR08, JPJ06, JR08, KD06, Leh02, MPOBD^{+}09, Nas01b, Sel06, STW^{+}09, SSDM06, SBM08, Tay06, WDH^{+}09, YCC^{+}06]. RNA- [YCC^{+}06]. RNA-binding [AS07, STW^{+}09]. RNA-guided [AJSL07]. RNA/Protein [Nas01b]. RNAi [Bah07]. RNAs [CMB^{+}01, Kun00, LRD04, NN07, OS03]. RNY [Leh02]. roads [DLM03]. Robust [CCC08, DMPS05, FK03c, RDF03, UHK01, WPE03, AWAB05, AB03, BFK08, BH03a, GKTN07, SR08, SSF09, Wan06]. robustly [IS08]. Robustness [ABKR07, CAS05, FK01c, MWB^{+}02, SYYI07, WW07, BRCB04, BWvK^{+}08, Est07, FV09, FAW06, GK06b, HMB^{+}08, KH09c, Ref04]. Rock [KNS05]. Rock-scissors-paper [KNS05]. Rodents [Går00, Jam04, PM03]. Rodgers [EK08]. Role [AWJ02, AB00, BRC07, BL00, Bra01, BHJ03, CPM^{+}09, CBH02, EM00, EMS02, GB01, HG02, JAA^{+}07, KK06, KK00b, Kun00, LZS^{+}08, LBJE03, MSGS08, NI06, NI07a, NC02, PD06, Sta00c, The00, UI02, WJ01, Aba09, Agi04, Ama04, AS07, AM06, Bon03, CEP05, CFCGCC03, Ded08, FPBM08, FDA^{+}09, FGMP08, FWE06, Gre05, GCC08, GPN05, HML09, JWH08, KSB07, KSEK09, KA03a, KM05, LB05, LJ09a, LRHB09, MAL03, MRGADD08, MM09, ML09b, MM03b, MPL06a, MJ07a, PBL06, Pic06, PC04, RM04, RGPB08, SES08, Sch03, SKR06b, SH08a, SRG^{+}03, SCS04b, SGGM05, TLZ05, VAAH05, WG06, XBGN05]. roles [DMPP09, Est07, HLA09, KF04, Mar09a, MBRRI08, RRH08, Yan08]. Root [LSMLB^{+}02, MAB00, Mur00, BCKE^{+}08, CKS04, Di 08, DFSD05, HML04, LBCL09, RF04b, SP07b]. Roots [SSKL01, BLL08, Bio08, Che07, Mig06, RF04a, RF04c]. Ropalidia [BIS^{+}07]. Rotation [FMP01, Nir02, NM06]. route [JMvdB09, PB07]. Routes [TU00]. RP [KA02]. RP-Effects [KA02]. RPE65 [GZG06]. RPs [KA02]. rRNA [AIK00, YCC^{+}06]. rRNA- [YCC^{+}06]. rRNAs [OS03]. ruffled [Kin04]. rufiventris [DPV00]. Rufous [DPV00]. Rufous-bellied [DPV00]. Rule [FI00, For00a, Sug02, VMW02, BM09, JD09, Row06, ST03, ZXWF08]. Rules [Bea00, BDMR06, BP08, BS04, HH04c, HH05b, KCS^{+}06, KGL09, MGL05, MS09a]. rumen [BKD^{+}06]. ruminoreticulum [VTC08a, VTC08b]. rumor [KSY^{+}08]. run [QD09, RDH09, RBS05]. Runaway [ND09]. Running [MKN02, BGF03, BG06, BB06b, BB07b, GSB05, HNTHA07, LW08a, RGG00, RBS05]. Rupture [FMP01, FHD09, ZHB04]. RuvAB [Xie07]. RuvAB-mediated [Xie07].

S [EK08, How09, CLW^{+}08, MCB07]. S4 [Gre05]. Saccharomyces [GKM^{+}00, Thi04]. saccular [KH07a, KH09b]. Sacculus [Koc00]. Safe [VRB01]. Safety [FMP01, LMVPM07]. sagittal [SH08b]. sagittal-plane [SH08b]. SAH [WZW^{+}07]. Saharan [MBH03]. Salamandridae [DOT02]. salinarium [CR02]. saline [NKC^{+}08]. Salivary [FY00]. Salmonella [AS00, Kee05, LGS^{+}09, LTLM09, PMB06, XBCF05, XFBC07]. Saltatory [Mal02]. Sample [MC01, Toy09]. samples [Eti09]. Sampling [BRCB04, Chi07, CH05b, LS04a, VFC^{+}09, Che06b, CWDM06, SPAH06, Sta09, WFGP04, WT09b]. sampling-based [Sta09]. Sampling-rate-dependent [VFC^{+}09]. Samuel [For09]. sanction [MCC^{+}09]. Sand [Yos03]. sapiens [SKR06a]. saplings [Che07]. Sarcomere [DSCD02]. Sarcoplasmic [AGT^{+}01]. SARS [CFCGCC03, HKC^{+}07, MPN^{+}05, NLM^{+}08, WR04]. SARS-CoV [NLM^{+}08]. Satellite [RPB03]. sativa [Mei05]. Satoshi [Gel07]. Saturation [CMW08]. scaffolds [LWRK07]. Scalar [MLA04]. Scale [CL05, Ken07, KBD06, PPEøP02, VG01, WR03, BNRW04, CLH07, FP03, FM07b, IiKY07, KC03, KHHS09, MS08a, NA02, Ots08, PAA05, SP00, SBZ^{+}08, XBGN05, ZLXY08]. Scale-dependence [WR03]. Scale-free [CL05, KBD06, FM07b, IiKY07, KHHS09]. Scales [Ell01, BH03b, GSRC^{+}06, GS08b, KSR07, RRS^{+}07]. Scaling [Cha01a, CMW02, NL03b, PW09b, SBvS06, Tay00, Tho05, TP05, WMLC02, BK05, BDMR06, Bio08, CY07, Dem06, HJ07, Hui09, Kai04, MGL05, MV06, MFI09, Niw04, Niw05, RA08, San03, Sav04]. scan [Raf02]. Scanning [BG00b, BPZ^{+}01, Cui07, RWP^{+}08, TR08]. scapularis [OBPH^{+}08]. Scarce [WSC02]. Scarring [CS00]. scatter [SMPM09]. scavengers [RH04]. scenarios [DRV^{+}06, HBWB08]. Scent [LM01b]. Schaefer [BDR08]. schedule [Wak05, YOYT07]. scheduling [ZT06]. scheme [DF09, LME06, LZ09b]. Schemes [LMT02, CA09]. Scheuring [LL01b]. Schizophrenia [Jam00, Rǎd08, Rǎd09]. School [Ano01h, GBK^{+}07a, Niw04]. school-formation [GBK^{+}07a]. Schools [IK02, GBK^{+}07a, Niw05, ZKH^{+}05]. Science [Win04]. scissors [KNS05]. scope [LHDvdM04]. score [ST05].

Scores [LLCM01]. scoring [PGLL07, SA07c]. screening [LT06]. Screens [WKL01]. Scrounger [Bea00, OT09]. Sea [Arm01, Kal00, Zac09]. Search [KMI02, VMW02, FP04a, HC07b, MD03, OI07, WDH^{+}09]. searches [BCV^{+}08, RB09a]. Searching [PB06, SCH05b]. seasonal [AB07a, HBOS05, ING04, PL09]. seasonality [GJ07]. Second [Ano09a, BP08, SSB^{+}07, Can07, ZXWF08]. Second-order [BP08]. Secondary [ERM00, BKKR08, CXZ^{+}09, DW08a, KSR07, WZ08]. Secretion [MW01, TMS00, BD08a, GCB^{+}07, JK04, Luc05, OHM05, Rot07, WTC09]. secretory [LKM06]. sedentary [EL09]. sedimentation [CC09a]. Seed [ARKL02, CR01, HH06c, KY00, LBF01, PM03, SA05a]. Seed-cache [PM03]. Seed-parasite [LBF01]. seeding [BEK^{+}03, CLHW07]. seeds [MA03b]. Seek [NS03b]. segetum [RRS^{+}07]. segment [AO03a, Han04, ZSZ^{+}06]. segmental [AD06a, HNTHA07]. Segmentation [KMH00, RPB03, XWD^{+}01, Cin03, GP08a, GK09, RGSFM07, UMI09]. Segmenting [ST03]. Segments [Lan00a, RMBM00, EPR07]. Segregation [BRC^{+}03]. SEIS [WC07a]. seismology [HB07]. seizure [RR08, UL06]. seizures [KRR09]. selectin [LCHK09]. SelectinE [VS08]. selecting [IS09]. Selection [DFP01, Di 00a, DPA03, Dus00, Dus06, FAW06, Fuk05, GBK^{+}07a, JP00, KW00b, KOK06, KK00c, LL01b, Loc08, LK03, MK01, MM02a, OLS^{+}02, OMT03, OKS01, RKNK01, Sta02, TL02, The00, WT01, Wel00b, WC02, WL02, YK03, YLL00, YL00, YFK05, ATO^{+}09, BP03, BdABA09, BGB08, BFR05, BH03b, CC06b, CT06, CSB^{+}09, Cor05, Dal06, DGZ07, Dic08, Fer09a, FR07, FR06, FL09, FZ07, FSS06a, Fuk04, Gab06, GLSW07, GGPFM08, GA08b, GCC08, HH05a, IMN04a, Jam04, JWH08, Kel07, KA03a, KTH09, KSN03, KK06, LJ09a, MCF04, MR03, Mas09, MHMGM08, MFM^{+}03, Mie05, MS09b, ND09, RBJ06, SH03, SC05, SW08a, Sch09, SW08b, SA05a, SP06, SBZ^{+}08, SSR04, She06, Smi08a, Smi08b, Smi08c, TAN09, TOA^{+}09, TPN07, TNP07, WS06, WT05, WT07, Wit03]. selection [WVA05b, dBWBP06, vV06, vVH07, vV09]. Selective [Kra01a, OAC03, LAG09, SK09a]. Selectivity [SSB^{+}02]. Self [BR00, BPZ^{+}01, BK01, CDDW02, CBC08, DS00b, FMI05, Gab06, GJ07, LL00, ORM03b, ORM03a, OI00, RMF08, Sac04a, AC04, CSS08, CS06, CPMG^{+}08, CMB^{+}01, For01, HS04, ILDP04, KG06a, KA03b, LRD04, MS07a, MB09, MCM^{+}09, MNMH^{+}08, MS05b, NLS08, NL03a, OB08, Oud05, QO08, SRS09, Sat04, VA06, WH07, XBT09]. self-attraction [SRS09]. Self-cluster [CDDW02]. self-consistency [CS06]. Self-deception [BK01]. Self-disturbance [BR00]. self-incompatibility [Sat04]. self-limiting [NL03a]. self-maintaining [MNMH^{+}08]. Self-maintenance [OI00]. Self-Organization [BPZ^{+}01, CBC08, FMI05, GJ07, RMF08, Sac04a, KG06a, LRD04, MB09, Oud05, QO08, WH07]. Self-Organized [ORM03a, ORM03b]. self-organizing [AC04, MCM^{+}09]. Self-other [Gab06]. self-promoting [CSS08]. self-recruitment [ILDP04]. self-regulating [NLS08]. Self-Regulation [LL00, XBT09]. self-renewal [CPMG^{+}08]. self-reparation [VA06]. Self-reproduction [OI00, MS07a]. self-sustained [KA03b]. self-sustaining [HS04, MS05b]. Self-synchronization [DS00b]. Self/Not [For01]. Self/Not-self [For01]. Selfish [EFW07, MKN02, VMW01, VMW02, Arc03, JBK04, NI06, NI07a, RV05]. Sel’kov [NM08]. Semantic [KM08]. Semelparity [RTK02]. Semi [CCGC02, MR07b, Ril00, SN07, GJ07, IPY07, TBB^{+}06]. semi-arid [GJ07]. semi-dilute [IPY07]. Semi-discrete [SN07]. Semi-log [MR07b]. Semi-quantitative [CCGC02, Ril00]. semi-stochastic [TBB^{+}06]. Semicircular [MV02a, MV02b, MV02c]. Semiconservative [TSS06]. sempervirens [DFCL08]. Senescence [MP09a, SK01, Ark05, KF04, PC04]. Senile [EKS02]. Sense [Ano01f, Tan01]. sensing [HSMM07, Kri09, KH08, NSI08, Nar06b, SMNL07, SN04]. sensitive [HH04b, Ped07, RA06]. sensitivities [SV05]. Sensitivity [Ace00, CZM09, HLS01, IS03, LG04, LTLM09, NAS07, PG01a, Vei03, Cal06, DGZ07, MHRK08, QD09, SHP09, TBB^{+}06]. Sensor [Bar01, SN05b]. sensors [DMPP09]. sensory [LAG09, Lew03, LB06, LFSG^{+}05, MDCC06, SA08a]. Separate [Yan08]. Separately [FGH01]. separating [Pen03]. Separation [TMI03a]. September [Ano00h, Ano00n, Ano01k, Ano01q, Ano02g, Ano02m, Ano03w, Ano03-29, Ano04-35, Ano04-39, Ano05-33, Ano05-38, Ano06-36, Ano06-41, Ano07-36, Ano07-42, Ano08-32, Ano08-38, Ano09-29, Ano09-35]. Sequence [AOH03, AøP03, BFH^{+}01, Cha00, FS09, JAJA07, MM02b, XWD^{+}01, BdBH09, BE08, FT07, FFTS09, JAV07, JQR08, Kur08b, LGK^{+}09, LGK^{+}12, LLS^{+}09, LZ09b, Mei05, MGDM08, PWZ09, QWQ07, Ste09, WDH^{+}09, YY07, YSW09, ZYD^{+}05]. Sequence-Based [AøP03]. sequence-dependent [WDH^{+}09]. sequence-structure [BdBH09]. Sequences [CDDW02, DVL^{+}00, DBHS00, GBCC01, Jav00, Leh02, NA02, OKS01, PAD00, WJMH00, XWD^{+}01, CRL06, DqLmW07, FmW08, For07a, GHC03, KMC^{+}07, LTW06, LDXW06, MK03, NA03, Pán08, PKS^{+}08, PGLL07, SKR06a, SKR06b, STW^{+}09, SBM08, TFYY03, VA04a, WLZ^{+}06, YZGW06, YAL04, ZC06, ZYW07, ZW03, ZYD^{+}05]. Sequencing [KSO06, RBS05]. Sequential [BCV00, FY00, FFHK09, HB09, SC09]. Sequential-Arrivals [BCV00]. sequentially [SH04]. sequestration [BFGD07, LGB03a, LGB03b]. Sequoia [DFCL08]. Series [HI00a, SZ01, WP01, AFD^{+}06, GXG03, HRSV06, KRGH07, Ref04, Wan06, Whi05]. serine [dCZJ^{+}04]. serpines [BKKR08]. Sessile [MSI01]. set [DZB^{+}04, TGN07]. setpoint [GGH^{+}05]. sets [HS04, RK07, Yan05, Yan06]. settling [McN06]. several [BRC07]. Severe [Sco00, CSD09, ND09]. severity [CY07, CXM^{+}09]. Sex [Bel06, BB02, CL00, FMD01, For00a, Gag00, Jam00, Jam01a, Jam01d, MMTS02, Ste02a, Tor00, Wel00b, XD02, YLL00, CSB^{+}09, CW07, CSM05, For07b, GRW03, Gra07, GI09, GT06b, Hel08, HMP04, Jam04, Jam06, Jam07, Jam08b, Jam08c, Jam09b, KAI08, LVH04, MSP03, MT06a, MT06b, MM09, MS03b, Mit04, PDM04, Pec06, SS03, Sat04, Wak05, WT05, Yam03, RHF07]. Sex- [YLL00]. sex-determining [Mit04]. sex-ratio [CSM05, WT05]. sex-reversed [GT06b]. sex-specific [CSB^{+}09]. sexes [Jam09a]. Sexual [CMB02, For00a, Jam01a, PDC02, TL02, YK03, ASD08, CC06b, CT06, Cor05, DBBC08, Fuk05, GOP09, GCC08, HI04, JE09, KI09, PDM04, SP06]. sexuality [YYY^{+}08]. sexually [GOP09, HK09, MM09, UH09]. shade [ZAdlPLM07]. Shannon [DPV00]. Shape [CR01, CACC02, Kaw01, KVMV01, LF01, PPGS03, PKA02, SB00, Abr09, AGZ^{+}06, BdABA09, GEK04, HBWB08, JQR08, KSG03, KH07b, PBZ06, PBZ08, RLCIB05, SBZ^{+}08, WO04, WFGP04]. shaped [BLZ07a, BLZ07b, LZ09a]. Shapes [JL01, CGK^{+}05, KSB07, Osb08]. shaping [Ded08, Yan08]. Shared [Bon03, CdOWS04, CJLS08]. Shared-enemy [Bon03]. Sharing [DS00a, MT06a, MT06b, Uit09]. sharp [Ros05]. sharpens [IS08]. Shear [Bar01, SKN^{+}03, CSON^{+}05]. Shedding [CB07b, IGHW07, LCHK09, USTG09]. sheds [KS08a]. shell [SS06b]. Shepherd [AKR09]. Shielding [AR08]. shift [IFN07]. Shifts [HCC00, YFKP03, BTCD07, SZ04, wTA09]. Shimojo [How09]. shock [SZ09]. Sholl [MR07b]. Shoot [PKL02, FLS^{+}04]. Short [MZK08, PE00, BRC^{+}03, SBM08]. short- [BRC^{+}03]. Short-Distance [PE00]. Shortening [SK01, SGD04]. Should [CMCH08, KV04, ITKL08, MH06b, OI04, OI05]. show [AA04]. Shows [CDDW02, PPEøP02, AS00]. shrub [FLS^{+}04]. Shuffling [MM02b]. shuttle [LZS^{+}08]. shuttles [NM08]. sialic [CPG09]. sialyl [VS08]. sialyloligosaccharide [SV03]. Sib [KY00]. Sibling [Tor00]. siblings [Row06]. side [BTA08, NGN^{+}04]. siderphore [FFME08]. Sign [GVK00, Kra01a, SPC02]. Sign-reversal [SPC02]. signa [Cor05]. Signal [CMB^{+}01, HVY^{+}07, KPS02, Kri09, Noe00, RGB06, SPC02, SSB06, AFD^{+}06, Arc09a, CRL06, De 03, FB08, HDNC04, Har06, JDMZ^{+}07, KTP09, KWGE04, KY03, MS03a, MCB07, NS05, NS09b, NSI08, PKS^{+}08, RM04, SP05a, SV07, SS09a, SYYI07, SPB06, SM06b, VdL03, Wil06a]. signal-transduction [NS05]. Signaling [Cum00, DT07a, FK01a, GSB01, RG03, AKLS05, BH03a, CLB05, DLF^{+}07, FWCN05, GP08a, GBK^{+}07b, HSL04, HH08, Jäg08, Ker04, Ker06, LHFH08, NSS^{+}08a, OAKC08, PP04, RWCK08, WL04, ZPP08]. signaling-induced [OAKC08]. Signalling [De 02a, SRMW00, Wes08, BSR06, HH04a, Hur06, RGSFM07, WWS^{+}06, YCS04]. Signals [VP01, HG08, JDMZ^{+}07, Jäg08, KCV05, MDCC06, MH06b, MR07a, PBHS05, PBR01, RK06, Rob03, SNA^{+}08, Thi04]. signatures [PCB07, Paw09a]. Significance [TMS00, Bon06, CMG06, IZGG05, LN03, WSM05]. Significant [PPEøP02, MWCS04]. silencing [DRMLS09]. Silico [EKS02, AOH03, BKKR08, CABB09, GKB03, WRKK09]. similar [PFRR08, SZC^{+}03]. similarities [Lap03]. Similarity [LBJE03, BM09, BD08b, SM09b]. Simple [ARW00, BE03, BHJ03, CY07, Går00, Gie06, Gro02, HS02a, HR00, ISC01, JJ01, JSV02, LJW02, OAKC08, SES08, Ten08, VA06, WSS07, AC04, ASMM08, AW06, ACK08, BST05, CHN08, DECEK06, Ded09, GSB05, GI09, HA07, Lew05, MGS08, MRF07, PNG03, San03, SW08b, SVGK07, SCS04b, SG06, Tak06, VdL03]. simpler [CSS06]. Simplest [BSRH02, BSRH03]. simplex [HHP05, NS09a, PBHS05]. Simplification [FSG00]. Simplified [BT06, TLA00, BPLS06, CMG06, LT08, NHTM09]. simplifying [DB08]. Simulate [FC01]. Simulated [KTH^{+}00, NO04, RV05, Shi06a, Wal00, öOW01, DBGM08, KD06, LBS06, Sta08, UKY^{+}09]. Simulating [AMD05, AMD06, BZM05, BLF^{+}09, MKLD02, RVMR01, SPB06, WR04, ZJG03, ZAD07, HTN04a, HTN04b]. Simulation [ANL01, BZ00, CC01c, CMS^{+}00, ELB02, FY00, GJE02, GLV02, Ker04, Ker06, KS01c, KA03b, MGS08, Rue07, SDL^{+}08, wTA09, TS02, WQ00, öWO01, ACK08, BMH07, BKMH07, BL08, BGB08, DGM05, DSU^{+}04, FPBM08, FVP^{+}07, GHHR03, HK05, HED06, LCL03, LS08b, LDW04, LDW05, LKW06, LME06, ML07a, ML07b, MSS05, PGF^{+}08, SBZ^{+}08, TSZ^{+}07, TNTJ08, Usa06, VN08, dQW07]. Simulations [KRNKH00, LP00, BS05c, CV08, GFWT04, GBK^{+}07a, HBB08, Lap03, LJ09b, LCHK09, Man06, NLM^{+}08, SKY09]. Simultaneous [AH00, GZL^{+}03, JBP02, SB09a, CH07, KR05]. Singapore [CFCGCC03]. Single [BB02, DSCD02, HS03, KK00a, Kut03, PH03, Sev06, TH03a, TGR^{+}00, TTKZ01, VLFN00, YK03, BGK09, Bra05, BTL08, CHR06, Cyt04, DHYHR09, FLG^{+}07, For07a, GFW^{+}09, GE05, GPN05, GZK06, Hua03, Kut05, Lei09, Lei10, MSGS08, MKA05, RvMK^{+}05, SRG^{+}03, SK09b, TOB08, VWR07]. single- [GZK06]. Single-Channel [TTKZ01, GFW^{+}09]. single-gene [Bra05, MKA05]. single-host [VWR07]. single-infection [SK09b]. Single-neuron [Sev06]. Single-solute [TH03a]. Single-species [BB02]. single-stranded [For07a]. single-substrate [GPN05, SRG^{+}03]. Single-turnover [TGR^{+}00]. Singular [MW07, FP03, GLE^{+}09]. singularities [DB05]. sink [Ama04]. Sinoatrial [öOW01]. sinuosity [Ben04b]. Sinus [öWO01]. sinusoidal [KCS^{+}06, WKRD09]. SIR [AB07a, KR07, KGG07]. SIRVS [ZT08]. Site [Kut03, TGR^{+}00, Jus08, WOLS07]. Sites [KMP03, Mac00, DL08, DHW^{+}09, HL06, HS07, KOK06, KmMK04, Lio09, OBL04, SN09, VSP06, Xie09, dQW07]. situ [BMH07, FBUL05]. sizable [Rot09]. Size [Ano01g, AROS07, CC01a, Cha03, Che03b, EH00, FI00, GTDA02, HI00a, Hay00, IKS00, Kaw01, Ken01, KMGV00, KVMV01, LST00, MGL03, NL03b, Now00, RH02, SS03, SH04, Sat04, Arc09b, Ari05, BPC08, BC09b, BR04, BG06, Bog04, BFA08, CFM04, Dus06, Fow09, Fuk04, GR08a, GBGAKD05, Här07a, IKD04, JG06, Lac09, LMH04, LBS06, LS08b, LHDvdM04, MV06, MFI09, MLPJ09, SH05, SWM09, SA05a, TK09a, TK09b, TH03b, WFGP04, YYY^{+}08, YHI03, dV06, vKPdR07]. size- [BPC08]. Size-dependent [FI00, SS03, Sat04]. size-independent [dV06]. size-invariant [MV06]. Size-number [SH04, SH05]. size-scaling [MFI09]. Size-structured [AROS07, vKPdR07]. size/distribution [LBS06]. size/number [SA05a]. Sizes [MC01, HRZ06, HP04, Niw03, Niw04, SA07b]. sizes-bias [HP04]. Skeletal [JMB00, Laz02, Mar03, FKAC06, HTCS07, JTGP06, LBQ^{+}05, LK08c, Mag04, PC09b]. skeletons [Zac09]. skew [MX08]. Skin [Ano01e, Mac01, SM01, BGE05a, BGE05b, BGE06a, BGE06b, LN05b, PGN08, SVN^{+}05]. Skylight [BH04]. Sleep [PR08, MKS^{+}09, Muz05, RKRW08]. slenderness [GMAH07]. sliding [FHD09]. slime [TTT09, TKN07]. SLLE [WYXC05]. slopes [Ait08]. SLOSS [Ova02]. slotted [SM06a]. Slow [Eic01, Ker00, CHN08, ES08, LBSP06, MSdlPS09, Ped07, Saf05, SW08a]. slow-fast [MSdlPS09]. slowing [CF09]. slowly [WCLL08]. slug [DO04, VW03]. Smad [NS09b]. Small [EN02, MS02, BNRW04, BOvD08, DCC^{+}08, GBK^{+}07a, HP04, LD09a, LRD04, LBSP06, MLWL06, MLPJ09, SK05, SBvS06, VdGN^{+}05]. small-scale [BNRW04]. small-size [MLPJ09]. small-world [SK05]. smallest [OBDJ06]. smallpox [NDE06]. SMB [Ano00e]. Smith [Ano06-54, Gav06, Har06, SS06d]. Smooth [HFGB02, BFG08b, BFG08a, FB08, KBT08, NMH07]. SMP [MRC09]. Snake [JLS01, Bak07]. snowdrift [SPS09]. SNPs [Ken07]. soaring [RH04]. Social [BW01, Bea00, BHOR01, KB02, NC02, OI06, Pep00, SMD00, Arc09b, BCH03, CLMMP06, DMH08, DW05, DBF07, EG07, HMND06, Hau06, KGL09, Kro08, MP09b, MM03b, NL04, OI07, OB08, PRT04, Sac04a, SLIL07, Tof06, Tul03, UM08, VN08, WS06, YY06]. Sociality [Gin00b, PSWK09]. Societies [GTDA02]. Society [Ano00e]. sodium [Dim05, Kur08a, SVG^{+}08, WCH^{+}09]. Soft [LDT09, Lap03]. Soft-cuticle [LDT09]. soil [BNRW04, BÅ03, BSJ04, WRNB04]. Soldier [CSS06, HYA02]. soldiers [AK04]. solid [AM04b, GA07, MF05, SGGM05, WBR09]. Soliton [Sin06]. Soliton/exciton [Sin06]. Solubility [DNS00, Kat03]. soluble [MRC09]. solute [MM03a, TH03a]. solutes [SCNP^{+}06]. Solution [CP03, DALP03, CTS^{+}08, GSB05, GC09, LMH04, LN03, LWC09, LHDvdM04, MZK08, MC06, TE05, VS08, Wax09]. Solutions [CT02, AB07a, BÅ03, MW07]. Solvent [GKTN07]. Soma [Zhi02]. Soma-to-Germline [Zhi02]. Somatic [Fra03, Kok04, MFM^{+}03, SK01, WW00, BFK08, BI09, SLL06, YKG^{+}05]. Some [BZ00, Dor03b, IMN04a, Jam01c, Jam08a, Mur00, Ort06, RMAI06b, Wil08a, CRB05, CMB^{+}01, Gel07, HH04b, Jam04, LY03b, Nar06a, TLC07, KV04]. Somite [KW00c]. Somitogenesis [CMS^{+}00, KMH00, Cin03]. Song [RKH08]. sons [Den08, Kan05, KV05, Kan06]. Sorting [CKJ^{+}02, SFV02, KD03, Pal08, UI04a, UI04b]. SOS [BKP09]. SOS-induced [BKP09]. sounds [Oud05]. Source [BR00, DO01, Ama04]. source-sink [Ama04]. sources [Bat06]. sp. [Bur09]. Space [AOH03, BFH^{+}01, Cha00, CP03, KKDV01, LF01, MSI01, Niw04, PB06, CHR06, DRV^{+}08, FT07, FFTS09, HTN04a, HTN04b, HQP^{+}09, JBF^{+}03, KB04, LF08, MK03, MDL04, PSPF07, RHH08, SP05a, SES08, TN04, TTN05, WFGP04, WW07]. space-filling [HTN04a, HTN04b]. Space-irrelevant [Niw04]. Space-limited [MSI01]. spaced [TN04]. Spaces [SSWF01]. spacing [ASMM08]. spanning [Byw09]. sparks [GS08a]. sparse [NSS^{+}08a]. Spatial [CC01a, CKJ^{+}02, FM02, Fra00b, FJBK05, GS02a, HDF04, Hau06, JJ01, KC05, KG06a, LNH08, LM01b, MY09, MKLD02, MI08a, MI08b, PPL^{+}00, PBR03, SLIL07, She06, SNA^{+}08, WK05, WRF01, AC07a, ACLW03, ACK09, ASD08, Cib08, DDJ06, GS08b, HAC^{+}09, Hui09, IKD04, JE09, JLCS08, KSB07, KUK07, KC03, KD03, Kri09, LMF08, Lio09, LFP^{+}05, LJL08, Mar04, MS09a, RKH^{+}06, SN09, SFVA09, SPVN06, Wak07, WMS08a, WKB07, WAL06, YHBW04]. Spatially [CLZZ02, Gre00, HS00a, Hie05, LSMB^{+}03, OSBH02, PL01, RSBY03, Sch02b, WMS08a, WZL00, BGF07, BGO08, CG06, CSM05, FLG^{+}09, Har07b, HM07, JI05, KB04, KBD06, Lio09, MS07a, MSH05, PWK03, SS06a, XBGN05, XFAS06]. Spatio [BGF07, Jac03, MGP^{+}08, SCH02c, FM04, PCS^{+}06]. Spatio-temporal [BGF07, Jac03, MGP^{+}08, SCH02c, FM04, PCS^{+}06]. Spatiotemporal [BR00, MPL06b, SRWL02, SIHH04, SSL08, YHM^{+}06]. spawning [MA03a]. Special [Ano06-54, CD00]. Specialists [Wah02, DW08b]. Specialization [KW00c, PPL^{+}00, RHF07]. specialized [Dru03]. Speciation [DM00, JP00, KRN01, PC09a, Bol06, For04, Gav06, RC09, RMAI06b, ZXWF08]. Speciation-rate [PC09a]. specie [BTCD07, HH06c, RRKF09, YD09]. Species [AP01, BH08, Dai02, FM02, For00a, HLH01, HS03, JSV02, KRN01, Krá01b, LS09, OKTS02, Ova02, Pel00, PPL^{+}00, SD02, AWO^{+}09, AJ05, Ari05, BLMV05a, BLMV05b, Bol06, BB02, CC06b, Chi07, CF09, CE05, CPR00, CGH01, CHR06, Dam04, Eng07, Fei08, GF09, Gos06, GEF09, GZK06, GT06b, Hui09, KS06, KDK02, KC05, KG06a, LJ06, LM08, MCM^{+}09, MT06a, MT06b, ML08a, ML08b, PCB07, PDC04, PCS^{+}06, PC09a, PM08, RDH09, RHM^{+}08, Sav04, SSLB07, SSR04, TH05, TBCD06, VDV00, WLHB07, YHI07, ZZLT07, vdBvdB09]. Species-area [Eng07, PPL^{+}00, LJ06, PC09a]. Specific [Fra00a, Kur08b, AMP06, BFGS07, CSB^{+}09, Cha01b, GZK06, HAC^{+}09, IvDHI08a, IvDHI08b, JAHKH09, LJK06, LPT05, MLMW01, Mei05, MJW00, PGLL07, Wil06b, vKdRP05]. specification [MKL09, RG06, Van06a]. Specificity [HI00b, FABdC04, KS08a, ZPP08, vdBWLS07]. spectra [BR04]. spectral [CKE06, HAC^{+}09, OR04]. Spectroscopy [KSM02, KHE06]. Spectrum [MLMW01, WMP03]. speech [Loc08, Oud05]. Speed [BR01, BGF03, BG06, Dus06, KS06, LRT04, MM08a, RP09a]. Speed-frequency [BR01]. speratus [MKN02]. Sperm [BP00, BP03, Blu07, Dus02, HH06a, JE09, SS00, SDS04, YOYT07]. Sperm-dependent [JE09, SDS04]. Spermatozoa [BI09, Nir02]. sphaerocarpa [FLS^{+}04]. sphere [Won05, ZSRB07]. spherical [LB06, Pic06, Wax06]. spheroid [KLN^{+}09, PA09b, PA09c]. spheroids [VHF06]. sphingolipid [AVSHV04]. Spider [SMD00, VCBV^{+}06]. Spiders [CIV09, FMH08]. Spike [OUPG09]. Spiking [ERM00, GTG01]. spillover [AWO^{+}09]. Spindle [Yos03]. spindles [Kok04]. spiral [GG09, HED06]. Spiralling [SW04b]. spiriferid [SKY09]. Splenectomy [Her00]. splicing [CHdV06]. Split [Eld00]. Splitting [Péc05, BFGS07, JTGP06]. Spongiform [CPC^{+}00]. Spontaneous [ERM00, Nar06b, PCA^{+}09, SN04, LH08, LH09, MHKS03, NSI08, SWC^{+}08]. sporadic [GWPW04, Wal07b]. Sporulation [RVMR01, MSS05]. Spots [KWGE04, DMW08, SIK03]. spp. [AIK00]. Spread [BR02, For02, LT06, NL04, Tas02, ADHM09, AWO^{+}09, EL09, GLE^{+}09, GBB08, ITLN09, MCK07, Mat06, MP09a, MLPJ09]. Spreading [BW01, CTS06, DSS08, FDA^{+}09, KRR09, MY09]. Spring [GSB05, SPG06, BB07b, RSB09, SH08b]. Spring-mass [GSB05, BB07b]. spring-mass-like [SH08b]. springtime [TNTJ08]. sprout [ASMM08, BGB08]. sprouting [JTGP06]. square [PKS^{+}08]. squid [LRT04]. squirmers [IPY07]. Src [CLB05]. St [YM04]. Stability [AM04a, CC01b, CGH01, CHR06, DC09, GMSW04, JJ01, LLCM01, OST09, Rou03, Sch00, SR02, TT02, TT03, UB01, ZZLT07, AF09, BRC07, BP08, Che03a, CL09, CK08, CG03, CH07, DI09, DDJ06, EBM07, Fow09, GSB05, GEF04, Her09, Jäg08, KGD09, Kon06, KH09c, LR07, LV08, MB09, MGHF06, Mic05, MPN07, Nei04, ON08, OQGC07, Paw09a, ROR05, RSS04, RRKF09, Sac07, SB09b, Sel06, SLL06, TO06, UD07, Van06b, WL03, ZJ05, TT01]. Stabilization [DSVBW07, VGBA06, HS08]. Stabilize [HB01]. stabilized [HI04]. Stabilizes [FLS01, FMD01]. Stabilizing [FS04a, RKNK01]. Stable [Ant02, BS01, Lan00a, LBSS02, PDA^{+}00, YLL00, AM04a, AD07a, BSR06, LW04, ML09a, ML12, MH05, MP05, SAC^{+}04b, TYW05, Uit09, YYY^{+}08]. stacked [WYC06]. Stage [Rev00, iTYU06, BKKR08, Di 06b, FMHW08, JHJ^{+}09b, Kal07, Pin00, SX06, vKdRP05]. Stage-dependent [iTYU06]. stage-specific [vKdRP05]. Stage-structured [Rev00]. staged [SHRR06]. Stages [DEM^{+}00b, FDA^{+}09, KB04, PGF^{+}08]. stalling [ITT09]. stand [ZAdlPLM07]. Standard [NA02, GKXS07, ZF06]. standing [MM03a, PB03]. standpoint [SA05a]. Staphylococcus [ABP^{+}00, MPM07, PWVvR09, MWD08]. star [MGDBdM08, MMUGD09, PBMU^{+}09]. star-graph [PBMU^{+}09]. Starting [FP04a]. STAT [SYYI07]. State [BNT^{+}00, EMS02, NWP07, SML02, WP01, WH01, AKLS05, Bra05, CP03, DS04, GRBR04, GC09, IS03, JAV07, LIK^{+}05, LWC09, LCL07b, MKT^{+}00, MH06a, MK06, MLWL06, Mur00, OQGC07, PBHS05, RGRB04, SF04, SSR04, TO06, Tao04b, TH03a, TE04, TE05, TE07, WFGP04, WW07, XT06, YLM03, Zah00]. State-Dependent [EMS02, NWP07, SSR04]. state-space [WW07]. States [EP00, CSD04, MGAR07, Moc05, Net09, SAC^{+}04b, TK05, vAR09, Tho00]. Static [dSKL09, HM07]. Statics [LD00]. stationarity [FHL^{+}06, Nås01a]. Stationary [Gie03, Gie06, Ref04]. Statistical [AH03a, Bas01, BDMP^{+}08, DP08, DOT02, Ish00, LBQ^{+}05, RC00, RKL02, TL00, WMLC02, dSKL09, BC09a, CB08a, Cui07, KS08b, MGT^{+}06, SVS04]. Statistical-Mechanical [Ish00]. Statistics [GBB08, HR08, TFYY03, BSJ04, CKE06, LL05, Tao04b, Van06a, XT06]. stature [SS03]. Status [DLGC02, Dic08]. status-based [Dic08]. Stay [EKIL01, IFN07]. Steady [CSD04, CP03, EP00, Mur00, SML02, XT06, AKLS05, Bra05, IS03, MH06a, Moc05, Tao04b, TH03a, TE04, TE05, TE07, WFGP04, YLM03, BNT^{+}00]. Steady-state [CP03, Mur00, XT06, AKLS05, Bra05, MH06a, Tao04b, WFGP04, BNT^{+}00]. Steady-states [CSD04]. Steep [BGP00]. Steepness [NM09]. Stem [FK01c, HVPN02, MLMW01, WW00, DBBW09, DBBW11, LWRK07, ØKRG04, OS03, QO08, RG06, Ros03, SRCDS08, Wod07]. Stems [Nik02, PBB03]. Stents [RMBM00]. step [BZ04, KSPA^{+}08, MM03a, NS05]. Steps [Orr03, Ros05, RA06, SYYI07]. stepwise [WRKK09]. sterile [TNTJ08]. Sterility [For00a]. stichotrichous [EPR07, PER03]. Sticking [WL02]. sticky [RSB09]. sticky-spring [RSB09]. stiffness [GMAH07, LW08a, Mil05, Pal08, TOB08]. stimulated [THL03]. stimulating [VGS^{+}05]. Stimulation [Has01a, KBT08]. Stimuli [CR02]. stimulus [AA04, Cyt04, KJ08, Rue07]. stirring [RB06]. Stochastic [CB05, CDB09, EN02, FPBM08, FSG00, FSGB^{+}02, GKXS07, GTG01, HS02b, HE08b, KOT07, Kom06, LHFH08, Lo07, LVV^{+}01, LP07c, Mal02, Man06, MNI06, MCN08, NSI08, PLB^{+}05, SSB^{+}07, SBPC05, SK01, SYSY02, SMPvdB08, TTR00, TSN05, TNP07, WT09b, XAP07, YNO09, YB07, YYI03, ZSS02, AH03b, AG06a, BM08, Bal04, CDFP04, DF09, Fen03, FL03, FCD^{+}05, Fuk04, GG07, GMY09, HHBY06, HRZ06, HSMM07, HC07a, HS07, IB06, JB06, KR07, KRB05b, LL05, LVL08a, LVL08b, LCL07a, LCL07b, LL09, MO07, MTS05, MPM07, MJ07a, ØØS06, OUPG09, PB09, PHL08, RRK06, RDYH09, RDF03, RBP^{+}09, SD06b, Sim08, Tan08, TZS07, TS08, TBB^{+}06, WOLS07, YB05, ZRSK07]. stochastic-dynamic [IB06]. stochastically [HP04]. Stochasticity [HBLG02, Lei09, Lei10, BB07a, EW09, HLA09, LPMC^{+}06, SCS04a, Ste04]. Stoichiometric [CBH02, MNMH^{+}08, MJW00, FP03, Gro04, IS03, NS09b]. Stoichiometry [AGT^{+}01, Gro02, BKD^{+}06, MK09]. stoichiometry-related [MK09]. stomatal [KRNG08, Kor09]. Stone [Bak07]. stony [Mig06]. Stopover [ELB02, WH00]. stopping [FP04a]. Storage [SE02b, SI02, FS04b]. store [Alp05]. straight [SI05]. straight-striped [SI05]. straightness [Ben04b]. strain [AB07a, CKM05, CW08a, CWDM06, Eam06]. Strains [GS00, RSBY03, RGG00, ABP^{+}00, BB03, HM05, TYW05, ZSF^{+}07]. Strand [ZSZ01, LAG07, MX08, PCH^{+}05, Wal07b, YDL06a]. Stranded [VO00, CMB^{+}01, For07a]. strandings [SPVN06]. strands [MX08, ZXWF08]. Strategic [AR05, NS03b]. Strategies [BS01, Dus01, ELB02, HP00b, ILDP04, KF03, KK00c, LP01, LBSS02, Nei01, PGM00, SD02, SCH02c, AM04a, ADHM09, ATO^{+}09, CEA07, CXM^{+}09, CC09b, EAC^{+}06, FYX^{+}09, FR06, FBM06, HGC03, HH04c, HH05b, Kom07, LTI08, LS07, LW04, MAC06, MWB05, Oht04, PBC09a, PLH05, RWCK08, STMH04, TY06, TAN09, Ten08, TBB^{+}06, VHF08, WR07a, WVA05a]. Strategy [ANT09, HS02a, HS00b, HYA02, MM00b, PDA^{+}00, SI02, TOA^{+}09, YLL00, AD07a, BL08, BRS^{+}09, CH05a, CLS08, CSS06, CC09b, HDZ^{+}07, HA09, IS09, Joh08, JITJ09, LYZ08, LPT05, MT06a, MT06b, RBW09, RGZ09, VCBV^{+}06, Yan05, Yan06, YHI07, ZT08]. Stratified [PKL02]. Stratonovich [Bra07]. Streak [PMMS01, BS05c]. Strength [Di 00a, Tay00, AGW^{+}06, AGW^{+}08, FHD09, Her09, KJJ07, RvMK^{+}05, Wes08]. Streptococcus [MCB07]. Streptozotocin [CCP^{+}00]. Stress [Bar01, GMAH07, HdGH07, RMBM00, SKN^{+}03, SK01, TT04, AJ08, CSON^{+}05, HG08, KF04, KH09a, MCMS06, RKH08, SB09c, WL03]. Stress-induced [Bar01, RMBM00]. Stress-modulated [HdGH07]. Stresses [Dor03b]. Stretch [FTEG02, KH09a, NMH07, SHPDL03]. stretch-induced [KH09a]. stretched [CL05]. stretching [LDW04]. striatal [SM09a]. striated [Nos06]. strict [ND09]. strict-and-severe [ND09]. Strikes [MD01]. Striking [CD02, HH07]. String [RPB03]. Strings [CD00, LW08b]. striped [SI05]. Stripes [Kor07, SIMK02, SIK03, KWGE04]. stroke [DBG06]. Strong [FZ04, Gin00b, CE05, MBJ09, NPN09, RBJ06]. strongest [BH04]. Strongly [AS00]. Structural [BWvK^{+}08, CLXC03, CLO^{+}02, FFTS09, FLS^{+}04, Kal00, LR07, LL01a, Mil05, MFB01, NSH^{+}03, SHP09, Tas02, APS08, BdBH09, CFLC06, CTZ^{+}06, CCZC08, Dru03, ELJ06, JAJH07, JAA^{+}07, JAJA07, JHJ09a, LZ09a, MBP07, PPRI08, SBM08, Tas05, Twa04, VGBA06, WBSY06, XWC08, Yag09, YPY^{+}09, ZDC08]. Structural-Functional [Kal00]. Structurally [Lan00a]. Structure [AH03a, AGCLMM03, BKE03, DSY01a, DSY01b, DHM01, FM02, GBLV03, GVK00, GZG06, Ish00, Kru02, LMT02, Moc08, Noe00, NJVA04, PPEøP02, SHF02, Tur02, Wax09, AJOK09, AB07a, ASD08, AKLS05, AP04, BMH07, BLMV05b, BdBH09, Bie06, BKE04, BKKR08, CSA07, CLMMP06, Che03a, DLF^{+}07, DMQ04, GBD06, GZG07, HDZ^{+}07, HMN09, Hol06, Izs05, JAV07, KJJ07, Ker04, Ker06, Kon06, KP08, KSR07, LFC04, LMVPM07, MGAR07, Mar04, MWCS04, MPD^{+}07, MGC04, MBRRI08, PAA07, Paw09a, PB09, Rak04, RG05, RMAI06b, SGS^{+}05, SAC^{+}04b, SFVA09, SM03, SX06, SMPvdB08, ST05, UM08, VS08, VN08, WKB07, WW07, YCC^{+}06, ZFVH05, ZAdlPLM07]. Structured [BR02, CL00, HS00a, KKK00, PL01, AHCN07, AP04, AROS07, DLB07, ELSFB07, FLG^{+}09, HM07, KBD06, Lio09, LYZ08, LLW09, OSK^{+}05, RNP04, RMP08, Rev00, SW09, SM09b, TL02, TOA^{+}09, TDW07, Vaz07, XBGN05, XFAS06, vKPdR07]. Structures [DOT02, PBT02, ZMQX01, CXZ^{+}09, DW08a, DGS09, HR08, JR08, MCK07, RRKF09, SB07, WZ08, Yun05]. Structuring [Di 05]. Student [LLCM01]. Studied [Di 01b, Laz03, BS05c]. Studies [CBH02, DGZ07, Jam01c, WQ00, AAK^{+}09, GZG07, MPOBD^{+}09, MGT^{+}06, WCLL08, WCH^{+}09]. Study [AS00, AM01, BPV01, BP01, CSM02, ELB02, GBD00, GLV02, GP00, GP01, IST01, KSM02, KmMK04, KMI02, MW01, NTK02, NA02, OSCK01, RB02c, TH03a, VLFN00, Wan00, YFKP03, öWO01, AHT^{+}07, AO05, BRC07, BL08, BM06, BG08, BBSLN06, DB08, DSA^{+}06, FDA^{+}09, FHL^{+}06, GABK08, Gie03, GXG03, GBGAKD05, HKC^{+}07, HJ07, Izs05, JA05, JR06, JK04, Ker04, Ker06, KG03, KMW03, LCL03, LIK^{+}05, LR07, LS08b, LB06, MS03a, MRJR09, Moc05, PDB08, PD06, PBB03, PLG^{+}06, Saf09, SHPDL03, SWI07, SBZ^{+}08, SV03, SM09a, TPP^{+}04, TNTJ08, Usa06, VCBV^{+}06, WW08, ZZW07]. Studying [WPE03, ZRSK07, SWN07]. Stx [EBM07]. Sub [Bur09, XWD^{+}01, CSM05, MBH03]. sub-individual-based [CSM05].

Sub-molecular [XWD^{+}01]. Sub-optimal [Bur09]. sub-Saharan [MBH03]. Subacute [CPC^{+}00]. Subcellular [Kut03, CL07a, CL07b, UW09, ZLLZ09]. SubChlo [DCL09]. subchloroplast [DCL09]. subcutaneous [CTB09]. Subdivided [WZL00, TH03b]. subfamily [ZCLZ07]. subgraph [ELJ06]. subgroup [SJ09]. Subject [Ano03-44, Ano03-45, Ano03-46, Ano04-61, Ano04-62, Ano04-57, Ano04-58, Ano04-59, Ano04-60, Ano05-57, Ano05-58, Ano05-59, Ano05-60, Ano05-61, Ano05-62, Ano06-55, Ano06-56, Ano06-57, Ano06-58, Ano06-59, Ano06-60, Ano07-55, Ano07-56, Ano07-57, Ano07-58, Ano07-59, Ano07-60, Ano08-52, Ano08-53, Ano08-54, Ano08-55, HH04b, MT06a, MT06b, MW09, SP06]. Subjected [RKNK01, NMH07]. subjective [Hur06]. subjectivity [TM06]. submission [MH06b]. submitochondria [hZzGqX^{+}09]. suboptimal [GWM04, LS08a]. Subpopulation [KTCD00]. Subpopulations [PL01]. subsets [FT09b]. subspaces [KP08]. Substance [KB02]. substitutable [RNP04]. Substitutes [SP02]. substitution [BW06, GKTN07, GMY09, MX08, MWCS04, WTL08, WW08]. substitutions [JS07]. Substrate [ADMZ02, MVM^{+}00, RGG00, VGCGM^{+}02, Wan00, Bat09, GPN05, Nar06a, SRG^{+}03, TE07, YLCW06]. Substrates [CCGC02, Ril00, BS05b, RNP04, TOB08]. substructures [KSS07]. subthreshold [SM09a]. subtilis [MBP07]. subtle [MWCS04]. Success [Dus00, Dus02, BCV^{+}08, CM03, LSH06, PDM04, RB09a]. Successful [MGP00]. succession [YNO09]. Successional [AP01]. such [Di 01b]. Suckling [DSK02]. Sudden [HLW00]. Sudden-death [HLW00]. Sufficient [CSP^{+}08, CSP^{+}09]. Sugar [DLGC02, Che07, LN03]. suggest [BBB^{+}07, OW07]. suggested [Nos06]. Suggests [LN02]. Suicide [VGCGM^{+}02]. sulfate [FWCN05]. sum [EAM07, HE08a]. summation [Agu08, PWK03]. Summer [Ano01h]. sunset [HHH06]. superba [GMMR07]. Supercritical [MAL03]. superfamily [MBBR06]. superoxide [GMFS06, KLK06]. superparasitism [HGC03]. Supersaturation [HHR01]. superspreading [GR08a]. Superstability [BB07a]. supertree [DMW08]. Suppes [MK01]. Supplementation [SFS^{+}01]. Supply [GBG01, RH08, TKM06]. supply-demand [RH08]. Support [CLXC03, DML02, DRLB05, Jam07, Nas01b, AS09, dCZJ^{+}04, CTZ^{+}06, JSA^{+}07, MRC09, MBBR06, ØKRG04, QLHL09, RBS05, Tan07a, YCC^{+}06, ZLLZ09, ZCLZ07]. supported [FHD09]. supporting [BBSLN08]. suppressed [Mas08]. suppresses [Mas09]. Suppression [LPLC00, CBR04, KRR09, LLC03, NKL06, SI09]. suppressor [IMKN05, KSN03, MI06]. suprachiasmatic [KA03b]. Surface [Bar01, Cum01, TM00, AW06, CMG06, FWLN04, Kal07, LCHK09, LMF08, Lew05, MD04, MGAD09a, MGAD09b, SPN06, WM04]. surfaces [Won05]. surgery [KSO06]. Surname [MZ02]. Surnames [SGS^{+}05]. Surprises [BGG06]. surprising [Sta08]. surrounding [LF08, TN04]. Survivability [SL00]. Survival [Eic01, JL01, Ker00, LVV^{+}01, MH09b, SS01, HF03a, Joh08, KJJB07, Mil08, SES08, WVA05a]. survival-of-the-flattest [SES08]. surviving [Lin04, ML09a, ML12]. survivorship [MGHF06]. susceptibility [MGT^{+}06, RMRG09, RGFP07, ZTKH09]. susceptible [MSDM06]. susceptible-infected [MSDM06]. Suspended [Mcn00, MN01]. Suspension [COS01, IPY07]. Suspension-Feeding [COS01]. sustain [WG03]. Sustainability [PM09, Eng07]. Sustainable [SD06a, SD06b, BDR08]. Sustained [BGP00, KR05, KGG07, KF04, KA03b, OCA08]. sustaining [HS04, MS05b]. SVIR [LTI08]. SVM [APS08, CRJC04]. swallowtail [OST09]. Swans [PE00]. Swarm [FSGB^{+}02]. Swarming [SD02]. swarms [YKiP^{+}09]. Swimming [Arm01, BG08, Dus06, Usa06, Ver04]. swirling [Zhu09]. Switch [CA00, MVM^{+}00, DHYHR09, LM09, OM08, RDYH09, SM09a]. Switches [KCA03, BH03a, CD05, MWS09, Thi04]. Switching [Fur02, MD01, øPVO02, CB07a, RGZ09, SSF09, TB04]. Switzerland [CAHH06]. symbiont [SH03]. symbionts [MKE^{+}09]. symbiosis [Ezo09]. Symbiotic [PBT02, MKE^{+}09]. Symmetric [KRNKH00, NA03, YY06, YM05a, ZWT^{+}08]. Symmetries [JS07]. Symmetry [Di 00c, Nas01b, Ber07, Cro00, OPN07, TIN06]. Sympathetic [HFGB02, KFG^{+}02, BFGB04, BFG05]. sympatric [Bol06, Gav06, RC09, Wak04]. symplast [MM03a, Pic06]. synapses [SGTF07]. Synaptic [Car02a, FLBB01, BFG08b, BFG08a, BFG09, FL03, MZK08, MS07b, PD06, SGTF07]. synaptotropic [Nie06]. Synchronization [BGO08, HBLG02, LN05b, CSP^{+}08, CSP^{+}09, DS00b, GBZ06, Hen04, HSMM07]. Synchronized [SI00, UHI02, Fen03]. Synchronous [GCYH01]. Synchrony [Kov02, YALT00, BEK^{+}03, FS06, MJ07b]. syndrome [CBB07]. synergists [BR08]. Synergy [HMND06, KDKS06]. synonymous [KOK06, Plu06, SWB06]. Synthase [BNT^{+}00, NM06]. Synthesis [AøP03, ANL01, HA00, Di 04, Jam08b, JEHK06, LFF06, LZGL03, MCMS06, SZ09]. synthetase [dFG08]. Synthetases [CDF00, KS04]. Synthetic [Tas02, CSW^{+}08, GGK05, Kam03]. syntrophic [KNT^{+}09, MKE^{+}09]. System [ADMZ02, Ano01b, Ano02a, ACK00, BE03, CLZZ02, CVH03, DFC^{+}02, FK01c, Kal00, KY02, Kra01a, Leh00, MV02a, MV02b, MV02c, OKTS02, Rec02, ST01, TMS00, AMC^{+}09, BRC07, Buc04, CLB05, CB07a, CRR08, CBF05, CD07, FMI05, GMK06, HK09, KLL07a, KA03a, LRD04, LKM06, LT08, MS07a, MDCC06, MD06, MC06, MCB07, MPL06b, PDM04, PDPL05, PCZL05, PCZL06, PR08, RW08, SH03, SYYI07, Sil09, SCS04b, SPE08, Tan07a, Tan07c, TG09a, VWR07, VGBA06, WTSN06, ZJ05, vAR09, vAGDR09]. System-theoretical [Kal00]. Systematic [FGH01, JPJ06, LP08, GABK08, PD04]. systematically [NA03]. Systemic [FP03, SLP00, Rǎd09]. Systems [Alb02, Cha01a, CMW02, FC01, GVK00, Jaz01, KT01, Kru02, LBF01, Mal00, Man01, MKLD02, Nak01, Nak03, PPR01, RC00, SCH02c, SPH03, TGP^{+}00, VW00, AC04, BGF07, BTL08, CCC08, CV08, CG03, DFSD05, EL09, FCP03, GV03, GMMGD^{+}08, GGH^{+}05, GN07, GCS07, GST08, HS04, Hor08, KM08, KBD06, KH08, LMM03, LSAA^{+}06, LP08, LFFT06, MB09, MHRK08, MvdO06, MPD^{+}07, MPN07, Pal08, PSPF07, QHF^{+}07, RG03, RS04, RF04b, RRR^{+}08, RSS04, Sac04a, Saf09, SV05, SD06a, SD06b, SHI05, SHI06b, SIK03, SI05, SA08a, SH07, TK05, TG09b, WSS07, Wil08a, WH07]. Szathmary [LL01b].

T [CLHW07, SRAL12, TG04, ACSY04, AWJ02, AB00, AB03, BdOP06, CSR^{+}05, CCP^{+}00, CDFP04, DdB03, FFD^{+}02, JAFW05, KSEK09, Koh07, Lan00a, LPLC00, LLC03, LGCL07, LHFH08, LVV^{+}01, MH07, ML04, Noe00, PD01, RWP^{+}08, SI09, SF04, SB05, SBPC05, SMPvdB08, VRB01, WSC02, WS02b, WJ01, YBV^{+}00, YCS04, vdBR04a, vdBWLS07, vLHH06]. T-bet [YCS04]. T-cell [AB00, AB03, FFD^{+}02, Lan00a, LLC03]. T-cell-Mediated [LPLC00]. T-cells [SB05]. T-tube [CCP^{+}00]. T. [ISC01]. T4 [RFH^{+}02]. tactics [AKdlPP06]. tag [Tan07a]. Tail [BG00b, Sac07]. Tailoring [Gar02]. Taken [Orr03, Jam08b, Ros05]. tale [PB03]. tales [Van06a]. talking [KYS06]. Talks [Ano09a]. tall [Den08, Kan05]. Tallgrass [BR00, GS02a]. Tamoxifen [TSC04]. Tandem [DBHS00, MKN02]. Tangled [CDHJ02]. tango [LHDvdM04]. Tape [SSA00]. Tapered [CCP^{+}00]. Target [Kut03, MLMW01, GABK08, MMUGD09, PWZ09, WL09]. targeting [MAC06]. Targets [WSC02]. Task [CC01c, DS00b, PG01a, TBS^{+}02, Wah02, JAHKH09]. task-specific [JAHKH09]. Taste [Mau02]. tat [IFN07, RON09]. Tautologous [Di 01b]. taxa [FT09b]. Taxon [OLS^{+}02]. TCA [NHTM09]. Teat [DSK02]. Tech [Ano01-33]. Technique [FGH01, GKNT09]. techniques [RDF03]. Telencephalic [DOT02]. Teleost [KMGV00]. telescoping [SH08b]. Telomerase [SHHD02, CY08]. telomerase-dimer [CY08]. Telomere [SK01, Ark05, GKB03, SGD04]. telomeres [DSVBW07]. telomeric [JM07]. Telomes [KRNKH00]. Temperature [ADMZ02, HRBL02, KI05b, LJW02, SRR08, Wel00a, WB06, GMK06, GB07, IB06, LMF03, LW07, MS03b, ROR05, RCS05, SZ09, THKU07, TPN07, TNP07]. temperature-compensated [THKU07]. temperatures [JAA^{+}07, JAJA07]. Template [ANL01, PER03]. Template-directed [ANL01]. Template-guided [PER03]. Templates [CB00, Möl01, Yag09]. Temporal [BKCR01, BHOR01, Kut05, WRF01, BGF07, BCRG04, FM04, Jac03, KUK07, Lie05, MGP^{+}08, MI08a, MI08b, PCS^{+}06, SGTF07, SCH02c, SWC^{+}08]. temporally [NS09a]. Tendency [SS01]. Tending [ORM03b]. tendinous [BL08]. tendon [LW08a]. Tensegrity [CLO^{+}02]. Tensile [KG02, ZCC01]. Tension [HVPN02, NB02]. tensor [Pie09b]. Term [HFH03, Ova02, AF09, BRC^{+}03, DI09, DHM01, DMQ04, For09, MZK08, Smi08d, UB01, vdBvdB09]. terminal [SV03]. termination [Bah07, Hau07, SWC^{+}08]. Termite [ORM03a, LB05, LBS06]. Termites [HYA02, MKN02, CSS06, LS08b]. terms [LSAA^{+}06]. Ternary [BNT^{+}00]. terrestrial [CM09, KSG03, MGC04]. Territoriality [LM01b]. Territory [BCV00, LS08b]. tessellation [BEF^{+}06]. Test [Gin00a, PDA^{+}00, RKL02, CSB^{+}09, PKS^{+}08, Zam03]. Testing [FHL^{+}06, Kai04, Mad00, PHP03, SS02, Cui07, KYZ^{+}08, SSM09]. testosterone [Gra07]. Tests [Di 03b, Hol06]. TGF [NS09b, TSC04]. TGF- [NS09b]. th [Ait08]. Th1 [YCS04]. Th2 [YCS04]. thalamic [OR05]. thalamocortical [CRR08, RR08, vAR09, vAGDR09]. thalamus [WR07b]. thalassemic [PBZ08]. Thaliana [MAB00, LMT05, Mei05]. Their [AGCLMM03, AWJ02, Kal00, AAL08, APL08, BP00, BIS^{+}07, BSJ04, CMB^{+}01, DLM08, Dru03, ES00, FDG02, HED06, HDW^{+}09, Jam09b, JE09, KCS^{+}06, Kok04, LS03, LTW06, MB09, MB06, MH05, MRA06, MNL^{+}07, NA03, PPRI08, PBB03, RAZ03, SLP00, WBH04, Wil08b, YAL04]. them [BR04, DBBW09, DBBW11]. Theophil [NDE06]. Theor [AGW^{+}08, AMD06, BZ05a, BBM04, BKE04, CSP^{+}09, DBBW11, EPJ^{+}11, FK03a, Gie06, GWM06, HTN04a, IvDHI08a, JCRJ07b, KCP09, Ker06, LH09, LGK^{+}12, Lei10, LVL08a, LW04, ML07a, ML12, MT06a, ML08b, MI08a, NI07a, NI04a, Paw09b, PCZL06, PA09b, RH09a, RMAI09, SHI06b, SI04c, SRAL12, SML08, TK09a, TQUN08, VGMM^{+}07a, VGMM^{+}08, TRM03a, TRM03b]. Theorem [YL00, Agu08, OP05, TG09b]. theorems [JR08]. Theoretic [CDDW02, BLRR08, DCL09, HB09, MSWH00, MGS09, RKF06]. Theoretical [Ano05h, AO05, BLZ07a, BSRH03, CPC^{+}00, CTS^{+}08, Fer09b, FGH01, GBD00, HS01, HH05b, LP00, LS07, MW01, MM03a, Nie06, Nie02, OI05, PBZ06, Paw07b, PBZ08, Paw09b, PH03, QASL08, SN06, SW08c, SP01, TT03, TOB08, Usa06, WH01, Arc07a, BM06, Bla04, BGF03, CSD04, CW08b, EHG03, GBD06, Hur06, JJEX09, Kal00, KH09b, LIK^{+}05, NMH07, NM06, Ots08, RKOS09, RAHO06, SWI07, Smi08d, THL03, Win06, WCLL08, MGAD09b]. Theories [PDA^{+}00, Arc07b]. Theory [AH00, AP01, Bea00, Cha01b, CV08, ETTV08, FK01c, GG01a, Gin00a, GAK^{+}06, HFY07, IST01, ISWT02, JL00, KK00a, PN00, RC00, Roy00, SLP00, WS03a, WP01, XFBC07, AK09, AW06, Apa09, BLL08, BdSFFDMC09, BHWB05, BD08a, CGK^{+}05, EH08, EAM07, FCP03, FZ07, Gam06, Gol08, GNH^{+}05, HE08a, HH04b, HC07b, Jäc07, KV05, KOT07, KV04, LN03, LN05a, Lan03, LZGL03, Luc05, MDD06, MH06b, MPN^{+}05, MV07, Muz05, OPN07, PT09, Pep04a, Pep04b, PA09a, QO08, RG03, Rob03, Rot07, SA05a, Sev06, SD06a, SD06b, SHI05, SHI06b, SM09b, TH06]. therapeutic [MAC06, RSC^{+}06, VHF08]. Therapy [PJ01, Wod01, DP04, FBM06, GSRC^{+}06, IZGG05, KT07, Ort06, PPD09, Smi08d, WBH04, ZT06, ZW07, dVG04, dVG06]. There [LBSS02, GKM^{+}00, GCC08, LW04]. Thermal [HD08, JLS01, Wel02, Bak07, Bat06, MGS08, NM09]. Thermodynamic [BBCQ04, LN03, LLCM01, Mic05, Pro03, Puj02, BS05b, GMK06, KRB05a, MB09, MB07, NGTB06]. thermodynamic-hemodynamic-pharmacokinetic [GMK06]. Thermodynamical [AH03b, BHHS01]. Thermodynamics [Alb02, Dem00a, Dem02, Smi08a, Smi08b, Smi08c, Kat03, LN05a]. Thermophile [Di 03b]. Thermophilic [Di 00b]. Thermoregulating [SB00]. Thermoregulation [SM01]. thermostable [VRAF06]. thermotactic [MGS08]. thermotaxis [NM09]. Thiol [FW00]. Thiol-disulfide [FW00]. Third [LN02, BP08, OBN07, Tak06, Ano01-33]. Thoma [BCRG04]. thorn [LY03a]. Thorns [LY01, LY03a]. thorny [LY03b]. Threat [Szá03]. Three [DOT02, GBG01, KTH^{+}00, PAH06, PAD00, SA07c, Van06a, WLFC08, BGE05a, BGE05b, Bye09, GKTN07, HTN04a, HTN04b, PDPL05, RR09, Sne03, TGV^{+}07, UCSZ07, ZAD07]. Three-Dimensional [GBG01, KTH^{+}00, WLFC08, HTN04a, HTN04b, TGV^{+}07, UCSZ07, ZAD07]. three-layer [BGE05b]. three-level [PDPL05]. Three-person [SA07c]. Threshold [OSBH02, BHC06, BB05, GWM05, GWM06, LL05, MLPJ09, Thi04, vdBR04a]. Thresholds [RPNH03, BdOP06, CSS08, FLG^{+}07, IMNT09, SSDM06]. Thrifty [Wel03]. Thrombin [MJW00]. thrombocytopenia [AM08]. Thrombocytopenic [Her00]. Throughput [WKL01, AC04]. Throwing [CC01c]. Thrush [DPV00]. thumb [HH04c, HH05b]. thylakoids [BRC^{+}03]. Thymine [MS01a, NTK02]. thymus [CLHW07]. thyroid [Leo07]. thyrotropin [Leo07]. thyrotropin-thyroid [Leo07]. tick [OBPH^{+}08, RPNH03]. tick-borne [RPNH03]. Tiling [ETTV08, Twa04, TH06]. TIM [XK07]. TIM-mediated [RCS05]. Time [ES00, GF02, HI00a, LJW02, PG01a, SZ01, YM05a, YM05b, ZCC01, AM04a, AFD^{+}06, BPC08, CY07, ELL04, GSRC^{+}06, GK04, GXG03, HCM^{+}07, HRSV06, HSF09, IPY07, Jam08b, KCV05, KRGH07, LKM06, Maz08, Mcn00, PWK03, Ref04, RSH^{+}06, RSG09, SV05, SG06, VCBV^{+}06, WE06, Wan06, XM09, YCC02, YM04, YD09]. time-delayed [RSG09]. Time-dependent [YM05a, YM05b, BPC08, SV05]. time-fractional [HCM^{+}07]. time-series [AFD^{+}06, GXG03]. time-varying [KCV05, PWK03]. Timely [FYX^{+}09]. Times [EMS02, TRM03c, BGK09, DBBW09, DBBW11, GBB08, HEH^{+}09, LCL07a, Ste04, TIN06, TRM03a, TRM03b, WVA05a]. Timing [CC01c, FB01, VP01, Ale09, RWCK08, SE07]. tiny [YOYT07]. tip [BGB08, KP06, TKM06]. tips [CKS04, SM06a]. Tissue [BSWM00, DNS00, FMP01, HHR01, KRNKH00, Laz03, MLMW01, PGLG01, TMI03a, WGH01, Wil06b, Agi04, Buc04, BOCF08, CPM^{+}09, Che06b, CE08, DBBW09, DBBW11, GC08, IZGG05, KC03, KSK08, KHE06, PC04, Rot04, SWC^{+}08, SJD^{+}09, SS05b, WBD^{+}09, ZTKH09]. tissue-engineering [WBD^{+}09]. Tissue-specific [MLMW01, Wil06b]. Tissues [ZM03, BDMR06, GMAH07, Lap03, MGL05]. Tit [IFN07, RON09]. Tit-for-tat [IFN07, RON09]. Titin [RSB09]. Titin-induced [RSB09]. Titration [MJW00]. TN [YSW09]. TNF [RWCK08]. together [Dal06]. toggle [DHYHR09]. Tolerance [ITM^{+}07, LLC03, MWB05, HBS07, Pep04a, Pep04b, ZAdlPLM07, ZCD05]. tolerance-stand [ZAdlPLM07]. tomography [BKM09, MKB03]. Too [MC01, LD09a]. took [AD07b, Di 05]. Tool [DVL^{+}00, OP05]. tools [BdSFFDMC09]. tooth [Osb08]. top [PT09]. top-down [PT09]. Topic [Ghi02]. Topological [KS01b, LY00, PP04, SMG^{+}03, WS03a, LS09, LZ09a, MGAR07, MLWL06, MGDM08, MGDBdM08, MMUGD09, PBMU^{+}09, SPAH06, YZHZ09]. Topology [Gru00, MY01, RWF01, SSB06, SSWF01, WRF01, AO03a, Amz04, CS06, GS06, GMdA^{+}07, JAV07, Kee04, MCF04]. TOPS [LW08b]. Tortuosity [GP00, Ben04b, TTN05, ZFVH05]. Total [GLK^{+}02, TE04]. tournaments [LS09]. toxic [BdSFFDMC09]. Toxicity [EM00, Bar05, CPG09, KLL08]. Toxin [CSM02, Cog07]. Toxin-producing [CSM02]. toxins [KJD05, MGL^{+}06b, SCS07]. Toxoplasma [KCP07, KCP09]. Trabecular [FK01b, FK03a, FH07, SB07]. Trace [Wal00]. Tracheid [AF02b]. Tracheid-level [AF02b]. trachomatis [WM04]. track [GNLEK07]. Tracking [FL09, VWR07, Cin06]. Tract [Kan01]. Traction [HS00c, HQP^{+}09, NOT04, YLCW06]. Tracts [LJW02, AAL08]. Trade [BH01, FS06, Ger09, PG01a, RGFP07, ALM09, BHWB05, HBWB08, KSB07, SH04, SH05, SA05a]. Trade-off [Ger09, PG01a, ALM09, BHWB05, HBWB08, KSB07, SH04, SH05, SA05a]. Trade-offs [BH01, RGFP07]. trading [CH07]. Traffic [LSD^{+}00, JSCN04]. trafficking [PW09b, Pic06]. Trail [KW00c, JR06]. trails [JSCN04, VTG^{+}06]. training [SdAC^{+}08]. Trait [KW00b, Pep00, CHR06, Cui07, CY09, HL05, LT09, LKW06, MM08a, SWN07]. Traits [CS02, DBG01, DPA03, TH00, AF09, Chu08, DHP06, DPA05, Jam04, Men07, MK09, MN07a]. trajectorial [SB07]. trajectorie [SB07]. Trajectories [BZ00, PE00, IS03, ILDP04, LDW09, UCSZ07]. trajectory [TIN07]. tramadol [DFP^{+}08]. trans [RM04]. trans-membrane [RM04]. transcellular [Gie06]. Transcription [KMP03, YYA09, Bah07, CZM09, IYGA08, PE04, RRK06, RG06, SZ04, Tan08, WOLS07, YMLK04]. transcription-translation [IYGA08]. Transcriptional [Fli05, LPMC^{+}06, TIM06, CXZ^{+}09, HJ05, LR07, SCS04a, vLBJK07]. transducer [THL03]. Transduction [KPS02, RWF01, SPC02, SNT03, WRF01, AFD^{+}06, Bak04, FB08, HVY^{+}07, HDNC04, MS03a, MCB07, NS05, NS09b, RM04, SYYI07, SPB06, SSB06, SM06b]. Transfer [FW01, Kan01, Kun00, MS01a, PDIS00, See00, Wil02, Alp05, DFW^{+}07, GMAH07, GHHR03, HKP07, KZ05, SFC^{+}09]. transfers [BS05b]. Transform [DVL^{+}00, QLHL09, ZLXY08]. transformation [KE09, PBZ06, PBZ08, SBGA04, Won05]. transformations [Pac09]. transforming [BKM09]. Transgenic [JL00, PGM00, LSMB^{+}03, Mar09b, MMV^{+}04, PGH^{+}04]. Transglutaminase [LGB02]. Transient [Kar03, WMS08a, CSON^{+}05, DSS08]. transients [LJ09b]. Transition [ACK00, DEM^{+}00b, LCL07b, MKT^{+}00, WP01, YN02, AO03b, CML08, CCV08, CLW^{+}08, DVC^{+}04, DB08, RBS05, Tlu07, WB06, Xie09, vdGMG^{+}09]. Transition-state [MKT^{+}00]. Transitions [TN01, Amz04, BL06b, CF09, DC09, JL06, SS06a, Wit03, ZZLT07]. transitive [NS03a]. Translation [BG00b, KJ02, Leh02, Nas01b, BB06b, GW06, Hau07, IYGA08, NN07, Tay06].

Translation/Replication [Leh02]. Translational [BG00b, Hau07, JPJ06, TG09b]. Translocation [AGCLMM03, BJD^{+}02, DO00]. Transmembrane [TM00]. Transmissible [CPC^{+}00]. Transmission [ACCC00, Car02a, CAHH06, HS00a, HB01, HLS01, NC02, AAEW09, AM09, BFG09, CHN08, Ded09, Fer09a, JMvdB09, KSY^{+}08, KGL09, MPM07, O’K05, PRT04, RPNH03, SH03, Sch03, TBB^{+}06, TBC^{+}08, XFBC07, Yam03]. transmitted [GOP09, JWB07, OBPH^{+}08]. Transmitter [HFGB02]. transmural [ZTKH09]. Transpiration [KRNKH00]. transplant [ØKRG04]. transplantation [DKLL05]. Transport [AGT^{+}01, DD02, GP00, GP01, GvO04, Hop02, Mcn00, MN01, NB02, RB02a, SP02, TH03a, TGP^{+}00, WPE03, Ban06, BTS08, CBF05, CTS06, FPM^{+}06, FWE06, GSNH08, HMN09, JTGP06, KRB05a, LT06, McN06, MS05a, MS06a, MS07b, MS08c, ML09b, PAH06, SSJ09, SCNP^{+}06, SRG^{+}03, Sin06, TKN07, Tho05, WC07a]. transport-related [CTS06, LT06, WC07a]. Transporters [Hop02, RBBH02, DB05, Pán08]. transporting [NM08]. Transposable [QA01]. transposition [Cud05]. transpositions [Kun05]. Transsexuals [GK00]. transverse [CMG06]. trauma [Che06b]. Traveling [LACL03, UMI09, MY09]. Travelling [DD02, HB02b, UL06]. treat [Ger09]. treated [PPD09]. Treatment [Gar02, RB02b, Sne03, TGP^{+}00, ADHM09, ABvdD^{+}08, BMD^{+}08, BRS^{+}09, DMO^{+}07, ECAV07, GMK06, GK06a, Kam03, KJJB07, KW07, LYZ08, MSK03, ØKRG04, PLSGG05, SHRR06, STK08, TSC04]. treatments [BSGT08, HSG05, KSO06]. Tree [AF02a, PKA02, SW03, SKN^{+}03, SMG^{+}03, AHCN07, BGE05a, BLL08, KTH09, MPL06a, Sac04a, SG07, SLH^{+}09, SSM09, SJ09, ZAdlPLM07]. Trees [Chu00, GBLV03, Had01, Kra01c, Kra02, SI00, AF09, BL06a, Cud05, FDR04, GCH^{+}07, LRHB09, MCF04, NI04c, PPRI08, SBZ^{+}08]. trehalose [Voi03]. tremor [HSG05, SHPDL03]. trends [ZKFW08]. tri [Bal04, PV08]. tri-frame [PV08]. tri-trophic [Bal04]. triads [dCZJ^{+}04]. Trichomes [SBH01]. Trifurcating [SMG^{+}03]. trigger [MPA^{+}08, dM09]. triggered [BPZ^{+}01, NBMS06]. Trimethylene [WP01]. trinucleotides [TFYY03, YSW09]. Tripedal [Kel01]. tripeptide [YDL06b]. triplet [Pat05]. Triplets [AIK00, NA03, WBSY06]. trisphosphate [GFW^{+}09, LGB03a, TLZ05, WMS08b]. Triterpene [Kal00]. tritrophic [KYS06]. Trivers [Kan05, Kan06, Kan07]. tRNA [CDF00, Che06a, Di 04, Di 06b, Di 08, Fol08, GBD06, KS04, SKR06a, SKR06b, YNL01, dFG08]. tRNAs [Dor03a, GS02c, JS07, NN07]. Trojan [CW07, GT06b]. Trophic [Har02, LBJE03, Bal04, HB02a, HB02b, LW07, MGM07, PDPL05, RYAI06, SH07]. Tropical [PPL^{+}00, PAA07]. trout [KRGH07]. true [TTT09, TKN07]. Truncated [PPR01]. trust [MM03b]. truthful [HH04a]. Trypanosoma [NO04, DLM08]. tryptophan [SZ04]. tube [CCP^{+}00]. Tuberculosis [ACCC00, ACCC02, BGMM08, CCM07, Ger09, GRH^{+}07, MK04, MSK03, SKI^{+}06, ZKFW08, RWCK08, SJGK04]. tubes [KGG08]. tubular [MCK07]. tubulin [Cib08]. tubulogenesis [NOT04]. tug [FR06]. Tumor [CAF03, DKS04, HFH03, Jac03, KTCD00, KTH^{+}00, MKLD02, PGLG01, AG06a, AD06b, CTB^{+}05, GT06a, GT08, GDD^{+}03, IMKN05, KLN^{+}09, KSN03, KM05, LS08a, LGCL07, ML07a, ML07b, MD06, MD03, MDD06, RSC^{+}06, SNA^{+}08, SRCDS08, VHF06, WBR08, WLFC08, ZAD07]. tumor-immune [MD06]. Tumorigenesis [MLMW01, PN02b, PC04, dT07b]. tumorigenic [KF04]. Tumors [IKS00, SA01, AM04b, AMD05, AMD06, BR06, CA09, GDC^{+}06, LGCL07, Lin07, MCK07, NL03a, OR05, SA05b, TCP05, dPGR06]. Tumour [BG00a, HS00c, KSB08, ASMM08, ABM03, AM06, AP09, BFGS07, BC06, DBBW09, DBBW11, DSU^{+}04, GA07, GA09, Lo07, MAC06, OBL04, PA09b, PA09c, PSJ04, Rej07, SGGM05, SGM08, WBR09]. tumour-associated [ASMM08]. tumour-induced [MAC06]. Tumours [MLMW01, KJJB07]. tuna [GNLEK07]. tuned [HHH06, MDCC06]. Tuning [DHYHR09, AJOK09, SGTF07, vdBR04a]. tunnel [LBS06]. tunneling [YNO09]. tunnelling [BGD^{+}06]. tunnels [LB05, LBS06]. Turbulence [LP00, LP01, Lew03, LB06, USTG09]. turbulence-initiated [USTG09]. Turbulent [Mcn00, MN01, McN06]. Turdus [DPV00]. Turelli [Sch02b]. turing [RM04, BGF07, Car02b, IK06, MSMKM06, SIK03, SI05]. Turing-Hopf [BGF07]. turing-type [RM04]. Turkey [NAV04]. Turn [JSA^{+}07, JHJ^{+}09b]. turning [BCV^{+}08, HNTHA07]. Turnover [FFD^{+}02, KDK02, BÅ03, KH09a, LFP^{+}05, PGMK^{+}03, SW08c, TGR^{+}00, Wod07, dT07b]. Turtles [Hay00, Sol01]. Tuscany [Ano02a]. twinfilin [MO07]. Two [Dus01, FY00, Fur02, Had01, KW00a, KK00a, LN02, MV02a, MV02b, MV02c, OLS^{+}02, PKA02, Rei02a, Sil02, Sol01, TG09a, VDV00, Zah00, ACR06, AO05, BH08, Bye09, CC08, CE08, CF09, Dam04, DBGM08, DBB09, DHW^{+}09, Gol07, HA09, HQP^{+}09, Iim07, JHJ^{+}09b, KNT^{+}09, KS06, KLL07b, KH08, LL08, LGCL07, LKM06, MM09, MBB08, MLWL06, MP09b, NL04, PB03, Rel04, SS06a, SB09a, SZLK^{+}05, SB06, SIK03, SI05, SX06, SG04, ST05, SCS07, Tak06, Tao04b, TYW05, TQUN07, TQUN08, TSN05, UW09, VWR07, VSP06, WTL08, WSS07, WLHB07, YD09, YKiP^{+}09, ZZLT07, vVH07]. Two-Component [Sil02, YKiP^{+}09]. two-deme [LL08]. two-dimensional [AO05, HQP^{+}09, Iim07, MBB08, SIK03, SI05]. Two-duct [MV02a, MV02b, MV02c]. two-gene [Tao04b, WSS07]. Two-locus [KW00a]. two-member [SS06a]. two-moment [UW09]. two-parameter [WTL08]. two-pathogen [VWR07]. two-person [MP09b]. two-phase [Rel04]. two-prey [SX06]. two-score [ST05]. two-sex [MM09]. Two-species [VDV00, CF09, Dam04, KS06, ZZLT07]. two-stage [JHJ^{+}09b]. Two-state [Zah00, MLWL06]. two-strategy [HA09]. Type [AB00, Hop02, LLW09, WS03a, AH03b, AJ08, BMT04, CC06a, GASA09, GABK08, KP06, KSEK09, LSS06, MKD^{+}05, NS09a, OAKC08, PBHS05, PCZL05, PCZL06, Raf02, RM04, TSRB08]. typeable [LKC07]. Types [Dus00, FLBB01, LMT02, AR08, CRJC04, ETH04, JSA^{+}07, JHJ^{+}09b, LVL08a, LVL08b, MS05a, PGLL07, RAK^{+}08, SYC06, WYXC05, WYC06, vVH07, SHF02]. typhimurium [AS00, Kee05]. typical [TBB^{+}06, TBC^{+}08]. Tyrannosaurus [HNTHA07]. Tyrosinase [CCGC02, Ril00]. tyrosine [HSL04].

U [DL08]. Uganda [CHCC^{+}04]. UK [TK09a, TK09b, TBB^{+}06, TBC^{+}08]. ulcerative [Agi04]. Ultimatum [PN01, Här07a, dSKL09]. ultradian [LKM06]. Ultrasonic [Ats01]. ultrasound [BLT03]. ultraviolet [BH04, BKP09]. unaffected [DDJ06]. unassuming [Rap08]. unavoidable [KCHP08]. Unbiased [BZ00]. uncertainties [VGBA06]. Uncertainty [HJ07, SP01, MHRK08, WA07]. Uncle [GK00]. uncorrelated [RBJ06]. Uncovering [RSSM06, TC09, WKL01]. Underdominance [DBG01]. underlie [MCM^{+}09]. Underlying [KMP03, SSWF01, CY09, JB04, KYZ^{+}08, KI04, KI05a, Luc05]. underpinning [GK04]. understand [Kau04]. Understanding [AM08, GG01c, MSH05, PJ01, PD00, QO08, XBCF05, BdSFFDMC09, KZ05, RG06, RRS06]. understory [Che07]. unequal [YB05]. Unexpected [BCL08, KCHP08]. unfocused [Has06]. unfolding [HQP08]. unfortunate [SB07]. Unhealthy [Wil08a]. uniaxial [Hol06]. Unified [Bro02, Chu00, CS02, GVB^{+}08, MV07, NM09]. unifies [DP08]. Uniform [SH03, SKN^{+}03]. uniformly [HH04b, TN04]. Unifying [BLL08, OR04, PN02a, ROR05, Lan03]. uninfested [KYS06]. uniquely [VL09]. uniqueness [MGL^{+}06b, OB08]. uniqueness-self-interest [OB08]. unisexual [HI04]. Unit [GG02, MPM07, Nos06]. Unit-restricted [GG02]. United [Tho00]. Unity [WS03a]. Universal [Cha01a, Di 00b, Di 01a, Di 03b, FH07, CML08, Di 06b, GDD^{+}03, GDC^{+}06]. universality [Di 01a, Kom04]. Universities [CMW02]. unlikely [Ped07]. unmineralized [FS04b]. unpalatable [SSR04]. unperturbed [DTG09]. Unpredictability [Has06]. unpredictable [SHI03]. unravel [FMHW08]. unreplicated [ZAB^{+}09]. Unsaturated [AF02a]. unscrambling [PER03]. Unstable [WS02a, SRCDS08]. unsteady [AO05, JA05]. unstructured [Sch05a]. Unusual [Dor03b, SKR06a]. unveils [HKP07]. unwrapping [Aba09]. updated [PGF^{+}08]. updating [CA09]. upon [PDA^{+}00]. Uptake [RBBH02, Bio08, CBB07, CRL06, DBR^{+}05, KLL07b, LWC09, MW09, RF04a, RF04b, RF04c, Smi08d]. upward [SW04b]. urban [AKS07]. urchin [Zac09]. ureteric [HIM09]. Usage [AS00, RB02b, AAEW09, CSW^{+}08, NGT05, Plu06, PB06, SWB06]. Use [DW08a, GKNT09, Möl01, PKL01, PH03, BRS^{+}09, CPG09, GT06b, NA03, OBL04, Pac09, PGM00, SLP00, YTAK01, vV05]. used [FGH01, Pal08, WFGP04]. Using [AF02b, BFH^{+}01, BE08, CFLC06, CT02, CTZ^{+}06, DBBW09, DBBW11, DOT02, DNS00, ELSFB07, GLE^{+}09, HAG03, HGB^{+}00, KTH^{+}00, LY02, LBJE03, PC02, PAD00, QLHL09, DPV00, SMG^{+}03, TS02, WYC06, hZzGqX^{+}09, ZCLZ07, AMW00, AIKP06, AKdlPP06, BL08, BG08, BLZ07a, BLZ07b, BMR08, CL07a, CC09a, DKS04, Di 01b, DL08, Dus06, ELJ06, HCM^{+}07, HGP^{+}07, HBB08, IB06, IvDH^{+}08, JSA^{+}07, JHJ^{+}09b, JWWS08, Kau04, KmMK04, LK08a, LLS^{+}09, Lin08, Mad00, MGS08, MMUGD09, NBT07, NAS07, PLH05, PM08, PB06, QWQ09, RCA09, RBW09, RA08, RW08, RRR04, SGTF07, SH04, SH05, Sch08b, SJGK04, SD06a, SD06b, SLHN06, THL03, TC09, TR08, VCGV07, WS02a, XWC08, YTL08, YYA09, ZLLZ09, ZW07]. Ussing [Luc05]. Usually [MC01]. utero [CSB^{+}09]. utility [Dic08]. UV [BH04, Möl02]. UV-Green [Möl02].

V [Ano04b, NM08, SPH03, RA06]. V-formations [SPH03]. V/K [RA06]. Vaccination [AWJ02, GG03a, Ste02a, ABvdD^{+}08, Ger09, GWM04, HDZ^{+}07, JWB^{+}09, KMW03, LTI08, NDE06, PLH05, TY06, Ten08, TP05, ZT08]. vaccine [ITLN09, ITI^{+}09, LKC07, RAHO06]. vaccine-resistance [ITLN09]. vaccines [Ano05h, BGMM08, Hof04, JAFW05, LGS^{+}09]. valid [TE04]. Validation [KS01c, DSU^{+}04, JAHKH09, Mag04]. validity [OP05, Pie09b, TE05]. valleys [GINT09]. Value [For01, LM01b, CWDM06, ELSFB07, FP03, Hur06]. valued [DHP06, Men07]. Values [WP01]. Varanus [See00]. Variability [Bon03, PKL01, RTK02, RDYH09, DPV00, TTKZ01, CW08b, GOP09, GS08b, HJ06, LPJB^{+}08, PSM06, WMS08a]. Variable [AGT^{+}01, Hay00, SM01, SR02, LBS06, SG04]. variables [Cud05, KHE06, PBB03]. Variance [MvdO06, BK05, BW06, LFFT06, MJ07a, NI04c, WTL08]. variance/covariance [LFFT06]. Variances [Cha01b, Jam09a]. Variant [NI04c, RG05]. Variation [CACC02, Gin00a, Hok00, JJ01, TTR00, BEK^{+}03, CS07, EBId09, GI09, HW05, HW06, HL06, JK03, Jam06, KF06, LS09, LKK07, VTL05, WVA05b, YZHZ09]. Variational [RH09a, RH09b]. Variations [CPR00, BG06, Bog04, Jam08b, Jam09b, TT04]. Varicosities [HFGB02, BFGB04, BFG05]. varied [LMT05]. Various [Kan01, MA03b, NA02, GEK04, KSY^{+}08, MS05a, SA07b, TY06]. varius [See00]. Varying [CLZZ02, GASA09, HP04, KCV05, NM08, PWK03, TOB08]. Vascular [Bar01, BR06, CCP^{+}00, GL01, HFGB02, Kra02, WBR09, Zam01, AM04b, BGE05a, FMI05, FI06, KRN03, NMH07, PA09a, PSJ04, PBB03, WBR08]. Vascularity [PGLG01]. vasculature [GT06a, Pie09a]. Vasomotion [GP01, PC09b]. Vector [CLXC03, KA02, MK03, AS09, dCZJ^{+}04, CTZ^{+}06, CBC^{+}09, DLM08, DRLB05, HRvdD08, ITI^{+}09, JSA^{+}07, MBBR06, QLHL09, RBW09, YCC^{+}06, ZLLZ09, ZCLZ07]. vector-borne [CBC^{+}09, HRvdD08]. vectored [DBBC08]. Vectors [LMT02, VSP06, WMP03]. Vegetation [RDL07, KvHMP09, LBCL09, LJL08, MS08a, MBRRI08]. Vegetative [ZM03]. VEGF [KS08a, SNA^{+}08]. VEGF-A [KS08a]. VEGF-C [KS08a]. VEGFR [KS08a]. VEGFR-1-binding [KS08a]. vehicles [OBL04]. veil [Chi07]. veil-line [Chi07]. veins [FM06]. Velcro [AJ08]. velocity [Alb08, DECEK06]. venation [CABB09]. Venous [WGH01, WBR09].

Ventilation [WGH00]. ventral [AGZ^{+}06, ZLN07]. ventricles [BBSLN06, BBSLN08]. ventricular [BBSLN08, BOCF08, Han04, Hop06, LAG07, OD03, ZTKH09]. vents [BBKGP08]. Verhulst [Nås01a]. verification [CTS^{+}08]. vermiform [BBB^{+}07]. versa [CB07b, MI08a, MI08b]. versus [Agi04, Gie03, GFWT04, Mat06, Niw03, SI05, WT09b, YHI07]. vertebrate [RH04, SNCM09, Sar04, UMI09]. vertex [HTN04a, HTN04b]. Vertical [FF02, HS01, YN02, MPN07, SH03, Yam03]. vertically [JWB07]. Vesicle [GBG01, BD08a, FSS06a, MS07a, Rot07, TKM06]. vesicles [MS07b]. Vessel [Dor03b, RMBM00, BR06, WBR09]. vessels [AM06, BFGB04, BFG05]. Via [GG03b, BR06, CC01a, FFTS09, Gag00, JLCS08, JP00, KB02, KGG07, LBS00, LSH06, PSPF07, SHRR06, TH06, WWS^{+}06, WS09]. Viability [NS03b, Tan07b]. vibrations [PT09]. Vibrio [KH08]. vice [CB07b, MI08a, MI08b]. vicinity [SNA^{+}08]. View [Nak01, Nak03, Nir02, Che06a, Lan03, LV08, NM09, SZC^{+}03]. viewed [GMMGDA05]. Viewing [LSH06]. Viewpoint [KRN01, GMMGD^{+}08]. Vigilance [PBR03, KHL07, Lin04, PBR01, RB08]. vigour [Arc09a]. violate [ZXWF08]. violations [Dic08]. Violaxanthin [LBS00]. Violent [Kan06]. viraemic [RPNH03]. Viral [ETTV08, Kom07, OSK^{+}05, SYSY02, ALM09, BST05, Bea06, Bel06, BB06a, BRND09, CMF08, DHRM08, DP04, GLE^{+}09, GB07, GCP04, HM05, KMT06, KTE08, NS09a, PPD09, RGFP07, RP09b, SH08a, STK08, TH06, WL07a, WJ01]. virion [GCP04]. virology [Twa04]. virotherapy [BCJ^{+}08]. virtual [SDL^{+}08]. Virulence [Fer09a, HS00a, AG06b, KB04, MCB07, O’K05, PLSGG05, PSWK09, SRN07, SK09b]. Virus [AB00, FPL03, IMN05, LN02, MY09, SCC^{+}00, BMD^{+}08, CRNP07, DLRP07, DBBC08, DHW^{+}09, FJBK05, GLE^{+}09, HI07, HHP05, HSC07, HM05, JMvdB09, KNWCB07, MW09, NS09a, PBHS05, PT09, SDL^{+}08, SCS04b, Twa04, XSD^{+}05]. virus-immune [FJBK05]. virus-vectored [DBBC08]. Viruses [CMB^{+}01, For02, Fra00b, JL00, BCJ^{+}08, Kom07, PW09a, RAZ03, SSDM06]. Visceral [Rec02, MA07b, Sil09]. Viscoelastic [HS00c, KS09]. Viscoelasticity [CLO^{+}02]. Viscosity [GD06, MW00, KNC^{+}04]. viscous [GA08b]. Vision [How09, CS08, HHH06]. visual [CS09, HLP06, How09, KWGE04, RP09a]. visualization [GDPMDS^{+}09]. Visualizing [DVL^{+}00]. vital [BSJ04]. Vitamin [EM00]. Vitro [SHHD02, AHT^{+}07, BCKE^{+}08, BMD^{+}08, DBBW09, DBBW11, FVP^{+}07, HM06, KA03a, NOT04, PA09b, PA09c, Ros03, ZW07]. vivo [DBBW09, DBBW11, DSU^{+}04, NB02, OLBM08]. vocabulary [Smi04]. Vocal [DPV00]. Voids [PPGS03]. Voilà [Noe00]. volatile [BKD^{+}06, ZTKH09]. Voltage [GJE02, SFC^{+}09, DBB09, Gre05, LSS06]. Voltage-Biased [GJE02]. voltage-dependent [DBB09]. Volterra [BM09, CG03, MC06, PDC02]. Volume [Ano03a, Ano03b, Ano03c, Ano03-44, Ano03-45, Ano03-46, YFKP03, AW06, Ano04h, Ano04i, Ano04-61, Ano04-62, BFG08a, GKTN07, HEH^{+}09, LCHB06].

Volumetric [DOT02]. volunteer [Arc09b]. vortex [JB04]. vortex-like [JB04]. Voting [BR09]. VRI [XK07]. VRI/PDP1 [XK07]. vs [RTK02, RWP^{+}08, SS03, SI02, SYSY02, Wod01, ZSS02]. vulnerability [Abr09, KJJ07]. vulnerable [CE05].

W [EK08]. waddling [KSPA^{+}08]. waist [CTB09]. waist-to-hip [CTB09]. Waiting [HSF09]. waking [RRR04]. Walk [FK03b, SFV02, AH03b, Akt04, BZ04, MLWL06, RBS05]. walk-to-run [RBS05]. Walking [BR01, FT07, Kel01, LW08a, RBS05]. Walks [Orr03, BNRW04, CH05b, PB06, RB09a, RBJ06, Ros05]. Wall [Dor03b, DWBB04]. Walls [Dor03b, Hol06, WL03]. Waning [FMLP06, GG03a, FGMP08]. War [ELL04, HG02, FR06, HGC03, vVH07]. Warburg [AP09]. ward [PFdP^{+}07]. warfare [SCS07]. warning [LY01, PBR01]. wars [Jam09b]. was [Di 01a, Di 03a, Di 03b, WD03]. wasp [BIS^{+}07]. Wasps [IST01, KB02]. Water [AF02a, AF02b, BHHS01, DLGC02, HVPN02, Kan01, YN02, Bio08, HML04, LN03, RF04a, RF04b, RF04c, WB06]. Watermelon [Kor07]. WATOR [DL00]. Wave [ACK00, AMV^{+}02, Sch02b, DSS08, FWLN04, GSM06, KS06, LBSP06, PGN08, SLHN06, SM06b, UMI09]. Wavelength [PH03]. Wavelet [DVL^{+}00, QLHL09, RAK^{+}08]. wavenumber [OR04]. Waves [ACK00, DD02, HB02b, GG09, LACL03, MY09, MPA^{+}08, RDC09, SI06, Tho05, UL06]. way [KJJ07, LSH06, ZJG03]. WBE [PA09a]. Weak [Arc09a, HB01, MGM07, CE05, LJ09a, RBJ06, WT07]. Weakly [GEF09]. Weapon [LY03a]. weaving [VCBV^{+}06]. Web [DHM01, Bal04, BRCB04, BLMV05b, DI09, DMQ04, GDS07, II06, KJJ07, KGD09, Kon06, LM08, Paw09a, PM08, PB09, RMAI06b, KM08, VCBV^{+}06]. Web-building [VCBV^{+}06]. webs [AP04, BDBR07, CSA07, Est07, FV09, GD08, MB07, RMAI06a, RMAI09, UD07, Van06b, CC01b, Har02, JSV02, LBJE03, MS02]. weevil [FHL^{+}06]. Weight [Jam01b, Mar03, TTR00, Wel02, AF09, CSS08, Hel08, Jam08c]. Well [LLCM01, Bea06, SH08b]. well-mixed [Bea06]. West [BDMR06, Cle07, MY09, MGL05]. whale [SPVN06]. Wheat [FMP01]. whether [BE08]. which [AGT^{+}01, Buc04, HFY07, PDA^{+}00, SML07, SML08, Tak06, Uit09]. Whirled [LY00]. white [VL09]. Who [AB04]. Whole [SML07, SML08, MSL04]. Whole-cell [SML07, SML08, MSL04]. Wholemount [Che00]. whom [AB04]. whose [RGG00]. wide [AA04, Cui07, GGK05, QHF^{+}07, ZSZ^{+}06]. width [KSPA^{+}08]. wild [BMT04, FGMP08, GT06b, Raf02, TNTJ08]. wild-type [BMT04, Raf02]. Will [CR01, NP05]. Willard [Kan05, Kan06, Kan07]. Wilson [MBB02]. win [IFN07]. win-stay [IFN07]. Wind [GA02, USTG09, WH00, MT06a, MT06b, MA03b, SS03]. wind-dispersal [MA03b]. wind-pollinated [MT06a, MT06b, SS03]. wind-pollination [MT06a, MT06b, USTG09]. Winds [WH00]. Winfree [Ost04, Rot04, TG04, Win04]. wing [BZM05, KKPB09, RWID^{+}08, SM06a]. wings [AO05, Sac05]. Wins [HS02a]. wise [APS08, AS09]. Within [AG06b, CPR00, FW01, Fra00b, LN02, SCH02c, ETH04, EW09, Fei08, FB06, GLE^{+}09, GCP04, IMN05, Jam08b, Jam09a, MB07, MKB03, NM06, PCB07, PMB06, RDF03, RGFP07, RDSB^{+}03, SNCM09, SSLB07, TBC^{+}08, UHK01, YZHZ09]. Within- [SCH02c]. within-herd [GLE^{+}09]. Within-host [AG06b, Fra00b, GCP04, RDF03, RGFP07]. Within-species [CPR00]. without [BLS^{+}09, KGD09, Par04]. Wnt [GP08a, RGSFM07]. Wolbachia [TYW05, VCGV07, ETH04, Sch02b]. Wolbachia-types [ETH04]. women [Hel08, Jam08c]. Wood [LY00, AF09, AJ08]. Word [Now00]. Work [KB02, RBS05, VA06]. Worker [SFV02, PRT04]. workers [BIS^{+}07, RHF07, Tof06]. working [PP08]. workspace [Wal07a]. World [DKD02, Leh02, MS02, Tho00, DS04, DLB07, DLM03, SK05, Tay06, VdGN^{+}05]. worlds [OBDJ06]. would [SS06d]. Wound [CS00, CLH07]. wound-healing [CLH07]. wrapping [Aba09, LDW05]. wrapping/unwrapping [Aba09]. WRCH [KS08b]. Wright [Wax09]. Wrightian [Win06]. wrinkling [CML08]. wrong [SG07, vV09]. wrongs [HMGK06].

X [How09, CS08, SP06]. X-linked [SP06]. X-ray [How09, CS08]. Xanthophyll [LBS00]. Xenopus [FWLN04, MKL09]. xuthus [OST09]. Xylella [VRAF06]. Xylem [FGH01, HVN07, HMN09, LN03, LMVPM07, SWRH03]. xylem-phloem [HMN09].

Y-linked [GMM09]. yaw [Sac07]. yawing [Sac05, SM06a]. year [YHI04]. Yeast [RvMK^{+}05, AVSHV04, BS05a, BB07a, GPB^{+}04, HK05, Hen04, Iro09, LR07, MCN08, RR09, SSK06, iTYU06, Voi03, WRG^{+}04, ZPP08]. yellow [Che07]. Yield [HCMF01, HCB^{+}02, BDR08, SPAH06, SPF08]. yields [ACK08]. young [PBB03, EKIL01, HYA02]. YY [GT06b].

Zealand [KMW03, TBR08]. Zeaxanthin [LBS00, DVC^{+}04]. Zebrafish [Moc02, TMI03b, ZLN07]. zebras [LY03b]. Zeeman [JDMZ^{+}07]. Zero [CY09, HNO03, Tho00, Che06b, EAM07, GR05, HE08a]. zero-energy-cost [GR05]. Zeroinflated [CY09]. Zero-Knowledge [HNO03]. zero-net-flux [Che06b]. zerosum [EAM07, HE08a]. zeros [CY09]. Zhabotinskii [SRR08]. Zigzagging [LS07]. Zipf [LY02]. zone [KS06]. zones [BFGD07, LL09, SPVN06]. zooplankton [BdSFFDMC09, LSH06, MPN07, SCS04b]. zymogens [VGMVR^{+}06].

References

Arocena:2004:NCM

[AA04] Miguel Arocena and Luis Acerenza. Necessary conditions for a minimal model of receptor to show adaptive response over a wide range of levels of stimulus. Journal of Theoretical Biology, 229(1):45–57, July 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001031.

Abatih:2009:IAU

[AAEW09] Emmanuel N. Abatih, Lis Alban, Annette K. Ersbøll, and Danilo M. Lo Fo Wong. Impact of antimicrobial usage on the transmission dynamics of antimicrobial resistant bacteria among pigs. Journal of Theoretical Biology, 256(4):561–573, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005444.

Arbeev:2009:GML

[AAK^{+}09] Konstantin G. Arbeev, Igor Akushevich, Alexander M. Kulminski, Liubov S. Arbeeva, Lucy Akushevich, Svetlana V. Ukraintseva, Irina V. Culminskaya, and Anatoli I. Yashin. Genetic model for longitudinal studies of aging, health, and longevity and its potential application to incomplete data. Journal of Theoretical Biology, 258(1):103–111, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000435.

Ababneh:2008:DTB

[AAL08] Anas M. Ababneh, Zaid Q. Ababneh, and Cathy C. Large. DNA a-tracts bending: Polarization effects on electrostatic interactions across their minor groove. Journal of Theoretical Biology, 252(4):742–749, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000878.

Asquith:2000:RCL

[AB00] Becca Asquith and Charles R. M. Bangham. The role of cytotoxic T lymphocytes in human T-cell lymphotropic virus type 1 infection. Journal of Theoretical Biology, 207(1):65–79, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921568.

Asquith:2003:DCF

[AB03] Becca Asquith and Charles R. M. Bangham. The dynamics of T-cell fratricide: application of a robust approach to mathematical modelling in immunology. Journal of Theoretical Biology, 222(1):53–69, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000134.

Allesina:2004:WDW

[AB04] Stefano Allesina and Antonio Bodini. Who dominates whom in the ecosystem? Energy flow bottlenecks and cascading extinctions. Journal of Theoretical Biology, 230(3):351–358, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002620.

Adams:2006:MRB

[AB06] Ben Adams and Michael Boots. Modelling the relationship between antibody-dependent enhancement and immunological distance with application to dengue. Journal of Theoretical Biology, 242(2):337–346, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600107X.

Adams:2007:IIC

[AB07a] Ben Adams and Michael Boots. The influence of immune cross-reaction on phase structure in resonant solutions of a multi-strain seasonal SIR model. Journal of Theoretical Biology, 248(1):202–211, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002160.

Aylward:2007:IIA

[AB07b] Nigel Aylward and N. Bofinger. Intramolecular interactions in aminoacyl cyclic-3’,5’-nucleotides. Journal of Theoretical Biology, 244(2):258–274, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003341.

Ababneh:2009:RPI

[Aba09] Anas M. Ababneh. The role of polarization interactions in the wrapping/unwrapping of nucleosomal DNA around the histone octamer: Implications to gene regulation. Journal of Theoretical Biology, 258(2):229–239, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900071X.

Aldana:2007:REG

[ABKR07] Maximino Aldana, Enrique Balleza, Stuart Kauffman, and Osbaldo Resendiz. Robustness and evolvability in genetic regulatory networks. Journal of Theoretical Biology, 245(3):433–448, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005170.

Alarcon:2003:CAM

[ABM03] T. Alarcón, H. M. Byrne, and P. K. Maini. A cellular automaton model for tumour growth in inhomogeneous environment. Journal of Theoretical Biology, 225(2):257–274, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002443.

Alarcon:2004:MME

[ABM04] T. Alarcón, H. M. Byrne, and P. K. Maini. A mathematical model of the effects of hypoxia on the cell-cycle of normal and cancer cells. Journal of Theoretical Biology, 229(3):395–411, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001493.

Alcaraz:2000:AAF

[ABP^{+}00] L. E. Alcaráz, S. E. Blanco, O. N. Puig, F. Tomás, and F. H. Ferretti. Antibacterial activity of flavonoids against methicillin-resistant staphylococcus aureus strains. Journal of Theoretical Biology, 205(2):231–240, July 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920629.

Antia:2003:MCR

[ABP^{+}03] Rustom Antia, Carl T. Bergstrom, Sergei S. Pilyugin, Susan M. Kaech, and Rafi Ahmed. Models of CD8+ responses: 1. What is the antigen-independent proliferation program. Journal of Theoretical Biology, 221(4):585–598, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303932085.

Anissimov:2002:CEO

[ABR02] Yuri G. Anissimov, Anthony J. Bracken, and Michael S. Roberts. Catheter effects in organ perfusion experiments. Journal of Theoretical Biology, 214(2):263–273, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924506.

Abrams:2009:ACP

[Abr09] Peter A. Abrams. Adaptive changes in prey vulnerability shape the response of predator populations to mortality. Journal of Theoretical Biology, 261(2):294–304, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003403.

Antal:2007:AIC

[ABTR07] T. Antal, K. B. Blagoev, S. A. Trugman, and S. Redner. Aging and immortality in a cell proliferation model. Journal of Theoretical Biology, 248(3):411–417, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002858.

Arino:2008:MIV

[ABvdD^{+}08] Julien Arino, Fred Brauer, P. van den Driessche, James Watmough, and Jianhong Wu. A model for influenza with vaccination and antiviral treatment. Journal of Theoretical Biology, 253(1):118–130, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000908.

Aach:2004:MMD

[AC04] John Aach and George M. Church. Mathematical models of diffusion-constrained polymerase chain reactions: basis of high-throughput nucleic acid assays and simple self-organizing systems. Journal of Theoretical Biology, 228(1):31–46, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004557.

Abedon:2007:BEG

[AC07a] Stephen T. Abedon and Rachel R. Culler. Bacteriophage evolution given spatial constraint. Journal of Theoretical Biology, 248(1):111–119, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000872.

Abedon:2007:OBP

[AC07b] Stephen T. Abedon and Rachel R. Culler. Optimizing bacteriophage plaque fecundity. Journal of Theoretical Biology, 249(3):582–592, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003815.

Aparicio:2000:TDT

[ACCC00] Juan P. Aparicio, Angel F. Capurro, and Carlos Castillo-Chavez. Transmission and dynamics of tuberculosis on generalized households. Journal of Theoretical Biology, 206 (3):327–341, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300921295.

Aparicio:2002:MDE

[ACCC02] Juan P. Aparicio, Angel F. Capurro, and Carlos Castillo-Chavez. Markers of disease evolution: The case of tuberculosis. Journal of Theoretical Biology, 215(2):227–237, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924890.

Accoto:2004:BAO

[ACD04] Dino Accoto, Piero Castrataro, and Paolo Dario. Biomechanical analysis of oligochaeta crawling. Journal of Theoretical Biology, 230(1):49–55, September 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001821.

Acerenza:2000:DLM

[Ace00] Luis Acerenza. Design of large metabolic responses. Constraints and sensitivity analysis. Journal of Theoretical Biology, 207(2):265–282, November 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921738.

Argentina:2000:HCW

[ACK00] M. Argentina, P. Coullet, and V. Krinsky. Head-on collisions of waves in an excitable FitzHugh–Nagumo system: a transition from wave annihilation to classical wave behavior. Journal of Theoretical Biology, 205(1):47–52, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920447.

Ackland:2004:MPD

[Ack04] G. J. Ackland. Maximization principles and daisyworld. Journal of Theoretical Biology, 227(1):121–128, March 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300393X.

Andrecut:2008:MCS

[ACK08] M. Andrecut, D. Cloud, and S. A. Kauffman. Monte Carlo simulation of a simple gene network yields new evolutionary insights. Journal of Theoretical Biology, 250(3):468–474, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005425.

Allstadt:2009:PSC

[ACK09] Andrew Allstadt, Thomas Caraco, and G. Korniss. Preemptive spatial competition under a reproduction-mortality constraint. Journal of Theoretical Biology, 258(4):537–549, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000769.

Ackland:2003:CDF

[ACL03] Graeme J. Ackland, Michael A. Clark, and Timothy M. Lenton. Catastrophic desert formation in daisyworld. Journal of Theoretical Biology, 223(1):39–44, July 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000699.

Aslanidi:2005:DCE

[ACLH05] O. V. Aslanidi, R. H. Clayton, J. L. Lambert, and A. V. Holden. Dynamical and cellular electrophysiological mechanisms of ECG changes during ischaemia. Journal of Theoretical Biology, 237(4):369–381, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001864.

Adams:2003:ODD

[ACLW03] B. Adams, J. Carr, T. M. Lenton, and A. White. Onedimensional daisyworld: spatial interactions and pattern formation. Journal of Theoretical Biology, 223(4):505–513, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001395.

Adimy:2006:POL

[ACR06] Mostafa Adimy, Fabien Crauste, and Shigui Ruan. Periodic oscillations in leukopoiesis models with two delays. Journal of Theoretical Biology, 242(2):288–299, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000920.

Allan:2004:CAI

[ACSY04] Martin J. Allan, Robin Callard, Jaroslav Stark, and Andrew Yates. Comparing antigen-independent mechanisms of T cell regulation. Journal of Theoretical Biology, 228(1):81–95, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004582.

Alves:2006:MSP

[AD06a] Filipa Alves and Rui Dilão. Modeling segmental patterning in drosophila: Maternal and gap genes. Journal of Theoretical Biology, 241(2):342–359, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005254.

Athale:2006:EER

[AD06b] Chaitanya A. Athale and Thomas S. Deisboeck. The effects of EGF-receptor density on multiscale tumor growth patterns. Journal of Theoretical Biology, 238(4):771–779, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002870.

Andre:2007:PRO

[AD07a] Jean-Baptiste André and Troy Day. Perfect reciprocity is the only evolutionarily stable strategy in the continuous iterated prisoner’s dilemma. Journal of Theoretical Biology, 247(1):11– 22, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000732.

Archetti:2007:EGC

[AD07b] Marco Archetti and Massimo Di Giulio. The evolution of the genetic code took place in an anaerobic environment. Journal of Theoretical Biology, 245(1):169–174, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004358.

Alexander:2009:CEM

[ADHM09] Murray E. Alexander, Sarah M. Dietrich, Yi Hua, and Seyed M. Moghadas. A comparative evaluation of modelling strategies for the effect of treatment and host interactions on the spread of drug resistance. Journal of Theoretical Biology, 259(2):253–263, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001362.

Andjus:2002:KPE

[ADMZ02] Radoslav K. Andjus, Zeljko Dzakula, Marina Marjanović, and Dragoslava Zivadinović. Kinetic properties of the enzymesubstrate system: a basis for immediate temperature compensation. Journal of Theoretical Biology, 217(1):33–46, July 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930201.

Aoki:2001:PEI

[AF01] Kazuhiro Aoki and Mitsuru Furusawa. Promotion of evolution by intracellular coexistence of mutator and normal DNA polymerases. Journal of Theoretical Biology, 209(2):213–222, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922574.

Aumann:2002:MTW

[AF02a] Craig A. Aumann and E. David Ford. Modeling tree water flow as an unsaturated flow through a porous medium. Journal of Theoretical Biology, 219(4):415–429, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930614.

Aumann:2002:PMD

[AF02b] Craig A. Aumann and E. David Ford. Parameterizing a model of Douglas fir water flow using a tracheid-level model. Journal of Theoretical Biology, 219(4):431–462, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930626.

Almeras:2009:BDL

[AF09] T. Alméras and M. Fournier. Biomechanical design and longterm stability of trees: Morphological and wood traits involved in the balance between weight increase and the gravitropic reaction. Journal of Theoretical Biology, 256(3):370–381, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005389.

Allen:2006:ADM

[AFD^{+}06] Edward E. Allen, Jacquelyn S. Fetrow, Larry W. Daniel, Stan J. Thomas, and David J. John. Algebraic dependency models of protein signal transduction networks from timeseries data. Journal of Theoretical Biology, 238(2):317–330, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002122.

Ahnert:2008:HMN

[AFZ08] Sebastian E. Ahnert, Thomas M. A. Fink, and Andrei Zinovyev. How much non-coding DNA do eukaryotes require? Journal of Theoretical Biology, 252(4):587–592, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930800057X.

Albano:2006:SMT

[AG06a] G. Albano and V. Giorno. A stochastic model in tumor growth. Journal of Theoretical Biology, 242(2):329–336, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000956.

Andre:2006:WHE

[AG06b] Jean-Baptiste André and Bernard Godelle. Within-host evolution and virulence in microparasites. Journal of Theoretical Biology, 241(2):402–409, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005291.

Aldana-Gonzalez:2003:TPP

[AGCLMM03] M. Aldana-González, G. Cocho, H. Larralde, and G. Martínez-Mekler. Translocation properties of primitive molecular machines and their relevance to the structure of the genetic code. Journal of Theoretical Biology, 220(1):27–45, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303931080.

Agius:2004:PDI

[Agi04] Lawrence M. Agius. A primary dysregulation in the immunoregulatory role of the intestinal mucosal epithelial cell in inflammatory bowel disease pathogenesis?: Biology of inflammatory response as tissue pattern entities in Crohn’s versus ulcerative colitis. Journal of Theoretical Biology, 227(2):219–228, March 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004193.

Alonso:2001:KAM

[AGT^{+}01] G. L. Alonso, D. A. González, D. Takara, M. A. Ostuni, and G. A. Sánchez. Kinetic analysis of a model of the sarcoplasmic reticulum Ca-ATPase, with variable stoichiometry, which enhances the amount and the rate of Ca transport. Journal of Theoretical Biology, 208(3):251–260, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921854.

Agutter:2008:FST

[Agu08] Paul S. Agutter. The flux-summation theorem and the ‘evolution of dominance’. Journal of Theoretical Biology, 254(4):821–825, October 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003937.

Ananthakrishnan:2006:QCA

[AGW^{+}06] Revathi Ananthakrishnan, Jochen Guck, Falk Wottawah, Stefan Schinkinger, Bryan Lincoln, Maren Romeyke, Tess Moon, and Josef Käs. Quantifying the contribution of actin networks to the elastic strength of fibroblasts. Journal of Theoretical Biology, 242(2):502–516, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001378. See corrigen-

dum [AGW^{+}08].

Ananthakrishnan:2008:CQC

[AGW^{+}08] Revathi Ananthakrishnan, Jochen Guck, Falk Wottawah, Stefan Schinkinger, Bryan Lincoln, Maren Romeyke, Tess Moon, and Josef Käs. Corrigendum to “Quantifying the contribution of actin networks to the elastic strength of fibroblasts” [J. Theor. Biol. 242 (2006) 502–516]. Journal of Theoretical Biology, 255(1):162, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004177. See [AGW^{+}06].

Aracena:2006:RNC

[AGZ^{+}06] Julio Aracena, Mauricio González, Alejandro Zuñiga, Marco A. Mendez, and Verónica Cambiazo. Regulatory network for cell shape changes during drosophila ventral furrow formation. Journal of Theoretical Biology, 239(1):49–62, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003139.

Aita:2000:TEM

[AH00] Takuyo Aita and Yuzuru Husimi. Theory of evolutionary molecular engineering through simultaneous accumulation of advantageous mutations. Journal of Theoretical Biology, 207(4):543–556, December 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921957.

Aita:2003:SFE

[AH03a] Takuyo Aita and Yuzuru Husimi. Statistical formulae of the energy distribution among a globular protein structure ensemble. Journal of Theoretical Biology, 220(1):107–121, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931584.

Aita:2003:TIA

[AH03b] Takuyo Aita and Yuzuru Husimi. Thermodynamical interpretation of an adaptive walk on a Mt. Fuji-type fitness landscape: Einstein relation-like formula holds in a stochastic evolution. Journal of Theoretical Biology, 225(2):215–228, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002406.

Aita:2008:FPF

[AH08] Takuyo Aita and Yuzuru Husimi. Fitting protein-folding free energy landscape for a certain conformation to an NK fitness landscape. Journal of Theoretical Biology, 253(1):151– 161, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000854.

Akutsu:2007:CBN

[AHCN07] Tatsuya Akutsu, Morihiro Hayashida, Wai-Ki Ching, and Michael K. Ng. Control of Boolean networks: Hardness results and algorithms for tree structured networks. Journal of Theoretical Biology, 244(4):670–679, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004334.

Aita:2007:ECP

[AHT^{+}07] Takuyo Aita, Yuuki Hayashi, Hitoshi Toyota, Yuzuru Husimi, Itaru Urabe, and Tetsuya Yomo. Extracting characteristic properties of fitness landscape from in vitro molecular evolution: a case study on infectivity of fd phage to E. coli. Journal of Theoretical Biology, 246(3):538–550, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005984.

Abella:2000:NTR

[AIK00] Charles A. Abella, Volodymyr N. Ivanov, and In S. Kim. Number of triplets in 16s rRNA gene related with pathogenicity of Bacillus spp. and Clostridium spp. Journal of Theoretical Biology, 205(4):581–586, August 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920903.

Assaraf:2006:CMA

[AIKP06] Yehuda G. Assaraf, Ilan Ifergan, Wisam N. Kadry, and Ron Y. Pinter. Computer modelling of antifolate inhibition of folate metabolism using hybrid functional petri nets. Journal of Theoretical Biology, 240(4):637–647, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004753.

Aita:2008:HDA

[Ait08] Takuyo Aita. Hierarchical distribution of ascending slopes, nearly neutral networks, highlands, and local optima at the d th order in an NK fitness landscape. Journal of Theoretical Biology, 254(2):252–263, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003019.

Anderson:2005:NPS

[AJ05] Paul E. Anderson and Henrik Jeldtoft Jensen. Network properties, species abundance and evolution in a model of evolutionary ecology. Journal of Theoretical Biology, 232(4):551–558, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004229.

Altaner:2008:MPI

[AJ08] C. M. Altaner and M. C. Jarvis. Modelling polymer interactions of the ‘molecular velcro’ type in wood under mechanical stress. Journal of Theoretical Biology, 253(3):434–445, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001240.

Abou-Jaoude:2009:SDF

[AJOK09] Wassim Abou-Jaoudé, Djomangan A. Ouattara, and Marcelle Kaufman. From structure to dynamics: Frequency tuning in the p53-Mdm2 network: I. Logical approach. Journal of Theoretical Biology, 258(4):561–577, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900068X.

Angeleska:2007:RGD

[AJSL07] Angela Angeleska, Natasa Jonoska, Masahico Saito, and Laura F. Landweber. RNA-guided DNA assembly. Journal of Theoretical Biology, 248(4):706–720, October 21, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002871.

Aoki:2004:HMS

[AK04] Shigeyuki Aoki and Utako Kurosu. How many soldiers are optimal for an aphid colony? Journal of Theoretical Biology, 230(3):313–317, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002528.

Allouche:2009:DAH

[AK09] Omri Allouche and Ronen Kadmon. Demographic analysis of Hubbell’s neutral theory of biodiversity. Journal of Theoretical Biology, 258(2):274–280, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800670X.

Auger:2006:BAP

[AKdlPP06] Pierre Auger, Bob W. Kooi, Rafael Bravo de la Parra, and Jean-Christophe Poggiale. Bifurcation analysis of a predatorprey model with predators using hawk and dove tactics. Journal of Theoretical Biology, 238(3):597–607, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002584.

Andrec:2005:ISG

[AKLS05] Michael Andrec, Boris N. Kholodenko, Ronald M. Levy, and Eduardo Sontag. Inference of signaling and gene regulatory networks by steady-state perturbation experiments: structure and accuracy. Journal of Theoretical Biology, 232(3):427–441, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004047.

Antal:2009:SMK

[AKR09] Tibor Antal, P. L. Krapivsky, and S. Redner. Shepherd model for knot-limited polymer ejection from a capsid. Journal of Theoretical Biology, 261(3):488–493, December 7, 2009.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003944.

Anderies:2007:LCR

[AKS07] John M. Anderies, Madhusudan Katti, and Eyal Shochat. Living in the city: Resource availability, predation, and bird population dynamics in urban areas. Journal of Theoretical Biology, 247(1):36–49, July 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000483.

Aktipis:2004:KWW

[Akt04] C. Athena Aktipis. Know when to walk away: contingent movement and the evolution of cooperation. Journal of Theoretical Biology, 231(2):249–260, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002930.

Alberty:2002:TSB

[Alb02] R. A. Alberty. Thermodynamics of systems of biochemical reactions. Journal of Theoretical Biology, 215(4):491–501, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925160.

Alberty:2008:DKP

[Alb08] Robert A. Alberty. Determination of kinetic parameters of enzyme-catalyzed reactions with a minimum number of velocity measurements. Journal of Theoretical Biology, 254(1):156–163, September 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002695.

Alerstam:2001:DBM

[Ale01] Thomas Alerstam. Detours in bird migration. Journal of Theoretical Biology, 209(3):319–331, April 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922660.

Alerstam:2009:FND

[Ale09] Thomas Alerstam. Flight by night or day? Optimal daily timing of bird migration. Journal of Theoretical Biology, 258(4):530–536, June 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000344.

Aguirre:2009:TBN

[ALM09] Jacobo Aguirre, Ester Lázaro, and Susanna C. Manrubia. A trade-off between neutrality and adaptability limits the optimization of viral quasispecies. Journal of Theoretical Biology, 261(1):148–155, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003488.

Alpert:2005:LET

[Alp05] Seymour S. Alpert. A limit on the energy transfer rate from the human fat store in hypophagia. Journal of Theoretical Biology, 233(1):1–13, March 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004175.

Agranovich:2006:CIG

[ALSM06] Alexandra Agranovich, Yoram Louzoun, Nadav Shnerb, and Sasi Moalem. Catalyst-induced growth with limited catalyst lifespan and competition. Journal of Theoretical Biology, 241(2):307–320, July 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005230.

Andersen:2001:RPC

[AM01] Louise Kold Andersen and Michael C. Mackey. Resonance in periodic chemotherapy: a case study of acute myelogenous leukemia. Journal of Theoretical Biology, 209(1):113–130, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922550.

Alboszta:2004:SES

[AM04a] Jan Alboszta and Jacek Mieķisz. Stability of evolutionarily stable strategies in discrete replicator dynamics with time delay. Journal of Theoretical Biology, 231(2):175–179, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002851.

Araujo:2004:NIV

[AM04b] R. P. Araujo and D. L. S. McElwain. New insights into vascular collapse and growth dynamics in solid tumors. Journal of Theoretical Biology, 228(3):335–346, June 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000384.

Araujo:2006:RMH

[AM06] R. P. Araujo and D. L. S. McElwain. The role of mechanical host-tumour interactions in the collapse of tumour blood vessels and tumour growth dynamics. Journal of Theoretical Biology, 238(4):817–827, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002912.

Apostu:2008:UCT

[AM08] Raluca Apostu and Michael C. Mackey. Understanding cyclical thrombocytopenia: a mathematical modeling approach. Journal of Theoretical Biology, 251(2):297–316, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006005.

Ajelli:2009:IBM

[AM09] Marco Ajelli and Stefano Merler. An individual-based model of hepatitis a transmission. Journal of Theoretical Biology, 259(3):478–488, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001465.

Amarasekare:2004:RDD

[Ama04] Priyanga Amarasekare. The role of density-dependent dispersal in source-sink dynamics. Journal of Theoretical Biology, 226(2):159–168, January 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003266.

Amarasekare:2006:PDC

[Ama06] Priyanga Amarasekare. Productivity, dispersal and the coexistence of intraguild predators and prey. Journal of Theoretical Biology, 243(1):121–133, November 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002487.

Auger:2009:EDA

[AMC^{+}09] Pierre Auger, Rachid Mchich, Tanmay Chowdhury, Gauthier Sallet, Maurice Tchuente, and Joydev Chattopadhyay. Effects of a disease affecting a predator on the dynamics of a predatorprey system. Journal of Theoretical Biology, 258(3):344–351, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308005705.

Athale:2005:SIM

[AMD05] Chaitanya Athale, Yuri Mansury, and Thomas S. Deisboeck. Simulating the impact of a molecular ‘decision-process’ on cellular phenotype and multicellular patterns in brain tumors. Journal of Theoretical Biology, 233(4):469–481, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004977. See corrigen-

dum [AMD06].

Athale:2006:CSI

[AMD06] Chaitanya Athale, Yuri Mansury, and Thomas S. Deisboeck. Corrigendum to “Simulating the impact of a molecular ‘decision-process’ on cellular phenotype and multicellular patterns in brain tumors”: [J. Theor. Biol. 233 (2005) 469– 481]. Journal of Theoretical Biology, 239(4):516–517, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003899. See [AMD05].

Alarcon:2006:MMF

[AMP06] Tomás Alarcón, Radu Marches, and Karen M. Page. Mathematical models of the fate of lymphoma B cells after antigen receptor ligation with specific antibodies. Journal of Theoretical Biology, 240(1):54–71, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500367X.

Ambrosini:2009:QMM

[AMS09] Roberto Ambrosini, Anders Pape Møller, and Nicola Saino. A quantitative measure of migratory connectivity. Journal of Theoretical Biology, 257(2):203–211, March 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006164.

Aslanidi:2002:MGI

[AMV^{+}02] O. V. Aslanidi, O. A. Mornev, M. Vesterager, M. P. Sørensen, and P. L. Christiansen. A model for glucose-induced wave propagation in pancreatic islets of langerhans. Journal of Theoretical Biology, 215(3):273–286, April 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925172.

Anderson:2000:NMA

[AMW00] Kevin R. Anderson, Neil H. Mendelson, and Joseph C. Watkins. A new mathematical approach predicts individual cell growth behavior using bacterial population information. Journal of Theoretical Biology, 202(1):87–94, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910512.

Amzallag:2004:ACB

[Amz04] G. N. Amzallag. Adaptive changes in bacteria: a consequence of nonlinear transitions in chromosome topology? Journal of Theoretical Biology, 229(3):361–369, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001468.

Arnaout:2000:CCA

[AN00] Ramy A. Arnaout and Martin A. Nowak. Competitive coexistence in antiviral immunity. Journal of Theoretical Biology, 204(3):431–441, June 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920277.

Aerts:2009:EIM

[AN09] Peter Aerts and Sandra Nauwelaerts. Environmentally induced mechanical feedback in locomotion: Frog performance as a model. Journal of Theoretical Biology, 261(3):372–378, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003324.

Assouline:2001:SNE

[ANL01] Shmuel Assouline, Shlomo Nir, and Noam Lahav. Simulation of non-enzymatic template-directed synthesis of oligonucleotides and peptides. Journal of Theoretical Biology, 208(1):117–125, January 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922057.

Anonymous:2000:ADO

[Ano00a] Anonymous. ANNOUNCEMENT: Digital object identifier. Journal of Theoretical Biology, 202(3):i, February 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930091083X.

Anonymous:2000:EP

[Ano00b] Anonymous. Editorial policy. Journal of Theoretical Biology, 204(4):479–480, June 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920873.

Anonymous:2000:Ea

[Ano00c] Anonymous. ERRATUM. Journal of Theoretical Biology, 203(2):187, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300920058.

Anonymous:2000:Eb

[Ano00d] Anonymous. ERRATUM. Journal of Theoretical Biology, 205(2):341, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300921088.

Anonymous:2000:ICMa

[Ano00e] Anonymous. International conference on mathematics in biology and annual meeting of the society for mathematical biology “SMB 2000”. Journal of Theoretical Biology, 204(1): 151, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920654.

Anonymous:2000:PNa

[Ano00f] Anonymous. Pages 1–127 (7 November 2000). Journal of Theoretical Biology, 207(1):??, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PMc

[Ano00g] Anonymous. Pages 1–151 (7 May 2000). Journal of Theoretical Biology, 204(1):??, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PSa

[Ano00h] Anonymous. Pages 1–168 (7 September 2000). Journal of Theoretical Biology, 206(1):??, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJe

[Ano00i] Anonymous. Pages 1–170 (7 July 2000). Journal of Theoretical Biology, 205(1):??, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PMa

[Ano00j] Anonymous. Pages 1–61 (7 March 2000). Journal of Theoretical Biology, 203(1):??, March 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJa

[Ano00k] Anonymous. Pages 1–94 (7 January 2000). Journal of Theoretical Biology, 202(1):??, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PNb

[Ano00l] Anonymous. Pages 129–297 (21 November 2000). Journal of Theoretical Biology, 207(2):??, November 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PMd

[Ano00m] Anonymous. Pages 153–309 (21 May 2000). Journal of Theoretical Biology, 204(2):??, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PSb

[Ano00n] Anonymous. Pages 169–322 (21 September 2000). Journal of Theoretical Biology, 206(2):??, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJf

[Ano00o] Anonymous. Pages 171–341 (21 July 2000). Journal of Theoretical Biology, 205(2):??, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PFa

[Ano00p] Anonymous. Pages 187–246 (7 February 2000). Journal of Theoretical Biology, 202(3):??, February 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PAa

[Ano00q] Anonymous. Pages 189–315 (7 April 2000). Journal of Theoretical Biology, 203(3):??, April 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PFb

[Ano00r] Anonymous. Pages 247–314 (21 February 2000). Journal of Theoretical Biology, 202(4):??, February 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PDa

[Ano00s] Anonymous. Pages 299–443 (7 December 2000). Journal of Theoretical Biology, 207(3):??, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJc

[Ano00t] Anonymous. Pages 311–478 (7 June 2000). Journal of Theoretical Biology, 204(3):??, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PAb

[Ano00u] Anonymous. Pages 317–454 (21 April 2000). Journal of Theoretical Biology, 203(4):??, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:POa

[Ano00v] Anonymous. Pages 323–447 (7 October 2000). Journal of Theoretical Biology, 206(3):??, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PAc

[Ano00w] Anonymous. Pages 343–514 (7 August 2000). Journal of Theoretical Biology, 205(3):??, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PDb

[Ano00x] Anonymous. Pages 445–578 (21 December 2000). Journal of Theoretical Biology, 207(4):??, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:POb

[Ano00y] Anonymous. Pages 449–619 (21 October 2000). Journal of Theoretical Biology, 206(4):??, October 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJd

[Ano00z] Anonymous. Pages 479–647 (21 June 2000). Journal of Theoretical Biology, 204(4):??, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PAd

[Ano00-27] Anonymous. Pages 515–666 (21 August 2000). Journal of Theoretical Biology, 205(4):??, August 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PMb

[Ano00-28] Anonymous. Pages 63–187 (21 March 2000). Journal of Theoretical Biology, 203(2):??, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:PJb

[Ano00-29] Anonymous. Pages 95–185 (21 January 2000). Journal of Theoretical Biology, 202(2):??, January 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2000:ICMb

[Ano00-30] Anonymous. The 6th International Conference on Mathematical Population Dynamics 3–8 June 2001 Marrakech, Morocco. Journal of Theoretical Biology, 207(3):443, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092215X.

Anonymous:2001:A

[Ano01a] Anonymous. ANNOUNCEMENT. Journal of Theoretical Biology, 209(3):381, April 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922702.

Anonymous:2001:ELA

[Ano01b] Anonymous. ERRATUM: a logical analysis of the drosophila gap-gene system. Journal of Theoretical Biology, 212(1):127, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923987.

Anonymous:2001:EMN

[Ano01c] Anonymous. ERRATUM: a model for a network of phosphorylation-dephosphorylation cycles displaying the dynamics of dominoes and clocks. Journal of Theoretical Biology, 212(4):565, October 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924385.

Anonymous:2001:EMJ

[Ano01d] Anonymous. Erratum: In memoriam: Joel E. Keiser. Journal of Theoretical Biology, 211(4):433, August 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923811.

Anonymous:2001:EAP

[Ano01e] Anonymous. ERRATUM: The antimicrobial properties of melanocytes, melanosomes and melanin and the evolution of black skin. Journal of Theoretical Biology, 212(1):128, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923999.

Anonymous:2001:ESC

[Ano01f] Anonymous. ERRATUM: The sense of conciousness. Journal of Theoretical Biology, 212(2):271, September 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924105.

Anonymous:2001:ESD

[Ano01g] Anonymous. ERRATUM: The size distribution of human hematagenous metastases. Journal of Theoretical Biology, 213(3):503, December 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924877.

Anonymous:2001:ESS

[Ano01h] Anonymous. ESMTB Summer school: Biology and mathematics of cells: Physiology, kinetics and evolution. Journal of Theoretical Biology, 209(4):503, April 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923069.

Anonymous:2001:PNa

[Ano01i] Anonymous. Pages 1–119 (7 November 2001). Journal of Theoretical Biology, 213(1):??, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJa

[Ano01j] Anonymous. Pages 1–125 (7 January 2001). Journal of Theoretical Biology, 208(1):??, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PSa

[Ano01k] Anonymous. Pages 1–128 (7 September 2001). Journal of Theoretical Biology, 212(1):??, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PMa

[Ano01l] Anonymous. Pages 1–130 (7 March 2001). Journal of Theoretical Biology, 209(1):??, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PMc

[Ano01m] Anonymous. Pages 1–132 (7 May 2001). Journal of Theoretical Biology, 210(1):??, May 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJe

[Ano01n] Anonymous. Pages 1–75 (7 July 2001). Journal of Theoretical Biology, 211(1):??, July 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PNb

[Ano01o] Anonymous. Pages 121–313 (21 November 2001). Journal of Theoretical Biology, 213(2):??, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJb

[Ano01p] Anonymous. Pages 127–249 (21 January 2001). Journal of Theoretical Biology, 208(2):??, January 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PSb

[Ano01q] Anonymous. Pages 129–271 (21 September 2001). Journal of Theoretical Biology, 212(2):??, September 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PMb

[Ano01r] Anonymous. Pages 131–255 (21 March 2001). Journal of Theoretical Biology, 209(2):??, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PMd

[Ano01s] Anonymous. Pages 133–263 (21 May 2001). Journal of Theoretical Biology, 210(2):??, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PAc

[Ano01t] Anonymous. Pages 187–295 (7 August 2001). Journal of Theoretical Biology, 211(3):??, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PFa

[Ano01u] Anonymous. Pages 251–386 (7 February 2001). Journal of Theoretical Biology, 208(3):??, February 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PAa

[Ano01v] Anonymous. Pages 257–381 (7 April 2001). Journal of Theoretical Biology, 209(3):??, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJc

[Ano01w] Anonymous. Pages 265–399 (7 June 2001). Journal of Theoretical Biology, 210(3):??, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:POa

[Ano01x] Anonymous. Pages 273–424 (7 October 2001). Journal of Theoretical Biology, 212(3):??, October 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PAd

[Ano01y] Anonymous. Pages 297–433 (21 August 2001). Journal of Theoretical Biology, 211(4):??, August 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PDa

[Ano01z] Anonymous. Pages 315–503 (7 December 2001). Journal of Theoretical Biology, 213(3):??, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PAb

[Ano01-27] Anonymous. Pages 383–503 (21 April 2001). Journal of Theoretical Biology, 209(4):??, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PFb

[Ano01-28] Anonymous. Pages 387–521 (21 February 2001). Journal of Theoretical Biology, 208(4):??, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJd

[Ano01-29] Anonymous. Pages 401–497 (21 June 2001). Journal of Theoretical Biology, 210(4):??, June 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:POb

[Ano01-30] Anonymous. Pages 425–565 (21 October 2001). Journal of Theoretical Biology, 212(4):??, October 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PDb

[Ano01-31] Anonymous. Pages 505–608 (21 December 2001). Journal of Theoretical Biology, 213(4):??, December 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:PJf

[Ano01-32] Anonymous. Pages 77–185 (21 July 2001). Journal of Theoretical Biology, 211(2):??, July 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2001:TGT

[Ano01-33] Anonymous. Third Georgia Tech–Emory International Conference on Bioinformatics. Journal of Theoretical Biology, 210(1):131–132, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923033.

Anonymous:2002:MEI

[Ano02a] Anonymous. Metabolic Engineering IV: Applied System Biology October 6–11, 2002 Il Ciocco, Castelvecchio Pascoli (Tuscany), Italy. Journal of Theoretical Biology, 216(2):255, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930292552X.

Anonymous:2002:PMc

[Ano02b] Anonymous. Pages 1–122 (7 May 2002). Journal of Theoretical Biology, 216(1):??, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PMa

[Ano02c] Anonymous. Pages 1–131 (7 March 2002). Journal of Theoretical Biology, 215(1):??, March 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJa

[Ano02d] Anonymous. Pages 1–134 (7 January 2002). Journal of Theoretical Biology, 214(1):??, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJe

[Ano02e] Anonymous. Pages 1–135 (7 July 2002). Journal of Theoretical Biology, 217(1):??, July 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PNa

[Ano02f] Anonymous. Pages 1–135 (7 November 2002). Journal of Theoretical Biology, 219(1):??, November 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PSa

[Ano02g] Anonymous. Pages 1–138 (7 September 2002). Journal of Theoretical Biology, 218(1):??, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PMd

[Ano02h] Anonymous. Pages 123–255 (21 May 2002). Journal of Theoretical Biology, 216(2):??, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PMb

[Ano02i] Anonymous. Pages 133–262 (21 March 2002). Journal of Theoretical Biology, 215(2):??, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJb

[Ano02j] Anonymous. Pages 135–328 (21 January 2002). Journal of Theoretical Biology, 214(2):??, January 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJf

[Ano02k] Anonymous. Pages 137–274 (21 July 2002). Journal of Theoretical Biology, 217(2):??, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PNb

[Ano02l] Anonymous. Pages 137–279 (21 November 2002). Journal of Theoretical Biology, 219(2):??, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PSb

[Ano02m] Anonymous. Pages 139–260 (21 September 2002). Journal of Theoretical Biology, 218(2):??, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJc

[Ano02n] Anonymous. Pages 257–385 (7 June 2002). Journal of Theoretical Biology, 216(3):??, June 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:POa

[Ano02o] Anonymous. Pages 261–394 (7 October 2002). Journal of Theoretical Biology, 218(3):??, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PAa

[Ano02p] Anonymous. Pages 263–384 (7 April 2002). Journal of Theoretical Biology, 215(3):??, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PAc

[Ano02q] Anonymous. Pages 275–411 (7 August 2002). Journal of Theoretical Biology, 217(3):??, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PDa

[Ano02r] Anonymous. Pages 281–413 (7 December 2002). Journal of Theoretical Biology, 219(3):??, December 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PFa

[Ano02s] Anonymous. Pages 329–518 (7 February 2002). Journal of Theoretical Biology, 214(3):??, February 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PAb

[Ano02t] Anonymous. Pages 385–501 (21 April 2002). Journal of Theoretical Biology, 215(4):??, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PJd

[Ano02u] Anonymous. Pages 387–512 (21 June 2002). Journal of Theoretical Biology, 216(4):??, June 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:POb

[Ano02v] Anonymous. Pages 395–565 (21 October 2002). Journal of Theoretical Biology, 218(4):??, October 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PAd

[Ano02w] Anonymous. Pages 413–545 (21 August 2002). Journal of Theoretical Biology, 217(4):??, August 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PDb

[Ano02x] Anonymous. Pages 415–554 (21 December 2002). Journal of Theoretical Biology, 219(4):??, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2002:PFb

[Ano02y] Anonymous. Pages 519–664 (21 February 2002). Journal of Theoretical Biology, 214(4):??, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:AIVa

[Ano03a] Anonymous. Author index for volume 222. Journal of Theoretical Biology, 222(4):I–II, June 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001802.

Anonymous:2003:AIVb

[Ano03b] Anonymous. Author index for volume 223. Journal of Theoretical Biology, 223(4):I–II, August 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002558.

Anonymous:2003:AIVc

[Ano03c] Anonymous. Author index for volume 224. Journal of Theoretical Biology, 224(4):I–II, October 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003151.

Anonymous:2003:IEBf

[Ano03d] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 223(2):ifc, July 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002078.

Anonymous:2003:IEBg

[Ano03e] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 223(3):ifc, August 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002121.

Anonymous:2003:IEBh

[Ano03f] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 223(4):ifc, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002510.

Anonymous:2003:IEBi

[Ano03g] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 224(1):ifc, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002704.

Anonymous:2003:IEBj

[Ano03h] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 224(2):ifc, September 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002844.

Anonymous:2003:IEBk

[Ano03i] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 224(3):ifc, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002960.

Anonymous:2003:IEBl

[Ano03j] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 224(4):ifc, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003114.

Anonymous:2003:IEBm

[Ano03k] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 225(1):ifc, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003436.

Anonymous:2003:IEBn

[Ano03l] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 225(2):ifc, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003485.

Anonymous:2003:IEBo

[Ano03m] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 225(3):ifc, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003643.

Anonymous:2003:IEBp

[Ano03n] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 225(4):ifc, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003710.

Anonymous:2003:IEBa

[Ano03o] Anonymous. IFC-editorial board. Journal of Theoretical Biology, 222(1):ifc, May 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001267.

Anonymous:2003:IEBb

[Ano03p] Anonymous. IFC-editorial board. Journal of Theoretical Biology, 222(2):ifc, May 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001310.

Anonymous:2003:IEBc

[Ano03q] Anonymous. IFC-editorial board. Journal of Theoretical Biology, 222(3):ifc, June 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001504.

Anonymous:2003:IEBd

[Ano03r] Anonymous. IFC-editorial board. Journal of Theoretical Biology, 222(4):ifc, June 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001760.

Anonymous:2003:IEBe

[Ano03s] Anonymous. IFC-editorial board. Journal of Theoretical Biology, 223(1):ifc, July 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300184X.

Anonymous:2003:PJd

[Ano03t] Anonymous. Pages 1–134 (7 July 2003). Journal of Theoretical Biology, 223(1):??, July 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PJa

[Ano03u] Anonymous. Pages 1–137 (7 January 2003). Journal of Theoretical Biology, 220(1):??, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PMc

[Ano03v] Anonymous. Pages 1–138 (7 May 2003). Journal of Theoretical Biology, 222(1):??, May 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PSa

[Ano03w] Anonymous. Pages 1–138 (7 September 2003). Journal of Theoretical Biology, 224(1):??, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PNa

[Ano03x] Anonymous. Pages 1–146 (7 November 2003). Journal of Theoretical Biology, 225(1):??, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PMa

[Ano03y] Anonymous. Pages 1–161 (7 March 2003). Journal of Theoretical Biology, 221(1):??, March 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PJe

[Ano03z] Anonymous. Pages 135–268 (21 July 2003). Journal of Theoretical Biology, 223(2):??, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PMd

[Ano03-27] Anonymous. Pages 139–272 (21 May 2003). Journal of Theoretical Biology, 222(2):??, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PJb

[Ano03-28] Anonymous. Pages 139–276 (21 January 2003). Journal of Theoretical Biology, 220(2):??, January 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PSb

[Ano03-29] Anonymous. Pages 139–276 (21 September 2003). Journal of Theoretical Biology, 224(2):??, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PNb

[Ano03-30] Anonymous. Pages 147–284 (21 November 2003). Journal of Theoretical Biology, 225(2):??, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PMb

[Ano03-31] Anonymous. Pages 163–307 (21 March 2003). Journal of Theoretical Biology, 221(2):??, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PAc

[Ano03-32] Anonymous. Pages 269–390 (7 August 2003). Journal of Theoretical Biology, 223(3):??, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PJc

[Ano03-33] Anonymous. Pages 273–406 (7 June 2003). Journal of Theoretical Biology, 222(3):??, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PFa

[Ano03-34] Anonymous. Pages 277–406 (7 February 2003). Journal of Theoretical Biology, 220(3):??, February 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PDa

[Ano03-35] Anonymous. Pages 285–418 (7 December 2003). Journal of Theoretical Biology, 225(3):??, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PAa

[Ano03-36] Anonymous. Pages 309–457 (7 April 2003). Journal of Theoretical Biology, 221(3):??, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PAd

[Ano03-37] Anonymous. Pages 391–536 (21 August 2003). Journal of Theoretical Biology, 223(4):??, August 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PFb

[Ano03-38] Anonymous. Pages 407–554 (21 February 2003). Journal of Theoretical Biology, 220(4):??, February 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PDb

[Ano03-39] Anonymous. Pages 419–544 (21 December 2003). Journal of Theoretical Biology, 225(4):??, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PAb

[Ano03-40] Anonymous. Pages 459–671 (21 April 2003). Journal of Theoretical Biology, 221(4):??, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PEEb

[Ano03-41] Anonymous. Pages EX1–EX2, 277–412 (7 October 2003). Journal of Theoretical Biology, 224(3):??, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PEEa

[Ano03-42] Anonymous. Pages EX1–EX2, 407–530 (21 June 2003). Journal of Theoretical Biology, 222(4):??, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:PEEc

[Ano03-43] Anonymous. Pages EX1–EX2, 413–556 (21 October 2003). Journal of Theoretical Biology, 224(4):??, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2003:SIVa

[Ano03-44] Anonymous. Subject index for volume 222. Journal of Theoretical Biology, 222(4):III–VII, June 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001814.

Anonymous:2003:SIVb

[Ano03-45] Anonymous. Subject index for volume 223. Journal of Theoretical Biology, 223(4):III–VII, August 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300256X.

Anonymous:2003:SIVc

[Ano03-46] Anonymous. Subject index for volume 224. Journal of Theoretical Biology, 224(4):III–VIII, October 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003163.

Anonymous:2004:Aa

[Ano04a] Anonymous. Announcement. Journal of Theoretical Biology, 227(3):449, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519304000670.

Anonymous:2004:AME

[Ano04b] Anonymous. Announcement: Matabolic engineering v: Genome to product. Journal of Theoretical Biology, 226 (2):251, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303004351.

Anonymous:2004:Ab

[Ano04c] Anonymous. (ATW). Journal of Theoretical Biology, 230(4): ??, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:AIa

[Ano04d] Anonymous. Author index. Journal of Theoretical Biology, 228(4):EX577–EX578, June 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400219X.

Anonymous:2004:AIb

[Ano04e] Anonymous. Author index. Journal of Theoretical Biology, 229(4):EX573–EX574, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002991.

Anonymous:2004:AIc

[Ano04f] Anonymous. Author index. Journal of Theoretical Biology, 230(4):603–604, October 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003662.

Anonymous:2004:AId

[Ano04g] Anonymous. Author index. Journal of Theoretical Biology, 231(4):615–616, December 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004412.

Anonymous:2004:AIVa

[Ano04h] Anonymous. Author index for volume 226. Journal of Theoretical Biology, 226(4):I–II, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400013X.

Anonymous:2004:AIVb

[Ano04i] Anonymous. Author index to volume 227. Journal of Theoretical Biology, 227(4):595–596, April 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000980.

Anonymous:2004:EB

[Ano04j]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 227(3):ifc, April 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400061X.

Anonymous:2004:IEBa

[Ano04k] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 226(1):ifc, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003850.

Anonymous:2004:IEBb

[Ano04l] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 226(2):ifc, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004028.

Anonymous:2004:IEBc

[Ano04m] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 226(3):ifc, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004119.

Anonymous:2004:IEBd

[Ano04n] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 226(4):ifc, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000074.

Anonymous:2004:IEBe

[Ano04o] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 227(1):ifc, March 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000190.

Anonymous:2004:IEBf

[Ano04p] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 227(2):ifc, March 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000426.

Anonymous:2004:IEBg

[Ano04q] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 227(4):ifc, April 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000815.

Anonymous:2004:IEBh

[Ano04r] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 228(1):ifc, May 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001171.

Anonymous:2004:IEBi

[Ano04s] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 228(2):ifc, May 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001262.

Anonymous:2004:IEBj

[Ano04t] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 228(3):ifc, June 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001614.

Anonymous:2004:IEBk

[Ano04u] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 228(4):ifc, June 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001687.

Anonymous:2004:IEBl

[Ano04v] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 229(1):ifc, July 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002103.

Anonymous:2004:IEBm

[Ano04w] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 229(2):ifc, July 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400222X.

Anonymous:2004:IEBn

[Ano04x] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 229(3):ifc, August 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400267X.

Anonymous:2004:IEBo

[Ano04y] Anonymous. IFC — Editorial Board. Journal of Theoretical Biology, 229(4):ifc, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002954.

Anonymous:2004:IEBp

[Ano04z] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 230(1):CO2, September 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400308X.

Anonymous:2004:IEBq

[Ano04-27] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 230(2):ifc, September 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003352.

Anonymous:2004:IEBr

[Ano04-28] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 230(3):ifc, October 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003558.

Anonymous:2004:IEB

[Ano04-29] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 230(4):ifc, October 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003625.

Anonymous:2004:IEBs

[Ano04-30] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 231(1):ifc, November 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003923.

Anonymous:2004:IEBt

[Ano04-31] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 231(2):ifc, November 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004072.

Anonymous:2004:IEBu

[Ano04-32] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 231(3):ifc, December 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004266.

Anonymous:2004:IEBv

[Ano04-33] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 231(4):ifc, December 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004370.

Anonymous:2004:PJa

[Ano04-34] Anonymous. Pages 1–124 (7 January 2004). Journal of Theoretical Biology, 226(1):??, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PSa

[Ano04-35] Anonymous. Pages 1–144 (7 September 2004). Journal of Theoretical Biology, 230(1):??, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PMb

[Ano04-36] Anonymous. Pages 1–148 (7 May 2004). Journal of Theoretical Biology, 228(1):??, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PN

[Ano04-37] Anonymous. Pages 1–152 (7 November 2004). Journal of Theoretical Biology, 231(1):??, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PJb

[Ano04-38] Anonymous. Pages 125–252 (21 January 2004). Journal of Theoretical Biology, 226(2):??, January 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PSb

[Ano04-39] Anonymous. Pages 145–288 (21 September 2004). Journal of Theoretical Biology, 230(2):??, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PMc

[Ano04-40] Anonymous. Pages 149–292 (21 May 2004). Journal of Theoretical Biology, 228(2):??, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PMa

[Ano04-41] Anonymous. Pages 159–298 (21 March 2004). Journal of Theoretical Biology, 227(2):??, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PF

[Ano04-42] Anonymous. Pages 253–372 (7 February 2004). Journal of Theoretical Biology, 226(3):??, February 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PAc

[Ano04-43] Anonymous. Pages 289–432 (7 August 2004). Journal of Theoretical Biology, 229(3):??, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PO

[Ano04-44] Anonymous. Pages 289–432 (7 October 2004). Journal of Theoretical Biology, 230(3):??, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PJc

[Ano04-45] Anonymous. Pages 293–436 (7 June 2004). Journal of Theoretical Biology, 228(3):??, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PAa

[Ano04-46] Anonymous. Pages 299–450 (7 April 2004). Journal of Theoretical Biology, 227(3):??, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PAd

[Ano04-47] Anonymous. Pages 433–572 (21 August 2004). Journal of Theoretical Biology, 229(4):??, August 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PAb

[Ano04-48] Anonymous. Pages 451–602 (21 April 2004). Journal of Theoretical Biology, 227(4):??, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEd

[Ano04-49] Anonymous. Pages EX1–EX2, 1–146 (7 July 2004). Journal of Theoretical Biology, 229(1):??, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEb

[Ano04-50] Anonymous. Pages EX1–EX2, 1–158 (7 March 2004). Journal of Theoretical Biology, 227(1):??, March 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEe

[Ano04-51] Anonymous. Pages EX1–EX2, 147–288 (21 July 2004). Journal of Theoretical Biology, 229(2):??, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEf

[Ano04-52] Anonymous. Pages EX1–EX2, 153–306 (21 November 2004). Journal of Theoretical Biology, 231(2):??, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEg

[Ano04-53] Anonymous. Pages EX1–EX2, 309–458 (7 December 2004). Journal of Theoretical Biology, 231(3):??, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEa

[Ano04-54] Anonymous. Pages EX1–EX2, 373–502 (21 February 2004). Journal of Theoretical Biology, 226(4):??, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEh

[Ano04-55] Anonymous. Pages EX1–EX2, 461–622 (21 December 2004). Journal of Theoretical Biology, 231(4):??, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:PEEc

[Ano04-56] Anonymous. Pages EX1–EX584, 437–576 (21 June 2004). Journal of Theoretical Biology, 228(4):??, June 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2004:SIa

[Ano04-57] Anonymous. Subject index. Journal of Theoretical Biology, 228(4):EX579–EX583, June 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002206.

Anonymous:2004:SIb

[Ano04-58] Anonymous. Subject index. Journal of Theoretical Biology, 229(4):EX575–EX580, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003017.

Anonymous:2004:SIc

[Ano04-59] Anonymous. Subject index. Journal of Theoretical Biology, 230(4):605–609, October 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003686.

Anonymous:2004:SId

[Ano04-60] Anonymous. Subject index. Journal of Theoretical Biology, 231(4):617–622, December 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004436.

Anonymous:2004:SIVa

[Ano04-61] Anonymous. Subject index for volume 226. Journal of Theoretical Biology, 226(4):III–VII, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000141.

Anonymous:2004:SIVb

[Ano04-62] Anonymous. Subject index to volume 227. Journal of Theoretical Biology, 227(4):597–602, April 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001018.

Anonymous:2005:A

[Ano05a] Anonymous. announcement. Journal of Theoretical Biology, 235(2):303, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305001414.

Anonymous:2005:AIa

[Ano05b] Anonymous. Author index. Journal of Theoretical Biology, 232(4):605–606, February 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400534X.

Anonymous:2005:AIb

[Ano05c] Anonymous. Author index. Journal of Theoretical Biology, 233(4):600–601, April 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000159.

Anonymous:2005:AIc

[Ano05d] Anonymous. Author index. Journal of Theoretical Biology, 234(4):605–606, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001025.

Anonymous:2005:AId

[Ano05e] Anonymous. Author index. Journal of Theoretical Biology, 235(4):597–598, August 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001979.

Anonymous:2005:AIe

[Ano05f] Anonymous. Author index. Journal of Theoretical Biology, 236(4):EX459–EX460, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003255.

Anonymous:2005:AIf

[Ano05g] Anonymous. Author index. Journal of Theoretical Biology, 237(4):EX441–EX442, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004601.

Anonymous:2005:EPA

[Ano05h] Anonymous. Erratum to: “Proteomic analyser with applications to diagnostics and vaccines” [Journal of Theoretical Biology 228 (2003) 459–465]. Journal of Theoretical Biology, 232(3):301, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002735. See [Hof04].

Anonymous:2005:IEBa

[Ano05i] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 232(1):ifc, January 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004540.

Anonymous:2005:IEBb

[Ano05j] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 232(2):ifc, January 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004886.

Anonymous:2005:IEBc

[Ano05k] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 232(3):CO2, February 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005089.

Anonymous:2005:IEBd

[Ano05l] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 232(4):CO2, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005302.

Anonymous:2005:IEBe

[Ano05m] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 233(1):CO2, March 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005508.

Anonymous:2005:IEBf

[Ano05n] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 233(2):CO2, March 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005594.

Anonymous:2005:IEBg

[Ano05o] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 233(3):CO2, April 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005892.

Anonymous:2005:IEBh

[Ano05p] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 233(4):CO2, April 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000111.

Anonymous:2005:IEBi

[Ano05q] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 234(1):CO2, May 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000299.

Anonymous:2005:IEBj

[Ano05r] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 234(2):CO2, May 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000585.

Anonymous:2005:IEBk

[Ano05s] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 234(3):CO2, June 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000858.

Anonymous:2005:IEBl

[Ano05t] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 234(4):CO2, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000986.

Anonymous:2005:IEBm

[Ano05u] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 235(1):CO2, July 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001098.

Anonymous:2005:IEBn

[Ano05v] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 235(2):CO2, July 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001359.

Anonymous:2005:IEBo

[Ano05w] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 235(3):CO2, August 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001657.

Anonymous:2005:IEBp

[Ano05x] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 235(4):CO2, August 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001931.

Anonymous:2005:IEBq

[Ano05y] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 236(1):CO2, September 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002274.

Anonymous:2005:IEBr

[Ano05z] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 236(2):CO2, September 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002602.

Anonymous:2005:IEBs

[Ano05-27] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 236(3):CO2, October 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002973.

Anonymous:2005:IEBt

[Ano05-28] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 236(4):CO2, October 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003218.

Anonymous:2005:IEBu

[Ano05-29] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 237(1):CO2, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003413.

Anonymous:2005:IEBv

[Ano05-30] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 237(2):CO2, November 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003917.

Anonymous:2005:IEBw

[Ano05-31] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 237(3):CO2, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004315.

Anonymous:2005:IEBx

[Ano05-32] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 237(4):CO2, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500456X.

Anonymous:2005:PSa

[Ano05-33] Anonymous. Pages 1–114 (7 September 2005). Journal of Theoretical Biology, 236(1):??, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PNa

[Ano05-34] Anonymous. Pages 1–116 (7 November 2005). Journal of Theoretical Biology, 237(1):??, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PJc

[Ano05-35] Anonymous. Pages 1–150 (7 July 2005). Journal of Theoretical Biology, 235(1):??, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PMa

[Ano05-36] Anonymous. Pages 1–150 (7 March 2005). Journal of Theoretical Biology, 233(1):??, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PMc

[Ano05-37] Anonymous. Pages 1–150 (7 May 2005). Journal of Theoretical Biology, 234(1):??, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PSb

[Ano05-38] Anonymous. Pages 115–228 (21 September 2005). Journal of Theoretical Biology, 236(2):??, September 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PNb

[Ano05-39] Anonymous. Pages 117–226 (21 November 2005). Journal of Theoretical Biology, 237(2):??, November 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PMd

[Ano05-40] Anonymous. Pages 151–298 (21 May 2005). Journal of Theoretical Biology, 234(2):??, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PMb

[Ano05-41] Anonymous. Pages 151–300 (21 March 2005). Journal of Theoretical Biology, 233(2):??, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PJd

[Ano05-42] Anonymous. Pages 151–304 (21 July 2005). Journal of Theoretical Biology, 235(2):??, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PDa

[Ano05-43] Anonymous. Pages 227–336 (7 December 2005). Journal of Theoretical Biology, 237(3):??, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:POa

[Ano05-44] Anonymous. Pages 229–348 (7 October 2005). Journal of Theoretical Biology, 236(3):??, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PJa

[Ano05-45] Anonymous. Pages 299–454 (7 June 2005). Journal of Theoretical Biology, 234(3):??, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PAa

[Ano05-46] Anonymous. Pages 301–450 (7 April 2005). Journal of Theoretical Biology, 233(3):??, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PFa

[Ano05-47] Anonymous. Pages 301–450 (7 February 2005). Journal of Theoretical Biology, 232(3):??, February 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PAc

[Ano05-48] Anonymous. Pages 305–450 (7 August 2005). Journal of Theoretical Biology, 235(3):??, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PDb

[Ano05-49] Anonymous. Pages 337–440 (21 December 2005). Journal of Theoretical Biology, 237(4):??, December 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:POb

[Ano05-50] Anonymous. Pages 349–458 (21 October 2005). Journal of Theoretical Biology, 236(4):??, October 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PAd

[Ano05-51] Anonymous. Pages 451–604 (21 August 2005). Journal of Theoretical Biology, 235(4):??, August 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PAb

[Ano05-52] Anonymous. Pages 451–608 (21 April 2005). Journal of Theoretical Biology, 233(4):??, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PFb

[Ano05-53] Anonymous. Pages 451–612 (21 February 2005). Journal of Theoretical Biology, 232(4):??, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PJb

[Ano05-54] Anonymous. Pages 455–612 (21 June 2005). Journal of Theoretical Biology, 234(4):??, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PEEa

[Ano05-55] Anonymous. Pages EX1–EX2, 1–150 (7 January 2005). Journal of Theoretical Biology, 232(1):??, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:PEEb

[Ano05-56] Anonymous. Pages EX1–EX2, 151–300 (21 January 2005). Journal of Theoretical Biology, 232(2):??, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2005:SIa

[Ano05-57] Anonymous. Subject index. Journal of Theoretical Biology, 232(4):607–612, February 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005363.

Anonymous:2005:SIb

[Ano05-58] Anonymous. Subject index. Journal of Theoretical Biology, 233(4):603–607, April 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000160.

Anonymous:2005:SIc

[Ano05-59] Anonymous. Subject index. Journal of Theoretical Biology, 234(4):607–612, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001049.

Anonymous:2005:SId

[Ano05-60] Anonymous. Subject index. Journal of Theoretical Biology, 235(4):599–603, August 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001992.

Anonymous:2005:SIe

[Ano05-61] Anonymous. Subject index. Journal of Theoretical Biology, 236(4):EX461–EX464, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003279.

Anonymous:2005:SIf

[Ano05-62] Anonymous. Subject index. Journal of Theoretical Biology, 237(4):EX443–EX446, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004625.

Anonymous:2006:AIa

[Ano06a] Anonymous. Author index. Journal of Theoretical Biology, 238(4):EX975–EX977, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600018X.

Anonymous:2006:AIb

[Ano06b] Anonymous. Author index. Journal of Theoretical Biology, 239(4):519–521, April 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001019.

Anonymous:2006:AIc

[Ano06c] Anonymous. Author index. Journal of Theoretical Biology, 240(4):673–675, June 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001846.

Anonymous:2006:AId

[Ano06d] Anonymous. Author index. Journal of Theoretical Biology, 241(4):968–970, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002840.

Anonymous:2006:AIe

[Ano06e] Anonymous. Author index. Journal of Theoretical Biology, 242(4):962–965, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003857.

Anonymous:2006:AIf

[Ano06f] Anonymous. Author index. Journal of Theoretical Biology, 243(4):608–610, December 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005017.

Anonymous:2006:EBa

[Ano06g] Anonymous. Editorial Board. Journal of Theoretical Biology, 243(1):CO2, November 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004140.

Anonymous:2006:EBb

[Ano06h] Anonymous. Editorial Board. Journal of Theoretical Biology, 243(2):CO2, November 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004590.

Anonymous:2006:EBc

[Ano06i] Anonymous. Editorial Board. Journal of Theoretical Biology, 243(3):CO2, December 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004772.

Anonymous:2006:EBd

[Ano06j] Anonymous. Editorial Board. Journal of Theoretical Biology, 243(4):CO2, December 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004978.

Anonymous:2006:IEBa

[Ano06k] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 238(1):CO2, January 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004844.

Anonymous:2006:IEBb

[Ano06l] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 238(2):CO2, January 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005345.

Anonymous:2006:IEBc

[Ano06m] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 238(3):CO2, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005503.

Anonymous:2006:IEBd

[Ano06n] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 238(4):CO2, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000142.

Anonymous:2006:IEBe

[Ano06o] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 239(1):CO2, March 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000464.

Anonymous:2006:IEBf

[Ano06p] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 239(2):CO2, March 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000683.

Anonymous:2006:IEBg

[Ano06q] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 239(3):CO2, April 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000841.

Anonymous:2006:IEBh

[Ano06r] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 239(4):CO2, April 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600097X.

Anonymous:2006:IEBi

[Ano06s] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 240(1):CO2, May 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600124X.

Anonymous:2006:IEBj

[Ano06t] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 240(2):CO2, May 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001421.

Anonymous:2006:IEBk

[Ano06u] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 240(3):CO2, June 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001494.

Anonymous:2006:IEBl

[Ano06v] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 240(4):CO2, June 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001809.

Anonymous:2006:IEBm

[Ano06w] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 241(1):CO2, July 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002062.

Anonymous:2006:IEBn

[Ano06x] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 241(2):CO2, July 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002347.

Anonymous:2006:IEBo

[Ano06y] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 241(3):CO2, August 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002633.

Anonymous:2006:IEBp

[Ano06z] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 241(4):CO2, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002803.

Anonymous:2006:IEBq

[Ano06-27] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 242(1):CO2, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003110.

Anonymous:2006:IEBr

[Ano06-28] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 242(2):CO2, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003365.

Anonymous:2006:IEBs

[Ano06-29] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 242(3):CO2, October 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003535.

Anonymous:2006:IEBt

[Ano06-30] Anonymous. IFC- editorial board. Journal of Theoretical Biology, 242(4):CO2, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600381X.

Anonymous:2006:PMa

[Ano06-31] Anonymous. Pages 1–128 (7 March 2006). Journal of Theoretical Biology, 239(1):??, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PMb

[Ano06-32] Anonymous. Pages 1–164 (7 May 2006). Journal of Theoretical Biology, 240(1):??, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PNa

[Ano06-33] Anonymous. Pages 1–170 (7 November 2006). Journal of Theoretical Biology, 243(1):??, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJe

[Ano06-34] Anonymous. Pages 1–172 (7 July 2006). Journal of Theoretical Biology, 241(1):??, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJa

[Ano06-35] Anonymous. Pages 1–244 (7 January 2006). Journal of Theoretical Biology, 238(1):??, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PSa

[Ano06-36] Anonymous. Pages 1–264 (7 September 2006). Journal of Theoretical Biology, 242(1):??, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PMc

[Ano06-37] Anonymous. Pages 165–336 (21 May 2006). Journal of Theoretical Biology, 240(2):??, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PNb

[Ano06-38] Anonymous. Pages 171–286 (21 November 2006). Journal of Theoretical Biology, 243(2):??, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJf

[Ano06-39] Anonymous. Pages 173–442 (21 July 2006). Journal of Theoretical Biology, 241(2):??, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJb

[Ano06-40] Anonymous. Pages 245–496 (21 January 2006). Journal of Theoretical Biology, 238(2):??, January 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PSb

[Ano06-41] Anonymous. Pages 265–528 (21 September 2006). Journal of Theoretical Biology, 242(2):??, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PDa

[Ano06-42] Anonymous. Pages 287–454 (7 December 2006). Journal of Theoretical Biology, 243(3):??, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PAa

[Ano06-43] Anonymous. Pages 289–400 (7 April 2006). Journal of Theoretical Biology, 239(3):??, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJc

[Ano06-44] Anonymous. Pages 337–510 (7 June 2006). Journal of Theoretical Biology, 240(3):??, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PAb

[Ano06-45] Anonymous. Pages 401–526 (21 April 2006). Journal of Theoretical Biology, 239(4):??, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PAc

[Ano06-46] Anonymous. Pages 443–706 (7 August 2006). Journal of Theoretical Biology, 241(3):??, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PDb

[Ano06-47] Anonymous. Pages 455–616 (21 December 2006). Journal of Theoretical Biology, 243(4):??, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PFa

[Ano06-48] Anonymous. Pages 497–736 (7 February 2006). Journal of Theoretical Biology, 238(3):??, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PJd

[Ano06-49] Anonymous. Pages 511–684 (21 June 2006). Journal of Theoretical Biology, 240(4):??, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:POa

[Ano06-50] Anonymous. Pages 529–800 (7 October 2006). Journal of Theoretical Biology, 242(3):??, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PAd

[Ano06-51] Anonymous. Pages 707–980 (21 August 2006). Journal of Theoretical Biology, 241(4):??, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:PFb

[Ano06-52] Anonymous. Pages 737–974 (21 February 2006). Journal of Theoretical Biology, 238(4):??, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:POb

[Ano06-53] Anonymous. Pages 801–976 (21 October 2006). Journal of Theoretical Biology, 242(4):??, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:SIM

[Ano06-54] Anonymous. Special issue in memory of John maynard Smith. Journal of Theoretical Biology, 239(2):??, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2006:SIa

[Ano06-55] Anonymous. Subject index. Journal of Theoretical Biology, 238(4):EX978–EX985, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000208.

Anonymous:2006:SIb

[Ano06-56] Anonymous. Subject index. Journal of Theoretical Biology, 239(4):522–526, April 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001020.

Anonymous:2006:SIc

[Ano06-57] Anonymous. Subject index. Journal of Theoretical Biology, 240(4):676–683, June 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001858.

Anonymous:2006:SId

[Ano06-58] Anonymous. Subject index. Journal of Theoretical Biology, 241(4):971–979, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002852.

Anonymous:2006:SIe

[Ano06-59] Anonymous. Subject index. Journal of Theoretical Biology, 242(4):966–975, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003869.

Anonymous:2006:SIf

[Ano06-60] Anonymous. Subject index. Journal of Theoretical Biology, 243(4):611–616, December 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005029.

Anonymous:2007:AIa

[Ano07a] Anonymous. Author index. Journal of Theoretical Biology, 244(4):745–747, February 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000136.

Anonymous:2007:AIb

[Ano07b] Anonymous. Author index. Journal of Theoretical Biology, 245(4):793–795, April 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001002.

Anonymous:2007:AIc

[Ano07c] Anonymous. Author index. Journal of Theoretical Biology, 246(4):762–764, June 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001804.

Anonymous:2007:AId

[Ano07d] Anonymous. Author index. Journal of Theoretical Biology, 247(4):863–865, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003050.

Anonymous:2007:AIe

[Ano07e] Anonymous. Author index. Journal of Theoretical Biology, 248(4):754–756, October 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700389X.

Anonymous:2007:AIf

[Ano07f] Anonymous. Author index. Journal of Theoretical Biology, 249(4):832–834, December 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005334.

Anonymous:2007:EBa

[Ano07g] Anonymous. Editorial Board. Journal of Theoretical Biology, 244(1):CO2, January 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005200.

Anonymous:2007:EBb

[Ano07h] Anonymous. Editorial Board. Journal of Theoretical Biology, 244(2):CO2, January 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005339.

Anonymous:2007:EBc

[Ano07i] Anonymous. Editorial Board. Journal of Theoretical Biology, 244(3):CO2, February 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005765.

Anonymous:2007:EBd

[Ano07j] Anonymous. Editorial Board. Journal of Theoretical Biology, 244(4):CO2, February 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000094.

Anonymous:2007:EBe

[Ano07k]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 245(1):CO2, March 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000343.

Anonymous:2007:EBf

[Ano07l]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 245(2):CO2, March 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000628.

Anonymous:2007:EBg

[Ano07m]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 245(3):ifc, April 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000768.

Anonymous:2007:EBh

[Ano07n]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 245(4):ifc, April 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000975.

Anonymous:2007:EBi

[Ano07o] Anonymous. Editorial Board. Journal of Theoretical Biology, 246(1):ifc, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519307001208.

Anonymous:2007:EBj

[Ano07p]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 246(2):ifc, May 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001488.

Anonymous:2007:EBk

[Ano07q]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 246(3):ifc, June 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001695.

Anonymous:2007:EBl

[Ano07r]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 246(4):ifc, June 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001774.

Anonymous:2007:EBm

[Ano07s] Anonymous. Editorial Board. Journal of Theoretical Biology, 247(1):ifc, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519307002093.

Anonymous:2007:EBn

[Ano07t]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 247(2):ifc, July 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002330.

Anonymous:2007:EBo

[Ano07u]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 247(3):ifc, August 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002536.

Anonymous:2007:EBp

[Ano07v]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 247(4):ifc, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003025.

Anonymous:2007:EBq

[Ano07w] Anonymous. Editorial Board. Journal of Theoretical Biology, 248(1):ifc, September 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003189.

Anonymous:2007:EBr

[Ano07x] Anonymous. Editorial Board. Journal of Theoretical Biology, 248(2):ifc, September 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003451.

Anonymous:2007:EBs

[Ano07y]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 248(3):ifc, October 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003724.

Anonymous:2007:EBt

[Ano07z]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 248(4):ifc, October 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003864.

Anonymous:2007:EBu

[Ano07-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 249(1):ifc, November 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004328.

Anonymous:2007:EBv

[Ano07-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 249(2):ifc, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004821.

Anonymous:2007:EBw

[Ano07-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 249(3):CO2, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004997.

Anonymous:2007:EBx

[Ano07-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 249(4):CO2, December 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005309.

Anonymous:2007:PJa

[Ano07-31] Anonymous. Pages 1–168 (7 January 2007). Journal of Theoretical Biology, 244(1):??, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PNa

[Ano07-32] Anonymous. Pages 1–180 (7 November 2007). Journal of Theoretical Biology, 249(1):??, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PMa

[Ano07-33] Anonymous. Pages 1–196 (7 March 2007). Journal of Theoretical Biology, 245(1):??, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PMc

[Ano07-34] Anonymous. Pages 1–204 (7 May 2007). Journal of Theoretical Biology, 246(1):??, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PJe

[Ano07-35] Anonymous. Pages 1–212 (7 July 2007). Journal of Theoretical Biology, 247(1):??, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PSa

[Ano07-36] Anonymous. Pages 1–224 (7 September 2007). Journal of Theoretical Biology, 248(1):??, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PJb

[Ano07-37] Anonymous. Pages 169–364 (21 January 2007). Journal of Theoretical Biology, 244(2):??, January 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PNb

[Ano07-38] Anonymous. Pages 181–410 (21 November 2007). Journal of Theoretical Biology, 249(2):??, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PMb

[Ano07-39] Anonymous. Pages 197–390 (21 March 2007). Journal of Theoretical Biology, 245(2):??, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PMd

[Ano07-40] Anonymous. Pages 205–394 (21 May 2007). Journal of Theoretical Biology, 246(2):??, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PJf

[Ano07-41] Anonymous. Pages 213–396 (21 July 2007). Journal of Theoretical Biology, 247(2):??, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PSb

[Ano07-42] Anonymous. Pages 225–400 (21 September 2007). Journal of Theoretical Biology, 248(2):??, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PFa

[Ano07-43] Anonymous. Pages 365–552 (7 February 2007). Journal of Theoretical Biology, 244(3):??, February 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PAa

[Ano07-44] Anonymous. Pages 391–600 (7 April 2007). Journal of Theoretical Biology, 245(3):??, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PJc

[Ano07-45] Anonymous. Pages 395–582 (7 June 2007). Journal of Theoretical Biology, 246(3):??, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PAc

[Ano07-46] Anonymous. Pages 397–582 (7 August 2007). Journal of Theoretical Biology, 247(3):??, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:POa

[Ano07-47] Anonymous. Pages 401–578 (7 October 2007). Journal of Theoretical Biology, 248(3):??, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PDa

[Ano07-48] Anonymous. Pages 411–640 (7 December 2007). Journal of Theoretical Biology, 249(3):??, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PFb

[Ano07-49] Anonymous. Pages 553–754 (21 February 2007). Journal of Theoretical Biology, 244(4):??, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:POb

[Ano07-50] Anonymous. Pages 579–764 (21 October 2007). Journal of Theoretical Biology, 248(4):??, October 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PJd

[Ano07-51] Anonymous. Pages 583–772 (21 June 2007). Journal of Theoretical Biology, 246(4):??, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PAd

[Ano07-52] Anonymous. Pages 583–872 (21 August 2007). Journal of Theoretical Biology, 247(4):??, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PAb

[Ano07-53] Anonymous. Pages 601–804 (21 April 2007). Journal of Theoretical Biology, 245(4):??, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:PDb

[Ano07-54] Anonymous. Pages 641–844 (21 December 2007). Journal of Theoretical Biology, 249(4):??, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2007:SIa

[Ano07-55] Anonymous. Subject index. Journal of Theoretical Biology, 244(4):748–754, February 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000148.

Anonymous:2007:SIb

[Ano07-56] Anonymous. Subject index. Journal of Theoretical Biology, 245(4):796–803, April 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001014.

Anonymous:2007:SIc

[Ano07-57] Anonymous. Subject index. Journal of Theoretical Biology, 246(4):765–771, June 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001816.

Anonymous:2007:SId

[Ano07-58] Anonymous. Subject index. Journal of Theoretical Biology, 247(4):866–872, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003062.

Anonymous:2007:SIe

[Ano07-59] Anonymous. Subject index. Journal of Theoretical Biology, 248(4):757–763, October 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003906.

Anonymous:2007:SIf

[Ano07-60] Anonymous. Subject index. Journal of Theoretical Biology, 249(4):835–843, December 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005346.

Anonymous:2008:AIa

[Ano08a] Anonymous. Author index. Journal of Theoretical Biology, 250(4):754–756, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307006418.

Anonymous:2008:AIb

[Ano08b] Anonymous. Author index. Journal of Theoretical Biology, 251(4):708–710, April 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001069.

Anonymous:2008:AIc

[Ano08c] Anonymous. Author index. Journal of Theoretical Biology, 252(4):776–778, June 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002257.

Anonymous:2008:AId

[Ano08d] Anonymous. Author index. Journal of Theoretical Biology, 253(4):861–863, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003470.

Anonymous:2008:EBa

[Ano08e]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 250(1):ifc, January 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005516.

Anonymous:2008:EBb

[Ano08f] Anonymous. Editorial Board. Journal of Theoretical Biology, 250(2):CO2, January 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700584X.

Anonymous:2008:EBc

[Ano08g] Anonymous. Editorial Board. Journal of Theoretical Biology, 250(3):CO2, February 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700611X.

Anonymous:2008:EBd

[Ano08h] Anonymous. Editorial Board. Journal of Theoretical Biology, 250(4):ifc, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307006388.

Anonymous:2008:EBe

[Ano08i]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 251(1):ifc, March 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000106.

Anonymous:2008:EBf

[Ano08j]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 251(2):ifc, March 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000398.

Anonymous:2008:EBg

[Ano08k]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 251(3):ifc, April 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000672.

Anonymous:2008:EBh

[Ano08l]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 251(4):ifc, April 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001033.

Anonymous:2008:EBi

[Ano08m] Anonymous. Editorial Board. Journal of Theoretical Biology, 252(1):ifc, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930800132X.

Anonymous:2008:EBj

[Ano08n]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 252(2):ifc, May 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001574.

Anonymous:2008:EBk

[Ano08o] Anonymous. Editorial Board. Journal of Theoretical Biology, 252(3):ifc, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519308001938.

Anonymous:2008:EBl

[Ano08p]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 252(4):ifc, June 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002221.

Anonymous:2008:EBm

[Ano08q] Anonymous. Editorial Board. Journal of Theoretical Biology, 253(1):ifc, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519308002531.

Anonymous:2008:EBn

[Ano08r]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 253(2):ifc, July 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002828.

Anonymous:2008:EBo

[Ano08s]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 253(3):ifc, August 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003123.

Anonymous:2008:EBp

[Ano08t]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 253(4):ifc, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003445.

Anonymous:2008:EBq

[Ano08u] Anonymous. Editorial Board. Journal of Theoretical Biology, 254(1):ifc, September 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003858.

Anonymous:2008:EBr

[Ano08v] Anonymous. Editorial Board. Journal of Theoretical Biology, 254(2):ifc, September 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004049.

Anonymous:2008:EBs

[Ano08w]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 254(3):ifc, October 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800430X.

Anonymous:2008:EBt

[Ano08x]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 254(4):ifc, October 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004554.

Anonymous:2008:EBu

[Ano08y] Anonymous. Editorial Board. Journal of Theoretical Biology, 255(1):ifc, November 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004785.

Anonymous:2008:EBv

[Ano08z] Anonymous. Editorial Board. Journal of Theoretical Biology, 255(2):ifc, November 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005249.

Anonymous:2008:EBw

[Ano08-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 255(3):ifc, December 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005511.

Anonymous:2008:EBx

[Ano08-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 255(4):ifc, December 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005638.

Anonymous:2008:PNa

[Ano08-29] Anonymous. Pages 1–162 (7 November 2008). Journal of Theoretical Biology, 255(1):??, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PMc

[Ano08-30] Anonymous. Pages 1–184 (7 May 2008). Journal of Theoretical Biology, 252(1):??, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PMa

[Ano08-31] Anonymous. Pages 1–194 (7 March 2008). Journal of Theoretical Biology, 251(1):??, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PSa

[Ano08-32] Anonymous. Pages 1–196 (7 September 2008). Journal of Theoretical Biology, 254(1):??, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PJa

[Ano08-33] Anonymous. Pages 1–220 (7 January 2008). Journal of Theoretical Biology, 250(1):??, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PJd

[Ano08-34] Anonymous. Pages 1–220 (7 July 2008). Journal of Theoretical Biology, 253(1):??, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PNb

[Ano08-35] Anonymous. Pages 163–268 (21 November 2008). Journal of Theoretical Biology, 255(2):??, November 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PMd

[Ano08-36] Anonymous. Pages 185–376 (21 May 2008). Journal of Theoretical Biology, 252(2):??, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PMb

[Ano08-37] Anonymous. Pages 195–388 (21 March 2008). Journal of Theoretical Biology, 251(2):??, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PSb

[Ano08-38] Anonymous. Pages 197–514 (21 September 2008). Journal of Theoretical Biology, 254(2):??, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PJb

[Ano08-39] Anonymous. Pages 221–388 (21 January 2008). Journal of Theoretical Biology, 250(2):??, January 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PJe

[Ano08-40] Anonymous. Pages 221–404 (21 July 2008). Journal of Theoretical Biology, 253(2):??, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PDa

[Ano08-41] Anonymous. Pages 269–344 (7 December 2008). Journal of Theoretical Biology, 255(3):??, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PDb

[Ano08-42] Anonymous. Pages 345–424 (21 December 2008). Journal of Theoretical Biology, 255(4):??, December 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PAa

[Ano08-43] Anonymous. Pages 389–552 (7 April 2008). Journal of Theoretical Biology, 251(3):??, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PFa

[Ano08-44] Anonymous. Pages 389–580 (7 February 2008). Journal of Theoretical Biology, 250(3):??, February 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PAc

[Ano08-45] Anonymous. Pages 405–628 (7 August 2008). Journal of Theoretical Biology, 253(3):??, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:POa

[Ano08-46] Anonymous. Pages 515–718 (7 October 2008). Journal of Theoretical Biology, 254(3):??, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PAb

[Ano08-47] Anonymous. Pages 553–716 (21 April 2008). Journal of Theoretical Biology, 251(4):??, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PFb

[Ano08-48] Anonymous. Pages 581–762 (21 February 2008). Journal of Theoretical Biology, 250(4):??, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PJc

[Ano08-49] Anonymous. Pages 587–786 (21 June 2008). Journal of Theoretical Biology, 252(4):??, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:PAd

[Ano08-50] Anonymous. Pages 629–872 (21 August 2008). Journal of Theoretical Biology, 253(4):??, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:POb

[Ano08-51] Anonymous. Pages 719–868 (21 October 2008). Journal of Theoretical Biology, 254(4):??, October 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2008:SIa

[Ano08-52] Anonymous. Subject index. Journal of Theoretical Biology, 250(4):757–761, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700642X.

Anonymous:2008:SIb

[Ano08-53] Anonymous. Subject index. Journal of Theoretical Biology, 251(4):711–716, April 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001070.

Anonymous:2008:SIc

[Ano08-54] Anonymous. Subject index. Journal of Theoretical Biology, 252(4):779–786, June 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002269.

Anonymous:2008:SId

[Ano08-55] Anonymous. Subject index. Journal of Theoretical Biology, 253(4):864–872, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003482.

Anonymous:2009:BTS

[Ano09a] Anonymous. Based on the talks at the second conference on computational and mathematical population dynamics. Journal of Theoretical Biology, 258(3):??, June 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:EBa

[Ano09b]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 256(1):ifc, January 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005894.

Anonymous:2009:EBb

[Ano09c]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 256(2):ifc, January 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800619X.

Anonymous:2009:EBc

[Ano09d]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 256(3):ifc, February 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006401.

Anonymous:2009:EBd

[Ano09e] Anonymous. Editorial Board. Journal of Theoretical Biology, 256(4):ifc, February 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006632.

Anonymous:2009:EBe

[Ano09f]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 257(1):ifc, March 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000149.

Anonymous:2009:EBf

[Ano09g]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 257(2):ifc, March 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900037X.

Anonymous:2009:EBg

[Ano09h]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 257(3):ifc, April 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000563.

Anonymous:2009:EBh

[Ano09i]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 257(4):ifc, April 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900085X.

Anonymous:2009:EBi

[Ano09j] Anonymous. Editorial Board. Journal of Theoretical Biology, 258(1):ifc, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309001003.

Anonymous:2009:EBj

[Ano09k]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 258(2):ifc, May 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001489.

Anonymous:2009:EBk

[Ano09l] Anonymous. Editorial Board. Journal of Theoretical Biology, 258(3):ifc, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309001726.

Anonymous:2009:EBl

[Ano09m]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 258(4):ifc, June 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001957.

Anonymous:2009:EBm

[Ano09n] Anonymous. Editorial Board. Journal of Theoretical Biology, 259(1):ifc, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309002057.

Anonymous:2009:EBn

[Ano09o]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 259(2):ifc, July 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002203.

Anonymous:2009:EBo

[Ano09p]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 259(3):ifc, August 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002768.

Anonymous:2009:EBp

[Ano09q]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 259(4):ifc, August 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002860.

Anonymous:2009:EBq

[Ano09r] Anonymous. Editorial Board. Journal of Theoretical Biology, 260(1):ifc, September 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003063.

Anonymous:2009:EBr

[Ano09s] Anonymous. Editorial Board. Journal of Theoretical Biology, 260(2):ifc, September 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003555.

Anonymous:2009:EBs

[Ano09t]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 260(3):ifc, October 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003786.

Anonymous:2009:EBt

[Ano09u]  Anonymous.  Editorial Board.  Journal of Theoreti-

cal Biology, 260(4):ifc, October 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003968.

Anonymous:2009:EBu

[Ano09v] Anonymous. Editorial Board. Journal of Theoretical Biology, 261(1):ifc, November 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004147.

Anonymous:2009:EBv

[Ano09w] Anonymous. Editorial Board. Journal of Theoretical Biology, 261(2):ifc, November 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004445.

Anonymous:2009:EBw

[Ano09x] Anonymous. Editorial Board. Journal of Theoretical Biology, 261(3):ifc, December 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004603.

Anonymous:2009:EBx

[Ano09y] Anonymous. Editorial Board. Journal of Theoretical Biology, 261(4):ifc, December 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004974.

Anonymous:2009:PJa

[Ano09z] Anonymous. Pages 1–148 (7 January 2009). Journal of Theoretical Biology, 256(1):??, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PMc

[Ano09-27] Anonymous. Pages 1–164 (7 May 2009). Journal of Theoretical Biology, 258(1):??, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PNa

[Ano09-28] Anonymous. Pages 1–164 (7 November 2009). Journal of Theoretical Biology, 261(1):??, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PSa

[Ano09-29] Anonymous. Pages 1–174 (7 September 2009). Journal of Theoretical Biology, 260(1):??, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PMa

[Ano09-30] Anonymous. Pages 1–190 (7 March 2009). Journal of Theoretical Biology, 257(1):??, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PJd

[Ano09-31] Anonymous. Pages 1–192 (7 July 2009). Journal of Theoretical Biology, 259(1):??, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PJb

[Ano09-32] Anonymous. Pages 149–304 (21 January 2009). Journal of Theoretical Biology, 256(2):??, January 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PMd

[Ano09-33] Anonymous. Pages 165–336 (21 May 2009). Journal of Theoretical Biology, 258(2):??, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PNb

[Ano09-34] Anonymous. Pages 165–360 (21 November 2009). Journal of Theoretical Biology, 261(2):??, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PSb

[Ano09-35] Anonymous. Pages 175–332 (21 September 2009). Journal of Theoretical Biology, 260(2):??, September 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PMb

[Ano09-36] Anonymous. Pages 191–344 (21 March 2009). Journal of Theoretical Biology, 257(2):??, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PJe

[Ano09-37] Anonymous. Pages 193–404 (21 July 2009). Journal of Theoretical Biology, 259(2):??, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PFa

[Ano09-38] Anonymous. Pages 305–484 (7 February 2009). Journal of Theoretical Biology, 256(3):??, February 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:POa

[Ano09-39] Anonymous. Pages 333–474 (7 October 2009). Journal of Theoretical Biology, 260(3):??, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PAa

[Ano09-40] Anonymous. Pages 345–526 (7 April 2009). Journal of Theoretical Biology, 257(3):??, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PDa

[Ano09-41] Anonymous. Pages 361–500 (7 December 2009). Journal of Theoretical Biology, 261(3):??, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PAc

[Ano09-42] Anonymous. Pages 405–658 (7 August 2009). Journal of Theoretical Biology, 259(3):??, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:POb

[Ano09-43] Anonymous. Pages 475–600 (21 October 2009). Journal of Theoretical Biology, 260(4):??, October 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PFb

[Ano09-44] Anonymous. Pages 485–662 (21 February 2009). Journal of Theoretical Biology, 256(4):??, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PJc

[Ano09-45] Anonymous. Pages 489–646 (21 June 2009). Journal of Theoretical Biology, 258(4):??, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PDb

[Ano09-46] Anonymous. Pages 501–654 (21 December 2009). Journal of Theoretical Biology, 261(4):??, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PAb

[Ano09-47] Anonymous. Pages 527–704 (21 April 2009). Journal of Theoretical Biology, 257(4):??, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2009:PAd

[Ano09-48] Anonymous. Pages 659–864 (21 August 2009). Journal of Theoretical Biology, 259(4):??, August 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anten:2002:ESL

[Ant02] Niels P. R. Anten. Evolutionarily stable leaf area production in plant populations. Journal of Theoretical Biology, 217(1):15– 32, July 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930225.

Antal:2009:SAG

[ANT09] Tibor Antal, Martin A. Nowak, and Arne Traulsen. Strategy abundance in 2×2 games for arbitrary mutation rates. Journal of Theoretical Biology, 257(2):340–344, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006127.

Amdam:2002:RAH

[AO02] Gro Vang Amdam and Stig W. Omholt. The regulatory anatomy of honeybee lifespan. Journal of Theoretical Biology, 216(2):209–228, May 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302925452.

Albert:2003:TRI

[AO03a] Réka Albert and Hans G. Othmer. The topology of the regulatory interactions predicts the expression pattern of the segment polarity genes in Drosophila melanogaster. Journal of Theoretical Biology, 223(1):1–18, July 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000353.

Amdam:2003:HBF

[AO03b] Gro Vang Amdam and Stig W. Omholt. The hive bee to forager transition in honeybee colonies: the double repressor hypothesis. Journal of Theoretical Biology, 223(4):451–464, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001218.

Azuma:2005:TST

[AO05] Akira Azuma and Masato Okamoto. Theoretical study on twodimensional aerodynamic characteristics of unsteady wings. Journal of Theoretical Biology, 234(1):67–78, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005405.

Aita:2003:SEN

[AOH03] Takuyo Aita, Motonori Ota, and Yuzuru Husimi. An in silico exploration of the neutral network in protein sequence space. Journal of Theoretical Biology, 221(4):599–613, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303932097.

Allen:2003:SBA

[AøP03] Timothy E. Allen and Bernhard ø. Palsson. Sequence-based analysis of metabolic demands for protein synthesis in prokaryotes. Journal of Theoretical Biology, 220(1):1–18, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303930876.

Amarasekare:2001:PDM

[AP01] Priyanga Amarasekare and Hugh Possingham. Patch dynamics and metapopulation theory: the case of successional species. Journal of Theoretical Biology, 209(3):333–344, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301922696.

Arii:2004:ENR

[AP04] K. Arii and L. Parrott. Emergence of non-random structure in local food webs generated from randomly structured regional webs. Journal of Theoretical Biology, 227(3):327–333, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004296.

Astanin:2009:MMW

[AP09] Sergey Astanin and Luigi Preziosi. Mathematical modelling of the Warburg effect in tumour cords. Journal of Theoretical Biology, 258(4):578–590, June 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000447.

Apaloo:2009:EMG

[Apa09] J. Apaloo. Evolutionary matrix games and optimization theory. Journal of Theoretical Biology, 257(1):84–89, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930800578X.

Araujo:2008:CDB

[APL08] R. P. Araujo, E. F. Petricoin, and L. A. Liotta. Critical dependence of blood-borne biomarker concentrations on the half-lives of their carrier proteins. Journal of Theoretical Biology, 253(3):616–622, August 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001884.

Armstrong:2006:CAM

[APS06] Nicola J. Armstrong, Kevin J. Painter, and Jonathan A. Sherratt. A continuum approach to modelling cell-cell adhesion. Journal of Theoretical Biology, 243(1):98–113, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002268.

Anand:2008:PPS

[APS08] Ashish Anand, Ganesan Pugalenthi, and P. N. Suganthan. Predicting protein structural class by SVM with class-wise optimized features and decision probabilities. Journal of Theoretical Biology, 253(2):375–380, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000982.

Alpern:2005:SMC

[AR05]  Steve Alpern and Diane Reyniers.  Strategic mating

with common preferences. Journal of Theoretical Biology, 237(4):337–354, December 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001487.

Aviram:2008:DTB

[AR08] Ira Aviram and Avinoam Rabinovitch. Dynamical types of bacteria and bacteriophages interaction: Shielding by debris. Journal of Theoretical Biology, 251(1):121–136, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005607.

Archetti:2000:OAC

[Arc00] Marco Archetti. The origin of autumn colours by coevolution. Journal of Theoretical Biology, 205(4):625–630, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920897.

Archetti:2003:SOR

[Arc03] Marco Archetti. A selfish origin for recombination. Journal of Theoretical Biology, 223(3):335–346, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001024.

Archetti:2005:ADM

[Arc05] Marco Archetti. Accumulation of deleterious mutations in hybridogenetic organisms. Journal of Theoretical Biology, 234(2):151–152, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519304005739.

Archetti:2007:ACN

[Arc07a] Marco Archetti. Autumn colours and the nutrient retranslocation hypothesis: a theoretical assessment. Journal of Theoretical Biology, 244(4):714–721, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004267.

Archetti:2007:CPE

[Arc07b] Marco Archetti. Colour preference as evidence for the theories on the evolution of autumn colours. Journal of Theoretical Biology, 245(3):595–596, April 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005443.

Archetti:2009:DVQ

[Arc09a] Marco Archetti. Decoupling vigour and quality in the autumn colours game: Weak individuals can signal, cheating can pay. Journal of Theoretical Biology, 256(3):479–484, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005493.

Archetti:2009:VDO

[Arc09b] Marco Archetti. The volunteer’s dilemma and the optimal size of a social group. Journal of Theoretical Biology, 261(3):475–480, December 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003890.

Arita:2005:RSM

[Ari05] Héctor T. Arita. Range size in mid-domain models of species diversity. Journal of Theoretical Biology, 232(1):119–126, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003753.

Arkus:2005:MMC

[Ark05] Natalie Arkus. A mathematical model of cellular apoptosis and senescence through the dynamics of telomere loss. Journal of Theoretical Biology, 235(1):13–32, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006150.

Aikio:2002:SBA

[ARKL02] Sami Aikio, Esa Ranta, Veijo Kaitala, and Per Lundberg. Seed bank in annuals: Competition between banker and non-banker morphs. Journal of Theoretical Biology, 217(3):341–349, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930341.

Armsworth:2001:DMS

[Arm01] Paul R. Armsworth. Directed motion in the sea: Efficient swimming by reef fish larvae. Journal of Theoretical Biology, 210(1):81–91, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922994.

Artzy-Randrup:2007:SSD

[AROS07] Yael Artzy-Randrup, Ronen Olinky, and Lewi Stone. Sizestructured demographic models of coral populations. Journal of Theoretical Biology, 245(3):482–497, April 7, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005108.

Anand:2008:MFG

[ARR08] M. Anand, K. Rajagopal, and K. R. Rajagopal. A model for the formation, growth, and lysis of clots in quiescent plasma. A comparison between the effects of antithrombin III deficiency and protein C deficiency. Journal of Theoretical Biology, 253(4):725–738, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800180X.

Aliev:2000:SNM

[ARW00] Rubin R. Aliev, William Richards, and John P. Wikswo. A simple nonlinear model of electrical activity in the intestine. Journal of Theoretical Biology, 204(1):21–28, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910695.

Alff-Steinberger:2000:CSM

[AS00] C. Alff-Steinberger. A comparative study of mutations in Escherichia coli and salmonella typhimurium shows that codon conservation is strongly correlated with codon usage. Journal of Theoretical Biology, 206(2):307–311, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092126X.

Anbarasu:2007:ERC

[AS07] Anand Anbarasu and Rao Sethumadhavan. Exploring the role of cation–π interactions in glycoproteins lipid-binding proteins and RNA-binding proteins. Journal of Theoretical Biology, 247(2):346–353, July 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700094X.

Anand:2009:MCC

[AS09] Ashish Anand and P. N. Suganthan. Multiclass cancer classification by support vector machines with class-wise optimized genes and probability estimates. Journal of Theoretical Biology, 259(3):533–540, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001854.

Ament:2008:ISS

[ASD08] Irene Ament, Stefan Scheu, and Barbara Drossel. Influence of spatial structure on the maintenance of sexual reproduction. Journal of Theoretical Biology, 254(3):520–528, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003366.

Aranda:2006:MIE

[ASMD06] Juan S. Aranda, Edgar Salgado, and Alejandro Muñoz-Diosdado. Multifractality in intracellular enzymatic reactions. Journal of Theoretical Biology, 240(2):209–217, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930500408X.

Addison-Smith:2008:SMM

[ASMM08] B. Addison-Smith, D. L. S. McElwain, and P. K. Maini. A simple mechanistic model of sprout spacing in tumour-associated angiogenesis. Journal of Theoretical Biology, 250(1):1–15, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003979.

Antal:2009:MSE

[ATO^{+}09] Tibor Antal, Arne Traulsen, Hisashi Ohtsuki, Corina E. Tarnita, and Martin A. Nowak. Mutation-selection equilibrium in games with multiple strategies. Journal of Theoretical Biology, 258(4):614–622, June 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000721.

Atsumi:2001:UMM

[Ats01] Tatsuo Atsumi. An ultrasonic motor model for bacterial flagellar motors. Journal of Theoretical Biology, 213(1):31–51, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924026.

Anderson:2005:IBM

[AV05] Alexander R. A. Anderson and Bakhtier N. Vasiev. An individual based model of rippling movement in a myxobacteria population. Journal of Theoretical Biology, 234(3):341–349, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930400582X.

Alvarez-Vasquez:2004:IKI

[AVSHV04] Fernando Alvarez-Vasquez, Kellie J. Sims, Yusuf A. Hannun, and Eberhard O. Voit. Integration of kinetic information on yeast sphingolipid metabolism in dynamical pathway models. Journal of Theoretical Biology, 226(3):265–291, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003291.

Almeida:2006:STP

[AW06] Paulo F. F. Almeida and Frederik W. Wiegel. A simple theory of peptide interactions on a membrane surface: Excluded volume and entropic order. Journal of Theoretical Biology, 238(2):269–278, January 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002468.

Aegerter-Wilmsen:2005:MRE

[AWAB05] Tinri Aegerter-Wilmsen, Christof M. Aegerter, and Ton Bisseling. Model for the robust establishment of precise proportions in the early drosophila embryo. Journal of Theoretical Biology, 234(1):13–19, May 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005247.

Altes:2002:DRC

[AWJ02] H. Korthals Altes, D. Wodarz, and V. A. A. Jansen. The dual role of CD4 T helper cells in the infection dynamics of HIV and their importance for vaccination. Journal of Theoretical Biology, 214(4):633–646, February 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192483X.

Allen:2009:HBM

[AWO^{+}09] Linda J. S. Allen, Curtis L. Wesley, Robert D. Owen, Douglas G. Goodin, David Koch, Colleen B. Jonsson, Yong-Kyu Chu, J. M. Shawn Hutchinson, and Robert L. Paige. A habitat-based model for the spread of hantavirus between reservoir and spillover species. Journal of Theoretical Biology, 260(4):510–522, October 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900321X.

Arino:2006:AFD

[AWW06] Julien Arino, Lin Wang, and Gail S. K. Wolkowicz. An alternative formulation for a delayed logistic equation. Journal of Theoretical Biology, 241(1):109–119, July 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500490X.

Bosatta:2003:ESC

[BÅ03] Ernesto Bosatta and Göran I. Ågren. Exact solutions to the continuous-quality equation for soil organic matter turnover. Journal of Theoretical Biology, 224(1):97–105, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001474.

Bahramian:2007:CTT

[Bah07] M. Bahman Bahramian. Coupling of transcription termination to RNAi. Journal of Theoretical Biology, 245(2):278–289, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005169.

Baker:2004:FET

[Bak04] Josh E. Baker. Free energy transduction in a chemical motor model. Journal of Theoretical Biology, 228(4):467–476, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000694.

Bakken:2007:CTM

[Bak07] George S. Bakken. A critique of thermal modeling of snake infrared reception by Jones, Lynn and stone (2001). Journal of Theoretical Biology, 244(1):167–168, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600333X.

Ballantyne:2004:RIH

[Bal04] Ford Ballantyne IV. The relative importance of herbivory and carnivory on the distribution of energy in a stochastic tri-trophic food web. Journal of Theoretical Biology, 226(3):349–357, February 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003576.

Banaji:2006:GME

[Ban06] Murad Banaji. A generic model of electron transport in mitochondria. Journal of Theoretical Biology, 243(4):501–516, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002980.

Barakat:2001:MSS

[Bar01] Abdul I. Barakat. A model for shear stress-induced deformation of a flow sensor on the surface of vascular endothelial cells. Journal of Theoretical Biology, 210(2):221–236, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301922908.

Bartosz:2005:FTC

[Bar05] Grzegorz Bartosz. Food toxicity contributes to the beneficial effects of calorie restriction. Journal of Theoretical Biology, 233(4):451–452, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005041.

Bascompte:2001:ASM

[Bas01] Jordi Bascompte. Aggregate statistical measures and metapopulation dynamics. Journal of Theoretical Biology, 209(3):373–379, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519301922751.

Battley:2006:STE

[Bat06] Edwin H. Battley. The sources of thermal energy exchange accompanying microbial catabolism. Journal of Theoretical Biology, 241(1):142–151, July 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500500X.

Battley:2009:EEB

[Bat09] Edwin H. Battley. Is electron equivalence between substrate and product preferable to C-mol equivalence in representations of microbial anabolism applicable to “origin of life” environmental conditions? Journal of Theoretical Biology, 260(2):267–275, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002562.

Boukal:2002:SSM

[BB02] David S. Boukal and Ludek Berec. Single-species models of the Allee effect: Extinction boundaries, sex ratios and mate encounters. Journal of Theoretical Biology, 218(3):375–394, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930845.

Boots:2003:BCE

[BB03] Michael Boots and Roger G. Bowers. Baseline criteria and the evolution of hosts and parasites: D_{0}, R_{0} and competition for resources between strains. Journal of Theoretical Biology, 223(3):361–365, August 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300105X.

Breban:2005:RTD

[BB05] Romulus Breban and Sally Blower. The reinfection threshold does not exist. Journal of Theoretical Biology, 235(2):151–152, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000822.

Bhattacharyya:2006:PCT

[BB06a] S. Bhattacharyya and D. K. Bhattacharya. Pest control through viral disease: Mathematical modeling and analysis. Journal of Theoretical Biology, 238(1):177–197, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002213.

Bullimore:2006:CFT

[BB06b] Sharon R. Bullimore and Jeremy F. Burn. Consequences of forward translation of the point of force application for the mechanics of running. Journal of Theoretical Biology, 238(1):211–219, January 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002134.

Braunewell:2007:SYC

[BB07a] Stefan Braunewell and Stefan Bornholdt. Superstability of the yeast cell-cycle dynamics: Ensuring causality in the presence of biochemical stochasticity. Journal of Theoretical Biology, 245(4):638–643, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005509.

Bullimore:2007:APS

[BB07b] Sharon R. Bullimore and Jeremy F. Burn. Ability of the planar spring-mass model to predict mechanical parameters in running humans. Journal of Theoretical Biology, 248(4):686–695, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002834.

Braunewell:2009:RRN

[BB09] Stefan Braunewell and Stefan Bornholdt. Reliability of regulatory networks and its evolution. Journal of Theoretical Biology, 258(4):502–512, June 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000836.

Bollinger:2007:BLB

[BBB^{+}07] R. Randal Bollinger, Andrew S. Barbas, Errol L. Bush, Shu S. Lin, and William Parker. Biofilms in the large bowel suggest an apparent function of the human vermiform appendix. Journal of Theoretical Biology, 249(4):826–831, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930700416X.

Beard:2004:TCB

[BBCQ04] Daniel A. Beard, Eric Babson, Edward Curtis, and Hong Qian. Thermodynamic constraints for biochemical networks. Journal of Theoretical Biology, 228(3):327–333, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000372.

Blumstein:2006:HDP

[BBD06] Daniel T. Blumstein, Avi Bitton, and Jose DaVeiga. How does the presence of predators influence the persistence of antipredator behavior? Journal of Theoretical Biology, 239(4):460–468, April 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003504.

Broom:2004:IID

[BBK04] Mark Broom, Carola Borries, and Andreas Koenig. Infanticide and infant defence by males-modelling the conditions in primate multi-male groups. Journal of Theoretical Biology, 231(2):261–270, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003066.

Bailly-Bechet:2008:MAI

[BBKGP08] Marc Bailly-Bechet, Michel Kerszberg, Françoise Gaill, and Florence Pradillon. A modeling approach of the influence of local hydrodynamic conditions on larval dispersal at hydrothermal vents. Journal of Theoretical Biology, 255(3):320–331, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004165.

Bernard:2003:OCN

[BBM03] Samuel Bernard, Jacques Bélair, and Michael C. Mackey. Oscillations in cyclical neutropenia: new evidence based on mathematical modeling. Journal of Theoretical Biology, 223(3):283–298, August 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000900. See erratum [UI04a].

Bernard:2004:EOC

[BBM04] Samuel Bernard, Jacques Bélair, and Micahel C. Mackey. Erratum to “Oscillations in cyclical neutropenia: new evidence based on mathematical modeling”: [J. Theor. Biol. 223, 283–298]. Journal of Theoretical Biology, 228(1):143, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000359. See [BBM03].

Barthelemy:2005:DPE

[BBPSV05] Marc Barthélemy, Alain Barrat, Romualdo Pastor-Satorras, and Alessandro Vespignani. Dynamical patterns of epidemic outbreaks in complex heterogeneous networks. Journal of Theoretical Biology, 235(2):275–288, July 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000251.

Bettendorff-Bakman:2006:FES

[BBSLN06] D.-E. Bettendorff-Bakman, P. Schmid, P. P. Lunkenheimer, and P. Niederer. A finite element study relating to the rapid filling phase of the human ventricles. Journal of Theoretical Biology, 238(2):303–316, January 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002110.

Bettendorff-Bakman:2008:DVA

[BBSLN08] D.-E. Bettendorff-Bakman, P. Schmid, P. P. Lunkenheimer, and P. Niederer. Diastolic ventricular aspiration: a mechanism supporting the rapid filling phase of the human ventricles. Journal of Theoretical Biology, 250(4):581–592, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005395.

Broom:2002:MDH

[BC02] M. Broom and C. Cannings. Modelling dominance hierarchy formation as a multi-player game. Journal of Theoretical Biology, 219(3):397–413, December 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931371.

Browning:2004:ECA

[BC04a] Lindsay Browning and Andrew M. Colman. Evolution of coordinated alternating reciprocity in repeated dyadic games. Journal of Theoretical Biology, 229(4):549–557, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400178X.

Burnley:2004:DPM

[BC04b] B. Tom Burnley and Jonathan P. L. Cox. Dynamic polyhedral models of globular proteins. Journal of Theoretical Biology, 229(2):197–208, July 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001341.

Bru:2006:EPG

[BC06] A. Brú and D. Casero. The effect of pressure on the growth of tumour cell colonies. Journal of Theoretical Biology, 243(2):171–180, November 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002177.

Barton:2009:ASP

[BC09a] N. H. Barton and J. B. Coe. On the application of statistical physics to evolutionary biology. Journal of Theoretical Biology, 259(2):317–324, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001222.

Bengtsson:2009:ESB

[BC09b]  Bengt O. Bengtsson and Nils Cronberg.  The effec-

tive size of bryophyte populations. Journal of Theoretical Biology, 258(1):121–126, May 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000046.

Bowles:2003:CEI

[BCH03] Samuel Bowles, Jung-Kyoo Choi, and Astrid Hopfensitz. The co-evolution of individual behaviors and social institutions. Journal of Theoretical Biology, 223(2):135–147, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000602.

Bajzer:2008:MCV

[BCJ^{+}08] Zeljko Bajzer, Thomas Carr, Kresimir Josić, Stephen J. Russell, and David Dingli. Modeling of cancer virotherapy with recombinant measles viruses. Journal of Theoretical Biology, 252(1):109–122, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000258.

Bastian:2008:MVG

[BCKE^{+}08] Peter Bastian, Andrés Chavarría-Krauser, Christian Engwer, Willi Jäger, Sven Marnach, and Mariya Ptashnyk. Modelling in vitro growth of dense root networks. Journal of Theoretical Biology, 254(1):99–109, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001902.

Brinkerhoff:2008:DLR

[BCL08] Christopher J. Brinkerhoff, Ji Sun Choi, and Jennifer J. Linderman. Diffusion-limited reactions in G-protein activation: Unexpected consequences of antagonist and agonist competition. Journal of Theoretical Biology, 251(4):561–569, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000052.

Burgos:2007:WNM

[BCP^{+}07] Enrique Burgos, Horacio Ceva, Roberto P. J. Perazzo, Mariano Devoto, Diego Medan, Martín Zimmermann, and Ana María Delbue. Why nestedness in mutualistic networks? Journal of Theoretical Biology, 249(2):307–313, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307003670.

Bernot:2004:AFM

[BCRG04] Gilles Bernot, Jean-Paul Comet, Adrien Richard, and Janine Guespin. Application of formal methods to biological regulatory networks: extending thomas’ asynchronous logical approach with temporal logic. Journal of Theoretical Biology, 229(3):339–347, August 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001444.

Broom:2000:SAM

[BCV00] M. Broom, C. Cannings, and G. T. Vickers. A sequentialarrivals model of territory acquisition II. Journal of Theoretical Biology, 207(3):389–403, December 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921878.

Bartumeus:2008:ITA

[BCV^{+}08] F. Bartumeus, J. Catalan, G. M. Viswanathan, E. P. Raposo, and M. G. E. da Luz. The influence of turning angles on the success of non-oriented animal searches. Journal of Theoretical Biology, 252(1):43–55, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000180.

Brusch:2008:CVT

[BD08a] Lutz Brusch and Andreas Deutsch. The coherence of the vesicle theory of protein secretion. Journal of Theoretical Biology, 252(2):370–373, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000088.

Bullimore:2008:CDS

[BD08b] Sharon R. Bullimore and J. Maxwell Donelan. Criteria for dynamic similarity in bouncing gaits. Journal of Theoretical Biology, 250(2):339–348, January 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004730.

Baptestini:2009:SCC

[BdABA09] Elizabeth M. Baptestini, Marcus A. M. de Aguiar, Daniel I. Bolnick, and Márcio S. Araújo. The shape of the competition and carrying capacity kernels affects the likelihood of disruptive selection. Journal of Theoretical Biology, 259(1):5– 11, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000733.

Benros:2009:ASS

[BdBH09] Cristina Benros, Alexandre G. de Brevern, and Serge Hazout. Analyzing the sequence-structure relationship of a library of local structural prototypes. Journal of Theoretical Biology, 256(2):215–226, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004657.

Bagdassarian:2007:BMF

[BDBR07] Carey K. Bagdassarian, Amy E. Dunham, Christopher G. Brown, and Daniel Rauscher. Biodiversity maintenance in food webs with regulatory environmental feedbacks. Journal of Theoretical Biology, 245(4):705–714, April 21, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005856.

Billoir:2008:SCW

[BDMP^{+}08] Elise Billoir, Marie Laure Delignette-Muller, Alexandre R. R. Péry, Olivier Geffard, and Sandrine Charles. Statistical cautions when estimating DEBtox parameters. Journal of Theoretical Biology, 254(1):55–64, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002440.

Banavar:2006:CRD

[BDMR06] Jayanth R. Banavar, John Damuth, Amos Maritan, and Andrea Rinaldo. Comment on “Revising the distributive networks models of West, Brown and Enquist (1997) and Banavar, Maritan and Rinaldo (1999): Metabolic inequity of living tissues provides clues for the observed allometric scaling rules” by Makarieva, Gorshkov and Li. Journal of Theoretical Biology, 239(3):391–393, April 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003371.

Burroughs:2006:RCA

[BdOP06] N. J. Burroughs, Bruno Miguel Paz Mendes de Oliveira, and Alberto Adrego Pinto. Regulatory T cell adjustment of quorum growth thresholds and the control of local immune responses. Journal of Theoretical Biology, 241(1):134–141, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004959.

Bousquet:2008:RMS

[BDR08] Nicolas Bousquet, Thierry Duchesne, and Louis-Paul Rivest. Redefining the maximum sustainable yield for the schaefer population model including multiplicative environmental noise. Journal of Theoretical Biology, 254(1):65–75, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002099.

Billoir:2009:DTM

[BdSFFDMC09] Elise Billoir, Aloysio da Silva Ferrão-Filho, Marie Laure Delignette-Muller, and Sandrine Charles. DEBtox theory and matrix population models as helpful tools in understanding the interaction between toxic cyanobacteria and zooplankton. Journal of Theoretical Biology, 258(3):380–388, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003688.

Bengtsson:2003:SMA

[BE03] Hans-Uno Bengtsson and Patrik Edén. A simple model for the arterial system. Journal of Theoretical Biology, 221(3):437– 443, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931985.

Bulashevska:2008:UBM

[BE08] Alla Bulashevska and Roland Eils. Using Bayesian multinomial classifier to predict whether a given protein sequence is intrinsically disordered. Journal of Theoretical Biology, 254(4):799–803, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002683.

Beauchamp:2000:LRS

[Bea00] Guy Beauchamp. Learning rules for social foragers: Implications for the producer–scrounger game and ideal free distribution theory. Journal of Theoretical Biology, 207(1):21–35, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921532.

Beauchemin:2006:PEW

[Bea06] Catherine Beauchemin. Probing the effects of the well-mixed assumption on viral infection dynamics. Journal of Theoretical Biology, 242(2):464–477, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001196.

Brennan:2003:CMD

[BEBV03] M. J. Brennan, S. J. Elliott, P. Bonello, and J. F. V. Vincent. The “click” mechanism in dipteran flight: if it exists, then what effect does it have? Journal of Theoretical Biology, 224(2):205–213, September 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001589.

Beil:2006:FRT

[BEF^{+}06] Michael Beil, Stefanie Eckel, Frank Fleischer, Hendrik Schmidt, Volker Schmidt, and Paul Walther. Fitting of random tessellation models to keratin filament networks. Journal of Theoretical Biology, 241(1):62–72, July 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004911.

Buonaccorsi:2003:MMS

[BEK^{+}03] John P. Buonaccorsi, Joseph Elkinton, Walt Koenig, Richard P. Duncan, Dave Kelly, and Victoria Sork. Measuring mast seeding behavior: relationships among population variation, individual variation and synchrony. Journal of Theoretical Biology, 224(1):107–114, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001486.

Bell:2006:SEC

[Bel06] Philip John Livingstone Bell. Sex and the eukaryotic cell cycle is consistent with a viral ancestry for the eukaryotic nucleus. Journal of Theoretical Biology, 243(1):54–63, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600213X.

Benhamou:2003:BNB

[Ben03] Simon Benhamou. Bicoordinate navigation based on nonorthogonal gradient fields. Journal of Theoretical Biology, 225(2):235–239, November 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300242X.

Bengtsson:2004:MEG

[Ben04a] Bengt O. Bengtsson. Modelling the evolution of genomes with integrated external and internal functions. Journal of Theoretical Biology, 231(2):271–278, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003030.

Benhamou:2004:HRE

[Ben04b] Simon Benhamou. How to reliably estimate the tortuosity of an animal’s path: straightness, sinuosity, or fractal dimension? Journal of Theoretical Biology, 229(2):209–220, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001353.

Berking:2003:MBH

[Ber03] Stefan Berking. A model for budding in hydra: pattern formation in concentric rings. Journal of Theoretical Biology, 222(1):37–52, May 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000122.

Berking:2007:GBS

[Ber07] Stefan Berking. Generation of bilateral symmetry in anthozoa: a model. Journal of Theoretical Biology, 246(3):477–490, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000227.

Beerenwinkel:2006:EDL

[BES06] Niko Beerenwinkel, Nicholas Eriksson, and Bernd Sturmfels. Evolution on distributive lattices. Journal of Theoretical Biology, 242(2):409–420, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001159.

Brennan:2008:FAR

[BFA08] Bernard J. Brennan, Samuel M. Flaxman, and Suzanne H. Alonzo. Female alternative reproductive behaviors: The effect of female group size on mate assessment and copying.

Journal of Theoretical Biology, 253(3):561–569, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930800177X.

Bennett:2005:QDC

[BFG05] M. R. Bennett, L. Farnell, and W. G. Gibson. A quantitative description of the contraction of blood vessels following the release of noradrenaline from sympathetic varicosities. Journal of Theoretical Biology, 234(1):107–122, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005430.

Bennett:2008:OBV

[BFG08a] M. R. Bennett, L. Farnell, and W. G. Gibson. Origins of blood volume change due to glutamatergic synaptic activity at astrocytes abutting on arteriolar smooth muscle cells. Journal of Theoretical Biology, 250(1):172–185, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004110.

Bennett:2008:OBC

[BFG08b] M. R. Bennett, L. Farnell, and W. G. Gibson. Origins of the BOLD changes due to synaptic activity at astrocytes abutting arteriolar smooth muscle. Journal of Theoretical Biology, 252(1):123–130, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000374.

Bennett:2009:PRL

[BFG09] Max R. Bennett, Les Farnell, and William G. Gibson. P2x7 regenerative-loop potentiation of glutamate synaptic transmission by microglia and astrocytes. Journal of Theoretical Biology, 261(1):1–16, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003373.

Bennett:2004:QDD

[BFGB04] M. R. Bennett, L. Farnell, W. G. Gibson, and D. Blair. A quantitative description of the diffusion of noradrenaline in the media of blood vessels following its release from sympathetic varicosities. Journal of Theoretical Biology, 226(3):359–372, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930300359X.

Bennett:2007:MCS

[BFGD07] M. R. Bennett, L. Farnell, W. G. Gibson, and P. Dickens. Mechanisms of calcium sequestration during facilitation at active zones of an amphibian neuromuscular junction. Journal of Theoretical Biology, 247(2):230–241, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001385.

Bertuzzi:2007:CRA

[BFGS07] A. Bertuzzi, A. Fasano, A. Gandolfi, and C. Sinisgalli. Cell resensitization after delivery of a cycle-specific anticancer drug and effect of dose splitting: Learning from tumour cords. Journal of Theoretical Biology, 244(3):388–399, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306003948.

Babajide:2001:EPS

[BFH^{+}01] Aderonke Babajide, Robert Farber, Ivo L. Hofacker, Jeff Inman, Alan S. Lapedes, and Peter F. Stadler. Exploring protein sequence space using knowledge-based potentials. Journal of Theoretical Biology, 212(1):35–46, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923434.

Bershad:2008:DAS

[BFK08] Anya K. Bershad, Miguel A. Fuentes, and David C. Krakauer. Developmental autonomy and somatic niche construction promotes robust cell fate decisions. Journal of Theoretical Biology, 254(2):408–416, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800249X.

Borrett:2007:FIE

[BFP07] Stuart R. Borrett, Brian D. Fath, and Bernard C. Patten. Functional integration of ecological networks through pathway proliferation. Journal of Theoretical Biology, 245(1):98–111, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004346.

Billiard:2005:MHP

[BFR05] Sylvain Billiard, Charlotte Faurie, and Michel Raymond. Maintenance of handedness polymorphism in humans: a frequency-dependent selection model. Journal of Theoretical Biology, 235(1):85–93, July 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304006204.

Bertuzzi:2000:CKT

[BG00a] Alessandro Bertuzzi and Alberto Gandolfi. Cell kinetics in a tumour cord. Journal of Theoretical Biology, 204(4):587–599, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910798.

Bi:2000:KPS

[BG00b] Xiping Bi and Dixie J. Goss. Kinetic proofreading scanning models for eukaryotic translational initiation: the cap and poly(a) tail dependency of translation. Journal of Theoretical Biology, 207(2):145–157, November 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921635.

Blanco:2006:BMS

[BG06] R. Ernesto Blanco and Rodolfo Gambini. A biomechanical model for size, speed and anatomical variations of the energetic costs of running mammals. Journal of Theoretical Biology, 241(1):49–61, July 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004789.

Bottani:2007:AMM

[BG07] Samuel Bottani and Basile Grammaticos. Analysis of a minimal model for p53 oscillations. Journal of Theoretical Biology, 249(2):235–245, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002044.

Bearon:2008:IBP

[BG08] R. N. Bearon and D. Grünbaum. From individual behaviour to population models: a case study using swimming algae. Journal of Theoretical Biology, 251(4):679–697, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000210.

Bentley:2008:ABS

[BGB08] Katie Bentley, Holger Gerhardt, and Paul A. Bates. Agentbased simulation of notch-mediated tip cell selection in angiogenic sprout initialisation. Journal of Theoretical Biology, 250(1):25–36, January 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004432.

Buhl:2006:GFT

[BGD^{+}06] Jérôme Buhl, Jacques Gautrais, Jean Louis Deneubourg, Pascale Kuntz, and Guy Theraulaz. The growth and form of tunnelling networks in ants. Journal of Theoretical Biology, 243(3):287–298, December 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002542.

Bauer:2005:FPM

[BGE05a] D. Bauer, R. Grebe, and A. Ehrlacher. First phase microcirculatory reaction to mechanical skin irritation: a three layer model of a compliant vascular tree. Journal of Theoretical Biology, 232(2):249–260, January 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003832.

Bauer:2005:TLC

[BGE05b] D. Bauer, R. Grebe, and A. Ehrlacher. A three-layer continuous model of porous media to describe the first phase of skin irritation. Journal of Theoretical Biology, 232(3):347–362, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003984.

Bauer:2006:NMMa

[BGE06a] D. Bauer, R. Grebe, and A. Ehrlacher. A new method to model change in cutaneous blood flow due to mechanical skin irritation: Part i: Comparison between experimental and numerical data. Journal of Theoretical Biology, 238(3):575–587, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500281X.

Bauer:2006:NMMb

[BGE06b] D. Bauer, R. Grebe, and A. Ehrlacher. A new method to model change in cutaneous blood flow due to mechanical skin irritation: Part II: Parameter identification procedure. Journal of Theoretical Biology, 238(3):588–596, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002821.

Blanco:2003:MMT

[BGF03] R. Ernesto Blanco, Rodolfo Gambini, and Richard A. Fariña. Mechanical model for theoretical determination of maximum running speed in mammals. Journal of Theoretical Biology, 222(1):117–125, May 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000195.

Baurmann:2007:ISE

[BGF07] Martin Baurmann, Thilo Gross, and Ulrike Feudel. Instabilities in spatially extended predator-prey systems: Spatiotemporal patterns in the neighborhood of Turing-Hopf bifurcations. Journal of Theoretical Biology, 245(2):220–229, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004280.

Baake:2006:SAL

[BGG06] Michael Baake, Uwe Grimm, and Robert Giegerich. Surprises in approximating Levenshtein distances. Journal of Theoretical Biology, 243(2):279–282, November 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002712.

Blavy:2009:MMH

[BGG^{+}09] P. Blavy, F. Gondret, H. Guillou, S. Lagarrigue, P. G. P. Martin, J. van Milgen, O. Radulescu, and A. Siegel. A minimal model for hepatic fatty acid balance during fasting: Application to PPAR alpha-deficient mice. Journal of Theoretical Biology, 261(2):266–278, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003312.

Baranyi:2009:PED

[BGK09] József Baranyi, Susan M. George, and Zoltán Kutalik. Parameter estimation for the distribution of single cell lag times. Journal of Theoretical Biology, 259(1):24–30, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001325.

Bhunu:2008:MEP

[BGMM08] C. P. Bhunu, W. Garira, Z. Mukandavire, and G. Magombedze. Modelling the effects of pre-exposure and post-exposure vaccines in tuberculosis control. Journal of Theoretical Biology, 254(3):633–649, October 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003093.

Bohn:2008:SCB

[BGO08] Andreas Bohn and Jordi García-Ojalvo. Synchronization of coupled biological oscillators under spatially heterogeneous environmental forcing. Journal of Theoretical Biology, 250(1):37–47, January 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004699.

Braiman:2000:FSS

[BGP00] Alex Braiman, Vladimir Gold’shtein, and Zvi Priel. Feasibility of a sustained steep Ca^{2+} gradient in the cytosol of electrically non-excitable cells. Journal of Theoretical Biology, 206(1):115–130, September 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921040.

Bowers:2001:CDT

[BH01] Roger G. Bowers and David E. Hodgkinson. Community dynamics, trade-offs, invasion criteria and the evolution of host resistance to microparasites. Journal of Theoretical Biology, 212(3):315–331, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923781.

Bluthgen:2003:HRS

[BH03a] Nils Blüthgen and Hanspeter Herzel. How robust are switches in intracellular signaling cascades? Journal of Theoretical Biology, 225(3):293–300, December 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002479.

Brown:2003:RMF

[BH03b] David H. Brown and Alan Hastings. Resistance may be futile: dispersal scales and selection for disease resistance in competing plants. Journal of Theoretical Biology, 222(3):373–388, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000444.

Barta:2004:WIA

[BH04] András Barta and Gábor Horváth. Why is it advantageous for animals to detect celestial polarization in the ultraviolet? Skylight polarization under clouds and canopies is strongest in the UV. Journal of Theoretical Biology, 226(4):429–437, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930300362X.

Babak:2008:SAD

[BH08] Petro Babak and Fangliang He. Species abundance distribution and dynamics in two locally coupled communities. Journal of Theoretical Biology, 253(4):739–748, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002142.

Bach:2006:EPC

[BHC06] L. A. Bach, T. Helvik, and F. B. Christiansen. The evolution of n-player cooperation-threshold games and ESS bifurcations. Journal of Theoretical Biology, 238(2):426–434, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002419.

Bakk:2001:TIP

[BHHS01] Audun Bakk, Johan S. Høye, Alex Hansen, and Kim Sneppen. Thermodynamical implications of a protein model with water interactions. Journal of Theoretical Biology, 210(3):367–373, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923112.

Bundschuh:2003:RDN

[BHJ03] R. Bundschuh, F. Hayot, and C. Jayaprakash. The role of dimerization in noise reduction of simple genetic networks. Journal of Theoretical Biology, 220(2):261–269, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930393164X.

Beshers:2001:SIR

[BHOR01] S. N. Beshers, Z. Y. Huang, Y. Oono, and G. E. Robinson. Social inhibition and the regulation of temporal polyethism in honey bees. Journal of Theoretical Biology, 213(3):461–479, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924270.

Bowers:2005:GTA

[BHWB05] Roger G. Bowers, Andrew Hoyle, Andrew White, and Michael Boots. The geometric theory of adaptive evolution: trade-off and invasion plots. Journal of Theoretical Biology, 233(3):363– 377, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004953.

Brodsky:2009:SSC

[BI09] Sergey V. Brodsky and Iouri Ivanov. Spermatozoa–somatic cell fusion — a mechanism for microchimerism formation. Journal of Theoretical Biology, 259(1):190–192, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001088.

Bottomley:2007:PBM

[BIB07] Christian Bottomley, Valerie Isham, and Maria-Gloria Basáñez. Population biology of multispecies helminth infection: Competition and coexistence. Journal of Theoretical Biology, 244(1):81–95, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003213.

Biegeleisen:2006:PSP

[Bie06] Ken Biegeleisen. The probable structure of the protamine-DNA complex. Journal of Theoretical Biology, 241(3):533–540, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005473.

Biondini:2008:ASL

[Bio08] Mario Biondini. Allometric scaling laws for water uptake by plant roots. Journal of Theoretical Biology, 251(1):35–59, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005784.

Bhadra:2007:HDW

[BIS^{+}07] Anindita Bhadra, Priya L. Iyer, A. Sumana, Sujata A. Deshpande, Saubhik Ghosh, and Raghavendra Gadagkar. How do workers of the primitively eusocial wasp ropalidia marginata detect the presence of their queens? Journal of Theoretical Biology, 246(3):574–582, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000240.

Brilot:2002:CCI

[BJ02] Ben O. Brilot and Rufus A. Johnstone. Cost, competition and information in communication between relatives. Journal of Theoretical Biology, 217(3):331–340, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930353.

Boswell:2002:FCN

[BJD^{+}02] Graeme P. Boswell, Helen Jacobs, Fordyce A. Davidson, Geoffrey M. Gadd, and Karl Ritz. Functional consequences of nutrient translocation in mycelial fungi. Journal of Theoretical Biology, 217(4):459–477, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930481.

Bjorklund:2002:OMC

[Bjö02] Mats Björklund. Optimal mate choice in a neural network. Journal of Theoretical Biology, 218(2):149–154, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930638.

Byrne:2001:SDE

[BK01] Christopher C. Byrne and Jeffrey A. Kurland. Selfdeception in an evolutionary game. Journal of Theoretical Biology, 212(4):457–480, October 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923902.

Ballantyne:2005:RCM

[BK05] Ford Ballantyne IV and Andrew J. Kerkhoff. Reproductive correlation and mean-variance scaling of reproductive output for a forest model. Journal of Theoretical Biology, 235(3):373–380, August 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000433.

Benison:2001:MTB

[BKCR01] Gregory Benison, Joel Keizer, Leo M. Chalupa, and David W. Robinson. Modeling temporal behavior of postnatal cat retinal ganglion cells. Journal of Theoretical Biology, 210(2):187–199, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922896.

Bannink:2006:ESV

[BKD^{+}06] A. Bannink, J. Kogut, J. Dijkstra, J. France, E. Kebreab, A. M. Van Vuuren, and S. Tamminga. Estimation of the stoichiometry of volatile fatty acid production in the rumen of lactating cows. Journal of Theoretical Biology, 238(1):36–51, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002444.

Bocharov:2003:MDL

[BKE03] Gennady Bocharov, Paul Klenerman, and Stephan Ehl. Modelling the dynamics of LCMV infection in mice: II. Compartmental structure and immunopathology. Journal of Theoretical Biology, 221(3):349–378, April 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931808. See erratum [BKE04].

Bocharov:2004:EMD

[BKE04] Gennady Bocharov, Paul Klenerman, and Stephan Ehl. Erratum to “Modelling the dynamics of LCMV infection in mice: II. Compartmental structure and immunopathology”: [J. Theor. Biol. 221, 349–378]. Journal of Theoretical Biology, 226(1):123, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003242. See [BKE03].

Boily:2000:CHF

[BKF00] P. Boily, P. H. Kvadsheim, and L. P. Folkow. Cutaneous heat flux models do not reliably predict metabolic rates of marine mammals. Journal of Theoretical Biology, 207(3):317–323, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921763.

Brylinski:2008:CSS

[BKKR08] Michał Brylinski, Leszek Konieczny, Andrzej Kononowicz, and Irena Roterman. Conservative secondary structure motifs already present in early-stage folding (in silico) as found in serpines family. Journal of Theoretical Biology, 251(2):275–285, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005474.

Bass:2009:BRT

[BKM09] Ludvik Bass, Susanne Keiding, and Ole L. Munk. Benefits and risks of transforming data from dynamic positron emission tomography, with an application to hepatic encephalopathy. Journal of Theoretical Biology, 256(4):632–636, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005729.

Bannayan:2007:GBM

[BKMH07] Mohammad Bannayan, Kazuhiko Kobayashi, Hassan Marashi, and Gerrit Hoogenboom. Gene-based modelling for rice: An opportunity to enhance the simulation of rice growth and development? Journal of Theoretical Biology, 249(3):593–605, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004092.

Belov:2009:MSI

[BKP09] Oleg V. Belov, Evgeny A. Krasavin, and Alexander Yu. Parkhomenko. Model of SOS-induced mutagenesis in bacteria Escherichia coli under ultraviolet irradiation. Journal of Theoretical Biology, 261(3):388–395, December 7, 2009.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003865.

Barilla:2000:ROD

[BL00] J. Barilla and M. Lokajícek. The role of oxygen in DNA damage by ionizing particles. Journal of Theoretical Biology, 207(3):405–414, December 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092188X.

Bohl:2006:IMM

[BL06a] Erich Bohl and Peter Lancaster. Implementation of a Markov model for phylogenetic trees. Journal of Theoretical Biology, 239(3):324–333, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003322.

Boyle:2006:FPE

[BL06b] Richard A. Boyle and Timothy M. Lenton. Fluctuation in the physical environment as a mechanism for reinforcing evolutionary transitions. Journal of Theoretical Biology, 242(4):832–843, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001925.

Barrett:2008:EAN

[BL08] R. S. Barrett and G. A. Lichtwark. Effect of altering neural, muscular and tendinous factors associated with aging on balance recovery using the ankle strategy: a simulation study. Journal of Theoretical Biology, 254(3):546–554, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003305.

Binder:2009:EPG

[BL09] Benjamin J. Binder and Kerry A. Landman. Exclusion processes on a growing domain. Journal of Theoretical Biology, 259(3):541–551, August 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900191X.

Blanchard:2004:QTA

[Bla04] Ray Blanchard. Quantitative and theoretical analyses of the relation between older brothers and homosexuality in men. Journal of Theoretical Biology, 230(2):173–187, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001924.

Beil:2009:SFK

[BLF^{+}09] Michael Beil, Sebastian Lück, Frank Fleischer, Stéphanie Portet, Wolfgang Arendt, and Volker Schmidt. Simulating the formation of keratin filament networks by a piecewise-deterministic Markov process. Journal of Theoretical Biology, 256(4):518–532, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005043.

Bejan:2008:UCT

[BLL08] A. Bejan, S. Lorente, and J. Lee. Unifying constructal theory of tree roots, canopies and forests. Journal of Theoretical Biology, 254(3):529–540, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003408.

Bastolla:2001:DPE

[BLMV01] U. Bastolla, M. Lässig, S. C. Manrubia, and A. Valleriani. Diversity patterns from ecological models at dynamical equilibrium. Journal of Theoretical Biology, 212(1):11–34, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923240.

Bastolla:2005:BMEa

[BLMV05a] Ugo Bastolla, Michael Lässig, Susanna C. Manrubia, and Angelo Valleriani. Biodiversity in model ecosystems, i: coexistence conditions for competing species. Journal of Theoretical Biology, 235(4):521–530, August 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000652.

Bastolla:2005:BMEb

[BLMV05b] Ugo Bastolla, Michael Lässig, Susanna C. Manrubia, and Angelo Valleriani. Biodiversity in model ecosystems, II: species assembly and food web structure. Journal of Theoretical Biology, 235(4):531–539, August 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000664.

Bascompte:2007:PMR

[BLOL07] Jordi Bascompte, Bartolo Luque, Jose Olarrea, and Lucas Lacasa. A probabilistic model of reserve design. Journal of Theoretical Biology, 247(1):205–211, July 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000896.

Broom:2008:GTM

[BLRR08] Mark Broom, Roger M. Luther, Graeme D. Ruxton, and Jan Rychtár. A game-theoretic model of kleptoparasitic behavior in polymorphic populations. Journal of Theoretical Biology, 255(1):81–91, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004104.

Bersini:2006:GBN

[BLS06] Hugues Bersini, Tom Lenaerts, and Francisco C. Santos. Growing biological networks: Beyond the gene-duplication model. Journal of Theoretical Biology, 241(3):488–505, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005448.

Backes:2009:CEI

[BLS^{+}09] Tracy M. Backes, Russell Latterman, Stephen A. Small, Steven Mattis, Gwyn Pauley, Emily Reilly, and Sharon R. Lubkin. Convergent extension by intercalation without mediolaterally fixed cell motion. Journal of Theoretical Biology, 256(2):180–186, January 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004414.

Brensing:2003:CDH

[BLT03] Karsten Brensing, Katrin Linke, and Dietmar Todt. Can dolphins heal by ultrasound? Journal of Theoretical Biology, 225(1):99–105, November 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300225X.

Blumenstiel:2007:SCC

[Blu07] Justin P. Blumenstiel. Sperm competition can drive a male-biased mutation rate. Journal of Theoretical Biology, 249(3):624–632, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004109.

Bebbington:2007:AMH

[BLZ07a] Mark Bebbington, Chin-Diew Lai, and Ricardas Zitikis. Addendum to “Modeling human mortality using mixtures of bathtub shaped failure distributions”: Journal of Theoretical Biology 245 (2007) 528–538. Journal of Theoretical Biology, 249(2):409–410, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004067.

Bebbington:2007:MHM

[BLZ07b] Mark Bebbington, Chin-Diew Lai, and Ricardas Zitikis. Modeling human mortality using mixtures of bathtub shaped failure distributions. Journal of Theoretical Biology, 245(3):528– 538, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005534.

Banavar:2003:CPF

[BM03a] Jayanth R. Banavar and Amos Maritan. A comment on the protein folds as platonic forms. Journal of Theoretical Biology, 223(2):263–265, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000985.

Bruist:2003:DCM

[BM03b] Michael F. Bruist and Eric Myers. Discrete and continuous mathematical models of DNA branch migration. Journal of Theoretical Biology, 220(2):139–156, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931213.

Bakk:2004:NBR

[BM04] Audun Bakk and Ralf Metzler. Nonspecific binding of the OR repressors CI and Cro of bacteriophage λ. Journal of Theoretical Biology, 231(4):525–533, December 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003285.

Barta:2006:TSD

[BM06] Efrath Barta and Alice Maroudas. A theoretical study of the distribution of insulin-like growth factor in human articular cartilage. Journal of Theoretical Biology, 241(3):628–638, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000063.

Bahi:2008:SMG

[BM08] Jacques M. Bahi and Christian J. Michel. A stochastic model of gene evolution with chaotic mutations. Journal of Theoretical Biology, 255(1):53–63, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003950.

Barabas:2009:WEB

[BM09] György Barabás and Géza Meszéna. When the exception becomes the rule: The disappearance of limiting similarity in the Lotka–Volterra model. Journal of Theoretical Biology, 258(1):89–94, May 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000034.

Beauchemin:2008:MAT

[BMD^{+}08] Catherine A. A. Beauchemin, James J. McSharry, George L. Drusano, Jack T. Nguyen, Gregory T. Went, Ruy M. Ribeiro, and Alan S. Perelson. Modeling amantadine treatment of influenza A virus in vitro. Journal of Theoretical Biology, 254(2):439–451, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002798.

Bankhead:2007:CAS

[BMH07] Armand Bankhead, Nancy S. Magnuson, and Robert B. Heckendorn. Cellular automaton simulation examining progenitor hierarchy structure effects on mammary ductal carcinoma in situ. Journal of Theoretical Biology, 246(3):491–498, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307000264.

BenHassen:2008:CIB

[BMR08] Hanen Ben Hassen, Afif Masmoudi, and Ahmed Rebai. Causal inference in biomolecular pathways using a Bayesian network approach and an implicit method. Journal of Theoretical Biology, 253(4):717–724, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002129.

Bunimovich-Mendrazitsky:2005:MPD

[BMS05] Svetlana Bunimovich-Mendrazitsky and Lewi Stone. Modeling polio as a disease of development. Journal of Theoretical Biology, 237(3):302–315, December 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001815.

Brumer:2006:GIQ

[BMS06] Yisroel Brumer, Franziska Michor, and Eugene I. Shakhnovich. Genetic instability and the quasispecies model. Journal of Theoretical Biology, 241(2):216–222, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005047.

Bain:2004:WTC

[BMT04] E. L. Bain, A. J. Millar, and M. S. Turner. The wildtype circadian period of neurospora is encoded in the residual network of the null frq mutants. Journal of Theoretical Biology, 229(3):413–420, August 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001511.

Behera:2004:PPC

[BN04] Narayan Behera and Vidyanand Nanjundiah. Phenotypic plasticity can potentiate rapid evolutionary change. Journal of Theoretical Biology, 226(2):177–184, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003308.

Bengtsson:2004:IWL

[BNRW04] Göran Bengtsson, Elna Nilsson, Tobias Rydén, and Magnus Wiktorsson. Irregular walks and loops combines in smallscale movement of a soil insect: implications for dispersal biology. Journal of Theoretical Biology, 231(2):299–306, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003169.

Broadley:2000:LRG

[BNT^{+}00] Kenneth J. Broadley, Paul H. J. Nederkoorn, Henk Timmerman, David Timms, and Robin H. Davies. The ligandreceptor-G-protein ternary complex as a GTP-synthase. Steady-state proton pumping and dose-response relationships for β-adrenoceptors. Journal of Theoretical Biology, 205(2):297–320, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300920678.

Bueno-Orovio:2008:MMH

[BOCF08] Alfonso Bueno-Orovio, Elizabeth M. Cherry, and Flavio H. Fenton. Minimal model for human ventricular action potentials in tissue. Journal of Theoretical Biology, 253(3):544–560, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001690.

Bodovitz:2008:NCC

[Bod08] Steven Bodovitz. The neural correlate of consciousness. Journal of Theoretical Biology, 254(3):594–598, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002038.

Bogaert:2004:PMH

[Bog04] Anthony F. Bogaert. The prevalence of male homosexuality: the effect of fraternal birth order and variations in family size. Journal of Theoretical Biology, 230(1):33–37, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001584.

Bokma:2006:ANN

[Bok06] Folmer Bokma. Artificial neural networks can learn to estimate extinction rates from molecular phylogenies. Journal of Theoretical Biology, 243(3):449–454, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002724.

Bolnick:2006:MSO

[Bol06] Daniel I. Bolnick. Multi-species outcomes in a common model of sympatric speciation. Journal of Theoretical Biology, 241(4):734–744, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000117.

Bonsall:2003:RVR

[Bon03] Michael B. Bonsall. The role of variability and risk on the persistence of shared-enemy, predator-prey assemblages. Journal of Theoretical Biology, 221(2):193–204, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303931754.

Bond:2004:CMP

[Bon04a] Alan H. Bond. A computational model for the primate neocortex based on its functional architecture. Journal of Theoretical Biology, 227(1):81–102, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003825.

Bond:2004:IPA

[Bon04b] Alan H. Bond. An information-processing analysis of the functional architecture of the primate neocortex. Journal of Theoretical Biology, 227(1):51–79, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003813.

Bonsall:2006:EAA

[Bon06] Michael B. Bonsall. The evolution of anisogamy: The adaptive significance of damage, repair and mortality. Journal of Theoretical Biology, 238(1):198–210, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002092.

Bruggeman:2008:EBS

[BOvD08] Frank J. Bruggeman, Ionela Oancea, and Roel van Driel. Exploring the behavior of small eukaryotic gene networks. Journal of Theoretical Biology, 252(3):482–487, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000891.

Ball:2000:SCG

[BP00] M. A. Ball and G. A. Parker. Sperm competition games: a comparison of loaded raffle models and their biological implications. Journal of Theoretical Biology, 206(4):487–506, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921428.

Baranyi:2001:PSB

[BP01] József Baranyi and Carmen Pin. A parallel study on bacterial growth and inactivation. Journal of Theoretical Biology, 210(3):327–336, June 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923124.

Ball:2003:SCG

[BP03] M. A. Ball and G. A. Parker. Sperm competition games: sperm selection by females. Journal of Theoretical Biology, 224(1):27–42, September 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001188.

Bomze:2008:OTR

[BP08] Immanuel Bomze and Christina Pawlowitsch. One-third rules with equality: Second-order evolutionary stability conditions in finite populations. Journal of Theoretical Biology, 254(3):616–620, October 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003032.

Ball:2008:ECL

[BPC08] M. A. Ball, G. A. Parker, and J. C. Chubb. The evolution of complex life cycles when parasite mortality is size- or timedependent. Journal of Theoretical Biology, 253(1):202–214, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000921.

Bertram:2006:SMM

[BPLS06] Richard Bertram, Morten Gram Pedersen, Dan S. Luciani, and Arthur Sherman. A simplified model for mitochondrial ATP production. Journal of Theoretical Biology, 243(4):575–586, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003225.

Bailon-Plaza:2001:MFS

[BPV01] Alicia Bailón-Plaza and Marjolein C. H. Van Der Meulen. A mathematical framework to study the effects of growth factor influences on fracture healing. Journal of Theoretical Biology, 212(2):191–209, September 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923720.

Bulychev:2001:LTP

[BPZ^{+}01] A. A. Bulychev, A. A. Polezhaev, S. V. Zykov, T. Yu. Pljusnina, G. Yu. Riznichenko, A. B. Rubin, W. Jantoß, V. S. Zykov, and S. C. Müller. Light-triggered pH banding profile in chara cells revealed with a scanning pH microprobe and its relation to self-organization phenomena. Journal of Theoretical Biology, 212(3):275–294, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923756.

Bascompte:2000:SDS

[BR00] Jordi Bascompte and Miguel A. Rodríguez. Self-disturbance as a source of spatiotemporal heterogeneity: the case of the tallgrass prairie. Journal of Theoretical Biology, 204(2):153–164, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920022.

Bertram:2001:MWS

[BR01] John E. A. Bertram and Andy Ruina. Multiple walking speed-frequency relations are predicted by constrained optimization. Journal of Theoretical Biology, 209(4):445–453, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922799.

Boyd:2002:GBN

[BR02] Robert Boyd and Peter J. Richerson. Group beneficial norms can spread rapidly in a structured population. Journal of Theoretical Biology, 215(3):287–296, April 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925159.

Benoit:2004:CMB

[BR04] Eric Benoît and Marie-Joëlle Rochet. A continuous model of biomass size spectra governed by predation and the effects of fishing on them. Journal of Theoretical Biology, 226(1):9–21, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930300290X.

Bartha:2006:VNR

[BR06] K. Bartha and H. Rieger. Vascular network remodeling via vessel cooption, regression and growth in tumors. Journal of Theoretical Biology, 241(4):903–918, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000373.

Bonsall:2008:LPN

[BR08] Michael B. Bonsall and Ben Raymond. Lethal pathogens, non-lethal synergists and the evolutionary ecology of resistance. Journal of Theoretical Biology, 254(2):339–349, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002944.

Boyd:2009:VYF

[BR09] Robert Boyd and Peter J. Richerson. Voting with your feet: Payoff biased migration and the evolution of group beneficial behavior. Journal of Theoretical Biology, 257(2):331–339, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006371.

Bray:2001:RHP

[Bra01] Bonnie Anderson Bray. The role of hyaluronan in the pulmonary alveolus. Journal of Theoretical Biology, 210(1):121– 130, May 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923057.

Brazhnik:2005:IGN

[Bra05] Paul Brazhnik. Inferring gene networks from steady-state response to single-gene perturbations. Journal of Theoretical Biology, 237(4):427–440, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001918.

Braumann:2007:HRE

[Bra07] Carlos A. Braumann. Harvesting in a random environment: Itô or Stratonovich calculus? Journal of Theoretical Biology, 244(3):424–432, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003961.

Borodich:2003:SPH

[BRC^{+}03] Andrei Borodich, Igor Rojdestvenski, Michael Cottam, Jan Anderson, and Gunnar Öquist. Segregation of the photosystems in higher plant thylakoids and short- and longterm regulation by a mesoscopic approach. Journal of Theoretical Biology, 225(4):431–441, December 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002765.

Bairagi:2007:RIS

[BRC07] N. Bairagi, P. K. Roy, and J. Chattopadhyay. Role of infection on the stability of a predator-prey system with several response functions — a comparative study. Journal of Theoretical Biology, 248(1):10–25, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002275.

Banasek-Richter:2004:SER

[BRCB04] Carolin Banasek-Richter, Marie-France Cattin, and Louis-Félix Bersier. Sampling effects and the robustness of quantitative and qualitative food-web descriptors. Journal of Theoretical Biology, 226(1):23–32, January 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003059.

Burg:2009:MMV

[BRND09] David Burg, Libin Rong, Avidan U. Neumann, and Harel Dahari. Mathematical modeling of viral kinetics under immune control during primary HIV-1 infection. Journal of Theoretical Biology, 259(4):751–759, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001702.

Broom:2002:UMD

[Bro02] M. Broom. A unified model of dominance hierarchy formation and maintenance. Journal of Theoretical Biology, 219(1):63–72, November 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302931097.

Billy:2009:PBM

[BRS^{+}09] Frédérique Billy, Benjamin Ribba, Olivier Saut, Hélène Morre-Trouilhet, Thierry Colin, Didier Bresch, Jean-Pierre Boissel, Emmanuel Grenier, and Jean-Pierre Flandrois. A pharmacologically based multiscale mathematical model of angiogenesis and its use in investigating the efficacy of a new cancer treatment strategy. Journal of Theoretical Biology, 260(4):545–562, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003026.

Braunheim:2000:NNM

[BS00] Benjamin B. Braunheim and Steven D. Schwartz. Neural network methods for identification and optimization of quantum mechanical features needed for bioactivity. Journal of Theoretical Biology, 206(1):27–45, September 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920988.

Binmore:2001:CMS

[BS01] Ken Binmore and Larry Samuelson. Can mixed strategies be stable in asymmetric games? Journal of Theoretical Biology, 210(1):1–14, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920836.

Brandt:2004:LRA

[BS04] Hannelore Brandt and Karl Sigmund. The logic of reprobation: assessment and action rules for indirect reciprocation. Journal of Theoretical Biology, 231(4):475–486, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002784.

Bahler:2005:LCR

[BS05a] Jürg Bähler and Sasa Svetina. A logical circuit for the regulation of fission yeast growth modes. Journal of Theoretical Biology, 237(2):210–218, November 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001633.

Bearne:2005:RIK

[BS05b] Stephen L. Bearne and Raymond J. Spiteri. Reduction of intrinsic kinetic and thermodynamic barriers for enzymecatalysed proton transfers from carbon acid substrates. Journal of Theoretical Biology, 233(4):563–571, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400520X.

Bodenstein:2005:FCP

[BS05c] L. Bodenstein and C. D. Stern. Formation of the chick primitive streak as studied in computer simulations. Journal of Theoretical Biology, 233(2):253–269, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004795.

Brandt:2006:GBD

[BS06] Hannelore Brandt and Karl Sigmund. The good, the bad and the discriminator-errors in direct and indirect reciprocity. Journal of Theoretical Biology, 239(2):183–194, March 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003796.

Banaie:2008:HDM

[BSGT08] Masood Banaie, Yashar Sarbaz, Shahriar Gharibzadeh, and Farzad Towhidkhah. Huntington’s disease: Modeling the gait disorder and proposing novel treatments. Journal of Theoretical Biology, 254(2):361–367, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002737.

Bruun:2004:EVS

[BSJ04] Sander Bruun, Johan Six, and Lars S. Jensen. Estimating vital statistics and age distributions of measurable soil organic carbon fractions based on their pathway of formation and radiocarbon content. Journal of Theoretical Biology, 230(2):241–250, September 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002085.

Broom:2006:ESD

[BSR06] M. Broom, M. P. Speed, and G. D. Ruxton. Evolutionarily stable defence and signalling of that defence. Journal of Theoretical Biology, 242(1):32–43, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000543.

Bywater:2002:CSM

[BSRH02] R. P. Bywater, A. Sørensen, P. Røgen, and P. G. Hjorth. Construction of the simplest model to explain complex receptor activation kinetics. Journal of Theoretical Biology, 218(2):139–147, September 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930365.

Bywater:2003:CSM

[BSRH03] R. P. Bywater, A. Sorensen, P. Rogen, and P. G. Hjorth. “Construction of the Simplest Model to Explain Complex Receptor Activation Kinetics” [Journal of Theoretical Biology 218 (2002) 139–147]. Journal of Theoretical Biology, 221(4):669, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932048.

Billings:2007:IMD

[BSS^{+}07] Lora Billings, Ira B. Schwartz, Leah B. Shaw, Marie McCrary, Donald S. Burke, and Derek A. T. Cummings. Instabilities in multiserotype disease models with antibody-dependent enhancement. Journal of Theoretical Biology, 246(1):18–27, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005923.

Beauchemin:2005:SCA

[BST05] Catherine Beauchemin, John Samuel, and Jack Tuszynski. A simple cellular automaton model for influenza A viral infections. Journal of Theoretical Biology, 232(2):223–234, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003819.

Boderke:2000:MDC

[BSWM00] Peter Boderke, Klaus Schittkowski, Manfred Wolf, and Hans P. Merkle. Modeling of diffusion and concurrent metabolism in cutaneous tissue. Journal of Theoretical Biology, 204(3):393–407, June 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930092023X.

Ben-Tal:2006:SMG

[BT06] Alona Ben-Tal. Simplified models for gas exchange in the human lungs. Journal of Theoretical Biology, 238(2):474–495, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002390.

Borup:2008:AMS

[BTA08] Mia T. Borup, Ala Trusina, and Anna M. C. Andersson. Aging mechanism as the “down side” of adaptation: a network approach. Journal of Theoretical Biology, 250(1):66–74, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004444.

Brooker:2007:MSR

[BTCD07] Rob W. Brooker, Justin M. J. Travis, Ewen J. Clark, and Calvin Dytham. Modelling species’ range shifts in a changing climate: The impacts of biotic interactions, dispersal distance and the rate of climate change. Journal of Theoretical Biology, 245(1):59–65, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004577.

Brinkerhoff:2008:CCC

[BTL08] Christopher J. Brinkerhoff, John R. Traynor, and Jennifer J. Linderman. Collision coupling, crosstalk, and compartmentalization in G-protein coupled receptor systems: Can a single model explain disparate results? Journal of Theoretical Biology, 255(3):278–286, December 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800413X.

Ben-Tal:2008:MCB

[BTS08] Alona Ben-Tal and Jeffrey C. Smith. A model for control of breathing in mammals: Coupling neural dynamics to peripheral gas exchange and transport. Journal of Theoretical Biology, 251(3):480–497, April 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006558.

Bucovaz:2004:PMW

[Buc04] E. T. Bucovaz. A proposed mechanism which may explain why the fetus is not considered to be foreign tissue by the maternal system during pregnancy. Journal of Theoretical Biology, 226(1):1–8, January 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002753.

Bull:2006:OMP

[Bul06] J. J. Bull. Optimality models of phage life history and parallels in disease evolution. Journal of Theoretical Biology, 241(4):928–938, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000415.

Bull:2008:OBM

[Bul08] J. J. Bull. The optimal burst of mutation to create a phenotype. Journal of Theoretical Biology, 254(3):667–673, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800310X.

Burgess:2009:SOP

[Bur09] Matthew G. Burgess. Sub-optimal pit construction in predatory ant lion larvae (Myrmeleon sp.). Journal of Theoretical Biology, 260(3):379–385, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002410.

Beardmore:2001:SDT

[BW01] Ivana Beardmore and K. A. Jane White. Spreading disease through social groupings in competition. Journal of Theoretical Biology, 212(2):253–269, September 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923689.

Bello:2006:NPS

[BW06] Y. Bello and D. Waxman. Near-periodic substitution and the genetic variance induced by environmental change. Journal of Theoretical Biology, 239(2):152–160, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003760.

Bruggeman:2002:MRA

[BWHK02] Frank J. Bruggeman, Hans V. Westerhoff, Jan B. Hoek, and Boris N. Kholodenko. Modular response analysis of cellular regulatory networks. Journal of Theoretical Biology, 218(4):507–520, October 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930961.

Berger-Wolf:2007:CAA

[BWMS07] Tanya Y. Berger-Wolf, Cristopher Moore, and Jared Saia. A computational approach to animal breeding. Journal of Theoretical Biology, 244(3):433–439, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600395X.

Behre:2008:SRM

[BWvK^{+}08] Jörn Behre, Thomas Wilhelm, Axel von Kamp, Eytan Ruppin, and Stefan Schuster. Structural robustness of metabolic networks with respect to multiple knockouts. Journal of Theoretical Biology, 252(3):433–441, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004729.

Byers:2005:CCE

[Bye05] John A. Byers. Chemical constraints on the evolution of olfactory communication channels of moths. Journal of Theoretical Biology, 235(2):199–206, July 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000081.

Byers:2009:MDF

[Bye09] John A. Byers. Modeling distributions of flying insects: Effective attraction radius of pheromone in two and three dimensions. Journal of Theoretical Biology, 256(1):81–89, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004712.

Bywater:2009:MSP

[Byw09] Robert P. Bywater. Membrane-spanning peptides and the origin of life. Journal of Theoretical Biology, 261(3):407–413, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003725.

Baker:2000:JBC

[BZ00] Mitchell B. Baker and Amos Zemel. Josef Bertrand catches some ants: Unbiased random trajectories for the simulation of position effects in groups. Journal of Theoretical Biology, 207(3):299–303, December 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921751.

Belotserkovskii:2004:AOD

[BZ04] Boris P. Belotserkovskii and David A. Zarling. Analysis of a one-dimensional random walk with irreversible losses at each step: applications for protein movement on DNA. Journal of Theoretical Biology, 226(2):195–203, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003321.

Bahramian:2005:EGI

[BZ05a] M. Bahman Bahramian and Helmut Zarbl. Erratum to “GENE impedence: a natural process for control of gene expression and the origin of RNA interference”: [J. Theor. Biol. 233 (2005) 301–314]. Journal of Theoretical Biology, 236(1):1, September 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500144X. See [BZ05b].

Bahramian:2005:GIN

[BZ05b] M. Bahman Bahramian and Helmut Zarbl. GENE impedance: a natural process for control of gene expression and the origin of RNA interference. Journal of Theoretical Biology, 233(3):301–314, April 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004783. See erratum [BZ05a].

Bajec:2005:SFW

[BZM05] Iztok Lebar Bajec, Nikolaj Zimic, and Miha Mraz. Simulating flocks on the wing: the fuzzy approach. Journal of Theoretical Biology, 233(2):199–220, March 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004746.

Cherry:2000:HMB

[CA00] Joshua L. Cherry and Frederick R. Adler. How to make a biological switch. Journal of Theoretical Biology, 203(2):117–133, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910683.

Chareyron:2009:MIC

[CA09] S. Chareyron and M. Alamir. Mixed immunotherapy and chemotherapy of tumors: Feedback design and model updating schemes. Journal of Theoretical Biology, 258(3):444– 454, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003536.

Corson:2009:SLV

[CABB09] Francis Corson, Mokhtar Adda-Bedia, and Arezki Boudaoud. In silico leaf venation networks: Growth and reorganization driven by mechanical forces. Journal of Theoretical Biology, 259(3):440–448, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002148.

Cubo:2002:HDT

[CACC02] J. Cubo, D. Azagra, A. Casinos, and J. Castanet. Heterochronic detection through a function for the ontogenetic variation of bone shape. Journal of Theoretical Biology, 215(1):57– 66, March 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924944.

Chen:2003:MGH

[CAF03] Wei-Yin Chen, Phanidhar R. Annamreddy, and L. T. Fan. Modeling growth of a heterogeneous tumor. Journal of Theoretical Biology, 221(2):205–227, March 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931845.

Chowell:2006:TDG

[CAHH06] G. Chowell, C. E. Ammon, N. W. Hengartner, and J. M. Hyman. Transmission dynamics of the great influenza pandemic of 1918 in Geneva, Switzerland: Assessing the effects of hypothetical interventions. Journal of Theoretical Biology, 241(2):193–204, July 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005096.

Calander:2006:PCC

[Cal06] Nils Calander. Propensity of a circadian clock to adjust to the 24h day–night light cycle and its sensitivity to molecular noise. Journal of Theoretical Biology, 241(4):716–724, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000099.

Campbell:2003:LBP

[Cam03] R. B. Campbell. A logistic branching process for population genetics. Journal of Theoretical Biology, 225(2):195– 203, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303002388.

Canessa:2007:MBM

[Can07] Enrique Canessa. Modeling of body mass index by Newton’s second law. Journal of Theoretical Biology, 248(4):646–656, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002901.

Cartling:2002:CNI

[Car02a] Bo Cartling. Control of neural information transmission by synaptic dynamics. Journal of Theoretical Biology, 214(2):275–292, January 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924622.

Cartwright:2002:LTP

[Car02b] Julyan H. E. Cartwright. Labyrinthine Turing pattern formation in the cerebral cortex. Journal of Theoretical Biology, 217(1):97–103, July 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930122.

Chaves:2005:RFB

[CAS05] Madalena Chaves, Réka Albert, and Eduardo D. Sontag. Robustness and fragility of Boolean models for genetic regulatory networks. Journal of Theoretical Biology, 235(3):431–449, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000500.

Cox:2000:GTA

[CB00] Melissa D. Cox and Guy B. Blanchard. Gaseous templates in ant nests. Journal of Theoretical Biology, 204(2):223–238, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920101.

Chen:2005:SMN

[CB05] Wei-Yin Chen and Sankar Bokka. Stochastic modeling of nonlinear epidemiology. Journal of Theoretical Biology, 234(4):455–470, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005843.

Chu:2007:OCE

[CB07a] Dominique Chu and Ian C. Blomfield. Orientational control is an efficient control mechanism for phase switching in the E. coli fim system. Journal of Theoretical Biology, 244(3):541–551, February 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600364X.

Clark:2007:SLP

[CB07b] Ben Clark and Seth Bullock. Shedding light on plant competition: Modelling the influence of plant morphology on light capture (and vice versa). Journal of Theoretical Biology, 244(2):208–217, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003286.

Calenge:2008:GFS

[CB08a] Clément Calenge and Mathieu Basille. A general framework for the statistical exploration of the ecological niche. Journal of Theoretical Biology, 252(4):674–685, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001008.

Cornish-Bowden:2008:RH

[CB08b] Athel Cornish-Bowden. Reinhart Heinrich (1946–2006). Journal of Theoretical Biology, 252(3):377–378, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002603.

Cornish-Bowden:2008:RHA

[CB08c] Athel Cornish-Bowden. Reinhart Heinrich (1946–2006): An annotated bibliography. Journal of Theoretical Biology, 252 (3):379–387, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519307005632.

Chakraborty:2007:MMP

[CBB07] Saikat Chakraborty, Vemuri Balakotaiah, and Akhil Bidani. Multiscale model for pulmonary oxygen uptake and its application to quantify hypoxemia in hepatopulmonary syndrome. Journal of Theoretical Biology, 244(2):190–207, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003067.

Clayton:2001:REC

[CBBH01] Richard H. Clayton, Andrew Bailey, Vadim N. Biktashev, and Arun V. Holden. Re-entrant cardiac arrhythmias in computational models of long QT myocardium. Journal of Theoretical Biology, 208(2):215–225, January 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922124.

Cornish-Bowden:2008:SOO

[CBC08] Athel Cornish-Bowden and María Luz Cárdenas. Selforganization at the origin of life. Journal of Theoretical Biology, 252(3):411–418, June 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003566.

Cosner:2009:EHM

[CBC^{+}09] C. Cosner, J. C. Beier, R. S. Cantrell, D. Impoinvil, L. Kapitanski, M. D. Potts, A. Troyo, and S. Ruan. The effects of human movement on the persistence of vector-borne diseases. Journal of Theoretical Biology, 258(4):550–560, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000757.

Craciun:2005:DSM

[CBF05] Gheorghe Craciun, Anthony Brown, and Avner Friedman. A dynamical system model of neurofilament transport in axons. Journal of Theoretical Biology, 237(3):316–322, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001827.

Cornish-Bowden:2002:RSA

[CBH02] Athel Cornish-Bowden and Jan-Hendrik S. Hofmeyr. The role of stoichiometric analysis in studies of metabolism: An example. Journal of Theoretical Biology, 216(2):179–191, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925476.

Childs:2004:PLD

[CBR04] Dylan Z. Childs, Michael B. Bonsall, and Mark Rees. Periodic local disturbance in host-parasitoid metapopulations: host suppression and parasitoid persistence. Journal of Theoretical Biology, 227(1):13–23, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002935.

Cantrell:2001:SHC

[CC01a] R. S. Cantrell and Chris Cosner. Spatial heterogeneity and critical patch size: Area effects via diffusion in closed environments. Journal of Theoretical Biology, 209(2):161–171, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930092246X.

Chen:2001:GSL

[CC01b] Xin Chen and Joel E. Cohen. Global stability, local stability and permanence in model food Webs. Journal of Theoretical Biology, 212(2):223–235, September 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923707.

Chowdhary:2001:BOT

[CC01c] Aasim G. Chowdhary and John H. Challis. The biomechanics of an overarm throwing task: a simulation model examination of optimal timing of muscle activations. Journal of Theoretical Biology, 211(1):39–53, July 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192332X.

Cai:2006:PMP

[CC06a] Yu-Dong Cai and Kuo-Chen Chou. Predicting membrane protein type by functional domain composition and pseudo-amino acid composition. Journal of Theoretical Biology, 238(2):395–400, January 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002547.

Castellano:2006:RSS

[CC06b] Sergio Castellano and Paolo Cermelli. Reconciling sexual selection to species recognition: a process-based model of mating decision. Journal of Theoretical Biology, 242(3):529–538, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001408.

Caputo:2008:DMT

[CC08] Michele Caputo and Cesare Cametti. Diffusion with memory in two cases of biological interest. Journal of Theoretical Biology, 254(3):697–703, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003330.

Chung:2009:ECS

[CC09a] C. A. Chung and Chih-Yuan Chen. The effect of cell sedimentation on measuring chondrocyte population migration using a Boyden chamber. Journal of Theoretical Biology, 261(4):610–625, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004111.

Crowder:2009:CRS

[CC09b] David W. Crowder and Yves Carrière. Comparing the refuge strategy for managing the evolution of insect resistance under different reproductive strategies. Journal of Theoretical Biology, 261(3):423–430, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003877.

Chen:2008:RMM

[CCC08] Bor-Sen Chen, Chia-Hung Chang, and Yung-Jen Chuang. Robust model matching control of immune systems under environmental disturbances: Dynamic game approach. Journal of Theoretical Biology, 253(4):824–837, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002087.

Cordonnier:2006:ECC

[CCF06] Thomas Cordonnier, Benoît Courbaud, and Alain Franc. The effect of colonization and competition processes on the relation between disturbance and diversity in plant communities. Journal of Theoretical Biology, 243(1):1–12, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002244.

Cabanes:2002:TKS

[CCGC02] Juana Cabanes, Soledad Chazarra, and Francisco García-Carmona. Tyrosinase kinetics: a semi-quantitative model of the mechanism of oxidation of monohydric and dihydric phenolic substrates — reply. Journal of Theoretical Biology, 214(2):321–325, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924634.

Colijn:2007:EHD

[CCM07] Caroline Colijn, Ted Cohen, and Megan Murray. Emergent heterogeneity in declining tuberculosis epidemics. Journal of Theoretical Biology, 247(4):765–774, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700197X.

Chang:2000:IVD

[CCP^{+}00] Kuo-Chu Chang, Tong-J Chen, Ying-I Peng, Te-Hei Li, and Yung-Zu Tseng. Impaired vascular dynamics in normotensive diabetic rats induced by streptozotocin: Tapered T-tube model analysis. Journal of Theoretical Biology, 204(3):371– 380, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920216.

Cloutier:2009:KMM

[CCT^{+}09] Mathieu Cloutier, Jingkui Chen, Frithjof Tatge, Virginie McMurray-Beaulieu, Michel Perrier, and Mario Jolicoeur. Kinetic metabolic modelling for the control of plant cells cytoplasmic phosphate. Journal of Theoretical Biology, 259 (1):118–131, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309000976.

Chattopadhyay:2008:PAT

[CCV08] Joydev Chattopadhyay, Samrat Chatterjee, and Ezio Venturino. Patchy agglomeration as a transition from monospecies to recurrent plankton blooms. Journal of Theoretical Biology, 253(2):289–295, July 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001215.

Chen:2008:PPS

[CCZC08] Chao Chen, Li-Xuan Chen, Xiao-Yong Zou, and Pei-Xiang Cai. Predicting protein structural class based on multi-features fusion. Journal of Theoretical Biology, 253(2):388–392, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001227.

Colosimo:2000:SFB

[CD00] Alfredo Colosimo and Aldo De Luca. Special factors in biological strings. Journal of Theoretical Biology, 204(1):29–46, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910774.

Cinquin:2002:PNF

[CD02] Olivier Cinquin and Jacques Demongeot. Positive and negative feedback: Striking a balance between necessary antagonists. Journal of Theoretical Biology, 216(2):229–241, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925440.

Cinquin:2005:HDS

[CD05] Olivier Cinquin and Jacques Demongeot. High-dimensional switches and the modelling of cellular differentiation. Journal of Theoretical Biology, 233(3):391–411, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004990.

Croisier:2007:CBA

[CD07] H. Croisier and P. C. Dauby. Continuation and bifurcation analysis of a periodically forced excitable system. Journal of Theoretical Biology, 246(3):430–448, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000318.

Clerc:2009:SMP

[CDB09] Thomas Clerc, Anthony C. Davison, and Louis-Félix Bersier. Stochastic modelling of prey depletion processes. Journal of Theoretical Biology, 259(3):523–532, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001866.

Cai:2002:ITA

[CDDW02] Yudong Cai, C. T. J. Dodson, Andrew J. Doig, and Olaf Wolkenhauer. Information-theoretic analysis of protein sequences shows that amino acids self-cluster. Journal of Theoretical Biology, 218(4):409–418, October 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930882.

Cavalcanti:2000:CAT

[CDF00] André R. O. Cavalcanti, Benício De Barros Neto, and Ricardo Ferreira. On the classes of aminoacyl-tRNA synthetases and the error minimization in the genetic code. Journal of Theoretical Biology, 204(1):15–20, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910828.

Chao:2004:SMC

[CDFP04] Dennis L. Chao, Miles P. Davenport, Stephanie Forrest, and Alan S. Perelson. A stochastic model of cytotoxic T cell responses. Journal of Theoretical Biology, 228(2):227–240, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000050.

Christensen:2002:TNM

[CDHJ02] Kim Christensen, Simone A. Di Collobiano, Matt Hall, and Henrik J. Jensen. Tangled nature: a model of evolutionary ecology. Journal of Theoretical Biology, 216(1):73–84, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925300.

Campos:2004:GPP

[CdOWS04] Paulo R. A. Campos, Viviane M. de Oliveira, Günter P. Wagner, and Peter F. Stadler. Gene phylogenies and protein-protein interactions: possible artifacts resulting from shared protein interaction partners. Journal of Theoretical Biology, 231(2):197–202, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002875.

Christianou:2005:KSV

[CE05] Maria Christianou and Bo Ebenman. Keystone species and vulnerable species in ecological communities: strong or weak interactors? Journal of Theoretical Biology, 235(1):95–103, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000020.

Cherry:2008:PTH

[CE08] Elizabeth M. Cherry and Steven J. Evans. Properties of two human atrial cell models in tissue: Restitution, memory, propagation, and reentry. Journal of Theoretical Biology, 254(3):674–690, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003391.

Cappuccio:2007:OII

[CEA07] Antonio Cappuccio, Moran Elishmereni, and Zvia Agur. Optimization of interleukin-21 immunotherapeutic strategies. Journal of Theoretical Biology, 248(2):259–266, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002469.

Canals:2005:ALR

[CEP05] R. M. Canals, L. S. Emeterio, and J. Peralta. Autotoxicity in lolium rigidum: analyzing the role of chemically mediated interactions in annual plant populations. Journal of Theoretical Biology, 235(3):402–407, August 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000470.

Cooper:2007:OFI

[CF07] William E. Cooper and William G. Frederick. Optimal flight initiation distance. Journal of Theoretical Biology, 244(1):59–67, January 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002943.

Chisholm:2009:CSI

[CF09] Ryan A. Chisholm and Elise Filotas. Critical slowing down as an indicator of transitions in two-species models. Journal of Theoretical Biology, 257(1):142–149, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005870.

Chowell:2003:SOO

[CFCGCC03] G. Chowell, P. W. Fenimore, M. A. Castillo-Garsow, and C. Castillo-Chavez. SARS outbreaks in Ontario, Hong Kong and Singapore: the role of diagnosis and isolation as a control mechanism. Journal of Theoretical Biology, 224(1):1–8, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002285.

Cai:2006:ULC

[CFLC06] Yu-Dong Cai, Kai-Yan Feng, Wen-Cong Lu, and Kuo-Chen Chou. Using LogitBoost classifier to predict protein structural classes. Journal of Theoretical Biology, 238(1):172–176, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002523.

Clarke:2004:NPG

[CFM04] Bertrand Clarke, Glenn Fawcett, and Jay E. Mittenthal. Netscan: a procedure for generating reaction networks by size. Journal of Theoretical Biology, 230(4):591–602, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002462.

Cressman:2003:ESL

[CG03] Ross Cressman and József Garay. Evolutionary stability in Lotka–Volterra systems. Journal of Theoretical Biology, 222(2):233–245, May 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000328.

Casagrandi:2006:IDP

[CG06] Renato Casagrandi and Marino Gatto. The intermediate dispersal principle in spatially explicit metapopulations. Journal of Theoretical Biology, 239(1):22–32, March 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003115.

Casellas:2008:ICD

[CGF^{+}08] E. Casellas, J. Gautrais, R. Fournier, S. Blanco, M. Combe, V. Fourcassié, G. Theraulaz, and C. Jost. From individual to collective displacements in heterogeneous environments. Journal of Theoretical Biology, 250(3):424–434, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004973.

Cressman:2001:ESC

[CGH01] Ross Cressman, József Garay, and Josef Hofbauer. Evolutionary stability concepts for N-species frequency-dependent interactions. Journal of Theoretical Biology, 211(1):1–10, July 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923215.

Clote:2005:RST

[CGK^{+}05] Peter Clote, Leszek Gasieniec, Roman Kolpakov, Evangelos Kranakis, and Danny Krizanc. On realizing shapes in the theory of RNA neutral networks. Journal of Theoretical Biology, 236(2):216–227, September 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000962.

Cummins:2007:IBA

[CGKC07] Bree Cummins, Tomás Gedeon, Isaac Klapper, and Ricardo Cortez. Interaction between arthropod filiform hairs in a fluid environment. Journal of Theoretical Biology, 247(2):266–280, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000707.

Christiansen:2005:QAE

[CGS05] Edmund Christiansen, Lars Garby, and Thorkild I. A. Sørensen. Quantitative analysis of the energy requirements for development of obesity. Journal of Theoretical Biology, 234(1):99–106, May 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005442.

Chen:2005:MRC

[CH05a] M. Keith Chen and Marc Hauser. Modeling reciprocation and cooperation in primates: evidence for a punishing strategy. Journal of Theoretical Biology, 235(1):5–12, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304006149.

Codling:2005:SRE

[CH05b] E. A. Codling and N. A. Hill. Sampling rate effects on measurements of correlated and biased random walks. Journal of Theoretical Biology, 233(4):573–588, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005211.

Crowley:2007:ESE

[CH07] Philip H. Crowley and Mary K. Hart. Evolutionary stability of egg trading and parceling in simultaneous hermaphrodites: The chalk bass revisited. Journal of Theoretical Biology, 246(3):420–429, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000471.

Chang:2000:TQM

[Cha00] Tylis Y. Chang. Towards a quantitative model of immunogenicity: Counting pathways in sequence space. Journal of Theoretical Biology, 206(2):255–278, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921222.

Changizi:2001:USL

[Cha01a] Mark A. Changizi. Universal scaling laws for hierarchical complexity in languages, organisms, behaviors and other combinatorial systems. Journal of Theoretical Biology, 211(3):277–295, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930192346X.

Charlesworth:2001:PAS

[Cha01b] Brian Charlesworth. Patterns of age-specific means and genetic variances of mortality rates predicted by the mutationaccumulation theory of ageing. Journal of Theoretical Biology, 210(1):47–65, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922969.

Changizi:2003:RBN

[Cha03] Mark A. Changizi. Relationship between number of muscles, behavioral repertoire size, and encephalization in mammals. Journal of Theoretical Biology, 220(2):157–168, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931250.

Chowell:2004:BRN

[CHCC^{+}04] G. Chowell, N. W. Hengartner, C. Castillo-Chavez, P. W. Fenimore, and J. M. Hyman. The basic reproductive number of Ebola and the effects of public health measures: the cases of Congo and Uganda. Journal of Theoretical Biology, 229(1):119–126, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001092.

Carrero:2006:MCS

[CHdV06] G. Carrero, M. J. Hendzel, and G. de Vries. Modelling the compartmentalization of splicing factors. Journal of Theoretical Biology, 239(3):298–312, April 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003309.

Chelvanayagam:2000:IDA

[Che00] David K. Chelvanayagam. Interpreting the distortion associated with a retinal wholemount. Journal of Theoretical Biology, 205(3):443–455, August 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920782.

Chechetkin:2003:BSS

[Che03a] V. R. Chechetkin. Block structure and stability of the genetic code. Journal of Theoretical Biology, 222(2):177–188, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000250.

Cherry:2003:GSO

[Che03b] Joshua L. Cherry. Genome size and operon content. Journal of Theoretical Biology, 221(3):401–410, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931948.

Chechetkin:2006:GCT

[Che06a] V. R. Chechetkin. Genetic code from tRNA point of view. Journal of Theoretical Biology, 242(4):922–934, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001962.

Chen:2006:ETT

[Che06b] Kevin C. Chen. Effects of tissue trauma on the characteristics of microdialysis zero-net-flux method sampling neurotransmitters. Journal of Theoretical Biology, 238(4):863–881, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002948.

Cheng:2007:LMR

[Che07] Song Cheng. Lorentzian model of roots for understory yellow birch and sugar maple saplings. Journal of Theoretical Biology, 246(2):309–322, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005959.

Chisholm:2007:SSA

[Chi07] Ryan A. Chisholm. Sampling species abundance distributions: Resolving the veil-line debate. Journal of Theoretical Biology, 247(4):600–607, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001889.

Csercsik:2008:SRK

[CHN08] Dávid Csercsik, Katalin M. Hangos, and György M. Nagy. A simple reaction kinetic model of rapid (g protein dependent) and slow (β-arrestin dependent) transmission. Journal of Theoretical Biology, 255(1):119–128, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003822.

Cressman:2006:SRE

[CHR06] Ross Cressman, Josef Hofbauer, and Frank Riedel. Stability of the replicator equation for a single species with a multi-dimensional continuous trait space. Journal of Theoretical Biology, 239(2):273–288, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003887.

Chuine:2000:UMB

[Chu00] Isabelle Chuine. A unified model for budburst of trees. Journal of Theoretical Biology, 207(3):337–347, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921787.

Chu:2008:EGL

[Chu08] Dominique Chu. The evolution of group-level pathogenic traits. Journal of Theoretical Biology, 253(2):355–362, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001306.

Cibert:2008:LAR

[Cib08] Christian Cibert. Are the local adjustments of the relative spatial frequencies of the dynein arms and the β-tubulin monomers involved in the regulation of the “9+2” axoneme? Journal of Theoretical Biology, 253(1):74–89, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000362.

Cinquin:2003:SCR

[Cin03] Olivier Cinquin. Is the somitogenesis clock really cellautonomous? A coupled-oscillator model of segmentation. Journal of Theoretical Biology, 224(4):459–468, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001930.

Cinquin:2006:FTM

[Cin06] Olivier Cinquin. Fast-tracking morphogen diffusion. Journal of Theoretical Biology, 238(3):532–540, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002560.

Chatterjee:2009:SBC

[CIV09] Samrat Chatterjee, Marco Isaia, and Ezio Venturino. Spiders as biological controllers in the agroecosystem. Journal of Theoretical Biology, 258(3):352–362, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006267.

Cuesta:2008:SRD

[CJLS08] J. A. Cuesta, R. Jiménez, H. Lugo, and A. Sánchez. The shared reward dilemma. Journal of Theoretical Biology, 251(2):253–263, March 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005826.

Christianou:2008:CDA

[CK08] Maria Christianou and Giorgos D. Kokkoris. Complexity does not affect stability in feasible model communities. Journal of Theoretical Biology, 253(1):162–169, July 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001148.

Campbell:2006:NMA

[CKE06] H. A. Campbell, J. Z. Klepacki, and S. Egginton. A new method in applying power spectral statistics to examine cardio-respiratory interactions in fish. Journal of Theoretical Biology, 241(2):410–419, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005308.

Couzin:2002:CMS

[CKJ^{+}02] Iain D. Couzin, Jens Krause, Richard James, Graeme D. Ruxton, and Nigel R. Franks. Collective memory and spatial sorting in animal groups. Journal of Theoretical Biology, 218(1):1–11, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930651.

Calvez:2005:CFM

[CKM05] V. Calvez, A. Korobeinikov, and P. K. Maini. Cluster formation for multi-strain infections with cross-immunity. Journal of Theoretical Biology, 233(1):75–83, March 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004631.

Chavarria-Krauser:2004:CGM

[CKS04] A. Chavarría-Krauser and U. Schurr. A cellular growth model for root tips. Journal of Theoretical Biology, 230(1):21–32, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001559.

Chickarmane:2007:ODA

[CKS07] Vijay Chickarmane, Boris N. Kholodenko, and Herbert M. Sauro. Oscillatory dynamics arising from competitive inhibition and multisite phosphorylation. Journal of Theoretical Biology, 244(1):68–76, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002049.

Courteau:2000:OSR

[CL00] Josiane Courteau and Sabin Lessard. Optimal sex ratios in structured populations. Journal of Theoretical Biology, 207(2):159–175, November 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930092160X.

Clarke:2005:SFN

[CL05] Geoff Clarke and Charles J. Lumsden. Scale-free neurodegeneration: cellular heterogeneity and the stretched exponential kinetics of cell death. Journal of Theoretical Biology, 233(4):515–525, April 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005168.

Chen:2007:PAP

[CL07a] Ying-Li Chen and Qian-Zhong Li. Prediction of apoptosis protein subcellular location using improved hybrid approach and pseudo-amino acid composition. Journal of Theoretical Biology, 248(2):377–381, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002500.

Chen:2007:PSL

[CL07b] Ying-Li Chen and Qian-Zhong Li. Prediction of the subcellular location of apoptosis proteins. Journal of Theoretical Biology, 245(4):775–783, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005522.

Chechetkin:2009:LSE

[CL09] V. R. Chechetkin and V. V. Lobzin. Local stability and evolution of the genetic code. Journal of Theoretical Biology, 261(4):643–653, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004123.

Caron-Lormier:2005:AOF

[CLB05] G. Caron-Lormier and H. Berry. Amplification and oscillations in the FAK/Src kinase system during integrin signaling. Journal of Theoretical Biology, 232(2):235–248, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003820.

Clemente:2007:WBE

[Cle07] Roberto Antonio Clemente. On West, Brown and Enquist’s metabolism allometry model. Journal of Theoretical Biology, 249(3):633–637, December 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003803.

Cai:2007:MSM

[CLH07] Anna Q. Cai, Kerry A. Landman, and Barry D. Hughes. Multi-scale modeling of a wound-healing cell migration assay. Journal of Theoretical Biology, 245(3):576–594, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004954.

Cai:2007:CDT

[CLHW07] Anna Q. Cai, Kerry A. Landman, Barry D. Hughes, and Colleen M. Witt. T cell development in the thymus: From periodic seeding to constant output. Journal of Theoretical Biology, 249(2):384–394, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700358X.

Coskun:2007:ACM

[CLM07] Huseyin Coskun, Yi Li, and Michael A. Mackey. Ameboid cell motility: a model and inverse problem, with an application to live cell imaging data. Journal of Theoretical Biology, 244(2):169–179, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003298.

Caron-Lormier:2006:BPA

[CLMMP06] G. Caron-Lormier, J. P. Masson, N. Ménard, and J. S. Pierre. A branching process, its application in biology: Influence of demographic parameters on the social structure in mammal groups. Journal of Theoretical Biology, 238(3):564–574, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002572.

Canadas:2002:CTM

[CLO^{+}02] Patrick Cañadas, Valerie M. Laurent, Christian Oddou, Daniel Isabey, and Sylvie Wendling. A cellular tensegrity model to analyse the structural viscoelasticity of the cytoskeleton. Journal of Theoretical Biology, 218(2):155–173, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930293064X.

Chi:2008:PFM

[CLS08] Xiao-Fei Chi, Xiang-Yang Lou, and Qing-Yao Shu. Progressive fine mapping in experimental populations: An improved strategy toward positional cloning. Journal of Theoretical Biology, 253(4):817–823, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002105.

Chauhan:2008:MMG

[CLW^{+}08] Anuradha Chauhan, Stefan Legewie, Pål O. Westermark, Stephan Lorenzen, and Hanspeter Herzel. A mesoscale model of G1/S phase transition in liver regeneration. Journal of Theoretical Biology, 252(3):465–473, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000349.

Cai:2003:SVM

[CLXC03] Yu-Dong Cai, Xiao-Jun Liu, Xue-Biao Xu, and Kuo-Chen Chou. Support vector machines for prediction of protein domain structural class. Journal of Theoretical Biology, 221(1):115–120, March 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931791.

Cai-Lin:2002:EDS

[CLZZ02] Xu Cai-Lin and Li Zi-Zhen. Effect of diffusion and spatially varying predation risk on the dynamics and equilibrium density of a predator–prey system. Journal of Theoretical Biology, 219(1):73–82, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931103.

Cox:2003:FMF

[CM03] Melissa D. Cox and Mary R. Myerscough. A flexible model of foraging by a honey bee colony: the effects of individual behaviour on foraging success. Journal of Theoretical Biology, 223(2):179–197, July 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000857.

Colijn:2005:MMHa

[CM05a] Caroline Colijn and Michael C. Mackey. A mathematical model of hematopoiesis — I. Periodic chronic myelogenous leukemia. Journal of Theoretical Biology, 237(2):117– 132, November 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305001542.

Colijn:2005:MMHb

[CM05b] Caroline Colijn and Michael C. Mackey. A mathematical model of hematopoiesis: II. Cyclical neutropenia. Journal of Theoretical Biology, 237(2):133–146, November 21, 2005.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001554.

Caruso:2009:EGE

[CM09] T. Caruso and M. Migliorini. Euclidean geometry explains why lengths allow precise body mass estimates in terrestrial invertebrates: The case of oribatid mites. Journal of Theoretical Biology, 256(3):436–440, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005122.

Cristillo:2001:DSR

[CMB^{+}01] A. D. Cristillo, J. R. Mortimer, I. H. Barrette, T. P. Lillicrap, and D. R. Forsdyke. Double-stranded RNA as a notself alarm signal: to evade, most viruses purine-load their RNAs, but some (HTLV-1, Epstein–Barr) pyrimidine-load. Journal of Theoretical Biology, 208(4):475–491, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922331.

Carrillo:2002:CSA

[CMB02] Claudia Carrillo, Michael Mogie, and Nicholas F. Britton. Coexistence of sexual and asexual conspecifics: a cellular automaton model. Journal of Theoretical Biology, 217(3):275–285, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930390.

Cotar:2008:SFP

[CMCH08] Codina Cotar, John M. McNamara, E. J. Collins, and Alasdair I. Houston. Should females prefer to mate with low-quality males? Journal of Theoretical Biology, 254(3):561–567, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003044.

Campos:2008:EDL

[CMF08] Daniel Campos, Vicenç Méndez, and Sergei Fedotov. The effects of distributed life cycles on the dynamics of viral infections. Journal of Theoretical Biology, 254(2):430–438, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002932.

Cassia-Moura:2006:CSH

[CMG06] R. Cassia-Moura and M. A. F. Gomes. A crumpled surface having transverse attractive interactions as a simplified model with biological significance. Journal of Theoretical Biology, 238(2):331–339, January 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002171.

Chaieb:2008:WWT

[CML08] Sahraoui Chaieb, Sárka Málková, and Jyotsana Lal. Why the wrinkling transition in partially polymerized membranes is not universal? Fractal-multifractal hierarchy. Journal of Theoretical Biology, 251(1):60–67, March 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004572.

Cassia-Moura:2000:DAC

[CMPL^{+}00] R. Cassia-Moura, A. Popescu, J. R. S. A. Lima, C. A. S. Andrade, L. S Ventura, K. S. A. Lima, and J. Rinzel. The dynamic activation of colicin ia channels in planar bilayer lipid membrane. Journal of Theoretical Biology, 206(2):235–241, September 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921246.

Collier:2000:CCM

[CMS^{+}00] J. R. Collier, D. Mcinerney, S. Schnell, P. K. Maini, D. J. Gavaghan, P. Houston, and C. D. Stern. A cell cycle model for somitogenesis: Mathematical formulation and numerical simulation. Journal of Theoretical Biology, 207(3):305–316, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921726.

Chiorino:2001:DRC

[CMTU01] G. Chiorino, J. A. J. Metz, D. Tomasoni, and P. Ubezio. Desynchronization rate in cell populations: Mathematical modeling and experimental data. Journal of Theoretical Biology, 208(2):185–199, January 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922136.

Changizi:2002:SDN

[CMW02] M. A. Changizi, M. A. Mcdannald, and D. Widders. Scaling of differentiation in networks: Nervous systems, organisms, ant colonies, ecosystems, businesses, universities, cities, electronic circuits, and legos. Journal of Theoretical Biology, 218(2):215–237, September 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930705.

Cui:2008:SRL

[CMW08] Jingan Cui, Xiaoxia Mu, and Hui Wan. Saturation recovery leads to multiple endemic equilibria and backward bifurcation. Journal of Theoretical Biology, 254(2):275–283, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002488.

Castro:2005:PCD

[COdAM05] Mauro A. A. Castro, Tor T. G. Onsten, Rita M. C. de Almeida, and José C. F. Moreira. Profiling cytogenetic diversity with entropy-based karyotypic analysis. Journal of Theoretical Biology, 234(4):487–495, June 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400606X.

Cogan:2006:EPF

[Cog06] N. G. Cogan. Effects of persister formation on bacterial response to dosing. Journal of Theoretical Biology, 238(3):694–703, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002754.

Cogan:2007:ITH

[Cog07] N. G. Cogan. Incorporating toxin hypothesis into a mathematical model of persister formation and dynamics. Journal of Theoretical Biology, 248(2):340–349, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002639.

Cooper:2001:RRM

[Coo01] Stephen Cooper. Revisiting the relationship of the mammalian g1 phase to cell differentiation. Journal of Theoretical Biology, 208(4):399–402, February 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922288.

Cordero:2005:EOS

[Cor05] Carlos Cordero. The evolutionary origin of signa in female lepidoptera: natural and sexual selection hypotheses. Journal of Theoretical Biology, 232(3):443–449, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004199.

Cheer:2001:CFD

[COS01] A. Y. Cheer, Y. Ogami, and S. L. Sanderson. Computational fluid dynamics in the oral cavity of ram suspensionfeeding fishes. Journal of Theoretical Biology, 210(4):463–474, June 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923252.

Covert:2003:CBM

[CP03] Markus W. Covert and Bernhard O. Palsson. Constraintsbased models: Regulation of gene expression reduces the steady-state solution space. Journal of Theoretical Biology, 221(3):309–325, April 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303930712.

Castiglione:2007:CIM

[CP07] F. Castiglione and B. Piccoli. Cancer immunotherapy, mathematical modeling and optimal control. Journal of Theoretical Biology, 247(4):723–732, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001853.

Chapron:2000:TAI

[CPC^{+}00] Yves Chapron, Jean-Michel Peyrin, Serge Crouzy, Alexandre Jaegly, and Dominique Dormont. Theoretical analysis of the implication of PrP in neuronal death during transmissible subacute spongiform encephalopathies: Hypothesis of a PrP oligomeric channel. Journal of Theoretical Biology, 204(1):103–111, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300910919.

Cowan:2009:EMA

[CPG09] Christopher B. Cowan, Dhara A. Patel, and Theresa A. Good. Exploring the mechanism of β-amyloid toxicity attenuation by multivalent sialic acid polymers through the use of mathematical models. Journal of Theoretical Biology, 258(2):189–197, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000538.

Cantalapiedra:2009:RME

[CPM^{+}09] Inma R. Cantalapiedra, Angelina Peñaranda, Lluis Mont, Josep Brugada, and Blas Echebarria. Reexcitation mechanisms in epicardial tissue: Role of I_{to} density heterogeneities and I_{Na} inactivation kinetics. Journal of Theoretical Biology, 259(4):850–859, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001933.

Crauste:2008:ASR

[CPMG^{+}08] Fabien Crauste, Laurent Pujo-Menjouet, Stéphane Génieys, Clément Molina, and Olivier Gandrillon. Adding self-renewal in committed erythroid progenitors improves the biological relevance of a mathematical model of erythropoiesis. Journal of Theoretical Biology, 250(2):322–338, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004778.

Cornillon:2000:AME

[CPR00] Pierre-André Cornillon, Dominique Pontier, and Marie-Joëlle Rochet. Autoregressive models for estimating phylogenetic and environmental effects: Accounting for within-species variations. Journal of Theoretical Biology, 202(4):247–256, February 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910408.

C:2003:QED

[CQLV^{+}03] Raúl Cañas C., Roberto A. Quiroz, Carlos León-Velarde, Adolfo Posadas, and Javier Osorio. Quantifying energy dissipation by grazing animals in harsh environments. Journal of Theoretical Biology, 225(3):351–359, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002601.

Cousens:2001:WWP

[CR01] R. D. Cousens and A. A. Rawlinson. When will plant morphology affect the shape of a seed dispersal “kernel”? Journal of Theoretical Biology, 211(3):229–238, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923410.

Casagrandi:2002:ALB

[CR02] Renato Casagrandi and Sergio Rinaldi. On the aperiodic locomotor behavior of halobacterium salinarium under periodic light stimuli. Journal of Theoretical Biology, 214(4):647–656, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924828.

Couturier:2005:HMP

[CRB05] Jacob P. Couturier and Robert S. Root-Bernstein. HIV may produce inhibitory microRNAs (miRNAs) that block production of CD28, CD4 and some interleukins. Journal of Theoretical Biology, 235(2):169–184, July 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000068.

Cai:2004:ASP

[CRJC04] Yu-Dong Cai, Pong-Wong Ricardo, Chih-Hung Jen, and Kuo-Chen Chou. Application of SVM to predict membrane protein types. Journal of Theoretical Biology, 226(4):373–376, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003400.

Chu:2006:ECM

[CRL06] Dominique Chu, Jonathan Rowe, and Hoong-Chien Lee. Evaluation of the current models for the evolution of bacterial DNA uptake signal sequences. Journal of Theoretical Biology, 238(1):157–166, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002420.

Ciupe:2007:MMA

[CRNP07] Stanca M. Ciupe, Ruy M. Ribeiro, Patrick W. Nelson, and Alan S. Perelson. Modeling the mechanisms of acute hepatitis B virus infection. Journal of Theoretical Biology, 247(1):23– 35, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000938.

Crowley:2000:HDM

[Cro00] Philip H. Crowley. Hawks, doves, and mixed-symmetry games. Journal of Theoretical Biology, 204(4):543–563, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092037X.

Clearwater:2008:MFM

[CRR08] J. M. Clearwater, C. J. Rennie, and P. A. Robinson. Mean field model of acetylcholine mediated dynamics in the thalamocortical system. Journal of Theoretical Biology, 255(3):287–298, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004232.

Cobbold:2000:MMN

[CS00] C. A. Cobbold and J. A. Sherratt. Mathematical modelling of nitric oxide activity in wound healing can explain keloid and hypertrophic scarring. Journal of Theoretical Biology, 204(2):257–288, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920125.

Corander:2002:UAJ

[CS02] Jukka Corander and Mikko J. Sillanpää. A unified approach to joint modeling of multiple quantitative and qualitative traits in gene mapping. Journal of Theoretical Biology, 218(4):435–446, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930900.

Corominas:2006:NTS

[CS06] Bernat Corominas and Ricard V. Solé. Network topology and self-consistency in language games. Journal of Theoretical Biology, 241(2):438–441, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005084.

Cooper:2007:MRM

[CS07] Stephen Cooper and Kerby Shedden. Microarrays and the relationship of mRNA variation to protein variation during the cell cycle. Journal of Theoretical Biology, 249(3):574–581, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004031.

Changizi:2008:XRV

[CS08] Mark A. Changizi and Shinsuke Shimojo. “X-ray vision” and the evolution of forward-facing eyes. Journal of Theoretical Biology, 254(4):756–767, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003585.

Changizi:2009:RHC

[CS09] Mark Changizi and Shinsuke Shimojo. Response to H. C. Howland, “Orbital orientation is not visual orientation”. Journal of Theoretical Biology, 257(3):524–525, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006759.

Camacho:2007:QAL

[CSA07] J. Camacho, D. B. Stouffer, and L. A. N. Amaral. Quantitative analysis of the local structure of food webs. Journal of Theoretical Biology, 246(2):260–268, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306006011.

Catalano:2009:SST

[CSB^{+}09] Ralph A. Catalano, Katherine Saxton, Tim Bruckner, Sidra Goldman, and Elizabeth Anderson. A sex-specific test of selection in utero. Journal of Theoretical Biology, 257(3):475–479, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006504.

Capps:2003:MNC

[CSC03] Graham J. Capps, David C. Samuels, and Patrick F. Chinnery. A model of the nuclear control of mitochondrial DNA replication. Journal of Theoretical Biology, 221(4):565–583, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932073.

Chaves:2004:SSR

[CSD04] Madalena Chaves, Eduardo D. Sontag, and Robert J. Dinerstein. Steady-states of receptor-ligand dynamics: a theoretical framework. Journal of Theoretical Biology, 227(3):413–428, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004399.

Cromer:2009:LRC

[CSD09] Deborah Cromer, Jaroslav Stark, and Miles P. Davenport. Low red cell production may protect against severe anemia during a malaria infection — Insights from modeling. Journal of Theoretical Biology, 257(4):533–542, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006553.

Chattopadhayay:2001:PMD

[CSFH^{+}01] Joydev Chattopadhayay, Ramrup Sarkar, Maria Elena Fritzsche-Hoballah, Ted C. J. Turlings, and Louis-Félix Bersier. Parasitoids may determine plant fitness — a mathematical model based on experimental data. Journal of Theoretical Biology, 212(3):295–302, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923744.

Chattopadhayay:2002:TPP

[CSM02] J. Chattopadhayay, R. R. Sarkar, and S. Mandal. Toxinproducing plankton may act as a biological control for planktonic blooms — field study and mathematical modelling. Journal of Theoretical Biology, 215(3):333–344, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192510X.

Crowley:2005:OCF

[CSM05] PhilipH. Crowley, Christopher R. Stieha, and D. Nicholas McLetchie. Overgrowth competition, fragmentation and sexratio dynamics: a spatially explicit, sub-individual-based model. Journal of Theoretical Biology, 233(1):25–42, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004473.

Civelekoglu-Scholey:2005:MCT

[CSON^{+}05] G. Civelekoglu-Scholey, A. Wayne Orr, I. Novak, J.-J. Meister, M. A. Schwartz, and A. Mogilner. Model of coupled transient changes of rac, rho, adhesions and stress fibers alignment in endothelial cells responding to shear stress. Journal of Theoretical Biology, 232(4):569–585, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004485.

Covert:2001:RGE

[CSP01] Markus W. Covert, Christophe H. Schilling, and Bernhard Palsson. Regulation of gene expression in flux balance models of metabolism. Journal of Theoretical Biology, 213(1):73–88, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924051.

Chalub:2006:EN

[CSP06] F. A. C. C. Chalub, F. C. Santos, and J. M. Pacheco. The evolution of norms. Journal of Theoretical Biology, 241(2):233– 240, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005114.

Carletti:2008:SCE

[CSP^{+}08] T. Carletti, R. Serra, I. Poli, M. Villani, and A. Filisetti. Sufficient conditions for emergent synchronization in protocell models. Journal of Theoretical Biology, 254(4):741–751, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003603. See corri-

gendum [CSP^{+}09].

Carletti:2009:CSC

[CSP^{+}09] T. Carletti, R. Serra, I. Poli, M. Villani, and A. Filisetti. Corrigendum to “Sufficient conditions for emergent synchronization in protocell models” [J. Theor. Biol. 254 (2008) 741–751]. Journal of Theoretical Biology, 257(4):704, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006735. See [CSP^{+}08].

Casal:2005:ABM

[CSR^{+}05] Arancha Casal, Cenk Sumen, Timothy E. Reddy, Mark S. Alber, and Peter P. Lee. Agent-based modeling of the context dependency in T cell recognition. Journal of Theoretical Biology, 236(4):376–391, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519305001281. See corrigendum [SRAL12].

Crosland:2006:SPS

[CSS06] Michael W. J. Crosland, John S. Scheibe, and Nan-Yao Su. Soldier production strategy in lower termites: a simpler alternative hypothesis. Journal of Theoretical Biology, 239(1):123– 127, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003176.

Christiansen:2008:FMA

[CSS08] Edmund Christiansen, Andrew Swann, and Thorkild I. A. Sørensen. Feedback models allowing estimation of thresholds for self-promoting body weight gain. Journal of Theoretical Biology, 254(4):731–736, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003561.

Cai:2008:OCU

[CSW^{+}08] Yujie Cai, Jun Sun, Jie Wang, Yanrui Ding, Na Tian, Xiangru Liao, and Wenbo Xu. Optimizing the codon usage of synthetic gene with QPSO algorithm. Journal of Theoretical Biology, 254(1):123–127, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002373.

Cateau:2002:KAM

[CT02] Hideyuki Câteau and Shigeru Tanaka. Kinetic analysis of multisite phosphorylation using analytic solutions to Michaelis– Menten equations. Journal of Theoretical Biology, 217(1):1– 14, July 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930249.

Castro:2006:AMT

[CT06] Laureano Castro and Miguel A. Toro. Assortative mating through a mechanism of sexual selection. Journal of Theoretical Biology, 243(3):386–392, December 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002931.

Castiglione:2005:CMI

[CTB^{+}05] Filippo Castiglione, Federico Toschi, Massimo Bernaschi, Sauro Succi, Roberta Benedetti, Brunangelo Falini, and Arcangelo Liso. Computational modeling of the immune response to tumor antigens. Journal of Theoretical Biology, 237(4):390–400, December 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001888.

Cornelissen:2009:PSF

[CTB09] Piers L. Cornelissen, Martin J. Toveé, and Melissa Bateson. Patterns of subcutaneous fat deposition and the relationship between body mass index and waist-to-hip ratio: Implications for models of physical attractiveness. Journal of Theoretical Biology, 256(3):343–350, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005031.

Cui:2006:SDT

[CTS06] Jing’an Cui, Yasuhiro Takeuchi, and Yasuhisa Saito. Spreading disease with transport-related infection. Journal of Theoretical Biology, 239(3):376–390, April 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500336X.

Chernyshev:2008:ELI

[CTS^{+}08] Andrey V. Chernyshev, Peter A. Tarasov, Konstantin A. Semianov, Vyacheslav M. Nekrasov, Alfons G. Hoekstra, and Valeri P. Maltsev. Erythrocyte lysis in isotonic solution of ammonium chloride: Theoretical modeling and experimental verification. Journal of Theoretical Biology, 251(1):93–107, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005061.

Chen:2006:UPA

[CTZ^{+}06] Chao Chen, Yuan-Xin Tian, Xiao-Yong Zou, Pei-Xiang Cai, and Jin-Yuan Mo. Using pseudo-amino acid composition and support vector machine to predict protein structural class.

Journal of Theoretical Biology, 243(3):444–448, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002748.

Cubo:2000:PHE

[Cub00] Jorge Cubo. Process heterochronies in endochondral ossification. Journal of Theoretical Biology, 205(3):343–353, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930092054X.

Cudennec:2005:TIA

[Cud05] C. Cudennec. On the transposition and the improvement of an allometric model from river networks to trees: identification of analogous objects, variables and modellings. Journal of Theoretical Biology, 233(2):151–157, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004692.

Cui:2007:SFG

[Cui07] Yuehua Cui. A statistical framework for genome-wide scanning and testing of imprinted quantitative trait loci. Journal of Theoretical Biology, 244(1):115–126, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003018.

Cummings:2000:MPF

[Cum00] Frederick W. Cummings. A model of pattern formation based on signaling pathways. Journal of Theoretical Biology, 207(1):107–116, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921672.

Cummings:2001:ISG

[Cum01] Frederick W. Cummings. The interaction of surface geometry with morphogens. Journal of Theoretical Biology, 212(3):303–313, October 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930192377X.

Currey:2004:IMP

[Cur04] John Currey. Incompatible mechanical properties in compact bone. Journal of Theoretical Biology, 231(4):569–580, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003418.

Colizza:2008:EMM

[CV08] Vittoria Colizza and Alessandro Vespignani. Epidemic modeling in metapopulation systems with heterogeneous coupling pattern: Theory and simulations. Journal of Theoretical Biology, 251(3):450–467, April 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005991.

Castillo:2003:CCP

[CVH03] Derik Castillo and Jorge X. Velasco-Hernández. Coexistence in a competitive parasitoid-host system. Journal of Theoretical Biology, 221(1):61–77, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931717.

Cotton:2007:ITS

[CW07] Samuel Cotton and Claus Wedekind. Introduction of Trojan sex chromosomes to boost population growth. Journal of Theoretical Biology, 249(1):153–161, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003438.

Caraco:2008:FLP

[CW08a] Thomas Caraco and Ing-Nang Wang. Free-living pathogens: Life-history constraints and strain competition. Journal of Theoretical Biology, 250(3):569–579, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005243.

Cunningham:2008:HHP

[CW08b] J. Paul Cunningham and Stuart A. West. How host plant variability influences the advantages to learning: a theoretical model for oviposition behaviour in lepidoptera. Journal of Theoretical Biology, 251(3):404–410, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005681.

Coen:2006:DMS

[CWDM06] Pietro G. Coen, Mark Wilks, Martino Dall’Antonia, and Michael Millar. Detection of multiple-strain carriers: The value of re-sampling. Journal of Theoretical Biology, 240 (1):98–103, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930500370X.

Crother:2007:IDC

[CWJ07] Brian I. Crother, Mary E. White, and Andrew D. Johnson. Inferring developmental constraint and constraint release: Primordial germ cell determination mechanisms as examples. Journal of Theoretical Biology, 248(2):322–330, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002767.

Chen:2009:BDL

[CXM^{+}09] Shu-Yan Chen, Jin Xu, Fernando T. Maestre, Cheng-Jin Chu, Gang Wang, and Sa Xiao. Beyond dual-lattice models: Incorporating plant strategies when modeling the interplay between facilitation and competition along environmental severity gradients. Journal of Theoretical Biology, 258(2):266–273, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000265.

Cao:2009:DEF

[CXZ^{+}09] Huai Cao, Hua-Zhen Xie, Wen Zhang, Kan Wang, Wei Li, and Ci-Quan Liu. Dynamic extended folding: Modeling the RNA secondary structures during co-transcriptional folding. Journal of Theoretical Biology, 261(1):93–99, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003336.

Chang:2007:SSL

[CY07] David B. Chang and Carl S. Young. Simple scaling laws for influenza A rise time, duration, and severity. Journal of Theoretical Biology, 246(4):621–635, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000690.

Czerlinski:2008:MTD

[CY08] George Czerlinski and Tjalling Ypma. Mechanisms of telomerase-dimer catalysis. Journal of Theoretical Biology, 250(3):512–523, February 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005231.

Cui:2009:ZIG

[CY09] Yuehua Cui and Wenzhao Yang. Zero-inflated generalized Poisson regression mixture model for mapping quantitative trait loci underlying count trait with many zeros. Journal of Theoretical Biology, 256(2):276–285, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005183.

Cytrynbaum:2004:PSS

[Cyt04] Eric N. Cytrynbaum. Periodic stimulus and the single cardiac cell-getting more out of 1d maps. Journal of Theoretical Biology, 229(1):69–83, July 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001055.

Cai:2005:PEF

[CZC05] Yu-Dong Cai, Guo-Ping Zhou, and Kuo-Chen Chou. Predicting enzyme family classes by hybridizing gene product composition and pseudo-amino acid composition. Journal of Theoretical Biology, 234(1):145–149, May 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400548X.

Chu:2009:MTF

[CZM09] Dominique Chu, Nicolae Radu Zabet, and Boris Mitavskiy. Models of transcription factor binding: Sensitivity of activation functions to model assumptions. Journal of Theoretical Biology, 257(3):419–429, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006310.

Daido:2002:PDC

[Dai02] Hiroaki Daido. Population dynamics of clock-controlled biological species: Models and why circadian rhythms are circadian. Journal of Theoretical Biology, 217(4):425–442, August 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930293050X.

Dale:2006:PNA

[Dal06] Trevor Dale. Protein and nucleic acid together: a mechanism for the emergence of biological selection. Journal of Theoretical Biology, 240(3):337–342, June 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004431.

Durkee:2003:FCB

[DALP03] J. W. Durkee, P. P. Antich, M. A. Lewis, and R. W. Parkey. A fully coupled binary biochemical reactivediffusion model with analytic solution. Journal of Theoretical Biology, 221(2):163–191, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931705.

Damgaard:2004:DDT

[Dam04] Christian Damgaard. Dynamics in a discrete two-species competition model: coexistence and over-compensation. Journal of Theoretical Biology, 227(2):197–203, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300417X.

Davis:2009:MGC

[Dav09] Brian K. Davis. On mapping the genetic code. Journal of Theoretical Biology, 259(4):860–862, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900215X.

Dawes:2009:MMC

[Daw09] Adriana T. Dawes. A mathematical model of α-catenin dimerization at adherens junctions in polarized epithelial cells. Journal of Theoretical Biology, 257(3):480–488, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006279.

Dietrich:2001:MD

[DB01] Axel Dietrich and Willem Been. Memory and DNA. Journal of Theoretical Biology, 208(2):145–149, January 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922070.

Daniel:2005:MTM

[DB05] R. W. Daniel and C. A. R. Boyd. Membrane transporters as machines: degenerate singularities as a requirement for function. Journal of Theoretical Biology, 232(2):217–221, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003807.

Davidich:2008:TDE

[DB08] Maria Davidich and Stefan Bornholdt. The transition from differential equations to Boolean networks: a case study in simplifying a regulatory network model. Journal of Theoretical Biology, 255(3):269–277, December 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003652.

Donnell:2009:MDT

[DBB09] Peter Donnell, Stephen A. Baigent, and Murad Banaji. Monotone dynamics of two cells dynamically coupled by a voltage-dependent gap junction. Journal of Theoretical Biology, 261(1):120–125, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003257.

Deredec:2008:NSM

[DBBC08] Anne Deredec, Ludek Berec, David S. Boukal, and Franck Courchamp. Are non-sexual models appropriate for predicting the impact of virus-vectored immunocontraception? Journal of Theoretical Biology, 250(2):281–290, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004705.

Daukste:2009:USC

[DBBW09] Liene Daukste, Britta Basse, Bruce C. Baguley, and David J. N. Wall. Using a stem cell and progeny model to illustrate the relationship between cell cycle times of in vivo human tumour cell tissue populations, in vitro primary cultures and the cell lines derived from them. Journal of Theoretical Biology, 260(4):563–571, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002963. See corrigen-

dum [DBBW11].

Daukste:2011:CUS

[DBBW11] Liene Daukste, Britta Basse, Bruce C. Baguley, and David J. N. Wall. Corrigendum to “Using a stem cell and progeny model to illustrate the relationship between cell cycle times of in vivo human tumour cell tissue populations, in vitro primary cultures and the cell lines derived from them” [J. Theor. Biol. 260 (2009) 563–571]. Journal of Theoretical Biology, 289(??):269, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004577. See [DBBW09].

deBivort:2007:MIM

[dBCP^{+}07] Benjamin Lovegren de Bivort, Chun-Chung Chen, Fabrizio Perretti, Giacomo Negro, Thomas M. Philip, and Yaneer Bar-Yam. Metabolic implications for the mechanism of mitochondrial endosymbiosis and human hereditary disorders. Journal of Theoretical Biology, 248(1):26–36, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001993.

Durand:2007:PGP

[DBF07] Eric Durand, Michael G. B. Blum, and Olivier François. Prediction of group patterns in social mammals based on a coalescent model. Journal of Theoretical Biology, 249(2):262–270, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700330X.

Davis:2001:EUA

[DBG01] Stephen Davis, Nicholas Bax, and Peter Grewe. Engineered underdominance allows efficient and economical introgression of traits into pest populations. Journal of Theoretical Biology, 212(1):83–98, September 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923574.

Dronne:2006:MMI

[DBG06] Marie-Aimée Dronne, Jean-Pierre Boissel, and Emmanuel Grenier. A mathematical model of ion movements in grey matter during a stroke. Journal of Theoretical Biology, 240(4):599– 615, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004728.

Dees:2008:PET

[DBGM08] Nathan D. Dees, Sonya Bahar, Ricardo Garcia, and Frank Moss. Patch exploitation in two dimensions: From Daphnia to simulated foragers. Journal of Theoretical Biology, 252(1):69– 76, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000477.

Dokholyan:2000:DDT

[DBHS00] Nikolay V. Dokholyan, Sergey V. Buldyrev, Shlomo Havlin, and H. Eugene Stanley. Distributions of dimeric tandem repeats in non-coding and coding DNA sequences. Journal of Theoretical Biology, 202(4):273–282, February 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910524.

Dellen:2005:COM

[DBR^{+}05] B. K. Dellen, M. J. Barber, M. L. Ristig, J. Hescheler, H. Sauer, and M. Wartenberg. [Ca^{2+}] oscillations in a model of energy-dependent Ca^{2+} uptake by the endoplasmic reticulum. Journal of Theoretical Biology, 237(3):279–290, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001797.

deBeaudrap:2006:MMF

[dBWBP06] P. de Beaudrap, G. Witten, G. Biltz, and E. Perrier. Mechanistic model of fuel selection in the muscle. Journal of Theoretical Biology, 242(1):151–163, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000750.

DellAcqua:2009:SPT

[DC09] Guido Dell’Acqua and Filippo Castiglione. Stability and phase transitions in a mathematical model of duchenne muscular dystrophy. Journal of Theoretical Biology, 260(2): 283–289, September 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309002501.

DeVendittis:2008:AMP

[DCC^{+}08] Emmanuele De Vendittis, Immacolata Castellano, Roberta Cotugno, Maria Rosaria Ruocco, Gennaro Raimo, and Mariorosario Masullo. Adaptation of model proteins from cold to hot environments involves continuous and small adjustments of average parameters related to amino acid composition. Journal of Theoretical Biology, 250(1):156–171, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004249.

Du:2009:SPP

[DCL09] Pufeng Du, Shengjiao Cao, and Yanda Li. SubChlo: Predicting protein subchloroplast locations with pseudo-amino acid composition and the evidence-theoretic k-nearest neighbor (ET-KNN) algorithm. Journal of Theoretical Biology, 261(2):330–335, November 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003683.

Delsanto:2008:MAC

[DCP^{+}08] P. P. Delsanto, C. A. Condat, N. Pugno, A. S. Gliozzi, and M. Griffa. A multilevel approach to cancer growth modeling. Journal of Theoretical Biology, 250(1):16–24, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004523.

Cai:2004:ICT

[dCZJ^{+}04] Yu dong Cai, Guo-Ping Zhou, Chin-Hung Jen, Shuo-Liang Lin, and Kuo-Chen Chou. Identify catalytic triads of serine hydrolases by support vector machines. Journal of Theoretical Biology, 228(4):551–557, June 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400075X.

Dobrolyubov:2002:PTT

[DD02] A. I. Dobrolyubov and G. Douchy. Peristaltic transport as the travelling deformation waves. Journal of Theoretical Biology, 219(1):55–61, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931073.

Dutilh:2003:DEC

[DdB03] Bas E. Dutilh and Rob J. de Boer. Decline in excision circles requires homeostatic renewal or homeostatic death of naive T cells. Journal of Theoretical Biology, 224(3):351–358, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001723.

Dey:2006:EMM

[DDJ06] Sutirth Dey, Sugat Dabholkar, and Amitabh Joshi. The effect of migration on metapopulation stability is qualitatively unaffected by demographic and spatial heterogeneity. Journal of Theoretical Biology, 238(1):78–84, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002158.

DePolavieja:2002:EDE

[De 02a] Gonzalo G. De Polavieja. Errors drive the evolution of biological signalling to costly codes. Journal of Theoretical Biology, 214(4):657–664, February 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924981.

DeSousa:2002:GFE

[De 02b] João Dinis De Sousa. Are genes for fetal over-exploitation outlaws? Journal of Theoretical Biology, 219(4):553–554, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931437.

DeJaegher:2003:EPH

[De 03] Kris De Jaegher. Error-proneness as a handicap signal. Journal of Theoretical Biology, 224(2):139–152, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001358.

Dawes:2006:AFB

[DECEK06] Adriana T. Dawes, G. Bard Ermentrout, Eric N. Cytrynbaum, and Leah Edelstein-Keshet. Actin filament branching and protrusion velocity in a simple 1d model of a motile cell. Journal of Theoretical Biology, 242(2):265–279, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000828.

Dediu:2008:RGB

[Ded08] Dan Dediu. The role of genetic biases in shaping the correlations between languages and genes. Journal of Theoretical Biology, 254(2):400–407, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002762.

Dediu:2009:GBT

[Ded09] Dan Dediu. Genetic biasing through cultural transmission: Do simple Bayesian models of language evolution generalise? Journal of Theoretical Biology, 259(3):552–561, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001647.

Demetrius:2000:TE

[Dem00a] Lloyd Demetrius. Thermodynamics and evolution. Journal of Theoretical Biology, 206(1):1–16, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921064.

Diebner:2000:MTA

[DEM^{+}00b] Hans H. Diebner, Martin Eichner, Louis Molineaux, William E. Collins, Geoffrey M. Jeffery, and Klaus Dietz. Modelling the transition of asexual blood stages of Plasmodium falciparum to gametocytes. Journal of Theoretical Biology, 202(2):113–127, January 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939991041X.

Demetrius:2002:TKP

[Dem02] Lloyd Demetrius. Thermodynamics and kinetics of protein folding: An evolutionary perspective. Journal of Theoretical Biology, 217(3):397–411, August 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930067.

Demetrius:2006:OAS

[Dem06] Lloyd Demetrius. The origin of allometric scaling laws in biology. Journal of Theoretical Biology, 243(4):455–467, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600227X.

Denny:2008:BTP

[Den08] Kevin Denny. Big and tall parents do not have more sons. Journal of Theoretical Biology, 250(4):752–753, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005620.

Dillon:2000:IMI

[DF00] Robert H. Dillon and Lisa J. Fauci. An integrative model of internal axoneme mechanics and external fluid dynamics in ciliary beating. Journal of Theoretical Biology, 207(3):415–430, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921829.

Dumas:2008:MOE

[DF08] André Dumas and James France. Modelling the ontogeny of ectotherms exhibiting indeterminate growth. Journal of Theoretical Biology, 254(1):76–81, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002403.

DiPatti:2009:SRS

[DF09] F. Di Patti and D. Fanelli. A stochastic reaction scheme for drug/metabolite interaction. Journal of Theoretical Biology, 259(2):382–388, July 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001398.

DeAtauri:2002:DCC

[DFC^{+}02] Pedro De Atauri, David A. Fell, Christophe Chassagnole, Dolors Magret, Jean Pierre Mazat, and Marta Cascante. Dependence of control coefficient distribution on the boundaries of a metabolic system: a generalized analysis of the effects of additional input and output reactions to a linear pathway. Journal of Theoretical Biology, 215(2):239–251, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924920.

Du:2008:HPM

[DFCL08] Ning Du, Jintu Fan, Shuo Chen, and Yang Liu. A hydraulicphotosynthetic model based on extended HLH and its application to Coast redwood (Sequoia sempervirens). Journal of Theoretical Biology, 253(2):393–400, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800129X.

deFarias:2008:ATS

[dFG08] Sávio Torres de Farias and Romeu Cardoso Guimarães. Aminoacyl-tRNA synthetase classes and groups in prokaryotes. Journal of Theoretical Biology, 250(2):221–229, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004547.

DeBoer:2001:RCD

[DFP01] Rob J. De Boer, António A. Freitas, and Alan S. Perelson. Resource competition determines selection of B cell repertoires. Journal of Theoretical Biology, 212(3):333–343, October 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923793.

DiPatti:2008:MPT

[DFP^{+}08] F. Di Patti, D. Fanelli, R. S. Pedersen, C. Giuliani, and F. Torricelli. Modelling the pharmacokinetics of tramadol: On the difference between CYP2D6 extensive and poor metabolizers. Journal of Theoretical Biology, 254(3):568–574, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003068.

Dupuy:2005:DBA

[DFSD05] L. Dupuy, T. Fourcaud, A. Stokes, and F. Danjon. A densitybased approach for the modelling of root architecture: application to maritime pine (Pinus pinaster Ait.) root systems. Journal of Theoretical Biology, 236(3):323–334, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001220.

Du:2007:IMH

[DFW^{+}07] Ning Du, Jintu Fan, Huijun Wu, Shuo Chen, and Yang Liu. An improved model of heat transfer through penguin feathers and down. Journal of Theoretical Biology, 248(4):727–735, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002998.

Dattoli:2009:CMD

[DGD^{+}09] G. Dattoli, C. Guiot, P. P. Delsanto, P. L. Ottaviani, S. Pagnutti, and T. S. Deisboeck. Cancer metabolism and the dynamics of metastasis. Journal of Theoretical Biology, 256(3):305–310, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005225.

Dasika:2005:DDE

[DGM05] Madhukar S. Dasika, Anshuman Gupta, and Costas D. Maranas. DEMSIM: a discrete event based mechanistic simulation platform for gene expression and regulation dynamics. Journal of Theoretical Biology, 232(1):55–69, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003509.

Dumont:2009:RCP

[DGS09] E. R. Dumont, I. R. Grosse, and G. J. Slater. Requirements for comparing the performance of finite element models of biological structures. Journal of Theoretical Biology, 256(1):96–103, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004189.

Demetrius:2007:DFI

[DGZ07] Lloyd Demetrius, Volker Matthias Gundlach, and Martin Ziehe. Darwinian fitness and the intensity of natural selection: Studies in sensitivity analysis. Journal of Theoretical Biology, 249(4):641–653, December 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004006.

Drossel:2001:IPP

[DHM01] Barbara Drossel, Paul G. Higgs, and Alan J. Mckane. The influence of predator–prey population dynamics on the longterm evolution of food web structure. Journal of Theoretical Biology, 208(1):91–107, January 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922033.

Dieckmann:2006:ADF

[DHP06] Ulf Dieckmann, Mikko Heino, and Kalle Parvinen. The adaptive dynamics of function-valued traits. Journal of Theoretical Biology, 241(2):370–389, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005266.

Dimitrov:2007:IMB

[DHR^{+}07] Dobromir T. Dimitrov, Thomas G. Hallam, Charles E. Rupprecht, Amy S. Turmelle, and Gary F. McCracken. Integrative models of bat rabies immunology, epizootiology and disease demography. Journal of Theoretical Biology, 245(3):498–509, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005315.

Dimitrov:2008:AMV

[DHRM08] Dobromir T. Dimitrov, Thomas G. Hallam, Charles E. Rupprecht, and Gary F. McCracken. Adaptive modeling of viral diseases in bats with a focus on rabies. Journal of Theoretical Biology, 255(1):69–80, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004219.

Du:2009:EAT

[DHW^{+}09] Qi-Shi Du, Ri-Bo Huang, Cheng-Hua Wang, Xiao-Ming Li, and Kuo-Chen Chou. Energetic analysis of the two controversial drug binding sites of the M2 proton channel in influenza A virus. Journal of Theoretical Biology, 259(1):159– 164, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900112X.

Dai:2009:TCD

[DHYHR09] Xiaofeng Dai, Shannon Healy, Olli Yli-Harja, and Andre S. Ribeiro. Tuning cell differentiation patterns and single cell dynamics by regulating proteins’ functionalities in a toggle switch. Journal of Theoretical Biology, 261(3):441–448, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003932.

DiGiulio:2000:GCO

[Di 00a] Massimo Di Giulio. Genetic code origin and the strength of natural selection. Journal of Theoretical Biology, 205 (4):659–661, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300921155.

DiGiulio:2000:UAL

[Di 00b] Massimo Di Giulio. The universal ancestor lived in a thermophilic or hyperthermophilic environment. Journal of Theoretical Biology, 203(3):203–213, April 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910865.

DiPaolo:2000:ESB

[Di 00c] Ezequiel A. Di Paolo. Ecological symmetry breaking can favour the evolution of altruism in an action-response game. Journal of Theoretical Biology, 203(2):135–152, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910786.

DiGiulio:2001:NUG

[Di 01a] Massimo Di Giulio. The non-universality of the genetic code: the universal ancestor was a progenote. Journal of Theoretical Biology, 209(3):345–349, April 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922714.

DiGiulio:2001:OGC

[Di 01b] Massimo Di Giulio. The origin of the genetic code cannot be studied using measurements based on the PAM matrix because this matrix reflects the code itself, making any such analyses tautologous. Journal of Theoretical Biology, 208(2):141–144, January 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922069.

DiGiulio:2003:ABD

[Di 03a] Massimo Di Giulio. The ancestor of the bacteria domain was a hyperthermophile. Journal of Theoretical Biology, 224(3):277–283, October 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001644.

DiGiulio:2003:UAW

[Di 03b] Massimo Di Giulio. The universal ancestor was a thermophile or a hyperthermophile: Tests and further evidence. Journal of Theoretical Biology, 221(3):425–436, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931973.

DiGiulio:2004:OTM

[Di 04] Massimo Di Giulio. The origin of the tRNA molecule: implications for the origin of protein synthesis. Journal of Theoretical Biology, 226(1):89–93, January 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003230.

DiGiulio:2005:SGC

[Di 05] Massimo Di Giulio. Structuring of the genetic code took place at acidic pH. Journal of Theoretical Biology, 237(2):219–226, November 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001736.

DiGiulio:2006:NEL

[Di 06a] Massimo Di Giulio. Nanoarchaeum equitans is a living fossil. Journal of Theoretical Biology, 242(1):257–260, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000555.

DiGiulio:2006:NMO

[Di 06b] Massimo Di Giulio. The non-monophyletic origin of the tRNA molecule and the origin of genes only after the evolutionary stage of the last universal common ancestor (LUCA). Journal of Theoretical Biology, 240(3):343–352, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004273.

DiGiulio:2008:PTG

[Di 08] Massimo Di Giulio. Permuted tRNA genes of cyanidioschyzon merolae, the origin of the tRNA molecule and the root of the eukarya domain. Journal of Theoretical Biology, 253(3):587–592, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002063.

DiGiulio:2009:MHO

[Di 09] Massimo Di Giulio. A methanogen hosted the origin of the genetic code. Journal of Theoretical Biology, 260(1):77–82, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002574.

DelBiancoFaria:2009:IBP

[DI09] Lucas Del Bianco Faria and Michel Iskin da Silveira Costa. The interplay between predator’s prey preference and environmental heterogeneity in food web long-term stability. Journal of Theoretical Biology, 258(3):339–343, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003251.

Dickson:2008:EUV

[Dic08] Eric S. Dickson. Expected utility violations evolve under status-based selection mechanisms. Journal of Theoretical Biology, 254(3):650–654, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800338X.

Dimitrov:2005:ISC

[Dim05] Alexander G. Dimitrov. Internodal sodium channels ensure active processes under myelin manifesting in depolarizing afterpotentials. Journal of Theoretical Biology, 235(4):451–462, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000512.

Dostalkova:2002:CPP

[DKD02] Iva Dostálková, Pavel Kindlmann, and Anthony F. G. Dixon. Are classical predator–prey models relevant to the real world? Journal of Theoretical Biology, 218(3):323–330, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930808.

DeConde:2005:PTD

[DKLL05] Rob DeConde, Peter S. Kim, Doron Levy, and Peter P. Lee. Post-transplantation dynamics of the immune response to chronic myelogenous leukemia. Journal of Theoretical Biology, 236(1):39–59, September 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000755.

Desper:2004:TCU

[DKS04] Richard Desper, Javed Khan, and Alejandro A. Schäffer. Tumor classification using phylogenetic methods on expression data. Journal of Theoretical Biology, 228(4):477–496, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000700.

Durrett:2000:LPF

[DL00] Rick Durrett and Simon Levin. Lessons on pattern formation from planet WATOR. Journal of Theoretical Biology, 205(2):201–214, July 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920617.

Du:2008:PCU

[DL08] Pufeng Du and Yanda Li. Prediction of C-to-U RNA editing sites in plant mitochondria using both biochemical and evolutionary information. Journal of Theoretical Biology, 253(3):579–586, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001823.

Donaldson:2007:EFR

[DLB07] Matina C. Donaldson, Michael Lachmann, and Carl T. Bergstrom. The evolution of functionally referential meaning in a structured world. Journal of Theoretical Biology, 246(2):225–233, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519307000021.

Diao:2007:CSH

[DLF^{+}07] Yuanbo Diao, Menglong Li, Zinan Feng, Jiajian Yin, and Yi Pan. The community structure of human cellular signaling network. Journal of Theoretical Biology, 247(4):608–615, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001890.

Daudet:2002:GMC

[DLGC02] F. A. Daudet, A. Lacointe, J. P. Gaudillère, and P. Cruiziat. Generalized Münch coupling between sugar and water fluxes for modelling carbon allocation as affected by water status. Journal of Theoretical Biology, 214(3):481–498, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924737.

Dworkin:2003:RRW

[DLM03] Jason P. Dworkin, Antonio Lazcano, and Stanley L. Miller. The roads to and from the RNA world. Journal of Theoretical Biology, 222(1):127–134, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000201.

Devillers:2008:ABM

[DLM08] Hugo Devillers, Jean Raymond Lobry, and Frédéric Menu. An agent-based model for predicting the prevalence of Trypanosoma cruzi I and II in their host and vector populations. Journal of Theoretical Biology, 255(3):307–315, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004463.

DOdorico:2008:BEI

[DLRL08] Paolo D’Odorico, Francesco Laio, Luca Ridolfi, and Manuel T. Lerdau. Biodiversity enhancement induced by environmental noise. Journal of Theoretical Biology, 255(3):332–337, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004682.

Dahari:2007:MHC

[DLRP07] Harel Dahari, Arthur Lo, Ruy M. Ribeiro, and Alan S. Perelson. Modeling hepatitis C virus dynamics: Liver regeneration and critical drug efficacy. Journal of Theoretical Biology, 247(2):371–381, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001117.

Drossel:2000:CSQ

[DM00] Barbara Drossel and Alan Mckane. Competitive speciation in quantitative genetic models. Journal of Theoretical Biology, 204(3):467–478, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920307.

Demongeot:2007:PCR

[DM07] Jacques Demongeot and Andrés Moreira. A possible circular RNA at the origin of life. Journal of Theoretical Biology, 249(2):314–324, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003281.

dOnofrio:2009:IRC

[dM09] Alberto d’Onofrio and Piero Manfredi. Information-related changes in contact patterns may trigger oscillations in the endemic prevalence of infectious diseases. Journal of Theoretical Biology, 256(3):473–478, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005377.

Davidson:2008:EHS

[DMH08] Ross S. Davidson, Glenn Marion, and Michael R. Hutchings. Effects of host social hierarchy on disease persistence. Journal of Theoretical Biology, 253(3):424–433, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001458.

Denton:2002:PFP

[DML02] Michael J. Denton, Craig J. Marshall, and Michael Legge. The protein folds as platonic forms: New support for the pre-Darwinian conception of evolution by natural law. Journal of Theoretical Biology, 219(3):325–342, December 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931280.

DAgata:2007:MAR

[DMO^{+}07] Erika M. C. D’Agata, Pierre Magal, Damien Olivier, Shigui Ruan, and Glenn F. Webb. Modeling antibiotic resistance in hospitals: The impact of minimizing treatment duration. Journal of Theoretical Biology, 249(3):487–499, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307003943.

Dayan:2009:GRR

[DMPP09] Frédéric Dayan, Marc Monticelli, Jacques Pouysségur, and Elisabeth Pécou. Gene regulation in response to graded hypoxia: The non-redundant roles of the oxygen sensors PHD and FIH in the HIF pathway. Journal of Theoretical Biology, 259(2):304–316, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001167.

Dette:2005:RED

[DMPS05] Holger Dette, Viatcheslav B. Melas, Andrey Pepelyshev, and Nikolay Strigul. Robust and efficient design of experiments for the Monod model. Journal of Theoretical Biology, 234(4):537– 550, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304006101.

Drossel:2004:INF

[DMQ04] Barbara Drossel, Alan J. McKane, and Christopher Quince. The impact of nonlinear functional responses on the longterm evolution of food web structure. Journal of Theoretical Biology, 229(4):539–548, August 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001778.

Davidson:2004:BAC

[DMW04] Peter L. Davidson, Peter D. Milburn, and Barry D. Wilson. Biological adaptive control model: a mechanical analogue of multi-factorial bone density adaptation. Journal of Theoretical Biology, 227(2):187–195, March 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004004.

Day:2008:EHS

[DMW08] William H. E. Day, F. R. McMorris, and Mark Wilkinson. Explosions and hot spots in supertree methods. Journal of Theoretical Biology, 253(2):345–348, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001501.

Derrida:2000:GPB

[DMZ00] Bernard Derrida, Susanna C. Manrubia, and Damián H. Zanette. On the genealogy of a population of biparental individuals. Journal of Theoretical Biology, 203(3):303–315, April 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910956.

Dowse:2000:EDC

[DNS00] Harold B. Dowse, Steven Norton, and Bruce D. Sidell. The estimation of the diffusion constant and solubility of O_{2} in tissue using kinetics. Journal of Theoretical Biology, 207(4):531–541, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921842.

Davidson:2000:TIO

[DO00] F. A. Davidson and S. Olsson. Translocation induced outgrowth of fungi in nutrient-free environments. Journal of Theoretical Biology, 205(1):73–84, July 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920459.

Danielsen:2001:DPH

[DO01] M. Danielsen and J. T. Ottesen. Describing the pumping heart as a pressure source. Journal of Theoretical Biology, 212(1):71–81, September 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923483.

Dallon:2004:HCM

[DO04] John C. Dallon and Hans G. Othmer. How cellular movement determines the collective force generated by the Dictyostelium discoideum slug. Journal of Theoretical Biology, 231(2):203–222, November 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002887.

Dorazi:2003:CTA

[Dor03a] Robert Dorazi. Can tRNAs act as antisense RNA? The case of mutA and dnaQ. Journal of Theoretical Biology, 225(3):383–388, December 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002686.

Doriot:2003:SUC

[Dor03b] Pierre-André Doriot. Some unusual considerations about vessel walls and wall stresses. Journal of Theoretical Biology, 221(1):133–141, March 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931821.

Dore:2002:UVI

[DOT02] Jean-Christophe Doré, Tiiu Ojasoo, and Michel Thireau. Using the volumetric indices of telencephalic structures to distinguish Salamandridae and Plethodontidae: Comparison of three statistical methods. Journal of Theoretical Biology, 214(3):427–439, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924671.

Dixit:2004:CPV

[DP04] Narendra M. Dixit and Alan S. Perelson. Complex patterns of viral load decay under antiretroviral therapy: influence of pharmacokinetics and intracellular delay. Journal of Theoretical Biology, 226(1):95–109, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003254.

Dewar:2008:SMU

[DP08] Roderick C. Dewar and Annabel Porté. Statistical mechanics unifies different ecological patterns. Journal of Theoretical Biology, 251(3):389–403, April 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006297.

Dugatkin:2003:EGB

[DPA03] Lee Alan Dugatkin, Michael Perlin, and Ronald Atlas. The evolution of group-beneficial traits in the absence of betweengroup selection. Journal of Theoretical Biology, 220(1):67–74, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931493.

Dugatkin:2005:ARE

[DPA05] Lee Alan Dugatkin, Michael Perlin, and Ronald Atlas. Antibiotic resistance and the evolution of group-beneficial traits. II: a metapopulation model. Journal of Theoretical Biology, 236(4):392–396, October 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500130X.

duPreez:2008:CAK

[dPCP^{+}08] F. B. du Preez, R. Conradie, G. P. Penkler, K. Holm, F. L. J. van Dooren, and J. L. Snoep. A comparative analysis of kinetic models of erythrocyte glycolysis. Journal of Theoretical Biology, 252(3):488–496, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004924.

dePillis:2006:MIC

[dPGR06] L. G. de Pillis, W. Gu, and A. E. Radunskaya. Mixed immunotherapy and chemotherapy of tumors: modeling, applications and biological interpretations. Journal of Theoretical Biology, 238(4):841–862, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002936.

Silva:2000:USE

[DPV00] Maria Luisa Da Silva, José Roberto C. Piqueira, and Jacques M. E. Vielliard. Using Shannon entropy on measuring the individual variability in the rufous-bellied thrush Turdus rufiventris vocal communication. Journal of Theoretical Biology, 207(1):57–64, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921556.

Dai:2007:CJM

[DqLmW07] Qi Dai, Xiao qing Liu, and Tian ming Wang. C(i,j) matrix: a better numerical characterization for graphical representations of biological sequences. Journal of Theoretical Biology, 247(1):103–109, July 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001075.

deQueiroz:2007:MNS

[dQW07] Marcio S. de Queiroz and Grover L. Waldrop. Modeling and numerical simulation of biotin carboxylase kinetics: Implications for half-sites reactivity. Journal of Theoretical Biology, 246(1):167–175, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005947.

Dubey:2005:SVM

[DRLB05] Anshul Dubey, Matthew J. Realff, Jay H. Lee, and Andreas S. Bommarius. Support vector machines for learning to identify the critical positions of a protein. Journal of Theoretical Biology, 234(3):351–361, June 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005855.

David-Rus:2009:IEC

[DRMLS09] Diana David-Rus, Swagatam Mukhopadhyay, Joel L. Lebowitz, and Anirvan M. Sengupta. Inheritance of epigenetic chromatin silencing. Journal of Theoretical Biology, 258(1):112– 120, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006577.

Drummond:2003:DSB

[Dru03] R. Drummond. Does a structural bridge exist between the DNA and the specialized cytoplasmic organelles during the early part of their development? A mechanism for the positioning of flagella and possibly other cytoplasmic organelles.

Journal of Theoretical Biology, 223(3):309–312, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001000.

Day:2006:RMM

[DRV^{+}06] Judy Day, Jonathan Rubin, Yoram Vodovotz, Carson C. Chow, Angela Reynolds, and Gilles Clermont. A reduced mathematical model of the acute inflammatory response II. Capturing scenarios of repeated endotoxin administration. Journal of Theoretical Biology, 242(1):237–256, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000804.

Daun:2008:EMA

[DRV^{+}08] Silvia Daun, Jonathan Rubin, Yoram Vodovotz, Anirban Roy, Robert Parker, and Gilles Clermont. An ensemble of models of the acute inflammatory response to bacterial lipopolysaccharide in rats: Results from parameter space reduction. Journal of Theoretical Biology, 253(4):843–853, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002166.

Dodds:2001:REL

[DRW01] P. S. Dodds, D. H. Rothman, and J. S. Weitz. Reexamination of the “3/4-law” of metabolism. Journal of Theoretical Biology, 209(1):9–27, March 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922380.

DeVries:2000:CSP

[DS00a] Gerda De Vries and Arthur Sherman. Channel sharing in pancreatic β-cells revisited: Enhancement of emergent bursting by noise. Journal of Theoretical Biology, 207(4):513–530, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921933.

Delgado:2000:SST

[DS00b] Jordi Delgado and Ricard V. Solé. Self-synchronization and task fulfilment in ant colonies. Journal of Theoretical Biology, 205(3):433–441, August 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920770.

Domokos:2004:DCS

[DS04] Gábor Domokos and István Scheuring. Discrete and continuous state population models in a noisy world. Journal of Theoretical Biology, 227(4):535–545, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300448X.

Deane:2005:QMF

[DS05] G. B. Deane and M. Dale Stokes. A quantitative model for flow-induced bioluminescence in dinoflagellates. Journal of Theoretical Biology, 237(2):147–169, November 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001578.

Downer:2006:IED

[DSA^{+}06] Joshua Downer, Joel R. Sevinsky, Natalie G. Ahn, Katheryn A. Resing, and M. D. Betterton. Incorporating expression data in metabolic modeling: a case study of lactate dehydrogenase. Journal of Theoretical Biology, 240(3):464–474, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004467.

Denoth:2002:SMF

[DSCD02] Jachen Denoth, Edgar Stüssi, Gabor Csucs, and Gaudenz Danuser. Single muscle fiber contraction is dictated by inter-sarcomere dynamics. Journal of Theoretical Biology, 216(1):101–122, May 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301925196.

Dostaikova:2002:BME

[DSK02] I. Dostáiková, M. Spinka, and P. Kindlmann. Begging for milk: Evolution of teat massaging in suckling pigs. Journal of Theoretical Biology, 215(3):321–332, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925111.

daSilva:2009:SFP

[dSKL09] Roberto da Silva, Gustavo A. Kellermann, and Luis C. Lamb. Statistical fluctuations in population bargaining in the ultimatum game: Static and evolutionary aspects. Journal of Theoretical Biology, 258(2):208–218, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000277.

Dahlem:2008:ECT

[DSS08] M. A. Dahlem, F. M. Schneider, and E. Schöll. Efficient control of transient wave forms to prevent spreading depolarizations. Journal of Theoretical Biology, 251(2):202–209, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005796.

Dionysiou:2004:FDS

[DSU^{+}04] Dimitra D. Dionysiou, Georgios S. Stamatakos, Nikolaos K. Uzunoglu, Konstantina S. Nikita, and Antigoni Marioli. A four-dimensional simulation model of tumour response to radiotherapy in vivo: parametric validation considering radiosensitivity, genetic profile and fractionation. Journal of Theoretical Biology, 230(1):1–20, September 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400133X.

Dyson:2007:STN

[DSVBW07] Janet Dyson, Eva Sánchez, Rosanna Villella-Bressan, and Glenn F. Webb. Stabilization of telomeres in nonlinear models of proliferating cell lines. Journal of Theoretical Biology, 244(3):400–408, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003997.

Dawson:2001:POFa

[DSY01a] Wayne Dawson, Kazuo Suzuki, and Kenji Yamamoto. A physical origin for functional domain structure in nucleic acids as evidenced by cross-linking entropy: I. Journal of Theoretical Biology, 213(3):359–386, December 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924361.

Dawson:2001:POFb

[DSY01b] Wayne Dawson, Kazuo Suzuki, and Kenji Yamamoto. A physical origin for functional domain structure in nucleic acids as evidenced by cross-linking entropy: II. Journal of Theoretical Biology, 213(3):387–412, December 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924373.

Derganc:2009:EMM

[DSZ09] Jure Derganc, Sasa Svetina, and Bostjan Zeks. Equilibrium mechanics of monolayered epithelium. Journal of Theoretical Biology, 260(3):333–339, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002987.

DeVylder:2006:HRL

[DT06] Bart De Vylder and Karl Tuyls. How to reach linguistic consensus: a proof of convergence for the naming game. Journal of Theoretical Biology, 242(4):818–831, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002189.

Dampier:2007:SPI

[DT07a] William Dampier and Aydin Tozeren. Signaling perturbations induced by invading H. pylori proteins in the host epithelial cells: a mathematical modeling approach. Journal of Theoretical Biology, 248(1):130–144, September 7, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001105.

dOnofrio:2007:NMM

[dT07b] Alberto d’Onofrio and Ian P. M. Tomlinson. A nonlinear mathematical model of cell turnover, differentiation and tumorigenesis in the intestinal crypt. Journal of Theoretical Biology, 244(3):367–374, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003985.

Darabos:2009:DUN

[DTG09] Ch. Darabos, M. Tomassini, and M. Giacobini. Dynamics of unperturbed and noisy generalized Boolean networks. Journal of Theoretical Biology, 260(4):531–544, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002823.

Dinh:2007:MPR

[DTM07] Anh-Tuan Dinh, Theo Theofanous, and Samir Mitragotri. Modeling of pattern regulation in melanophores. Journal of Theoretical Biology, 244(1):141–153, January 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003250.

Dusenbery:2000:SHG

[Dus00] David B. Dusenbery. Selection for high gamete encounter rates explains the success of male and female mating types. Journal of Theoretical Biology, 202(1):1–10, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910172.

Dusenbery:2001:PBS

[Dus01] David B. Dusenbery. Performance of basic strategies for following gradients in two dimensions. Journal of Theoretical Biology, 208(3):345–360, February 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922240.

Dusenbery:2002:EME

[Dus02] David B. Dusenbery. Ecological models explaining the success of distinctive sperm and eggs (oogamy). Journal of Theoretical Biology, 219(1):1–7, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931000.

Dusenbery:2006:SHG

[Dus06] David B. Dusenbery. Selection for high gamete encounter rates explains the evolution of anisogamy using plausible assumptions about size relationships of swimming speed and duration. Journal of Theoretical Biology, 241(1):33–38, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004777.

deVladar:2006:DDS

[dV06] Harold P. de Vladar. Density-dependence as a sizeindependent regulatory mechanism. Journal of Theoretical Biology, 238(2):245–256, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500216X.

DHaese:2004:NPQ

[DVC^{+}04] David D’Haese, Karine Vandermeiren, Roland Julien Caubergs, Yves Guisez, Ludwig De Temmerman, and Nele Horemans. Non-photochemical quenching kinetics during the dark to light transition in relation to the formation of antheraxanthin and zeaxanthin. Journal of Theoretical Biology, 227(2):175–186, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003990.

deVladar:2004:DRC

[dVG04] Harold P. de Vladar and Jorge A. González. Dynamic response of cancer under the influence of immunological activity and therapy. Journal of Theoretical Biology, 227(3):335–348, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004302.

deVladar:2006:DRC

[dVG06] Harold P. de Vladar and Jorge A. González. Dynamic response of cancer under the influence of immunological activity and therapy. Journal of Theoretical Biology, 240(1):162–163, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005023.

Dodin:2000:FWT

[DVL^{+}00] Guy Dodin, Pierre Vandergheynst, Patrick Levoir, Christine Cordier, and Laurence Marcourt. Fourier and wavelet transform analysis, a tool for visualizing regular patterns in DNA sequences. Journal of Theoretical Biology, 206(3):323–326, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921271.

Dodds:2005:GMS

[DW05] P. S. Dodds and D. J. Watts. A generalized model of social and biological contagion. Journal of Theoretical Biology, 232(4):587–604, February 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004515.

Dai:2008:ULR

[DW08a] Qi Dai and Tianming Wang. Use of linear regression model to compare RNA secondary structures. Journal of Theoretical Biology, 253(4):854–860, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002075.

DOrazio:2008:IDL

[DW08b] A. E. D’Orazio and T. A. Waite. Incomplete division of labor: Error-prone multitaskers coexist with specialists. Journal of Theoretical Biology, 250(3):449–460, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004766.

Driessen:2004:CMC

[DWBB04] N. J. B. Driessen, W. Wilson, C. V. C. Bouten, and F. P. T. Baaijens. A computational model for collagen fibre remodelling in the arterial wall. Journal of Theoretical Biology, 226(1):53–64, January 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003199.

Desai:2004:ICB

[DZB^{+}04] Dinakar Desai, Kunbo Zhang, Sailen Barik, Alok Srivastava, Mark E. Bolander, and Gobinda Sarkar. Intragenic codon bias in a set of mouse and human genes. Journal of Theoretical Biology, 230(2):215–225, September 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002061.

Enderling:2006:MMR

[EAC^{+}06] Heiko Enderling, Alexander R. A. Anderson, Mark A. J. Chaplain, Alastair J. Munro, and Jayant S. Vaidya. Mathematical modelling of radiotherapy strategies for early breast cancer. Journal of Theoretical Biology, 241(1):158–171, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004996.

Eames:2006:PDS

[Eam06] Ken T. D. Eames. Partnership dynamics and strain competition. Journal of Theoretical Biology, 243(2):205–213, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600258X.

Etienne:2007:ZSA

[EAM07] Rampal S. Etienne, David Alonso, and Alan J. McKane. The zero-sum assumption in neutral biodiversity theory. Journal of Theoretical Biology, 248(3):522–536, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002883.

Esposito:2009:NAC

[EBId09] S. Esposito, V. Botte, D. Iudicone, and M. Ribera d’Alcala’. Numerical analysis of cumulative impact of phytoplankton photoresponses to light variation on carbon assimilation. Journal of Theoretical Biology, 261(3):361–371, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003476.

Evans:2007:MSS

[EBM07] Thomas Evans, Roger G. Bowers, and Martin Mortimer. Modelling the stability of stx lysogens. Journal of Theoretical Biology, 248(2):241–250, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002664.

Enderling:2007:MMB

[ECAV07] Heiko Enderling, Mark A. J. Chaplain, Alexander R. A. Anderson, and Jayant S. Vaidya. A mathematical model of breast cancer development, local treatment and recurrence. Journal of Theoretical Biology, 246(2):245–259, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005728.

Eldakar:2007:SPA

[EFW07] Omar Tonsi Eldakar, Dene Leo Farrell, and David Sloan Wilson. Selfish punishment: Altruism can be maintained by competition among cheaters. Journal of Theoretical Biology, 249(2):198–205, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003554.

Enquist:2007:ESL

[EG07] Magnus Enquist and Stefano Ghirlanda. Evolution of social learning does not explain the origin of human cumulative culture. Journal of Theoretical Biology, 246(1):129–135, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005911.

Etienne:2000:OSN

[EH00] Rampal S. Etienne and J. A. P. Heesterbeek. On optimal size and number of reserves for metapopulation persistence. Journal of Theoretical Biology, 203(1):33–50, March 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910603.

Ennemoser:2008:RHE

[EH08] A. Ennemoser and N. Hausmann. Regulation of human eye growth by the energy density of the radiation in the retina — a biophysical theory in ophthalmology. Journal of Theoretical Biology, 253(2):302–309, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001434.

Ellermeyer:2003:TEI

[EHG03] Sean Ellermeyer, Jerald Hendrix, and Nariman Ghoochan. A theoretical and empirical investigation of delayed growth response in the continuous culture of bacteria. Journal of Theoretical Biology, 222(4):485–494, June 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000638.

Eichler:2001:SPS

[Eic01] D. Eichler. Slow proliferation as a survival advantage: an attempt to resolve a paradox. Journal of Theoretical Biology, 210(3):319–325, June 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923082.

Edlund:2007:HWL

[EK07] Lena Edlund and Evelyn Korn. Hermaphroditism: What’s not to like? Journal of Theoretical Biology, 245(3):520–527, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005297.

Edlund:2008:REH

[EK08] Lena Edlund and Evelyn Korn. Reply to “The evolution of hermaphroditism” by Matthew S. Grober and Edmund W. Rodgers. Journal of Theoretical Biology, 251(1):193, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005644.

Edelstein-Keshet:2001:MPA

[EKIL01] Leah Edelstein-Keshet, Aliza Israel, and Peter Lansdorp. Modelling perspectives on aging: Can mathematics help us stay Young? Journal of Theoretical Biology, 213(4):509–525, December 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924294.

Edelstein-Keshet:2002:EFA

[EKS02] Leah Edelstein-Keshet and Athan Spiros. Exploring the formation of Alzheimer’s disease senile plaques in silico. Journal of Theoretical Biology, 216(3):301–326, June 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925403.

Eriksson:2005:CDM

[EL05] Anders Eriksson and Kristian Lindgren. Cooperation driven by mutations in multi-person Prisoner’s dilemma. Journal of Theoretical Biology, 232(3):399–409, February 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004023.

Ezanno:2009:MMS

[EL09] Pauline Ezanno and Matthieu Lesnoff. A metapopulation model for the spread and persistence of contagious bovine pleuropneumonia (CBPP) in African sedentary mixed crop-livestock systems. Journal of Theoretical Biology, 256(4):493–503, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005158.

Erni:2002:SSP

[ELB02] Birgit Erni, Felix Liechti, and Bruno Bruderer. Stopover strategies in passerine bird migration: a simulation study. Journal of Theoretical Biology, 219(4):479–493, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302931383.

Elder:2000:SGO

[Eld00] David Elder. Split gene origin and periodic introns. Journal of Theoretical Biology, 207(4):455–472, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921611.

Eom:2006:ELS

[ELJ06] Young-Ho Eom, Soojin Lee, and Hawoong Jeong. Exploring local structural organization of metabolic networks using subgraph patterns. Journal of Theoretical Biology, 241(4):823–829, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000312.

Ellner:2001:PAL

[Ell01] Stephen P. Ellner. Pair approximation for lattice models with multiple interaction scales. Journal of Theoretical Biology, 210(4):435–447, June 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923227.

Eriksson:2004:WAI

[ELL04] Anders Eriksson, Kristian Lindgren, and Torbjörn Lundh. War of attrition with implicit time cost. Journal of Theoretical Biology, 230(3):319–332, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002541.

Engen:2007:URV

[ELSFB07] Steinar Engen, Russell Lande, Bernt-Erik Sæther, and Marco Festa-Bianchet. Using reproductive value to estimate key parameters in density-independent age-structured populations. Journal of Theoretical Biology, 244(2):308–317, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306003432.

Erickson:2000:PRE

[EM00] Jason M. Erickson and Anthony R. Mawson. Possible role of endogenous retinoid (vitamin a) toxicity in the pathophysiology of primary biliary cirrhosis. Journal of Theoretical Biology, 206(1):47–54, September 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921027.

Eurich:2002:RID

[EMS02] Christian W. Eurich, Michael C. Mackey, and Helmut Schwegler. Recurrent inhibitory dynamics: The role of statedependent distributions of conduction delay times. Journal of Theoretical Biology, 216(1):31–50, May 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925348.

Etienne:2002:NES

[EN02] Rampal S. Etienne and C. J. Nagelkerke. Non-equilibria in small metapopulations: Comparing the deterministic Levins model with its stochastic counterpart. Journal of Theoretical Biology, 219(4):463–478, December 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931358.

Etchegoin:2003:MPD

[EN03] P. Etchegoin and M. Nöllmann. A model for protein– DNA interaction dynamics. Journal of Theoretical Biology, 220(2):233–239, January 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931602.

Engen:2007:HCL

[Eng07] Steinar Engen. Heterogeneous communities with lognormal species abundance distribution: Species-area curves and sustainability. Journal of Theoretical Biology, 249(4):791–803, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004250.

Edwards:2000:MSS

[EP00] J. S. Edwards and B. O. Palsson. Multiple steady states in kinetic models of red cell metabolism. Journal of Theoretical Biology, 207(1):125–127, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921659.

Evans:2009:MMD

[EPJ^{+}09] C. J. Evans, R. M. Phillips, P. F. Jones, P. M. Loadman, B. D. Sleeman, C. J. Twelves, and S. W. Smye. A mathematical model of doxorubicin penetration through multicellular layers. Journal of Theoretical Biology, 257(4):598–608, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006048. See corri-

gendum [EPJ^{+}11].

Evans:2011:CMM

[EPJ^{+}11] C. J. Evans, R. M. Phillips, P. F. Jones, P. M. Loadman, B. D. Sleeman, C. J. Twelves, and S. W. Smye. Corrigendum to “A mathematical model of doxorubicin penetration through multicellular layers” [J. Theor. Biol. 257 (2009) 598–608]. Journal of Theoretical Biology, 273(1):235, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006855. See [EPJ^{+}09].

Ehrenfeucht:2007:MOI

[EPR07] A. Ehrenfeucht, D. M. Prescott, and G. Rozenberg. A model for the origin of internal eliminated segments (IESs) and gene rearrangement in stichotrichous ciliates. Journal of Theoretical Biology, 244(1):108–114, January 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002591.

Enns-Ruttan:2000:SSS

[ERM00] Jennifer S. Enns-Ruttan and Robert M. Miura. Spontaneous secondary spiking in excitable cells. Journal of Theoretical Biology, 205(2):181–199, July 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920563.

Engen:2000:PTQ

[ES00] Steinar Engen and Bernt-Erik Sæther. Predicting the time to quasi-extinction for populations far below their carrying capacity. Journal of Theoretical Biology, 205(4):649–658, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920940.

Eshel:2001:MAC

[ES01a] Ilan Eshel and Emilia Sansone. Multiple asymmetry and concord resolutions of a conflict. Journal of Theoretical Biology, 213(2):209–222, November 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924129.

Eshel:2001:P

[ES01b] Ilan Eshel and Avner Shaked. Partnership. Journal of Theoretical Biology, 208(4):457–474, February 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092232X.

Eberl:2008:EBS

[ES08] Hermann J. Eberl and Rangarajan Sudarsan. Exposure of biofilms to slow flow fields: The convective contribution to growth and disinfection. Journal of Theoretical Biology, 253(4):788–807, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001860.

El-Samad:2002:CHP

[ESGK02] H. El-Samad, J. P. Goff, and M. Khammash. Calcium homeostasis and parturient hypocalcemia: An integral feedback perspective. Journal of Theoretical Biology, 214(1):17–29, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924221.

Estrada:2007:FWR

[Est07] Ernesto Estrada. Food webs robustness to biodiversity loss: The roles of connectance, expansibility and degree distribution. Journal of Theoretical Biology, 244(2):296–307, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003444.

Engelstadter:2004:IDD

[ETH04] Jan Engelstädter, Arndt Telschow, and Peter Hammerstein. Infection dynamics of different Wolbachia-types within one host population. Journal of Theoretical Biology, 231(3):345–355, December 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003200.

Etienne:2009:MLE

[Eti09] Rampal S. Etienne. Maximum likelihood estimation of neutral model parameters for multiple samples with different degrees of dispersal limitation. Journal of Theoretical Biology, 257(3):510–514, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006516.

ElSawy:2008:DIV

[ETTV08] K. M. ElSawy, A. Taormina, R. Twarock, and L. Vaughan. Dynamical implications of viral tiling theory. Journal of Theoretical Biology, 252(2):357–369, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000544.

Ezoe:2009:ESA

[EW09] Hideo Ezoe and Nami Washizu. Effect of stochasticity in the availability of pollinators on the resource allocation within a flower. Journal of Theoretical Biology, 258(4):630–636, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000745.

Enden:2005:MDI

[EZM05] Giora Enden, You H. Zhang, and José C. Merchuk. A model of the dynamics of insect cell infection at low multiplicity of infection. Journal of Theoretical Biology, 237(3):257–264, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001773.

Ezoe:2009:DLM

[Ezo09] Hideo Ezoe. Dual lattice model of the evolution of facultative symbiosis with continuous Prisoner’s dilemma game. Journal of Theoretical Biology, 259(4):744–750, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001908.

Flores:2004:GME

[FABdC04] Luis E. Flores, Eduardo J. Aguilar, Valmir C. Barbosa, and Luís Alfredo V. de Carvalho. A graph model for the evolution of specificity in humoral immunity. Journal of Theoretical Biology, 229(3):311–325, August 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001420.

Forster:2006:SMR

[FAW06] Robert Forster, Christoph Adami, and Claus O. Wilke. Selection for mutational robustness in finite populations. Journal of Theoretical Biology, 243(2):181–190, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002694.

Fogel:2000:DEP

[FB00] David B. Fogel and Hans-Georg Beyer. Do evolutionary processes minimize expected losses? Journal of Theoretical Biology, 207(1):117–123, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921660.

Furlong:2001:IPP

[FB01] Rebecca F. Furlong and John F. Y. Brookfield. Inference of past population expansion from the timing of coalescence events in a gene genealogy. Journal of Theoretical Biology, 209(1):75–86, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922471.

Fortier:2006:ECD

[FB06] Pierre A. Fortier and Maimouna Bagna. Estimating conductances of dual-recorded neurons within a network of coupled cells. Journal of Theoretical Biology, 240(3):501–510, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004480.

Fajmut:2008:MKP

[FB08] Ales Fajmut and Milan Brumen. MLC-kinase/phosphatase control of Ca^{2+} signal transduction in airway smooth muscles. Journal of Theoretical Biology, 252(3):474–481, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004912.

Foley:2006:CEG

[FBM06] Catherine Foley, Samuel Bernard, and Michael C. Mackey. Cost-effective g-CSF therapy strategies for cyclical neutropenia: Mathematical modelling based hypotheses. Journal of Theoretical Biology, 238(4):754–763, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002791.

Franks:2005:BIM

[FBUL05] S. J. Franks, H. M. Byrne, J. C. E. Underwood, and C. E. Lewis. Biological inferences from a mathematical model of comedo ductal carcinoma in situ of the breast. Journal of Theoretical Biology, 232(4):523–543, February 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004205.

Fennell:2001:DSM

[FC01] Dean A. Fennell and Finbarr E. Cotter. A dynamical systems model to simulate the perturbation kinetics of gene expression by antisense oligonucleotides. Journal of Theoretical Biology, 209(1):103–112, March 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922501.

Foray:2005:RRR

[FCD^{+}05] Nicolas Foray, Anne-Marie Charvet, David Duchemin, Vincent Favaudon, and Daniel Lavalette. The repair rate of radiation-induced DNA damage: a stochastic interpretation based on the gamma function. Journal of Theoretical Biology, 236(4):448–458, October 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001475.

Fath:2003:RCE

[FCP03] Brian D. Fath, Heriberto Cabezas, and Christopher W. Pawlowski. Regime changes in ecological systems: an information theory approach. Journal of Theoretical Biology, 222(4):517–530, June 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000675.

Florence:2009:REP

[FDA^{+}09] Gerson Florence, Markus A. Dahlem, Antônio-Carlos G. Almeida, José W. M. Bassani, and Jürgen Kurths. The role of extracellular potassium dynamics in the different stages of ictal bursting and spreading depression: a computational study. Journal of Theoretical Biology, 258(2):219–228, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000502.

Ferdy:2002:EMB

[FDG02] Jean-Baptiste Ferdy, Laurence Després, and Bernard Godelle. Evolution of mutualism between globeflowers and their pollinating flies. Journal of Theoretical Biology, 217(2):219–234, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930183.

Fleurant:2004:AMT

[FDR04] C. Fleurant, J. Duchesne, and P. Raimbault. An allometric model for trees. Journal of Theoretical Biology, 227(1):137– 147, March 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003953.

Feigel:2008:ECE

[Fei08] Alexander Feigel. Essential conditions for evolution of communication within a species. Journal of Theoretical Biology, 254(4):768–774, October 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003615.

Feng:2003:ECS

[Fen03] Jianfeng Feng. Effects of correlated and synchronized stochastic inputs to leaky integrator neuronal model. Journal of Theoretical Biology, 222(2):151–162, May 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000225.

Ferdy:2009:VTM

[Fer09a] Jean-Baptiste Ferdy. Virulence and transmission modes in metapopulations: When group selection increases virulence. Journal of Theoretical Biology, 256(2):286–296, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308005080.

Fernando:2009:TCM

[Fer09b] W. J. N. Fernando. Theoretical considerations and modeling of chemical inactivation of microorganisms: Inactivation of giardia cysts by free chlorine. Journal of Theoretical Biology, 259(2):297–303, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001283.

Flynn:2002:MEV

[FF02] Kevin J. Flynn and Michael J. R. Fasham. A modelling exploration of vertical migration by phytoplankton. Journal of Theoretical Biology, 218(4):471–484, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930936.

Fraser:2002:ADC

[FFD^{+}02] Christophe Fraser, Neil M. Ferguson, Frank De Wolf, Azra C. Ghani, Geoff P. Garnett, and Roy M. Anderson. Antigen-driven T-cell turnover. Journal of Theoretical Biology, 219(2):177–192, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930857.

Foster:2009:MSB

[FFHK09] David V. Foster, Jacob G. Foster, Sui Huang, and Stuart A. Kauffman. A model of sequential branching in hierarchical cell fate determination. Journal of Theoretical Biology, 260(4):589–597, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900318X.

Fukui:2007:ECE

[FFIS07] Shin Fukui, Takema Fukatsu, Takashi Ikegami, and Masakazu Shimada. Endosymbiosis as a compact ecosystem with material cycling: Parasitism or mutualism? Journal of Theoretical Biology, 246(4):746–754, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000689.

Fgaier:2008:PMS

[FFME08] Hedia Fgaier, Balazs Feher, Robin C. McKellar, and Hermann J. Eberl. Predictive modeling of siderphore production by Pseudomonas fluorescens under iron limitation. Journal of Theoretical Biology, 251(2):348–362, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005978.

Forsten:2000:PMR

[FFN00] Kimberly E. Forsten, Michael Fannon, and Matthew A. Nugent. Potential mechanisms for the regulation of growth factor binding by heparin. Journal of Theoretical Biology, 205(2):215–230, July 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920642.

Frenkel:2009:SRF

[FFTS09] Zakharia M. Frenkel, Zeev M. Frenkel, Edward N. Trifonov, and Sagi Snir. Structural relatedness via flow networks in protein sequence space. Journal of Theoretical Biology, 260(3):438–444, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003154.

Fishman:2001:TAS

[FGH01] Svetlana Fishman, Michel Génard, and Jean-Gérard Huguet. Theoretical analysis of systematic errors introduced by a pedicel–girdling technique used to estimate separately the xylem and phloem flows. Journal of Theoretical Biology, 213(3):435–446, December 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924427.

Fouchet:2008:IMR

[FGMP08] David Fouchet, Jean-Sébastien Guitton, Stéphane Marchandeau, and Dominique Pontier. Impact of myxomatosis in relation to local persistence in wild rabbit populations: The role of waning immunity and the reproductive period. Journal of Theoretical Biology, 250(4):593–605, February 21, 2008.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005413.

Fritsch:2007:UMP

[FH07] Andreas Fritsch and Christian Hellmich. ‘universal’ microstructural patterns in cortical and trabecular, extracellular and extravascular bone materials: Micromechanicsbased prediction of anisotropic elasticity. Journal of Theoretical Biology, 244(4):597–620, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004206.

Fritsch:2009:DSB

[FHD09] Andreas Fritsch, Christian Hellmich, and Luc Dormieux. Ductile sliding between mineral crystals followed by rupture of collagen crosslinks: Experimentally supported micromechanical explanation of bone strength. Journal of Theoretical Biology, 260(2):230–252, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002392.

Fushing:2006:TMN

[FHL^{+}06] Hsieh Fushing, Chii-Ruey Hwang, Hou-Cheng Lee, Yen-Chiu Lan, and Shwu-Bin Horng. Testing and mapping nonstationarity in animal behavioral processes: a case study on an individual female bean weevil. Journal of Theoretical Biology, 238(4):805–816, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002900.

Finkel:2000:MSD

[FI00] Zoe V. Finkel and Andrew J. Irwin. Modeling size-dependent photosynthesis: Light absorption and the allometric rule. Journal of Theoretical Biology, 204(3):361–369, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920204.

Feugier:2006:HCC

[FI06] François G. Feugier and Yoh Iwasa. How canalization can make loops: a new model of reticulated leaf vascular pattern formation. Journal of Theoretical Biology, 243(2):235–244, November 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002207.

Finkel:2006:MEO

[Fin06] Robert W. Finkel. Model for energy oscillations in cells. Journal of Theoretical Biology, 238(2):286–289, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002043.

Fishman:2003:IRA

[Fis03] Michael A. Fishman. Indirect reciprocity among imperfect individuals. Journal of Theoretical Biology, 225(3):285–292, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002467.

Fishman:2006:IDE

[Fis06] Michael A. Fishman. Involuntary defection and the evolutionary origins of empathy. Journal of Theoretical Biology, 242(4):873–879, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001937.

Funk:2005:SMV

[FJBK05] Georg A. Funk, Vincent A. A. Jansen, Sebastian Bonhoeffer, and Timothy Killingback. Spatial models of virus-immune dynamics. Journal of Theoretical Biology, 233(2):221–236, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004758.

Fall:2001:MMI

[FK01a] Christopher P. Fall and Joel E. Keizer. Mitochondrial modulation of intracellular Ca^{2+} signaling. Journal of Theoretical Biology, 210(2):151–165, May 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922926.

Fox:2001:TEB

[FK01b] Julia C. Fox and Tony M. Keaveny. Trabecular eccentricity and bone adaptation. Journal of Theoretical Biology, 212(2):211–221, September 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923719. See erratum [FK03a].

Furusawa:2001:TRI

[FK01c] Chikara Furusawa and Kunihiko Kaneko. Theory of robustness of irreversible differentiation in a stem cell system: Chaos hypothesis. Journal of Theoretical Biology, 209(4):395–416, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922647.

Fox:2003:ETE

[FK03a] Julic C. Fox and Tony M. Keaveny. Erratum to “Trabecular eccentricity and bone adaptation”: [J. Theor. Biol. 212, 211–221]. Journal of Theoretical Biology, 223(2):267, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000663. See [FK01b].

Fujitani:2003:ARW

[FK03b] Youhei Fujitani and Ichizo Kobayashi. Asymmetric random walk in a reaction intermediate of homologous recombination. Journal of Theoretical Biology, 220(3):359–370, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931675.

Furusawa:2003:RDC

[FK03c] Chikara Furusawa and Kunihiko Kaneko. Robust development as a consequence of generated positional information. Journal of Theoretical Biology, 224(4):413–435, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001899.

Field:2006:FCG

[FKAC06] Mark L. Field, Omar Khan, Jayasimha Abbaraju, and Joseph F. Clark. Functional compartmentation of glycogen phosphorylase with creatine kinase and Ca^{2+} ATPase in skeletal muscle. Journal of Theoretical Biology, 238(2):257–268, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002195.

Fuhrmann:2007:ICA

[FKS07] Jan Fuhrmann, Josef Käs, and Angela Stevens. Initiation of cytoskeletal asymmetry for cell polarization and movement. Journal of Theoretical Biology, 249(2):278–288, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003967.

Feng:2003:RBN

[FL03]  Jianfeng Feng and Guibin Li.  The relationship be-

tween neuronal calcium concentration and firing rate during stochastic synaptic inputs. Journal of Theoretical Biology, 223(3):367–375, August 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001152.

Flaxman:2009:TPT

[FL09] Samuel M. Flaxman and Yuan Lou. Tracking prey or tracking the prey’s resource? Mechanisms of movement and optimal habitat selection by predators. Journal of Theoretical Biology, 256(2):187–200, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004992.

Feng:2001:BBF

[FLBB01] Jianfeng Feng, Guibin Li, David Brown, and Hilary Buxton. Balance between four types of synaptic input for the integrateand-fire model. Journal of Theoretical Biology, 209(1):61–73, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922446.

Feng:2007:ETQ

[FLG^{+}07] Xiaoli Feng, Yuxiao Li, Jianzhong Gu, Yizhong Zhuo, and Huijie Yang. Error thresholds for quasispecies on single peak Gaussian-distributed fitness landscapes. Journal of Theoretical Biology, 246(1):28–32, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005832.

Fouchet:2009:EMS

[FLG^{+}09] David Fouchet, Jacques Le Pendu, Jean-Sébastien Guitton, Micheline Guiserix, Stéphane Marchandeau, and Dominique Pontier. Evolution of microparasites in spatially and genetically structured host populations: The example of RHDV infecting rabbits. Journal of Theoretical Biology, 257(2):212–227, March 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006085.

Flickinger:2005:TFC

[Fli05] R. A. Flickinger. Transcriptional frequency and cell determination. Journal of Theoretical Biology, 232(2):151–156, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002504.

Fishman:2001:HSR

[FLS01] Michael A. Fishman, Arnon Lotem, and Lewi Stone. Heterogeneity stabilizes reciprocal altruism interactions. Journal of Theoretical Biology, 209(1):87–95, March 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922483.

Fungairino:2004:SCS

[FLS^{+}04] S. G. Fungairiño, F. López, J. M. Serrano, F. J. Acosta, and P. de las Heras. Structural cost of shoot modules and its implications on plant potential fitness in a Mediterranean perennial shrub, retama sphaerocarpa (l.) boiss. Journal of Theoretical Biology, 227(1):129–136, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003941.

Franco-Lara:2007:AFL

[FLWB07] Ezequiel Franco-Lara and Dirk Weuster-Botz. Application of fuzzy-logic models for metabolic control analysis. Journal of Theoretical Biology, 245(3):391–399, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005145.

Ferdy:2002:AES

[FM02] Jean-Baptiste Ferdy and Jane Molofsky. Allee effect, spatial structure and species coexistence. Journal of Theoretical Biology, 217(4):413–424, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930511.

Frey:2003:CCA

[FM03] Gabriel Frey and Christian J. Michel. Circular codes in archaeal genomes. Journal of Theoretical Biology, 223(4): 413–431, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930300119X.

Filipe:2004:EDM

[FM04] J. A. N. Filipe and M. M. Maule. Effects of dispersal mechanisms on spatio-temporal development of epidemics. Journal of Theoretical Biology, 226(2):125–141, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002789.

Fujita:2006:PFL

[FM06] Hironori Fujita and Atsushi Mochizuki. Pattern formation of leaf veins by the positive feedback regulation between auxin flow and auxin efflux carrier. Journal of Theoretical Biology, 241(3):541–551, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005485.

Faugeras:2007:MFP

[FM07a] Blaise Faugeras and Olivier Maury. Modeling fish population movements: From an individual-based representation to an advection-diffusion equation. Journal of Theoretical Biology, 247(4):837–848, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001944.

Fortuna:2007:DSF

[FM07b] Miguel A. Fortuna and Carlos J. Melián. Do scale-free regulatory networks allow more expression than random ones? Journal of Theoretical Biology, 247(2):331–336, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001312.

Foley:2009:MMG

[FM09] Catherine Foley and Michael C. Mackey. Mathematical model for g-CSF administration after chemotherapy. Journal of Theoretical Biology, 257(1):27–44, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005006.

Flatt:2001:BSS

[FMD01] Thomas Flatt, Nicolas Maire, and Michael Doebeli. A bit of sex stabilizes host–parasite dynamics. Journal of Theoretical Biology, 212(3):345–354, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192380X.

Fletcher:2000:MMA

[FMF^{+}00] Lynda Fletcher, Elizabeth Mcfarlane, Keith Fletcher, Paul A. A. Grant, and Iain C. Campbell. A mathematical model for assessing changes in neurofilament protein levels in neurites and cell bodies of differentiating neuroblastoma cells. Journal of Theoretical Biology, 205(2):241–252, July 21, 2000.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920599.

Fromhage:2008:MEM

[FMH08] Lutz Fromhage, John M. McNamara, and Alasdair I. Houston. A model for the evolutionary maintenance of monogyny in spiders. Journal of Theoretical Biology, 250(3):524–531, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004948.

Frey:2008:HQM

[FMHW08] Simone Frey, Thomas Millat, Stefan Hohmann, and Olaf Wolkenhauer. How quantitative measures unravel design principles in multi-stage phosphorylation cascades. Journal of Theoretical Biology, 254(1):27–36, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002191.

Feugier:2005:SOV

[FMI05] Francois G. Feugier, A. Mochizuki, and Y. Iwasa. Selforganization of the vascular system in plant leaves: Interdependent dynamics of auxin flux and carrier proteins. Journal of Theoretical Biology, 236(4):366–375, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001268.

Fouchet:2006:WMI

[FMLP06] D. Fouchet, S. Marchandeau, M. Langlais, and D. Pontier. Waning of maternal immunity and the impact of diseases: The example of myxomatosis in natural rabbit populations. Journal of Theoretical Biology, 242(1):81–89, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000592.

Farquhar:2001:RSF

[FMP01] Tony Farquhar and Helen Meyer-Phillips. Relative safety factors against global buckling, anchorage rotation, and tissue rupture in wheat. Journal of Theoretical Biology, 211(1):55– 65, July 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923306.

Feng:2008:CPP

[FmW08] Jie Feng and Tian ming Wang. Characterization of protein primary sequences based on partial ordering. Journal of Theoretical Biology, 254(4):752–755, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003639.

Foltan:2008:TGG

[Fol08] Jaromir S. Foltan. tRNA genes and the genetic code. Journal of Theoretical Biology, 253(3):469–482, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001197.

Forsdyke:2000:HRH

[For00a] D. R. Forsdyke. Haldane’s rule: Hybrid sterility affects the heterogametic sex first because sexual differentiation is on the path to species differentiation. Journal of Theoretical Biology, 204(3):443–452, June 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920289.

Forst:2000:MEC

[For00b] Christian V. Forst. Molecular evolution of catalysis. Journal of Theoretical Biology, 205(3):409–431, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920769.

Forsdyke:2001:AVP

[For01] D. R. Forsdyke. Adaptive value of polymorphism in intracellular self/not-self discrimination? Journal of Theoretical Biology, 210(4):425–434, June 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923203.

Fort:2002:CAS

[For02] Joaquim Fort. A comment on amplification and spread of viruses in a growing plaque. Journal of Theoretical Biology, 214(3):515–518, February 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924695.

Forsdyke:2004:CSR

[For04] Donald R. Forsdyke. Chromosomal speciation: a reply. Journal of Theoretical Biology, 230(2):189–196, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001936.

Forsdyke:2007:CFE

[For07a] Donald R. Forsdyke. Calculation of folding energies of singlestranded nucleic acid sequences: Conceptual issues. Journal of Theoretical Biology, 248(4):745–753, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003359.

Forsdyke:2007:MSI

[For07b] Donald R. Forsdyke. Molecular sex: The importance of base composition rather than homology when nucleic acids hybridize. Journal of Theoretical Biology, 249(2):325– 330, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930700361X.

Forsdyke:2009:SBH

[For09] Donald R. Forsdyke. Samuel Butler and human long term memory: Is the cupboard bare? Journal of Theoretical Biology, 258(1):156–164, May 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000514.

Fowler:2009:ICS

[Fow09] Mike S. Fowler. Increasing community size and connectance can increase stability in competitive communities. Journal of Theoretical Biology, 258(2):179–188, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000253.

Fox:2005:DME

[Fox05] Richard Fox. Directed molecular evolution by machine learning and the influence of nonlinear interactions. Journal of Theoretical Biology, 234(2):187–199, May 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005697.

Famili:2003:SMR

[FP03] Iman Famili and Bernhard O. Palsson. Systemic metabolic reactions are obtained by singular value decomposition of genome-scale stoichiometric matrices. Journal of Theoretical Biology, 224(1):87–96, September 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001462.

Forger:2004:SSR

[FP04a] Daniel B. Forger and David Paydarfar. Starting, stopping, and resetting biological oscillators: in search of optimum perturbations. Journal of Theoretical Biology, 230(4):521–532, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002413.

Forger:2004:MBC

[FP04b] Daniel B. Forger and Charles S. Peskin. Model based conjectures on mammalian clock controversies. Journal of Theoretical Biology, 230(4):533–539, October 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002425.

Ficici:2007:EDF

[FP07] Sevan G. Ficici and Jordan B. Pollack. Evolutionary dynamics of finite populations in games with polymorphic fitness equilibria. Journal of Theoretical Biology, 247(3):426–441, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001087.

Fass:2008:SSA

[FPBM08] Joseph Fass, Chi Pak, James Bamburg, and Alex Mogilner. Stochastic simulation of actin dynamics reveals the role of annealing and fragmentation. Journal of Theoretical Biology, 252(1):173–183, May 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000076.

Fath:2001:CEG

[FPC01] Brian D. Fath, Bernard C. Patten, and Jae S. Choi. Complementarity of ecological goal functions. Journal of Theoretical Biology, 208(4):493–506, February 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922343.

Fromont:2003:DPC

[FPL03] Emmanuelle Fromont, Dominique Pontier, and Michel Langlais. Disease propagation in connected host populations with density-dependent dynamics: the case of the feline leukemia virus. Journal of Theoretical Biology, 223(4):465–475, August 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300122X.

Forest:2006:MAT

[FPM^{+}06] Loïc Forest, Fernando Padilla, Salomé Martínez, Jacques Demongeot, and Jaime San Martín. Modelling of auxin transport affected by gravity and differential radial growth. Journal of Theoretical Biology, 241(2):241–251, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005126.

Frieden:2001:PGI

[FPS01] B. R. Frieden, A. Plastino, and B. H. Soffer. Population genetics from an information perspective. Journal of Theoretical Biology, 208(1):49–64, January 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921994.

Faller:2008:DPD

[FPS08] Beáta Faller, Fabio Pardi, and Mike Steel. Distribution of phylogenetic diversity under random extinction. Journal of Theoretical Biology, 251(2):286–296, March 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006029.

Fowler:2002:PDC

[FR02] M. S. Fowler and G. D. Ruxton. Population dynamic consequences of Allee effects. Journal of Theoretical Biology, 215(1):39–46, March 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924865.

Flaxman:2006:PCS

[FR06] Samuel M. Flaxman and H. Kern Reeve. Putting competition strategies into ideal free distribution models: Habitat selection as a tug of war. Journal of Theoretical Biology, 243(4):587–593, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600292X.

Fernando:2007:NSC

[FR07] Chrisantha Fernando and Jonathan Rowe. Natural selection in chemical evolution. Journal of Theoretical Biology, 247(1):152–167, July 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000604.

Frank:2000:SNS

[Fra00a] Steven A. Frank. Specific and non-specific defense against parasitic attack. Journal of Theoretical Biology, 202(4):283–304, February 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910548.

Frank:2000:WHS

[Fra00b] Steven A. Frank. Within-host spatial dynamics of viruses and defective interfering particles. Journal of Theoretical Biology, 206(2):279–290, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921209.

Frank:2002:IRP

[Fra02] Steven A. Frank. Immune response to parasitic attack: Evolution of a pulsed character. Journal of Theoretical Biology, 219(3):281–290, December 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930293122X.

Frank:2003:SMC

[Fra03] Steven A. Frank. Somatic mosaicism and cancer: inference based on a conditional Luria–Delbrück distribution. Journal of Theoretical Biology, 223(4):405–412, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001176.

Frank:2008:EDR

[Fra08] Steven A. Frank. Evolutionary dynamics of redundant regulatory control. Journal of Theoretical Biology, 255(1):64–68, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003962.

Flatt:2004:SFI

[FS04a] Thomas Flatt and István Scheuring. Stabilizing factors interact in promoting host-parasite coexistence. Journal of Theoretical Biology, 228(2):241–249, May 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000177.

Freeman:2004:EES

[FS04b] Joseph W. Freeman and Frederick H. Silver. Elastic energy storage in unmineralized and mineralized extracellular matrices (ECMs): a comparison between molecular modeling and experimental measurements. Journal of Theoretical Biology, 229(3):371–381, August 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400147X.

Fishman:2006:TPB

[FS06] Michael A. Fishman and Lewi Stone. Trade up polygyny and breeding synchrony in avian populations. Journal of Theoretical Biology, 238(1):104–110, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002146.

Franks:2007:EMM

[FS07] Daniel W. Franks and Thomas N. Sherratt. The evolution of multicomponent mimicry. Journal of Theoretical Biology, 244(4):631–639, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004292.

Fischer:2009:SLB

[FS09] Mareike Fischer and Mike Steel. Sequence length bounds for resolving a deep phylogenetic divergence. Journal of Theoretical Biology, 256(2):247–252, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005110.

Fatade:2000:RMG

[FSG00] Ayowumi Fatade, Jerry Snowhite, and Michael E. Green. A resonance model gives the response to membrane potential for an ion channel: II. Simplification of the calculation, and prediction of stochastic resonance. Journal of Theoretical Biology, 206(3):387–393, October 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921374.

Freund:2002:BSR

[FSGB^{+}02] Jan A. Freund, Lutz Schimansky-Geier, Beatrix Beisner, Alexander Neiman, David F. Russell, Tatyana Yakusheva, and Frank Moss. Behavioral stochastic resonance: How the noise from a Daphnia swarm enhances individual prey capture by juvenile paddlefish. Journal of Theoretical Biology, 214(1):71–83, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924452.

Fishman:2003:FAT

[FSL03] Michael A. Fishman, Lewi Stone, and Arnon Lotem. Fertility assurance through extrapair fertilizations and male paternity defense. Journal of Theoretical Biology, 221(1):103–114, March 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931742.

Fontanari:2006:CEP

[FSS06a] José F. Fontanari, Mauro Santos, and Eörs Szathmáry. Coexistence and error propagation in pre-biotic vesicle models: a group selection approach. Journal of Theoretical Biology, 239(2):247–256, March 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003863.

Fritzsch:2006:AMG

[FSS06b] Guido Fritzsch, Martin Schlegel, and Peter F. Stadler. Alignments of mitochondrial genome arrangements: Applications to metazoan phylogeny. Journal of Theoretical Biology, 240(4):511–520, June 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004492.

Frenkel:2007:WTP

[FT07] Zakharia M. Frenkel and Edward N. Trifonov. Walking through protein sequence space. Journal of Theoretical Biology, 244(1):77–80, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003328.

Finnoff:2009:PCE

[FT09a] David Finnoff and John Tschirhart. Plant competition and exclusion with optimizing individuals. Journal of Theoretical Biology, 261(2):227–237, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003646.

Fischer:2009:MPS

[FT09b] Mareike Fischer and Bhalchandra D. Thatte. Maximum parsimony on subsets of taxa. Journal of Theoretical Biology, 260(2):290–293, September 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002689.

Faahraeus:2002:KCC

[FTEG02] C. Fåhraeus, S. Theander, A. Edman, and W. Grampp. The K–Cl cotransporter in the lobster stretch receptor neurone — a kinetic analysis. Journal of Theoretical Biology, 217(3):287–309, August 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930389.

Fard:2007:HPH

[FTG07] Pooria Jafari Moghaddam Fard, Mohammad Reza Tajvidi, and Shahriar Gharibzadeh. High-pressure hydrocephalus: a novel analytical modeling approach. Journal of Theoretical Biology, 248(3):401–410, October 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700286X.

Fukshansky:2004:SMM

[Fuk04] Anna Fukshansky. A stochastic model for the mutationselection balance in an infinite asexual population with a genome of fixed size. Journal of Theoretical Biology, 231(4):557–562, December 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003327.

Fukshansky:2005:SMS

[Fuk05] Anna Fukshansky. Selection, mutation and sexual reproduction in an infinite haploid population with a genome of finite length. Journal of Theoretical Biology, 237(3):238–244, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500175X.

Furuichi:2002:DBP

[Fur02] Nozomu Furuichi. Dynamics between a predator and a prey switching two kinds of escape motions. Journal of Theoretical Biology, 217(2):159–166, July 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930274.

Fedor:2009:RKI

[FV09] Anna Fedor and Vera Vasas. The robustness of keystone indices in food webs. Journal of Theoretical Biology, 260(3):372–378, October 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003142.

Ferrer:2007:IBM

[FVP^{+}07] Jordi Ferrer, Jaume Vidal, Clara Prats, Joaquim Valls, Esperanza Herreros, Daniel Lopez, Antoni Giró, and Domingo Gargallo. Individual-based model and simulation of Plasmodium falciparum infected erythrocyte in vitro cultures. Journal of Theoretical Biology, 248(3):448–459, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002718.

Feughelman:2000:TDI

[FW00] M. Feughelman and B. K. Willis. Thiol-disulfide interchange a potential key to conformational change associated with amyloid fibril formation. Journal of Theoretical Biology, 206(2):313–315, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092112X.

Farmery:2001:MME

[FW01] A. D. Farmery and J. P. Whiteley. A mathematical model of electron transfer within the mitochondrial respiratory cytochromes. Journal of Theoretical Biology, 213(2):197–207, November 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924117.

Forsten-Williams:2005:KFB

[FWCN05] Kimberly Forsten-Williams, Ceres C. Chua, and Matthew A. Nugent. The kinetics of FGF-2 binding to heparan sulfate proteoglycans and MAP kinase signaling. Journal of Theoretical Biology, 233(4):483–499, April 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004989.

Fricks:2006:NAI

[FWE06] John Fricks, Hongyun Wang, and Timothy C. Elston. A numerical algorithm for investigating the role of the motorcargo linkage in molecular motor-driven transport. Journal of Theoretical Biology, 239(1):33–48, March 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003127.

Fall:2004:CRP

[FWLN04] Christopher P. Fall, John M. Wagner, Leslie M. Loew, and Richard Nuccitelli. Cortically restricted production of IP3 leads to propagation of the fertilization Ca^{2+} wave along the cell surface in a model of the xenopus egg. Journal of Theoretical Biology, 231(4):487–496, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002929.

Fujita-Yoshigaki:2000:SRE

[FY00] Junko Fujita-Yoshigaki. Simulation of regulated exocytosis of amylase from salivary parotid acinar cells by a consecutive reaction model comprising two sequential first-order reactions. Journal of Theoretical Biology, 204(2):165–177, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920095.

Feng:2009:TIO

[FYX^{+}09] Zhilan Feng, Yiding Yang, Dashun Xu, Pei Zhang, and … John W. Glasser. Timely identification of optimal control strategies for emerging infectious diseases. Journal of Theoretical Biology, 259(1):165–171, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001118.

Fletcher:2004:SAC

[FZ04] Jeffrey A. Fletcher and Martin Zwick. Strong altruism can evolve in randomly formed groups. Journal of Theoretical Biology, 228(3):303–313, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000335.

Fletcher:2007:EAG

[FZ07] Jeffrey A. Fletcher and Martin Zwick. The evolution of altruism: Game theory in multilevel selection and inclusive fitness. Journal of Theoretical Biology, 245(1):26–36, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004541.

Green:2002:PEW

[GA02] Martin Green and Thomas Alerstam. The problem of estimating wind drift in migrating birds. Journal of Theoretical Biology, 218(4):485–496, October 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930948.

Gerlee:2007:EHC

[GA07] P. Gerlee and A. R. A. Anderson. An evolutionary hybrid cellular automaton model of solid tumour growth. Journal of Theoretical Biology, 246(4):583–603, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000549.

Gerlee:2008:HCA

[GA08a] P. Gerlee and A. R. A. Anderson. A hybrid cellular automaton model of clonal evolution in cancer: The emergence of the glycolytic phenotype. Journal of Theoretical Biology, 250(4):705–722, February 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005267.

Grafen:2008:NSA

[GA08b] Alan Grafen and Marco Archetti. Natural selection of altruism in inelastic viscous homogeneous populations. Journal of Theoretical Biology, 252(4):694–710, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000350.

Gerlee:2009:ECM

[GA09] P. Gerlee and A. R. A. Anderson. Evolution of cell motility in an individual-based model of tumour growth. Journal of Theoretical Biology, 259(1):67–83, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001106.

Gonzalez:2002:KMB

[GAA02] Ramon González, Barbara A. Andrews, and Juan A. Asenjo. Kinetic model of BiP- and PDI-mediated protein folding and assembly. Journal of Theoretical Biology, 214(4):529–537, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924786.

Gabora:2006:SOO

[Gab06] Liane Gabora. Self-other organization: Why early life did not evolve through natural selection. Journal of Theoretical Biology, 241(3):443–450, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005321.

Gerber:2008:DED

[GABK08] Susanne Gerber, Heike Aßmus, Barbara Bakker, and Edda Klipp. Drug-efficacy depends on the inhibitor type and the target position in a metabolic network — a systematic study. Journal of Theoretical Biology, 252(3):442–455, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004511.

Gagen:2000:IPM

[Gag00] M. J. Gagen. Information processing in multigame environments modeling the evolution of sex via punctuated equilibria. Journal of Theoretical Biology, 206(1):55–72, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921039.

Giuggioli:2006:THR

[GAK^{+}06] L. Giuggioli, G. Abramson, V. M. Kenkre, R. R. Parmenter, and T. L. Yates. Theory of home range estimation from displacement measurements of animal populations. Journal of Theoretical Biology, 240(1):126–135, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003735.

Gamble:2006:ACA

[Gam06] Wilbert Gamble. Atherosclerosis: The carbonic anhydrase, carbon dioxide, calcium concerted theory. Journal of Theoretical Biology, 239(1):16–21, March 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003103.

Gaarding:2000:SMI

[Går00] Lars Gårding. A simple model for the interplay of predators, rodents and food. Journal of Theoretical Biology, 206(1):73–80, September 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921003.

Gardner:2002:MMD

[Gar02] Shea N. Gardner. Modeling multi-drug chemotherapy: Tailoring treatment to individuals. Journal of Theoretical Biology, 214(2):181–207, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924592.

Garte:2004:FPH

[Gar04] Seymour Garte. Fractal properties of the human genome. Journal of Theoretical Biology, 230(2):251–260, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002474.

Garay:2009:CDA

[Gar09] József Garay. Cooperation in defence against a predator. Journal of Theoretical Biology, 257(1):45–51, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006036.

Geli:2009:MTB

[GASA09] Patricia Geli, Mikael Andersson, Åke Svensson, and Dan I. Andersson. A multi-type branching model with varying environment for bacterial dynamics with postantibiotic effect. Journal of Theoretical Biology, 256(1):58–64, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308004980.

Gandhi:2007:ALB

[GAT07] Nikhil Gandhi, Gonen Ashkenasy, and Emmanuel Tannenbaum. Associative learning in biochemical networks. Journal of Theoretical Biology, 249(1):58–66, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003293.

Gavrilets:2006:MSM

[Gav06] Sergey Gavrilets. The maynard Smith model of sympatric speciation. Journal of Theoretical Biology, 239(2):172–182, March 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003784.

Ghosh:2001:RHC

[GB01] Subhendu Ghosh and Amal K. Bera. The role of H^{+} concentration on the capacitance of a membrane channel protein. Journal of Theoretical Biology, 208(3):383–384, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922008.

Ganusov:2002:EIP

[GB02] Vitaly V. Ganusov and Anatoly V. Brilkov. Estimating the instability parameters of plasmid-bearing cells. I. Chemostat culture. Journal of Theoretical Biology, 219(2):193–205, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931012.

Ghosh:2007:ETV

[GB07] S. Ghosh and S. Bhattacharyya. Effect of temperature on viral infection and its control: a mathematical approach. Journal of Theoretical Biology, 247(1):50–63, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000847.

Gautreau:2008:GDS

[GBB08] Aurélien Gautreau, Alain Barrat, and Marc Barthélemy. Global disease spread: Statistics and estimation of arrival times. Journal of Theoretical Biology, 251(3):509–522, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930700608X.

Guedon:2001:PAB

[GBCC01] Y. Guédon, D. Barthélémy, Y. Caraglio, and E. Costes. Pattern analysis in branching and axillary flowering sequences. Journal of Theoretical Biology, 212(4):481–520, October 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923926.

Gall:2000:ALG

[GBD00] David Gall, Erika Baus, and Geneviève Dupont. Activation of the liver glycogen phosphorylase by Ca^{2+} oscillations: a theoretical study. Journal of Theoretical Biology, 207(4):445–454, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921398.

Galindo:2006:TSG

[GBD06] Johan F. Galindo, Clara I. Bermúdez, and Edgar E. Daza. tRNA structure from a graph and quantum theoretical perspective. Journal of Theoretical Biology, 240(4):574–582, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004698.

Gierz:2001:TDM

[GBG01] Gerhard Gierz and Salomon Bartnicki-Garcia. A threedimensional model of fungal morphogenesis based on the vesicle supply center concept. Journal of Theoretical Biology, 208(2):151–164, January 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922094.

Gomez-Benito:2005:IFG

[GBGAKD05] M. J. Gómez-Benito, J. M. García-Aznar, J. H. Kuiper, and M. Doblaré. Influence of fracture gap size on the pattern of long bone healing: a computational study. Journal of Theoretical Biology, 235(1):105–119, July 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000032.

Guimaraes:2007:ISF

[GBK^{+}07a] Paulo. R. Guimarães, Roberta M. Bonaldo, João P. Krajewski, Paulo Guimarães, Aluísio Pinheiro, Joseph Powers, and Sérgio Furtado dos Reis. Investigating small fish schools: Selection of school-formation models by means of general linear models and numerical simulations. Journal of Theoretical Biology, 245(4):784–789, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005601.

Gupta:2007:BNA

[GBK^{+}07b] Simone Gupta, Siddharth S. Bisht, Ritushree Kukreti, Sanjeev Jain, and Samir K. Brahmachari. Boolean network analysis of a neurotransmitter signaling pathway. Journal of Theoretical Biology, 244(3):463–469, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003675.

Grasman:2003:MSA

[GBLV03] J. Grasman, J. W. Brascamp, J. L. Van Leeuwen, and B. Van Putten. The multifractal structure of arterial trees. Journal of Theoretical Biology, 220(1):75–82, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931511.

Grabska-Barwinska:2006:MER

[GBZ06] Agnieszka Grabska-Barwińska and Jaroslaw Zygierewicz. A model of event-related EEG synchronization changes in beta and gamma frequency bands. Journal of Theoretical Biology, 238(4):901–913, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003036.

Gerisch:2008:MMC

[GC08] A. Gerisch and M. A. J. Chaplain. Mathematical modelling of cancer cell invasion of tissue: Local and nonlocal models and the effect of adhesion. Journal of Theoretical Biology, 250(4):684–704, February 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005097.

Guo:2009:PSC

[GC09] Hongjian Guo and Lansun Chen. Periodic solution of a chemostat model with Monod growth rate and impulsive state feedback control. Journal of Theoretical Biology, 260(4):502–509, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003208.

Gin:2007:MMF

[GCB^{+}07] Elan Gin, Edmund J. Crampin, David A. Brown, Trevor J. Shuttleworth, David I. Yule, and James Sneyd. A mathematical model of fluid secretion from a parotid acinar cell. Journal of Theoretical Biology, 248(1):64–80, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002032.

Gualla:2008:TRA

[GCC08] Filippo Gualla, Paolo Cermelli, and Sergio Castellano. Is there a role for amplifiers in sexual selection? Journal of Theoretical Biology, 252(2):255–271, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000830.

Guedon:2007:AGC

[GCH^{+}07] Yann Guédon, Yves Caraglio, Patrick Heuret, Emilie Lebarbier, and Céline Meredieu. Analyzing growth components in trees. Journal of Theoretical Biology, 248(3):418–447, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002706.

Gilchrist:2004:OWH

[GCP04] Michael A. Gilchrist, Daniel Coombs, and Alan S. Perelson. Optimizing within-host viral fitness: infected cell lifespan and virion production rate. Journal of Theoretical Biology, 229(2):281–288, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400150X.

Gross:2007:MEC

[GCS07] Jeffrey D. Gross, I. Constantinidis, and A. Sambanis. Modeling of encapsulated cell systems. Journal of Theoretical Biology, 244(3):500–510, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003651.

Grover:2001:SCB

[GCYH01] N. B. Grover, C. Coustère-Yakir, and C. E. Helmstetter. Synchronous cultures from the baby machine: Anatomy of a model. Journal of Theoretical Biology, 212(3):391–398, October 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923860.

Grazi:2006:VIP

[GD06] Enrico Grazi and Cristina Di Bona. Viscosity as an inseparable partner of muscle contraction. Journal of Theoretical Biology, 242(4):853–861, October 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002190.

Guill:2008:ECE

[GD08] Christian Guill and Barbara Drossel. Emergence of complexity in evolving niche-model food webs. Journal of Theoretical Biology, 251(1):108–120, March 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005760.

Guiot:2006:DEP

[GDC^{+}06] Caterina Guiot, Pier Paolo Delsanto, Alberto Carpinteri, Nicola Pugno, Yuri Mansury, and Thomas S. Deisboeck. The dynamic evolution of the power exponent in a universal growth model of tumors. Journal of Theoretical Biology, 240(3):459– 463, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004455.

Guiot:2003:DTG

[GDD^{+}03] Caterina Guiot, Piero Giorgio Degiorgis, Pier Paolo Delsanto, Pietro Gabriele, and Thomas S. Deisboeck. Does tumor growth follow a “universal law”? Journal of Theoretical Biology, 225(2):147–151, November 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002212.

Gonzalez-Diaz:2009:GLG

[GDPMDS^{+}09] H. González-Díaz, L. G. Pérez-Montoto, A. Duardo-Sanchez, E. Paniagua, S. Vázquez-Prieto, R. Vilas, M. A. Dea-Ayuela, F. Bolas-Fernández, C. R. Munteanu, J. Dorado, J. Costas, and F. M. Ubeira. Generalized lattice graphs for 2d-visualization of biological information. Journal of Theoretical Biology, 261(1):136–147, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003385.

Garcia-Domingo:2007:FWC

[GDS07] Josep L. Garcia-Domingo and Joan Saldaña. Food-web complexity emerging from ecological dynamics on adaptive networks. Journal of Theoretical Biology, 247(4):819–826, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001932.

Goedecke:2005:MOM

[GE05] D. Michael Goedecke and Timothy C. Elston. A model for the oscillatory motion of single dynein molecules. Journal of Theoretical Biology, 232(1):27–39, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003479.

Gross:2004:EFS

[GEF04] Thilo Gross, Wolfgang Ebenhöh, and Ulrike Feudel. Enrichment and foodchain stability: the impact of different forms of predator-prey interaction. Journal of Theoretical Biology, 227(3):349–358, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004314.

Gross:2009:INW

[GEF09] Thilo Gross, Andrew M. Edwards, and Ulrike Feudel. The invisible niche: Weakly density-dependent mortality and the coexistence of species. Journal of Theoretical Biology, 258(1):148–155, May 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000320.

Grover:2004:BSM

[GEK04] N. B. Grover, E. Eidelstein, and L. J. H. Koppes. Bacterial shape maintenance: an evaluation of various models. Journal of Theoretical Biology, 227(4):547–559, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004491.

Gelman:2007:LER

[Gel07] Andrew Gelman. Letter to the editors regarding some papers of Dr. satoshi kanazawa. Journal of Theoretical Biology, 245(3):597–599, April 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005467.

Gernhard:2008:CRP

[Ger08] Tanja Gernhard. The conditioned reconstructed process. Journal of Theoretical Biology, 253(4):769–778, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001811.

Gerberry:2009:TBB

[Ger09] David J. Gerberry. Trade-off between BCG vaccination and the ability to detect and treat latent tuberculosis. Journal of Theoretical Biology, 261(4):548–560, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004093.

Guy:2002:PMP

[GF02] Robert D. Guy and Aaron L. Fogelson. Probabilistic modeling of platelet aggregation: Effects of activation time and receptor occupancy. Journal of Theoretical Biology, 219(1):33–53, November 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931061.

Gonzalez-Forero:2009:RAB

[GF09] Mauricio González-Forero. Removing ambiguity from the biological species concept. Journal of Theoretical Biology, 256(1):76–80, January 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800489X.

Gin:2009:MCM

[GFW^{+}09] Elan Gin, Martin Falcke, Larry E. Wagner, David I. Yule, and James Sneyd. Markov chain Monte Carlo fitting of singlechannel data from inositol trisphosphate receptors. Journal of Theoretical Biology, 257(3):460–474, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006565.

Gopalakrishnan:2004:LIC

[GFWT04] Manoj Gopalakrishnan, Kimberly Forsten-Williams, and Uwe C. Täuber. Ligand-induced coupling versus receptor preassociation: cellular automaton simulations of FGF-2 binding. Journal of Theoretical Biology, 227(2):239–251, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004211.

Gavrilov:2001:RTA

[GG01a] Leonid A. Gavrilov and Natalia S. Gavrilova. The reliability theory of aging and longevity. Journal of Theoretical Biology, 213(4):527–545, December 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924300.

Gonze:2001:MNP

[GG01b] Didier Gonze and Albert Goldbeter. A model for a network of phosphorylation-dephosphorylation cycles displaying the dynamics of dominoes and clocks. Journal of Theoretical Biology, 210(2):167–186, May 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092294X.

Gorlov:2001:CBA

[GG01c] Ivan P. Gorlov and Olga Yu. Gorlova. Cost-benefit analysis of recombination and its application for understanding of chiasma interference. Journal of Theoretical Biology, 213(1):1–8, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923975.

Gourbiere:2002:CBU

[GG02] Sébastien Gourbiere and Ffrançois Gourbiere. Competition between unit-restricted fungi: a metapopulation model. Journal of Theoretical Biology, 217(3):351–368, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930293033X.

Glass:2003:ADC

[GG03a] K. Glass and B. T. Grenfell. Antibody dynamics in childhood diseases: Waning and boosting of immunity and the impact of vaccination. Journal of Theoretical Biology, 221(1):121–131, March 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930393181X.

Gracheva:2003:ICI

[GG03b] M. E. Gracheva and J. D. Gunton. Intercellular communication via intracellular calcium oscillations. Journal of Theoretical Biology, 221(4):513–518, April 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303932012.

Gaspard:2007:SCF

[GG07] P. Gaspard and E. Gerritsma. The stochastic chemomechanics of the f1 -ATPase molecular motor. Journal of Theoretical Biology, 247(4):672–686, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001579.

Grosberg:2009:CMH

[GG09] Anna Grosberg and Morteza Gharib. Computational models of heart pumping efficiencies based on contraction waves in spiral elastic bands. Journal of Theoretical Biology, 257(3):359–370, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006103.

Gibson:2005:MSC

[GGH^{+}05] A. St Clair Gibson, J. H. Goedecke, Y. X. Harley, L. J. Myers, M. I. Lambert, T. D. Noakesa, and E. V. Lambert. Metabolic setpoint control mechanisms in different physiological systems at rest and during exercise. Journal of Theoretical Biology, 236(1):60–72, September 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000767.

Ghim:2005:LSL

[GGK05] Cheol-Min Ghim, Kwang-Il Goh, and Byungnam Kahng. Lethality and synthetic lethality in the genome-wide metabolic network of Escherichia coli. Journal of Theoretical Biology, 237(4):401–411, December 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500189X.

Gomez-Gardenes:2008:NSC

[GGPFM08] Jesús Gómez-Gardeñes, Julia Poncela, Luis Mario Floría, and Yamir Moreno. Natural selection of cooperation and degree hierarchy in heterogeneous populations. Journal of Theoretical Biology, 253(2):296–301, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001203.

Geris:2008:ABF

[GGV^{+}08] Liesbet Geris, Alf Gerisch, Jos Vander Sloten, Rüdiger Weiner, and Hans Van Oosterwyck. Angiogenesis in bone fracture healing: a bioregulatory model. Journal of Theoretical Biology, 251(1):137–158, March 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700567X.

Ghorai:2002:ABC

[GH02] S. Ghorai and N. A. Hill. Axisymmetric bioconvection in a cylinder. Journal of Theoretical Biology, 219(2):137– 152, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519302930778.

Goh:2003:CLC

[GHA03] K. L. Goh, D. W. L. Hukins, and R. M. Aspden. Critical length of collagen fibrils in extracellular matrix. Journal of Theoretical Biology, 223(2):259–261, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000833.

Guedon:2003:CMB

[GHC03] Y. Guédon, P. Heuret, and E. Costes. Comparison methods for branching and axillary flowering sequences. Journal of Theoretical Biology, 225(3):301–325, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002492.

Gregson:2003:PNL

[GHHR03] Andrew M. Gregson, Adam G. Hart, Michael Holcombe, and Francis L. W. Ratnieks. Partial nectar loads as a cause of multiple nectar transfer in the honey bee (Apis mellifera): a simulation model. Journal of Theoretical Biology, 222(1):1– 8, May 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302004873.

Ghirlanda:2002:IGP

[Ghi02] Stefano Ghirlanda. Intensity generalization: Physiology and modelling of a neglected topic. Journal of Theoretical Biology, 214(3):389–404, February 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192452X.

Grant:2009:SMA

[GI09] Valerie J. Grant and R. J. Irwin. A simple model for adaptive variation in the sex ratios of mammalian offspring. Journal of Theoretical Biology, 258(1):38–42, May 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000125.

Giersch:2003:SDG

[Gie03] Christoph Giersch. Stationary diffusion gradients associated with photosynthetic carbon flux — a study of compartmental versus diffusion-reaction models. Journal of Theoretical Biology, 224(3):385–397, October 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001826. See addendum [Gie06].

Giersch:2006:SAE

[Gie06] Christoph Giersch. Simple analytical expression for the extent of transcellular diffusion gradients. Addendum to “Stationary diffusion gradients … ” [J. Theor. Biol. 224, 385–397 (2003)]. Journal of Theoretical Biology, 242(1):261–262, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001342. See [Gie03].

Gingerich:2000:AGN

[Gin00a] Philip D. Gingerich. Arithmetic or geometric normality of biological variation: an empirical test of theory. Journal of Theoretical Biology, 204(2):201–221, May 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920083.

Gintis:2000:SRH

[Gin00b] Herbert Gintis. Strong reciprocity and human sociality. Journal of Theoretical Biology, 206(2):169–179, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921118.

Gintis:2003:HGA

[Gin03] Herbert Gintis. The Hitchhiker’s guide to altruism: Geneculture coevolution, and the internalization of norms. Journal of Theoretical Biology, 220(4):407–418, February 21, 2003.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931043.

Gokhale:2009:PEA

[GINT09] Chaitanya S. Gokhale, Yoh Iwasa, Martin A. Nowak, and Arne Traulsen. The pace of evolution across fitness valleys. Journal of Theoretical Biology, 259(3):613–620, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001805.

Guttal:2007:SOP

[GJ07] Vishwesha Guttal and C. Jayaprakash. Self-organization and productivity in semi-arid ecosystems: Implications of seasonality in rainfall. Journal of Theoretical Biology, 248(3):490–500, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002627.

Gardner:2002:EMS

[GJE02] Carl L. Gardner, Joseph W. Jerome, and Robert S. Eisenberg. Electrodiffusion model simulation of rectangular current pulses in a voltage-biased biological channel. Journal of Theoretical Biology, 219(3):291–299, December 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931231.

Green:2000:DMA

[GK00] Richard Green and E. B. Keverne. The disparate maternal aunt–uncle ratio in male transsexuals: an explanation invoking genomic imprinting. Journal of Theoretical Biology, 202(1):55–63, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910391.

Geritz:2004:MUD

[GK04] Stefan A. H. Geritz and Éva Kisdi. On the mechanistic underpinning of discrete-time population models with complex dynamics. Journal of Theoretical Biology, 228(2):261–269, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930400030X.

Gaohua:2006:MMI

[GK06a] Lu Gaohua and Hidenori Kimura. A mathematical model of intracranial pressure dynamics for brain hypothermia treatment. Journal of Theoretical Biology, 238(4):882–900, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500295X.

Gardner:2006:REM

[GK06b] Andy Gardner and Alex T. Kalinka. Recombination and the evolution of mutational robustness. Journal of Theoretical Biology, 241(4):707–715, August 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000087.

Gonzalez:2009:HBP

[GK09] Aitor González and Ryoichiro Kageyama. Hopf bifurcation in the presomitic mesoderm during the mouse segmentation. Journal of Theoretical Biology, 259(1):176–189, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000691.

Golubev:2003:SIC

[GKB03] A. Golubev, S. Khrustalev, and A. Butov. An in silico investigation into the causes of telomere length heterogeneity and its implications for the Hayflick limit. Journal of Theoretical Biology, 225(2):153–170, November 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002297.

Green:2006:PIB

[GKK06] Darren M. Green, Istvan Z. Kiss, and Rowland R. Kao. Parameterization of individual-based models: Comparisons with deterministic mean-field models. Journal of Theoretical Biology, 239(3):289–297, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003292.

Gierlik:2000:TRA

[GKM^{+}00] Agnieszka Gierlik, Maria Kowalczuk, Paweł Mackiewicz, Mirosław R. Dudek, and Stanisław Cebrat. Is there replicationassociated mutational pressure in the saccharomyces cerevisiae genome? Journal of Theoretical Biology, 202(4):305–314, February 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910627.

Georgiou:2009:UFC

[GKNT09] D. N. Georgiou, T. E. Karakasidis, J. J. Nieto, and A. Torres. Use of fuzzy clustering technique and matrices to classify amino acids and its impact to Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 257(1):17–26, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005808.

Gin:2006:BAC

[GKS06] Elan Gin, Vivien Kirk, and James Sneyd. A bifurcation analysis of calcium buffering. Journal of Theoretical Biology, 242(1):1–15, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000439.

Goodarzi:2007:SAR

[GKTN07] Hani Goodarzi, Ali Katanforoush, Noorossadat Torabi, and Hamed Shateri Najafabadi. Solvent accessibility, residue charge and residue volume, the three ingredients of a robust amino acid substitution matrix. Journal of Theoretical Biology, 245(4):715–725, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600573X.

Geneser:2007:SMM

[GKXS07] Sarah E. Geneser, Robert M. Kirby, Dongbin Xiu, and Frank B. Sachse. Stochastic Markovian modeling of electrophysiology of ion channels: Reconstruction of standard deviations in macroscopic currents. Journal of Theoretical Biology, 245(4):627–637, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004875.

Gafiychuk:2001:PVN

[GL01] V. V. Gafiychuk and I. A. Lubashevsky. On the principles of the vascular network branching. Journal of Theoretical Biology, 212(1):1–9, September 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922775.

Gosselin:2009:API

[GL09] Frédéric Gosselin and Jean-Dominique Lebreton. Asymptotic properties of infinite Leslie matrices. Journal of Theoretical Biology, 256(2):157–163, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004906.

Gaucel:2009:USP

[GLE^{+}09] Sébastien Gaucel, Béatrice Laroche, Pauline Ezanno, Elisabeta Vergu, and Suzanne Touzeau. Using singular perturbations to reduce an epidemiological model: Application to bovine viral diarrhoea virus within-herd spread. Journal of Theoretical Biology, 258(3):426–436, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004220.

Gravenor:2002:MET

[GLK^{+}02] Mike B. Gravenor, Alun L. Lloyd, Peter G. Kremsner, Michel A. Missinou, Mike English, Kevin Marsh, and Dominic Kwiatkowski. A model for estimating total parasite load in falciparum malaria patients. Journal of Theoretical Biology, 217(2):137–148, July 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930304.

Gasieniec:2007:RPS

[GLSW07] Leszek Gasieniec, Cindy Y. Li, Paul Sant, and Prudence W. H. Wong. Randomized probe selection algorithm for microarray design. Journal of Theoretical Biology, 248(3):512–521, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002779.

Ginovart:2002:IIB

[GLV02] Marta Ginovart, Daniel López, and Joaquim Valls. INDISIM, an individual-based discrete simulation model to study bacterial cultures. Journal of Theoretical Biology, 214(2):305–319, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930192466X.

Goh:2007:STC

[GMAH07] Kheng Lim Goh, Judith R. Meakin, Richard M. Aspden, and David W. L. Hukins. Stress transfer in collagen fibrils reinforcing connective tissues: Effects of collagen fibril slenderness and relative stiffness. Journal of Theoretical Biology, 245(2):305–311, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004838.

Guimaraes:2007:BMD

[GMdA^{+}07] Paulo R. Guimarães, Glauco Machado, Marcus A. M. de Aguiar, Pedro Jordano, Jordi Bascompte, Aluísio Pinheiro, and Sérgio Furtado dos Reis. Build-up mechanisms determining the topology of mutualistic networks. Journal of Theoretical Biology, 249(2):181–189, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003797.

Gardner:2006:WDS

[GMFS06] Rui Gardner, Pedro Moradas-Ferreira, and Armindo Salvador. Why does superoxide dismutase overexpression often increase hydrogen peroxide concentrations? An alternative explanation. Journal of Theoretical Biology, 242(3):798–800, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001615.

Gaohua:2006:IMT

[GMK06] Lu Gaohua, Tsuyoshi Maekawa, and Hidenori Kimura. An integrated model of thermodynamic-hemodynamic-pharmacokinetic system and its application on decoupling control of intracranial temperature and pressure in brain hypothermia treatment. Journal of Theoretical Biology, 242(1):16–31, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000440.

Gonzalez:2009:BBP

[GMM09] Miguel González, Rodrigo Martínez, and Manuel Mota. Bisexual branching processes to model extinction conditions for y-linked genes. Journal of Theoretical Biology, 258(3):478– 488, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005730.

Garcia-Manso:2008:MPP

[GMMGD^{+}08] J. M. García-Manso, J. M. Martín-González, M. E. Da Silva-Grigoletto, D. Vaamonde, P. Benito, and J. Calderón. Male powerlifting performance described from the viewpoint of complex systems. Journal of Theoretical Biology, 251(3):498–508, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307006509.

Garcia-Manso:2005:MLD

[GMMGDA05] Juan M. García-Manso, Juan M. Martín-González, Nancy Dávila, and Enrique Arriaza. Middle and long distance athletics races viewed from the perspective of complexity. Journal of Theoretical Biology, 233(2):191–198, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004734.

Gal:2007:CDR

[GMMR07] József Gál, Taeko Miyazaki, and Victor Benno Meyer-Rochow. Computational determination of refractive index distribution in the crystalline cones of the compound eye of Antarctic krill (Euphausia superba). Journal of Theoretical Biology, 244(2):318–325, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003420.

Giesl:2004:SAE

[GMSW04] Peter Giesl, Dorothea Meisel, Jürgen Scheurle, and Heiko Wagner. Stability analysis of the elbow with a load. Journal of Theoretical Biology, 228(1):115–125, May 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004612.

Gronau:2009:TOD

[GMY09] Ilan Gronau, Shlomo Moran, and Irad Yavneh. Towards optimal distance functions for stochastic substitution models. Journal of Theoretical Biology, 260(2):294–307, September 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002537.

Greenman:2007:EFI

[GN07] J. V. Greenman and R. A. Norman. Environmental forcing, invasion and control of ecological and epidemiological systems. Journal of Theoretical Biology, 247(3):492–506, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001531.

Goodarzi:2005:OGC

[GNH^{+}05] Hani Goodarzi, Hamed Shateri Najafabadi, Kasra Hassani, Hamed Ahmadi Nejad, and Noorossadat Torabi. On the optimality of the genetic code, with the consideration of coevolution theory by comparison of prominent cost measure matrices. Journal of Theoretical Biology, 235(3):318–325, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000378.

Gutenkunst:2007:IRD

[GNLEK07] Ryan Gutenkunst, Nathaniel Newlands, Molly Lutcavage, and Leah Edelstein-Keshet. Inferring resource distributions from Atlantic bluefin tuna movements: An analysis based on net displacement and length of track. Journal of Theoretical Biology, 245(2):243–257, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004905.

Golubski:2007:IEB

[Gol07] Antonio J. Golubski. Interactive effects of builders and exploiters on environmental quality and the outcome of competition between the two. Journal of Theoretical Biology, 249(1):46–57, November 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002974.

Goldsmith:2008:AEI

[Gol08] Theodore C. Goldsmith. Aging, evolvability, and the individual benefit requirement; medical implications of aging theory controversies. Journal of Theoretical Biology, 252(4):764–768, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800101X.

Golubev:2009:HCG

[Gol09] A. Golubev. How could the Gompertz-makeham law evolve. Journal of Theoretical Biology, 258(1):1–17, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000241.

Gouveia-Oliveira:2009:HVN

[GOP09] Rodrigo Gouveia-Oliveira and Anders Gorm Pedersen. Higher variability in the number of sexual partners in males can contribute to a higher prevalence of sexually transmitted diseases in females. Journal of Theoretical Biology, 261(1):100–106, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002707.

Gosselin:2006:ADE

[Gos06] Frédéric Gosselin. An assessment of the dependence of evenness indices on species richness. Journal of Theoretical Biology, 242(3):591–597, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001676.

Goldman:2000:CSE

[GP00] Daniel Goldman and Aleksander S. Popel. A computational study of the effect of capillary network anastomoses and tortuosity on oxygen transport. Journal of Theoretical Biology, 206(2):181–194, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921131.

Goldman:2001:CSE

[GP01] Daniel Goldman and Aleksander S. Popel. A computational study of the effect of vasomotion on oxygen transport from capillary networks. Journal of Theoretical Biology, 209(2):189–199, March 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922549.

Goldbeter:2008:MSC

[GP08a] Albert Goldbeter and Olivier Pourquié. Modeling the segmentation clock as a network of coupled oscillations in the notch, wnt and FGF signaling pathways. Journal of Theoretical Biology, 252(3):574–585, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000179.

Grazi:2008:POP

[GP08b] Enrico Grazi and Sara Pozzati. Protein osmotic pressure modulates actin filament length distribution. Journal of Theoretical Biology, 251(3):411–420, April 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006303.

Gillespie:2004:MMA

[GPB^{+}04] Colin S. Gillespie, Carole J. Proctor, Richard J. Boys, Daryl P. Shanley, Darren J. Wilkinson, and Thomas B. L. Kirkwood. A mathematical model of ageing in yeast. Journal of Theoretical Biology, 229(2):189–196, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001237.

Gravner:2007:PFL

[GPG07] Janko Gravner, Damien Pitman, and Sergey Gavrilets. Percolation on fitness landscapes: Effects of correlation, phenotype, and incompatibilities. Journal of Theoretical Biology, 248(4):627–645, October 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003335.

Greer:2006:MAD

[GPMW06] Meredith L. Greer, Laurent Pujo-Menjouet, and Glenn F. Webb. A mathematical analysis of the dynamics of prion proliferation. Journal of Theoretical Biology, 242(3):598–606, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001652.

Gupta:2005:DSS

[GPN05] Shakti Gupta, S. S. Pilyugin, and Atul Narang. The dynamics of single-substrate continuous cultures: The role of ribosomes. Journal of Theoretical Biology, 232(4):467–490, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004151.

Gomes:2005:FLB

[GR05] Mario W. Gomes and Andy L. Ruina. A five-link 2d brachiating ape model with life-like zero-energy-cost motions. Journal of Theoretical Biology, 237(3):265–278, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001785.

Garske:2008:ESE

[GR08a] T. Garske and C. J. Rhodes. The effect of superspreading on epidemic outbreak size distributions. Journal of Theoretical Biology, 253(2):228–237, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001124.

Grober:2008:EH

[GR08b] Matthew S. Grober and Edmund W. Rodgers. The evolution of hermaphroditism. Journal of Theoretical Biology, 251(1):190– 192, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005449.

Grafen:2002:FFL

[Gra02] Alan Grafen. A first formal link between the price equation and an optimization program. Journal of Theoretical Biology, 217(1):75–91, July 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930158.

Grafen:2006:OIF

[Gra06] Alan Grafen. Optimization of inclusive fitness. Journal of Theoretical Biology, 238(3):541–563, February 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002559.

Grant:2007:CMT

[Gra07] Valerie J. Grant. Could maternal testosterone levels govern mammalian sex ratio deviations? Journal of Theoretical Biology, 246(4):708–719, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000720.

Guilford:2004:PED

[GRBR04] Tim Guilford, Stephen Roberts, Dora Biro, and Iead Rezek. Positional entropy during pigeon homing II: navigational interpretation of Bayesian latent state models. Journal of Theoretical Biology, 227(1):25–38, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003060.

Gregorius:2000:ASH

[Gre00] Hans-Rolf Gregorius. Adaptation to spatially heterogeneous modifying and adaptive environments. Journal of Theoretical Biology, 205(1):121–131, July 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920496.

Green:2005:PRP

[Gre05] Michael E. Green. A possible role for phosphate in complexing the arginines of S4 in voltage gated channels. Journal of Theoretical Biology, 233(3):337–341, April 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004849.

Gregorius:2009:RD

[Gre09] Hans-Rolf Gregorius. Relational diversity. Journal of Theoretical Biology, 257(1):150–158, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005857.

Gonze:2002:MEF

[GRG02] Didier Gonze, Marc R. Roussel, and Albert Goldbeter. A model for the enhancement of fitness in cyanobacteria based on resonance of a circadian oscillator with the external light– dark cycle. Journal of Theoretical Biology, 214(4):577–597, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924762.

Gomes:2007:IPI

[GRH^{+}07] M. Gabriela M. Gomes, Paula Rodrigues, Frank M. Hilker, Natalia B. Mantilla-Beniers, Marion Muehlen, Ana Cristina Paulo, and Graham F. Medley. Implications of partial immunity on the prospects for tuberculosis control by post-exposure interventions. Journal of Theoretical Biology, 248(4):608–617, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002846.

Grover:2002:SHC

[Gro02] James P. Grover. Stoichiometry, herbivory and competition for nutrients: Simple models based on planktonic ecosystems. Journal of Theoretical Biology, 214(4):599–618, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924889.

Grover:2004:PCN

[Gro04] James P. Grover. Predation, competition, and nutrient recycling: a stoichiometric approach with multiple nutrients. Journal of Theoretical Biology, 229(1):31–43, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400102X.

Gruzdev:2000:DTH

[Gru00] Alexey D. Gruzdev. DNA topology in heterochromatin (a hypothesis). Journal of Theoretical Biology, 207(2):255– 264, November 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300921702.

Gardner:2003:EME

[GRW03] A. Gardner, S. E. Reece, and S. A. West. Even more extreme fertility insurance and the sex ratios of protozoan blood parasites. Journal of Theoretical Biology, 223(4):515–521, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001425.

Goldberg:2000:BGH

[GS00] Billi Goldberg and Raphael B. Stricker. Bridging the gap: Human diploid cell strains and the origin of AIDS. Journal of Theoretical Biology, 204(4):497–503, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920332.

Gamarra:2002:BDR

[GS02a] Javier G. P. Gamarra and Ricard V. Solé. Biomass-diversity responses and spatial dependencies in disturbed tallgrass prairies. Journal of Theoretical Biology, 215(4):469–480, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301925202.

Gilchrist:2002:MHP

[GS02b] Michael A. Gilchrist and Akira Sasaki. Modeling host–parasite coevolution: a nested approach based on mechanistic models. Journal of Theoretical Biology, 218(3):289–308, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930766.

Gimenez:2002:DPT

[GS02c] Juan A. Martínez Giménez and Rafael Tabarés Seisdedos. On the dimerization of the primitive tRNAs: Implications in the origin of genetic code. Journal of Theoretical Biology, 217(4):493–498, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930444.

Gaillard:2006:DTG

[GS06] Claire Gaillard and François Strauss. DNA topology and genome organization in higher eukaryotes: a model. Journal of Theoretical Biology, 243(4):604–607, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002906.

Groff:2008:CDI

[GS08a] Jeffrey R. Groff and Gregory D. Smith. Calcium-dependent inactivation and the dynamics of calcium puffs and sparks. Journal of Theoretical Biology, 253(3):483–499, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001537.

Guichard:2008:IEC

[GS08b] Frederic Guichard and Robin Steenweg. Intrinsic and extrinsic causes of spatial variability across scales in a metacommunity. Journal of Theoretical Biology, 250(1):113–124, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930700447X.

Gintis:2001:CSC

[GSB01] Herbert Gintis, Eric Alden Smith, and Samuel Bowles. Costly signaling and cooperation. Journal of Theoretical Biology, 213(1):103–119, November 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924063.

Geyer:2005:SMR

[GSB05] Hartmut Geyer, Andre Seyfarth, and Reinhard Blickhan. Spring-mass running: simple approximate solution and application to gait stability. Journal of Theoretical Biology, 232(3):315–328, February 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003881.

Gnacadja:2007:MIR

[GSG^{+}07] Gilles Gnacadja, Alex Shoshitaishvili, Michael J. Gresser, Brian Varnum, David Balaban, Mark Durst, Chris Vezina, and Yu Li. Monotonicity of interleukin-1 receptor-ligand binding with respect to antagonist in the presence of decoy receptor. Journal of Theoretical Biology, 244(3):478–488, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003249.

Godula:2006:EWE

[GSM06] Tomas Godula, Hana Sevcikova, and John H. Merkin. The emergence of wave emitting centres in an excitable medium. Journal of Theoretical Biology, 240(1):136–148, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003747.

Gunji:2008:MMC

[GSNH08] Yukio-Pegio Gunji, Tomohiro Shirakawa, Takayuki Niizato, and Taichi Haruna. Minimal model of a cell connecting amoebic motion and adaptive transport networks. Journal of Theoretical Biology, 253(4):659–667, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002026.

Garcia:2006:HDD

[GSRC^{+}06] José A. García, Luis E. Soto-Ramírez, Germinal Cocho, Tzipe Govezensky, and Marco V. José. HIV-1 dynamics at different time scales under antiretroviral therapy. Journal of Theoretical Biology, 238(1):220–229, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002250.

Guo:2008:HAB

[GST08] Zaiyi Guo, Peter M. A. Sloot, and Joc Cing Tay. A hybrid agent-based approach for modeling microbiological systems. Journal of Theoretical Biology, 255(2):163–175, November 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004141.

Goriely:2003:BMH

[GT03] Alain Goriely and Michael Tabor. Biomechanical models of hyphal growth in actinomycetes. Journal of Theoretical Biology, 222(2):211–218, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000298.

Gevertz:2006:MEV

[GT06a] Jana L. Gevertz and Salvatore Torquato. Modeling the effects of vasculature evolution on early brain tumor growth. Journal of Theoretical Biology, 243(4):517–531, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002918.

Gutierrez:2006:MDE

[GT06b] Juan B. Gutierrez and John L. Teem. A model describing the effect of sex-reversed YY fish in an established wild population: The use of a Trojan Y chromosome to cause extinction of an introduced exotic species. Journal of Theoretical Biology, 241(2):333–341, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005242.

Guebel:2008:CMO

[GT08] Daniel V. Guebel and Néstor V. Torres. A computer model of oxygen dynamics in human colon mucosa: Implications in normal physiology and early tumor development. Journal of Theoretical Biology, 250(3):389–409, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004626.

Gautrais:2002:EPC

[GTDA02] Jacques Gautrais, Guy Theraulaz, Jean-Louis Deneubourg, and Carl Anderson. Emergent polyethism as a consequence of increased colony size in insect societies. Journal of Theoretical Biology, 215(3):363–373, April 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925068.

Gracheva:2001:SEI

[GTG01] M. E. Gracheva, R. Toral, and J. D. Gunton. Stochastic effects in intercellular calcium spiking in hepatocytes. Journal of Theoretical Biology, 212(1):111–125, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923628.

Garay:2003:CEE

[GV03] József Garay and Zoltán Varga. Coincidence of ESAD and ESS in dominant-recessive hereditary systems. Journal of Theoretical Biology, 222(3):297–305, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000274.

Gastaldo:2008:IRR

[GVB^{+}08] Jérôme Gastaldo, Muriel Viau, Michael Bouchot, Aurélie Joubert, Anne-Marie Charvet, and Nicolas Foray. Induction and repair rate of DNA damage: a unified model for describing effects of external and internal irradiation and contamination with heavy metals. Journal of Theoretical Biology, 251(1):68– 81, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005401.

Gilad:2007:MMP

[GvHP^{+}07] E. Gilad, J. von Hardenberg, A. Provenzale, M. Shachak, and E. Meron. A mathematical model of plants as ecosystem engineers. Journal of Theoretical Biology, 244(4):680–691, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003493.

Grassly:2000:EPS

[GVK00] Nicholas C. Grassly, Arndt Von Haeseler, and David C. Krakauer. Error, population structure and the origin of diverse sign systems. Journal of Theoretical Biology, 206(3):369–378, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921337.

Graham:2004:TLE

[GvO04] Bruce P. Graham and Arjen van Ooyen. Transport limited effects in a model of dendritic branching. Journal of Theoretical Biology, 230(3):421–432, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002802.

Gilchrist:2006:MPT

[GW06] Michael A. Gilchrist and Andreas Wagner. A model of protein translation including codon bias, nonsense errors, and ribosome recycling. Journal of Theoretical Biology, 239(4):417– 434, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003395.

Gomes:2004:IRV

[GWM04] M. Gabriela M. Gomes, Lisa J. White, and Graham F. Medley. Infection, reinfection, and vaccination under suboptimal immune protection: epidemiological perspectives. Journal of Theoretical Biology, 228(4):539–549, June 21, 2004.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000748.

Gomes:2005:RT

[GWM05] M. Gabriela M. Gomes, Lisa J. White, and Graham F. Medley. The reinfection threshold. Journal of Theoretical Biology, 236(1):111–113, September 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000834. See erratum [GWM06].

Gomes:2006:ERT

[GWM06] M. Gabriela M. Gomes, Lisa J. White, and Graham F. Medley. Erratum to “The reinfection threshold”: [J. Theor. Biol. 236 (2005) 111–113]. Journal of Theoretical Biology, 239(4):518, April 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000397. See [GWM05].

Garcion:2004:RMS

[GWPW04] Emmanuel Garcion, Bradley Wallace, Laurent Pelletier, and Didier Wion. RNA mutagenesis and sporadic prion diseases. Journal of Theoretical Biology, 230(2):271–274, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002553.

Glass:2003:ITS

[GXG03] K. Glass, Y. Xia, and B. T. Grenfell. Interpreting timeseries analyses for continuous-time biological models-measles as a case study. Journal of Theoretical Biology, 223(1):19– 25, July 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000316.

Guo:2004:HFI

[GZ04] Guanq-Qin Guo and Guo-Chang Zheng. Hypotheses for the functions of intercellular bridges in male germ cell development and its cellular mechanisms. Journal of Theoretical Biology, 229(1):139–146, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001110.

Guo:2006:SPR

[GZG06] Hao Guo, Chong Zheng, and Elizabeth R. Gaillard. Structure prediction of the RPE65 protein. Journal of Theoretical Biology, 242(1):117–122, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000622.

Guo:2007:CSS

[GZG07] Hao Guo, Chong Zheng, and Elizabeth R. Gaillard. Computational studies for the structure and function of mRPE65. Journal of Theoretical Biology, 245(2):312–318, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004693.

Gurarie:2006:DRS

[GZK06] D. Gurarie, P. A. Zimmerman, and C. H. King. Dynamic regulation of single- and mixed-species malaria infection: Insights to specific and non-specific mechanisms of control. Journal of Theoretical Biology, 240(2):185–199, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004182.

Goltsev:2003:SAP

[GZL^{+}03] V. Goltsev, I. Zaharieva, P. Lambrev, I. Yordanov, and R. Strasser. Simultaneous analysis of prompt and delayed chlorophyll a fluorescence in leaves during the induction period of dark to light adaptation. Journal of Theoretical Biology, 225(2):171–183, November 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002364.

Hud:2000:IMS

[HA00] Nicholas V. Hud and Frank A. L. Anet. Intercalation-mediated synthesis and replication: a new approach to the origin of life. Journal of Theoretical Biology, 205(4):543–562, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920848.

Hampton:2007:SMM

[HA07] Marshall Hampton and Matthew T. Andrews. A simple molecular mathematical model of mammalian hibernation. Journal of Theoretical Biology, 247(2):297–302, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001099.

Hashimoto:2009:FPE

[HA09] Koh Hashimoto and Kazuyuki Aihara. Fixation probabilities in evolutionary game dynamics with a two-strategy game in finite diploid populations. Journal of Theoretical Biology, 258(4):637–645, June 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000666.

Hanin:2009:CSS

[HAC^{+}09] Leonid Hanin, Saria S. Awadalla, Paul Cox, Galina Glazko, and Andrei Yakovlev. Chromosome-specific spatial periodicities in gene expression revealed by spectral analysis. Journal of Theoretical Biology, 256(3):333–342, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005365.

Hadany:2001:CBT

[Had01] Lilach Hadany. A conflict between two evolutionary levels in trees. Journal of Theoretical Biology, 208(4):507–521, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922367.

Haeri:2003:MPU

[HAG03] M. Haeri, D. Asemani, and Sh. Gharibzadeh. Modeling of pain using artificial neural networks. Journal of Theoretical Biology, 220(3):277–284, February 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931304.

Han:2001:PIR

[Han01] Bo-Ping Han. Photosynthesis-irradiance response at physiological level: a mechanistic model. Journal of Theoretical Biology, 213(2):121–127, November 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924130.

Han:2002:MMA

[Han02] Bo-Ping Han. A mechanistic model of algal photoinhibition induced by photodamage to photosystem — II. Journal of Theoretical Biology, 214(4):519–527, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924683.

Han:2004:EBS

[Han04] Hai-Chao Han. An echocardiogram-based 16-segment model for predicting left ventricular ejection fraction improvement. Journal of Theoretical Biology, 228(1):7–15, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004533.

Harms:2001:CBP

[Har01] William Harms. Cooperative boundary populations: the evolution of cooperation on mortality risk gradients. Journal of Theoretical Biology, 213(2):299–313, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924245.

Hart:2002:IPI

[Har02] Deborah R. Hart. Intraguild predation, invertebrate predators, and trophic cascades in lake food Webs. Journal of Theoretical Biology, 218(1):111–128, September 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930535.

Harper:2006:MSA

[Har06] D. G. C. Harper. Maynard Smith: Amplifying the reasons for signal reliability. Journal of Theoretical Biology, 239(2):203– 209, March 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003814.

Hardling:2007:FEU

[Här07a] Roger Härdling. Fairness evolution in the ultimatum game is a function of reward size. Journal of Theoretical Biology, 246(4):720–724, June 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000586.

Hartmann:2007:LBS

[Har07b] Dirk Hartmann. Limit behaviour of spatially growing cell cultures. Journal of Theoretical Biology, 244(3):409–415, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003936.

Haschke:2001:BMD

[Has01a] Heinz Haschke. A biokinetic model to describe consequences of inhibition/stimulation in DNA-proofreading and repair — 1. Development of the model. Journal of Theoretical Biology, 212(4):425–456, October 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300923944.

Hastings:2001:MJK

[Has01b] Dr A. Hastings. In memoriam: Joel E. Keizer. Journal of Theoretical Biology, 210(2):133–139, May 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192329X.

Hashimoto:2006:UIU

[Has06] Koh Hashimoto. Unpredictability induced by unfocused games in evolutionary game dynamics. Journal of Theoretical Biology, 241(3):669–675, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600004X.

Hauert:2006:SES

[Hau06] Christoph Hauert. Spatial effects in social dilemmas. Journal of Theoretical Biology, 240(4):627–636, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004765.

Hauryliuk:2007:ETM

[Hau07] Vasili Hauryliuk. Evolution of translational machinery: Could translation termination come into being before elongation? Journal of Theoretical Biology, 248(3):574–578, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002810.

Havsteen:2004:CPC

[Hav04] Bent H. Havsteen. Control of the pathogenic conformational change of the prion protein by an attractor of low order. Journal of Theoretical Biology, 231(1):39–48, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001948.

Hays:2000:IVR

[Hay00] Graeme C. Hays. The implications of variable remigration intervals for the assessment of population size in marine turtles. Journal of Theoretical Biology, 206(2):221–227, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921167.

Hayes:2001:HHM

[Hay01] Christopher Hayes. Hypothesis: Hemimelia and the maintenance of pentadactyly? Journal of Theoretical Biology, 209(4):493–495, April 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922842.

Henrich:2001:WPP

[HB01] Joseph Henrich and Robert Boyd. Why people punish defectors: Weak conformist transmission can stabilize costly enforcement of norms in cooperative dilemmas. Journal of Theoretical Biology, 208(1):79–89, January 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922021.

Hillier:2002:BTM

[HB02a] Jonathan G. Hillier and A. Nicholas E. Birch. Bi-trophic mathematical model for pest adaptation to a resistant crop. Journal of Theoretical Biology, 215(3):305–319, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925135.

Hillier:2002:TWR

[HB02b] Jonathan G. Hillier and A. Nicholas E. Birch. Travelling waves of resistance in a Bi-trophic pest adaptation model. Journal of Theoretical Biology, 219(4):507–519, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302931413.

Hinow:2007:MSI

[HB07] Peter Hinow and Erik M. Boczko. Molecular seismology: An inverse problem in nanobiology. Journal of Theoretical Biology, 246(1):145–158, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600556X.

Harrison:2009:GTM

[HB09] M. D. Harrison and M. Broom. A game-theoretic model of interspecific brood parasitism with sequential decisions. Journal of Theoretical Biology, 256(4):504–517, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308004529.

Housaindokht:2008:ALB

[HBB08] M. R. Housaindokht, M. R. Bozorgmehr, and M. Bahrololoom. Analysis of ligand binding to proteins using molecular dynamics simulations. Journal of Theoretical Biology, 254(2): 294–300, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930800218X.

Halloy:2002:FAM

[HBLG02] J. Halloy, B. A. Bernard, G. Loussouarn, and A. Goldbeter. The follicular automaton model: Effect of stochasticity and of synchronization of hair cycles. Journal of Theoretical Biology, 214(3):469–479, February 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924749.

Huppert:2005:MSP

[HBOS05] Amit Huppert, Bernd Blasius, Ronen Olinky, and Lewi Stone. A model for seasonal phytoplankton blooms. Journal of Theoretical Biology, 236(3):276–290, October 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001219.

Hart:2007:HAI

[HBS07] Emma Hart, Hugues Bersini, and Francisco C. Santos. How affinity influences tolerance in an idiotypic network. Journal of Theoretical Biology, 249(3):422–436, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003530.

Hoyle:2008:ITS

[HBWB08] Andrew Hoyle, Roger G. Bowers, Andrew White, and Michael Boots. The influence of trade-off shape on evolutionary behaviour in classical ecological scenarios. Journal of Theoretical Biology, 250(3):498–511, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700495X.

Hotton:2007:AFD

[HC07a] Scott Hotton and Mike E. Colvin. Analytic formulas for discrete stochastic models of cell populations with both differentiation and de-differentiation. Journal of Theoretical Biology, 245(4):610–626, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005479.

Hoyle:2007:STM

[HC07b] Martin Hoyle and James E. Cresswell. A search theory model of patch-to-patch forager movement with application to pollinator-mediated gene flow. Journal of Theoretical Biology, 248(1):154–163, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002056.

Hanigan:2002:MMM

[HCB^{+}02] M. D. Hanigan, L. A. Crompton, B. J. Bequette, J. A. N. Mills, and J. France. Modelling mammary metabolism in the dairy cow to predict milk constituent yield, with emphasis on amino acid metabolism and milk protein production: Model evaluation. Journal of Theoretical Biology, 217(3):311–330, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930377.

Hansen:2000:GCM

[HCC00] Thomas F. Hansen, Ashley J. R. Carter, and Chi-Hua Chiu. Gene conversion may aid adaptive peak shifts. Journal of Theoretical Biology, 207(4):495–511, December 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921891.

Handorf:2008:EPM

[HCEK08] Thomas Handorf, Nils Christian, Oliver Ebenhöh, and Daniel Kahn. An environmental perspective on metabolism. Journal of Theoretical Biology, 252(3):530–537, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005437.

Hapca:2007:MNM

[HCM^{+}07] Simona Hapca, John W. Crawford, Keith MacMillan, Mike J. Wilson, and Iain M. Young. Modelling nematode movement using time-fractional dynamics. Journal of Theoretical Biology, 248(1):212–224, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002214.

Hanigan:2001:MMM

[HCMF01] M. D. Hanigan, L. A. Crompton, J. A. Metcalf, and J. France. Modelling mammary metabolism in the dairy cow to predict milk constituent yield, with emphasis on amino acid metabolism and milk protein production: Model construction. Journal of Theoretical Biology, 213(2):223–239, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924178.

Hanigan:2004:IMA

[HCR^{+}04] M. D. Hanigan, L. A. Crompton, C. K. Reynolds, D. Wray-Cahen, M. A. Lomax, and J. France. An integrative model of amino acid metabolism in the liver of the lactating dairy cow. Journal of Theoretical Biology, 228(2):271–289, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000311.

Humphrey:2008:TEC

[HD08] J. A. C. Humphrey and E. S. Dykes. Thermal energy conduction in a honey bee comb due to cell-heating bees. Journal of Theoretical Biology, 250(1):194–208, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004559.

Hagenaars:2004:SHP

[HDF04] T. J. Hagenaars, C. A. Donnelly, and N. M. Ferguson. Spatial heterogeneity and the persistence of infectious diseases. Journal of Theoretical Biology, 229(3):349–359, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001456.

Hariton:2007:SMC

[HdGH07] I. Hariton, G. deBotton, T. C. Gasser, and G. A. Holzapfel. Stress-modulated collagen fiber remodeling in a human carotid bifurcation. Journal of Theoretical Biology, 248(3):460–470, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002792.

Hauert:2002:RDO

[HDHS02] Christoph Hauert, Silvia De Monte, Josef Hofbauer, and Karl Sigmund. Replicator dynamics for optional public good games. Journal of Theoretical Biology, 218(2):187–194, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930675.

Hancock:2004:NER

[HDNC04] John T. Hancock, Radhika Desikan, Steven J. Neill, and Andrew R. Cross. New equations for redox and nano-signal transduction. Journal of Theoretical Biology, 226(1):65–68, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003205.

Haurie:2000:MCN

[HDRM00] Caroline Haurie, David C. Dale, Ryszard Rudnicki, and Michael C. Mackey. Modeling complex neutrophil dynamics in the grey collie. Journal of Theoretical Biology, 204(4):505– 519, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920344.

Huang:2009:PCD

[HDW^{+}09] Ri-Bo Huang, Qi-Shi Du, Yu-Tuo Wei, Zong-Wen Pang, Hang Wei, and Kuo-Chen Chou. Physics and chemistrydriven artificial neural network for predicting bioactivity of peptides and proteins and their design. Journal of Theoretical Biology, 256(3):428–435, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004505.

Hartvigsen:2007:NSV

[HDZ^{+}07] G. Hartvigsen, J. M. Dresch, A. L. Zielinski, A. J. Macula, and C. C. Leary. Network structure, and vaccination strategy and effort interact to affect the dynamics of influenza epidemics. Journal of Theoretical Biology, 246(2):205–213, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005972.

Haegeman:2008:RZS

[HE08a] Bart Haegeman and Rampal S. Etienne. Relaxing the zerosum assumption in neutral biodiversity theory. Journal of Theoretical Biology, 252(2):288–294, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000301.

Hilborn:2008:SCO

[HE08b] Robert C. Hilborn and Jessie D. Erwin. Stochastic coherence in an oscillatory gene circuit model. Journal of Theoretical Biology, 253(2):349–354, July 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001264.

Hellwig:2006:CPM

[HED06] Heino Hellwig, Ralph Engelmann, and Oliver Deussen. Contact pressure models for spiral phyllotaxis and their computer simulation. Journal of Theoretical Biology, 240(3):489–500, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004479.

Halter:2009:CVD

[HEH^{+}09] Michael Halter, John T. Elliott, Joseph B. Hubbard, Alessandro Tona, and Anne L. Plant. Cell volume distributions reveal cell growth rates and division times. Journal of Theoretical Biology, 257(1):124–130, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005742.

Helle:2008:HWB

[Hel08] Samuli Helle. Height, weight, body mass index and offspring sex at birth in contemporary Finnish women. Journal of Theoretical Biology, 252(4):773–775, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000994.

Henson:2004:MSY

[Hen04] Michael A. Henson. Modeling the synchronization of yeast respiratory oscillations. Journal of Theoretical Biology, 231(3):443–458, December 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003303.

Hersh:2000:MAR

[Her00] Jeff Hersh. Mathematical analysis of the relative contributions of decreased production and increased peripheral destruction in idiopathic thrombocytopenic purpura and implications in splenectomy. Journal of Theoretical Biology, 203(2):153–162, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910725.

Hernandez:2009:DNS

[Her09] Maria-Josefina Hernandez. Disentangling nature, strength and stability issues in the characterization of population interactions. Journal of Theoretical Biology, 261(1):107–119, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003038.

Hardersen:2003:IDS

[HF03a] Sönke Hardersen and Chris Frampton. The influence of differential survival on the distribution of fluctuating asymmetry — a modelling approach. Journal of Theoretical Biology, 224(4):479–482, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001954.

Heithaus:2003:ODU

[HF03b] Michael R. Heithaus and Alejandro Frid. Optimal diving under the risk of predation. Journal of Theoretical Biology, 223(1):79–92, July 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000730.

Henery:2002:PFV

[HFGB02] R. Henery, L. Farnell, W. G. Gibson, and M. R. Bennett. Potential fields in vascular smooth muscle generated by transmitter release from sympathetic varicosities. Journal of Theoretical Biology, 218(4):531–548, October 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930985.

Hahnfeldt:2003:MLT

[HFH03] Philip Hahnfeldt, Judah Folkman, and Lynn Hlatky. Minimizing long-term tumor burden: The logic for metronomic chemotherapeutic dosing and its antiangiogenic basis. Journal of Theoretical Biology, 220(4):545–554, February 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931626.

Hanage:2006:IHR

[HFS06] William P. Hanage, Christophe Fraser, and Brian G. Spratt. The impact of homologous recombination on the generation of diversity in bacteria. Journal of Theoretical Biology, 239(2):210–219, March 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003826.

Hayashi:2007:TGO

[HFY07] Motoyuki Hayashi, Noboru Fujita, and Atsushi Yamauchi. Theory of grazing optimization in which herbivory improves photosynthetic ability. Journal of Theoretical Biology, 248(2):367–376, September 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002494.

Haccou:2002:EAG

[HG02] Patsy Haccou and Olivier Glaizot. The ESS in an asymmetric generalized war of attrition with mistakes in role perception. Journal of Theoretical Biology, 214(3):329–349, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924543.

Hatier:2008:FAM

[HG08] Jean-Hugues B. Hatier and Kevin S. Gould. Foliar anthocyanins as modulators of stress signals. Journal of Theoretical Biology, 253(3):625–627, August 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800204X.

Holste:2000:OCP

[HGB^{+}00] Dirk Holste, Ivo Grosse, Sergey V. Buldyrev, H. Eugene Stanley, and Hanspeter Herzel. Optimization of coding potentials using positional dependence of nucleotide frequencies. Journal of Theoretical Biology, 206(4):525–537, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921441.

Haccou:2003:PLS

[HGC03] Patsy Haccou, Olivier Glaizot, and Chris Cannings. Patch leaving strategies and superparasitism: an asymmetric generalized war of attrition. Journal of Theoretical Biology, 225(1):77–89, November 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002236.

Huthmacher:2008:CAP

[HGH08] Carola Huthmacher, Christoph Gille, and Hermann-Georg Holzhütter. A computational analysis of protein interactions in metabolic networks reveals novel enzyme pairs potentially involved in metabolic channeling. Journal of Theoretical Biology, 252(3):456–464, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004742.

Higgins:2007:MCM

[HGP^{+}07] Erin R. Higgins, Pranay Goel, Jose L. Puglisi, Donald M. Bers, Mark Cannell, and James Sneyd. Modelling calcium microdomains using homogenisation. Journal of Theoretical Biology, 247(4):623–644, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001294.

Hely:2001:CMD

[HGV01] Tim A. Hely, Bruce Graham, and Arjen Van Ooyen. A computational model of dendrite elongation and branching based on MAP2 phosphorylation. Journal of Theoretical Biology, 210(3):375–384, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923148.

Hausken:2004:TSH

[HH04a] Kjell Hausken and Jack Hirshleifer. The truthful signalling hypothesis: an explicit general equilibrium model. Journal of Theoretical Biology, 228(4):497–511, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000712.

Hegedus:2004:HWU

[HH04b] Ramón Hegedüs and Gábor Horváth. How and why are uniformly polarization-sensitive retinae subject to polarizationrelated artefacts? Correction of some errors in the theory of polarization-induced false colours. Journal of Theoretical Biology, 230(1):77–87, September 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001869.

Hutchinson:2004:MCW

[HH04c] John M. C. Hutchinson and Konrad Halupka. Mate choice when males are in patches: optimal strategies and good rules of thumb. Journal of Theoretical Biology, 231(1):129–151, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002826. See erra-

tum [HH05b].

Hu:2005:BSP

[HH05a] Xin-Sheng Hu and Fangliang He. Background selection and population differentiation. Journal of Theoretical Biology, 235(2):207–219, July 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000093.

Hutchinson:2005:EMC

[HH05b] John M. C. Hutchinson and Konard Halupka. Erratum to “Mate choice when males are in patches: optimal strategies and good rules of thumb” [Journal of Theoretical Biology 231 (2004) 129–151]. Journal of Theoretical Biology, 232(3):303, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004357. See [HH04c].

Hodgson:2006:SCP

[HH06a] D. J. Hodgson and D. J. Hosken. Sperm competition promotes the exploitation of rival ejaculates. Journal of Theoretical Biology, 243(2):230–234, November 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600275X.

Hruschka:2006:FCE

[HH06b] Daniel J. Hruschka and Joseph Henrich. Friendship, cliquishness, and the emergence of cooperation. Journal of Theoretical Biology, 239(1):1–15, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003097.

Hu:2006:SPF

[HH06c] Xin-Sheng Hu and Fangliang He. Seed and pollen flow in expanding a species’ range. Journal of Theoretical Biology, 240(4):662–672, June 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004741.

Hall:2007:MII

[HH07] Richard J. Hall and Alan Hastings. Minimizing invader impacts: Striking the right balance between removal and restoration. Journal of Theoretical Biology, 249(3):437–444, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004213.

Hofbauer:2008:FCB

[HH08] Josef Hofbauer and Simon M. Huttegger. Feasibility of communication in binary signaling games. Journal of Theoretical Biology, 254(4):843–849, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003548.

Hanin:2006:CSM

[HHBY06] Leonid Hanin, Ollivier Hyrien, Joel Bedford, and Andrei Yakovlev. A comprehensive stochastic model of irradiated cell populations in culture. Journal of Theoretical Biology, 239(4):401–416, April 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003383.

Hegedus:2006:WDD

[HHH06] Ramón Hegedüs, Ákos Horváth, and Gábor Horváth. Why do dusk-active cockchafers detect polarization in the green? The polarization vision in melolontha melolontha is tuned to the high polarized intensity of downwelling light under canopies during sunset. Journal of Theoretical Biology, 238(1):230–244, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002511.

Halford:2005:MAD

[HHP05] William P. Halford, Keith J. Halford, and Amy T. Pierce. Mathematical analysis demonstrates that interferons-β and -γ interact in a multiplicative manner to disrupt herpes simplex virus replication. Journal of Theoretical Biology, 234(3):439– 454, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304006058.

Houser:2001:CDI

[HHR01] D. S. Houser, R. Howard, and S. Ridgway. Can divinginduced tissue nitrogen supersaturation increase the chance of acoustically driven bubble growth in marine mammals? Journal of Theoretical Biology, 213(2):183–195, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924154.

Hakoyama:2000:ERD

[HI00a] Hiroshi Hakoyama and Yoh Iwasa. Extinction risk of a densitydependent population estimated from a time series of population size. Journal of Theoretical Biology, 204(3):337–359, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920198.

Harada:2000:ESI

[HI00b] Kouji Harada and Takashi Ikegami. Evolution of specificity in an immune network. Journal of Theoretical Biology, 203(4):439–449, April 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920010.

Hakoyama:2004:CSU

[HI04] Hiroshi Hakoyama and Yoh Iwasa. Coexistence of a sexual and an unisexual form stabilized by parasites. Journal of Theoretical Biology, 226(2):185–194, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300331X.

Hakoyama:2005:ERM

[HI05] Hiroshi Hakoyama and Yoh Iwasa. Extinction risk of a meta-population: aggregation approach. Journal of Theoretical Biology, 232(2):203–216, January 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003790.

Haeno:2007:PRE

[HI07] Hiroshi Haeno and Yoh Iwasa. Probability of resistance evolution for exponentially growing virus in the host. Journal of Theoretical Biology, 246(2):323–331, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000215.

Hiebeler:2005:SCD

[Hie05] David Hiebeler. Spatially correlated disturbances in a locally dispersing population model. Journal of Theoretical Biology, 232(1):143–149, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003789.

Hirashima:2009:DMB

[HIM09] Tsuyoshi Hirashima, Yoh Iwasa, and Yoshihiro Morishita. Dynamic modeling of branching morphogenesis of ureteric bud in early kidney development. Journal of Theoretical Biology, 259(1):58–66, July 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001258.

Hakoyama:2000:CRF

[HIN00] Hiroshi Hakoyama, Yoh Iwasa, and Junko Nakanishi. Comparing risk factors for population extinction. Journal of Theoretical Biology, 204(3):327–336, June 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920186.

Hayot:2005:FLM

[HJ05] F. Hayot and C. Jayaprakash. A feedforward loop motif in transcriptional regulation: induction and repression. Journal of Theoretical Biology, 234(1):133–143, May 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005466.

Hayot:2006:NOC

[HJ06] F. Hayot and C. Jayaprakash. NF-κB oscillations and cell-to-cell variability. Journal of Theoretical Biology, 240(4):583–591, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004704.

Hui:2007:UAE

[HJ07] Dafeng Hui and Robert B. Jackson. Uncertainty in allometric exponent estimation: a case study in scaling metabolic rate with body mass. Journal of Theoretical Biology, 249(1):168–177, November 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700327X.

Hanzalek:2005:DPA

[HK05] Petr Hanzálek and Jaroslav Kypr. DNA in phosphorus atom representation: the heteronomous double helices of poly(dA)· poly(dT) and poly(dG)· poly(dC) and simulation of the yeast genome and of a human chromosome DNA. Journal of Theoretical Biology, 232(1):83–91, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003522.

Hellweger:2007:IBM

[HK07] Ferdi L. Hellweger and Ehsan Kianirad. Individual-based modeling of phytoplankton: Evaluating approaches for applying the cell quota model. Journal of Theoretical Biology, 249(3):554–565, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004043.

Hardling:2009:DSA

[HK09] Roger Härdling and Kristina Karlsson. The dynamics of sexually antagonistic coevolution and the complex influences of mating system and genetic correlation. Journal of Theoretical Biology, 260(2):276–282, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002446.

Hsieh:2007:IQS

[HKC^{+}07] Ying-Hen Hsieh, Chwan-Chuan King, Cathy W. S. Chen, Mei-Shang Ho, Sze-Bi Hsu, and Yi-Chun Wu. Impact of quarantine on the 2003 SARS outbreak: a retrospective modeling study. Journal of Theoretical Biology, 244(4):729–736, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600422X.

Holcman:2007:MHI

[HKP07] D. Holcman, V. Kasatkin, and A. Prochiantz. Modeling homeoprotein intercellular transfer unveils a parsimonious mechanism for gradient and boundary formation in early brain development. Journal of Theoretical Biology, 249(3):503–517, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003591.

Hagedorn:2000:PIG

[HL00] Thomas R. Hagedorn and Laura F. Landweber. Phylogenetic invariants and geometry. Journal of Theoretical Biology, 205(3):365–376, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930092071X.

Hanlon:2005:CMD

[HL05] Phil Hanlon and Andy Lorenz. A computational method to detect epistatic effects contributing to a quantitative trait. Journal of Theoretical Biology, 235(3):350–364, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930500041X.

Hu:2006:AGV

[HL06] Xin-Sheng Hu and Bailian Li. Additive genetic variation and the distribution of QTN effects among sites. Journal of Theoretical Biology, 243(1):76–85, November 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002463.

Handel:2009:ARC

[HLA09] Andreas Handel, Ira M. Longini, and Rustom Antia. Antiviral resistance and the control of pandemic influenza: The roles of stochasticity, evolution and model details. Journal of Theoretical Biology, 256(1):117–125, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800492X.

Harrison:2001:DMC

[HLH01] Mary Ann Harrison, Ying-Cheng Lai, and Robert D. Holt. Dynamical mechanism for coexistence of dispersing species. Journal of Theoretical Biology, 213(1):53–72, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930192404X.

Howe:2006:CBE

[HLP06] Catherine Q. Howe, R. Beau Lotto, and Dale Purves. Comparison of Bayesian and empirical ranking approaches to visual perception. Journal of Theoretical Biology, 241(4):866–875, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000300.

Hyman:2001:ISM

[HLS01] James M. Hyman, Jia Li, and E. Ann Stanley. The initialization and sensitivity of multigroup models for the transmission of HIV. Journal of Theoretical Biology, 208(2):227–249, January 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922148.

Hirsch:2000:MRC

[HLW00] H. R. Hirsch, X. Liu, and T. M. Witten. Mortality-rate crossovers and maximum lifespan in advantaged and disadvantaged populations: Accelerated-mortality and sudden-death models. Journal of Theoretical Biology, 205(2):171–180, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920630.

Hanegraaf:2001:DMC

[HM01] P. P. F. Hanegraaf and E. B. Muller. The dynamics of the macromolecular composition of biomass. Journal of Theoretical Biology, 212(2):237–251, September 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923690.

Hewer:2005:PID

[HM05] R. Hewer and D. Meyer. Peptide immunogens designed to enhance immune responses against human immunodeficiency virus (HIV) mutant strains: a plausible means of preventing viral persistence. Journal of Theoretical Biology, 233(1):85–90, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004643.

Hartmann:2006:MVL

[HM06] Dirk Hartmann and Takashi Miura. Modelling in vitro lung branching morphogenesis during development. Journal of Theoretical Biology, 242(4):862–872, October 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001950.

Hiebeler:2007:ESD

[HM07] David E. Hiebeler and Benjamin R. Morin. The effect of static and dynamic spatially structured disturbances on a locally dispersing population. Journal of Theoretical Biology, 246(1):136–144, May 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005935.

Hardway:2008:MPR

[HMB^{+}08] Heather Hardway, Bibhash Mukhopadhyay, Timothy Burke, Theron James Hitchman, and Robin Forman. Modeling the precision and robustness of hunchback border during drosophila embryonic development. Journal of Theoretical Biology, 254(2):390–399, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002671.

Hancock:2006:MMW

[HMGK06] Penelope A. Hancock, E. J. Milner-Gulland, and Matthew J. Keeling. Modelling the many-wrongs principle: The navigational advantages of aggregation in nomadic foragers. Journal of Theoretical Biology, 240(2):302–310, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004236.

Ho:2004:OMP

[HML04] Melissa D. Ho, Bryan C. McCannon, and Jonathan P. Lynch. Optimization modeling of plant root architecture for water and phosphorus acquisition. Journal of Theoretical Biology, 226(3):331–340, February 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003539.

Handel:2009:ERI

[HML09] Andreas Handel, Elisa Margolis, and Bruce R. Levin. Exploring the role of the immune response in preventing antibiotic resistance. Journal of Theoretical Biology, 256(4):655–662, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005754.

Holtta:2009:LPF

[HMN09] T. Hölttä, M. Mencuccini, and E. Nikinmaa. Linking phloem function to structure: Analysis with a coupled xylem-phloem transport model. Journal of Theoretical Biology, 259(2):325– 337, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001441.

Hauert:2006:SDC

[HMND06] Christoph Hauert, Franziska Michor, Martin A. Nowak, and Michael Doebeli. Synergy and discounting of cooperation in social dilemmas. Journal of Theoretical Biology, 239(2):195– 202, March 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003802.

Hurley:2004:MES

[HMP04] M. A. Hurley, P. Matthiessen, and A. D. Pickering. A model for environmental sex reversal in fish. Journal of Theoretical Biology, 227(2):159–165, March 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003977.

Huck:2003:ZKC

[HNO03] Steffen Huck, Hans-Theo Normann, and Jörg Oechssler. Zero-knowledge cooperation in dilemma games. Journal of Theoretical Biology, 220(1):47–54, January 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931298.

Hutchinson:2007:IME

[HNTHA07] John R. Hutchinson, Victor Ng-Thow-Hing, and Frank C. Anderson. A 3d interactive method for estimating body segmental parameters in animals: Application to the turning and running performance of Tyrannosaurus rex. Journal of Theoretical Biology, 246(4):660–680, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000550.

Hoffmann:2004:PAA

[Hof04] Geoffrey W. Hoffmann. Proteomic analyser with applications to diagnostics and vaccines. Journal of Theoretical Biology, 228(4):459–465, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000682. See erratum [Ano05h].

Hogeweg:2000:EMM

[Hog00] P. Hogeweg. Evolving mechanisms of morphogenesis: on the interplay between differential adhesion and cell differentiation. Journal of Theoretical Biology, 203(4):317–333, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910877.

Hokkanen:2000:CPV

[Hok00] J. E. I. Hokkanen. Chaotic or periodic variation? Looking at crustacea hearts. Journal of Theoretical Biology, 203(4):451– 454, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910968.

Holzapfel:2006:DMM

[Hol06] Gerhard A. Holzapfel. Determination of material models for arterial walls from uniaxial extension tests and histological structure. Journal of Theoretical Biology, 238(2):290–302, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002080.

Hopfer:2002:MDT

[Hop02] Ulrich Hopfer. A Maxwell’s demon type of membrane transport: Possibility for active transport by ABC-type transporters? Journal of Theoretical Biology, 214(4):539–547, February 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924798.

Hopenfeld:2006:MAA

[Hop06] B. Hopenfeld. A mathematical analysis of the action potential plateau duration of a human ventricular myocyte. Journal of Theoretical Biology, 240(2):311–322, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500425X.

Horie:2008:OFB

[Hor08] Ryota Horie. An optimization framework of biological dynamical systems. Journal of Theoretical Biology, 253(1):45– 54, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000957.

Hotton:2003:FCP

[Hot03] Scott Hotton. Finding the center of a phyllotactic pattern. Journal of Theoretical Biology, 225(1):15–32, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002170.

Howland:2009:OOV

[How09] H. C. Howland. Orbital orientation is not visual orientation: a comment on “X-ray Vision and the evolution of forwardfacing eyes” by M. A. Changizi and S. Shimojo. Journal of Theoretical Biology, 257(3):522–523, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006589.

Hayflick:2000:BGR

[HP00a] Leonard Hayflick and Stanley Plotkin. Bridging the gap — a reply. Journal of Theoretical Biology, 207(4):577–578, December 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921908.

Hill:2000:CSP

[HP00b] N. A. Hill and L. A. Plumpton. Control strategies for the polarotactic orientation of the microorganism euglena gracilis. Journal of Theoretical Biology, 203(4):357–365, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910907.

Huhn:2004:ICS

[HP04] Manfred Hühn and Hans-Peter Piepho. Inbreeding coefficients for stochastically varying small population sizes-bias of calculation based on effective numbers. Journal of Theoretical Biology, 226(4):467–475, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003795.

Horova:2009:HFC

[HPZ09] Ivana Horová, Zdenek Pospísil, and Jirí Zelinka. Hazard function for cancer patients and cancer cell dynamics. Journal of Theoretical Biology, 258(3):437–443, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003081.

He:2008:NPI

[HQP08] Yun He, Sanbo Qin, and Xian-Ming Pan. A novel procedure for identification of the folding/unfolding patterns of dimeric proteins. Journal of Theoretical Biology, 250(3):461–467, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005085.

Huang:2009:CTF

[HQP^{+}09] Jianyong Huang, Lei Qin, Xiaoling Peng, Tao Zhu, Chunyang Xiong, Youyi Zhang, and Jing Fang. Cellular traction force recovery: An optimal filtering approach in two-dimensional Fourier space. Journal of Theoretical Biology, 259(4):811–819, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002306.

Hurst:2000:DDD

[HR00] Laurence D. Hurst and James P. Randerson. Dosage, deletions and dominance: Simple models of the evolution of gene expression. Journal of Theoretical Biology, 205(4):641–647, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920952.

Huang:2008:SCR

[HR08] Fenix W. D. Huang and Christian M. Reidys. Statistics of canonical RNA pseudoknot structures. Journal of Theoretical Biology, 253(3):570–578, August 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001768.

Hutt:2002:PCP

[HRBL02] M.-Th. Hütt, U. Rascher, F. Beck, and U. Lüttge. Period-2 cycles and 2:1 phase locking in a biological clock driven by temperature pulses. Journal of Theoretical Biology, 217(3):383–390, August 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930316.

Harnos:2006:AEI

[HRSV06] Andrea Harnos, Jenő Reiczigel, Norbert Solymosi, and Gábor Vattay. Analysis of the effect of immunization in rabies time series. Journal of Theoretical Biology, 240(1):72–77, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003656.

Hosack:2008:CVB

[HRvdD08] Geoffrey R. Hosack, Philippe A. Rossignol, and P. van den Driessche. The control of vector-borne disease epidemics. Journal of Theoretical Biology, 255(1):16–25, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003998.

Hanin:2006:SMS

[HRZ06] Leonid Hanin, Jason Rose, and Marco Zaider. A stochastic model for the sizes of detectable metastases. Journal of Theoretical Biology, 243(3):407–417, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002967.

Haraguchi:2000:EPV

[HS00a] Yoshihiro Haraguchi and Akira Sasaki. The evolution of parasite virulence and transmission rate in a spatially structured population. Journal of Theoretical Biology, 203(2):85– 96, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910652.

Hazel:2000:ICS

[HS00b] Wade Hazel and Richard Smock. Inheritance in the conditional strategy revisited. Journal of Theoretical Biology, 204(2):307– 309, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920137.

Holmes:2000:MMT

[HS00c] M. J. Holmes and B. D. Sleeman. A mathematical model of tumour angiogenesis incorporating cellular traction and viscoelastic effects. Journal of Theoretical Biology, 202(2):95–112, January 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251939991038X.

Han:2001:CMD

[HS01] Bo-Ping Han and Milan Straskraba. Control mechanisms of diel vertical migration: Theoretical assumptions. Journal of Theoretical Biology, 210(3):305–318, June 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923070.

Hauert:2002:SAS

[HS02a] Christoph Hauert and Olaf Stenull. Simple adaptive strategy wins the Prisoner’s dilemma. Journal of Theoretical Biology, 218(3):261–272, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930729.

Hernandez-Suarez:2002:MCA

[HS02b] Carlos M. Hernández-Suárez. A Markov chain approach to calculate R_{0} in stochastic epidemic models. Journal of Theoretical Biology, 215(1):83–93, March 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924968.

Hug:2003:MNM

[HS03] Hubert Hug and Rainer Schuler. Measurement of the number of molecules of a single mRNA species in a complex mRNA preparation. Journal of Theoretical Biology, 221(4):615–624, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932115.

Hordijk:2004:DAS

[HS04] Wim Hordijk and Mike Steel. Detecting autocatalytic, selfsustaining sets in chemical reaction systems. Journal of Theoretical Biology, 227(4):451–461, April 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004417.

Huertas:2007:DLD

[HS07] Marco A. Huertas and Gregory D. Smith. The dynamics of luminal depletion and the stochastic gating of Ca^{2+}- activated Ca^{2+} channels and release sites. Journal of Theoretical Biology, 246(2):332–354, May 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000070.

Hilker:2008:DIS

[HS08] Frank M. Hilker and Kirsten Schmitz. Disease-induced stabilization of predator-prey oscillations. Journal of Theoretical Biology, 255(3):299–306, December 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004268.

Hoffman:2001:CCP

[HSB01] A. Hoffman, Lee M. Spetner, and M. Burke. Cessation of cell proliferation by adjustment of cell redox potential. Journal of Theoretical Biology, 211(4):403–407, August 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923562.

Hancioglu:2007:DMH

[HSC07] Baris Hancioglu, David Swigon, and Gilles Clermont. A dynamical model of human immune response to influenza A virus infection. Journal of Theoretical Biology, 246(1):70–86, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005820.

Higgins:2009:WTD

[HSF09] Erin R. Higgins, Heiko Schmidle, and Martin Falcke. Waiting time distributions for clusters of IP3 receptors. Journal of Theoretical Biology, 259(2):338–349, July 21, 2009.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001295.

Haeri:2005:MPT

[HSG05] Mohammad Haeri, Yashar Sarbaz, and Shahriar Gharibzadeh. Modeling the Parkinson’s tremor and its treatments. Journal of Theoretical Biology, 236(3):311–322, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001232.

Hwang:2008:PIN

[HSK^{+}08] Sohyun Hwang, Seung-Woo Son, Sang Cheol Kim, Young Joo Kim, Hawoong Jeong, and Doheon Lee. A protein interaction network associated with asthma. Journal of Theoretical Biology, 252(4):722–731, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000647.

Haugh:2004:CRP

[HSL04] Jason M. Haugh, Ian C. Schneider, and Jodee M. Lewis. On the cross-regulation of protein tyrosine phosphatases and receptor tyrosine kinases in intracellular signaling. Journal of Theoretical Biology, 230(1):119–132, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001894.

Hong:2007:ENI

[HSMM07] Dawei Hong, William M. Saidel, Shushuang Man, and Joseph V. Martin. Extracellular noise-induced stochastic synchronization in heterogeneous quorum sensing network. Journal of Theoretical Biology, 245(4):726–736, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005662.

Heino:2007:BFB

[HTCS07] Jenni Heino, Knarik Tunyan, Daniela Calvetti, and Erkki Somersalo. Bayesian flux balance analysis applied to a skeletal muscle metabolic model. Journal of Theoretical Biology, 248(1):91–110, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001841.

Honda:2004:ETD

[HTN04a] Hisao Honda, Masaharu Tanemura, and Tatsuzo Nagai. Erratum to “A three-dimensional vertex dynamics cell model of space-filling polyhedra simulating cell behavior in a cell aggregate”: [J. Theor. Biol. 226 (2004) 439–453]. Journal of Theoretical Biology, 228(1):147, May 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000591. See [HTN04b].

Honda:2004:TDV

[HTN04b] Hisao Honda, Masaharu Tanemura, and Tatsuzo Nagai. A three-dimensional vertex dynamics cell model of space-filling polyhedra simulating cell behavior in a cell aggregate. Journal of Theoretical Biology, 226(4):439–453, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003680. See [HTN04a].

Huang:2003:ECB

[Hua03] Xinyi Huang. Equilibrium competition binding assay: inhibition mechanism from a single dose response. Journal of Theoretical Biology, 225(3):369–376, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002650.

Hui:2009:SPS

[Hui09] Cang Hui. On the scaling patterns of species spatial distribution and association. Journal of Theoretical Biology, 261(3):481–487, December 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003853.

Hulbert:2005:IFA

[Hul05] A. J. Hulbert. On the importance of fatty acid composition of membranes for aging. Journal of Theoretical Biology, 234(2):277–288, May 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005788.

Hurd:2006:RHP

[Hur06] Peter L. Hurd. Resource holding potential, subjective resource value, and game theoretical models of aggressiveness signalling. Journal of Theoretical Biology, 241(3):639–648, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000026.

Holtta:2007:MBG

[HVN07] T. Hölttä, T. Vesala, and E. Nikinmaa. A model of bubble growth leading to xylem conduit embolism. Journal of Theoretical Biology, 249(1):111–123, November 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002743.

Holtta:2002:RBE

[HVPN02] T. Hölttä, T. Vesala, M. Perämäki, and E. Nikinmaa. Relationships between embolism, stem water tension, and diameter changes. Journal of Theoretical Biology, 215(1):23–38, March 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924853.

Han:2007:STN

[HVY^{+}07] Zhangang Han, Thomas M. Vondriska, Ling Yang, W. Robb MacLellan, James N. Weiss, and Zhilin Qu. Signal transduction network motifs and biological memory. Journal of Theoretical Biology, 246(4):755–761, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000525.

Hofmeyr:2001:BCP

[HW01] Jan-Hendrik S. Hofmeyr and Hans V. Westerhoff. Building the cellular puzzle: Control in multi-level reaction networks. Journal of Theoretical Biology, 208(3):261–285, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922161.

Heffernan:2005:MCE

[HW05] Jane M. Heffernan and Lindi M. Wahl. Monte Carlo estimates of natural variation in HIV infection. Journal of Theoretical Biology, 236(2):137–153, September 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000913.

Heffernan:2006:NVH

[HW06] Jane M. Heffernan and Lindi M. Wahl. Natural variation in HIV infection: Monte Carlo estimates that include CD8 effector cells. Journal of Theoretical Biology, 243(2):191–204, November 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002311.

Howard:2009:GCA

[HW09] Gregory J. Howard and Thomas F. Webster. Generalized concentration addition: a method for examining mixtures containing partial agonists. Journal of Theoretical Biology, 259(3):469–477, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001386.

Heinz:2007:IIM

[HWCF07] Simone K. Heinz, Christian Wissel, Larissa Conradt, and Karin Frank. Integrating individual movement behaviour into dispersal functions. Journal of Theoretical Biology, 245(4):601–609, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519306005704.

Horiuchi:2002:SPS

[HYA02] Shiro Horiuchi, Norio Yamamura, and Takuya Abe. Soldier production strategy in lower termites: From Young instars or old instars? Journal of Theoretical Biology, 218(2):195–205, September 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930687.

Zeng:2009:UAC

[hZzGqX^{+}09] Yu hong Zeng, Yan zhi Guo, Rong quan Xiao, Li Yang, Le zheng Yu, and Meng long Li. Using the augmented Chou’s pseudo amino acid composition for predicting protein submitochondria locations based on auto covariance approach. Journal of Theoretical Biology, 259(2):366–372, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001301.

Indic:2006:CAD

[IB06] Premananda Indic and Emery N. Brown. Characterizing the amplitude dynamics of the human core-temperature circadian rhythm using a stochastic-dynamic model. Journal of Theoretical Biology, 239(4):499–506, April 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003553.

Imhof:2007:TTW

[IFN07] Lorens A. Imhof, Drew Fudenberg, and Martin A. Nowak. Tit-for-tat or win-stay, lose-shift? Journal of Theoretical Biology, 247(3):574–580, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001439.

Ivanek:2007:MCA

[IGHW07] Renata Ivanek, Yrjö T. Gröhn, Alphina Jui-Jung Ho, and Martin Wiedmann. Markov chain approach to analyze the dynamics of pathogen fecal shedding — example of Listeria monocytogenes shedding in a herd of dairy cattle. Journal of Theoretical Biology, 245(1):44–58, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004553.

Iglic:2006:CIA

[IHV^{+}06] Ales Iglic, Henry Hägerstrand, Peter Veranic, Ana Plemenitas, and Veronika Kralj-Iglic. Curvature-induced accumulation of anisotropic membrane components and raft formation in cylindrical membrane protrusions. Journal of Theoretical Biology, 240(3):368–373, June 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004248.

Ito:2006:FWF

[II06] Hiroshi C. Ito and Takashi Ikegami. Food-web formation with recursive evolutionary branching. Journal of Theoretical Biology, 238(1):1–10, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002055.

Iguchi:2007:BDK

[IiKY07] Kazumoto Iguchi, Shu ichi Kinoshita, and Hiroaki S. Yamada. Boolean dynamics of Kauffman models with a scale-free network. Journal of Theoretical Biology, 247(1):138–151, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307000860.

Iima:2007:TDA

[Iim07] Makoto Iima. A two-dimensional aerodynamic model of freely flying insects. Journal of Theoretical Biology, 247(4):657–671, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001178.

Inada:2002:OFM

[IK02] Yoshinobu Inada and Keiji Kawachi. Order and flexibility in the motion of fish schools. Journal of Theoretical Biology, 214(3):371–387, February 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192449X.

Ishihara:2006:TPP

[IK06] Shuji Ishihara and Kunihiko Kaneko. Turing pattern with proportion preservation. Journal of Theoretical Biology, 238(3):683–693, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002742.

Ifti:2004:ENS

[IKD04] Margarita Ifti, Timothy Killingback, and Michael Doebeli. Effects of neighbourhood size and connectivity on the spatial continuous Prisoner’s dilemma. Journal of Theoretical Biology, 231(1):97–106, November 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002759.

Kinoshita:2009:IPB

[iKIY09] Shu ichi Kinoshita, Kazumoto Iguchi, and Hiroaki S. Yamada. Intrinsic properties of Boolean dynamics in complex networks. Journal of Theoretical Biology, 256(3):351–369, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005456.

Iwata:2000:DMG

[IKS00] K. Iwata, K. Kawasaki, and N. Shigesada. A dynamical model for the growth and size distribution of multiple metastatic tumors. Journal of Theoretical Biology, 203(2):177–186, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910750.

Irisson:2004:STC

[ILDP04] Jean-Olivier Irisson, Anselme LeVan, Michel De Lara, and Serge Planes. Strategies and trajectories of coral reef fish larvae optimizing self-recruitment. Journal of Theoretical Biology, 227(2):205–218, March 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004181.

Iber:2002:MMG

[IM02] Dagmar Iber and Philip K. Maini. A mathematical model for germinal centre kinetics and affinity maturation. Journal of Theoretical Biology, 219(2):153–175, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930791.

Iwasa:2005:PGT

[IMKN05] Yoh Iwasa, Franziska Michor, Natalia L. Komarova, and Martin A. Nowak. Population genetics of tumor suppressor genes. Journal of Theoretical Biology, 233(1):15–23, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004242.

Iwasa:2004:SBP

[IMN04a] Yoh Iwasa, Franziska Michor, and Martin Nowak. Some basic properties of immune selection. Journal of Theoretical Biology, 229(2):179–188, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001158.

Iwasa:2004:EDI

[IMN04b] Yoh Iwasa, Franziska Michor, and Martin A. Nowak. Evolutionary dynamics of invasion and escape. Journal of Theoretical Biology, 226(2):205–214, January 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003333.

Iwasa:2005:VEW

[IMN05] Yoh Iwasa, Franziska Michor, and Martin A. Nowak. Virus evolution within patients increases pathogenicity. Journal of Theoretical Biology, 232(1):17–26, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003467.

Iwami:2009:IIH

[IMNT09] Shingo Iwami, Tomoyuki Miura, Shinji Nakaoka, and Yasuhiro Takeuchi. Immune impairment in HIV infection: Existence of risky and immunodeficiency thresholds. Journal of Theoretical Biology, 260(4):490–501, October 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003002.

Ireland:2004:ESH

[ING04] J. M. Ireland, R. A. Norman, and J. V. Greenman. The effect of seasonal host birth rates on population dynamics: the importance of resonance. Journal of Theoretical Biology, 231(2):229–238, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002905.

Ivanchenko:2008:PNM

[INSR08] Mikhail V. Ivanchenko, Thomas Nowotny, Allen I. Selverston, and Mikhail I. Rabinovich. Pacemaker and network mechanisms of rhythm generation: Cooperation and competition. Journal of Theoretical Biology, 253(3):452–461, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002002.

Ishikawa:2007:ORT

[IPY07] T. Ishikawa, T. J. Pedley, and T. Yamaguchi. Orientational relaxation time of bottom-heavy squirmers in a semi-dilute suspension. Journal of Theoretical Biology, 249(2):296–306, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003608.

Irons:2009:LAB

[Iro09] D. J. Irons. Logical analysis of the budding yeast cell cycle. Journal of Theoretical Biology, 257(4):543–559, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006619.

Ingalls:2003:SAS

[IS03] Brian P. Ingalls and Herbert M. Sauro. Sensitivity analysis of stoichiometric networks: an extension of metabolic control analysis to non-steady state trajectories. Journal of Theoretical Biology, 222(1):23–36, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000110.

Ishihara:2008:MIN

[IS08] Shuji Ishihara and Tatsuo Shibata. Mutual interaction in network motifs robustly sharpens gene expression in developmental processes. Journal of Theoretical Biology, 252(1):131–144, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000489.

Ito:2009:ODS

[IS09] Kiyoshi Ito and Satoki Sakai. Optimal defense strategy against herbivory in plants: Conditions selecting for induced defense, constitutive defense, and no-defense. Journal of Theoretical Biology, 260(3):453–459, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003130.

Isasi:2001:SMI

[ISC01] S. Cossy Isasi, G. J. Sibona, and C. A. Condat. A simple model for the interaction between T. cruzi and its antibodies during Chagas infection. Journal of Theoretical Biology, 208(1):1–13, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930092191X.

Ishii:2000:SMM

[Ish00] Haruhiko Ishii. A statistical-mechanical model for regulation of long-range chromatin structure and gene expression. Journal of Theoretical Biology, 203(3):215–228, April 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910816.

Imai:2001:ISC

[IST01] Hirotami T. Imai, Yoko Satta, and Naoyuki Takahata. Integrative study on chromosome evolution of mammals, ants and wasps based on the minimum interaction theory. Journal of Theoretical Biology, 210(4):475–497, June 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923276.

Imhof:2004:ODE

[ISW04] Lorens A. Imhof, Dale Song, and Weng Kee Wong. Optimal design of experiments with anticipated pattern of missing observations. Journal of Theoretical Biology, 228(2):251–260, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000293.

Imai:2002:EHC

[ISWT02] Hirotami T. Imai, Yoko Satta, Masayasu Wada, and Naoyuki Takahata. Estimation of the highest chromosome number of eukaryotes based on the minimum interaction theory. Journal of Theoretical Biology, 217(1):61–74, July 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930293016X.

Iwami:2009:MDV

[ITI^{+}09] Shingo Iwami, Yasuhiro Takeuchi, Kentaro Iwamoto, Yoshimi Naruo, and Masahiro Yasukawa. A mathematical design of vector vaccine against autoimmune disease. Journal of Theoretical Biology, 256(3):382–392, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005146.

Iwami:2008:PAI

[ITKL08] Shingo Iwami, Yasuhiro Takeuchi, Andrei Korobeinikov, and Xianning Liu. Prevention of avian influenza epidemic: What policy should we choose? Journal of Theoretical Biology, 252(4):732–741, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000817.

Iwami:2009:AFP

[ITL09] Shingo Iwami, Yasuhiro Takeuchi, and Xianning Liu. Avian flu pandemic: Can we prevent it? Journal of Theoretical Biology, 257(1):181–190, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005997.

Iwami:2009:GSV

[ITLN09] Shingo Iwami, Yasuhiro Takeuchi, Xianning Liu, and Shinji Nakaoka. A geographical spread of vaccine-resistance in avian influenza epidemics. Journal of Theoretical Biology, 259(2):219–228, July 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001453.

Iwami:2007:DPA

[ITM^{+}07] Shingo Iwami, Yasuhiro Takeuchi, Yoshiharu Miura, Toru Sasaki, and Tsuyoshi Kajiwara. Dynamical properties of autoimmune disease models: Tolerance, flare-up, dormancy. Journal of Theoretical Biology, 246(4):646–659, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000501.

Imafuku:2009:HSP

[ITT09] Yasuhiro Imafuku, Neil Thomas, and Katsuhisa Tawada. Hopping and stalling of processive molecular motors. Journal of Theoretical Biology, 261(1):43–49, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003233.

Tainaka:2006:SDD

[iTYU06] Kei ichi Tainaka, Jin Yoshimura, and Takashi Ushimaru. Stage-dependent density effect in the cell cycle of budding yeast. Journal of Theoretical Biology, 242(3):736–742, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001688.

Ilyina:2002:MML

[IV02] Ekaterina A. Ilyina and Yuri B. Vakhtin. Molecular model for limb development. Journal of Theoretical Biology, 216(2):159– 177, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930292549X.

Isaksson:2008:DMI

[IvDH^{+}08] Hanna Isaksson, Corrinus C. van Donkelaar, Rik Huiskes, Jiang Yao, and Keita Ito. Determining the most important cellular characteristics for fracture healing using design of experiments methods. Journal of Theoretical Biology, 255(1):26–39, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003986.

Isaksson:2008:CMR

[IvDHI08a] Hanna Isaksson, Corrinus C. van Donkelaar, Rik Huiskes, and Keita Ito. Corrigendum to “A mechano-regulatory bone-healing model incorporating cell-phenotype specific activity”: [J. Theor. Biol. 252 (2008) 230–246]. Journal of Theoretical Biology, 254(3):717, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800355X. See [IvDHI08b].

Isaksson:2008:MRB

[IvDHI08b] Hanna Isaksson, Corrinus C. van Donkelaar, Rik Huiskes, and Keita Ito. A mechano-regulatory bone-healing model incorporating cell-phenotype specific activity. Journal of Theoretical Biology, 252(2):230–246, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519308000507. See corrigendum [IvDHI08a].

Ichinose:2008:RTT

[IYGA08] N. Ichinose, T. Yada, O. Gotoh, and K. Aihara. Reconstruction of transcription-translation dynamics with a model of gene networks. Journal of Theoretical Biology, 255(4):378–386, December 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004670.

Israeli:2005:PPS

[IZGG05] David Israeli, Simin Ziaei, Patrick Gonin, and Luis Garcia. A proposal for the physiological significance of mdr1 and Bcrp1/ Abcg2 gene expression in normal tissue regeneration and after cancer therapy. Journal of Theoretical Biology, 232(1):41–45, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003480.

Izsak:2005:PSF

[Izs05] János Izsák. A pilot study on the frequency structure of histological neoplasm diagnoses in rats. Journal of Theoretical Biology, 236(4):427–437, October 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001335.

Jabbarzadeh:2005:CRM

[JA05] Ehsan Jabbarzadeh and Cameron F. Abrams. Chemotaxis and random motility in unsteady chemoattractant fields: a computational study. Journal of Theoretical Biology, 235(2):221–232, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000214.

Jahandideh:2007:PPR

[JAA^{+}07] Samad Jahandideh, Ebrahim Barzegari Asadabadi, Parviz Abdolmaleki, Mina Jahandideh, and Somayyeh Hoseini. Protein psychrophilicity: Role of residual structural properties in adaptation of proteins to low temperatures. Journal of Theoretical Biology, 248(4):721–726, October 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002986.

Jackson:2003:IAM

[Jac03] Trachette L. Jackson. Intracellular accumulation and mechanism of action of doxorubicin in a spatio-temporal tumor model. Journal of Theoretical Biology, 220(2):201–213, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931560.

Jackle:2007:CTR

[Jäc07] Josef Jäckle. The causal theory of the resting potential of cells. Journal of Theoretical Biology, 249(3):445–463, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003384.

Jansen:2005:CBC

[JAFW05] Vincent A. A. Jansen, Hester Korthals Altes, Georg A. Funk, and Dominik Wodarz. Contrasting B cell- and T cell-based protective vaccines. Journal of Theoretical Biology, 234(1):39– 48, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005272.

Jager:2008:ESC

[Jäg08] Gerhard Jäger. Evolutionary stability conditions for signaling games with costly signals. Journal of Theoretical Biology, 253(1):131–141, July 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000945.

Jinha:2009:TSV

[JAHKH09] A. Jinha, R. Ait-Haddou, M. Kaya, and W. Herzog. A taskspecific validation of homogeneous non-linear optimisation approaches. Journal of Theoretical Biology, 259(4):695–700, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001829.

Jahandideh:2007:SSP

[JAJA07] Samad Jahandideh, Parviz Abdolmaleki, Mina Jahandideh, and Ebrahim Barzegari Asadabadi. Sequence and structural parameters enhancing adaptation of proteins to low temperatures. Journal of Theoretical Biology, 246(1):159–166, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005698.

Jahandideh:2007:NHM

[JAJH07] Samad Jahandideh, Parviz Abdolmaleki, Mina Jahandideh, and Sayyed Hamed Sadat Hayatshahi. Novel hybrid method for the evaluation of parameters contributing in determination of protein structural classes. Journal of Theoretical Biology, 244(2):275–281, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003626.

James:2000:HHC

[Jam00] William H. James. The hypothesized hormonal control of offspring sex ratio: Evidence from families ascertained by schizophrenia and epilepsy. Journal of Theoretical Biology, 206(3):445–447, October 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921453.

James:2001:FLR

[Jam01a] William H. James. Finger-length ratios, sexual orientation and offspring sex ratios. Journal of Theoretical Biology, 212(3):273–274, October 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923768.

James:2001:HBW

[Jam01b]  William H. James.  Handedness, birth weight, mor-

tality and Barker’s hypothesis. Journal of Theoretical Biology, 210(3):345–346, June 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923094.

James:2001:HGH

[Jam01c] William H. James. Hypothesis: Gonadal hormones act as confounders in epidemiological studies of the associations between some behavioural risk factors and some pathological conditions. Journal of Theoretical Biology, 209(1):97–102, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922495.

James:2001:PSR

[Jam01d] William H. James. A prediction on the sex ratios of piglets. Journal of Theoretical Biology, 209(1):1–2, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922355.

James:2004:HHS

[Jam04] William. H. James. Hypotheses on how selection for some traits in rodents led to correlated responses in offspring sex ratios. Journal of Theoretical Biology, 228(1):1–6, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003801.

James:2006:PCA

[Jam06] William H. James. Possible constraints on adaptive variation in sex ratio at birth in humans and other primates. Journal of Theoretical Biology, 238(2):383–394, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002249.

James:2007:SHO

[Jam07] William H. James. Support for a hypothesis on the offspring sex ratios of hepatitis B carriers. Journal of Theoretical Biology, 247(2):395, July 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000835.

James:2008:SCP

[Jam08a] William H. James. Some comments on the paper of Grant (2007). Journal of Theoretical Biology, 253(2):401–404, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001239.

James:2008:VHS

[Jam08b] William H. James. The variations of human sex ratio at birth with time of conception within the cycle, coital rate around the time of conception, duration of time taken to achieve conception, and duration of gestation: a synthesis. Journal of Theoretical Biology, 255(2):199–204, November 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003731.

James:2008:WAW

[Jam08c] William H. James. Women’s age, weight, parity and offspring sex ratio: a comment on the paper of helle. Journal of Theoretical Biology, 254(3):716, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003524.

James:2009:VDC

[Jam09a] William H. James. The variances of the distributions of the combinations of the sexes within mammalian litters: Notes to mark the centenary of the problem. Journal of Theoretical Biology, 259(2):264–268, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001337.

James:2009:VHS

[Jam09b] William H. James. The variations of human sex ratio at birth during and after wars, and their potential explanations. Journal of Theoretical Biology, 257(1):116–123, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308004967.

Javorszky:2000:IBS

[Jav00] Karl Javorszky. Interaction between sequences and mixtures. Journal of Theoretical Biology, 205(4):663–666, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920885.

Jha:2007:PSD

[JAV07] Anupam Nath Jha, G. K. Ananthasuresh, and Saraswathi Vishveshwara. Protein sequence design based on the topology of the native state structure. Journal of Theoretical Biology, 248(1):81–90, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002007.

Jazwinski:2001:IAA

[Jaz01] S. Michał Jazwinski. Integrative approaches to aging biological systems. Journal of Theoretical Biology, 213(4):505–507, December 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924397.

Jalife:2004:MMG

[JB04] José Jalife and Omer Berenfeld. Molecular mechanisms and global dynamics of fibrillation: an integrative approach to the underlying basis of vortex-like reentry. Journal of Theoretical Biology, 230(4):475–487, October 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002383.

Jeune:2006:SAP

[JB06] Bernard Jeune and Denis Barabé. A stochastic approach to phyllotactic patterns analysis. Journal of Theoretical Biology, 238(1):52–59, January 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002663.

Janssen:2008:ECA

[JB08] Marco A. Janssen and Clint Bushman. Evolution of cooperation and altruistic punishment when retaliation is possible. Journal of Theoretical Biology, 254(3):541–545, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003299.

Jeanson:2003:MAM

[JBF^{+}03] Raphaël Jeanson, Stéphane Blanco, Richard Fournier, Jean-Louis Deneubourg, Vincent Fourcassié, and Guy Theraulaz. A model of animal movements in a bounded space. Journal of Theoretical Biology, 225(4):443–451, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002777.

Jahandideh:2008:EPC

[JBJ^{+}08] Mina Jahandideh, Seyyed Mohsen Hosseini Barkooie, Samad Jahandideh, Parviz Abdolmaleki, Mohammad Mehdi Movahedi, Somayyeh Hoseini, Ebrahim Barzegari Asadabadi, Fatemeh Javani Jouni, Zahra Karami, and Nader Hodjati Firoozabadi. Elucidating the protein cold-adaptation: Investigation of the parameters enhancing protein psychrophilicity. Journal of Theoretical Biology, 255(1):113–118, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308004001.

James:2004:GMS

[JBK04] R. James, P. G. Bennett, and J. Krause. Geometry for mutualistic and selfish herds: the limited domain of danger. Journal of Theoretical Biology, 228(1):107–113, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004600.

Jones:2002:FCR

[JBP02] M. E. Jones, M. N. Berry, and J. W. Phillips. Futile cycles revisited: a Markov chain model of simultaneous glycolysis and gluconeogenesis. Journal of Theoretical Biology, 217(4):509–523, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930420.

Jo:2009:BEI

[JCK09] Junghyo Jo, Moo Young Choi, and Duk-Su Koh. Beneficial effects of intercellular interactions between pancreatic islet cells in blood glucose regulation. Journal of Theoretical Biology, 257(2):312–319, March 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006346.

Jolly:2007:MBM

[JCRJ07a] C. Jolly, A. J. L. Cook, J. Raferty, and M. E. Jones. Measuring bidirectional mutation. Journal of Theoretical Biology, 246(2):269–277, May 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519306005996. See corrigendum [JCRJ07b].

Jolly:2007:CMB

[JCRJ07b] C. Jolly, A. J. L. Cook, J. Raftery, and M. E. Jones. Corrigendum to “Measuring bidirectional mutation”: [J. Theor. Biol. 246 (2006) 269–277]. Journal of Theoretical Biology, 247(4):862, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700269X. See [JCRJ07a].

Joffe:2003:NMQ

[JCW^{+}03] Tracey H. Joffe, Timothy J. Cole, Jonathan C. K. Wells, Michael P. Geary, Charles H. Rodeck, and Peter C. Hindmarsh. A new method for quantifying encephalization in the growing individual. Journal of Theoretical Biology, 225(3):361–367, December 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002637.

Ju:2009:REF

[JD09] Shu Ju and Donald L. DeAngelis. The R^{∗} rule and energy flux in a plant-nutrient ecosystem. Journal of Theoretical Biology, 256(3):326–332, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005171.

Jafarnia-Dabanloo:2007:MZM

[JDMZ^{+}07] N. Jafarnia-Dabanloo, D. C. McLernon, H. Zhang, A. Ayatollahi, and V. Johari-Majd. A modified Zeeman model for producing HRV signals and its application to ECG signal generation. Journal of Theoretical Biology, 244(2):180–189, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003481.

Jonsson:2001:CLH

[JE01] Annie Jonsson and Bo Ebenman. Are certain life histories particularly prone to local extinction? Journal of Theoretical Biology, 209(4):455–463, April 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922805.

Janko:2009:SDP

[JE09] K. Janko and J. Eisner. Sperm-dependent parthenogens delay the spatial expansion of their sexual hosts. Journal of Theoretical Biology, 261(3):431–440, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003762.

Jesse:2008:FCM

[JEDH08] M. Jesse, P. Ezanno, S. Davis, and J. A. P. Heesterbeek. A fully coupled, mechanistic model for infectious disease dynamics in a metapopulation: Movement and epidemic duration. Journal of Theoretical Biology, 254(2):331–338, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002968.

Joyce:2006:MOD

[JEHK06] Andrew R. Joyce, Keith Easterling, Stephen G. Holtzman, and Michael J. Kuhar. Modeling the onset of drug dependence: a consideration of the requirement for protein synthesis. Journal of Theoretical Biology, 240(4):531–537, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004546.

Jeschke:2006:DDE

[Jes06] Jonathan M. Jeschke. Density-dependent effects of prey defenses and predator offenses. Journal of Theoretical Biology, 242(4):900–907, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002153.

Jaeger:2001:COM

[JG01] Johannes Jaeger and Brian C. Goodwin. A cellular oscillator model for periodic pattern formation. Journal of Theoretical Biology, 213(2):171–181, November 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924142.

Janssen:2006:DPC

[JG06] Marco A. Janssen and Robert L. Goldstone. Dynamicpersistence of cooperation in public good games when group size is dynamic. Journal of Theoretical Biology, 243(1):134–142, November 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002578.

Janda:2008:DMP

[JG08] Jozef Janda and Gernot Geginat. A deterministic model for the processing and presentation of bacteria-derived antigenic peptides. Journal of Theoretical Biology, 250(3):532–546, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005218.

Jahandideh:2009:TMA

[JHJ09a] Samad Jahandideh, Somayyeh Hoseini, and Mina Jahandideh. Toward more accuracy in protein structural bioinformatics: Have novel hybrid modeling procedures been born? Journal of Theoretical Biology, 256(1):147, January 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004955.

Jahandideh:2009:TTP

[JHJ^{+}09b] Samad Jahandideh, Somayyeh Hoseini, Mina Jahandideh, Afsaneh Hoseini, and Fatemeh Miri Disfani. γ-turn types prediction in proteins using the two-stage hybrid neural discriminant model. Journal of Theoretical Biology, 259(3):517–522, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900188X.

Joron:2005:EMM

[JI05] Mathieu Joron and Yoh Iwasa. The evolution of a Müllerian mimic in a spatially distributed community. Journal of Theoretical Biology, 237(1):87–103, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001608.

Jung:2009:OCS

[JITJ09] Eunok Jung, Shingo Iwami, Yasuhiro Takeuchi, and Tae-Chang Jo. Optimal control strategy for prevention of avian influenza pandemic. Journal of Theoretical Biology, 260(2):220–229, September 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002550.

Jaggi:2001:ISV

[JJ01] Sidharth Jaggi and Amitabh Joshi. Incorporating spatial variation in density enhances the stability of simple population dynamics models. Journal of Theoretical Biology, 209(2):249–255, March 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922604.

Jeger:2009:GTM

[JJEX09] M. J. Jeger, P. Jeffries, Y. Elad, and X.-M. Xu. A generic theoretical model for biological control of foliar plant diseases. Journal of Theoretical Biology, 256(2):201–214, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005092.

Jagers:2003:RVC

[JK03] Peter Jagers and Fima Klebaner. Random variation and concentration effects in PCR. Journal of Theoretical Biology, 224(3):299–304, October 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001668.

Joseph:2004:MSH

[JK04] Ian M. Joseph and Denise Kirschner. A model for the study of helicobacter pylori interaction with human gastric acid secretion. Journal of Theoretical Biology, 228(1):55–80, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004570.

Jafri:2006:MMM

[JK06] M. Saleet Jafri and Malgorzata Kotulska. Modeling the mechanism of metabolic oscillations in ischemic cardiac myocytes. Journal of Theoretical Biology, 242(4):801–817, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001949.

Ji:2000:HTP

[JL00] Fengmin Ji and Liaofu Luo. A hypercycle theory of proliferation of viruses and resistance to the viruses of transgenic plant. Journal of Theoretical Biology, 204(3):453–465, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920253.

Judas:2001:CII

[JL01] L. Judas and M. Lokajícek. Cell inactivation by ionizing particles and the shapes of survival curves. Journal of Theoretical Biology, 210(1):15–21, May 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922830.

Jablonka:2006:EIM

[JL06] Eva Jablonka and Marion J. Lamb. The evolution of information in the major transitions. Journal of Theoretical Biology, 239(2):236–246, March 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003851.

Jimenez:2008:ERC

[JLCS08] Raúl Jiménez, Haydee Lugo, José A. Cuesta, and Angel Sánchez. Emergence and resilience of cooperation in the spatial prisoner’s dilemma via a reward mechanism. Journal of Theoretical Biology, 250(3):475–483, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004961.

Jones:2001:TMS

[JLS01] B. S. Jones, W. F. Lynn, and M. O. Stone. Thermal modeling of snake infrared reception: Evidence for limited detection range. Journal of Theoretical Biology, 209(2):201–211, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922562.

Jabbour:2007:HDT

[JM07] M. N. Jabbour and G. T. Matioli. Homotopic dynamics of telomeric loops in mammalian chromatids. Journal of Theoretical Biology, 249(3):500–502, December 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004158.

Johansson:2000:FEM

[JMB00] T. Johansson, P. Meier, and R. Blickhan. A finite-element model for the mechanical analysis of skeletal muscles. Journal of Theoretical Biology, 206(1):131–149, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092109X.

Jeger:2009:ETR

[JMvdB09] Michael J. Jeger, Laurence V. Madden, and Frank van den Bosch. The effect of transmission route on plant virus epidemic development and disease control. Journal of Theoretical Biology, 258(2):198–207, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000101.

Jiang-Ning:2004:COC

[JNWJWB04] Song Jiang-Ning, Li Wei-Jiang, and Xu Wen-Bo. Cooperativity of the oxidization of cysteines in globular proteins. Journal of Theoretical Biology, 231(1):85–95, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002747.

Johnson:2008:MBF

[Joh08] Leah R. Johnson. Microcolony and biofilm formation as a survival strategy for bacteria. Journal of Theoretical Biology, 251(1):24–34, March 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005188.

Johnson:2000:RSP

[JP00] Norman A. Johnson and Adam H. Porter. Rapid speciation via parallel, directional selection on regulatory genetic pathways. Journal of Theoretical Biology, 205(4):527–542, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920708.

Jeffries:2006:SDG

[JPJ06] Clark Jeffries, Diana O. Perkins, and Michael Jarstfer. Systematic discovery of the grammar of translational inhibition by RNA hairpins. Journal of Theoretical Biology, 241(2):205– 215, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005102.

Jin:2008:NNS

[JQR08] Emma Y. Jin, Jing Qin, and Christian M. Reidys. Neutral networks of sequence to shape maps. Journal of Theoretical Biology, 250(3):484–497, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004286.

Johnson:2006:MES

[JR06] Katie Johnson and Louis F. Rossi. A mathematical and experimental study of ant foraging trail dynamics. Journal of Theoretical Biology, 241(2):360–369, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005278.

Jin:2008:CLL

[JR08] Emma Y. Jin and Christian M. Reidys. Central and local limit theorems for RNA structures. Journal of Theoretical Biology, 250(3):547–559, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004493.

Jestin:2007:SBS

[JS07] Jean-Luc Jestin and Christophe Soulé. Symmetries by base substitutions in the genetic code predict 2’ or 3’ aminoacylation of tRNAs. Journal of Theoretical Biology, 247(2):391–394, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001130.

Jahandideh:2007:TTP

[JSA^{+}07] Samad Jahandideh, Amir Sabet Sarvestani, Parviz Abdolmaleki, Mina Jahandideh, and Mahdyar Barfeie. γ-turn types prediction in proteins using the support vector machines. Journal of Theoretical Biology, 249(4):785–790, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004201.

John:2004:CET

[JSCN04] Alexander John, Andreas Schadschneider, Debashish Chowdhury, and Katsuhiro Nishinari. Collective effects in traffic on bi-directional ant trails. Journal of Theoretical Biology, 231(2):279–285, November 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003042.

Jonker:2002:PIC

[JST^{+}02] Catholijn M. Jonker, Jacky L. Snoep, Jan Treur, Hans V. Westerhoff, and Wouter C. A. Wijngaards. Putting intentions into cell biochemistry: An artificial intelligence perspective. Journal of Theoretical Biology, 214(1):105–134, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924440.

Jonker:2008:BMC

[JST^{+}08] C. M. Jonker, J. L. Snoep, J. Treur, H. V. Westerhoff, and W. C. A. Wijngaards. BDI-modelling of complex intracellular dynamics. Journal of Theoretical Biology, 251(1):1–23, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005127.

Jordan:2002:SPE

[JSV02] Ferenc Jordán, István Scheuring, and Gábor Vida. Species positions and extinction dynamics in simple food Webs. Journal of Theoretical Biology, 215(4):441–448, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925238.

Ji:2006:CMO

[JTGP06] James W. Ji, Nikolaos M. Tsoukias, Daniel Goldman, and Aleksander S. Popel. A computational model of oxygen transport in skeletal muscle for sprouting and splitting modes of angiogenesis. Journal of Theoretical Biology, 241(1):94–108, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005059.

Juska:2008:MMM

[Jus08] Alfonsas Juska. Minimal models of multi-site ligand-binding kinetics. Journal of Theoretical Biology, 255(4):396–403, December 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004499.

Jazwinski:2001:PRC

[JW01] S. Michał Jazwinski and Jarosław Wawryn. Profiles of random change during aging contain hidden information about longevity and the aging process. Journal of Theoretical Biology, 213(4):599–608, December 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924348.

Jones:2007:IPV

[JWB07] E. O. Jones, A. White, and M. Boots. Interference and the persistence of vertically transmitted parasites. Journal of Theoretical Biology, 246(1):10–17, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005686.

Joshi:2009:ICV

[JWB^{+}09] Badal Joshi, Xueying Wang, Sayanti Banerjee, Haiyan Tian, Anastasios Matzavinos, and Mark A. J. Chaplain. On immunotherapies and cancer vaccination protocols: a mathematical modelling approach. Journal of Theoretical Biology, 259(4):820–827, August 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002161.

Jeger:2008:ACR

[JWH08] Michael J. Jeger, Pieter J. Wijngaarden, and Rolf F. Hoekstra. Adaptation to the cost of resistance in a haploid clonally reproducing organism: The role of mutation, migration and selection. Journal of Theoretical Biology, 252(4):621–632, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800088X.

Ji:2008:IPL

[JWWS08] Xiaofeng Ji, Yuejun Wang, Haiying Wang, and Mi Sun. Identification of protein latent periodicities using recurrent correlation analysis. Journal of Theoretical Biology, 255(3):316–319, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004451.

Ji:2009:EDF

[JZST09] Ting Ji, Boyu Zhang, Yuehua Sun, and Yi Tao. Evolutionary dynamics of fearfulness and boldness. Journal of Theoretical Biology, 256(4):637–643, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005201.

Koroleva:2002:FCR

[KA02] S. V. Koroleva and I. P. Ashmarin. Functional continuum of regulatory peptides (RPs): Vector model of RP-effects representation. Journal of Theoretical Biology, 216(3):257–271, June 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925555.

Kirner:2003:PSC

[KA03a] T. Kirner and J. Ackermann. Pattern selection in a cooperative biochemical in vitro amplification system: the role of parasites. Journal of Theoretical Biology, 224(4):539–549, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002029.

Kunz:2003:SCR

[KA03b] Hanspeter Kunz and Peter Achermann. Simulation of circadian rhythm generation in the suprachiasmatic nucleus with locally coupled self-sustained oscillators. Journal of Theoretical Biology, 224(1):63–78, September 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001413.

Kaitaniemi:2004:TAS

[Kai04] Pekka Kaitaniemi. Testing the allometric scaling laws. Journal of Theoretical Biology, 228(2):149–153, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004624.

Kamimura:2008:CPG

[KAI08] Yoshitaka Kamimura, Jun Abe, and Hiroshi Ito. The continuous public goods game and the evolution of cooperative sex ratios. Journal of Theoretical Biology, 252(2):277–287, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000805.

Kalinin:2000:STH

[Kal00] Vladimir I. Kalinin. System-theoretical (holistic) approach to the modelling of structural-functional relationships of biomolecules and their evolution: an example of triterpene glycosides from sea cucumbers (Echinodermata, Holothurioidea). Journal of Theoretical Biology, 206(1):151–168, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921106.

Kalapos:2007:ERC

[Kal07] Miklós Péter Kalapos. The energetics of the reductive citric acid cycle in the pyrite-pulled surface metabolism in the early stage of evolution. Journal of Theoretical Biology, 248(2):251–258, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002202.

Kamb:2003:MLF

[Kam03] Alexander Kamb. Mutation load, functional overlap, and synthetic lethality in the evolution and treatment of cancer. Journal of Theoretical Biology, 223(2):205–213, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000870.

Kandjov:2001:HWR

[Kan01] Iwajlo M. Kandjov. Heat and water rate transfer processes in the human respiratory tract at various altitudes. Journal of Theoretical Biology, 208(3):287–293, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922173.

Kanazawa:2005:BTP

[Kan05] Satoshi Kanazawa. Big and tall parents have more sons: Further generalizations of the Trivers–Willard hypothesis. Journal of Theoretical Biology, 235(4):583–590, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000706.

Kanazawa:2006:VMM

[Kan06] Satoshi Kanazawa. Violent men have more sons: Further evidence for the generalized Trivers–Willard hypothesis (gTWH). Journal of Theoretical Biology, 239(4):450–459, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003498.

Kanazawa:2007:BPM

[Kan07] Satoshi Kanazawa. Beautiful parents have more daughters: a further implication of the generalized Trivers–Willard hypothesis (gTWH). Journal of Theoretical Biology, 244(1):133–140, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003079.

Kao:2006:EPT

[Kao06] Rowland R. Kao. Evolution of pathogens towards low R_{0} in heterogeneous populations. Journal of Theoretical Biology, 242(3):634–642, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600138X.

Kargacin:2003:RCB

[Kar03] Gary J. Kargacin. Responses of Ca^{2+}-binding proteins to localized, transient changes in intracellular [Ca^{2+}]. Journal of Theoretical Biology, 221(2):245–258, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931870.

Katz:2003:SAT

[Kat03] Y. Katz. Solubility, adsorption, and the thermodynamics of the cutoff effect. Journal of Theoretical Biology, 225(3):341–349, December 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002595.

Kauffman:2004:PUE

[Kau04] Stuart Kauffman. A proposal for using the ensemble approach to understand genetic regulatory networks. Journal of Theoretical Biology, 230(4):581–590, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002450.

Kawata:2001:INS

[Kaw01] Masakado Kawata. The influence of neighborhood size and habitat shape on the accumulation of deleterious mutations. Journal of Theoretical Biology, 211(3):187–199, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301922684.

Karsai:2002:OWN

[KB02] István Karsai and Gábor Balázsi. Organization of work via a natural substance: Regulation of nest construction in social wasps. Journal of Theoretical Biology, 218(4):549–565, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930997.

Kamo:2004:CPS

[KB04] Masashi Kamo and Michael Boots. The curse of the pharaoh in space: free-living infectious stages and the evolution of virulence in spatially explicit populations. Journal of Theoretical Biology, 231(3):435–441, December 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003261.

Killingback:2006:SFE

[KBD06] Timothy Killingback, Hendrik J. Blok, and Michael Doebeli. Scale-free extinction dynamics in spatially structured host-parasitoid systems. Journal of Theoretical Biology, 241(4):745–750, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000129.

Kapela:2008:MMC

[KBT08] Adam Kapela, Anastasios Bezerianos, and Nikolaos M. Tsoukias. A mathematical model of Ca^{2+} dynamics in rat mesenteric smooth muscle cell: Agonist and NO stimulation. Journal of Theoretical Biology, 253(2):238–260, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001173.

Keener:2003:ESS

[KC03] James P. Keener and Eric Cytrynbaum. The effect of spatial scale of resistive inhomogeneity on defibrillation of cardiac tissue. Journal of Theoretical Biology, 223(2):233–248, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000894.

Korniss:2005:SDI

[KC05] Gyorgy Korniss and Thomas Caraco. Spatial dynamics of invasion: the geometry of introduced species. Journal of Theoretical Biology, 233(1):137–150, March 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004709.

Konuma:2007:ECD

[KC07] Junji Konuma and Satoshi Chiba. Ecological character displacement caused by reproductive interference. Journal of Theoretical Biology, 247(2):354–364, July 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700118X.

Kobayashi:2003:MGS

[KCA03] Tetsuya Kobayashi, Luonan Chen, and Kazuyuki Aihara. Modeling genetic switches with positive feedback loops. Journal of Theoretical Biology, 221(3):379–399, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931900.

Kuo:2008:UCU

[KCHP08] Chih-Horng Kuo, Vanessa Corby-Harris, and Daniel E. L. Promislow. The unavoidable costs and unexpected benefits of parasitism: Population and metapopulation models of parasite-mediated competition. Journal of Theoretical Biology, 250(2):244–256, January 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004900.

Kafsack:2007:KMT

[KCP07] Björn F. C. Kafsack, Vern B. Carruthers, and Fernando J. Pineda. Kinetic modeling of toxoplasma gondii invasion. Journal of Theoretical Biology, 249(4):817–825, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004183.

Kafsack:2009:EKM

[KCP09] Björn F. C. Kafsack, Vern B. Carruthers, and Fernando J. Pineda. Erratum to “Kinetic modeling of Toxoplasma gondii invasion” [J. Theor. Biol. 249 (2007) 817–825]. Journal of Theoretical Biology, 260(4):598, October 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002586.

Karbowski:2006:CRT

[KCS^{+}06] Jan Karbowski, Christopher J. Cronin, Adeline Seah, Jane E. Mendel, Daniel Cleary, and Paul W. Sternberg. Conservation rules, their breakdown, and optimality in caenorhabditis sinusoidal locomotion. Journal of Theoretical Biology, 242(3):652–669, October 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001627.

Kier:2002:CAM

[KCT02] Lemont B. Kier, Chao-Kun Cheng, and Bernard Testa. A cellular automata model of ligand passage over a protein hydrodynamic landscape. Journal of Theoretical Biology, 215(4):415–426, April 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301925251.

Keenan:2005:CMT

[KCV05] Daniel M. Keenan, Somesh Chattopadhyay, and Johannes D. Veldhuis. Composite model of time-varying appearance and disappearance of neurohormone pulse signals in blood. Journal of Theoretical Biology, 236(3):242–255, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001177.

Kumar:2004:DAI

[KCVC04] Rukmini Kumar, Gilles Clermont, Yoram Vodovotz, and Carson C. Chow. The dynamics of acute inflammation. Journal of Theoretical Biology, 230(2):145–155, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001845.

Kesmir:2003:SMG

[KD03] Can Kesmir and Rob J. De Boer. A spatial model of germinal center reactions: cellular adhesion based sorting of B cells results in efficient affinity maturation. Journal of Theoretical Biology, 222(1):9–22, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000109.

Kupczok:2006:DSR

[KD06]  Anne Kupczok and Peter Dittrich.  Determinants of

simulated RNA evolution. Journal of Theoretical Biology, 238(3):726–735, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002778.

Kooijman:2002:LIM

[KDK02] S. A. L. M. Kooijman, H. A. Dijkstra, and B. W. Kooi. Light-induced mass turnover in a mono-species community of mixotrophs. Journal of Theoretical Biology, 214(2):233–254, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924580.

Kawaguchi:2006:CEM

[KDKS06] Isao Kawaguchi, Masahiro Doi, Shizuko Kakinuma, and Yoshiya Shimada. Combined effect of multiple carcinogens and synergy index. Journal of Theoretical Biology, 243(1):143–151, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002232.

Kerkhoff:2009:MNW

[KE09] Andrew J. Kerkhoff and Brian J. Enquist. Multiplicative by nature: Why logarithmic transformation is necessary in allometry. Journal of Theoretical Biology, 257(3):519–521, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006681.

Keeling:2000:MMM

[Kee00] Matt J. Keeling. Multiplicative moments and measures of persistence in ecology. Journal of Theoretical Biology, 205(2):269– 281, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920666.

Keener:2004:TD

[Kee04] James P. Keener. The topology of defibrillation. Journal of Theoretical Biology, 230(4):459–473, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002371.

Keener:2005:MLC

[Kee05] J. P. Keener. A model for length control of flagellar hooks of salmonella typhimurium. Journal of Theoretical Biology, 234(2):263–275, May 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005764.

Kelly:2001:TKW

[Kel01] Robert E. Kelly. Tripedal knuckle-walking: a proposal for the evolution of human locomotion and handedness. Journal of Theoretical Biology, 213(3):333–358, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192421X.

Kelly:2007:MSB

[Kel07] John K. Kelly. Mutation-selection balance in mixed mating populations. Journal of Theoretical Biology, 246(2):355–365, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306006023.

Kendal:2001:SDH

[Ken01] Wayne S. Kendal. The size distribution of human hematogenous metastases. Journal of Theoretical Biology, 211(1):29– 38, July 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923264.

Kendal:2002:FDN

[Ken02] Wayne S. Kendal. A frequency distribution for the number of hematogenous organ metastases. Journal of Theoretical Biology, 217(2):203–218, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930213.

Kendal:2007:SIC

[Ken07] Wayne S. Kendal. Scale invariant correlations between genes and SNPs on human chromosome 1 reveal potential evolutionary mechanisms. Journal of Theoretical Biology, 245(2):329–340, March 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004851.

Kerszberg:2000:SSR

[Ker00] Michel Kerszberg. The survival of slow reproducers. Journal of Theoretical Biology, 206(1):81–89, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921015.

Keranen:2004:SSE

[Ker04] Soile V. E. Keränen. Simulation study on effects of signaling network structure on the developmental increase in complexity. Journal of Theoretical Biology, 231(1):3–21, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001249. See erra-

tum [Ker06].

Keranen:2006:ESS

[Ker06] Soile V. E. Keränen. Erratum to: “Simulation study on effects of signaling network structure on the developmental increase in complexity”: [J. Theor. Biol. 231 (2004) 3– 21]. Journal of Theoretical Biology, 242(1):263, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001780. See [Ker04].

Kerr:2003:CCN

[KF03] Benjamin Kerr and Marcus W. Feldman. Carving the cognitive niche: Optimal learning strategies in homogeneous and heterogeneous environments. Journal of Theoretical Biology, 220(2):169–188, January 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931468.

Karpinets:2004:MDT

[KF04] Tatiana V. Karpinets and Brent D. Foy. Model of the developing tumorigenic phenotype in mammalian cells and the roles of sustained stress and replicative senescence. Journal of Theoretical Biology, 227(2):253–264, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004223.

Kaneko:2006:ERB

[KF06] Kunihiko Kaneko and Chikara Furusawa. An evolutionary relationship between genetic variation and phenotypic fluctuation. Journal of Theoretical Biology, 240(1):78–86, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003681.

Kearns:2002:QCF

[KFG^{+}02] J. Kearns, L. Farnell, W. G. Gibson, Y. Q. Lin, and M. R. Bennett. Quantal current fields around individual boutons in sympathetic ganglia. Journal of Theoretical Biology, 214(2):135–146, January 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924208.

Keeling:2000:IBP

[KG00] Matt J. Keeling and Bryan T. Grenfell. Individual-based perspectives on R_{0}. Journal of Theoretical Biology, 203(1):51–61, March 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910640.

Kotha:2002:MTM

[KG02] S. P. Kotha and N. Guzelsu. Modeling the tensile mechanical behavior of bone along the longitudinal direction. Journal of Theoretical Biology, 219(2):269–279, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302931206.

Khan:2003:SPP

[KG03] Q. J. A. Khan and A. F. Ghaleb. A study of preypredator relations for mammals. Journal of Theoretical Biology, 223(2):171–178, July 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000845.

Krone:2006:SSO

[KG06a] Stephen M. Krone and Yongtao Guan. Spatial selforganization in a cyclic resource-species model. Journal of Theoretical Biology, 241(1):14–25, July 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004807.

Kurosawa:2006:ACO

[KG06b] Gen Kurosawa and Albert Goldbeter. Amplitude of circadian oscillations entrained by 24-h light-dark cycles. Journal of Theoretical Biology, 242(2):478–488, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001214.

Kosiol:2004:NCM

[KGB04] Carolin Kosiol, Nick Goldman, and Nigel H. Buttimore. A new criterion and method for amino acid classification. Journal of Theoretical Biology, 228(1):97–106, May 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004594.

Kartascheff:2009:PCS

[KGD09] Boris Kartascheff, Christian Guill, and Barbara Drossel. Positive complexity-stability relations in food web models without foraging adaptation. Journal of Theoretical Biology, 259(1):12–23, July 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001210.

Kuske:2007:SOC

[KGG07] Rachel Kuske, Luis F. Gordillo, and Priscilla Greenwood. Sustained oscillations via coherence resonance in SIR. Journal of Theoretical Biology, 245(3):459–469, April 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004930.

Kroeger:2008:MCD

[KGG08] Jens H. Kroeger, Anja Geitmann, and Martin Grant. Model for calcium dependent oscillatory growth in pollen tubes. Journal of Theoretical Biology, 253(2):363–374, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001100.

Kendal:2009:ESL

[KGL09] Jeremy Kendal, Luc-Alain Giraldeau, and Kevin Laland. The evolution of social learning rules: Payoff-biased and frequency-dependent biased transmission. Journal of Theoretical Biology, 260(2):210–219, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002288.

Kroon:2007:MSC

[KH07a] Martin Kroon and Gerhard A. Holzapfel. A model for saccular cerebral aneurysm growth by collagen fibre remodelling. Journal of Theoretical Biology, 247(4):775–787, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001142.

Kwan:2007:OFS

[KH07b] Maxine Kwan and Mont Hubbard. Optimal foot shape for a passive dynamic biped. Journal of Theoretical Biology, 248(2):331–339, September 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002317.

Kuttler:2008:ITQ

[KH08] Christina Kuttler and Burkhard A. Hense. Interplay of two quorum sensing regulation systems of vibrio fischeri. Journal of Theoretical Biology, 251(1):167–180, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005747.

Kaunas:2009:KMS

[KH09a] Roland Kaunas and Hui-Ju Hsu. A kinematic model of stretch-induced stress fiber turnover and reorientation. Journal of Theoretical Biology, 257(2):320–330, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006115.

Kroon:2009:TMF

[KH09b] Martin Kroon and Gerhard A. Holzapfel. A theoretical model for fibroblast-controlled growth of saccular cerebral aneurysms. Journal of Theoretical Biology, 257(1):73–83, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005602.

Kukillaya:2009:MIL

[KH09c] Raghavendra P. Kukillaya and Philip Holmes. A model for insect locomotion in the horizontal plane: Feedforward activation of fast muscles, stability, and robustness. Journal of Theoretical Biology, 261(2):210–226, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003464.

Kocsis:2006:MME

[KHE06] Laszlo Kocsis, Peter Herman, and Andras Eke. Mathematical model for the estimation of hemodynamic and oxygenation variables by tissue spectroscopy. Journal of Theoretical Biology, 241(2):262–275, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005199.

Koganezawa:2009:MES

[KHHS09] Masayuki Koganezawa, Hiroaki Hara, Yoshinori Hayakawa, and Ichiro Shimada. Memory effects on scale-free dynamics in foraging drosophila. Journal of Theoretical Biology, 260(3):353–358, October 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002847.

Kimbrell:2007:IVI

[KHL07] Tristan Kimbrell, Robert D. Holt, and Per Lundberg. The influence of vigilance on intraguild predation. Journal of Theoretical Biology, 249(2):218–234, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003694.

Khrennikov:2004:PPR

[Khr04] Andrei Khrennikov. Probabilistic pathway representation of cognitive information. Journal of Theoretical Biology, 231(4):597–613, December 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003443.

Krishnan:2004:MFD

[KI04] J. Krishnan and P. A. Iglesias. A modeling framework describing the enzyme regulation of membrane lipids underlying gradient perception in dictyostelium cells. Journal of Theoretical Biology, 229(1):85–99, July 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001067.

Krishnan:2005:MFD

[KI05a] J. Krishnan and P. A. Iglesias. A modelling framework describing the enzyme regulation of membrane lipids underlying gradient perception in dictyostelium cells II: Input-output analysis. Journal of Theoretical Biology, 235(4):504–520, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000640.

Kurosawa:2005:TCC

[KI05b] Gen Kurosawa and Yoh Iwasa. Temperature compensation in circadian clock models. Journal of Theoretical Biology, 233(4):453–468, April 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004837.

Kimura:2009:RDM

[KI09] Mariko Kimura and Yasuo Ihara. Replicator-dynamics models of sexual conflict. Journal of Theoretical Biology, 260(1):90–97, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002628.

Komoto:2003:GDB

[KiHM^{+}03] Atsushi Komoto, Ken ichi Hanaki, Shinya Maenosono, Joe Yuichiro Wakano, Yukio Yamaguchi, and Kenji Yamamoto. Growth dynamics of bacillus circulans colony. Journal of Theoretical Biology, 225(1):91–97, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002248.

Kimmel:2007:OAP

[Kim07] Marek Kimmel. Ovide arino, a personal memory. Journal of Theoretical Biology, 244(3):365–366, February 7, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004097.

King:2004:ADD

[Kin04] Michael R. King. Apparent 2-D diffusivity in a ruffled cell membrane. Journal of Theoretical Biology, 227(3):323–326, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004284.

King:2007:DBC

[Kin07] Michael R. King. Do blood capillaries exhibit optimal bumpiness? Journal of Theoretical Biology, 249(1):178–180, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003372.

Krakauer:2002:RQD

[KJ02] David C. Krakauer and Vincent A. A. Jansen. Red Queen dynamics of protein translation. Journal of Theoretical Biology, 218(1):97–109, September 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930547.

Komarova:2008:PHC

[KJ08] Natalia L. Komarova and Kimberly A. Jameson. Population heterogeneity and color stimulus heterogeneity in agent-based color categorization. Journal of Theoretical Biology, 253(4):680–700, August 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001549.

Kent:2005:MMC

[KJD05] Adam Kent, Susanne Plesner Jensen, and C. Patrick Doncaster. Model of microtine cycles caused by lethal toxins in non-preferred food plants. Journal of Theoretical Biology, 234(4):593–604, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304006137.

Kong:2008:NMM

[KJD08] Dong Kong, Baohua Ji, and Lanhong Dai. Nonlinear mechanical modeling of cell adhesion. Journal of Theoretical Biology, 250(1):75–84, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004560.

Karlsson:2007:FWS

[KJJ07] Patrik Karlsson, Tomas Jonsson, and Annie Jonsson. Food web structure and interaction strength pave the way for vulnerability to extinction. Journal of Theoretical Biology, 249(1):77–92, November 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003396.

Kirkby:2007:MMT

[KJJB07] Norman F. Kirkby, Sarah J. Jefferies, Raj Jena, and Neil G. Burnet. A mathematical model of the treatment and survival of patients with high-grade brain tumours. Journal of Theoretical Biology, 245(1):112–124, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004085.

Kobayashi:2000:STE

[KK00a] Yutaka Kobayashi and Kihachiro Kikuzawa. A single theory explains two empirical laws applicable to plant populations. Journal of Theoretical Biology, 205(2):253–260, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920587.

Kowald:2000:ADM

[KK00b] Axel Kowald and Thomas B. L. Kirkwood. Accumulation of defective mitochondria through delayed degradation of damaged organelles and its possible role in the ageing of post-mitotic and dividing cells. Journal of Theoretical Biology, 202(2):145–160, January 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910469.

Kraines:2000:NSM

[KK00c] David P. Kraines and Vivian Y. Kraines. Natural selection of memory-one strategies for the iterated Prisoner’s dilemma. Journal of Theoretical Biology, 203(4):335–355, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910890.

Kondrashov:2006:RSF

[KK06] Fyodor A. Kondrashov and Alexey S. Kondrashov. Role of selection in fixation of gene duplications. Journal of Theoretical Biology, 239(2):141–151, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003759.

Khrennikov:2009:ACP

[KK09] A. Yu. Khrennikov and S. V. Kozyrev. 2-adic clustering of the PAM matrix. Journal of Theoretical Biology, 261(3):396–406, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003713.

Kurbel:2001:MGG

[KKDV01] Sven Kurbel, Beatrica Kurbel, Branko Dmitrović, and Aleksandar Vcev. A model of the gastric gland ejection cycle: Low ejection fractions require reduction of the glandular dead space. Journal of Theoretical Biology, 210(3):337–343, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923136.

Kelpin:2000:NMP

[KKK00] F. D. L. Kelpin, M. A. Kirkilionis, and B. W. Kooi. Numerical methods and parameter estimation of a structured population model with discrete events in the life history. Journal of Theoretical Biology, 207(2):217–230, November 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921684.

Kim:2009:ECD

[KKPB09] Won-Kap Kim, Jin Hwan Ko, Hoon Cheol Park, and Doyoung Byun. Effects of corrugation of the dragonfly wing on gliding performance. Journal of Theoretical Biology, 260(4):523–530, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003282.

Kesseler:2007:MMH

[KKR^{+}07] Kevin J. Kesseler, William K. Kaufmann, Joyce T. Reardon, Timothy C. Elston, and Aziz Sancar. A mathematical model for human nucleotide excision repair: Damage recognition by random order assembly and kinetic proofreading. Journal of Theoretical Biology, 249(2):361–375, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003529.

Kazmierczak:2009:RKA

[KL09] Bogdan Kazmierczak and Tomasz Lipniacki. Regulation of kinase activity by diffusion and feedback. Journal of Theoretical Biology, 259(2):291–296, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001131.

Klinke:2006:RPM

[Kli06] David J. Klinke. The ratio of p40 monomer to dimer is an important determinant of IL-12 bioactivity. Journal of Theoretical Biology, 240(2):323–335, May 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004261.

Kowald:2006:APM

[KLK06] Axel Kowald, Hans Lehrach, and Edda Klipp. Alternative pathways as mechanism for the negative effects associated with overexpression of superoxide dismutase. Journal of Theoretical Biology, 238(4):828–840, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002924.

Kim:2007:MRM

[KLL07a] Peter S. Kim, Peter P. Lee, and Doron Levy. Modeling regulation mechanisms in the immune system. Journal of Theoretical Biology, 246(1):33–69, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005625.

Klausmeier:2007:MFU

[KLL07b] C. A. Klausmeier, E. Litchman, and S. A. Levin. A model of flexible uptake of two essential resources. Journal of Theoretical Biology, 246(2):278–289, May 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000033.

Kulkarni:2008:ACL

[KLL08] Kedar Kulkarni, Peter Larsen, and Andreas A. Linninger. Assessing chronic liver toxicity based on relative gene expression data. Journal of Theoretical Biology, 254(2):308–318, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002725.

Kim:2009:MMP

[KLN^{+}09] Yangjin Kim, Sean Lawler, Michał O. Nowicki, E. Antonio Chiocca, and Avner Friedman. A mathematical model for pattern formation of glioma cells outside the tumor spheroid core. Journal of Theoretical Biology, 260(3):359–371, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900304X.

Komarova:2005:REP

[KM05] Natalia L. Komarova and Vladimir Mironov. On the role of endothelial progenitor cells in tumor neovascularization. Journal of Theoretical Biology, 235(3):338–349, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305000408.

Kell:2008:MMR

[KM08] Douglas B. Kell and Pedro Mendes. The markup is the model: Reasoning about systems biology models in the semantic Web era. Journal of Theoretical Biology, 252(3):538– 543, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930700519X.

Knibbe:2007:ECB

[KMC^{+}07] Carole Knibbe, Olivier Mazet, Fabien Chaudier, Jean-Michel Fayard, and Guillaume Beslon. Evolutionary coupling between the deleteriousness of gene mutations and the amount of non-coding sequences. Journal of Theoretical Biology, 244(4):621–630, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004061.

Krogh-Madsen:2004:ADP

[KMGDG04] T. Krogh-Madsen, L. Glass, E. J. Doedel, and M. R. Guevara. Apparent discontinuities in the phase-resetting response of cardiac pacemakers. Journal of Theoretical Biology, 230(4):499–519, October 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002401.

Kranenbarg:2000:PCB

[KMGV00] S. Kranenbarg, M. Muller, J. L. W. Gielen, and J. H. G. Verhagen. Physical constraints on body size in teleost embryos. Journal of Theoretical Biology, 204(1):113–133, May 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910932.

Kaern:2000:SSD

[KMH00] Mads Kærn, Michael Menzinger, and Axel Hunding. Segmentation and somitogenesis derived from phase dynamics in growing oscillatory media. Journal of Theoretical Biology, 207(4):473–493, December 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921830.

Kurosawa:2002:CSC

[KMI02] Gen Kurosawa, Atsushi Mochizuki, and Yoh Iwasa. Comparative study of circadian clock models, in search of processes promoting oscillation. Journal of Theoretical Biology, 216(2):193– 208, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925464.

Krishnamachari:2004:SDB

[KmMK04] A. Krishnamachari, Vijnan moy Mandal, and Karmeshu. Study of DNA binding sites using the Rényi parametric entropy measure. Journal of Theoretical Biology, 227(3):429– 436, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004405.

Kim:2003:BPU

[KMP03] Jan T. Kim, Thomas Martinetz, and Daniel Polani. Bioinformatic principles underlying the information content of transcription factor binding sites. Journal of Theoretical Biology, 220(4):529–544, February 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931535.

Keef:2006:MEA

[KMT06] T. Keef, C. Micheletti, and R. Twarock. Master equation approach to the assembly of viral capsids. Journal of Theoretical Biology, 242(3):713–721, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001731.

Korobeinikov:2003:EEV

[KMW03] A. Korobeinikov, P. K. Maini, and W. J. Walker. Estimation of effective vaccination rate: pertussis in new Zealand as a case study. Journal of Theoretical Biology, 224(2):269–275, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001632.

Komarova:2003:LDF

[KN03] Natalia L. Komarova and Martin A. Nowak. Language dynamics in finite populations. Journal of Theoretical Biology, 221(3):445–457, April 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931997.

Kucken:2005:FF

[KN05] Michael Kücken and Alan C. Newell. Fingerprint formation. Journal of Theoretical Biology, 235(1):71–83, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304006198.

Katsnelson:2004:IVM

[KNC^{+}04] Leonid B. Katsnelson, Larissa V. Nikitina, Denis Chemla, Olga Solovyova, Catherine Coirault, Yves Lecarpentier, and Vladimir S. Markhasin. Influence of viscosity on myocardium mechanical activity: a mathematical model. Journal of Theoretical Biology, 230(3):385–405, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002656.

Komarova:2001:EDG

[KNN01] Natalia L. Komarova, Partha Niyogi, and Martin A. Nowak. The evolutionary dynamics of grammar acquisition. Journal of Theoretical Biology, 209(1):43–59, March 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922409.

Karolyi:2005:RSP

[KNS05] György Károlyi, Zoltán Neufeld, and István Scheuring. Rockscissors-paper game in a chaotic flow: The effect of dispersion on the cyclic competition of microorganisms. Journal of Theoretical Biology, 236(1):12–20, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500072X.

Katsuyama:2009:CCB

[KNT^{+}09] Chie Katsuyama, Shinji Nakaoka, Yasuhiro Takeuchi, Kanako Tago, Masahito Hayatsu, and Kenji Kato. Complementary cooperation between two syntrophic bacteria in pesticide degradation. Journal of Theoretical Biology, 256(4):644–654, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005481.

Kepler:2007:DMI

[KNWCB07] G. M. Kepler, H. K. Nguyen, J. Webster-Cyriaque, and H. T. Banks. A dynamic model for induced reactivation of latent virus. Journal of Theoretical Biology, 244(3):451–462, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003614.

Koch:2000:LDP

[Koc00] Arthur L. Koch. Length distribution of the peptidoglycan chains in the sacculus of Escherichia coli. Journal of Theoretical Biology, 204(4):533–541, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920393.

Kohler:2007:MMD

[Koh07] Brynja Kohler. Mathematically modeling dynamics of T cell responses: Predictions concerning the generation of memory cells. Journal of Theoretical Biology, 245(4):669–676, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004929.

Kokkorogiannis:2004:SID

[Kok04] Theodoros Kokkorogiannis. Somatic and intramuscular distribution of muscle spindles and their relation to muscular angiotypes. Journal of Theoretical Biology, 229(2):263–280, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001390.

Kondrashov:2006:SFN

[KOK06] Fyodor A. Kondrashov, Aleksey Y. Ogurtsov, and Alexey S. Kondrashov. Selection in favor of nucleotides G and C diversifies evolution rates and levels of polymorphism at mammalian synonymous sites. Journal of Theoretical Biology, 240(4):616– 626, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500473X.

Komarova:2004:RME

[Kom04] Natalia L. Komarova. Replicator-mutator equation, universality property and population dynamics of learning. Journal of Theoretical Biology, 230(2):227–239, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002073.

Komarova:2006:SMD

[Kom06] Natalia Komarova. Stochastic modeling of drug resistance in cancer. Journal of Theoretical Biology, 239(3):351–366, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003346.

Komarova:2007:VRS

[Kom07] Natalia L. Komarova. Viral reproductive strategies: How can lytic viruses be evolutionarily competitive? Journal of Theoretical Biology, 249(4):766–784, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004298.

Kondoh:2003:HFR

[Kon03] Michio Kondoh. Habitat fragmentation resulting in overgrazing by herbivores. Journal of Theoretical Biology, 225 (4):453–460, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303002790.

Kondoh:2006:DFA

[Kon06] Michio Kondoh. Does foraging adaptation create the positive complexity-stability relationship in realistic food-web structure? Journal of Theoretical Biology, 238(3):646–651, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002869.

Kong:2009:CCL

[Kon09] Yong Kong. Calculating complexity of large randomized libraries. Journal of Theoretical Biology, 259(3):641–645, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001684.

Korzeniewski:2001:CFD

[Kor01] Bernard Korzeniewski. Cybernetic formulation of the definition of life. Journal of Theoretical Biology, 209(3):275–286, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922623.

Korn:2007:WSC

[Kor07] Robert W. Korn. Watermelon stripes. A case for the clonal mosaic model in plants. Journal of Theoretical Biology, 247(4):859–861, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001920.

Korn:2009:NHS

[Kor09] Robert W. Korn. A new hypothesis for stomatal placement in arabidopsis. Journal of Theoretical Biology, 260(1):172–174, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002343.

Kimura:2007:SAM

[KOT07] H. Kimura, H. Okano, and R. J. Tanaka. Stochastic approach to molecular interactions and computational theory of metabolic and genetic regulations. Journal of Theoretical Biology, 248(4):590–607, October 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002962.

Kovac:2002:SHO

[Kov02] Vladimír Kovác. Synchrony and heterochrony in ontogeny (of fish). Journal of Theoretical Biology, 217(4):499–507, August 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930432.

Karagiannis:2006:DMC

[KP06] Emmanouil D. Karagiannis and Aleksander S. Popel. Distinct modes of collagen type I proteolysis by matrix metalloproteinase (MMP) 2 and membrane type I MMP during the migration of a tip endothelial cell: Insights from a computational model. Journal of Theoretical Biology, 238(1):124–145, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002225.

Kuchel:2008:ISI

[KP08] Philip W. Kuchel and David J. Philp. Isotopomer subspaces as indicators of metabolic-pathway structure. Journal of Theoretical Biology, 252(3):391–401, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002512.

Krakauer:2002:MDG

[KPS02] David C. Krakauer, Karen M. Page, and Stuart Sealfon. Module dynamics of the GnRH signal transduction network. Journal of Theoretical Biology, 218(4):457–470, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930924.

Kar:2005:SSG

[KR05] Sandip Kar and Deb Shankar Ray. Sustained simultaneous glycolytic and insulin oscillations in β-cells. Journal of Theoretical Biology, 237(1):58–66, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001529.

Kenah:2007:NBA

[KR07] Eben Kenah and James M. Robins. Network-based analysis of stochastic SIR epidemic models with random and proportionate mixing. Journal of Theoretical Biology, 249(4):706–722, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004304.

Krakauer:2001:SIP

[Kra01a] David C. Krakauer. Selective imitation for a private sign system. Journal of Theoretical Biology, 213(2):145–157, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924087.

Kral:2001:SO

[Krá01b] Petr Král. Species orthogonalization. Journal of Theoretical Biology, 212(3):355–366, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923823.

Kramer:2001:MMA

[Kra01c] Eric M. Kramer. A mathematical model of auxin-mediated radial growth in trees. Journal of Theoretical Biology, 208(4):387–397, February 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922203.

Kramer:2002:MMP

[Kra02] Eric M. Kramer. A mathematical model of pattern formation in the vascular cambium of trees. Journal of Theoretical Biology, 216(2):147–158, May 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925518.

Kjelstrup:2005:ATK

[KRB05a] S. Kjelstrup, J. M. Rubi, and D. Bedeaux. Active transport: a kinetic description based on thermodynamic grounds. Journal of Theoretical Biology, 234(1):7–12, May 7, 2005.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005235.

Kutalik:2005:CBS

[KRB05b] Zoltán Kutalik, Moe Razaz, and József Baranyi. Connection between stochastic and deterministic modelling of microbial growth. Journal of Theoretical Biology, 232(2):285–299, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003868.

King:2007:NTS

[KRGH07] J. W. King, K. Ruohonen, D. J. Grove, and A. Hammerstein. Nonlinear time series analysis of food intake in the dab and the rainbow trout. Journal of Theoretical Biology, 245(4):749–762, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005157.

Krishnan:2009:SPT

[Kri09] J. Krishnan. Signal processing through a generalized module of adaptation and spatial sensing. Journal of Theoretical Biology, 259(1):31–43, July 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000800.

Kliman:2001:DGC

[KRN01] Richard M. Kliman, Bryan T. Rogers, and Mohamed A. F. Noor. Differences in (g+c) content between species: a commentary on Forsdyke’s “chromosomal viewpoint” of speciation. Journal of Theoretical Biology, 209(2):131–140, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922422.

Konrad:2003:DGB

[KRN03] W. Konrad and A. Roth-Nebelsick. The dynamics of gas bubbles in conduits of vascular plants and implications for embolism repair. Journal of Theoretical Biology, 224(1):43–61, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001383.

Konrad:2008:MSD

[KRNG08] W. Konrad, A. Roth-Nebelsick, and M. Grein. Modelling of stomatal density response to atmospheric CO2. Journal of Theoretical Biology, 253(4):638–658, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001677.

Konrad:2000:TAE

[KRNKH00] W. Konrad, A. Roth-Nebelsick, H. Kerp, and H. Hass. Transpiration and assimilation of early Devonian land plants with axially symmetric telomes — simulations on the tissue level. Journal of Theoretical Biology, 206(1):91–107, September 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921076.

Kronauer:2008:GIK

[Kro08] Daniel J. C. Kronauer. Genomic imprinting and kinship in the social hymenoptera: What are the predictions? Journal of Theoretical Biology, 254(4):737–740, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003317.

Kim:2009:DES

[KRR09] J. W. Kim, J. A. Roberts, and P. A. Robinson. Dynamics of epileptic seizures: Evolution, spreading, and suppression. Journal of Theoretical Biology, 257(4):527–532, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006383.

Krupa:2002:NER

[Kru02] Boris Krupa. On the number of experiments required to find the causal structure of complex systems. Journal of Theoretical Biology, 219(2):257–267, November 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930293119X.

Kaandorp:2001:MMR

[KS01a] Jaap A. Kaandorp and Peter M. A. Sloot. Morphological models of radiate accretive growth and the influence of hydrodynamics. Journal of Theoretical Biology, 209(3):257–274, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922611.

Karasev:2001:TNG

[KS01b] V. A. Karasev and V. E. Stefanov. Topological nature of the genetic code. Journal of Theoretical Biology, 209(3):303–317, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922659.

Kleinstein:2001:TQS

[KS01c] Steven H. Kleinstein and Jaswinder Pal Singh. Toward quantitative simulation of germinal center dynamics: Biological and modeling insights from experimental validation. Journal of Theoretical Biology, 211(3):253–275, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923446.

Kliman:2003:PMP

[KS03] Harvey J. Kliman and Lee Segel. The placenta may predict the baby. Journal of Theoretical Biology, 225(1):143–145, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002480.

Klipcan:2004:AAB

[KS04] Liron Klipcan and Mark Safro. Amino acid biogenesis, evolution of the genetic code and aminoacyl-tRNA synthetases. Journal of Theoretical Biology, 228(3):389–396, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000505.

Kawaguchi:2006:WSI

[KS06] Isao Kawaguchi and Akira Sasaki. The wave speed of intergradation zone in two-species lattice Müllerian mimicry model. Journal of Theoretical Biology, 243(4):594–603, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002888.

Kasap:2008:CVB

[KS08a] Murat Kasap and Ali Sazci. The comparison of VEGFR-1-binding domain of VEGF-A with modelled VEGF-C sheds light on receptor specificity. Journal of Theoretical Biology, 253(3):446–451, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001422.

Kubrycht:2008:LHM

[KS08b] Jaroslav Kubrycht and Karel Sigler. Length of the hypermutation motif DGYW/WRCH in the focus of statistical limits. Implications for a double-motif or extended motif recognition models. Journal of Theoretical Biology, 255(1):8–15, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004098.

Kim:2009:CMF

[KS09] Jin Seob Kim and Sean X. Sun. Continuum modeling of forces in growing viscoelastic cytoskeletal networks. Journal of Theoretical Biology, 256(4):596–606, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800547X.

Kamo:2007:RTS

[KSB07] Masashi Kamo, Akira Sasaki, and Mike Boots. The role of trade-off shapes in the evolution of parasites in spatial host populations: An approximate analytical approach. Journal of Theoretical Biology, 244(4):588–596, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003663.

Kelly:2008:TGM

[KSB08] Catherine Kelly, Kieran Smallbone, and Michael Brady. Tumour glycolysis: The many faces of HIF. Journal of Theoretical Biology, 254(2):508–513, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002701.

Khadra:2009:RLA

[KSEK09] Anmar Khadra, Pere Santamaria, and Leah Edelstein-Keshet. The role of low avidity T cells in the protection against type 1 diabetes: a modeling investigation. Journal of Theoretical Biology, 256(1):126–141, January 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004918.

Kokshenev:2003:LBA

[KSG03] V. B. Kokshenev, J. K. L. Silva, and G. J. M. Garcia. Longbone allometry of terrestrial mammals and the geometricshape and elastic-force constraints of bone evolution. Journal of Theoretical Biology, 224(4):551–556, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001905.

Kim:2008:CMA

[KSK08] Jaeyeon Kim, Gerald M. Saidel, and Satish C. Kalhan. A computational model of adipose tissue metabolism: Evidence for intracellular compartmentation and differential activation of lipases. Journal of Theoretical Biology, 251(3):523–540, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307006273.

Kargol:2002:ANR

[KSM02] Armin Kargol, Bryan Smith, and Mark M. Millonas. Applications of nonequilibrium response spectroscopy to the study of channel gating. Experimental design and optimization. Journal of Theoretical Biology, 218(2):239–258, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930730.

Komarova:2003:MSN

[KSN03] Natalia L. Komarova, Anirvan Sengupta, and Martin A. Nowak. Mutation-selection networks of cancer initiation: tumor suppressor genes and chromosomal instability. Journal of Theoretical Biology, 223(4):433–450, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001206.

Kohandel:2006:MMO

[KSO06] M. Kohandel, S. Sivaloganathan, and A. Oza. Mathematical modeling of ovarian cancer treatments: Sequencing of surgery and chemotherapy. Journal of Theoretical Biology, 242(1):62–68, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000579.

Kurz:2008:PWG

[KSPA^{+}08] Max J. Kurz, Melissa Scott-Pandorf, Chris Arellano, Diane Olsen, and Greg Whitaker. The penguin waddling gait pattern has a more consistent step width than step length. Journal of Theoretical Biology, 252(2):272–276, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000246.

Kurgan:2007:NSB

[KSR07] Lukasz A. Kurgan, Wojciech Stach, and Jishou Ruan. Novel scales based on hydrophobicity indices for secondary protein structure. Journal of Theoretical Biology, 248(2):354–366, September 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002482.

Koschutzki:2007:RNE

[KSS07] Dirk Koschützki, Henning Schwöbbermeyer, and Falk Schreiber. Ranking of network elements based on functional substructures. Journal of Theoretical Biology, 248(3):471–479, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002780.

Kaufman:2007:NNC

[KST07] M. Kaufman, C. Soulé, and R. Thomas. A new necessary condition on interaction graphs for multistationarity. Journal of Theoretical Biology, 248(4):675–685, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002950.

Kawachi:2008:RTM

[KSY^{+}08] Kazuki Kawachi, Motohide Seki, Hiraku Yoshida, Yohei Otake, Katsuhide Warashina, and Hiroshi Ueda. A rumor transmission model with various contact interactions. Journal of Theoretical Biology, 253(1):55–60, July 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005954.

Kon:2001:EEH

[KT01] Ryusuke Kon and Yasuhiro Takeuchi. The effect of evolution on host-parasitoid systems. Journal of Theoretical Biology, 209(3):287–302, April 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922635.

Koeslag:2003:ECC

[KT03] Johan H. Koeslag and Elmarie Terblanche. Evolution of cooperation: cooperation defeats defection in the cornfield model. Journal of Theoretical Biology, 224(3):399–410, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001887.

Kysela:2007:OBM

[KT07] David T. Kysela and Paul E. Turner. Optimal bacteriophage mutation rates for phage therapy. Journal of Theoretical Biology, 249(3):411–421, December 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003827.

Kansal:2000:ESC

[KTCD00] A. R. Kansal, S. Torquato, E. A. Chiocca, and T. S. Deisboeck. Emergence of a subpopulation in a computational model of tumor growth. Journal of Theoretical Biology, 207(3):431–441, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921866.

Keef:2008:BVC

[KTE08] T. Keef, R. Twarock, and K. M. Elsawy. Blueprints for viral capsids in the family of polyomaviridae. Journal of Theoretical Biology, 253(4):808–816, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001847.

Kansal:2000:SBT

[KTH^{+}00] A. R. Kansal, S. Torquato, G. R. Harsh, E. A. Chiocca, and T. S. Deisboeck. Simulated brain tumor growth dynamics using a three-dimensional cellular automaton. Journal of Theoretical Biology, 203(4):367–382, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920009.

Klimek:2009:PTL

[KTH09] Peter Klimek, Stefan Thurner, and Rudolf Hanel. Pruning the tree of life: k-core percolation as selection mechanism. Journal of Theoretical Biology, 256(1):142–146, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004979.

Katsaloulis:2009:LRC

[KTP09] P. Katsaloulis, T. Theoharis, and A. Provata. Long range clustering of oligonucleotides containing the CG signal. Journal of Theoretical Biology, 258(1):18–26, May 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000228.

Kato:2007:ETS

[KUK07] Satoshi Kato, Jotaro Urabe, and Masakado Kawata. Effects of temporal and spatial heterogeneities created by consumerdriven nutrient recycling on algal diversity. Journal of Theoretical Biology, 245(2):364–377, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004887.

Kunisawa:2000:FRM

[Kun00] Takashi Kunisawa. Functional role of mycobacteriophage transfer RNAs. Journal of Theoretical Biology, 205(1):167– 170, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920575.

Kunisawa:2001:GAP

[Kun01] Takashi Kunisawa. Gene arrangements and phylogeny in the class proteobacteria. Journal of Theoretical Biology, 213(1):9–19, November 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923963.

Kunisawa:2003:GAB

[Kun03] Takashi Kunisawa. Gene arrangements and branching orders of gram-positive bacteria. Journal of Theoretical Biology, 222(4):495–503, June 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300064X.

Kunisawa:2005:BOA

[Kun05] Takashi Kunisawa. Branching orders among α-proteobacteria and mitochondria inferred from gene transpositions. Journal of Theoretical Biology, 234(1):1–5, May 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005260.

Kunisawa:2006:DMB

[Kun06] Takashi Kunisawa. Dichotomy of major bacterial phyla inferred from gene arrangement comparisons. Journal of Theoretical Biology, 239(3):367–375, April 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003358.

Kurochkina:2007:AAC

[Kur07] N. Kurochkina. Amino acid composition of parallel helixhelix interfaces. Journal of Theoretical Biology, 247(1):110– 121, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000677.

Kurbel:2008:EOS

[Kur08a] Sven Kurbel. Are extracellular osmolality and sodium concentration determined by Donnan effects of intracellular protein charges and of pumped sodium? Journal of Theoretical Biology, 252(4):769–772, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000866.

Kurochkina:2008:SSC

[Kur08b] N. Kurochkina. Specific sequence combinations at parallel and antiparallel helix-helix interfaces. Journal of Theoretical Biology, 255(2):188–198, November 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004426.

Kuthan:2003:MMS

[Kut03] Hartmut Kuthan. A mathematical model of single target site location by Brownian movement in subcellular compartments. Journal of Theoretical Biology, 221(1):79–87, March 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303931729.

Kuthan:2005:TFN

[Kut05] Hartmut Kuthan. Temporal fluctuation of nuclear pore complex localization by single diffusing mRNP complexes. Journal of Theoretical Biology, 236(3):256–262, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001207.

Krivan:2004:SHP

[KV04] Vlastimil Krivan and Ivo Vrkoc. Should “handled” prey be considered? Some consequences for functional response, predator-prey dynamics and optimal foraging theory. Journal of Theoretical Biology, 227(2):167–174, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003989.

Kanazawa:2005:EMS

[KV05] Satoshi Kanazawa and Griet Vandermassen. Engineers have more sons, nurses have more daughters: an evolutionary psychological extension of Baron–Cohen’s extreme male brain theory of autism. Journal of Theoretical Biology, 233(4):589–599, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005223.

Kletter:2009:PVR

[KvHMP09] A. Y. Kletter, J. von Hardenberg, E. Meron, and A. Provenzale. Patterned vegetation and rainfall intermittency. Journal of Theoretical Biology, 256(4):574–583, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005614.

Kranenbarg:2001:CFC

[KVMV01] S. Kranenbarg, J. H. G. Verhagen, M. Muller, and J. L. Van Leeuwen. Consequences of forced convection for the constraints on size and shape in embryos. Journal of Theoretical Biology, 212(4):521–533, October 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923914.

Kelly:2000:MEN

[KW00a] John K. Kelly and Michael J. Wade. Molecular evolution near a two-locus balanced polymorphism. Journal of Theoretical Biology, 204(1):83–101, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920034.

Kelly:2000:PRS

[KW00b] John K. Kelly and Scott Williamson. Predicting response to selection on a quantitative trait: a comparison between models for mixed-mating populations. Journal of Theoretical Biology, 207(1):37–56, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921544.

Kerszberg:2000:CTM

[KW00c] Michel Kerszberg and Lewis Wolpert. A clock and trail model for somite formation, specialization and polarization. Journal of Theoretical Biology, 205(3):505–510, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092085X.

Krakovska:2007:ODT

[KW07] O. Krakovska and L. M. Wahl. Optimal drug treatment regimens for HIV depend on adherence. Journal of Theoretical Biology, 246(3):499–509, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306006047.

Kenward:2004:SSE

[KWGE04] Benjamin Kenward, Carl-Adam Wachtmeister, Stefano Ghirlanda, and Magnus Enquist. Spots and stripes: the evolution of repetition in visual signal form. Journal of Theoretical Biology, 230(3):407–419, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002814.

Kobayashi:2000:ESD

[KY00] Yutaka Kobayashi and Norio Yamamura. Evolution of seed dormancy due to sib competition: Effect of dispersal and inbreeding. Journal of Theoretical Biology, 202(1):11–24, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910275.

Kaneko:2002:KOH

[KY02] Kunihiko Kaneko and Tetsuya Yomo. On a kinetic origin of heredity: Minority control in a replicating system with mutually catalytic molecules. Journal of Theoretical Biology, 214(4):563–576, February 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924816.

Kobayashi:2003:ESE

[KY03] Yutaka Kobayashi and Norio Yamamura. Evolution of signal emission by non-infested plants growing near infested plants to avoid future risk. Journal of Theoretical Biology, 223(4):489–503, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001243.

Kobayashi:2006:ETP

[KYS06] Yutaka Kobayashi, Norio Yamamura, and Maurice W. Sabelis. Evolution of talking plants in a tritrophic context: Conditions for uninfested plants to attract predators prior to herbivore attack. Journal of Theoretical Biology, 243(3):361–374, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002220.

Kang:2008:EBA

[KYZ^{+}08] Guolian Kang, Weihua Yue, Jifeng Zhang, Yuehua Cui, Yijun Zuo, and Dai Zhang. An entropy-based approach for testing genetic epistasis underlying complex diseases. Journal of Theoretical Biology, 250(2):362–374, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700478X.

Koslowski:2005:TQU

[KZ05] Thorsten Koslowski and Fabian Zehender. Towards a quantitative understanding of horizontal gene transfer: a kinetic model. Journal of Theoretical Biology, 237(1):23–29, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001499.

Lacks:2001:RBE

[Lac01] Daniel J. Lacks. Regressive biological evolution due to environmental change. Journal of Theoretical Biology, 209(4):487– 491, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922829.

Lachance:2009:IPS

[Lac09] Joseph Lachance. Inbreeding, pedigree size, and the most recent common ancestor of humanity. Journal of Theoretical Biology, 261(2):238–247, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003701.

Lin:2003:TWM

[LACL03] Juan Lin, Viggo Andreasen, Renato Casagrandi, and Simon A. Levin. Traveling waves in a model of influenza A drift. Journal of Theoretical Biology, 222(4):437–445, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000560.

Lett:2003:MFP

[LAD03] Christophe Lett, Pierre Auger, and Rafael Bravo De La Parra. Migration frequency and the persistence of host–parasitoid interactions. Journal of Theoretical Biology, 221(4):639–654, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932139.

Lopez:2007:RHG

[LAG07] Alejandro López, Humberto Arce, and Michael R. Guevara. Rhythms of high-grade block in an ionic model of a strand of regionally ischemic ventricular muscle. Journal of Theoretical Biology, 249(1):29–45, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002937.

Lemasson:2009:CMA

[LAG09] B. H. Lemasson, J. J. Anderson, and R. A. Goodwin. Collective motion in animal groups from a neurobiological perspective: The adaptive benefits of dynamic sensory loads and selective attention. Journal of Theoretical Biology, 261(4):501–510, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003658.

Lalam:2006:ERE

[Lal06] Nadia Lalam. Estimation of the reaction efficiency in polymerase chain reaction. Journal of Theoretical Biology, 242(4):947–953, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002281.

Landry:2000:HCE

[Lan00a] Samuel J. Landry. Helper T-cell epitope immunodominance associated with structurally stable segments of hen egg lysozyme and HIV gp120. Journal of Theoretical Biology, 203(3):189– 201, April 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910561.

Lange:2000:MIC

[Lan00b]  Klaus Lange.  Microvillar ion channels: Cytoskeletal

modulation of ion fluxes. Journal of Theoretical Biology, 206(4):561–584, October 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921465.

Lane:2003:UVA

[Lan03] Nick Lane. A unifying view of ageing and disease: the doubleagent theory. Journal of Theoretical Biology, 225(4):531–540, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003047.

Lappa:2003:GFD

[Lap03] Marcello Lappa. The growth and the fluid dynamics of protein crystals and soft organic tissues: models and simulations, similarities and differences. Journal of Theoretical Biology, 224(2):225–240, September 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001607.

Lazenby:2002:SBF

[Laz02] Richard A. Lazenby. Skeletal biology, functional asymmetry and the origins of “handedness”. Journal of Theoretical Biology, 218(1):129–138, September 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930523.

Lazar:2003:CFR

[Laz03] Dusan Lazár. Chlorophyll a fluorescence rise induced by high light illumination of dark-adapted plant tissue studied by means of a model of photosystem II and considering photosystem II heterogeneity. Journal of Theoretical Biology, 220(4):469–503, February 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931407.

Ladley:2005:RLC

[LB05] Dan Ladley and Seth Bullock. The role of logistic constraints in termite construction of chambers and tunnels. Journal of Theoretical Biology, 234(4):551–564, June 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006113.

Lewis:2006:PPR

[LB06] D. M. Lewis and S. I. Bala. Plankton predation rates in turbulence: a study of the limitations imposed on a predator with a non-spherical field of sensory perception. Journal of Theoretical Biology, 242(1):44–61, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000567.

Le:2007:EDC

[LB07] Stephen Le and Robert Boyd. Evolutionary dynamics of the continuous iterated Prisoner’s dilemma. Journal of Theoretical Biology, 245(2):258–267, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004255.

Lefever:2009:DGV

[LBCL09] René Lefever, Nicolas Barbier, Pierre Couteron, and Olivier Lejeune. Deeply gapped vegetation patterns: On crown/ root allometry, criticality and desertification. Journal of Theoretical Biology, 261(2):194–209, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003427.

Law:2001:MEP

[LBF01] Richard Law, Judith L. Bronstein, and Régis Ferrière. On mutualists and exploiters: Plant-insect coevolution in pollinating seed-parasite systems. Journal of Theoretical Biology, 212(3):373–389, October 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923835.

Luczkovich:2003:DMT

[LBJE03] Joseph J. Luczkovich, Stephen P. Borgatti, Jeffrey C. Johnson, and Martin G. Everett. Defining and measuring trophic role similarity in food Webs using regular equivalence. Journal of Theoretical Biology, 220(3):303–321, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930393147X.

Lecarpentier:2005:SMM

[LBQ^{+}05] Y. Lecarpentier, F.-X. Blanc, J. Quillard, J.-L. Hébert, X. Krokidis, and C. Coirault. Statistical mechanics of myosin molecular motors in skeletal muscles. Journal of Theoretical Biology, 235(3):381–392, August 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000445.

Latowski:2000:MMD

[LBS00] Dariusz Latowski, Kvetoslava Burda, and Kazimierz Strzalka. A mathematical model describing kinetics of conversion of violaxanthin to zeaxanthin via intermediate antheraxanthin by the xanthophyll cycle enzyme violaxanthin Deepoxidase. Journal of Theoretical Biology, 206(4):507–514, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921416.

Lee:2006:FER

[LBS06] S.-H. Lee, P. Bardunias, and N.-Y. Su. Food encounter rates of simulated termite tunnels with variable food size/ distribution pattern and tunnel branch length. Journal of Theoretical Biology, 243(4):493–500, December 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003304.

Lin:2006:MSW

[LBSP06] Anita Shu-Han Lin, Martin L. Buist, Nicolas P. Smith, and Andrew J. Pullan. Modelling slow wave activity in the small intestine. Journal of Theoretical Biology, 242(2):356–362, September 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001093.

Lorberbaum:2002:TRN

[LBSS02] Jeffrey P. Lorberbaum, Daryl E. Bohning, Ananda Shastri, and Lauren E. Sine. Are there really no evolutionarily stable strategies in the iterated prisoner’s dilemma? Journal of Theoretical Biology, 214(2):155–169, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924555. See erratum [LW04].

Laoudj-Chenivesse:2003:FDD

[LCBTP03] Dalila Laoudj-Chenivesse, René Bennes, and Emmanuel Tronel-Peyroz. On the flexibility and the dynamics of DNA. Journal of Theoretical Biology, 225(4):541–543, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002571.

Luo:2006:MCM

[LCHB06] C. Luo, J. W. Clark, T. A. Heming, and A. Bidani. A macrophage cell model for pH and volume regulation. Journal of Theoretical Biology, 238(2):449–463, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002365.

Lee:2009:ADS

[LCHK09] Dooyoung Lee, Kelly E. Caputo, Daniel A. Hammer, and Michael R. King. Adhesive dynamics simulations of the mechanical shedding of l-selectin from the neutrophil surface. Journal of Theoretical Biology, 260(1):27–30, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002355.

Lagomarsino:2003:SSD

[LCL03] M. C. Lagomarsino, F. Capuani, and C. P. Lowe. A simulation study of the dynamics of a driven filament in an Aristotelian fluid. Journal of Theoretical Biology, 224(2):215–224, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001590.

Lomasko:2007:OHSa

[LCL07a] Tatiana Lomasko, Geoff Clarke, and Charles J. Lumsden. One-hit stochastic decline in a mechanochemical model of cytoskeleton-induced neuron death i: Cell-fate arrival times. Journal of Theoretical Biology, 249(1):1–17, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930700272X.

Lomasko:2007:OHSb

[LCL07b] Tatiana Lomasko, Geoff Clarke, and Charles J. Lumsden. One-hit stochastic decline in a mechanochemical model of cytoskeleton-induced neuron death II: Transition state metastability. Journal of Theoretical Biology, 249(1):18–28, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002731.

Leon:2000:CSJ

[LD00] Jesús Alberto León and José Renato De Nóbrega. Comparative statics of joint reproductive allocation. Journal of Theoretical Biology, 205(4):563–579, August 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920861.

Lange:2009:BAE

[LD09a] Alexander Lange and Reuven Dukas. Bayesian approximations and extensions: Optimal decisions for small brains and possibly big ones too. Journal of Theoretical Biology, 259(3):503–516, August 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001271.

Legendre:2009:ICR

[LD09b] Pierre Legendre and Yves Desdevises. Independent contrasts and regression through the origin. Journal of Theoretical Biology, 259(4):727–743, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001921.

Lof:2009:OMA

[LdGH09] Marjolein E. Lof, Maarten de Gee, and Lia Hemerik. Odor-mediated aggregation enhances the colonization ability of Drosophila melanogaster. Journal of Theoretical Biology, 258(3):363–370, June 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004207.

Lin:2009:SCB

[LDT09] Huai-Ti Lin, A. Luis Dorfmann, and Barry A. Trimmer. Soft-cuticle biomechanics: a constitutive model of anisotropy for caterpillar integument. Journal of Theoretical Biology, 256(3):447–457, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005420.

Li:2004:BDS

[LDW04] Wei Li, Shuo-Xing Dou, and Peng-Ye Wang. Brownian dynamics simulation of nucleosome formation and disruption under stretching. Journal of Theoretical Biology, 230(3):375–383, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002644.

Li:2005:HOI

[LDW05] Wei Li, Shuo-Xing Dou, and Peng-Ye Wang. The histone octamer influences the wrapping direction of DNA on it: Brownian dynamics simulation of the nucleosome chirality. Journal of Theoretical Biology, 235(3):365–372, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305000421.

Li:2009:FMH

[LDW09] Ning Li, Kiranmoy Das, and Rongling Wu. Functional mapping of human growth trajectories. Journal of Theoretical Biology, 261(1):33–42, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003269.

Liu:2006:PCN

[LDXW06] Xiao Qing Liu, Qi Dai, Zhilong Xiu, and Tianming Wang. PNN-curve: a new 2d graphical representation of DNA sequences and its application. Journal of Theoretical Biology, 243(4):555–561, December 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003092.

Lehmann:2000:PCC

[Leh00] Jean Lehmann. Physico-chemical constraints connected with the coding properties of the genetic system. Journal of Theoretical Biology, 202(2):129–144, January 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910457.

Lehmann:2002:ASD

[Leh02] Jean Lehmann. Amplification of the sequences displaying the pattern RNY in the RNA world: the translation → translation/replication hypothesis. Journal of Theoretical Biology, 219(4):521–537, December 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931425.

Lei:2009:SSG

[Lei09] Jinzhi Lei. Stochasticity in single gene expression with both intrinsic noise and fluctuation in kinetic parameters. Journal of Theoretical Biology, 256(4):485–492, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005778. See corrigendum [Lei10].

Lei:2010:CSS

[Lei10] Jinzhi Lei. Corrigendum to: “Stochasticity in single gene expression with both intrinsic noise and fluctuation in kinetic parameters” [J. Theor. Biol. 256 (2009) 485–492]. Journal of Theoretical Biology, 262(2):381, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004718. See [Lei09].

Lenbury:2001:ICC

[Len01] Yongwimon Lenbury. International conference on computational mathematics and modeling. Journal of Theoretical Biology, 212(4):563, October 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924099.

Leow:2007:MMP

[Leo07] Melvin Khee-Shing Leow. A mathematical model of pituitarythyroid interaction to provide an insight into the nature of the thyrotropin-thyroid hormone relationship. Journal of Theoretical Biology, 248(2):275–287, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002470.

Lewis:2003:PER

[Lew03] D. M. Lewis. Planktonic encounter rates in homogeneous isotropic turbulence: the case of predators with limited fields of sensory perception. Journal of Theoretical Biology, 222(1):73–97, May 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000158.

Lewis:2005:SMP

[Lew05] D. M. Lewis. A simple model of plankton population dynamics coupled with a LES of the surface mixed layer. Journal of Theoretical Biology, 234(4):565–591, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006125.

Lapedes:2001:GSS

[LF01] Alan Lapedes and Robert Farber. The geometry of shape space: Application to influenza. Journal of Theoretical Biology, 212(1):57–69, September 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923471.

Lu:2008:QIC

[LF08] Yi Lu and Stephen J. Freeland. A quantitative investigation of the chemical space surrounding amino acid alphabet formation. Journal of Theoretical Biology, 250(2):349–361, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004936.

Leon:2004:GMF

[LFC04] Kalet León, Jose Faro, and Jorge Carneiro. A general mathematical framework to model generation structure in a population of asynchronously dividing cells. Journal of Theoretical Biology, 229(4):455–476, August 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001547.

Li:2006:PHI

[LFF06] Xu-Qing Li, Ping Fan, and Jinghong Fan. Polarity and hydrophobicity interactions in protein synthesis process. Journal of Theoretical Biology, 240(1):87–97, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003693.

Lopez-Fanjul:2006:EGD

[LFFT06] Carlos López-Fanjul, Almudena Fernández, and Miguel A. Toro. The effect of genetic drift on the variance/covariance components generated by multilocus additive× additive epistatic systems. Journal of Theoretical Biology, 239(2):161– 171, March 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003772.

Laurence:2000:BEE

[LFLM00] Jocelyn A. Laurence, Peter W. French, Robyn A. Lindner, and David R. McKenzie. Biological effects of electromagnetic fields — mechanisms for the effects of pulsed microwave radiation on protein conformation. Journal of Theoretical Biology, 206(2):291–298, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921234.

Little:2005:MRI

[LFP^{+}05] M. P. Little, J. A. N. Filipe, K. M. Prise, M. Folkard, and O. V. Belyakov. A model for radiation-induced bystander effects, with allowance for spatial position and the effects of cell turnover. Journal of Theoretical Biology, 232(3):329–338, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003893.

Liepelt:2005:IPN

[LFSG^{+}05] Steffen Liepelt, Jan A. Freund, Lutz Schimansky-Geier, Alexander Neiman, and David F. Russell. Information processing in noisy burster models of sensory neurons. Journal of Theoretical Biology, 237(1):30–40, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001505.

Lopez-Fidalgo:2002:DIM

[LFW02] J. López-Fidalgo and Weng Kee Wong. Design issues for the Michaelis–Menten model. Journal of Theoretical Biology, 215(1):1–11, March 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930192497X.

Leloup:2004:MMC

[LG04] Jean-Christophe Leloup and Albert Goldbeter. Modeling the mammalian circadian clock: Sensitivity analysis and multiplicity of oscillatory mechanisms. Journal of Theoretical Biology, 230(4):541–562, October 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002437.

Larreta-Garde:2002:MEM

[LGB02]  Véronique Larreta-Garde and Hugues Berry.  Model-

ing extracellular matrix degradation balance with proteinase/transglutaminase cycle. Journal of Theoretical Biology, 217(1):105–124, July 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930109.

Lemon:2003:MRAa

[LGB03a] G. Lemon, W. G. Gibson, and M. R. Bennett. Metabotropic receptor activation, desensitization and sequestration — I: modelling calcium and inositol 1,4,5-trisphosphate dynamics following receptor activation. Journal of Theoretical Biology, 223(1):93–111, July 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000791.

Lemon:2003:MRAb

[LGB03b] G. Lemon, W. G. Gibson, and M. R. Bennett. Metabotropic receptor activation, desensitization and sequestration — II: modelling the dynamics of the pleckstrin homology domain. Journal of Theoretical Biology, 223(1):113–129, July 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000808.

Leon:2007:HRC

[LGCL07] Kalet Leon, Karina Garcia, Jorge Carneiro, and Agustin Lage. How regulatory CD25 + CD4 + T cells impinge on tumor immunobiology? On the existence of two alternative dynamical classes of tumors. Journal of Theoretical Biology, 247(1):122– 137, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000598.

Lema:2000:DMC

[LGE00] Martín A. Lema, Diego A. Golombek, and Julián Echave. Delay model of the circadian pacemaker. Journal of Theoretical Biology, 204(4):565–573, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920381.

Lee:2009:MSE

[LGK^{+}09] Ha Youn Lee, Elena E. Giorgi, Brandon F. Keele, Brian Gaschen, Gayathri S. Athreya, Jesus F. Salazar-Gonzalez, Kimmy T. Pham, Paul A. Goepfert, J. Michael Kilby, Michael S. Saag, Eric L. Delwart, Michael P. Busch, Beatrice H. Hahn, George M. Shaw, Bette T. Korber, Tanmoy Bhattacharya, and Alan S. Perelson. Modeling sequence evolution in acute HIV-1 infection. Journal of Theoretical Biology, 261(2):341–360, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519309003506. See corrigendum [LGK^{+}12].

Lee:2012:CMS

[LGK^{+}12] Ha Youn Lee, Elena E. Giorgi, Brandon F. Keele, Brian Gaschen, Gayathri S. Athreya, Jesus F. Salazar-Gonzalez, Kimmy T. Pham, Paul A. Goepfert, J. Michael Kilby, Michael S. Saag, Eric L. Delwart, Michael P. Busch, Beatrice H. Hahn, George M. Shaw, Bette T. Korber, Tanmoy Bhattacharya, and Alan S. Perelson. Corrigendum to “Modeling sequence evolution in acute HIV-1 infection” [J. Theor. Biol. 261 (2009) 341–360]. Journal of Theoretical Biology, 297(??):187, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006199. See [LGK^{+}09].

Lu:2009:API

[LGS^{+}09] Zhao Lu, Yrjo T. Grohn, Rebecca L. Smith, David R. Wolfgang, Jo Ann S. Van Kessel, and Ynte H. Schukken. Assessing the potential impact of salmonella vaccines in an endemically infected dairy herd. Journal of Theoretical Biology, 259(4):770–784, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002033.

Longuet-Higgins:2001:GCD

[LH01] Michael S. Longuet-Higgins. Geometrical constraints on the development of a diatom. Journal of Theoretical Biology, 210(1):101–105, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930192301X.

Lindner:2006:EDO

[LH06] Patrick Felix Oliver Lindner and Bernd Hitzmann. Experimental design for optimal parameter estimation of an enzyme kinetic process based on the analysis of the Fisher information matrix. Journal of Theoretical Biology, 238(1):111–123, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002183.

Lavrentovich:2008:MMS

[LH08] Maxim Lavrentovich and Sheryl Hemkin. A mathematical model of spontaneous calcium(II) oscillations in astrocytes. Journal of Theoretical Biology, 251(4):553–560, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006510.

Lavrentovich:2009:CMM

[LH09] Maxim Lavrentovich and Sheryl Hemkin. Corrigendum to “A mathematical model of spontaneous calcium(II) oscillations in astrocytes” [J. Theor. Biol. 251 (2008) 553–560]. Journal of Theoretical Biology, 260(2):332, September 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002641.

Lee:2001:NLC

[LHD^{+}01] C. T. Lee, M. F. Hoopes, J. Diehl, W. Gilliland, G. Huxel, E. V. Leaver, K. Mccann, J. Umbanhowar, and A. Mogilner. Non-local concepts and models in biology. Journal of Theoretical Biology, 210(2):201–219, May 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922872.

Luttikhuizen:2004:GSO

[LHDvdM04] P. C. Luttikhuizen, P. J. C. Honkoop, J. Drent, and J. van der Meer. A general solution for optimal egg size during external fertilization, extended scope for intermediate optimal egg size and the introduction of Don ottavio ‘tango’. Journal of Theoretical Biology, 231(3):333–343, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003194.

Lipniacki:2008:SEB

[LHFH08] Tomasz Lipniacki, Beata Hat, James R. Faeder, and William S. Hlavacek. Stochastic effects and bistability in T cell receptor signaling. Journal of Theoretical Biology, 254(1):110–122, September 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002208.

Liebermeister:2005:RTP

[Lie05] Wolfram Liebermeister. Response to temporal parameter fluctuations in biochemical networks. Journal of Theoretical Biology, 234(3):423–438, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304006022.

Lazar:2005:TSE

[LIK^{+}05] Dusan Lazár, Petr Ilík, Jerzy Kruk, Kazimierz Strzalka, and Jan Naus. A theoretical study on effect of the initial redox state of cytochrome b 559 on maximal chlorophyll fluorescence level (F M): implications for photoinhibition of photosystem II. Journal of Theoretical Biology, 233(2):287–300, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004862.

Line:2001:MMF

[Lin01] Sergio R. P. Line. Molecular morphogenetic fields in the development of human dentition. Journal of Theoretical Biology, 211(1):67–75, July 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923331.

Lind:2004:WDP

[Lin04] Johan Lind. What determines probability of surviving predator attacks in bird migration?: the relative importance of vigilance and fuel load. Journal of Theoretical Biology, 231(2):223–227, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002899.

Lin:2007:MMP

[Lin07] Shili Lin. Mixture modeling of progression pathways of heterogeneous breast tumors. Journal of Theoretical Biology, 249(2):254–261, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003931.

Lin:2008:MMD

[Lin08] Hao Lin. The modified Mahalanobis discriminant for predicting outer membrane proteins by using Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 252(2):350– 356, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000556.

Lion:2009:RSS

[Lio09] Sébastien Lion. Relatedness in spatially structured populations with empty sites: An approach based on spatial moment equations. Journal of Theoretical Biology, 260(1):121–131, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002409.

Little:2007:MCC

[Lit07] Mark P. Little. A multi-compartment cell repopulation model allowing for inter-compartmental migration following radiation exposure, applied to leukaemia. Journal of Theoretical Biology, 245(1):83–97, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004504.

Lawson:2006:SAR

[LJ06] Daniel Lawson and Henrik Jeldtoft Jensen. The speciesarea relationship and evolution. Journal of Theoretical Biology, 241(3):590–600, August 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005576.

Lawson:2009:RWS

[LJ09a] Daniel John Lawson and Henrik Jeldtoft Jensen. The role of weak selection and high mutation rates in nearly neutral evolution. Journal of Theoretical Biology, 257(4):696–703, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006747.

Lazar:2009:AAF

[LJ09b] Dusan Lazár and Jirí Jablonský. On the approaches applied in formulation of a kinetic model of photosystem II: Different approaches lead to different simulations of the chlorophyll a fluorescence transients. Journal of Theoretical Biology, 257(2):260–269, March 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006152.

Lawson:2006:DPI

[LJK06] Daniel Lawson, Henrik Jeldtoft Jensen, and Kunihiko Kaneko. Diversity as a product of inter-specific interactions. Journal of Theoretical Biology, 243(3):299–307, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002992.

Liu:2008:NIS

[LJL08] Quan-Xing Liu, Zhen Jin, and Bai-Lian Li. Numerical investigation of spatial pattern in a vegetation model with feedback function. Journal of Theoretical Biology, 254(2):350–360, September 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002476.

Lenaerts:2005:ELG

[LJTD05] Tom Lenaerts, Bart Jansen, Karl Tuyls, and Bart De Vylder. The evolutionary language game: An orthogonal approach. Journal of Theoretical Biology, 235(4):566–582, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930500069X.

Logan:2002:LTT

[LJW02] J. David Logan, Anthony Joern, and William Wolesensky. Location, time, and temperature dependence of digestion in simple animal tracts. Journal of Theoretical Biology, 216(1):5– 18, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925415.

Lyubich:2003:MFP

[LK03] Yuri Lyubich and Valery Kirzhner. Mathematical frameworks for phenotypical selection and epistasis. Journal of Theoretical Biology, 221(4):625–638, April 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303932127.

Lao:2006:CMI

[LK06] Bert J. Lao and Daniel T. Kamei. A compartmental model of iron regulation in the mouse. Journal of Theoretical Biology, 243(4):542–554, December 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002700.

Lao:2008:ICM

[LK08a] Bert J. Lao and Daniel T. Kamei. Investigation of cellular movement in the prostate epithelium using an agent-based model. Journal of Theoretical Biology, 250(4):642–654, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005486.

Lee:2008:CIC

[LK08b] Young-Seon Lee and James P. Keener. A calcium-induced calcium release mechanism mediated by calsequestrin. Journal of Theoretical Biology, 253(4):668–679, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001781.

Locke:2008:DCE

[LK08c] B. R. Locke and S. T. Kinsey. Diffusional constraints on energy metabolism in skeletal muscle. Journal of Theoretical Biology, 254(2):417–429, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800307X.

Lin:2007:MMC

[LKC07] Ximin Lin, James S. Koopman, and Stephen E. Chick. Mathematical model comparisons of potential non-typeable haemophilus influenzae vaccine effects. Journal of Theoretical Biology, 245(1):66–76, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600467X.

Lehmann:2007:MCE

[LKK07] Laurent Lehmann, Lukas F. Keller, and Hanna Kokko. Mate choice evolution, dominance effects, and the maintenance of genetic variation. Journal of Theoretical Biology, 244(2):282–295, January 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003316.

Li:2006:MGI

[LKM06] Jiaxu Li, Yang Kuang, and Clinton C. Mason. Modeling the glucose-insulin regulatory system and ultradian insulin secretory oscillations with two explicit time delays. Journal of Theoretical Biology, 242(3):722–735, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001391.

Lecarpentier:2008:IEP

[LKM^{+}08] Y. Lecarpentier, X. Krokidis, P. Martin, T. Pineau, J.-L. Hébert, J. Quillard, M. Cortes-Morichetti, and C. Coirault. Increased entropy production in diaphragm muscle of PPAR α knockout mice. Journal of Theoretical Biology, 250(1):92–102, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004262.

Li:2006:IQT

[LKW06] Hongying Li, Bong-Rae Kim, and Rongling Wu. Identification of quantitative trait nucleotides that regulate cancer growth: a simulation approach. Journal of Theoretical Biology, 242(2):426–439, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001172.

Lenton:2000:DDC

[LL00] Timothy M. Lenton and James E. Lovelock. Daisyworld is Darwinian: Constraints on adaptation are important for planetary self-regulation. Journal of Theoretical Biology, 206(1):109–114, September 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921052.

Li:2001:PSC

[LL01a] Qian-Zhong Li and Zhi-Qing Lu. The prediction of the structural class of protein: Application of the measure of diversity. Journal of Theoretical Biology, 213(3):493–502, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924415.

Lifson:2001:CDS

[LL01b] Shneior Lifson and Hanna Lifson. Coexistence and Darwinian selection among replicators: Response to the preceding paper by scheuring and szathmary. Journal of Theoretical Biology, 212(1):107–109, September 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923616.

Lindner:2005:EED

[LL05] Benjamin Lindner and André Longtin. Effect of an exponentially decaying threshold on the firing statistics of a stochastic integrate-and-fire neuron. Journal of Theoretical Biology, 232(4):505–521, February 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004187.

Li:2006:RPP

[LL06] Qian-Zhong Li and Hao Lin. The recognition and prediction of σ70 promoters in Escherichia coli K-12. Journal of Theoretical Biology, 242(1):135–141, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000646.

Lewis:2007:EDE

[LL07] Hannah M. Lewis and Richard Law. Effects of dynamics on ecological networks. Journal of Theoretical Biology, 247(1):64– 76, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000744.

Ladret:2008:EGD

[LL08] Véronique Ladret and Sabin Lessard. Evolutionary game dynamics in a finite asymmetric two-deme population and emergence of cooperation. Journal of Theoretical Biology, 255(1):137–151, November 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003767.

Lomasko:2009:OHS

[LL09] Tatiana Lomasko and Charles J. Lumsden. One-hit stochastic decline in a mechanochemical model of cytoskeleton-induced neuron death III: Diffusion pulse death zones. Journal of Theoretical Biology, 256(1):104–116, January 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004438.

Leon:2003:TIM

[LLC03] Kalet León, Agustin Lage, and Jorge Carneiro. Tolerance and immunity in a mathematical model of T-cell mediated suppression. Journal of Theoretical Biology, 225(1):107–126, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002261.

Le:2001:LTS

[LLCM01] Shu-Yun Le, Wei-Min Liu, Jih-H. Chen, and Jacob V. Maizel. Local thermodynamic stability scores are well represented by a non-central Student’s t distribution. Journal of Theoretical Biology, 210(4):411–423, June 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923185.

Loeuille:2002:CPH

[LLF02] Nicolas Loeuille, Michel Loreaul, and Réegis Ferrière. Consequences of plant–herbivore coevolution on the dynamics and functioning of ecosystems. Journal of Theoretical Biology, 217(3):369–381, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930328.

Lou:2008:MMP

[LLO08] Chunbo Lou, Zhengyan Li, and Qi Ouyang. A molecular model for persister in E. coli. Journal of Theoretical Biology, 255(2):205–209, November 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004013.

Li:2009:PFC

[LLS^{+}09] Xi Li, Bo Liao, Yu Shu, Qingguang Zeng, and Jiawei Luo. Protein functional class prediction using global encoding of amino acid sequence. Journal of Theoretical Biology, 261(2):290–293, November 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900335X.

Logan:2009:TIF

[LLW09] J. David Logan, Glenn Ledder, and William Wolesensky. Type II functional response for continuous, physiologically structured models. Journal of Theoretical Biology, 259(2):373–381, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900160X.

Lacan:2001:ACC

[LM01a] Jérôme Lacan and Christian J. Michel. Analysis of a circular code model. Journal of Theoretical Biology, 213(2):159–170, November 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924166.

Lewis:2001:EAM

[LM01b] Mark A. Lewis and Paul Moorcroft. ESS analysis of mechanistic models for territoriality: the value of scent marks in spatial resource partitioning. Journal of Theoretical Biology, 210(4):449–461, June 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923239.

Leise:2007:MMD

[LM07] Tanya L. Leise and Emily E. Moin. A mathematical model of the drosophila circadian clock with emphasis on posttranslational mechanisms. Journal of Theoretical Biology, 248(1):48–63, September 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001956.

Lugo:2008:CSE

[LM08] Carlos A. Lugo and Alan J. McKane. The characteristics of species in an evolutionary food web model. Journal of Theoretical Biology, 252(4):649–661, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000933.

Laschov:2009:MML

[LM09] Dmitriy Laschov and Michael Margaliot. Mathematical modeling of the lambda switch: a fuzzy logic approach. Journal of Theoretical Biology, 260(4):475–489, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003014.

Liang:2006:FES

[LME06] Y. Liang, R. M. McMeeking, and A. G. Evans. A finite element simulation scheme for biological muscular hydrostats. Journal of Theoretical Biology, 242(1):142–150, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000749.

Laurence:2003:AHE

[LMF03] Jocelyn A. Laurence, David R. McKenzie, and Kenneth R. Foster. Application of the heat equation to the calculation of temperature rises from pulsed microwave exposure. Journal of Theoretical Biology, 222(3):403–405, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000420.

Leiderman:2008:ESI

[LMF08] Karin M. Leiderman, Laura A. Miller, and Aaron L. Fogelson. The effects of spatial inhomogeneities on flow through the endothelial surface layer. Journal of Theoretical Biology, 252(2):313–325, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800026X.

Lamboy:2004:NMS

[LMH04] Warren F. Lamboy and Gabriel Moreno-Hagelsieb. A new method of solution for the occupancy problem and its application to operon size prediction. Journal of Theoretical Biology, 227(3):315–322, April 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004272.

Letelier:2003:AS

[LMM03] Juan Carlos Letelier, Gonzalo Marín, and Jorge Mpodozis. Autopoietic and (M, R) systems. Journal of Theoretical Biology, 222(2):261–272, May 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000341.

Lu:2008:ICJ

[LMS^{+}08] Z. Lu, R. M. Mitchell, R. L. Smith, J. S. Van Kessel, P. P. Chapagain, Y. H. Schukken, and Y. T. Grohn. The importance of culling in Johne’s disease control. Journal of Theoretical Biology, 254(1):135–146, September 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002427.

Lin:2002:AAE

[LMT02] Kuang Lin, Alex C. W. May, and William R. Taylor. Amino acid encoding schemes from protein structure alignments: Multi-dimensional vectors to describe residue types. Journal of Theoretical Biology, 216(3):361–365, June 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925123.

Locke:2005:MGN

[LMT05] J. C. W. Locke, A. J. Millar, and M. S. Turner. Modelling genetic networks with noisy and varied experimental data: the circadian clock in arabidopsis thaliana. Journal of Theoretical Biology, 234(3):383–393, June 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005879.

Loepfe:2007:RXN

[LMVPM07] Lasse Loepfe, Jordi Martinez-Vilalta, Josep Piñol, and Maurizio Mencuccini. The relevance of xylem network structure for plant hydraulic efficiency and safety. Journal of Theoretical Biology, 247(4):788–803, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001592.

Leisner:2002:TPC

[LN02] S. M. Leisner and D. A. Neher. Third position codon composition suggests two classes of genes within the cauliflower mosaic virus genome. Journal of Theoretical Biology, 217(2):195–201, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930237.

Lampinen:2003:TAI

[LN03] Markku J. Lampinen and Tuula Noponen. Thermodynamic analysis of the interaction of the xylem water and phloem sugar solution and its significance for the cohesion theory. Journal of Theoretical Biology, 224(3):285–298, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001656.

Lampinen:2005:EDT

[LN05a] Markku J. Lampinen and Tuula Noponen. Electric dipole theory and thermodynamics of actomyosin molecular motor in muscle contraction. Journal of Theoretical Biology, 236(4):397–421, October 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001293.

Laptev:2005:SOP

[LN05b] M. V. Laptev and N. K. Nikulin. Synchronization of oscillations of proliferation of keratinocytes in psoriatic skin by external periodic force: a mathematical model. Journal of Theoretical Biology, 235(4):485–494, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500055X.

Langer:2008:SIC

[LNH08] Philipp Langer, Martin A. Nowak, and Christoph Hauert. Spatial invasion of cooperation. Journal of Theoretical Biology, 250(4):634–641, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005498.

LeComber:2006:GPA

[LNRR06] Steven C. Le Comber, Barry Nicholls, D. Kim Rossmo, and Paul A. Racey. Geographic profiling and animal foraging. Journal of Theoretical Biology, 240(2):233–240, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004157.

Lo:2007:SGM

[Lo07] C. F. Lo. Stochastic Gompertz model of tumour cell growth. Journal of Theoretical Biology, 248(2):317–321, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002184.

Locke:2008:CCS

[Loc08] John L. Locke. Cost and complexity: Selection for speech and language. Journal of Theoretical Biology, 251(4):640–652, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000027.

Lorimer:2006:CPR

[Lor06] John W. Lorimer. Curved paths in raptor flight: Deterministic models. Journal of Theoretical Biology, 242(4):880–889, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001366.

Lewis:2000:PCR

[LP00] D. M. Lewis and T. J. Pedley. Planktonic contact rates in homogeneous isotropic turbulence: Theoretical predictions and kinematic simulations. Journal of Theoretical Biology, 205(3):377–408, August 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920733.

Lewis:2001:ITP

[LP01] D. M. Lewis and T. J. Pedley. The influence of turbulence on plankton predation strategies. Journal of Theoretical Biology, 210(3):347–365, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923100.

Lan:2007:ECP

[LP07a] Yueheng Lan and Garegin A. Papoian. Evolution of complex probability distributions in enzyme cascades. Journal of Theoretical Biology, 248(3):537–545, October 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002895.

Llaneras:2007:IAD

[LP07b] F. Llaneras and J. Picó. An interval approach for dealing with flux distributions and elementary modes activity patterns. Journal of Theoretical Biology, 246(2):290–308, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930600600X.

Lv:2007:SBM

[LP07c] Qiming Lv and Jonathan W. Pitchford. Stochastic von Bertalanffy models, with applications to fish recruitment. Journal of Theoretical Biology, 244(4):640–655, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004103.

Livnat:2008:SML

[LP08] Adi Livnat and Nicholas Pippenger. Systematic mistakes are likely in bounded optimal decision-making systems. Journal of Theoretical Biology, 250(3):410–423, February 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004754.

Lipniacki:2004:MMN

[LPB^{+}04] Tomasz Lipniacki, Paweł Paszek, Allan R. Brasier, Bruce Luxon, and Marek Kimmel. Mathematical model of NF-κB regulatory module. Journal of Theoretical Biology, 228(2):195– 215, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000037.

Lee:2006:DPF

[LPC06] S.-H. Lee, H. K. Pak, and T.-S. Chon. Dynamics of preyflock escaping behavior in response to predator’s attack. Journal of Theoretical Biology, 240(2):250–259, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004121.

Li:2008:APV

[LPJB^{+}08] J. Li, C. E. Petit-Jetté, D. Gohore Bi, F. Fenneteau, J. R. E. del Castillo, and F. Nekka. Assessing pharmacokinetic variability directly induced by drug intake behaviour through development of a feeding behaviour-pharmacokinetic model. Journal of Theoretical Biology, 251(3):468–479, April 7, 2008.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006054.

Leon:2000:MCM

[LPLC00] Kalet León, Rolando Peréz, Agustin Lage, and Jorge Carneiro. Modelling T-cell-mediated suppression dependent on interactions in multicellular conjugates. Journal of Theoretical Biology, 207(2):231–254, November 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921696.

Lipniacki:2006:TSG

[LPMC^{+}06] Tomasz Lipniacki, Paweł Paszek, Anna Marciniak-Czochra, Allan R. Brasier, and Marek Kimmel. Transcriptional stochasticity in gene expression. Journal of Theoretical Biology, 238(2):348–367, January 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500250X.

Ludovisi:2005:SED

[LPT05] Alessandro Ludovisi, Piera Pandolfi, and Maria Illuminata Taticchi. The strategy of ecosystem development: specific dissipation as an indicator of ecosystem maturity. Journal of Theoretical Biology, 235(1):33–43, July 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006162.

Lee:2007:CST

[LR07] Deok-Sun Lee and Heiko Rieger. Comparative study of the transcriptional regulatory networks of E. coli and yeast: Structural characteristics leading to marginal dynamic stability. Journal of Theoretical Biology, 248(4):618–626, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307003165.

Lau:2006:EDA

[LRB^{+}06] Kam-Keung Lau, Stephen Roberts, Dora Biro, Robin Freeman, Jessica Meade, and Tim Guilford. An edge-detection approach to investigating pigeon navigation. Journal of Theoretical Biology, 239(1):71–78, March 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003152.

Lehmann:2004:FSR

[LRD04] Jean Lehmann, Bruno Riedo, and Giovanni Dietler. Folding of small RNAs displaying the GNC base-pattern: implications for the self-organization of the genetic system. Journal of Theoretical Biology, 227(3):381–395, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300434X.

Lyles:2009:RLE

[LRHB09] Danielle Lyles, Todd S. Rosenstock, Alan Hastings, and Patrick H. Brown. The role of large environmental noise in masting: General model and example from pistachio trees. Journal of Theoretical Biology, 259(4):701–713, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001817.

Lindsay:2004:NDB

[LRT04] K. A. Lindsay, J. R. Rosenberg, and G. Tucker. A note on the discrepancy between the predicted and observed speed of the propagated action potential in the squid giant axon. Journal of Theoretical Biology, 230(1):39–48, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400181X.

Lai:2003:MMD

[LS03] Yinglei Lai and Fengzhu Sun. Microsatellite mutations during the polymerase chain reaction: mean field approximations and their applications. Journal of Theoretical Biology, 224(1):127–137, September 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001553.

Lai:2004:SDM

[LS04a] Yinglei Lai and Fengzhu Sun. Sampling distribution for microsatellites amplified by PCR: mean field approximation and its applications to genotyping. Journal of Theoretical Biology, 228(2):185–194, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000025.

Laubenbacher:2004:CAA

[LS04b] Reinhard Laubenbacher and Brandilyn Stigler. A computational algebra approach to the reverse engineering of gene regulatory networks. Journal of Theoretical Biology, 229(4):523–537, August 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001754.

Llobera:2007:ZTI

[LS07] M. Llobera and T. J. Sluckin. Zigzagging: Theoretical insights on climbing strategies. Journal of Theoretical Biology, 249(2):206–217, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003542.

Ledzewicz:2008:OSP

[LS08a] Urszula Ledzewicz and Heinz Schättler. Optimal and suboptimal protocols for a class of mathematical models of tumor anti-angiogenesis. Journal of Theoretical Biology, 252(2):295– 312, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000611.

Lee:2008:SST

[LS08b] Sang-Hee Lee and Nan-Yao Su. A simulation study of territory size distribution of mangrove termites on Atlantic coast of Panama. Journal of Theoretical Biology, 253(3):518–523, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002014.

Laird:2009:SCI

[LS09] Robert A. Laird and Brandon S. Schamp. Species coexistence, intransitivity, and topological variation in competitive tournaments. Journal of Theoretical Biology, 256(1):90–95, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004864.

Letelier:2006:OIM

[LSAA^{+}06] Juan-Carlos Letelier, Jorge Soto-Andrade, Flavio Guíñez Abarzúa, Athel Cornish-Bowden, and María Luz Cárdenas. Organizational invariance and metabolic closure: Analysis in terms of (M, R) systems. Journal of Theoretical Biology, 238(4):949–961, February 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003073.

Lesnik:2000:RTC

[LSD^{+}00] T. Lesnik, J. Solomovici, A. Deana, R. Ehrlich, and C. Reiss. Ribosome traffic in E. coli and regulation of gene expression. Journal of Theoretical Biology, 202(2):175–185, January 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910470.

Lopes:2007:IAM

[LSD^{+}07] Christelle Lopes, Thierry Spataro, Christophe Doursat, Laurent Lapchin, and Roger Arditi. An implicit approach to model plant infestation by insect pests. Journal of Theoretical Biology, 248(1):164–178, September 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002445.

Liu:2006:VDG

[LSH06] Shun-Hui Liu, Song Sun, and Bo-Ping Han. Viewing DVM via general behaviors of zooplankton: a way bridging the success of individual and population. Journal of Theoretical Biology, 238(2):435–448, January 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002377.

Loos:2003:SEM

[LSMB^{+}03] Christine Loos, Ralf Seppelt, Sara Meier-Bethke, Joachim Schiemann, and Otto Richter. Spatially explicit modelling of transgenic maize pollen dispersal and cross-pollination. Journal of Theoretical Biology, 225(2):241–255, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002431.

Lopez-Sanchez:2002:MCP

[LSMLB^{+}02] J. López-Sánchez, A. Murciano, R. Lahoz-Beltrá, J. Zamora, N. I. Giménez-Abián, J. F. López-Sáez, C. De La Torre, and J. L. Cánovas. Modelling complex populations formed by proliferating, quiescent and quasi-quiescent cells: Application to plant root meristems. Journal of Theoretical Biology, 215(2):201–213, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301925056.

Linninger:2008:PCE

[LSMZ08] Andreas A. Linninger, Mahadevabharath R. Somayaji, Megan Mekarski, and Libin Zhang. Prediction of convectionenhanced drug delivery to the human brain. Journal of Theoretical Biology, 250(1):125–138, January 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004225.

Lee:2006:HHT

[LSS06] Jack Lee, Bruce Smaill, and Nicolas Smith. Hodgkin–Huxley type ion channel characterization: An improved method of voltage clamp experiment parameter estimation. Journal of Theoretical Biology, 242(1):123–134, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000634.

Lleonart:2000:RAE

[LST00] Jordi Lleonart, Jordi Salat, and Gabriel J. Torres. Removing allometric effects of body size in morphological analysis. Journal of Theoretical Biology, 205(1):85–93, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920435.

Liu:2006:SDT

[LT06] Xianning Liu and Yasuhiro Takeuchi. Spread of disease with transport-related infection and entry screening. Journal of Theoretical Biology, 242(2):517–528, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001330.

Liu:2008:MSR

[LT08] Weijiu Liu and Fusheng Tang. Modeling a simplified regulatory system of blood glucose at molecular levels. Journal of Theoretical Biology, 252(4):608–620, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000829.

Lampert:2009:MAT

[LT09] Adam Lampert and Tsvi Tlusty. Mutability as an altruistic trait in finite asexual populations. Journal of Theoretical Biology, 261(3):414–422, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900407X.

Lemaire:2004:MIB

[LTG^{+}04] Vincent Lemaire, Frank L. Tobin, Larry D. Greller, Carolyn R. Cho, and Larry J. Suva. Modeling the interactions between osteoblast and osteoclast activities in bone remodeling. Journal of Theoretical Biology, 229(3):293–309, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001407.

Liu:2008:SEM

[LTI08] Xianning Liu, Yasuhiro Takeuchi, and Shingo Iwami. SVIR epidemic models with vaccination strategies. Journal of Theoretical Biology, 253(1):1–11, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005115.

Lurette:2009:SAI

[LTLM09] Amandine Lurette, Suzanne Touzeau, Matieyendou Lamboni, and Hervé Monod. Sensitivity analysis to identify key parameters influencing salmonella infection dynamics in a pig batch. Journal of Theoretical Biology, 258(1):43–52, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000307.

Li:2006:DGD

[LTW06] Chun Li, Nannan Tang, and Jun Wang. Directed graphs of DNA sequences and their numerical characterization. Journal of Theoretical Biology, 241(2):173–177, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005060.

Lucas:2005:ATU

[Luc05] M. L. Lucas. Amendments to the theory underlying ussing chamber data of chloride ion secretion after bacterial enterotoxin exposure. Journal of Theoretical Biology, 234(1):21–37, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005259.

Lopez-Villalta:2008:MVD

[LV08] Julián Simón López-Villalta. A metabolic view of the diversitystability relationship. Journal of Theoretical Biology, 252 (1):39–42, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519308000283.

Lopes:2005:KMD

[LVABV05] Francisco J. P. Lopes, Carlos E. Vanario-Alonso, Paulo M. Bisch, and Fernando M. C. Vieira. A kinetic mechanism for drosophila bicoid cooperative binding. Journal of Theoretical Biology, 235(2):185–198, July 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500007X.

Leimar:2004:ACE

[LVH04] Olof Leimar, Tom J. M. Van Dooren, and Peter Hammerstein. Adaptation and constraint in the evolution of environmental sex determination. Journal of Theoretical Biology, 227(4):561– 570, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004508.

Little:2008:ESC

[LVL08a] Mark P. Little, Paolo Vineis, and Guangquan Li. Erratum to “A stochastic carcinogenesis model incorporating multiple types of genomic instability fitted to colon cancer data” [J. Theor. Biol. 254 (2008) 229–238]. Journal of Theoretical Biology, 255(2):268, November 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004530. See [LVL08b].

Little:2008:SCM

[LVL08b] Mark P. Little, Paolo Vineis, and Guangquan Li. A stochastic carcinogenesis model incorporating multiple types of genomic instability fitted to colon cancer data. Journal of Theoretical Biology, 254(2):229–238, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002749. See erratum [LVL08a].

Luciani:2001:SMC

[LVV^{+}01] F. Luciani, S. Valensin, R. Vescovini, P. Sansoni, F. Fagnoni, C. Franceschi, M. Bonafè, and G. Turchetti. A stochastic model for CD8 + T cell dynamics in human immunosenescence: Implications for survival and longevity. Journal of Theoretical Biology, 213(4):587–597, December 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924336.

Lorberbaum:2004:ETR

[LW04] Jeffrey P. Lorberbaum and W. Reynolds Williams. Erratum to “Are there really no evolutionarily stable strategies in the iterated prisoner’s dilemma?”: [J. Theor. Biol. 214, 155–169]. Journal of Theoretical Biology, 227(4):593, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000347. See [LBSS02].

Logan:2007:IME

[LW07] J. David Logan and William Wolesensky. An index to measure the effects of temperature change on trophic interactions. Journal of Theoretical Biology, 246(2):366–376, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005571.

Lichtwark:2008:OMF

[LW08a] G. A. Lichtwark and A. M. Wilson. Optimal muscle fascicle length and tendon stiffness for maximising gastrocnemius efficiency during human walking and running. Journal of Theoretical Biology, 252(4):662–673, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000325.

Liu:2008:CTS

[LW08b] Liwei Liu and Tianming Wang. Comparison of TOPS strings based on LZ complexity. Journal of Theoretical Biology, 251(1):159–166, March 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005759.

Li:2009:PSC

[LWC09] Zuxiong Li, Tieying Wang, and Lansun Chen. Periodic solution of a chemostat model with Beddington-DeAnglis uptake function and impulsive state feedback control. Journal of Theoretical Biology, 261(1):23–32, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003348.

Lee:2008:ERD

[LWFP08] DongYoub Lee, Anthony S. Wexler, Michelle V. Fanucchi, and Charles G. Plopper. Expiration rate drives human airway design. Journal of Theoretical Biology, 253(2):381–387, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001276.

Lemon:2007:MMH

[LWRK07] Greg Lemon, Sarah L. Waters, Felicity R. A. J. Rose, and John R. King. Mathematical modelling of human mesenchymal stem cell proliferation and differentiation inside artificial porous scaffolds. Journal of Theoretical Biology, 249(3):543–553, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003992.

Lev-Yadun:2000:WGW

[LY00] Simcha Lev-Yadun. Whirled grain in wood and topological defects. Journal of Theoretical Biology, 205(3):511–514, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920915.

Lev-Yadun:2001:AWC

[LY01] Simcha Lev-Yadun. Aposematic (warning) coloration associated with thorns in higher plants. Journal of Theoretical Biology, 210(3):385–388, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930192315X.

Li:2002:ZLI

[LY02] Wentian Li and Yaning Yang. Zipf’s law in importance of genes for cancer classification using microarray data. Journal of Theoretical Biology, 219(4):539–551, December 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302931450.

Lev-Yadun:2003:WTA

[LY03a] Simcha Lev-Yadun. Weapon (thorn) automimicry and mimicry of aposematic colorful thorns in plants. Journal of Theoretical Biology, 224(2):183–188, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001565.

Lev-Yadun:2003:WDS

[LY03b] Simcha Lev-Yadun. Why do some thorny plants resemble green zebras? Journal of Theoretical Biology, 224(4):483–489, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001966.

Liu:2008:GBA

[LYZ08] Helong Liu, Jingyuan Yu, and Guangtian Zhu. Global behaviour of an age-infection-structured HIV model with impulsive drug-treatment strategy. Journal of Theoretical Biology, 253(4):749–754, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002051.

Liu:2009:DSF

[LZ09a] Xin Liu and Ya-Pu Zhao. Donut-shaped fingerprint in homologous polypeptide relationships — a topological feature related to pathogenic structural changes in conformational disease. Journal of Theoretical Biology, 258(2):294–301, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000708.

Liu:2009:SMS

[LZ09b] Xin Liu and Ya-Pu Zhao. A scheme for multiple sequence alignment optimization — an improvement based on family representative mechanics features. Journal of Theoretical Biology, 261(4):593–597, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004081.

Long:2003:PTO

[LZGL03] Charles A. Long, G. P. Zhang, Thomas F. George, and Claudine F. Long. Physical theory, origin of flight, and a synthesis proposed for birds. Journal of Theoretical Biology, 224(1):9–26, September 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001164.

Lu:2008:RMA

[LZS^{+}08] Ming Lu, Lufang Zhou, William C. Stanley, Marco E. Cabrera, Gerald M. Saidel, and Xin Yu. Role of the malate-aspartate shuttle on the metabolic response to myocardial ischemia. Journal of Theoretical Biology, 254(2):466–475, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002804.

Millar:2003:KMP

[MA03a] Russell B. Millar and Marti J. Anderson. The kinetics of monospermic and polyspermic fertilization in freespawning marine invertebrates. Journal of Theoretical Biology, 224(1):79–85, September 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001450.

Minami:2003:VFM

[MA03b] Shizuka Minami and Akira Azuma. Various flying modes of wind-dispersal seeds. Journal of Theoretical Biology, 225(1): 1–14, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303002169.

Morishita:2004:NRT

[MA04] Yoshihiro Morishita and Kazuyuki Aihara. Noise-reduction through interaction in gene expression and biochemical reaction processes. Journal of Theoretical Biology, 228(3):315–325, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000360.

Miftahof:2007:DIP

[MA07a] R. Miftahof and N. Akhmadeev. Dynamics of intestinal propulsion. Journal of Theoretical Biology, 246(2):377–393, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000239.

Miftahof:2007:NBV

[MA07b] R. Miftahof and N. R. Akhmadeev. Neurochemical bases of visceral nociception: Mathematical model. Journal of Theoretical Biology, 249(2):343–360, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003633.

Mendoza:2000:GRR

[MAB00] Luis Mendoza and Elena R. Alvarez-Buylla. Genetic regulation of root hair development in arabidopsis thaliana: a network model. Journal of Theoretical Biology, 204(3):311–326, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920149.

Macfadyen:2000:RCD

[Mac00] Leah P. Macfadyen. Regulation of competence development in haemophilus influenzae: Proposed competence regulatory elements are CRP-binding sites. Journal of Theoretical Biology, 207(3):349–359, December 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921799.

Mackintosh:2001:APM

[Mac01] James A. Mackintosh. The antimicrobial properties of melanocytes, melanosomes and melanin and the evolution of black skin. Journal of Theoretical Biology, 211(2):101–113, July 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923318.

McDougall:2006:MMD

[MAC06] Steven R. McDougall, Alexander R. A. Anderson, and Mark A. J. Chaplain. Mathematical modelling of dynamic adaptive tumour-induced angiogenesis: Clinical implications and therapeutic targeting strategies. Journal of Theoretical Biology, 241(3):564–589, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005564.

Maddison:2000:TCC

[Mad00] Wayne P. Maddison. Testing character correlation using pairwise comparisons on a phylogeny. Journal of Theoretical Biology, 202(3):195–204, February 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910500.

Maganaris:2004:PMM

[Mag04] Constantinos N. Maganaris. A predictive model of momentangle characteristics in human skeletal muscle: Application and validation in muscles across the ankle joint. Journal of Theoretical Biology, 230(1):89–98, September 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001870.

Millor:2006:IDC

[MAHD06] J. Millor, J. M. Amé, J. Halloy, and J. L. Deneubourg. Individual discrimination capability and collective decisionmaking. Journal of Theoretical Biology, 239(3):313–323, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003310.

Malchow:2000:MIP

[Mal00] Horst Malchow. Motional instabilities in prey–predator systems. Journal of Theoretical Biology, 204(4):639–647, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920745.

Maly:2002:SMP

[Mal02] Ivan V. Maly. A stochastic model for patterning of the cytoplasm by the saltatory movement. Journal of Theoretical Biology, 216(1):59–71, May 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302925312.

Manrubia:2003:SBP

[MAL03] Susanna C. Manrubia, María Arribas, and Ester Lázaro. Supercritical branching processes and the role of fluctuations under exponential population growth. Journal of Theoretical Biology, 225(4):497–505, December 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002947.

Mangel:2001:CAS

[Man01] Marc Mangel. Complex adaptive systems, aging and longevity. Journal of Theoretical Biology, 213(4):559–571, December 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924312.

Mantzaris:2006:SDS

[Man06] Nikos V. Mantzaris. Stochastic and deterministic simulations of heterogeneous cell population dynamics. Journal of Theoretical Biology, 241(3):690–706, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000075.

Martin:2003:FDR

[Mar03] R. Bruce Martin. Fatigue damage, remodeling, and the minimization of skeletal weight. Journal of Theoretical Biology, 220(2):271–276, January 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931481.

Martin:2004:MCS

[Mar04] A. P. Martin. A Malthusian curb on spatial structure in microorganism populations. Journal of Theoretical Biology, 230(3):343–349, October 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002619.

Marshall:2009:DGR

[Mar09a] James A. R. Marshall. The donation game with roles played between relatives. Journal of Theoretical Biology, 260(3):386–391, October 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003178.

Marshall:2009:EGD

[Mar09b] John M. Marshall. The effect of gene drive on containment of transgenic mosquitoes. Journal of Theoretical Biology, 258(2):250–265, May 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000496.

Masuda:2003:EMM

[Mas03] Tadashi Masuda. An electromechanical model of myosin molecular motors. Journal of Theoretical Biology, 225(4): 507–515, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303003023.

Masuda:2008:ODE

[Mas08] Naoki Masuda. Oscillatory dynamics in evolutionary games are suppressed by heterogeneous adaptation rates of players. Journal of Theoretical Biology, 251(1):181–189, March 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005693.

Masuda:2009:DCN

[Mas09] Naoki Masuda. Directionality of contact networks suppresses selection pressure in evolutionary dynamics. Journal of Theoretical Biology, 258(2):323–334, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000459.

Matthaus:2006:DVN

[Mat06] Franziska Matthäus. Diffusion versus network models as descriptions for the spread of prion diseases in the brain. Journal of Theoretical Biology, 240(1):104–113, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003711.

Maurin:2002:MMC

[Mau02] Francis Maurin. A mathematical model of chemoreception for odours and taste. Journal of Theoretical Biology, 215(3):297– 303, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925147.

Mazat:2008:FTM

[Maz08] Jean-Pierre Mazat. The first time I met Reinhart Heinrich …. Journal of Theoretical Biology, 252(3):568, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006352.

Mcleod:2000:DDM

[MB00] Barry Mcleod and Nigel Burroughs. Defective deletion mutant amplification. Journal of Theoretical Biology, 206(4):449–464, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921283.

McKenzie:2005:IMP

[MB05] F. Ellis McKenzie and William H. Bossert. An integrated model of Plasmodium falciparum dynamics. Journal of Theoretical Biology, 232(3):411–426, February 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004035.

McGraw:2006:IBM

[MB06] James E. McGraw and John F. Y. Brookfield. The interaction between mobile DNAs and their hosts in a fluctuating environment. Journal of Theoretical Biology, 243(1):13–23, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002256.

Meysman:2007:TPF

[MB07] Filip J. R. Meysman and Stijn Bruers. A thermodynamic perspective on food webs: Quantifying entropy production within detrital-based ecosystems. Journal of Theoretical Biology, 249(1):124–139, November 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003414.

Marakushev:2009:PTA

[MB09] Sergey A. Marakushev and Ol’ga V. Belonogova. The parageneses thermodynamic analysis of chemoautotrophic CO 2 fixation archaic cycle components, their stability and selforganization in hydrothermal systems. Journal of Theoretical Biology, 257(4):588–597, April 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006061.

Monteiro:2002:ARW

[MBB02] L. H. A. Monteiro, M. A. Bussab, and J. G. Chaui Berlinck. Analytical results on a Wilson–Cowan neuronal network modified model. Journal of Theoretical Biology, 219(1):83–91, November 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931115.

Meekangvan:2006:NDM

[MBB^{+}06] Preeda Meekangvan, Alan A. Barhorst, Thomas D. Burton, Sankar Chatterjee, and Lawrence Schovanec. Nonlinear dynamical model and response of avian cranial kinesis. Journal of Theoretical Biology, 240(1):32–47, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003668.

Menconi:2008:DCG

[MBB08] Giulia Menconi, Vieri Benci, and Marcello Buiatti. Data compression and genomes: a two-dimensional life domain map. Journal of Theoretical Biology, 253(2):281–288, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001252.

Mondal:2006:PAA

[MBBR06] Sukanta Mondal, Rajasekaran Bhavna, Rajasekaran Mohan Babu, and Suryanarayanarao Ramakumar. Pseudo amino acid composition and multi-class support vector machines approach for conotoxin superfamily classification. Journal of Theoretical Biology, 243(2):252–260, November 21, 2006.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600261X.

Mitchell:2008:CPP

[MBD08] G. Mitchell, J. P. Bobbitt, and S. Devries. Cerebral perfusion pressure in giraffe: Modelling the effects of head-raising and -lowering. Journal of Theoretical Biology, 252(1):98–108, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000313.

Manning:2003:MHS

[MBH03] John T. Manning, Peter E. Bundred, and P. Henzi. Melanin and HIV in sub-Saharan Africa. Journal of Theoretical Biology, 223(1):131–133, July 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000705.

Munkemuller:2009:HRP

[MBJ09] Tamara Münkemüller, Harald Bugmann, and Karin Johst. Hutchinson revisited: Patterns of density regulation and the coexistence of strong competitors. Journal of Theoretical Biology, 259(1):109–117, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001209.

Marashi:2007:APD

[MBP07] Sayed-Amir Marashi, Reza Behrouzi, and Hamid Pezeshk. Adaptation of proteins to different environments: a comparison of proteome structural properties in bacillus subtilis and Escherichia coli. Journal of Theoretical Biology, 244(1):127–132, January 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003262.

Muneepeerakul:2008:PVB

[MBRRI08] Rachata Muneepeerakul, Enrico Bertuzzo, Andrea Rinaldo, and Ignacio Rodriguez-Iturbe. Patterns of vegetation biodiversity: The roles of dispersal directionality and river network structure. Journal of Theoretical Biology, 252(2):221– 229, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000490.

Mogie:2001:SSU

[MC01] Michael Mogie and Moya Cousins. Are sample sizes usually at least an order of magnitude too low for reliable estimates of leaf asymmetry? Journal of Theoretical Biology, 211(2):181– 185, July 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923380.

Meng:2006:APS

[MC06] Xinzhu Meng and Lansun Chen. Almost periodic solution of non-autonomous Lotka–Volterra predator-prey dispersal system with delays. Journal of Theoretical Biology, 243(4):562–574, December 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002979.

Mitrophanov:2007:CSP

[MCB07] Alexander Y. Mitrophanov, Gordon Churchward, and Mark Borodovsky. Control of Streptococcus pyogenes virulence: Modeling of the CovR/S signal transduction system. Journal of Theoretical Biology, 246(1):113–128, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005510.

Marco:2009:EME

[MCC^{+}09] Diana E. Marco, Juan P. Carbajal, Sergio Cannas, Rebeca Pérez-Arnedo, Ángeles Hidalgo-Perea, José Olivares, José E. Ruiz-Sainz, and Juan Sanjuán. An experimental and modelling exploration of the host-sanction hypothesis in legume-rhizobia mutualism. Journal of Theoretical Biology, 259(3):423–433, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001428.

Mcelreath:2003:REC

[Mce03] Richard Mcelreath. Reputation and the evolution of conflict. Journal of Theoretical Biology, 220(3):345–357, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931663.

Maia:2004:EST

[MCF04] Leonardo P. Maia, Alexandre Colato, and José F. Fontanari. Effect of selection on the topology of genealogical trees. Journal of Theoretical Biology, 226(3):315–320, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003394.

Marciniak-Czochra:2007:RDA

[MCK07] Anna Marciniak-Czochra and Marek Kimmel. Reactiondiffusion approach to modeling of the spread of early tumors along linear or tubular structures. Journal of Theoretical Biology, 244(3):375–387, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003924.

Marco:2009:CED

[MCM^{+}09] Diana E. Marco, Sergio A. Cannas, Marcelo A. Montemurro, Bo Hu, and Shi-Yuan Cheng. Comparable ecological dynamics underlie early cancer invasion and species dispersal, involving self-organizing processes. Journal of Theoretical Biology, 256(1):65–75, January 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004888.

Mendoza-Cozatl:2006:CGP

[MCMS06] David G. Mendoza-Cózatl and Rafael Moreno-Sánchez. Control of glutathione and phytochelatin synthesis under cadmium stress. Pathway modeling for plants. Journal of Theoretical Biology, 238(4):919–936, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500305X.

Mcnair:2000:TTS

[Mcn00] James N. Mcnair. Turbulent transport of suspended particles and dispersing benthic organisms: The hitting-time distribution for the local exchange model. Journal of Theoretical Biology, 202(3):231–246, February 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910573.

McNair:2006:PSL

[McN06] James N. McNair. Probabilistic settling in the local exchange model of turbulent particle transport. Journal of Theoretical Biology, 241(2):420–437, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500531X.

Mura:2008:SPN

[MCN08] Ivan Mura and Attila Csikász-Nagy. Stochastic Petri net extension of a yeast cell cycle model. Journal of Theoretical Biology, 254(4):850–860, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800372X.

Mittenthal:2001:NMA

[MCWF01] Jay E. Mittenthal, Bertrand Clarke, Thomas G. Waddell, and Glenn Fawcett. A new method for assembling metabolic networks, with application to the Krebs citric acid cycle. Journal of Theoretical Biology, 208(3):361–382, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922252.

Mylius:2001:RSB

[MD01] Sido D. Mylius and Odo Diekmann. The resident strikes back: Invader-induced switching of resident attractor. Journal of Theoretical Biology, 211(4):297–311, August 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923495.

Mansury:2003:ISP

[MD03] Yuri Mansury and Thomas S. Deisboeck. The impact of “search precision” in an agent-based tumor model. Journal of Theoretical Biology, 224(3):325–337, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001693.

Ma:2004:KAA

[MD04] Huilian Ma and Richard B. Dickinson. Kinetic analysis of the attachment of a biological particle to a surface by macromolecular binding. Journal of Theoretical Biology, 226(2):237–250, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003369.

Mallet:2006:CAM

[MD06] D. G. Mallet and L. G. De Pillis. A cellular automata model of tumor-immune system interactions. Journal of Theoretical Biology, 239(3):334–350, April 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003334.

Mohanty:2008:MBM

[MD08] Utkala Mohanty and Narendra M. Dixit. Mechanism-based model of the pharmacokinetics of enfuvirtide, an HIV fusion inhibitor. Journal of Theoretical Biology, 251(3):541–551, April 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307006546.

Magal:2006:HCC

[MDCC06] Christelle Magal, Olivier Dangles, Philippe Caparroy, and Jérôme Casas. Hair canopy of cricket sensory system tuned to predator signals. Journal of Theoretical Biology, 241(3):459–466, August 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305005412.

Mansury:2006:EGT

[MDD06] Yuri Mansury, Mark Diggory, and Thomas S. Deisboeck. Evolutionary game theory in an agent-based brain tumor model: Exploring the ‘genotype-phenotype’ link. Journal of Theoretical Biology, 238(1):146–156, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002456.

Murrell:2004:MCP

[MDL04] David J. Murrell, Ulf Dieckmann, and Richard Law. On moment closures for population dynamics in continuous space. Journal of Theoretical Biology, 229(3):421–432, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001523.

Meister:2005:CBL

[Mei05] Armin Meister. Calculation of binding length of base-specific DNA dyes by comparison of sequence and flow cytometric data. Application to Oryza sativa and Arabidopsis thaliana. Journal of Theoretical Biology, 232(1):93–97, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003534.

Mengel:2007:EFV

[Men07] Friederike Mengel. The evolution of function-valued traits for conditional cooperation. Journal of Theoretical Biology, 245(3):564–575, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519306005285.

Miftahof:2005:IPS

[MF05] R. Miftahof and E. Fedotov. Intestinal propulsion of a solid non-deformable bolus. Journal of Theoretical Biology, 235(1):57–70, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519304006186.

Mogilner:2001:SCN

[MFB01] A. Mogilner, A. J. Fisher, and R. J. Baskin. Structural changes in the neck linker of kinesin explain the load dependence of the Motor’s mechanical cycle. Journal of Theoretical Biology, 211(2):143–157, July 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923367.

Mei:2009:LNA

[MFI09] Zhi-Ping Mei, Zoe V. Finkel, and Andrew J. Irwin. Light and nutrient availability affect the size-scaling of growth in phytoplankton. Journal of Theoretical Biology, 259(3):582–588, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001842.

Michor:2003:SSA

[MFM^{+}03] Franziska Michor, Steven A. Frank, Robert M. May, Yoh Iwasa, and Martin A. Nowak. Somatic selection for and against cancer. Journal of Theoretical Biology, 225(3):377–382, December 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002674.

Monro:2009:MCR

[MG09] Helen C. Monro and Eamonn A. Gaffney. Modelling chemotherapy resistance in palliation and failed cure. Journal of Theoretical Biology, 257(2):292–302, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006334.

Moreo:2009:BISa

[MGAD09a] Pedro Moreo, José Manuel García-Aznar, and Manuel Doblaré. Bone ingrowth on the surface of endosseous implants. Part 1: Mathematical model. Journal of Theoretical Biology, 260(1):1–12, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004025.

Moreo:2009:BISb

[MGAD09b] Pedro Moreo, José Manuel García-Aznar, and Manuel Doblaré. Bone ingrowth on the surface of endosseous implants. Part 2: Theoretical and numerical analysis. Journal of Theoretical Biology, 260(1):13–26, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900246X.

Martelli:2007:AST

[MGAR07] Carlotta Martelli, Andrea Giansanti, Ivan Arisi, and Vittorio Rosato. Asymptotic states and topological structure of an activation-deactivation chemical network. Journal of Theoretical Biology, 245(3):423–432, April 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004711.

Mendoza:2004:ECS

[MGC04] Manuel Mendoza, Brian Goodwin, and Carlos Criado. Emergence of community structure in terrestrial mammaldominated ecosystems. Journal of Theoretical Biology, 230 (2):203–214, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930400205X.

Munteanu:2008:NRP

[MGDBdM08] Cristian Robert Munteanu, Humberto González-Díaz, Fernanda Borges, and Alexandre Lopes de Magalhães. Natural/random protein classification models based on star network topological indices. Journal of Theoretical Biology, 254(4):775–783, October 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800369X.

McDonald-Gibson:2008:ERC

[MGDDH08] J. McDonald-Gibson, J. G. Dyke, E. A. Di Paolo, and I. R. Harvey. Environmental regulation can arise under minimal assumptions. Journal of Theoretical Biology, 251(4):653–666, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307006534.

Munteanu:2008:ENE

[MGDM08] Cristian Robert Munteanu, Humberto González-Díaz, and Alexandre L. Magalhães. Enzymes/non-enzymes classification model complexity based on composition, sequence, 3d and topological indices. Journal of Theoretical Biology, 254(2):476–482, September 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003020.

Mesterton-Gibbons:2006:EDS

[MGHF06] Mike Mesterton-Gibbons, Ian C. W. Hardy, and Jeremy Field. The effect of differential survivorship on the stability of reproductive queueing. Journal of Theoretical Biology, 242(3):699–712, October 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002001.

Makarieva:2003:NMR

[MGL03] Anastassia M. Makarieva, Victor G. Gorshkov, and Bai-Lian Li. A note on metabolic rate dependence on body size in plants and animals. Journal of Theoretical Biology, 221(2):301–307, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931857.

Makarieva:2005:RDN

[MGL05] Anastassia M. Makarieva, Victor G. Gorshkov, and Bai-Lian Li. Revising the distributive networks models of West, Brown and Enquist (1997) and Banavar, Maritan and Rinaldo (1999): Metabolic inequity of living tissues provides clues for the observed allometric scaling rules. Journal of Theoretical Biology, 237(3):291–301, December 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001803.

Makarieva:2006:DNM

[MGL06a] Anastassia M. Makarieva, Victor G. Gorshkov, and Bai-Lian Li. Distributive network model of Banavar, Damuth, Maritan and Rinaldo (2002): Critique and perspective. Journal of Theoretical Biology, 239(3):394–397, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003589.

Medvinsky:2006:IPR

[MGL^{+}06b] Alexander B. Medvinsky, Maria M. Gonik, Bai-Lian Li, Vassili V. Velkov, and Horst Malchow. Invasion of pests resistant to Bt toxins can lead to inherent non-uniqueness in genetically modified bt-plant dynamics: Mathematical modeling. Journal of Theoretical Biology, 242(3):539–546, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001639.

Maser:2007:WTI

[MGM07] Gabriel L. Maser, Frédéric Guichard, and Kevin S. McCann. Weak trophic interactions and the balance of enriched metacommunities. Journal of Theoretical Biology, 247(2):337–345, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001063.

Monteiro:2000:CSE

[MGP00] L. H. A. Monteiro, C. H. O. Gonçalves, and J. R. C. Piqueira. A condition for successful escape of a mutant after primary HIV infection. Journal of Theoretical Biology, 203(4):399–406, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910920.

Marhl:2008:STM

[MGP^{+}08] Marko Marhl, Marko Gosak, Matjaz Perc, C. Jane Dixon, and Anne K. Green. Spatio-temporal modelling explains the effect of reduced plasma membrane Ca^{2+} efflux on intracellular Ca^{2+} oscillations in hepatocytes. Journal of Theoretical Biology, 252(3):419–426, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005656.

Matsuoka:2008:SCE

[MGS08] Tomohiro Matsuoka, Sohei Gomi, and Ryuzo Shingai. Simulation of C. elegans thermotactic behavior in a linear thermal gradient using a simple phenomenological motility model. Journal of Theoretical Biology, 250(2):230–243, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004791.

Mesterton-Gibbons:2009:NIG

[MGS09] Mike Mesterton-Gibbons and Tom N. Sherratt. Neighbor intervention: a game-theoretic model. Journal of Theoretical Biology, 256(2):263–275, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005195.

Moore:2006:FCF

[MGT^{+}06] Jason H. Moore, Joshua C. Gilbert, Chia-Ti Tsai, Fu-Tien Chiang, Todd Holden, Nate Barney, and Bill C. White. A flexible computational framework for detecting, characterizing, and interpreting statistical patterns of epistasis in genetic studies of human disease susceptibility. Journal of Theoretical Biology, 241(2):252–261, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005217.

Meyer-Hermann:2002:MMG

[MH02] Michael Meyer-Hermann. A mathematical model for the germinal center morphology and affinity maturation. Journal of Theoretical Biology, 216(3):273–300, June 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925506.

McNamara:2005:IAK

[MH05] John M. McNamara and Alasdair I. Houston. If animals know their own fighting ability, the evolutionarily stable level of fighting is reduced. Journal of Theoretical Biology, 232(1):1–6, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003431.

Malygin:2006:PAC

[MH06a] Ernst G. Malygin and Stanley Hattman. A probabilistic approach to compact steady-state kinetic equations for enzymic reactions. Journal of Theoretical Biology, 242(3):627–633, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001354.

Matsumura:2006:WSS

[MH06b] Shuichi Matsumura and Thomas J. Hayden. When should signals of submission be given? — A game theory model. Journal of Theoretical Biology, 240(3):425–433, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004406.

Meyer-Hermann:2007:COB

[MH07] Michael E. Meyer-Hermann. Are T cells at the origin of B cell lymphomas? Journal of Theoretical Biology, 244(4):656–669, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004073.

Melendez-Hevia:2008:MRH

[MH08] Enrique Meléndez-Hevia. Memories of Reinhart Heinrich. Journal of Theoretical Biology, 252(3):544–545, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003360.

Marin:2009:BND

[MH09a] Ignacio Marín and Sergio Hoyas. Basic networks: Definition and applications. Journal of Theoretical Biology, 258(1):53– 59, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000319.

Mikaberidze:2009:SBM

[MH09b] Alexey Mikaberidze and Masudul Haque. Survival benefits in mimicry: a quantitative framework. Journal of Theoretical Biology, 259(3):462–468, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000952.

Meyer-Hermann:2001:RPD

[MHDOG01] Michael Meyer-Hermann, Andreas Deutsch, and Michal Or-Guil. Recycling probability and dynamical properties of germinal center reactions. Journal of Theoretical Biology, 210(3):265–285, June 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922970.

Merks:2003:MCG

[MHKS03] Roeland Merks, Alfons Hoekstra, Jaap Kaandorp, and Peter Sloot. Models of coral growth: spontaneous branching, compactification and the Laplacian growth assumption. Journal of Theoretical Biology, 224(2):153–166, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001401.

Merks:2004:POM

[MHKS04] Roeland M. H. Merks, Alfons G. Hoekstra, Jaap A. Kaandorp, and Peter M. A. Sloot. Polyp oriented modelling of coral growth. Journal of Theoretical Biology, 228(4):559–576, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000761.

Melendez-Hevia:2008:PCC

[MHMGM08] Enrique Meléndez-Hevia, Nancy Montero-Gómez, and Francisco Montero. From prebiotic chemistry to cellular metabolism — the chemical evolution of metabolism before Darwinian natural selection. Journal of Theoretical Biology, 252(3):505–519, June 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700570X.

Marino:2008:MPG

[MHRK08] Simeone Marino, Ian B. Hogue, Christian J. Ray, and Denise E. Kirschner. A methodology for performing global uncertainty and sensitivity analysis in systems biology. Journal of Theoretical Biology, 254(1):178–196, September 7, 2008.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001896.

Michor:2006:DMS

[MI06] Franziska Michor and Yoh Iwasa. Dynamics of metastasis suppressor gene inactivation. Journal of Theoretical Biology, 241(3):676–689, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000051.

Muko:2008:ESH

[MI08a] Soyoka Muko and Yoh Iwasa. Erratum to “Spatial heterogeneity of mortality and temporal fluctuation in fertility promote coexistence but not vice versa: a random-community approach” [J. Theor. Biol. 253 (2008) 593–600]. Journal of Theoretical Biology, 254(4):868, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003974. See [MI08b].

Muko:2008:SHM

[MI08b] Soyoka Muko and Yoh Iwasa. Spatial heterogeneity of mortality and temporal fluctuation in fertility promote coexistence but not vice versa: a random-community approach. Journal of Theoretical Biology, 253(3):593–600, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001859. See erratum [MI08a].

Michaelian:2005:TSE

[Mic05] K. Michaelian. Thermodynamic stability of ecosystems. Journal of Theoretical Biology, 237(3):323–335, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001839.

Michel:2007:EPP

[Mic07] Christian J. Michel. Evolution probabilities and phylogenetic distance of dinucleotides. Journal of Theoretical Biology, 249(2):271–277, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003682.

Miekisz:2005:ESE

[Mie05] Jacek Mieķisz. Equilibrium selection in evolutionary games with random matching of players. Journal of Theoretical Biology, 232(1):47–53, January 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003492.

Miguel:2006:CPF

[Mig06] António F. Miguel. Constructal pattern formation in stony corals, bacterial colonies and plant roots under different hydrodynamics conditions. Journal of Theoretical Biology, 242(4):954–961, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001998.

Miller:2005:SDR

[Mil05] Laura A. Miller. Structural dynamics and resonance in plants with nonlinear stiffness. Journal of Theoretical Biology, 234(4):511–524, June 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304006083.

Milne:2008:NDS

[Mil08] Eugene M. G. Milne. The natural distribution of survival. Journal of Theoretical Biology, 255(2):223–236, November 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003755.

Milner:2009:E

[Mil09] Fabio A. Milner. Editorial. Journal of Theoretical Biology, 258(3):337–338, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309001696.

Mittwoch:2004:EAS

[Mit04] Ursula Mittwoch. The elusive action of sex-determining genes: mitochondria to the rescue? Journal of Theoretical Biology, 228(3):359–365, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000402.

Mullowney:2007:RVC

[MJ07a] Paul Mullowney and Alex James. The role of variance in capped-rate stochastic growth models with external mortality. Journal of Theoretical Biology, 244(2):228–238, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306003274.

Munkemuller:2007:HDI

[MJ07b] Tamara Münkemüller and Karin Johst. How does intraspecific density regulation influence metapopulation synchrony and persistence? Journal of Theoretical Biology, 245(3):553– 563, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600511X.

Munro:2000:CNS

[MJW00] Peter D. Munro, Craig M. Jackson, and Donald J. Winzor. Consequences of the non-specific binding of a protein to a linear polymer: Reconciliation of stoichiometric and equilibrium titration data for the thrombin–heparin interaction. Journal of Theoretical Biology, 203(4):407–418, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300910993.

Magalhaes:2001:SPG

[MK01] João Carlos Marques Magalhães and Décio Krause. Suppes predicate for genetics and natural selection. Journal of Theoretical Biology, 209(2):141–153, March 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922434.

Muller:2003:VSC

[MK03] H.-M. Müller and S. E. Koonin. Vector space classification of DNA sequences. Journal of Theoretical Biology, 223(2):161– 169, July 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000821.

Marino:2004:HIR

[MK04] Simeone Marino and Denise E. Kirschner. The human immune response to mycobacterium tuberculosis in lung and lymph node. Journal of Theoretical Biology, 227(4):463–486, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004429.

Masuda:2006:MSE

[MK06] Naoki Masuda and Norio Konno. Multi-state epidemic processes on complex networks. Journal of Theoretical Biology, 243(1):64–75, November 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002438.

Mizuno:2009:EES

[MK09] Akiko N. Mizuno and Masakado Kawata. The effects of the evolution of stoichiometry-related traits on population dynamics in plankton communities. Journal of Theoretical Biology, 259(2):209–218, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000939.

Morishita:2005:EPM

[MKA05] Yoshihiro Morishita, Tetsuya J. Kobayashi, and Kazuyuki Aihara. Evaluation of the performance of mechanisms for noise attenuation in a single-gene expression. Journal of Theoretical Biology, 235(2):241–264, July 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000238.

Munk:2003:CWC

[MKB03] Ole Lajord Munk, Susanne Keiding, and Ludvik Bass. Capillaries within compartments: microvascular interpretation of dynamic positron emission tomography data. Journal of Theoretical Biology, 225(1):127–141, November 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002273.

Maree:2005:QMD

[MKD^{+}05] Athanasius F. M. Marée, Mitsuhiro Komba, Cheryl Dyck, Marek Łabecki, Diane T. Finegood, and Leah Edelstein-Keshet. Quantifying macrophage defects in type 1 diabetes. Journal of Theoretical Biology, 233(4):533–551, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005181.

Muller:2009:DEB

[MKE^{+}09] Erik B. Muller, Sebastiaan A. L. M. Kooijman, Peter J. Edmunds, Francis J. Doyle, and Roger M. Nisbet. Dynamic energy budgets in syntrophic symbiotic relationships between heterotrophic hosts and photoautotrophic symbionts. Journal of Theoretical Biology, 259(1):44–57, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900109X.

Middleton:2009:BMM

[MKL09] A. M. Middleton, J. R. King, and M. Loose. Bistability in a model of mesoderm and anterior mesendoderm specification in xenopus laevis. Journal of Theoretical Biology, 260(1):41–55, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002380.

Mansury:2002:EPT

[MKLD02] Yuri Mansury, Mark Kimura, Jose Lobo, and Thomas S. Deisboeck. Emerging patterns in tumor systems: Simulating the dynamics of multicellular clusters with an agentbased spatial agglomeration model. Journal of Theoretical Biology, 219(3):343–370, December 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931310.

Matsuura:2002:HTR

[MKN02] Kenji Matsuura, Eizi Kuno, and Takayoshi Nishida. Homosexual tandem running as selfish herd in reticulitermes speratus: Novel antipredatory behavior in termites. Journal of Theoretical Biology, 214(1):63–70, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924476.

McCauley:2009:NMM

[MKS^{+}09] Peter McCauley, Leonid V. Kalachev, Amber D. Smith, Gregory Belenky, David F. Dinges, and Hans P. A. Van Dongen. A new mathematical model for the homeostatic effects of sleep loss on neurobehavioral performance. Journal of Theoretical Biology, 256(2):227–239, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004761.

Ma:2000:TSE

[MKT^{+}00] Buyong Ma, Sandeep Kumar, Chung-Jung Tsai, Zengjian Hu, and Ruth Nussinov. Transition-state ensemble in enzyme catalysis: Possibility, reality, or necessity? Journal of Theoretical Biology, 203(4):383–397, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930091097X.

Moore:2004:MMC

[ML04] Helen Moore and Natasha K. Li. A mathematical model for chronic myelogenous leukemia (CML) and T cell interaction. Journal of Theoretical Biology, 227(4):513–523, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004454.

Macklin:2007:ENS

[ML07a] Paul Macklin and John Lowengrub. Erratum to “Nonlinear simulation of the effect of microenvironment on tumor growth”: [J. Theor. Biol. 245 (2007) 677–704]. Journal of Theoretical Biology, 247(3):581, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001609. See [ML07b].

Macklin:2007:NSE

[ML07b] Paul Macklin and John Lowengrub. Nonlinear simulation of the effect of microenvironment on tumor growth. Journal of Theoretical Biology, 245(4):677–704, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005649. See erratum [ML07a].

Morozov:2008:APM

[ML08a] Andrew Morozov and Bai-Lian Li. Abundance patterns in multi-species communities exposed to habitat destruction. Journal of Theoretical Biology, 251(4):593–605, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000234. See corrigen-

dum [ML08b].

Morozov:2008:CAP

[ML08b] Andrew Morozov and Bai-Lian Li. Corrigendum to “Abundance patterns in multi-species communities exposed to habitat destruction”: [J. Theor. Biol. 251 (2008) 593–605]. Journal of Theoretical Biology, 253(3):628, August 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002361. See [ML08a].

Mai:2009:BNB

[ML09a] Zhongxing Mai and Haiyan Liu. Boolean network-based analysis of the apoptosis network: Irreversible apoptosis and stable surviving. Journal of Theoretical Biology, 259(4):760–769, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001878. See corri-

gendum [ML12].

Mitchell:2009:QER

[ML09b] Cassie S. Mitchell and Robert H. Lee. A quantitative examination of the role of cargo-exerted forces in axonal transport. Journal of Theoretical Biology, 257(3):430–437, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006528.

Mai:2012:CBN

[ML12] Zhongxing Mai and Haiyan Liu. Corrigendum to “Boolean network-based analysis of the apoptosis network: Irreversible apoptosis and stable surviving” [J. Theor. Biol. 259 (4) (2009) 760–769]. Journal of Theoretical Biology, 313(??):217, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200416X. See [ML09a].

Maquet:2004:SMA

[MLA04] Jean Maquet, Christophe Letellier, and Luis A. Aguirre. Scalar modeling and analysis of a 3d biochemical reaction model. Journal of Theoretical Biology, 228(3):421–430, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000530.

Liu:2006:PDA

[mLLS^{+}06] Wei min Liu, Rui Li, James Z. Sun, Jing Wang, Julie Tsai, Wei Wen, Alexander Kohlmann, and P. Mickey Williams. PQN and DQN: Algorithms for expression microarrays. Journal of Theoretical Biology, 243(2):273–278, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002530.

Mao:2001:ETS

[MLMW01] J. H. Mao, K. A. Lindsay, R. J. Mairs, and T. E. Wheldon. The effect of tissue-specific growth patterns of target stem cells on the spectrum of tumours resulting from multistage tumorigenesis. Journal of Theoretical Biology, 210(1):93–100, May 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923008.

Moslonka-Lefebvre:2009:DSS

[MLPJ09] Mathieu Moslonka-Lefebvre, Marco Pautasso, and Mike J. Jeger. Disease spread in small-size directed networks: Epidemic threshold, correlation between links to and from nodes, and clustering. Journal of Theoretical Biology, 260(3):402–411, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002677.

Mossio:2009:CEC

[MLS09] Matteo Mossio, Giuseppe Longo, and John Stewart. A computable expression of closure to efficient causation. Journal of Theoretical Biology, 257(3):489–498, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006498.

Mi:2006:RBF

[MLWL06] Dong Mi, G. R. Liu, J.-S. Wang, and Z. R. Li. Relationships between the folding rate constant and the topological parameters of small two-state proteins based on general random walk model. Journal of Theoretical Biology, 241(1):152– 157, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004960.

Moore:2000:MDM

[MM00a] Gregory L. Moore and Costas D. Maranas. Modeling DNA mutation and recombination for directed evolution experiments. Journal of Theoretical Biology, 205(3):483–503, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920824.

Mouritsen:2000:MEM

[MM00b] Henrik Mouritsen and Ole Mouritsen. A mathematical expectation model for bird navigation based on the clock-and-compass strategy. Journal of Theoretical Biology, 207(2):283–291, November 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921714.

Manevitz:2002:MPR

[MM02a] Larry M. Manevitz and Shimon Marom. Modeling the process of rate selection in neuronal activity. Journal of Theoretical Biology, 216(3):337–343, June 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925397.

Moore:2002:PSR

[MM02b] Gregory L. Moore and Costas D. Maranas. Predicting out-of-sequence reassembly in DNA shuffling. Journal of Theoretical Biology, 219(1):9–17, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931024.

Mizuno:2003:TES

[MM03a] Shoji Mizuno and Akiko Mizuno. Theoretical examination of the standing osmotic gradient response to a step change of the solute influx into the symplast. Journal of Theoretical Biology, 225(3):413–417, December 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002649.

Mohtashemi:2003:EIR

[MM03b] Mojdeh Mohtashemi and Lik Mui. Evolution of indirect reciprocity by social information: the role of trust and reputation in evolution of altruism. Journal of Theoretical Biology, 223(4):523–531, August 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001437.

Mahdjoub:2008:PDT

[MM08a] Tewfik Mahdjoub and Frédéric Menu. Prolonged diapause: a trait increasing invasion speed? Journal of Theoretical Biology, 251(2):317–330, March 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006091.

Momiji:2008:DDH

[MM08b] Hiroshi Momiji and Nicholas A. M. Monk. Dissecting the dynamics of the Hes1 genetic oscillator. Journal of Theoretical Biology, 254(4):784–798, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003573.

Maxin:2009:RSA

[MM09] Daniel Maxin and Fabio Augusto Milner. The role of sexually abstained groups in two-sex demographic and epidemic logistic models with non-linear mortality. Journal of Theoretical Biology, 258(3):389–402, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004517.

Manning:2002:DRO

[MMTS02] J. T. Manning, S. Martin, R. L. Trivers, and M. Soler. 2nd to 4th digit ratio and offspring sex ratio. Journal of Theoretical Biology, 217(1):93–95, July 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930146.

Munteanu:2009:MTQ

[MMUGD09] Cristian Robert Munteanu, Alexandre L. Magalhães, Eugenio Uriarte, and Humberto González-Díaz. Multi-target QPDR classification model for human breast and colon cancerrelated proteins using star graph topological indices. Journal of Theoretical Biology, 257(2):303–311, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006140.

Medvinsky:2004:MIR

[MMV^{+}04] Alexander B. Medvinsky, Andrew Y. Morozov, Vassili V. Velkov, Bai-Lian Li, Mikhail S. Sokolov, and Horst Malchow. Modeling the invasion of recessive bt-resistant insects: An impact on transgenic plants. Journal of Theoretical Biology, 231(1):121–127, November 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002796.

Mcnair:2001:TTS

[MN01] James N. Mcnair and J. Denis Newbold. Turbulent transport of suspended particles and dispersing benthic organisms: the hitting-distance problem for the local exchange model. Journal of Theoretical Biology, 209(3):351–369, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301922738.

Mougi:2006:EMR

[MN06] Akihiko Mougi and Kinya Nishimura. Evolution of the maturation rate collapses competitive coexistence. Journal of Theoretical Biology, 241(3):467–476, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005424.

Mougi:2007:ELH

[MN07a] Akihiko Mougi and Kinya Nishimura. Evolution of lifehistory traits collapses competitive coexistence. Journal of Theoretical Biology, 248(3):552–559, October 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002676.

Mougi:2007:RPE

[MN07b] Akihiko Mougi and Kinya Nishimura. A resolution of the paradox of enrichment. Journal of Theoretical Biology, 248(1):194–201, September 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001877.

Michor:2006:SDM

[MNI06] Franziska Michor, Martin A. Nowak, and Yoh Iwasa. Stochastic dynamics of metastasis formation. Journal of Theoretical Biology, 240(4):521–530, June 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004534.

Moon:2007:HAG

[MNL^{+}07] Sung Joon Moon, B. Nabet, Naomi E. Leonard, Simon A. Levin, and I. G. Kevrekidis. Heterogeneous animal group models and their group-level alignment dynamics: An equationfree approach. Journal of Theoretical Biology, 246(1):100–112, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005741.

Montero:2008:SAS

[MNMH^{+}08] Francisco Montero, Juan Carlos Nuño, Enrique Meléndez-Hevia, Felix Olasagasti, Sara Vázquez, and Federico Morán. Stoichiometric analysis of self-maintaining metabolisms. Journal of Theoretical Biology, 252(3):427–432, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005966.

Meyers:2006:PED

[MNP06] Lauren Ancel Meyers, M. E. J. Newman, and Babak Pourbohloul. Predicting epidemics on directed contact networks. Journal of Theoretical Biology, 240(3):400–418, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004418.

Matzavinos:2007:SAA

[MO07] Anastasios Matzavinos and Hans G. Othmer. A stochastic analysis of actin polymerization in the presence of twinfilin and gelsolin. Journal of Theoretical Biology, 249(4):723–736, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930700402X.

Manapat:2009:OD

[MOBN09] Michael Manapat, Hisashi Ohtsuki, Reinhard Bürger, and Martin A. Nowak. Originator dynamics. Journal of Theoretical Biology, 256(4):586–595, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800516X.

Mochizuki:2002:PFC

[Moc02] Atsushi Mochizuki. Pattern formation of the cone mosaic in the zebrafish retina: a cell rearrangement model. Journal of Theoretical Biology, 215(3):345–361, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925081.

Mochizuki:2005:ASN

[Moc05] Atsushi Mochizuki. An analytical study of the number of steady states in gene regulatory networks. Journal of Theoretical Biology, 236(3):291–310, October 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001244.

Mochizuki:2008:SRN

[Moc08] Atsushi Mochizuki. Structure of regulatory networks and diversity of gene expression patterns. Journal of Theoretical Biology, 250(2):307–321, January 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004481.

Mohle:2000:APP

[Möh00] M. Möhle. Ancestral processes in population genetics-the coalescent. Journal of Theoretical Biology, 204(4):629–638, June 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920320.

Moller:2001:DIU

[Möl01] Ralf Möller. Do insects use templates or parameters for landmark navigation? Journal of Theoretical Biology, 210(1):33– 45, May 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922957.

Moller:2002:ICE

[Möl02] Ralf Möller. Insects could exploit UV-green contrast for landmark navigation. Journal of Theoretical Biology, 214(4):619–631, February 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924841.

Miekisz:2005:PDS

[MP05] Jacek Mieķisz and Tadeusz Platkowski. Population dynamics with a stable efficient equilibrium. Journal of Theoretical Biology, 237(4):363–368, December 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001852.

Mitteldorf:2009:SAL

[MP09a] Josh Mitteldorf and John Pepper. Senescence as an adaptation to limit the spread of disease. Journal of Theoretical Biology, 260(2):186–195, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002422.

Mogielski:2009:MDI

[MP09b] Krzysztof Mogielski and Tadeusz Platkowski. A mechanism of dynamical interactions for two-person social dilemmas. Journal of Theoretical Biology, 260(1):145–150, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002598.

Merler:2008:CCT

[MPA^{+}08] Stefano Merler, Piero Poletti, Marco Ajelli, Bruno Caprile, and Piero Manfredi. Coinfection can trigger multiple pandemic waves. Journal of Theoretical Biology, 254(2):499– 507, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519308003056.

Medan:2007:AAN

[MPD^{+}07] Diego Medan, Roberto P. J. Perazzo, Mariano Devoto, Enrique Burgos, Martín G. Zimmermann, Horacio Ceva, and Ana M. Delbue. Analysis and assembling of network structure in mutualistic systems. Journal of Theoretical Biology, 246(3):510– 521, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000045.

Moorcroft:2006:PRN

[MPL06a] P. R. Moorcroft, S. W. Pacala, and M. A. Lewis. Potential role of natural enemies during tree range expansions following climate change. Journal of Theoretical Biology, 241(3):601–616, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005588.

Morozov:2006:SCP

[MPL06b] Andrew Morozov, Sergei Petrovskii, and Bai-Lian Li. Spatiotemporal complexity of patchy invasion in a predatorprey system with the Allee effect. Journal of Theoretical Biology, 238(1):18–35, January 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002237.

McBryde:2007:SMM

[MPM07] E. S. McBryde, A. N. Pettitt, and D. L. S. McElwain. A stochastic mathematical model of methicillin resistant staphylococcus aureus transmission in an intensive care unit: Predicting the impact of interventions. Journal of Theoretical Biology, 245(3):470–481, April 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005480.

Meyers:2005:NTS

[MPN^{+}05] Lauren Ancel Meyers, Babak Pourbohloul, M. E. J. Newman, Danuta M. Skowronski, and Robert C. Brunham. Network theory and SARS: predicting outbreak diversity. Journal of Theoretical Biology, 232(1):71–81, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003510.

Morozov:2007:TRP

[MPN07] Andrew Yu. Morozov, Sergei V. Petrovskii, and Nikolay P. Nezlin. Towards resolving the paradox of enrichment: The impact of zooplankton vertical migrations on plankton systems stability. Journal of Theoretical Biology, 248(3):501–511, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002615.

Marrero-Ponce:2009:NBI

[MPOBD^{+}09] Yovani Marrero-Ponce, Sadiel E. Ortega-Broche, Yunaimy Echeverría Díaz, Ysaias J. Alvarado, Nestor Cubillan, Gladys Casas Cardoso, Francisco Torrens, and Facundo Pérez-Giménez. Nucleotide’s bilinear indices: Novel bio-macromolecular descriptors for bioinformatics studies of nucleic acids. I. Prediction of paromomycin’s affinity constant with HIV-1 Ψ-RNA packaging region. Journal of Theoretical Biology, 259(2):229–241, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000940.

Marshall:2003:KSM

[MR03] James A. R. Marshall and Jonathan E. Rowe. Kin selection may inhibit the evolution of reciprocation. Journal of Theoretical Biology, 222(3):331–335, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000390.

Merilaita:2007:ASR

[MR07a] Sami Merilaita and Graeme D. Ruxton. Aposematic signals and the relationship between conspicuousness and distinctiveness. Journal of Theoretical Biology, 245(2):268–277, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005133.

Milosevic:2007:SAN

[MR07b] Nebojsa T. Milosević and Dusan Ristanović. The sholl analysis of neuronal cell images: Semi-log or log-log method? Journal of Theoretical Biology, 245(1):130–140, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004322.

Merkle:2006:EPI

[MRA06] Tobias Merkle, Martin Rost, and Wolfgang Alt. Egocentric path integration models and their application to desert arthropods. Journal of Theoretical Biology, 240(3):385–399, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500442X.

Merkey:2009:MHS

[MRC09] Brian V. Merkey, Bruce E. Rittmann, and David L. Chopp. Modeling how soluble microbial products (SMP) support heterotrophic bacteria in autotroph-based biofilms. Journal of Theoretical Biology, 259(4):670–683, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002379.

Morozov:2007:SMD

[MRF07] Andrew Yu. Morozov, Cécile Robin, and Alain Franc. A simple model for the dynamics of a host-parasite-hyperparasite interaction. Journal of Theoretical Biology, 249(2):246–253, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002251.

Martinez-Reina:2008:RBD

[MRGADD08] J. Martínez-Reina, J. M. García-Aznar, J. Domínguez, and M. Doblaré. On the role of bone damage in calcium homeostasis. Journal of Theoretical Biology, 254(3):704–712, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003007.

Milosevic:2009:QAD

[MRJR09] Nebojsa T. Milosević, Dusan Ristanović, Herbert F. Jelinek, and Katarina Rajković. Quantitative analysis of dendritic morphology of the alpha and delta retinal ganglion cells in the rat: a cell classification study. Journal of Theoretical Biology, 259(1):142–150, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001192.

Medvedev:2001:DRM

[MS01a] Dmitry Medvedev and Alexei A. Stuchebrukhov. DNA repair mechanism by photolyase: Electron transfer path from the photolyase catalytic cofactor FADH^{−} to DNA thymine dimer. Journal of Theoretical Biology, 210(2):237–248, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930192291X.

Miura:2001:MME

[MS01b] Tomoyuki Miura and Pierre Sonigo. A mathematical model for experimental gene evolution. Journal of Theoretical Biology, 209(4):497–502, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922866.

Montoya:2002:SWP

[MS02] Jose M. Montoya and Ricard V. Solé. Small world patterns in food Webs. Journal of Theoretical Biology, 214(3):405–412, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924609.

Marhl:2003:UWC

[MS03a] Marko Marhl and Stefan Schuster. Under what conditions signal transduction pathways are highly flexible in response to external forcing? A case study on calcium oscillations. Journal of Theoretical Biology, 224(4):491–500, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001991.

McLachlan:2003:HMC

[MS03b] John C. McLachlan and Helen Storey. Hot male: can sex in humans be modified by temperature? Journal of Theoretical Biology, 222(1):71–72, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000146.

Melkikh:2005:MAT

[MS05a] Alexey V. Melkikh and Vladimir D. Seleznev. Models of active transport of ions in biomembranes of various types of cells. Journal of Theoretical Biology, 234(3):403–412, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304006009.

Mossel:2005:RBN

[MS05b] Elchanan Mossel and Mike Steel. Random biochemical networks: the probability of self-sustaining autocatalysis. Journal of Theoretical Biology, 233(3):327–336, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004825.

Melkikh:2006:MAT

[MS06a] Alexey V. Melkikh and Vladimir D. Seleznev. Model of active transport of ions in biomembranes based on ATP-dependent change of height of diffusion barriers to ions. Journal of Theoretical Biology, 242(3):617–626, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001573.

Munteanu:2006:PDC

[MS06b] Andreea Munteanu and Ricard V. Solé. Phenotypic diversity and chaos in a minimal cell model. Journal of Theoretical Biology, 240(3):434–442, June 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004510.

Macia:2007:PSR

[MS07a] Javier Macía and Ricard V. Solé. Protocell self-reproduction in a spatially extended metabolism-vesicle system. Journal of Theoretical Biology, 245(3):400–410, April 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005121.

Melkikh:2007:MAT

[MS07b] Alexey V. Melkikh and Vladimir D. Seleznev. Models of active transport of neurotransmitters in synaptic vesicles. Journal of Theoretical Biology, 248(2):350–353, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002640.

Manor:2008:FCV

[MS08a] Alon Manor and Nadav M. Shnerb. Facilitation, competition, and vegetation patchiness: From scale free distribution to patterns. Journal of Theoretical Biology, 253(4):838–842, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001914.

Matsuoka:2008:EBN

[MS08b] Tsutomu Matsuoka and Hiromi Seno. Ecological balance in the native population dynamics may cause the paradox of pest control with harvesting. Journal of Theoretical Biology, 252(1):87–97, May 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000295.

Melkikh:2008:MAT

[MS08c] A. V. Melkikh and M. I. Sutormina. Model of active transport of ions in cardiac cell. Journal of Theoretical Biology, 252(2):247–254, May 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000568.

Moyano:2009:ELR

[MS09a] Luis G. Moyano and Angel Sánchez. Evolving learning rules and emergence of cooperation in spatial prisoner’s dilemma. Journal of Theoretical Biology, 259(1):84–95, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000988.

Munteanu:2009:MNL

[MS09b] Andreea Munteanu and Peter F. Stadler. Mutate now, die later. Evolutionary dynamics with delayed selection. Journal of Theoretical Biology, 260(3):412–421, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002999.

Marva:2009:RSF

[MSdlPS09] M. Marvá, E. Sánchez, R. Bravo de la Parra, and L. Sanz. Reduction of slow-fast discrete models coupling migration and demography. Journal of Theoretical Biology, 258(3):371–379, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003664.

Mohtashemi:2006:SIM

[MSDM06] Mojdeh Mohtashemi, Peter Szolovits, James Dunyak, and Kenneth D. Mandl. A susceptible-infected model of early detection of respiratory infection outbreaks on a background of influenza. Journal of Theoretical Biology, 241(4):954–963, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000531.

Martis:2008:RCI

[MSGS08] Remila L. Martis, Sameer Kumar Singh, M. Michael Gromiha, and C. Santhosh. Role of cation-π interactions in single chain ‘all-alpha’ proteins. Journal of Theoretical Biology, 250(4):655–662, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005206.

Marion:2005:UFB

[MSH05] Glenn Marion, David L. Swain, and Mike R. Hutchings. Understanding foraging behaviour in spatially heterogeneous environments. Journal of Theoretical Biology, 232(1):127–142, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003765.

Muko:2001:DMS

[MSI01] Soyoka Muko, Kazuhiko Sakai, and Yoh Iwasa. Dynamics of marine sessile organisms with space-limited growth and recruitment: Application to corals. Journal of Theoretical Biology, 210(1):67–80, May 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922982.

Murphy:2003:TTH

[MSK03] Brian M. Murphy, Benjamin H. Singer, and Denise Kirschner. On treatment of tuberculosis in heterogeneous populations. Journal of Theoretical Biology, 223(4):391–404, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000389.

Morgan:2004:FWC

[MSL04] Jeffrey J. Morgan, Ivan V. Surovtsev, and Paul A. Lindahl. A framework for whole-cell mathematical modeling. Journal of Theoretical Biology, 231(4):581–596, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400342X.

Miura:2006:MMP

[MSMKM06] Takashi Miura, Kohei Shiota, Gillian Morriss-Kay, and Philip K. Maini. Mixed-mode pattern in doublefoot mutant mouse limb–Turing reaction–diffusion model on a growing domain during limb development. Journal of Theoretical Biology, 240(4):562–573, June 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004686.

MacCarthy:2003:EPD

[MSP03] Thomas MacCarthy, Robert Seymour, and Andrew Pomiankowski. The evolutionary potential of the drosophila sex determination gene network. Journal of Theoretical Biology, 225(4):461–468, December 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002820.

Marwan:2005:RRN

[MSS05] Wolfgang Marwan, Arumugam Sujatha, and Christine Starostzik. Reconstructing the regulatory network controlling commitment and sporulation in Physarum polycephalum based on hierarchical Petri net modelling and simulation. Journal of Theoretical Biology, 236(4):349–365, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930500127X.

Moulton:2007:OPD

[MSS07] Vincent Moulton, Charles Semple, and Mike Steel. Optimizing phylogenetic diversity under constraints. Journal of Theoretical Biology, 246(1):186–194, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600590X.

Murabito:2009:CEM

[MSSS09] Ettore Murabito, Evangelos Simeonidis, Kieran Smallbone, and Jonathan Swinton. Capturing the essence of a metabolic network: a flux balance analysis approach. Journal of Theoretical Biology, 260(3):445–452, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002732.

Mcnamara:2000:DGT

[MSWH00] John M. Mcnamara, Tamás Székely, James N. Webb, and Alasdair I. Houston. A dynamic game-theoretic model of parental care. Journal of Theoretical Biology, 205(4):605–623, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920939.

Masaka:2006:CFS

[MT06a] Kazuhiko Masaka and Takenori Takada. Corrigendum to “Floral sex ratio strategy in wind-pollinated monoecious species subject to wind-pollination efficiency and competitive sharing among male flowers as a game”: [J. Theor. Biol. 240 (2006) 114–125]. Journal of Theoretical Biology, 242(1):264, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002013. See [MT06b].

Masaka:2006:FSR

[MT06b] Kazuhiko Masaka and Takenori Takada. Floral sex ratio strategy in wind-pollinated monoecious species subject to wind-pollination efficiency and competitive sharing among male flowers as a game. Journal of Theoretical Biology, 240(1):114–125, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003723. See [MT06a].

Mazzag:2005:ERC

[MTS05] Borbala Mazzag, Christopher J. Tignanelli, and Gregory D. Smith. The effect of residual Ca^{2+} on the stochastic gating of Ca^{2+}-regulated Ca^{2+} channel models. Journal of Theoretical Biology, 235(1):121–150, July 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000044.

Murphy:2000:SCM

[Mur00] Ricardo Murphy. Some compartmental models of the root: Steady-state behavior. Journal of Theoretical Biology, 207(4):557–576, December 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921945.

Murugan:2007:ADA

[Mur07] Rajamanickam Murugan. Action at distance along the DNA: a random jump model on enhancer action. Journal of Theoretical Biology, 248(4):696–705, October 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002822.

Muzur:2005:TIT

[Muz05] Amir Muzur. Toward an integrative theory of sleep and dreaming. Journal of Theoretical Biology, 233(1):103–118, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004667.

Muller:2002:OMPa

[MV02a] M. Muller and J. H. G. Verhagen. Optimization of the mechanical performance of a two-duct semicircular Duct system — Part 1: Dynamics and Duct dimensions. Journal of Theoretical Biology, 216(4):409–424, June 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930031.

Muller:2002:OMPb

[MV02b] M. Muller and J. H. G. Verhagen. Optimization of the mechanical performance of a two-duct semicircular Duct system — Part 2: Excitation of endolymph movements. Journal of Theoretical Biology, 216(4):425–442, June 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930043.

Muller:2002:OMPc

[MV02c] M. Muller and J. H. G. Verhagen. Optimization of the mechanical performance of a two-duct semicircular Duct system — Part 3: The positioning of the ducts in the head. Journal of Theoretical Biology, 216(4):443–459, June 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930055.

Makela:2006:QPS

[MV06] Annikki Mäkelä and Harry T. Valentine. The quarterpower scaling model does not imply size-invariant hydraulic resistance in plants. Journal of Theoretical Biology, 243(2):283–285, November 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002451.

Murado:2007:NHD

[MV07] M. A. Murado and J. A. Vázquez. The notion of hormesis and the dose-response theory: a unified approach. Journal of Theoretical Biology, 244(3):489–499, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004036.

McClay:2006:VCA

[MvdO06] Joseph L. McClay and Edwin J. C. G. van den Oord. Variance component analysis of polymorphic metabolic systems.

Journal of Theoretical Biology, 240(1):149–159, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004145.

Martinov:2000:SSN

[MVM^{+}00] Michael V. Martinov, Victor M. Vitvitsky, Eugene V. Mosharov, Ruma Banerjee, and Fazoil I. Ataullakhanov. A substrate switch: a new mode of regulation in the methionine metabolic pathway. Journal of Theoretical Biology, 204(4):521–532, June 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920356.

Michod:2006:LHE

[MVS^{+}06] Richard E. Michod, Yannick Viossat, Cristian A. Solari, Mathilde Hurand, and Aurora M. Nedelcu. Life-history evolution and the origin of multicellularity. Journal of Theoretical Biology, 239(2):257–272, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003875.

Mitteldorf:2000:PVE

[MW00] Joshua Mitteldorf and David Sloan Wilson. Population viscosity and the evolution of altruism. Journal of Theoretical Biology, 204(4):481–496, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920071.

Marique:2001:CKC

[MW01] T. Marique and J. Wérenne. Control of 92 kDa collagenase secretion in mammalian cells by modulation of AP-1 activity: an experimentally based theoretical study. Journal of Theoretical Biology, 209(1):3–8, March 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922379.

McKane:2007:SSD

[MW07] A. J. McKane and D. Waxman. Singular solutions of the diffusion equation of population genetics. Journal of Theoretical Biology, 247(4):849–858, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001981.

Menge:2009:DNE

[MW09] Duncan N. L. Menge and Joshua S. Weitz. Dangerous nutrients: Evolution of phytoplankton resource uptake subject to virus attack. Journal of Theoretical Biology, 257(1):104–115, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005596.

Morohashi:2002:RMP

[MWB^{+}02] Mineo Morohashi, Amanda E. Winn, Mark T. Borisuk, Hamid Bolouri, John Doyle, and Hiroaki Kitano. Robustness as a measure of plausibility in models of biochemical networks. Journal of Theoretical Biology, 216(1):19–30, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925373.

Miller:2005:EHR

[MWB05] M. R. Miller, A. White, and M. Boots. The evolution of host resistance: Tolerance and control as distinct strategies. Journal of Theoretical Biology, 236(2):198–207, September 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305000949.

McClellan:2004:GCE

[MWCS04] David A. McClellan, David G. Whiting, Ryan Christensen, and Joshua Sailsbery. Genetic codes as evolutionary filters: subtle differences in the structure of genetic codes result in significant differences in patterns of nucleotide substitution. Journal of Theoretical Biology, 226(4):393–400, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003552.

Murphy:2008:CMA

[MWD08] James T. Murphy, Ray Walshe, and Marc Devocelle. A computational model of antibiotic-resistance mechanisms in Methicillin-Resistant Staphylococcus aureus (MRSA). Journal of Theoretical Biology, 254(2):284–293, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002956.

Muneepeerakul:2007:NMM

[MWL^{+}07] Rachata Muneepeerakul, Joshua S. Weitz, Simon A. Levin, Andrea Rinaldo, and Ignacio Rodriguez-Iturbe. A neutral metapopulation model of biodiversity in river networks. Journal of Theoretical Biology, 245(2):351–363, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004747.

Macia:2009:WCS

[MWS09] Javier Macía, Stefanie Widder, and Ricard Solé. Why are cellular switches Boolean? General conditions for multistable genetic circuits. Journal of Theoretical Biology, 261(1):126–135, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003221.

Marin:2008:GSP

[MX08] Antonio Marín and Xuhua Xia. GC skew in protein-coding genes between the leading and lagging strands in bacterial genomes: New substitution models incorporating strand bias. Journal of Theoretical Biology, 253(3):508–513, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001793.

Majewska:2001:TGJ

[MY01] Ania Majewska and Rafael Yuste. Topology of gap junction networks in C. elegans. Journal of Theoretical Biology, 212(2):155–167, September 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923641.

Maidana:2009:SSW

[MY09] Norberto Aníbal Maidana and Hyun Mo Yang. Spatial spreading of West Nile virus described by traveling waves. Journal of Theoretical Biology, 258(3):403–417, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006322.

Manrubia:2002:BBB

[MZ02] Susanna C. Manrubia and Damián H. Zanette. At the boundary between biological and cultural evolution: The origin of surname distributions. Journal of Theoretical Biology, 216(4):461–477, June 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930293002X.

Mazurkiewicz:2008:STS

[MZK08] Joanna Mazurkiewicz, Jarosław Zygierewicz, and Mikolaj Korzyński. Short term synaptic depression model — analytical solution and analysis. Journal of Theoretical Biology, 254(1):82–88, September 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002506.

Nikolaou:2002:SMS

[NA02] Christoforos Nikolaou and Yannis Almirantis. A study of the middle-scale nucleotide clustering in DNA sequences of various origin and functionality, by means of a method based on a modified standard deviation. Journal of Theoretical Biology, 217(4):479–492, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930456.

Nikolaou:2003:MSC

[NA03] Christoforos Nikolaou and Yannis Almirantis. Mutually symmetric and complementary triplets: differences in their use distinguish systematically between coding and noncoding genomic sequences. Journal of Theoretical Biology, 223(4):477–487, August 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001231.

Ninio:2004:HBC

[NA04] Jacques Ninio and Sebastian Amigorena. How B cells and dendritic cells may cooperate in antigen purification. Journal of Theoretical Biology, 231(3):309–317, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003170.

Nakajima:2001:PEB

[Nak01] Toshiyuki Nakajima. Probabilities of encounters between objects in biological systems: Meta-observer view. Journal of Theoretical Biology, 211(4):347–363, August 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923549.

Nakajima:2003:PEB

[Nak03] Toshiyuki Nakajima. Probabilities of encounters between objects in biological systems 2: Cognizer view. Journal of Theoretical Biology, 221(1):39–51, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931687.

Nagarajan:2004:MGN

[NAP04] Radhakrishnan Nagarajan, Jane E. Aubin, and Charlotte A. Peterson. Modeling genetic networks from clonal analysis. Journal of Theoretical Biology, 230(3):359–373, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002632.

Narang:2006:CAS

[Nar06a] Atul Narang. Comparative analysis of some models of gene regulation in mixed-substrate microbial growth. Journal of Theoretical Biology, 242(2):489–501, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001226.

Narang:2006:SPE

[Nar06b] Atul Narang. Spontaneous polarization in eukaryotic gradient sensing: a mathematical model based on mutual inhibition of frontness and backness pathways. Journal of Theoretical Biology, 240(4):538–553, June 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004662.

Narang:2007:EDL

[Nar07] Atul Narang. Effect of DNA looping on the induction kinetics of the lac operon. Journal of Theoretical Biology, 247(4):695–712, August 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001464.

Naasell:2001:EQS

[Nås01a] Ingemar Nåsell. Extinction and quasi-stationarity in the Verhulst logistic model. Journal of Theoretical Biology, 211(1):11– 27, July 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923288.

Nashimoto:2001:RPS

[Nas01b] Masayuki Nashimoto. The RNA/Protein symmetry hypothesis: Experimental support for reverse translation of primitive proteins. Journal of Theoretical Biology, 209(2):181–187, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922537.

Nikolaev:2007:SCA

[NAS07] Evgeni V. Nikolaev, Jordan C. Atlas, and Michael L. Shuler. Sensitivity and control analysis of periodically forced reaction networks using the Green’s function method. Journal of Theoretical Biology, 247(3):442–461, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000859.

Nemeth:2004:CGP

[NAV04] Zs. I. Németh, L. Albert, and Sz. Varga. Characterization of germination with physical parameters: correlation of relative mass and density as an indicator function of the germination of European Turkey oak acorn. Journal of Theoretical Biology, 231(2):167–174, November 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400284X.

Nicosia:2002:MME

[NB02] Mark A. Nicosia and James G. Brasseur. A mathematical model for estimating muscle tension in vivo during esophageal bolus transport. Journal of Theoretical Biology, 219(2):235–255, November 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302931188.

Ndiaye:2006:RTD

[NBMS06] P. I. Ndiaye, D. J. Bicout, B. Mondet, and P. Sabatier. Rainfall triggered dynamics of aedes mosquito aggressiveness. Journal of Theoretical Biology, 243(2):222–229, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930600244X.

Nakamaru:2007:DDF

[NBT07] Mayuko Nakamaru, Yayoi Beppu, and Kazuki Tsuji. Does disturbance favor dispersal? An analysis of ant migration using the colony-based lattice model. Journal of Theoretical Biology, 248(2):288–300, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002196.

Naug:2002:RCO

[NC02] Dhruba Naug and Scott Camazine. The role of colony organization on pathogen transmission in social insects. Journal of Theoretical Biology, 215(4):427–439, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192524X.

Nalwaya:2004:AEN

[ND04] Nitesh Nalwaya and William M. Deen. Analysis of the effects of nitric oxide and oxygen on nitric oxide production by macrophages. Journal of Theoretical Biology, 226(4):409–419, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003606.

Nakamaru:2009:RSC

[ND09] Mayuko Nakamaru and Ulf Dieckmann. Runaway selection for cooperation and strict-and-severe punishment. Journal of Theoretical Biology, 257(1):1–8, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800475X.

Nicolis:2008:CDM

[NDD08] S. C. Nicolis, E. Despland, and A. Dussutour. Collective decision-making and behavioral polymorphism in group living organisms. Journal of Theoretical Biology, 254(3):580–586, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003354.

Nishiura:2006:ENR

[NDE06] Hiroshi Nishiura, Klaus Dietz, and Martin Eichner. The earliest notes on the reproduction number in relation to herd immunity: Theophil lotz and smallpox vaccination. Journal of Theoretical Biology, 241(4):964–967, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000257.

Nunes:2006:LCB

[NdGG06] A. Nunes, M. M. Telo da Gama, and M. G. M. Gomes. Localized contacts between hosts reduce pathogen diversity. Journal of Theoretical Biology, 241(3):477–487, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305005436.

Neill:2001:OUN

[Nei01] Daniel B. Neill. Optimality under noise: Higher memory strategies for the alternating Prisoner’s dilemma. Journal of Theoretical Biology, 211(2):159–180, July 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923379.

Neill:2004:ESL

[Nei04] Daniel B. Neill. Evolutionary stability for large populations. Journal of Theoretical Biology, 227(3):397–401, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004363.

Nettle:2009:EML

[Net09] Daniel Nettle. An evolutionary model of low mood states. Journal of Theoretical Biology, 257(1):100–103, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005717.

Nikoloski:2008:MNN

[NGMS08] Zoran Nikoloski, Sergio Grimbs, Patrick May, and Joachim Selbig. Metabolic networks are NP-hard to reconstruct. Journal of Theoretical Biology, 254(4):807–816, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003706.

Nevo:2004:AID

[NGN^{+}04] Uri Nevo, Ido Golding, Avidan U. Neumann, Michał Schwartz, and Solange Akselrod. Autoimmunity as an immune defense against degenerative processes: a primary mathematical model illustrating the bright side of autoimmunity. Journal of Theoretical Biology, 227(4):583–592, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004521.

Najafabadi:2005:OCU

[NGT05] Hamed Shateri Najafabadi, Hani Goodarzi, and Noorossadat Torabi. Optimality of codon usage in Escherichia coli due to load minimization. Journal of Theoretical Biology, 237(2):203–209, November 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001621.

Najafabadi:2006:ANN

[NGTB06] Hamed Shateri Najafabadi, Hani Goodarzi, Noorossadat Torabi, and Setareh Sadat Banihosseini. Applying a neural network to predict the thermodynamic parameters for an expanded nearest-neighbor model. Journal of Theoretical Biology, 238(3):657–665, February 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002729.

Nazaret:2009:MEM

[NHTM09] Christine Nazaret, Margit Heiske, Kevin Thurley, and Jean-Pierre Mazat. Mitochondrial energetic metabolism: a simplified model of TCA cycle with ATP production. Journal of Theoretical Biology, 258(3):455–464, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004931.

Nishimura:2004:EEC

[NI04a] Kinya Nishimura and Yutaka Isoda. Erratum to “Evolution of cannibalism: referring to costs of cannibalism”: [J. Theor. Biol. 226, 291–302]. Journal of Theoretical Biology, 228(2):291, May 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000578. See [NI04b].

Nishimura:2004:ECR

[NI04b] Kinya Nishimura and Yutaka Isoda. Evolution of cannibalism: referring to costs of cannibalism. Journal of Theoretical Biology, 226(3):293–302, February 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003370. See erratum [NI04a].

Nishimura:2004:VET

[NI04c] Kinya Nishimura and Yutaka Isoda. Variant evolutionary trees under phenotypic variance. Journal of Theoretical Biology, 226(1):79–87, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303003229.

Nakamaru:2006:CAP

[NI06] Mayuko Nakamaru and Yoh Iwasa. The coevolution of altruism and punishment: Role of the selfish punisher. Journal of Theoretical Biology, 240(3):475–488, June 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004509. See corrigendum [NI07a].

Nakamaru:2007:CCA

[NI07a] Mayuko Nakamaru and Yoh Iwasa. Corrigendum to “The coevolution of altruism and punishment: Role of the selfish punisher”: [J. Theor. Biol. 240 (2006) 475–488]. Journal of Theoretical Biology, 245(1):193–196, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005716. See [NI06].

Nishiura:2007:DEC

[NI07b] Hiroshi Nishiura and Hisashi Inaba. Discussion: Emergence of the concept of the basic reproduction number from mathematical demography. Journal of Theoretical Biology, 244(2):357–364, January 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600347X.

Nielsen:2002:EEG

[Nie02] Bjørn Gilbert Nielsen. Entropic elasticity in the generation of muscle force — a theoretical model. Journal of Theoretical Biology, 219(1):99–119, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931139.

Niell:2006:TAS

[Nie06] Cristopher M. Niell. Theoretical analysis of a synaptotropic dendrite growth mechanism. Journal of Theoretical Biology, 241(1):39–48, July 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004984.

Niklas:2002:ABP

[Nik02] Karl J. Niklas. On the allometry of biomass partitioning and light harvesting for plants with leafless stems. Journal of Theoretical Biology, 217(1):47–52, July 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930195.

Nir:2002:FDM

[Nir02] Avinoam Nir. The fertilization dance: a mechanical view of the egg rotation during the initial spermatozoa–ovum interaction. Journal of Theoretical Biology, 214(2):171–179, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924646.

Nishii:2006:AEE

[Nis06]  Jun Nishii.  An analytical estimation of the energy

cost for legged locomotion. Journal of Theoretical Biology, 238(3):636–645, February 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002857.

Niwa:2003:PLV

[Niw03] Hiro-Sato Niwa. Power-law versus exponential distributions of animal group sizes. Journal of Theoretical Biology, 224(4):451–457, October 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001929.

Niwa:2004:SIS

[Niw04] Hiro-Sato Niwa. Space-irrelevant scaling law for fish school sizes. Journal of Theoretical Biology, 228(3):347–357, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000396.

Niwa:2005:PLS

[Niw05] Hiro-Sato Niwa. Power-law scaling in dimension-to-biomass relationship of fish schools. Journal of Theoretical Biology, 235(3):419–430, August 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000494.

Nolting:2004:ECP

[NJVA04] Bengt Nölting, Dennis Jülich, Winfried Vonau, and Karl Andert. Evolutionary computer programming of protein folding and structure predictions. Journal of Theoretical Biology, 229(1):13–18, July 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000785.

Nakajima:2008:RDB

[NK08] Akihiko Nakajima and Kunihiko Kaneko. Regulative differentiation as bifurcation of interacting cell population. Journal of Theoretical Biology, 253(4):779–787, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001835.

Neimark:2008:BCM

[NKC^{+}08] Matthew A. Neimark, Angelos-Aristeidis Konstas, Jae H. Choi, Andrew F. Laine, and John Pile-Spellman. Brain cooling maintenance with cooling cap following induction with intracarotid cold saline infusion: a quantitative model. Journal of Theoretical Biology, 253(2):333–344, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001513.

Na:2006:MMH

[NKL06] Dokyun Na, Dongsup Kim, and Doheon Lee. Mathematical modeling of humoral immune response suppression by passively administered antibodies in mice. Journal of Theoretical Biology, 241(4):830–851, August 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000324.

Newman:2003:MMS

[NL03a] William I. Newman and Jorge A. Lazareff. A mathematical model for self-limiting brain tumors. Journal of Theoretical Biology, 222(3):361–371, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000432.

Nikolic:2003:SSM

[NL03b] Djordje L. Nikolic and Daniel J. Lacks. Size scaling of mutation avalanches in a model for protein evolution. Journal of Theoretical Biology, 221(2):259–267, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931882.

Nakamaru:2004:STL

[NL04] Mayuko Nakamaru and Simon A. Levin. Spread of two linked social norms on complex interaction networks. Journal of Theoretical Biology, 230(1):57–64, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001833.

Nukoolkarn:2008:MDS

[NLM^{+}08] Veena Nukoolkarn, Vannajan Sanghiran Lee, Maturos Malaisree, Ornjira Aruksakulwong, and Supot Hannongbua. Molecular dynamic simulations analysis of ritronavir and lopinavir as SARS-CoV 3CL pro inhibitors. Journal of Theoretical Biology, 254(4):861–867, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003718.

Nedorezov:2008:AIS

[NLS08] Lev V. Nedorezov, Bernhard L. Löhr, and Dinara L. Sadykova. Assessing the importance of self-regulating mechanisms in diamondback moth population dynamics: Application of discrete mathematical models. Journal of Theoretical Biology, 254(3):587–593, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003421.

Nartsissov:2006:ARB

[NM06] Yaroslav R. Nartsissov and Elena V. Mashkovtseva. Application of rigid body mechanics to theoretical description of rotation within F_{0}F_{1}-ATP synthase. Journal of Theoretical Biology, 242(2):300–308, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000907.

Nazaret:2008:OPR

[NM08] Christine Nazaret and Jean-Pierre Mazat. An old paper revisited: “A mathematical model of carbohydrate energy metabolism. Interaction between glycolysis, the Krebs cycle and the H-transporting shuttles at varying ATPases load” by V. V. Dynnik, R. Heinrich and E. E. Sel’kov. Journal of Theoretical Biology, 252(3):520–529, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000064.

Nakazato:2009:STG

[NM09] Kenichi Nakazato and Atsushi Mochizuki. Steepness of thermal gradient is essential to obtain a unified view of thermotaxis in C. elegans. Journal of Theoretical Biology, 260(1):56–65, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002549.

Na:2007:TMF

[NMH07] S. Na, G. A. Meininger, and J. D. Humphrey. A theoretical model for f-actin remodeling in vascular smooth muscle cells subjected to cyclic stretch. Journal of Theoretical Biology, 246(1):87–99, May 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005583.

Nowak:2006:GIC

[NMI06] Martin A. Nowak, Franziska Michor, and Yoh Iwasa. Genetic instability and clonal expansion. Journal of Theoretical Biology, 241(1):26–32, July 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004972.

Nagano:2007:MTB

[NN07] Kozo Nagano and Nozomi Nagano. Mechanism of translation based on intersubunit complementarities of ribosomal RNAs and tRNAs. Journal of Theoretical Biology, 245(4):644–668, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005455.

Navid:2004:SCD

[NO04] Ali Navid and Peter J. Ortoleva. Simulated complex dynamics of glycolysis in the protozoan parasite trypanosoma brucei. Journal of Theoretical Biology, 228(4):449–458, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000566.

Noest:2000:DLF

[Noe00] André J. Noest. Designing lymphocyte functional structure for optimal signal detection: Voilà, T cells. Journal of Theoretical Biology, 207(2):195–216, November 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921647.

Nosaka:2006:GCI

[Nos06] Michiko Nosaka. Geometrical correspondence identified and a new interaction unit suggested in striated muscle. Journal of Theoretical Biology, 238(2):464–473, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002389.

Namy:2004:CCP

[NOT04] Patrick Namy, Jacques Ohayon, and Philippe Tracqui. Critical conditions for pattern formation and in vitro tubulogenesis driven by cellular traction fields. Journal of Theoretical Biology, 227(1):103–120, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003916.

Nowak:2000:BRR

[Now00] Martin A. Nowak. The basic reproductive ratio of a word, the maximum size of a lexicon. Journal of Theoretical Biology, 204(2):179–189, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300910853.

Nowlan:2005:EMB

[NP05] Niamh C. Nowlan and Patrick J. Prendergast. Evolution of mechanoregulation of bone growth will lead to non-optimal bone phenotypes. Journal of Theoretical Biology, 235(3):408–418, August 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000482.

Narang:2007:BGR

[NP07] Atul Narang and Sergei S. Pilyugin. Bacterial gene regulation in diauxic and non-diauxic growth. Journal of Theoretical Biology, 244(2):326–348, January 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600350X.

Neff:2008:MCN

[NP08] Bryan D. Neff and Trevor E. Pitcher. Mate choice for nonadditive genetic benefits: a resolution to the lek paradox. Journal of Theoretical Biology, 254(1):147–155, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002658.

Noel:2009:DIC

[NPN09] Jason T. Noel, Sergei S. Pilyugin, and Atul Narang. The diffusive influx and carrier efflux have a strong effect on the bistability of the lac operon in Escherichia coli. Journal of Theoretical Biology, 256(1):14–28, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004748.

Nakamaru:2003:CTI

[NS03a] Mayuko Nakamaru and Akira Sasaki. Can transitive inference evolve in animals playing the hawk-dove game? Journal of Theoretical Biology, 222(4):461–470, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000596.

Noldeke:2003:SCH

[NS03b] Georg Nöldeke and Larry Samuelson. Strategic choice handicaps when females seek high male net viability. Journal of Theoretical Biology, 221(1):53–59, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931699.

Nakabayashi:2005:OPS

[NS05] Jun Nakabayashi and Akira Sasaki. Optimal phosphorylation step number of intracellular signal-transduction pathway. Journal of Theoretical Biology, 233(3):413–421, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005028.

Nakabayashi:2006:MMA

[NS06] Jun Nakabayashi and Akira Sasaki. A mathematical model for apoptosome assembly: The optimal cytochrome c /Apaf-1 ratio. Journal of Theoretical Biology, 242(2):280–287, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000932.

Nishimura:2007:MRR

[NS07] Shin I. Nishimura and Masaki Sasai. Modulation of the reaction rate of regulating protein induces large morphological and motional change of amoebic cell. Journal of Theoretical Biology, 245(2):230–237, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004516.

Nakabayashi:2009:FTO

[NS09a] Jun Nakabayashi and Akira Sasaki. The function of temporally ordered viral gene expression in the intracellular replication of herpes simplex virus type 1 (HSV-1). Journal of Theoretical Biology, 261(1):156–164, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003439.

Nakabayashi:2009:MMS

[NS09b] Jun Nakabayashi and Akira Sasaki. A mathematical model of the stoichiometric control of smad complex formation in TGF-β signal transduction pathway. Journal of Theoretical Biology, 259(2):389–403, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001416.

Nolting:2008:PFF

[NSG08] Bengt Nölting, Neema Salimi, and Ulrich Guth. Protein folding forces. Journal of Theoretical Biology, 251(2):331–347, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005814.

Nolting:2003:SDR

[NSH^{+}03] Bengt Nölting, Wolfram Schälike, Patrick Hampel, Florian Grundig, Siegfried Gantert, Nicole Sips, Wolfhard Bandlow, and Phoebe X. Qi. Structural determinants of the rate of protein folding. Journal of Theoretical Biology, 223(3):299–307, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000912.

Naoki:2008:SCS

[NSI08] Honda Naoki, Yuichi Sakumura, and Shin Ishii. Stochastic control of spontaneous signal generation for gradient sensing in chemotaxis. Journal of Theoretical Biology, 255(2):259–266, November 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004153.

Nakayama:2002:PDM

[NSM02] Yuu Nakayama, Hiromi Seno, and Hiroyuki Matsuda. A population dynamic model for facultative agamosperms. Journal of Theoretical Biology, 215(2):253–262, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923732.

Napoletani:2008:ASR

[NSS^{+}08a] D. Napoletani, T. Sauer, D. C. Struppa, E. Petricoin, and L. Liotta. Augmented sparse reconstruction of protein signaling networks. Journal of Theoretical Biology, 255(1):40–52, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003925.

Newell:2008:PCC

[NSS08b] Alan C. Newell, Patrick D. Shipman, and Zhiying Sun. Phyllotaxis: Cooperation and competition between mechanical and biochemical processes. Journal of Theoretical Biology, 251(3):421–439, April 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307006030.

Novak:2004:MRP

[NT04] Béla Novák and John J. Tyson. A model for restriction point control of the mammalian cell cycle. Journal of Theoretical Biology, 230(4):563–579, October 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002449.

Nakayama:2002:IGS

[NTK02] Naofumi Nakayama, Syunsuke Tanaka, and Osamu Kikuchi. Ab initio GB study of methylation reaction of adenine, cytosine, guanine, and thymine by methanediazonium ion. Journal of Theoretical Biology, 215(1):13–22, March 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925007.

Nakayama:2006:FCE

[NTU06] Hideki Nakayama, Hiroto Tanaka, and Toshimitsu Ushio. The formulation of the control of an expression pattern in a gene network by propositional calculus. Journal of Theoretical Biology, 240(3):443–450, June 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004522.

Nevai:2007:SDC

[NWP07] Andrew L. Nevai, Thomas A. Waite, and Kevin M. Passino. State-dependent choice and ecological rationality. Journal of Theoretical Biology, 247(3):471–479, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001452.

Nakaoka:2009:EPC

[NWT09] Shinji Nakaoka, Wendi Wang, and Yasuhiro Takeuchi. Effect of parental care and aggregation on population dynamics. Journal of Theoretical Biology, 260(1):161–171, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002720.

Ofria:2003:SPG

[OAC03] Charles Ofria, Christoph Adami, and Travis C. Collier. Selective pressures on genomes in molecular evolution. Journal of Theoretical Biology, 222(4):477–483, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000626.

Okazaki:2008:SCM

[OAKC08] Noriaki Okazaki, Ryoji Asano, Takashi Kinoshita, and Hiroshi Chuman. Simple computational models of type I/type II cells in Fas signaling-induced apoptosis. Journal of Theoretical Biology, 250(4):621–633, February 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005255.

Orrell:2004:CIE

[OB04] David Orrell and Hamid Bolouri. Control of internal and external noise in genetic regulatory networks. Journal of Theoretical Biology, 230(3):301–312, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002516.

Okada:2008:HUS

[OB08] Daijiro Okada and Paul M. Bingham. Human uniquenessself-interest and social cooperation. Journal of Theoretical Biology, 253(2):261–270, July 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000970.

Olesen:2006:SAW

[OBDJ06] Jens M. Olesen, Jordi Bascompte, Yoko L. Dupont, and Pedro Jordano. The smallest of all worlds: Pollination networks. Journal of Theoretical Biology, 240(2):270–276, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004170.

Owen:2004:MMU

[OBL04] Markus R. Owen, Helen M. Byrne, and Claire E. Lewis. Mathematical modelling of the use of macrophages as vehicles for drug delivery to hypoxic tumour sites. Journal of Theoretical Biology, 226(4):377–391, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003412.

Ohtsuki:2007:OTL

[OBN07] Hisashi Ohtsuki, Pedro Bordalo, and Martin A. Nowak. The one-third law of evolutionary dynamics. Journal of Theoretical Biology, 249(2):289–295, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003311.

Ogden:2008:PEC

[OBPH^{+}08] N. H. Ogden, M. Bigras-Poulin, K. Hanincová, A. Maarouf, C. J. O’Callaghan, and K. Kurtenbach. Projected effects of climate change on tick phenology and fitness of pathogens transmitted by the North American tick ixodes scapularis. Journal of Theoretical Biology, 254(3):621–632, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003329.

Ortega:2008:KPR

[OCA08] Fernando Ortega, Marta Cascante, and Luis Acerenza. Kinetic properties required for sustained or paradoxical control of metabolic fluxes under large changes in enzyme activities. Journal of Theoretical Biology, 252(3):569–573, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005140.

Oda:2000:CBO

[OCP^{+}00] G. A. Oda, I. L. Caldas, J. R. C. Piqueira, J. M. Waterhouse, and M. D. Marques. Coupled biological oscillators in a cave insect. Journal of Theoretical Biology, 206(4):515–524, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921519.

Ottesen:2003:MVC

[OD03] J. T. Ottesen and M. Danielsen. Modeling ventricular contraction with heart rate changes. Journal of Theoretical Biology, 222(3):337–346, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000407.

Otaki:2001:LAM

[OF01] Joji M. Otaki and Stuart Firestein. Length analyses of mammalian G-protein-coupled receptors. Journal of Theoretical Biology, 211(2):77–100, July 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922726.

OBrien:2005:MPG

[OF05] Ricky O’Brien and Nev Fowkes. Modification patterns in germinating barley-malting II. Journal of Theoretical Biology, 233(3):315–325, April 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004813.

Overgaard:2005:IMM

[OHM05] R. V. Overgaard, J. E. Henriksen, and H. Madsen. Insights to the minimal model of insulin secretion through a mean-field beta cell model. Journal of Theoretical Biology, 237(4):382–389, December 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001876.

Ohtsuki:2004:RSI

[Oht04] Hisashi Ohtsuki. Reactive strategies in indirect reciprocity. Journal of Theoretical Biology, 227(3):299–314, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004260.

Ono:2000:SMS

[OI00] Naoaki Ono and Takashi Ikegami. Self-maintenance and selfreproduction in an abstract cell model. Journal of Theoretical Biology, 206(2):243–253, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921210.

Ohtsuki:2004:HSW

[OI04] Hisashi Ohtsuki and Yoh Iwasa. How should we define goodness? — reputation dynamics in indirect reciprocity. Journal of Theoretical Biology, 231(1):107–120, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002772. See erratum [OI05].

Ohtsuki:2005:EHS

[OI05] Hisashi Ohtsuki and Yoh Iwasa. Erratum to: “How should we define goodness? — reputation dynamics in indirect reciprocity” [Journal of Theoretical Biology 231 (2004) 107– 120]. Journal of Theoretical Biology, 232(4):451, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400462X. See [OI04].

Ohtsuki:2006:LES

[OI06] Hisashi Ohtsuki and Yoh Iwasa. The leading eight: Social norms that can maintain cooperation by indirect reciprocity. Journal of Theoretical Biology, 239(4):435–444, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003474.

Ohtsuki:2007:GAE

[OI07] Hisashi Ohtsuki and Yoh Iwasa. Global analyses of evolutionary dynamics and exhaustive search for social norms that maintain cooperation by reputation. Journal of Theoretical Biology, 244(3):518–531, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003900.

OKeefe:2005:EVP

[O’K05] Kara J. O’Keefe. The evolution of virulence in pathogens with frequency-dependent transmission. Journal of Theoretical Biology, 233(1):55–64, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004503.

Ostby:2004:MMG

[ØKRG04] Ivar Østby, Gunnar Kvalheim, Leiv S. Rusten, and Per Grøttum. Mathematical modeling of granulocyte reconstitution after high-dose chemotherapy with stem cell support: effect of post-transplant g-CSF treatment. Journal of Theoretical Biology, 231(1):69–83, November 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002607.

Otsuka:2001:ISE

[OKS01] Jinya Otsuka, Yosuke Kawai, and Nobuyoshi Sugaya. The influence of selection on the evolutionary distance estimated from the base changes observed between homologous nucleotide sequences. Journal of Theoretical Biology, 213(2):129–144, November 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924075.

Ohsawa:2002:RHD

[OKTS02] Kyoko Ohsawa, Kohkichi Kawasaki, Fugo Takasu, and Nanako Shigesada. Recurrent habitat disturbance and species diversity in a multiple-competitive species system. Journal of Theoretical Biology, 216(2):123–138, May 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925543.

OToole:2008:MMD

[OLBM08] Matthew O’Toole, Robert Latham, Rehan M. Baqri, and Kyle E. Miller. Modeling mitochondrial dynamics during in vivo axonal elongation. Journal of Theoretical Biology, 255(4):369–377, December 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004724.

OBrien:2002:TIA

[OLS^{+}02] Michael J. O’Brien, R. Lee Lyman, Youssef Saab, Elias Saab, John Darwent, and Daniel S. Glover. Two issues in archaeological phylogenetics: Taxon construction and outgroup selection. Journal of Theoretical Biology, 215(2):133–150, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925488.

Ochab-Marcinek:2008:PAR

[OM08] Anna Ochab-Marcinek. Predicting the asymmetric response of a genetic switch to noise. Journal of Theoretical Biology, 254(1):37–44, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002154.

Ono:2003:EGS

[OMT03] Seiji Ono, Kazuharu Misawa, and Kazuki Tsuji. Effect of group selection on the evolution of altruistic behavior. Journal of Theoretical Biology, 220(1):55–66, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931444.

Ohtsuki:2006:REG

[ON06] Hisashi Ohtsuki and Martin A. Nowak. The replicator equation on graphs. Journal of Theoretical Biology, 243(1):86–97, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002426.

Ohtsuki:2007:DRG

[ON07] Hisashi Ohtsuki and Martin A. Nowak. Direct reciprocity on graphs. Journal of Theoretical Biology, 247(3):462–470, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001300.

Ohtsuki:2008:ESG

[ON08] Hisashi Ohtsuki and Martin A. Nowak. Evolutionary stability on graphs. Journal of Theoretical Biology, 251(4):698–707, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000192.

Ogawa:2006:PMG

[OOHI06] Naoko Ogawa, Hiromasa Oku, Koichi Hashimoto, and Masatoshi Ishikawa. A physical model for galvanotaxis of paramecium cell. Journal of Theoretical Biology, 242(2): 314–328, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519306000944.

Ostby:2006:SMC

[ØØS06] Ivar Østby, Leiv Øyehaug, and Harald B. Steen. A stochastic model of cancer initiation including a bystander effect.

Journal of Theoretical Biology, 241(4):751–764, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000269.

ostborn:2001:SSE

[öOW01] Per östborn, Gunnar Ohlén, and Björn Wohlfart. Simulated sinoatrial exit blocks explained by circle map analysis. Journal of Theoretical Biology, 211(3):219–227, August 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923409.

Ozer:2005:CTV

[OP05] Mahmut Ozer and Ivo Provaznik. A comparative tool for the validity of rate kinetics in ion channels by Onsager reciprocity theorem. Journal of Theoretical Biology, 233(2):237–243, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930400476X.

Ohtsuki:2007:EGT

[OPN07] Hisashi Ohtsuki, Jorge M. Pacheco, and Martin A. Nowak. Evolutionary graph theory: Breaking the symmetry between interaction and replacement. Journal of Theoretical Biology, 246(4):681–694, June 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000562.

oyehaug:2002:RBM

[øPVO02] Leiv øyehaug, Erik Plahte, Dag I. Våge, and Stig W. Omholt. The regulatory basis of melanogenic switching. Journal of Theoretical Biology, 215(4):449–468, April 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925214.

Olivares-Quiroz:2007:PNS

[OQGC07] L. Olivares-Quiroz and L. S. Garcia-Colin. Protein’s native state stability in a chemically induced denaturation mechanism. Journal of Theoretical Biology, 246(2):214–224, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005844.

OConnor:2004:UIS

[OR04] S. C. O’Connor and P. A. Robinson. Unifying and interpreting the spectral wavenumber content of EEGs, ECoGs, and ERPs. Journal of Theoretical Biology, 231(3):397–412, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003248.

OConnor:2005:AEA

[OR05] S. C. O’Connor and P. A. Robinson. Analysis of the electroencephalographic activity associated with thalamic tumors. Journal of Theoretical Biology, 233(2):271–286, March 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004801.

OToole:2003:SOCb

[ORM03a] D. V. O’Toole, P. A. Robinson, and M. R. Myerscough. Self-organized criticality and emergent oscillations in models of termite architecture with crowding. Journal of Theoretical Biology, 221(1):15–27, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303930475.

OToole:2003:SOCa

[ORM03b] D. V. O’Toole, P. A. Robinson, and M. R. Myerscough. Self-organized criticality in ant brood tending. Journal of Theoretical Biology, 221(1):1–14, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303930463.

Orr:2003:MMN

[Orr03] H. Allen Orr. A minimum on the mean number of steps taken in adaptive walks. Journal of Theoretical Biology, 220(2):241–247, January 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931614.

Orr:2006:DFE

[Orr06] H. Allen Orr. The distribution of fitness effects among beneficial mutations in Fisher’s geometric model of adaptation. Journal of Theoretical Biology, 238(2):279–285, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002031.

Ortega:2006:SOA

[Ort06] Horacio Ortega. Some observations on the article “dynamic response of cancer under the influence of immunological activity and therapy”. Journal of Theoretical Biology, 240(1):160– 161, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004819.

Otsuka:2003:AFB

[OS03] Jinya Otsuka and Nobuyoshi Sugaya. Advanced formulation of base pair changes in the stem regions of ribosomal RNAs; its application to mitochondrial rRNAs for resolving the phylogeny of animals. Journal of Theoretical Biology, 222(4):447– 460, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000572.

Ohtsuki:2006:EDC

[OS06] Akiko Ohtsuki and Akira Sasaki. Epidemiology and diseasecontrol under gene-for-gene plant-pathogen interaction. Journal of Theoretical Biology, 238(4):780–794, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002882.

Osborn:2008:MGR

[Osb08] Jeffrey. W. Osborn. A model of growth restraints to explain the development and evolution of tooth shapes in mammals. Journal of Theoretical Biology, 255(3):338–343, December 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004633.

Ovaskainen:2002:MME

[OSBH02] Otso Ovaskainen, Kazunori Sato, Jordi Bascompte, and Ilkka Hanski. Metapopulation models for extinction threshold in spatially correlated landscapes. Journal of Theoretical Biology, 215(1):95–108, March 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301925020.

Olofsson:2001:AGB

[OSCK01] Peter Olofsson, Otto Schwalb, Ranajit Chakraborty, and Marek Kimmel. An application of a general branching process in the study of the genetics of aging. Journal of Theoretical Biology, 213(4):547–557, December 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192435X.

Overgaard:2008:NCC

[OSJ08] Morten Overgaard, Kristian Sandberg, and Mads Jensen. The neural correlate of consciousness? Journal of Theoretical Biology, 254(3):713–715, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003378.

Orive:2005:VII

[OSK^{+}05] Maria E. Orive, Miles N. Stearns, John K. Kelly, Michael Barfield, Marilyn S. Smith, and Robert D. Holt. Viral infection in internally structured hosts. I. Conditions for persistent infection. Journal of Theoretical Biology, 232(4):453–466, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004059.

Oster:2004:CPM

[Ost04] George Oster. Clocks and patterns in myxobacteria: a remembrance of Art Winfree. Journal of Theoretical Biology, 230(4):451–458, October 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400236X.

Okamoto:2009:SAG

[OST09] Makoto Okamoto, Shigeru Sunada, and Hiroshi Tokutake. Stability analysis of gliding flight of a swallowtail butterfly papilio xuthus. Journal of Theoretical Biology, 257(2):191–202, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006000.

Ohtsuka:2009:PDP

[OT09] Yasunori Ohtsuka and Yukihiko Toquenaga. The patch distributed producer-scrounger game. Journal of Theoretical Biology, 260(2):261–266, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002604.

Otsuka:2008:TAL

[Ots08] Jinya Otsuka. A theoretical approach to the large-scale evolution of multicellularity and cell differentiation. Journal of Theoretical Biology, 255(1):129–136, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003627.

Oudeyer:2005:SOS

[Oud05] Pierre-Yves Oudeyer. The self-organization of speech sounds. Journal of Theoretical Biology, 233(3):435–449, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005053.

Ozer:2009:SLJ

[OUPG09] Mahmut Ozer, Muhammet Uzuntarla, Matjaz Perc, and Lyle J. Graham. Spike latency and jitter of neuronal membrane patches with stochastic Hodgkin–Huxley channels. Journal of Theoretical Biology, 261(1):83–92, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003191.

Ovaskainen:2002:LTP

[Ova02] Otso Ovaskainen. Long-term persistence of species and the SLOSS problem. Journal of Theoretical Biology, 218 (4):419–433, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519302930894.

OHely:2007:MLL

[OW07] Martin O’Hely and Leesa Wockner. Models for loosely linked gene duplicates suggest lengthy persistence of both copies. Journal of Theoretical Biology, 246(4):636–645, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000574.

ostborn:2001:ARP

[öWO01] Per östborn, Björn Wohlfart, and Gunnar Ohlén. Arrhythmia as a result of poor intercellular coupling in the sinus node: a simulation study. Journal of Theoretical Biology, 211(3):201–217, August 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923392.

Otaki:2003:LAD

[OY03] Joji M. Otaki and Haruhiko Yamamoto. Length analyses of drosophila odorant receptors. Journal of Theoretical Biology, 223(1):27–37, July 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000687.

Petit:2009:PPT

[PA09a] Giai Petit and Tommaso Anfodillo. Plant physiology in theory and practice: An analysis of the WBE model for vascular plants. Journal of Theoretical Biology, 259(1):1–4, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001179.

Piotrowska:2009:CQC

[PA09b] Monika Joanna Piotrowska and Simon D. Angus. Corrigendum to “A quantitative cellular automaton model of in vitro multicellular spheroid tumour growth” [J. Theor. Biol. 258 (2009) 165–178]. Journal of Theoretical Biology, 259(4):864, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002525.

Piotrowska:2009:QCA

[PA09c] Monika Joanna Piotrowska and Simon D. Angus. A quantitative cellular automaton model of in vitro multicellular spheroid tumour growth. Journal of Theoretical Biology, 258(2):165– 178, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000678. See corri-

gendum [PA09b].

Pagnutti:2005:LGG

[PAA05] C. Pagnutti, M. Anand, and M. Azzouz. Lattice geometry, gap formation and scale invariance in forests. Journal of Theoretical Biology, 236(1):79–87, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000780.

Pagnutti:2007:PLI

[PAA07] C. Pagnutti, M. Azzouz, and M. Anand. Propagation of local interactions create global gap structure and dynamics in a tropical rainforest. Journal of Theoretical Biology, 247(1):168– 181, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000719.

Packard:2009:ULT

[Pac09] Gary C. Packard. On the use of logarithmic transformations in allometric analyses. Journal of Theoretical Biology, 257(3):515–518, April 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005468.

Powell:2000:FOA

[PAD00] David R. Powell, Lloyd Allison, and Trevor I. Dix. Fast, optimal alignment of three sequences using linear gap costs. Journal of Theoretical Biology, 207(3):325–336, December 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921775.

Pokhilko:2006:MMM

[PAH06] Alexandra V. Pokhilko, Fazoil I. Ataullakhanov, and Ekhson L. Holmuhamedov. Mathematical model of mitochondrial ionic homeostasis: Three modes of Ca^{2+} transport. Journal of Theoretical Biology, 243(1):152–169, November 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002219.

Palsson:2001:EDM

[Pál01] Snæbjörn Pálsson. The effects of deleterious mutations in cyclically parthenogenetic organisms. Journal of Theoretical Biology, 208(2):201–214, January 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922112.

Palsson:2008:DMU

[Pal08] Eirikur Palsson. A 3-D model used to explore how cell adhesion and stiffness affect cell sorting and movement in multicellular systems. Journal of Theoretical Biology, 254(1):1–13, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002385.

Panek:2008:CCA

[Pán08] Josef Pánek. A comparative computational analysis of protein sequences and literature mining classify ‘orphan’ neurotransmitter transporters. Journal of Theoretical Biology, 254(2):301–307, September 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002993.

Paradis:2004:CER

[Par04] Emmanuel Paradis. Can extinction rates be estimated without fossils? Journal of Theoretical Biology, 229(1):19–30, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304000797.

Patel:2005:TGC

[Pat05] Apoorva Patel. The triplet genetic code had a doublet predecessor. Journal of Theoretical Biology, 233(4):527–532, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400517X.

Portet:2003:OCN

[PAVS03] Stéphanie Portet, Ovide Arino, Jany Vassy, and Damien Schoëvaërt. Organization of the cytokeratin network in an epithelial cell. Journal of Theoretical Biology, 223(3):313–333, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001012.

Pawlowitsch:2007:FPC

[Paw07a] Christina Pawlowitsch. Finite populations choose an optimal language. Journal of Theoretical Biology, 249(3):606–616, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003840.

Pawlowski:2007:MMC

[Paw07b] Piotr H. Pawlowski. Mechanokinetic model of cell membrane: Theoretical analysis of plasmalemma homeostasis, growth and division. Journal of Theoretical Biology, 249(1):67–76, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003268. See corri-

gendum [Paw09b].

Pawar:2009:CAS

[Paw09a] Samraat Pawar. Community assembly, stability and signatures of dynamical constraints on food web structure. Journal of Theoretical Biology, 259(3):601–612, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001660.

Pawlowski:2009:CMM

[Paw09b] Piotr H. Pawlowski. Corrigendum to “Mechanokinetic model of cell membrane: Theoretical analysis of plasmalemma homeostasis, growth and division”: [J. Theor. Biol. 249 (2007)

67–76]. Journal of Theoretical Biology, 257(3):526, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930800653X. See [Paw07b].

Penney:2002:MNB

[PB02] M. S. Penney and N. F. Britton. Modelling natural burst firing in nigral dopamine neurons. Journal of Theoretical Biology, 219(2):207–223, November 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931164.

Panchanathan:2003:TTD

[PB03] Karthik Panchanathan and Robert Boyd. A tale of two defectors: the importance of standing for evolution of indirect reciprocity. Journal of Theoretical Biology, 224(1):115–126, September 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303001541.

Prasad:2006:SPN

[PB06] B. R. Guru Prasad and Renee M. Borges. Searching on patch networks using correlated random walks: Space usage and optimal foraging predictions using Markov chain models. Journal of Theoretical Biology, 240(2):241–249, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004091.

Purcell:2007:FIR

[PB07] Jessica Purcell and Anders Brodin. Factors influencing route choice by avian migrants: a dynamic programming model of Pacific brant migration. Journal of Theoretical Biology, 249(4):804–816, December 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004080.

Peischl:2008:EDU

[PB08] Stephan Peischl and Reinhard Bürger. Evolution of dominance under frequency-dependent intraspecific competition. Journal of Theoretical Biology, 251(2):210–226, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005735.

Powell:2009:ESP

[PB09] Craig R. Powell and Richard P. Boland. The effects of stochastic population dynamics on food web structure. Journal of Theoretical Biology, 257(1):170–180, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800581X.

Powers:2003:FOD

[PBB03] S. J. Powers, P. Brain, and P. W. Barlow. First-order differential equation models with estimable parameters as functions of environmental variables and their application to a study of vascular development in young hybrid aspen stems. Journal of Theoretical Biology, 222(2):219–232, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000304.

Parker:2009:GGI

[PBC09a] G. A. Parker, M. A. Ball, and J. C. Chubb. To grow or not to grow? Intermediate and paratenic hosts as helminth life cycle strategies. Journal of Theoretical Biology, 258(1):135– 147, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000071.

Parker:2009:WDL

[PBC09b] G. A. Parker, M. A. Ball, and J. C. Chubb. Why do larval helminths avoid the gut of intermediate hosts? Journal of Theoretical Biology, 260(3):460–473, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002811.

Poitrineau:2004:JED

[PBH04] Karine Poitrineau, Sam P. Brown, and Michael E. Hochberg. The joint evolution of defence and inducibility against natural enemies. Journal of Theoretical Biology, 231(3):389–396, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003236.

Panet:2005:ICD

[PBHS05] Amos Panet, Efrat Braun, Alik Honigman, and Israel Steiner. Involvement of cellular death signals in the reactivation of herpes simplex virus type 1 and lambda bacteriophage from a latent state. Journal of Theoretical Biology, 236(1):88–94, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000792.

Pedersen:2006:DRM

[PBL06] Lee G. Pedersen, Lee Bartolotti, and Leping Li. Deuterium and its role in the machinery of evolution. Journal of Theoretical Biology, 238(4):914–918, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003048.

Perez-Bello:2009:AFP

[PBMU^{+}09] Alcides Perez-Bello, Cristian Robert Munteanu, Florencio M. Ubeira, Alexandre Lopes De Magalhães, Eugenio Uriarte, and Humberto González-Díaz. Alignment-free prediction of mycobacterial DNA promoters based on pseudo-folding lattice network or star-graph topological indices. Journal of Theoretical Biology, 256(3):458–466, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005055.

Proctor:2001:MAV

[PBR01] C. J. Proctor, M. Broom, and G. D. Ruxton. Modelling antipredator vigilance and flight response in group foragers when warning signals are ambiguous. Journal of Theoretical Biology, 211(4):409–417, August 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923537.

Proctor:2003:CBS

[PBR03] Carole J. Proctor, Mark Broom, and Graeme D. Ruxton. A communication-based spatial model of antipredator vigilance. Journal of Theoretical Biology, 220(1):123–137, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931596.

Provorov:2002:DGE

[PBT02] Nikolai A. Provorov, Alex Yu. Borisov, and Igor A. Tikhonovich. Developmental genetics and evolution of symbiotic structures in nitrogen-fixing nodules and arbuscular mycorrhiza. Journal of Theoretical Biology, 214(2):215–232, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924531.

Peletier:2009:IPP

[PBvdG09] Lambertus A. Peletier, Neil Benson, and Piet H. van der Graaf. Impact of plasma-protein binding on receptor occupancy: An analytical description. Journal of Theoretical Biology, 256(2):253–262, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004700.

Pawlowski:2006:TMR

[PBZ06] Piotr H. Pawlowski, Beata Burzyńska, and Piotr Zielenkiewicz. Theoretical model of reticulocyte to erythrocyte shape transformation. Journal of Theoretical Biology, 243(1):24–38, November 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002499.

Pawlowski:2008:TMT

[PBZ08] Piotr H. Pawlowski, Beata Burzyńska, and Piotr Zielenkiewicz. Theoretical model of thalassemic erythrocyte shape transformation. Journal of Theoretical Biology, 254(3):575–579, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003275.

Paradis:2002:ACD

[PC02] Emmanuel Paradis and Julien Claude. Analysis of comparative data using generalized estimating equations. Journal of Theoretical Biology, 218(2):175–185, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930663.

Poyatos:2004:NNR

[PC04] Juan F. Poyatos and Amancio Carnero. Non-neutral role of replicative senescence in tissue homeostasis and tumorigenesis. Journal of Theoretical Biology, 230(3):333–341, October 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002590.

Pigolotti:2009:SRD

[PC09a] Simone Pigolotti and Massimo Cencini. Speciation-rate dependence in species-area relationships. Journal of Theoretical Biology, 260(1):83–89, September 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002483.

Pradhan:2009:DVD

[PC09b] Ranjan K. Pradhan and V. S. Chakravarthy. Desynchronized vasomotion and desynchronized fiber activation pattern enhance oxygenation in a model of skeletal muscle. Journal of Theoretical Biology, 259(2):242–252, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000782.

Poletti:2009:SBC

[PCA^{+}09] Piero Poletti, Bruno Caprile, Marco Ajelli, Andrea Pugliese, and Stefano Merler. Spontaneous behavioural changes in response to epidemics. Journal of Theoretical Biology, 260(1):31–40, September 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309002021.

Pal:2007:NMW

[PCB07] Amita Pal, Arunabha Chakrabarti, and Jolly Basak. New motifs within the NB-ARC domain of R proteins: Probable mechanisms of integration of geminiviral signatures within the host species of Fabaceae family and implications in conferring disease resistance. Journal of Theoretical Biology, 246(3):564– 573, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000306.

P:2005:CBS

[PCH^{+}05] J. Barral P., L. Cantini, A. Hasmy, J. Jiménez, and A. Marcano. Correlation between strand asymmetry and phylogeny in mitochondrial DNA. Journal of Theoretical Biology, 236(4):422–426, October 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001311.

Pritchard:2005:GME

[PCK^{+}05] Leighton Pritchard, Dave Corne, Douglas Kell, Jem Rowland, and Mike Winson. A general model of error-prone PCR. Journal of Theoretical Biology, 234(4):497–509, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304006071.

Pearce:2006:MST

[PCS^{+}06] Ian G. Pearce, Mark A. J. Chaplain, Peter G. Schofield, Alexander R. A. Anderson, and Stephen F. Hubbard. Modelling the spatio-temporal dynamics of multi-species hostparasitoid interactions: Heterogeneous patterns and ecological implications. Journal of Theoretical Biology, 241(4):876–886, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600035X.

Peercy:2006:KMO

[PCSL^{+}06] Bradford E. Peercy, Steven J. Cox, Sagit Shalel-Levanon, Ka-Yiu San, and George Bennett. A kinetic model of oxygen regulation of cytochrome production in Escherichia coli. Journal of Theoretical Biology, 242(3):547–563, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001603.

Pei:2005:EPO

[PCZL05] Yongzhen Pei, Lansun Chen, Qingrui Zhang, and Changguo Li. Extinction and permanence of one-prey multi-predators of Holling type II function response system with impulsive biological control. Journal of Theoretical Biology, 235(4):495–503, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000561. See corri-

gendum [PCZL06].

Pei:2006:CEP

[PCZL06] Yongzhen Pei, Lansun Chen, Qingrui Zhang, and Changguo Li. Corrigendum to “Extinction and permanence of one-prey multi-predators of Holling type II function response system with impulsive biological control”: [J. Theor. Biol. 235 (2005) 495–503]. Journal of Theoretical Biology, 239(3):400, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003632. See [PCZL05].

Pritchard:2000:DPL

[PD00] Leighton Pritchard and Mark J. Dufton. Do proteins learn to evolve? The Hopfield network as a basis for the understanding of protein evolution. Journal of Theoretical Biology, 202(1):77–86, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910433.

Powell:2001:RAB

[PD01] P. D. Powell and J. C. Demartini. Retroviral antibody binding of the MHC class II molecule: a biochemical influence on CD4 T cell differentiation in HIV infection? Journal of Theoretical Biology, 209(2):155–160, March 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922458.

Petermann:2004:CAE

[PD04] Thomas Petermann and Paolo De Los Rios. Cluster approximations for epidemic processes: a systematic description of correlations beyond the pair level. Journal of Theoretical Biology, 229(1):1–11, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000773.

Park:2006:RPL

[PD06] Eun-Hyoung Park and Dominique M. Durand. Role of potassium lateral diffusion in non-synaptic epilepsy: a computational study. Journal of Theoretical Biology, 238(3):666–682, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002730.

Powell:2000:ESS

[PDA^{+}00] P. D. Powell, J. C. Demartini, P. Azari, L. A. Stargell, L. Cordain, and A. Tucker. Evolutionary stable strategy: a test for theories of retroviral pathology which are based upon the concept of molecular mimicry. Journal of Theoretical Biology, 202(3):213–229, February 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251939991055X.

Pappu:2008:CSL

[PDB08] Vijay Pappu, Sai K. Doddi, and Prosenjit Bagchi. A computational study of leukocyte adhesion and its effect on flow pattern in microvessels. Journal of Theoretical Biology, 254(2):483–498, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800266X.

Pound:2002:LVM

[PDC02] Graeme E. Pound, C. Patrick Doncaster, and Simon J. Cox. A Lotka–Volterra model of coexistence between a sexual population and multiple asexual clones. Journal of Theoretical Biology, 217(4):535–545, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930407.

Pavoine:2004:DAS

[PDC04] Sandrine Pavoine, Anne-Béatrice Dufour, and Daniel Chessel. From dissimilarities among species to dissimilarities among communities: a double principal coordinate analysis. Journal of Theoretical Biology, 228(4):523–537, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000736.

Peluso:2000:PME

[PDIS00] Andrea Peluso, Mariangela Di Donato, Roberto Improta, and Gloria Anna Ada Saracino. A plausible mechanism of electron transfer between quinones in photosynthetic reaction centers. Journal of Theoretical Biology, 207(1):101–105, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921593.

Pearson:2004:MIS

[PDM04] G. A. Pearson, Scott Dillery, and J. R. Meyer. Modeling intra-sexual competition in a sex pheromone system: how much can female movement affect female mating success? Journal of Theoretical Biology, 231(4):549–555, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003315.

Peet:2005:CDT

[PDPL05] Alison B. Peet, Peter A. Deutsch, and Enrique Peacock-López. Complex dynamics in a three-level trophic system with intraspecies interaction. Journal of Theoretical Biology, 232(4):491–503, February 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004163.

Powell:2000:OTS

[PE00] James A. Powell and Katharina A. M. Engelhardt. Optimal trajectories for the short-distance foraging flights of swans. Journal of Theoretical Biology, 204(3):415–430, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920265.

Pirone:2004:FTF

[PE04] Jason R. Pirone and Timothy C. Elston. Fluctuations in transcription factor binding can explain the graded and binary responses observed in inducible gene expression. Journal of Theoretical Biology, 226(1):111–121, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003278.

Pecou:2005:SDL

[Péc05] Elisabeth Pécou. Splitting the dynamics of large biochemical interaction networks. Journal of Theoretical Biology, 232(3):375–384, February 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930400400X.

Peck:2006:ASI

[Pec06] Joel R. Peck. Altruism, sex, and inbreeding when the genotype-phenotype map is additive. Journal of Theoretical Biology, 239(2):130–140, March 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003516.

Pedersen:2005:CNE

[Ped05] Morten Gram Pedersen. A comment on noise enhanced bursting in pancreatic β-cells. Journal of Theoretical Biology, 235(1):1–3, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305000524.

Pedersen:2007:PBH

[Ped07] Morten Gram Pedersen. Phantom bursting is highly sensitive to noise and unlikely to account for slow bursting in βcells: Considerations in favor of metabolically driven oscillations. Journal of Theoretical Biology, 248(2):391–400, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002755.

Pelletier:2000:MAO

[Pel00] Jon D. Pelletier. Model assessments of the optimal design of nature reserves for maximizing species longevity. Journal of Theoretical Biology, 202(1):25–32, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910305.

Pennycuick:2003:CEH

[Pen03] C. J. Pennycuick. The concept of energy height in animal locomotion: separating mechanics from physiology. Journal of Theoretical Biology, 224(2):189–203, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001577.

Pepper:2000:RTG

[Pep00] John W. Pepper. Relatedness in trait group models of social evolution. Journal of Theoretical Biology, 206(3):355–368, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921325.

Peper:2004:TDTa

[Pep04a] Abraham Peper. A theory of drug tolerance and dependence i: a conceptual analysis. Journal of Theoretical Biology, 229(4):477–490, August 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001560.

Peper:2004:TDTb

[Pep04b] Abraham Peper. A theory of drug tolerance and dependence II: the mathematical model. Journal of Theoretical Biology, 229(4):491–500, August 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001572.

Prescott:2003:TGR

[PER03] David M. Prescott, A. Ehrenfeucht, and G. Rozenberg. Template-guided recombination for IES elimination and unscrambling of genes in stichotrichous ciliates. Journal of Theoretical Biology, 222(3):323–330, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000377.

Piqueira:2007:BQB

[PFdP^{+}07] J. R. C. Piqueira, G. M. Fagali, M. de Pinho, R. F. Sanches, C. M. Del-Ben, and A. W. Zuardi. Bits and q-bits in a psychiatric ward. Journal of Theoretical Biology, 247(1):182–185, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000884.

Piazza:2008:DMM

[PFRR08] Matthew Piazza, Xiao-Jiang Feng, Joshua D. Rabinowitz, and Herschel Rabitz. Diverse metabolic model parameters generate similar methionine cycle dynamics. Journal of Theoretical Biology, 251(4):628–639, April 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006364.

Porcher:2000:QDP

[PG00] Emmanuelle Porcher and Marino Gatto. Quantifying the dynamics of prion infection: a bifurcation analysis of Laurent’s model. Journal of Theoretical Biology, 205(2):283–296, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092068X.

Pereira:2001:TTA

[PG01a] Henrique M. Pereira and Deborah M. Gordon. A trade-off in task allocation between sensitivity to the environment and response time. Journal of Theoretical Biology, 208(2):165–184, January 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922100.

Pittendrigh:2001:PRC

[PG01b] B. R. Pittendrigh and P. J. Gaffney. Pesticide resistance: Can we make it a renewable resource? Journal of Theoretical Biology, 211(4):365–375, August 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923598.

Prats:2008:AIS

[PGF^{+}08] Clara Prats, Antoni Giró, Jordi Ferrer, Daniel López, and Josep Vives-Rego. Analysis and IbM simulation of the stages in bacterial lag phase: Basis for an updated definition. Journal of Theoretical Biology, 252(1):56–68, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000337.

Pittendrigh:2004:ARC

[PGH^{+}04] Barry R. Pittendrigh, Patrick J. Gaffney, Joseph E. Huesing, David W. Onstad, Richard T. Roush, and Larry L. Murdock. “active” refuges can inhibit the evolution of resistance in insects towards transgenic insect-resistant plants. Journal of Theoretical Biology, 231(4):461–474, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400253X.

Patel:2001:CAM

[PGLG01] Aalpen A. Patel, Edward T. Gawlinski, Susan K. Lemieux, and Robert A. Gatenby. A cellular automaton model of early tumor growth and invasion: The effects of native tissue vascularity and increased anaerobic tumor metabolism. Journal of Theoretical Biology, 213(3):315–331, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923859.

Pu:2007:PMP

[PGLL07] Xian Pu, Jian Guo, Howard Leung, and Yuanlie Lin. Prediction of membrane protein types from sequences and position-specific scoring matrices. Journal of Theoretical Biology, 247(2):259–265, July 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700032X.

Pittendrigh:2000:DMN

[PGM00] B. R. Pittendrigh, P. Gaffney, and L. L. Murdock. Deterministic modeling of negative cross-resistance strategies for use in transgenic host-plant resistance. Journal of Theoretical Biology, 204(1):135–150, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092006X.

Pilyugin:2003:RMQ

[PGMK^{+}03] Sergei S. Pilyugin, Vitaly V. Ganusov, Kaja Murali-Krishna, Rafi Ahmed, and Rustom Antia. The rescaling method for quantifying the turnover of cell populations. Journal of Theoretical Biology, 225(2):275–283, November 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002455.

Pliquett:2008:PHW

[PGN08] Uwe Pliquett, Ch. Gusbeth, and Richard Nuccitelli. A propagating heat wave model of skin electroporation. Journal of Theoretical Biology, 251(2):195–201, March 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005930.

Pojoga:2003:TEA

[PH03] L. H. Pojoga and R. H. Hilderman. A theoretical and experimental approach to the use of single wavelength calcium indicators. Journal of Theoretical Biology, 221(4):543–557, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932036.

Polynikis:2009:CDO

[PHdB09] A. Polynikis, S. J. Hogan, and M. di Bernardo. Comparing different ODE modelling approaches for gene regulatory networks. Journal of Theoretical Biology, 261(4):511–530, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900352X.

Perkins:2004:IMG

[PHG04] Theodore J. Perkins, Mike Hallett, and Leon Glass. Inferring models of gene expression dynamics. Journal of Theoretical Biology, 230(3):289–299, October 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002486.

Philipson:2002:FMB

[Phi02] Lars Philipson. Functional modules of the brain. Journal of Theoretical Biology, 215(1):109–119, March 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925019.

Puszynski:2008:OBS

[PHL08] Krzysztof Puszyński, Beata Hat, and Tomasz Lipniacki. Oscillations and bistability in the stochastic model of p53 regulation. Journal of Theoretical Biology, 254(2):452–465, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002919.

Penny:2003:TFE

[PHP03] David Penny, Michael D. Hendy, and Anthony M. Poole. Testing fundamental evolutionary hypotheses. Journal of Theoretical Biology, 223(3):377–385, August 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000997.

Pickard:2006:AMS

[Pic06] William F. Pickard. Absorption by a moving spherical organelle in a heterogeneous cytoplasm: Implications for the role of trafficking in a symplast. Journal of Theoretical Biology, 240(2):288–301, May 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500411X.

Pietak:2009:DLR

[Pie09a] Alexis Mari Pietak. Describing long-range patterns in leaf vasculature by metaphoric fields. Journal of Theoretical Biology, 261(2):279–289, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003695.

Pietruszka:2009:GPV

[Pie09b] M. Pietruszka. General proof of the validity of a new tensor equation of plant growth. Journal of Theoretical Biology, 256(4):584–585, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005821.

Pinsky:2000:MSM

[Pin00] Paul F. Pinsky. A multi-stage model of adenoma development. Journal of Theoretical Biology, 207(2):129–143, November 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921489.

Payne:2001:UBT

[PJ01] Robert J. H. Payne and Vincent A. A. Jansen. Understanding bacteriophage therapy as a density-dependent kinetic process. Journal of Theoretical Biology, 208(1):37–48, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921982.

Purvis:2002:EPT

[PKA02] Andy Purvis, Aris Katzourakis, and Paul-Michael Agapow. Evaluating phylogenetic tree shape: Two modifications to fusco and Cronk’s method. Journal of Theoretical Biology, 214(1):99–103, January 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924439.

Pertoldi:2001:NME

[PKL01] Cino Pertoldi, Torsten Nygaard Kristensen, and Volker Loeschcke. A new method for estimating environmental variability for clonal organisms, and the use of fluctuating asymmetry as an indicator of developmental instability. Journal of Theoretical Biology, 210(4):407–410, June 21, 2001.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923173.

Pineda-Krch:2002:CLD

[PKL02] Mario Pineda-Krch and Kari Lehtilä. Cell lineage dynamics in stratified shoot apical meristems. Journal of Theoretical Biology, 219(4):495–505, December 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931395.

Pirhaji:2008:PCS

[PKS^{+}08] Leila Pirhaji, Mehdi Kargar, Armita Sheari, Hadi Poormohammadi, Mehdi Sadeghi, Hamid Pezeshk, and Changiz Eslahchi. The performances of the chi-square test and complexity measures for signal recognition in biological sequences. Journal of Theoretical Biology, 251(2):380–387, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307005802.

Powell:2000:MEA

[PKW^{+}00] James Powell, Bruce Kennedy, Peter White, Barbara Bentz, Jesse Logan, and David Roberts. Mathematical elements of attack risk analysis for mountain pine beetles. Journal of Theoretical Biology, 204(4):601–620, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910804.

Petrovskii:2001:ICB

[PL01] Sergei Petrovskii and Bai-Lian Li. Increased coupling between subpopulations in a spatially structured environment can lead to population outbreaks. Journal of Theoretical Biology, 212(4):549–562, October 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923938.

Pitzer:2009:ERB

[PL09] Virginia E. Pitzer and Marc Lipsitch. Exploring the relationship between incidence and the average age of infection during seasonal epidemics. Journal of Theoretical Biology, 260(2):175–185, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002690.

Paszek:2005:SEM

[PLB^{+}05] Paweł Paszek, Tomasz Lipniacki, Allan R. Brasier, Bing Tian, David E. Nowak, and Marek Kimmel. Stochastic effects of multiple regulators on expression profiles in eukaryotes. Journal of Theoretical Biology, 233(3):423–433, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930400503X.

Prats:2006:IBM

[PLG^{+}06] Clara Prats, Daniel López, Antoni Giró, Jordi Ferrer, and Joaquim Valls. Individual-based modelling of bacterial cultures to study the microscopic causes of the lag phase. Journal of Theoretical Biology, 241(4):939–953, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000427.

Patel:2005:FOV

[PLH05] Rajan Patel, Ira M. Longini, and M. Elizabeth Halloran. Finding optimal vaccination strategies for pandemic influenza using genetic algorithms. Journal of Theoretical Biology, 234(2):201– 212, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005703.

Porco:2005:ETP

[PLSGG05] Travis C. Porco, James O. Lloyd-Smith, Kimber L. Gross, and Alison P. Galvani. The effect of treatment on pathogen virulence. Journal of Theoretical Biology, 233(1):91–102, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304004655.

Pluhar:2006:AAP

[Plu06] Wolfgang Pluhar. AT2–AT3-profiling: a new look at synonymous codon usage. Journal of Theoretical Biology, 243(3):308–321, December 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002955.

Price:2003:SCE

[PM03] Mary V. Price and John E. Mittler. Seed-cache exchange promotes coexistence and coupled consumer oscillations: a model of desert rodents as resource processors. Journal of Theoretical Biology, 223(2):215–231, July 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000882.

Petie:2007:CFP

[PM07] Ronald Petie and Mees Muller. Curvature facilitates prey fixation in predatory insect claws. Journal of Theoretical Biology, 244(4):565–575, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600405X.

Powell:2008:PSA

[PM08] Craig R. Powell and Alan J. McKane. Predicting the species abundance distribution using a model food web. Journal of Theoretical Biology, 255(4):387–395, December 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004645.

Preece:2009:SDH

[PM09] T. Preece and Y. Mao. Sustainability of dioecious and hermaphrodite populations on a lattice. Journal of Theoretical Biology, 261(2):336–340, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003737.

Prevost:2006:MSI

[PMB06] K. Prévost, P. Magal, and C. Beaumont. A model of salmonella infection within industrial house hens. Journal of Theoretical Biology, 242(3):755–763, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600169X.

Page:2001:MPS

[PMMS01] Karen M. Page, Philip K. Maini, Nicholas A. M. Monk, and Claudio D. Stern. A model of primitive streak initiation in the chick embryo. Journal of Theoretical Biology, 208(4):419–438, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930092229X.

Plotkin:2000:LEI

[PN00] Joshua B. Plotkin and Martin A. Nowak. Language evolution and information theory. Journal of Theoretical Biology, 205(1):147–159, July 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920538.

Page:2001:GAD

[PN01] Karen M. Page and Martin A. Nowak. A generalized adaptive dynamics framework can describe the evolutionary ultimatum game. Journal of Theoretical Biology, 209(2):173–179, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922513.

Page:2002:UED

[PN02a] Karen M. Page and Martin A. Nowak. Unifying evolutionary dynamics. Journal of Theoretical Biology, 219(1):93–98, November 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931127.

Plotkin:2002:DEA

[PN02b] Joshua B. Plotkin and Martin A. Nowak. The different effects of apoptosis and DNA repair on tumorigenesis. Journal of Theoretical Biology, 214(3):453–467, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924713.

Perry:2003:FCC

[PNG03] Laura G. Perry, Claudia Neuhauser, and Susan M. Galatowitsch. Founder control and coexistence in a simple model of asymmetric competition for light. Journal of Theoretical Biology, 222(4):425–436, June 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000559.

Pohlmeyer:2008:GCR

[Poh08] Roland Pohlmeyer. The genetic code revisited. Journal of Theoretical Biology, 253(3):623–624, August 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002117.

Papin:2004:TAM

[PP04] Jason A. Papin and Bernhard O. Palsson. Topological analysis of mass-balanced signaling networks: a framework to obtain network properties including crosstalk. Journal of Theoretical Biology, 227(2):283–297, March 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004259.

Popescu:2008:WPL

[PP08] Dumitru Popescu and Alin Gabriel Popescu. The working of a pulsatory liposome. Journal of Theoretical Biology, 254(3):515–519, October 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003640.

Percus:2009:AVB

[PPD09] Jerrome K. Percus, Ora E. Percus, and Michele Di Mascio. The amplitudes of viral blips in HIV-1 infected patients treated with antiretroviral therapy are power-law distributed. Journal of Theoretical Biology, 257(3):454–459, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000022.

Papin:2002:GSM

[PPEøP02] Jason A. Papin, Nathan D. Price, Jeremy S. Edwards, and Bernhard ø. Palsson. The genome-scale metabolic extreme pathway structure in haemophilus influenzae shows significant network redundancy. Journal of Theoretical Biology, 215(1):67–82, March 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924993.

Pradhan:2003:SAL

[PPGS03] Gauri R. Pradhan, Sagar A. Pandit, Anil D. Gangal, and V. Sitaramam. Shape anisotropy of lipid molecules and voids. Journal of Theoretical Biology, 220(2):189–199, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303931559.

Plotkin:2000:SAC

[PPL^{+}00] Joshua B. Plotkin, Matthew D. Potts, Nandi Leslie, N. Manokaran, James Lafrankie, and Peter S. Ashton. Speciesarea curves, spatial aggregation, and habitat specialization in tropical forests. Journal of Theoretical Biology, 207(1):81–99, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921581.

Paar:2001:LBT

[PPR01] V. Paar, N. Pavin, and M. Rosandić. Link between truncated fractals and coupled oscillators in biological systems. Journal of Theoretical Biology, 212(1):47–56, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923343.

Plucinski:2008:CFA

[PPRI08] Mateusz Pluciński, Szymon Pluciński, and Ignacio Rodríguez-Iturbe. Consequences of the fractal architecture of trees on their structural measures. Journal of Theoretical Biology, 251(1):82–92, March 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005462.

Phillips:2008:SDQ

[PR08] A. J. K. Phillips and P. A. Robinson. Sleep deprivation in a quantitative physiologically based model of the ascending arousal system. Journal of Theoretical Biology, 255(4):413–423, December 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800444X.

Pross:2003:DFL

[Pro03] Addy Pross. The driving force for Life’s emergence: Kinetic and thermodynamic considerations. Journal of Theoretical Biology, 220(3):393–406, February 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930393178X.

Price:2003:NBA

[PRP^{+}03] Nathan D. Price, Jennifer L. Reed, Jason A. Papin, Sharon J. Wiback, and Bernhard O. Palsson. Network-based analysis of metabolic regulation in the human red blood cell. Journal of Theoretical Biology, 225(2):185–194, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002376.

Pie:2004:NAA

[PRT04] Marcio R. Pie, Rebeca B. Rosengaus, and James F. A. Traniello. Nest architecture, activity pattern, worker density and the dynamics of disease transmission in social insects. Journal of Theoretical Biology, 226(1):45–51, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003096.

Painter:2003:MMI

[PS03] Kevin J. Painter and Jonathan A. Sherratt. Modelling the movement of interacting cell populations. Journal of Theoretical Biology, 225(3):327–339, December 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002583.

Plank:2004:MMT

[PSJ04] M. J. Plank, B. D. Sleeman, and P. F. Jones. A mathematical model of tumour angiogenesis, regulated by vascular endothelial growth factor and the angiopoietins. Journal of Theoretical Biology, 229(4):435–454, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001535.

Piqueira:2006:BMM

[PSM06] J. R. C. Piqueira, F. A. Serboncini, and L. H. A. Monteiro. Biological models: Measuring variability with classical and quantum information. Journal of Theoretical Biology, 242(2):309–313, September 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000919.

Poolman:2007:MDM

[PSPF07] Mark G. Poolman, Cristiana Sebu, Michael K. Pidcock, and David A. Fell. Modular decomposition of metabolic systems via null-space analysis. Journal of Theoretical Biology, 249(4):691–705, December 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003700.

Pastor-Satorras:2003:EPI

[PSSS03] Romualdo Pastor-Satorras, Eric Smith, and Ricard V. Solé. Evolving protein interaction networks through gene duplication. Journal of Theoretical Biology, 222(2):199–210, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000286.

Pidaparti:2009:OHB

[PSSY09] Ramana M. Pidaparti, David V. Svintradze, YingFeng Shan, and Hiroki Yokota. Optimization of hydrogen bonds for combined DNA/collagen complex. Journal of Theoretical Biology, 256(2):149–156, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004669.

Prado:2009:CCH

[PSWK09] Federico Prado, Alyssa Sheih, Jevin D. West, and Benjamin Kerr. Coevolutionary cycling of host sociality and pathogen virulence in contact networks. Journal of Theoretical Biology, 261(4):561–569, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003920.

Peeters:2009:GTI

[PT09] Kasper Peeters and Anne Taormina. Group theory of icosahedral virus capsid vibrations: a top-down approach. Journal of Theoretical Biology, 256(4):607–624, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005432.

Pacheco:2009:ACM

[PTD09] Jorge M. Pacheco, Arne Traulsen, and David Dingli. The allometry of chronic myeloid leukemia. Journal of Theoretical Biology, 259(3):635–640, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001623.

Plaetzer:2005:MEE

[PTFF05] Kristjan Plaetzer, S. Randall Thomas, Renate Falkner, and Gernot Falkner. The microbial experience of environmental phosphate fluctuations. An essay on the possibility of putting intentions into cell biochemistry. Journal of Theoretical Biology, 235(4):540–554, August 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000676.

Pacheco:2006:ALE

[PTN06] Jorge M. Pacheco, Arne Traulsen, and Martin A. Nowak. Active linking in evolutionary games. Journal of Theoretical Biology, 243(3):437–443, December 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002736.

Pacheco:2008:RGD

[PTON08] Jorge M. Pacheco, Arne Traulsen, Hisashi Ohtsuki, and Martin A. Nowak. Repeated games and direct reciprocity under active linking. Journal of Theoretical Biology, 250(4):723–731, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005450.

Pujol:2002:CMT

[Puj02] Toni Pujol. The consequence of maximum thermodynamic efficiency in daisyworld. Journal of Theoretical Biology, 217(1):53–60, July 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930171.

Provorov:2000:PGR

[PV00] Nikolai A. Provorov and Nikolai I. Vorobyov. Population genetics of rhizobia: Construction and analysis of an “infection and release” model. Journal of Theoretical Biology, 205(1):105–119, July 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920514.

Pienaar:2008:TFM

[PV08] Elsje Pienaar and Hendrik J. Viljoen. The tri-frame model. Journal of Theoretical Biology, 251(4):616–627, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006200.

Patwa:2009:IHC

[PW09a] Z. Patwa and L. M. Wahl. The impact of host-cell dynamics on the fixation probability for lytic viruses. Journal of Theoretical Biology, 259(4):799–810, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002173.

Perelson:2009:SAL

[PW09b] Alan S. Perelson and Frederik W. Wiegel. Scaling aspects of lymphocyte trafficking. Journal of Theoretical Biology, 257(1):9–16, March 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005869.

Perumal:2009:DAC

[PWG09] Thanneer M. Perumal, Yan Wu, and Rudiyanto Gunawan. Dynamical analysis of cellular networks based on the Green’s function matrix. Journal of Theoretical Biology, 261(2):248–259, November 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900349X.

Peletier:2003:CSH

[PWK03] Mark A. Peletier, Hans V. Westerhoff, and Boris N. Kholodenko. Control of spatially heterogeneous and time-varying cellular reaction networks: a new summation law. Journal of Theoretical Biology, 225(4):477–487, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002893.

Pienaar:2009:MCR

[PWVvR09] Elsje Pienaar, Scott E. Whitney, Hendrik J. Viljoen, and Nicolaas F. J. van Rensburg. A model of the complex response of staphylococcus aureus to methicillin. Journal of Theoretical Biology, 257(3):438–445, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800636X.

Pei:2009:PFM

[PWZ09] Yunfei Pei, Xi Wang, and Xuegong Zhang. Predicting the fate of microRNA target genes based on sequence features. Journal of Theoretical Biology, 261(1):17–22, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003361.

Quesneville:2001:GAB

[QA01] Hadi Quesneville and Dominique Anxolabéhère. Genetic algorithm-based model of evolutionary dynamics of class II transposable elements. Journal of Theoretical Biology, 213(1):21–30, November 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924014.

Quince:2008:BGFa

[QASL08] Christopher Quince, Peter A. Abrams, Brian J. Shuter, and Nigel P. Lester. Biphasic growth in fish i: Theoretical foundations. Journal of Theoretical Biology, 254(2):197–206, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002774.

Qian:2000:CFB

[QB00] Hong Qian and James B. Bassingthwaighte. A class of flow bifurcation models with lognormal distribution and fractal dispersion. Journal of Theoretical Biology, 205(2):261–268, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920605.

Quayle:2006:MEG

[QB06] A. P. Quayle and S. Bullock. Modelling the evolution of genetic regulatory networks. Journal of Theoretical Biology, 238(4):737–753, February 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500278X.

Qian:2009:LRS

[QD09] Xiaoning Qian and Edward R. Dougherty. On the longrun sensitivity of probabilistic Boolean networks. Journal of Theoretical Biology, 257(4):560–577, April 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006590.

Qu:2007:COP

[QHF^{+}07] K. Qu, A. Abi Haidar, J. Fan, L. Ensman, K. Tuncay, M. Jolly, and P. Ortoleva. Cancer onset and progression: a genomewide, nonlinear dynamical systems perspective on onconetworks. Journal of Theoretical Biology, 246(2):234–244, May 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005613.

Qiu:2009:USV

[QLHL09] Jian-Ding Qiu, San-Hua Luo, Jian-Hua Huang, and Ru-Ping Liang. Using support vector machines to distinguish enzymes: Approached by incorporating wavelet transform. Journal of Theoretical Biology, 256(4):625–631, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308005584.

Qu:2008:USC

[QO08] K. Qu and P. Ortoleva. Understanding stem cell differentiation through self-organization theory. Journal of Theoretical Biology, 250(4):606–620, February 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005152.

Quince:2008:BGFb

[QSAL08] Christopher Quince, Brian J. Shuter, Peter A. Abrams, and Nigel P. Lester. Biphasic growth in fish II: Empirical assessment. Journal of Theoretical Biology, 254(2):207–214, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002786.

Qi:2007:NGR

[QWQ07] Xiao-Qin Qi, Jie Wen, and Zhao-Hui Qi. New 3d graphical representation of DNA sequence based on dual nucleotides. Journal of Theoretical Biology, 249(4):681–690, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004122.

Qi:2009:CAD

[QWQ09] Zhao-Hui Qi, Jian-Min Wang, and Xiao-Qin Qi. Classification analysis of dual nucleotides using dimension reduction. Journal of Theoretical Biology, 260(1):104–109, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002367.

Qian:2006:MBD

[QXDW06] Jun Qian, Ping Xie, Shuo-Xing Dou, and Peng-Ye Wang. A model for biased diffusion of collagenase along collagen fibrils. Journal of Theoretical Biology, 243(3):322–327, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930600230X.

Ruszczycky:2006:IVK

[RA06] Mark W. Ruszczycky and Vernon E. Anderson. Interpretation of V/K isotope effects for enzymatic reactions exhibiting multiple isotopically sensitive steps. Journal of Theoretical Biology, 243(3):328–342, December 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002608.

Rakshit:2008:AAM

[RA08] Sourav Rakshit and G. K. Ananthasuresh. An amino acid map of inter-residue contact energies using metric multi-dimensional scaling. Journal of Theoretical Biology, 250(2):291–297, January 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700450X.

Radulescu:2008:SPG

[Rǎd08] Anca Rǎdulescu. Schizophrenia — a parameters’ game? Journal of Theoretical Biology, 254(1):89–98, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002415.

Radulescu:2009:MEM

[Rǎd09] Anca Rǎdulescu. A multi-etiology model of systemic degeneration in schizophrenia. Journal of Theoretical Biology, 259(2):269–279, July 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001246.

Raeymaekers:2002:DBN

[Rae02] L. Raeymaekers. Dynamics of Boolean networks controlled by biologically meaningful functions. Journal of Theoretical Biology, 218(3):331–341, October 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930293081X.

Raffa:2002:HMR

[Raf02] Robert B. Raffa. Harmonic mean relationship between affinity for wild-type receptors and Alanine-scan mutants. Journal of Theoretical Biology, 218(2):207–214, September 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930699.

Ruiz:2006:TAC

[RAHO06] Lina María Ruiz, Gemma Armengol, Edwin Habeych, and Sergio Orduz. A theoretical analysis of codon adaptation index of the boophilus microplus bm86 gene directed to the optimization of a DNA vaccine. Journal of Theoretical Biology, 239(4):445–449, April 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003486.

Rakocevic:2004:HSG

[Rak04] Miloje M. Rakocević. A harmonic structure of the genetic code. Journal of Theoretical Biology, 229(2):221–234, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001365.

Rezaei:2008:PMP

[RAK^{+}08] Mohammad Ali Rezaei, Parviz Abdolmaleki, Zahra Karami, Ebrahim Barzegari Asadabadi, Mohammad Amin Sherafat, Hamid Abrishami-Moghaddam, Marziyeh Fadaie, and Mohammad Forouzanfar. Prediction of membrane protein types by means of wavelet analysis and cascaded neural networks. Journal of Theoretical Biology, 254(4):817–820, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308003597.

Rapoport:2008:RHU

[Rap08] Tom A. Rapoport. Reinhart Heinrich — an unassuming intellectual giant. Journal of Theoretical Biology, 252(3):388– 390, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003001.

Rabinovitch:2003:BDE

[RAZ03] Avinoam Rabinovitch, Ira Aviram, and Arieh Zaritsky. Bacterial debris — an ecological mechanism for coexistence of bacteria and their viruses. Journal of Theoretical Biology, 224(3):377–383, October 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001747.

Ramakrishnan:2002:IBM

[RB02a] Vivek Ramakrishnan and David D. Busath. An inverting basket model for AE1 transport. Journal of Theoretical Biology, 215(2):215–226, March 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924907.

Regoes:2002:HCU

[RB02b] Roland R. Regoes and Sebastian Bonhoeffer. HIV coreceptor usage and drug treatment. Journal of Theoretical Biology, 217(4):443–457, August 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930493.

Root-Bernstein:2002:MCI

[RB02c] Robert Root-Bernstein. Molecular complementarity III. Peptide complementarity as a basis for peptide receptor evolution: a bioinformatic case study of insulin, glucagon and gastrin. Journal of Theoretical Biology, 218(1):71–84, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930560.

Richards:2006:IDS

[RB06] Kelvin J. Richards and Stuart J. Brentnall. The impact of diffusion and stirring on the dynamics of interacting populations. Journal of Theoretical Biology, 238(2):340–347, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500247X.

Ruxton:2008:AGA

[RB08] Graeme D. Ruxton and Guy Beauchamp. The application of genetic algorithms in behavioural ecology, illustrated with a model of anti-predator vigilance. Journal of Theoretical Biology, 250(3):435–448, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005176.

Reynolds:2009:OSR

[RB09a] A. M. Reynolds and F. Bartumeus. Optimising the success of random destructive searches: Lévy walks can outperform ballistic motions. Journal of Theoretical Biology, 260(1):98–103, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002495.

Rohlf:2009:MCR

[RB09b] Thimo Rohlf and Stefan Bornholdt. Morphogenesis by coupled regulatory networks: Reliable control of positional information and proportion regulation. Journal of Theoretical Biology, 261(2):176–193, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003397.

Root-Bernstein:2002:DNR

[RBBH02] Robert Root-Bernstein, Julia V. Busik, and Douglas N. Henry. Are diabetic neuropathy, retinopathy and nephropathy caused by hyperglycemic exclusion of dehydroascorbate uptake by glucose transporters? Journal of Theoretical Biology, 216(3):345– 359, June 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930292535X.

Rokyta:2006:PAW

[RBJ06] Darin R. Rokyta, Craig J. Beisel, and Paul Joyce. Properties of adaptive walks on uncorrelated landscapes under strong selection and weak mutation. Journal of Theoretical Biology, 243(1):114–120, November 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002475.

Ratel:2001:PCD

[RBN^{+}01] David Ratel, Sylvie Boisseau, Valéry Nasser, François Berger, and Didier Wion. Programmed cell death or cell death programme? That is the question. Journal of Theoretical Biology, 208(3):385–386, February 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922185.

Russo:2009:LPS

[RBP^{+}09] Tommaso Russo, Paolo Baldi, Antonio Parisi, Giuseppe Magnifico, Stefano Mariani, and Stefano Cataudella. Lévy processes and stochastic von Bertalanffy models of growth, with application to fish population analysis. Journal of Theoretical Biology, 258(4):521–529, June 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000290.

Ruina:2005:CME

[RBS05] Andy Ruina, John E. A. Bertram, and Manoj Srinivasan. A collisional model of the energetic cost of support work qualitatively explains leg sequencing in walking and galloping, pseudo-elastic leg behavior in running and the walk-to-run transition. Journal of Theoretical Biology, 237(2):170–192, November 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305001591.

Rafikov:2009:OCS

[RBW09] M. Rafikov, L. Bevilacqua, and A. P. P. Wyse. Optimal control strategy of malaria vector using genetically modified mosquitoes. Journal of Theoretical Biology, 258(3):418–425, June 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308004190.

Rojdestvenski:2000:MSP

[RC00] Igor Rojdestvenski and Michael G. Cottam. Mapping of statistical physics to information theory with application to biological systems. Journal of Theoretical Biology, 202(1):43–54, January 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910421.

Ribeiro:2009:ELA

[RC09] Fabiano Ribeiro and Nestor Caticha. Emergence and loss of assortative mating in sympatric speciation. Journal of Theoretical Biology, 258(3):465–477, June 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006292.

Radulescu:2009:HEL

[RCA09] Anca Rǎdulescu, Kingsley Cox, and Paul Adams. Hebbian errors in learning: An analysis using the Oja model. Journal of Theoretical Biology, 258(4):489–501, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000423.

Ruoff:2005:PTM

[RCS05] Peter Ruoff, Melinda K. Christensen, and Vijay K. Sharma. PER/TIM-mediated amplification, gene dosage effects and temperature compensation in an interlocking-feedback loop model of the drosophila circadian clock. Journal of Theoretical Biology, 237(1):41–57, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001517.

Reinecke:2007:CMM

[RD07] Isabel Reinecke and Peter Deuflhard. A complex mathematical model of the human menstrual cycle. Journal of Theoretical Biology, 247(2):303–330, July 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001166.

Rios-Doria:2009:QAL

[RDC09] D. Rios-Doria and G. Chowell. Qualitative analysis of the level of cross-protection between epidemic waves of the 1918–1919 influenza pandemic. Journal of Theoretical Biology, 261(4):584–592, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003919.

Riley:2003:RPE

[RDF03] Steven Riley, Christl A. Donnelly, and Neil M. Ferguson. Robust parameter estimation techniques for stochastic within-host macroparasite models. Journal of Theoretical Biology, 225(4):419–430, December 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002662.

Rapaport:2009:LRC

[RDH09] Alain Rapaport, Denis Dochain, and Jérôme Harmand. Long run coexistence in the chemostat with multiple species. Journal of Theoretical Biology, 257(2):252–259, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006073.

Ridolfi:2007:VDI

[RDL07] Luca Ridolfi, Paolo D’Odorico, and Francesco Laio. Vegetation dynamics induced by phreatophyte-aquifer interactions. Journal of Theoretical Biology, 248(2):301–310, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002172.

Ruiz-Diez:2003:DIW

[RDSB^{+}03] Beatriz Ruiz-Díez, Patricia Sánchez, Fernando Baquero, José L. Martínez, and Alfonso Navas. Differential interactions within the Caenorhabditis elegans–Pseudomonas aeruginosa pathogenesis model. Journal of Theoretical Biology, 225(4):469–476, December 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002881.

Ribeiro:2009:VDD

[RDYH09] Andre S. Ribeiro, Xiaofeng Dai, and Olli Yli-Harja. Variability of the distribution of differentiation pathway choices regulated by a multipotent delayed stochastic switch. Journal of Theoretical Biology, 260(1):66–76, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002513.

Reinhold:2004:IMP

[RE04] K. Reinhold and L. Engqvist. Inaccurate mimicry and predator ecology. Journal of Theoretical Biology, 229(4):433–434, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001808.

Recordati:2002:VNS

[Rec02] Giorgio Recordati. The visceral nervous system and its environments. Journal of Theoretical Biology, 214(2):293–304, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924610.

Refinetti:2004:NST

[Ref04] Roberto Refinetti. Non-stationary time series and the robustness of circadian rhythms. Journal of Theoretical Biology, 227(4):571–581, April 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300451X.

Reick:2002:TNC

[Rei02a] Christian H. Reick. Two notions of conspicuity and the classification of phyllotaxis. Journal of Theoretical Biology, 215(3):263–271, April 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302925270.

Reilly:2002:MPD

[Rei02b] W. Ian Reilly. Magnetic position determination by homing pigeons? Journal of Theoretical Biology, 218(1):47–54, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930584.

Rejniak:2007:IBF

[Rej07] Katarzyna A. Rejniak. An immersed boundary framework for modelling the growth of individual cells: An application to the early tumour development. Journal of Theoretical Biology, 247(1):186–204, July 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000951.

Reluga:2004:TPE

[Rel04] Tim Reluga. A two-phase epidemic driven by diffusion. Journal of Theoretical Biology, 229(2):249–261, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001389.

Retsky:2004:LE

[Ret04] Michael Retsky. Letter to the editor. Journal of Theoretical Biology, 229(3):289, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001596.

Revilla:2000:RCS

[Rev00] Tomás A. Revilla. Resource competition in stage-structured populations. Journal of Theoretical Biology, 204(2):289–298, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920174.

Revilla:2002:EIP

[Rev02] Tomás Revilla. Effects of intraguild predation on resource competition. Journal of Theoretical Biology, 214(1):49–62, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924488.

Roose:2004:EMW

[RF04a] T. Roose and A. C. Fowler. Erratum to “A model for water uptake by plant roots”. Journal of Theoretical Biology, 229(3):291, August 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001766. See [RF04c].

Roose:2004:MMW

[RF04b] T. Roose and A. C. Fowler. A mathematical model for water and nutrient uptake by plant root systems. Journal of Theoretical Biology, 228(2):173–184, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300465X.

Roose:2004:MWU

[RF04c] T. Roose and A. C. Fowler. A model for water uptake by plant roots. Journal of Theoretical Biology, 228(2):155–171, May 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004648. See erra-

tum [RF04a].

Rabinovitch:2002:BTD

[RFH^{+}02] Avinoam Rabinovitch, Itzhak Fishov, Hilla Hadas, Monica Einav, and Arieh Zaritsky. Bacteriophage T4 development in Escherichia coli is growth rate dependent. Journal of Theoretical Biology, 216(1):1–4, May 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925439.

Rezaei-Ghaleh:2003:SBS

[RG03] N. Rezaei-Ghaleh. Signaling in biological systems: insights from communication theory. Journal of Theoretical Biology, 224(3):411–412, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001711.

Recker:2005:MPP

[RG05] Mario Recker and Sunetra Gupta. A model for pathogen population structure with cross-protection depending on the extent of overlap in antigenic variant repertoires. Journal of Theoretical Biology, 232(3):363–373, February 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003996.

Roeder:2006:TUL

[RG06] Ingo Roeder and Ingmar Glauche. Towards an understanding of lineage specification in hematopoietic stem cells: a mathematical model for the interaction of transcription factors GATA-1 and PU.1. Journal of Theoretical Biology, 241(4):852–865, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000348.

Rost:2006:SAI

[RGB06] Evelyn Rost, Ralf Geske, and Michael Baake. Signal analysis of impulse response functions in MR and CT measurements of cerebral blood flow. Journal of Theoretical Biology, 240(3):451–458, June 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004443.

Radde:2008:MFL

[RGF^{+}08] Nicole Radde, Jutta Gebert, Ulrich Faigle, Rainer Schrader, and Karin Schnetz. Modeling feedback loops in the h-NS-mediated regulation of the Escherichia coli bgl operon. Journal of Theoretical Biology, 250(2):298–306, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004602.

Rong:2007:MWH

[RGFP07] Libin Rong, Michael A. Gilchrist, Zhilan Feng, and Alan S. Perelson. Modeling within-host HIV-1 dynamics and the evolution of drug resistance: Trade-offs between viral enzyme function and drug susceptibility. Journal of Theoretical Biology, 247(4):804–818, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001968.

Roussel:2000:MDF

[RGG00] Marc R. Roussel, Didier Gonze, and Albert Goldbeter. Modeling the differential fitness of cyanobacterial strains whose circadian oscillators have different free-running periods: Comparing the mutual inhibition and substrate depletion hypotheses. Journal of Theoretical Biology, 205(2):321–340, July 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920721.

Rosandic:2008:RAH

[RGPB08] Marija Rosandić, Matko Gluncić, Vladimir Paar, and Ivan Basar. The role of alphoid higher order repeats (HORs) in the centromere folding. Journal of Theoretical Biology, 254(3):555–560, October 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002981.

Roberts:2004:PED

[RGRB04] Stephen Roberts, Tim Guilford, Iead Rezek, and Dora Biro. Positional entropy during pigeon homing i: application of Bayesian latent state modelling. Journal of Theoretical Biology, 227(1):39–50, March 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003072.

Rodriguez-Gonzalez:2007:SCM

[RGSFM07] J. G. Rodríguez-González, M. Santillán, A. C. Fowler, and Michael C. Mackey. The segmentation clock in mice: Interaction between the wnt and notch signalling pathways. Journal of Theoretical Biology, 248(1):37–47, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002226.

Rodriguez:2005:MAR

[RGW^{+}05] Mauricio Rodríguez, Theresa A. Good, Melinda E. Wales, Jean P. Hua, and James R. Wild. Modeling allosteric regulation of de novo pyrimidine biosynthesis in Escherichia coli. Journal of Theoretical Biology, 234(3):299–310, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005776.

Risau-Gusman:2009:CSC

[RGZ09] Sebastián Risau-Gusman and Damián H. Zanette. Contact switching as a control strategy for epidemic outbreaks. Journal of Theoretical Biology, 257(1):52–60, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005766.

Reed:2002:SDL

[RH02] William J. Reed and Barry D. Hughes. On the size distribution of live genera. Journal of Theoretical Biology, 217(1):125– 135, July 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930092.

Ruxton:2004:OVS

[RH04] Graeme D. Ruxton and David C. Houston. Obligate vertebrate scavengers must be large soaring fliers. Journal of Theoretical Biology, 228(3):431–436, June 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000542.

Rohwer:2008:ICR

[RH08] Johann M. Rohwer and Jan-Hendrik S. Hofmeyr. Identifying and characterising regulatory metabolites with generalised supply-demand analysis. Journal of Theoretical Biology, 252(3):546–554, June 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005280.

Rhodes:2009:CVD

[RH09a] C. J. Rhodes and T. D. Hollingsworth. Corrigendum to “Variational data assimilation with epidemic models” [J. Theor. Biol. 258 (2009) 591–602]. Journal of Theoretical Biology, 260(4):599, October 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003294.

Rhodes:2009:VDA

[RH09b] C. J. Rhodes and T. D. Hollingsworth. Variational data assimilation with epidemic models. Journal of Theoretical Biology, 258(4):591–602, June 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000794.

Ratnieks:2007:MIS

[RHF07] Francis L. W. Ratnieks, Heikki Helanterä, and Kevin R. Foster. Are mistakes inevitable? Sex allocation specialization by workers can reduce the genetic information needed to assess queen mating frequency. Journal of Theoretical Biology, 244(3):470–477, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004048.

Rands:2000:PPC

[RHG00] Sean A. Rands, Alasdair I. Houston, and Catherine E. Gasson. Prey processing in central place foragers. Journal of Theoretical Biology, 202(2):161–174, January 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910482.

Roy:2008:GLM

[RHH08] Manojit Roy, Karin Harding, and Robert D. Holt. Generalizing Levins metapopulation model in explicit space: Models of intermediate complexity. Journal of Theoretical Biology, 255(1):152–161, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003834.

Redding:2008:EDS

[RHM^{+}08] David W. Redding, Klaas Hartmann, Aki Mimoto, Drago Bokal, Matt DeVos, and Arne Ø. Mooers. Evolutionarily distinctive species often capture more phylogenetic diversity than expected. Journal of Theoretical Biology, 251(4):606–615, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307006285.

Ricotta:2003:PEM

[Ric03] Carlo Ricotta. On parametric evenness measures. Journal of Theoretical Biology, 222(2):189–197, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000262.

Riley:2000:TKS

[Ril00] Patrick A. Riley. Tyrosinase kinetics: a semi-quantitative model of the mechanism of oxidation of monohydric and dihydric phenolic substrates. Journal of Theoretical Biology, 203(1):1–12, March 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519399910615.

Riley:2002:RLC

[Ril02] Patrick A. Riley. Response to letter from cabanes et al. Journal of Theoretical Biology, 214(2):327–328, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924774.

Rashidi:2005:NMA

[RK05] Armin Rashidi and Iman Khodarahmi. Nonlinear modeling of the atrioventricular node physiology in atrial fibrillation. Journal of Theoretical Biology, 232(4):545–549, February 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004217.

Ray:2006:RMA

[RK06] J. Christian J. Ray and Denise E. Kirschner. Requirement for multiple activation signals by anti-inflammatory feedback in macrophages. Journal of Theoretical Biology, 241(2):276–294, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005187.

Ribeiro:2007:NAE

[RK07] Andre S. Ribeiro and Stuart A. Kauffman. Noisy attractors and ergodic sets in models of gene regulatory networks. Journal of Theoretical Biology, 247(4):743–755, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002020.

Rashidi:2006:MMC

[RKF06] Armin Rashidi, Iman Khodarahmi, and Marc D. Feldman. Mathematical modeling of the course and prognosis of factitious disorders: a game-theoretic approach. Journal of Theoretical Biology, 240(1):48–53, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003644.

Rothnie:2006:FRM

[RKH^{+}06] P. Rothnie, D. Kabaso, P. R. Hof, B. I. Henry, and S. L. Wearne. Functionally relevant measures of spatial complexity in neuronal dendritic arbors. Journal of Theoretical Biology, 238(3):505–526, February 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002353.

Ritchie:2008:SLI

[RKH08] Graham R. S. Ritchie, Simon Kirby, and David J. C. Hawkey. Song learning as an indicator mechanism: Modelling the developmental stress hypothesis. Journal of Theoretical Biology, 251(4):570–583, April 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006078.

Rosenwald:2002:TSM

[RKL02] Shai Rosenwald, Ran Kafri, and Doron Lancet. Test of a statistical model for molecular recognition in biological repertoires. Journal of Theoretical Biology, 216(3):327–336, June 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925385.

Ryndin:2001:PMP

[RKNK01] A. Ryndin, V. Kirzhner, E. Nevo, and A. Korol. Polymorphism maintenance in populations with mixed random mating and apomixis subjected to stabilizing and cyclical selection. Journal of Theoretical Biology, 212(2):169–181, September 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923653.

Rodrigues:2009:GTM

[RKOS09] António Rodrigues, Heinz Koeppl, Hisashi Ohtsuki, and Akiko Satake. A game theoretical model of deforestation in human-environment relationships. Journal of Theoretical Biology, 258(1):127–134, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000083.

Roy:2008:NMA

[RKRW08] Sandip Roy, James M. Krueger, David M. Rector, and Yan Wan. A network model for activity-dependent sleep regulation. Journal of Theoretical Biology, 253(3):462–468, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001689.

Ramo:2006:PAC

[RKYH06] P. Rämö, J. Kesseli, and O. Yli-Harja. Perturbation avalanches and criticality in gene regulatory networks. Journal of Theoretical Biology, 242(1):164–170, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000762.

Rolland-Lagan:2005:CMI

[RLCIB05] Anne-Gaëlle Rolland-Lagan, Enrico Coen, Stephen J. Impey, and J. Andrew Bangham. A computational method for inferring growth parameters and shape changes during development based on clonal analysis. Journal of Theoretical Biology, 232(2):157–177, January 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003054.

Rauch:2004:RTM

[RM04] Erik M. Rauch and Mark M. Millonas. The role of transmembrane signal transduction in turing-type cellular pattern formation. Journal of Theoretical Biology, 226(4):401–407, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003588.

Rossberg:2006:FWE

[RMAI06a] A. G. Rossberg, H. Matsuda, T. Amemiya, and K. Itoh. Food webs: Experts consuming families of experts. Journal of Theoretical Biology, 241(3):552–563, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519305005606. See corrigendum [RMAI09].

Rossberg:2006:SPS

[RMAI06b] A. G. Rossberg, H. Matsuda, T. Amemiya, and K. Itoh. Some properties of the speciation model for food-web structure — mechanisms for degree distributions and intervality. Journal of Theoretical Biology, 238(2):401–415, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002432.

Rossberg:2009:CFW

[RMAI09] A. G. Rossberg, H. Matsuda, T. Amemiya, and K. Itoh. Corrigendum to “Food webs: Experts consuming families of experts” [J. Theor. Biol. 241 (2006) 552–563]. Journal of Theoretical Biology, 258(2):335, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900023X. See [RMAI06a].

Rachev:2000:MSI

[RMBM00] A. Rachev, E. Manoach, J. Berry, and J. E. Moore. A model of stress-induced geometrical remodeling of vessel segments adjacent to stents and artery/graft anastomoses. Journal of Theoretical Biology, 206(3):429–443, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930092143X.

Reichenbach:2008:SOM

[RMF08] Tobias Reichenbach, Mauro Mobilia, and Erwin Frey. Selforganization of mobile populations in cyclic competition. Journal of Theoretical Biology, 254(2):368–383, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002464.

Reluga:2008:BBM

[RMP08] Timothy C. Reluga, Jan Medlock, and Alan S. Perelson. Backward bifurcations and multiple equilibria in epidemic models with structured immunity. Journal of Theoretical Biology, 252(1):155–165, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000271.

Rodrigues:2009:HSI

[RMRG09] Paula Rodrigues, Alessandro Margheri, Carlota Rebelo, and M. Gabriela M. Gomes. Heterogeneity in susceptibility to infection can explain high reinfection rates. Journal of Theoretical Biology, 259(2):280–290, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900126X.

Raine:2007:LDB

[RN07] D. J. Raine and V. Norris. Lipid domain boundaries as prebiotic catalysts of peptide bond formation. Journal of Theoretical Biology, 246(1):176–185, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005893.

Reeves:2004:GMC

[RNP04] Gregory T. Reeves, Atul Narang, and Sergei S. Pilyugin. Growth of mixed cultures on mixtures of substitutable substrates: the operating diagram for a structured model. Journal of Theoretical Biology, 226(2):143–157, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002807.

Ranganath:2004:MIM

[RNS04] R. M. Ranganath, R. Nijagunaiah, and G. Shubhakara. Meiosis-II and mitosis have different molecular formats. Journal of Theoretical Biology, 227(3):445–447, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004375.

Reed:2004:MMM

[RNSU04] Michael C. Reed, H. Frederik Nijhout, Rachel Sparks, and Cornelia M. Ulrich. A mathematical model of the methionine cycle. Journal of Theoretical Biology, 226(1):33–43, January 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003084.

Robinson:2003:NTC

[Rob03] P. A. Robinson. Neurophysical theory of coherence and correlations of electroencephalographic and electrocorticographic signals. Journal of Theoretical Biology, 222(2):163–175, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000237.

Roes:2007:OOC

[Roe07] Frans L. Roes. Outside opportunities and costs incurred by others. Journal of Theoretical Biology, 247(2):365–370, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001324.

Roes:2008:NAA

[Roe08] Frans L. Roes. A note on ants and aphids. Journal of Theoretical Biology, 250(2):388, January 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004808.

Rand:2009:DRC

[RON09] David G. Rand, Hisashi Ohtsuki, and Martin A. Nowak. Direct reciprocity with costly punishment: Generous tit-for-tat prevails. Journal of Theoretical Biology, 256(1):45–57, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004876.

Ratkowsky:2005:UTE

[ROR05] David A. Ratkowsky, June Olley, and Tom Ross. Unifying temperature effects on the growth rate of bacteria and the stability of globular proteins. Journal of Theoretical Biology, 233(3):351–362, April 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004941.

Rosenberger:2003:CCL

[Ros03] R. F. Rosenberger. The choice of cell lineages during the in vitro differentiation of mammalian embryonic stem cells. Journal of Theoretical Biology, 223(3):387–389, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001048.

Rosenberg:2005:SMM

[Ros05] Noah A. Rosenberg. A sharp minimum on the mean number of steps taken in adaptive walks. Journal of Theoretical Biology, 237(1):17–22, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001463.

Roth:2004:AWB

[Rot04] Bradley J. Roth. Art Winfree and the bidomain model of cardiac tissue. Journal of Theoretical Biology, 230(4):445–449, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003540.

Rothman:2007:IVT

[Rot07] Stephen Rothman. The incoherence of the vesicle theory of protein secretion. Journal of Theoretical Biology, 245(1):150– 160, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004723.

Rothman:2008:RLP

[Rot08] Stephen Rothman. Reply to the letter of Professors Brusch and Deutsch. Journal of Theoretical Biology, 252(2):374–375, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000519.

Rothschild:2009:ERS

[Rot09] Casey G. Rothschild. The evolution of reciprocity in sizable human groups. Journal of Theoretical Biology, 257(4):609–617, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000095.

Rousset:2003:MDC

[Rou03] François Rousset. A minimal derivation of convergence stability measures. Journal of Theoretical Biology, 221(4):665–668, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932103.

Rowthorn:2006:EAB

[Row06] Robert Rowthorn. The evolution of altruism between siblings: Hamilton’s rule revisited. Journal of Theoretical Biology, 241(4):774–790, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000270.

Roy:2000:LTG

[Roy00] Duncan Roy. Learning and the theory of games. Journal of Theoretical Biology, 204(3):409–414, June 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920228.

Riabinina:2009:MVD

[RP09a] Olena Riabinina and Andrew O. Philippides. A model of visual detection of angular speed for bees. Journal of Theoretical Biology, 257(1):61–72, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005791.

Rong:2009:MHP

[RP09b] Libin Rong and Alan S. Perelson. Modeling HIV persistence, the latent reservoir, and viral blips. Journal of Theoretical Biology, 260(2):308–331, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002665.

Rosandic:2003:KSS

[RPB03] M. Rosandić, V. Paar, and I. Basar. Key-string segmentation algorithm and higher-order repeat 16mer (54 copies) in human alpha satellite DNA in chromosome 7. Journal of Theoretical Biology, 221(1):29–37, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931651.

Rosa:2003:TDP

[RPNH03] Roberto Rosà, Andrea Pugliese, Rachel Norman, and Peter J. Hudson. Thresholds for disease persistence in models for tick-borne infections including non-viraemic transmission, extended feeding and tick aggregation. Journal of Theoretical Biology, 224(3):359–376, October 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001735.

Roberts:2008:MAS

[RR08] James A. Roberts and Peter A. Robinson. Modeling absence seizure dynamics: Implications for basic mechanisms and measurement of thalamocortical and corticothalamic latencies. Journal of Theoretical Biology, 253(1):189–201, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001185.

Rensing:2009:HCY

[RR09] Ludger Rensing and Peter Ruoff. How can yeast cells decide between three activated MAP kinase pathways? A model approach. Journal of Theoretical Biology, 257(4):578–587, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309000198.

Reynolds:2006:RMM

[RRC^{+}06] Angela Reynolds, Jonathan Rubin, Gilles Clermont, Judy Day, Yoram Vodovotz, and G. Bard Ermentrout. A reduced mathematical model of the acute inflammatory response: I. Derivation of model and analysis of anti-inflammation. Journal of Theoretical Biology, 242(1):220–236, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000816.

Robinson:2008:ABM

[RRH08] Elva J. H. Robinson, Francis L. W. Ratnieks, and M. Holcombe. An agent-based model to investigate the roles of attractive and repellent pheromones in ant decision making during foraging. Journal of Theoretical Biology, 255(2):250–258, November 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004281.

Rempala:2006:SMG

[RRK06] Grzegorz A. Rempala, Kenneth S. Ramos, and Ted Kalbfleisch. A stochastic model of gene transcription: An application to L1 retrotransposition events. Journal of Theoretical Biology, 242(1):101–116, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000610.

Ruokolainen:2009:CSU

[RRKF09] Lasse Ruokolainen, Esa Ranta, Veijo Kaitala, and Mike S. Fowler. Community stability under different correlation structures of species’ environmental responses. Journal of Theoretical Biology, 261(3):379–387, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003750.

Rowe:2004:ENP

[RRR04] Donald. L. Rowe, Peter. A. Robinson, and Christopher. J. Rennie. Estimation of neurophysiological parameters from the waking EEG using a biophysical model of brain dynamics. Journal of Theoretical Biology, 231(3):413–433, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930400325X.

Rossi:2008:DPF

[RRR^{+}08] Federico Rossi, Sandra Ristori, Mauro Rustici, Nadia Marchettini, and Enzo Tiezzi. Dynamics of pattern formation in biomimetic systems. Journal of Theoretical Biology, 255(4):404–412, December 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004487.

Romanyukha:2006:ECI

[RRS06] Alexei A. Romanyukha, Sergey G. Rudnev, and Igor A. Sidorov. Energy cost of infection burden: An approach to understanding the dynamics of host-pathogen interactions. Journal of Theoretical Biology, 241(1):1–13, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004790.

Reynolds:2007:AFP

[RRS^{+}07] A. M. Reynolds, D. R. Reynolds, A. D. Smith, G. P. Svensson, and C. Löfstedt. Appetitive flight patterns of male agrotis segetum moths over landscape scales. Journal of Theoretical Biology, 245(1):141–149, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004759.

Roberts:2004:AII

[RS04] Alan Roberts and Lewi Stone. Advantageous indirect interactions in systems of competition. Journal of Theoretical Biology, 228(3):367–375, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304000487.

Rode:2009:TIF

[RSB09] Christian Rode, Tobias Siebert, and Reinhard Blickhan. Titininduced force enhancement and force depression: a ‘stickyspring’ mechanism in muscle contractions? Journal of Theoretical Biology, 259(2):350–360, July 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001143.

Rauch:2003:DGS

[RSBY03] Erik M. Rauch, Hiroki Sayama, and Yaneer Bar-Yam. Dynamics and genealogy of strains in spatially extended host– pathogen models. Journal of Theoretical Biology, 221(4):655– 664, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931274.

Ribba:2006:MMM

[RSC^{+}06] B. Ribba, O. Saut, T. Colin, D. Bresch, E. Grenier, and J. P. Boissel. A multiscale mathematical model of avascular tumor growth to investigate the therapeutic benefit of anti-invasive agents. Journal of Theoretical Biology, 243(4):532–541, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003043.

Radmacher:2001:GMO

[RSD^{+}01] Michael D. Radmacher, Richard Simon, Richard Desper, Raymond Taetle, Alejandro A. Schäffer, and Mark A. Nelson. Graph models of oncogenesis with an application to melanoma. Journal of Theoretical Biology, 212(4):535–548, October 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923951.

Rudiger:2009:DEE

[RSG09] S. Rüdiger and L. Schimansky-Geier. Dynamics of excitable elements with time-delayed coupling. Journal of Theoretical Biology, 259(1):96–100, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000484.

Rocha:2006:CTD

[RSH^{+}06] Eduardo P. C. Rocha, John Maynard Smith, Laurence D. Hurst, Matthew T. G. Holden, Jessica E. Cooper, Noel H. Smith, and Edward J. Feil. Comparisons of dN/dS are time dependent for closely related bacterial genomes. Journal of Theoretical Biology, 239(2):226–235, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500384X.

Rozdilsky:2004:EIC

[RSS04] Ian D. Rozdilsky, Lewi Stone, and Andrew Solow. The effects of interaction compartments on stability for competitive systems. Journal of Theoretical Biology, 227(2):277–282, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004247.

Rand:2006:UDP

[RSSM06] D. A. Rand, B. V. Shulgin, J. D. Salazar, and A. J. Millar. Uncovering the design principles of circadian clocks: Mathematical analysis of flexibility and evolutionary goals. Journal of Theoretical Biology, 238(3):616–635, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002845.

Ranta:2002:EVS

[RTK02] Esa Ranta, David Tesar, and Veijo Kaitala. Environmental variability and semelparity vs. iteroparity as life histories. Journal of Theoretical Biology, 217(3):391–396, August 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302930298.

Ruan:2001:OPM

[Rua01] Shigui Ruan. Oscillations in plankton models with nutrient recycling. Journal of Theoretical Biology, 208(1):15–26, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921969.

Ruebenbauer:2007:SNR

[Rue07] Agnieszka Ruebenbauer. Simulation of the neural response to the odour stimulus. Journal of Theoretical Biology, 248(2):311–316, September 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002287.

Reluga:2005:SES

[RV05] Timothy C. Reluga and Steven Viscido. Simulated evolution of selfish herd behavior. Journal of Theoretical Biology, 234(2):213–225, May 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005715.

Reijenga:2005:YGO

[RvMK^{+}05] Karin A. Reijenga, Yvonne M. G. A. van Megen, Bob W. Kooi, Barbara M. Bakker, Jacky L. Snoep, Henk W. van Verseveld, and Hans V. Westerhoff. Yeast glycolytic oscillations that are not controlled by a single oscillophore: a new definition of oscillophore strength. Journal of Theoretical Biology, 232(3):385–398, February 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004011.

Ruoff:2001:GMS

[RVMR01] Peter Ruoff, Merete Vinsjevik, Christian Monnerjahn, and Ludger Rensing. The Goodwin model: Simulating the effect of light pulses on the circadian sporulation rhythm of neurospora crassa. Journal of Theoretical Biology, 209(1):29–42, March 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922392.

Ronkko:2008:MCE

[RW08] Mauno Rönkkö and Garry Wong. Modeling the C. elegans nematode and its environment using a particle system. Journal of Theoretical Biology, 253(2):316–322, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001550.

Ray:2008:TTI

[RWCK08] J. Christian J. Ray, Jian Wang, John Chan, and Denise E. Kirschner. The timing of TNF and IFN-γ signaling affects macrophage activation strategies during Mycobacterium tuberculosis infection. Journal of Theoretical Biology, 252(1):24–38, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000222.

Reinhardt:2001:TIT

[RWF01] M. Reinhardt, D. Wenzler, and L. Fukshansky. Topology of an intracellular transduction chain (phototropism of phycomyces ): 2. Mathematical model of pattern transduction. Journal of Theoretical Biology, 211(4):333–345, August 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301923501.

Riskin:2008:QCB

[RWID^{+}08] Daniel K. Riskin, David J. Willis, José Iriarte-Díaz, Tyson L. Hedrick, Mykhaylo Kostandov, Jian Chen, David H. Laidlaw, Kenneth S. Breuer, and Sharon M. Swartz. Quantifying the complexity of bat wing kinematics. Journal of Theoretical Biology, 254(3):604–615, October 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800324X.

Robinson:2008:NRE

[RWK08] P. A. Robinson, H. Wu, and J. W. Kim. Neural rate equations for bursting dynamics derived from conductance-based equations. Journal of Theoretical Biology, 250(4):663–672, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005164.

Riggs:2008:CRV

[RWP^{+}08] Thomas Riggs, Adrienne Walts, Nicolas Perry, Laura Bickle, Jennifer N. Lynch, Amy Myers, Joanne Flynn, Jennifer J. Linderman, Mark J. Miller, and Denise E. Kirschner. A comparison of random vs. chemotaxis-driven contacts of T cells with dendritic cells during repertoire scanning. Journal of Theoretical Biology, 250(4):732–751, February 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005073.

Rennie:2000:MCE

[RWR00] Christopher J. Rennie, James J. Wright, and Peter A. Robinson. Mechanisms of cortical electrical activity and emergence of gamma rhythm. Journal of Theoretical Biology, 205(1):17– 35, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930092040X.

Rossberg:2006:ETL

[RYAI06] A. G. Rossberg, K. Yanagi, T. Amemiya, and K. Itoh. Estimating trophic link density from quantitative but incomplete diet data. Journal of Theoretical Biology, 243(2):261– 272, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519306002566.

Rabinovitch:2003:DMI

[RZF03] Avinoam Rabinovitch, Arieh Zaritsky, and Mario Feingold. DNA-membrane interactions can localize bacterial cell center. Journal of Theoretical Biology, 225(4):493–496, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002923.

Subramanian:2001:PHB

[SA01] Balakrishna Subramanian and David E. Axelrod. Progression of heterogeneous breast tumors. Journal of Theoretical Biology, 210(1):107–119, May 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923021.

Seto:2005:DID

[SA05a] Mayumi Seto and Tasuku Akagi. Daisyworld inhabited with daisies incorporating a seed size/number trade-off: the mechanism of negative feedback on selection from a standpoint of the competition theory. Journal of Theoretical Biology, 234(2):167–172, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304005673.

Sontag:2005:EPP

[SA05b] Laura Sontag and David E. Axelrod. Evaluation of pathways for progression of heterogeneous breast tumors. Journal of Theoretical Biology, 232(2):179–189, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400373X.

Seto:2007:CIR

[SA07a] Mayumi Seto and Tasuku Akagi. Coexistence introducing regulation of environmental conditions. Journal of Theoretical Biology, 248(2):267–274, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002263.

Suzuki:2007:EIR

[SA07b] Shinsuke Suzuki and Eizo Akiyama. Evolution of indirect reciprocity in groups of various sizes and comparison with direct reciprocity. Journal of Theoretical Biology, 245(3):539–552, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005303.

Suzuki:2007:TPG

[SA07c] Shinsuke Suzuki and Eizo Akiyama. Three-person game facilitates indirect reciprocity under image scoring. Journal of Theoretical Biology, 249(1):93–100, November 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003426.

Sornborger:2008:EFS

[SA08a] Andrew T. Sornborger and Malcolm R. Adams. The evolution of fidelity in sensory systems. Journal of Theoretical Biology, 253(1):142–150, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800115X.

Suzuki:2008:COE

[SA08b] Shinsuke Suzuki and Eizo Akiyama. Chaos, oscillation and the evolution of indirect reciprocity in n-person games. Journal of Theoretical Biology, 252(4):686–693, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000581.

Sachs:2004:SOT

[Sac04a] Tsvi Sachs. Self-organization of tree form: a model for complex social systems. Journal of Theoretical Biology, 230(2):197–202, September 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002048.

Shurin:2004:ASS

[SAC^{+}04b] Jonathan B. Shurin, Priyanga Amarasekare, Jonathan M. Chase, Robert D. Holt, Martha F. Hoopes, and Mathew A. Leibold. Alternative stable states and regional community structure. Journal of Theoretical Biology, 227(3):359–368, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004326.

Sachs:2005:AYM

[Sac05] Gottfried Sachs. Aerodynamic yawing moment characteristics of bird wings. Journal of Theoretical Biology, 234(4):471–478, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304006034.

Sachs:2007:TEY

[Sac07] Gottfried Sachs. Tail effects on yaw stability in birds. Journal of Theoretical Biology, 249(3):464–472, December 7, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003402.

Saftenku:2005:MSG

[Saf05] E. È. Saftenku. Modeling of slow glutamate diffusion and AMPA receptor activation in the cerebellar glomerulus. Journal of Theoretical Biology, 234(3):363–382, June 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005867.

Saftenku:2009:CSN

[Saf09] E. È. Saftenku. Computational study of non-homogeneous distribution of Ca^{2+} handling systems in cerebellar granule cells. Journal of Theoretical Biology, 257(2):228–244, March 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006280.

Santillan:2003:ASL

[San03] Moisés Santillán. Allometric scaling law in a simple oxygen exchanging network: possible implications on the biological allometric scaling laws. Journal of Theoretical Biology, 223(2):249–257, July 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000973.

Saroff:2004:PEB

[Sar04] Harry A. Saroff. pH effects on the binding of oxygen to non-vertebrate monomeric hemoglobins. A linked function model. Journal of Theoretical Biology, 229(1):113–118, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001080.

Sato:2004:SDS

[Sat04] Toshihiko Sato. Size-dependent sex allocation in hermaphroditic plants: the effects of resource pool and self-incompatibility. Journal of Theoretical Biology, 227(2):265–275, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004235.

Savage:2004:IAS

[Sav04] Van M. Savage. Improved approximations to scaling relationships for species, populations, and ecosystems across latitudinal and elevational gradients. Journal of Theoretical Biology, 227(4):525–534, April 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004466.

Sumpter:2000:SDT

[SB00] D. J. T. Sumpter and D. S. Broomhead. Shape and dynamics of thermoregulating honey bee clusters. Journal of Theoretical Biology, 204(1):1–14, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910639.

Scherer:2005:EMM

[SB05] Almut Scherer and Sebastian Bonhoeffer. Epitope downmodulation as a mechanism for the coexistence of competing T-cells. Journal of Theoretical Biology, 233(3):379–390, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004965.

Sen:2006:MIT

[SB06] Pali Sen and Denis Bell. A model for the interaction of two chemicals. Journal of Theoretical Biology, 238(3):652–656, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002717.

Skedros:2007:MAT

[SB07] John G. Skedros and Sidney L. Baucom. Mathematical analysis of trabecular ‘trajectories’ in apparent trajectorial structures: The unfortunate historical emphasis on the human proximal femur. Journal of Theoretical Biology, 244(1):15–45, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002773.

Schenker:2009:SIE

[SB09a] Patricia Schenker and Antonio Baici. Simultaneous interaction of enzymes with two modifiers: Reappraisal of kinetic models and new paradigms. Journal of Theoretical Biology, 261(2):318–329, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003440.

Schmitt:2009:DSL

[SB09b] J. Schmitt and S. Bonnono. Dynamics and stability of lateral plane locomotion on inclines. Journal of Theoretical Biology, 261(4):598–609, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003907.

Shuryak:2009:MIB

[SB09c] Igor Shuryak and David J. Brenner. A model of interactions between radiation-induced oxidative stress, protein and DNA damage in deinococcus radiodurans. Journal of Theoretical Biology, 261(2):305–317, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003671.

Spencer:2004:ODE

[SBGA04] Sabrina L. Spencer, Matthew J. Berryman, José A. García, and Derek Abbott. An ordinary differential equation model for the multistep transformation to cancer. Journal of Theoretical Biology, 231(4):515–524, December 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003273.

Schreuder:2001:MIN

[SBH01] Maarten D. J. Schreuder, Carol A. Brewer, and Carlton Heine. Modelled influences of non-exchanging trichomes on leaf boundary layers and gas exchange. Journal of Theoretical Biology, 210(1):23–32, May 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922854.

Santangelo:2007:PDC

[SBI07] Giovanni Santangelo, Lorenzo Bramanti, and Mimmo Iannelli. Population dynamics and conservation biology of the over-exploited Mediterranean red coral. Journal of Theoretical Biology, 244(3):416–423, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003973.

Stich:2008:SRP

[SBM08] Michael Stich, Carlos Briones, and Susanna C. Manrubia. On the structural repertoire of pools of short, random RNA sequences. Journal of Theoretical Biology, 252(4):750–763, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000842.

Sepulveda:2005:SMC

[SBPC05] Nuno Sepúlveda, Laurent Boucontet, Pablo Pereira, and Jorge Carneiro. Stochastic modeling of T cell receptor γ gene rearrangement. Journal of Theoretical Biology, 234(2):153–165, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005478.

Scholz:2006:SJG

[SBvS06] Melanie N. Scholz, Maarten F. Bobbert, and A. J. Knoek van Soest. Scaling and jumping: Gravity loses grip on small jumpers. Journal of Theoretical Biology, 240(4):554–561, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004674.

Shahaf:2008:ADS

[SBZ^{+}08] Gitit Shahaf, Michał Barak, Neta S. Zuckerman, Naamah Swerdlin, Malka Gorfine, and Ramit Mehr. Antigen-driven selection in germinal centers as reflected by the shape characteristics of immunoglobulin gene lineage trees: a large-scale simulation study. Journal of Theoretical Biology, 255(2):210–222, November 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004116.

Sarkar:2003:OPB

[SC03] R. R. Sarkar and J. Chattopadhayay. Occurrence of planktonic blooms under environmental fluctuations and its possible control mechanism-mathematical models and experimental observations. Journal of Theoretical Biology, 224(4):501–516, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002005.

Sanchez:2005:AMA

[SC05] Angel Sánchez and José A. Cuesta. Altruism may arise from individual selection. Journal of Theoretical Biology, 235(2):233–240, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305000226.

Shen:2009:PPF

[SC09] Hong-Bin Shen and Kuo-Chen Chou. Predicting protein fold pattern with functional domain and sequential evolution information. Journal of Theoretical Biology, 256(3):441–446, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005213.

Stafford:2000:MPV

[SCC^{+}00] Max A. Stafford, Lawrence Corey, Yunzhen Cao, Eric S. Daar, David D. Ho, and Alan S. Perelson. Modeling plasma virus concentration during primary HIV infection. Journal of Theoretical Biology, 203(3):285–301, April 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910762.

Salazar-Ciudad:2000:GNC

[SCGFS00] Isaac Salazar-Ciudad, Jordi Garcia-Fernández, and Ricard V. Solé. Gene networks capable of pattern formation: From induction to reaction–diffusion. Journal of Theoretical Biology, 205(4):587–603, August 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920927.

Scheuring:2000:HDS

[Sch00] István Scheuring. Habitat diversity and stability in a metapopulation. Journal of Theoretical Biology, 206(2):229– 233, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519300921179.

Schlink:2002:PMA

[Sch02a] Uwe Schlink. On the probability model for asthma attacks. Journal of Theoretical Biology, 215(4):405–414, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925269.

Schofield:2002:SEM

[Sch02b] Peter Schofield. Spatially explicit models of Turelli–Hoffmann Wolbachia invasive wave fronts. Journal of Theoretical Biology, 215(1):121–131, March 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924932.

Schofield:2002:MMH

[SCH02c] Peter Schofield, Mark Chaplain, and Stephen Hubbard. Mathematical modelling of host–parasitoid systems: Effects of chemically mediated parasitoid foraging strategies on withinand between-generation spatio-temporal dynamics. Journal of Theoretical Biology, 214(1):31–47, January 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192418X.

Schinazi:2003:RRT

[Sch03] Rinaldo B. Schinazi. On the role of reinfection in the transmission of infectious diseases. Journal of Theoretical Biology, 225(1):59–63, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002200.

Scheuring:2005:ICP

[Sch05a] István Scheuring. The iterated continuous prisoner’s dilemma game cannot explain the evolution of interspecific mutualism in unstructured populations. Journal of Theoretical Biology, 232(1):99–104, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003728.

Schofield:2005:ESL

[SCH05b] Peter Schofield, Mark Chaplain, and Stephen Hubbard. Evolution of searching and life history characteristics in individualbased models of host-parasitoid-microbe associations. Journal of Theoretical Biology, 237(1):1–16, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001451.

Schreiber:2006:PDP

[Sch06] Sebastian J. Schreiber. Persistence despite perturbations for interacting populations. Journal of Theoretical Biology, 242(4):844–852, October 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600172X.

Schank:2008:DLK

[Sch08a] Jeffrey C. Schank. The development of locomotor kinematics in neonatal rats: An agent-based modeling analysis in group and individual contexts. Journal of Theoretical Biology, 254(4):826–842, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003743.

Schmid:2008:PUD

[Sch08b] Ralf Schmid. Pitfalls of using out of date databases. Comment on Guo et al. and Guo et al. Journal of Theoretical Biology, 255(2):267, November 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004475.

Scheuring:2009:EGC

[Sch09] István Scheuring. Evolution of generous cooperative norms by cultural group selection. Journal of Theoretical Biology, 257(3):397–407, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519308006607.

Sumner:2008:MIP

[SCJJ08] J. G. Sumner, M. A. Charleston, L. S. Jermiin, and P. D. Jarvis. Markov invariants, plethysms, and phylogenetics. Journal of Theoretical Biology, 253(3):601–615, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001720.

Schley:2006:MER

[SCNP^{+}06] D. Schley, R. Carare-Nnadi, C. P. Please, V. H. Perry, and R. O. Weller. Mechanisms to explain the reverse perivascular transport of solutes out of the brain. Journal of Theoretical Biology, 238(4):962–974, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003085.

Scott:2000:EEH

[Sco00] Christopher B. Scott. Energy expenditure of heavy to severe exercise and recovery. Journal of Theoretical Biology, 207(2):293–297, November 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930092174X.

Simpson:2004:FDC

[SCS04a] Michael L. Simpson, Chris D. Cox, and Gary S. Sayler. Frequency domain chemical Langevin analysis of stochasticity in gene transcriptional regulation. Journal of Theoretical Biology, 229(3):383–394, August 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001481.

Singh:2004:RVI

[SCS04b] Brajendra K. Singh, Joydev Chattopadhyay, and Somdatta Sinha. The role of virus infection in a simple phytoplankton zooplankton system. Journal of Theoretical Biology, 231(2):153–166, November 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002838.

Szabo:2007:CAB

[SCS07] Péter Szabó, Tamás Czárán, and György Szabó. Competing associations in bacterial warfare with two toxins. Journal of Theoretical Biology, 248(4):736–744, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003153.

Saffre:2002:SSC

[SD02] F. Saffre and J. L. Deneubourg. Swarming strategies for cooperative species. Journal of Theoretical Biology, 214(3):441–451, February 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924725.

Sole:2004:ECC

[SD04] Ricard V. Solé and Thomas S. Deisboeck. An error catastrophe in cancer? Journal of Theoretical Biology, 228(1):47– 54, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004569.

Shastri:2006:SEMa

[SD06a] Y. Shastri and U. Diwekar. Sustainable ecosystem management using optimal control theory: Part 1 (deterministic systems). Journal of Theoretical Biology, 241(3):506–521, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500545X.

Shastri:2006:SEMb

[SD06b] Y. Shastri and U. Diwekar. Sustainable ecosystem management using optimal control theory: Part 2 (stochastic systems). Journal of Theoretical Biology, 241(3):522–532, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005461.

Selivanov:2008:CEM

[SdAC^{+}08] V. A. Selivanov, P. de Atauri, J. J. Centelles, J. Cadefau, J. Parra, R. Cussó, J. Carreras, and M. Cascante. The changes in the energy metabolism of human muscle induced by training. Journal of Theoretical Biology, 252(3):402–410, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004717.

Spencer:2004:PAM

[SDBH04] Matthew Spencer, Elizabeth A. Davidson, Adrian C. Barbrook, and Christopher J. Howe. Phylogenetics of artificial manuscripts. Journal of Theoretical Biology, 227(4):503–511, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004442.

Shapiro:2008:VLE

[SDL^{+}08] M. Shapiro, K. A. Duca, K. Lee, E. Delgado-Eckert, J. Hawkins, A. S. Jarrah, R. Laubenbacher, N. F. Polys, V. Hadinoto, and D. A. Thorley-Lawson. A virtual look at Epstein-Barr virus infection: Simulation mechanism. Journal of Theoretical Biology, 252(4):633–648, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000465.

Schley:2004:PMS

[SDS04] David Schley, C. Patrick Doncaster, and Tim Sluckin. Population models of sperm-dependent parthenogenesis. Journal of Theoretical Biology, 229(4):559–572, August 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001791.

Sambelashvili:2002:PER

[SE02a] Alexandre Sambelashvili and Igor R. Efimov. The pinwheel experiment re-revisited. Journal of Theoretical Biology, 214(2):147–153, January 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924567.

Shertzer:2002:ESE

[SE02b] Kyle W. Shertzer and Stephen P. Ellner. Energy storage and the evolution of population dynamics. Journal of Theoretical Biology, 215(2):183–200, March 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930192507X.

Seno:2007:MMO

[SE07] Hiromi Seno and Hiromi Endo. A mathematical model on the optimal timing of offspring desertion. Journal of Theoretical Biology, 246(3):555–563, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000276.

Seebacher:2000:HTM

[See00] F. Seebacher. Heat transfer in a microvascular network: the effect of heart rate on heating and cooling in reptiles (Pogona barbata and Varanus varius). Journal of Theoretical Biology, 203(2):97–109, March 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910676.

Seligmann:2006:EPA

[Sel06] Hervé Seligmann. Error propagation across levels of organization: From chemical stability of ribosomal RNA to developmental stability. Journal of Theoretical Biology, 242(1):69–80, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000580.

Seligmann:2007:CMR

[Sel07] Hervé Seligmann. Cost minimization of ribosomal frameshifts. Journal of Theoretical Biology, 249(1):162–167, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307003347.

Sardanyes:2008:SQM

[SES08] Josep Sardanyés, Santiago F. Elena, and Ricard V. Solé. Simple quasispecies models for the survival-of-the-flattest effect: The role of space. Journal of Theoretical Biology, 250(3):560–568, February 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700522X.

Sevush:2006:SNT

[Sev06] Steven Sevush. Single-neuron theory of consciousness. Journal of Theoretical Biology, 238(3):704–725, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002766.

Sarkar:2004:LAP

[SF04] A. Sarkar and B. R. Franza. A logical analysis of the process of T cell activation: different consequences depending on the state of CD28 engagement. Journal of Theoretical Biology, 226(4):455–466, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003692.

Silvestre:2008:ICH

[SF08a] Daniel A. M. M. Silvestre and José F. Fontanari. The information capacity of hypercycles. Journal of Theoretical Biology, 254(4):804–806, October 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003779.

Silvestre:2008:PMI

[SF08b] Daniel A. M. M. Silvestre and José F. Fontanari. Package models and the information crisis of prebiotic evolution. Journal of Theoretical Biology, 252(2):326–337, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308000635.

Sun:2009:VFD

[SFC^{+}09] Sean X. Sun, Brenda Farrell, Matthew S. Chana, George Oster, William E. Brownell, and Alexander A. Spector. Voltage and frequency dependence of prestin-associated charge transfer. Journal of Theoretical Biology, 260(1):137–144, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002318.

Swihart:2001:EHD

[SFS^{+}01] Robert K. Swihart, Zhilan Feng, Norman A. Slade, Doran M. Mason, and Thomas M. Gehring. Effects of habitat destruction and resource supplementation in a predator–prey metapopulation model. Journal of Theoretical Biology, 210(3):287–303, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923045.

Sendova-Franks:2002:RWM

[SFV02] Ana B. Sendova-Franks and Jan Van Lent. Random walk models of worker sorting in ant colonies. Journal of Theoretical Biology, 217(2):255–274, July 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930110.

Sicardi:2009:RMS

[SFVA09] Estrella A. Sicardi, Hugo Fort, Mendeli H. Vainstein, and Jeferson J. Arenzon. Random mobility and spatial structure often enhance cooperation. Journal of Theoretical Biology, 256(2):240–246, January 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308004943.

Sriram:2004:TVD

[SG04] K. Sriram and M. S. Gopinathan. A two variable delay model for the circadian rhythm of neurospora crassa. Journal of Theoretical Biology, 231(1):23–38, November 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001419.

Srividhya:2006:STD

[SG06] J. Srividhya and M. S. Gopinathan. A simple time delay model for eukaryotic cell cycle. Journal of Theoretical Biology, 241(3):617–627, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500559X.

Schaefer:2007:MEL

[SG07] H. Martin Schaefer and Kevin S. Gould. Modelling the evolution of leaf colouration with binary assumptions is barking up the wrong tree. Journal of Theoretical Biology, 249(3):638–639, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003839.

Sabouri-Ghomi:2008:ABP

[SGCK^{+}08] Mohsen Sabouri-Ghomi, Andrea Ciliberto, Sandip Kar, Bela Novak, and John J. Tyson. Antagonism and bistability in protein interaction networks. Journal of Theoretical Biology, 250(1):209–218, January 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004171.

Sidorov:2004:KMT

[SGD04] Igor A. Sidorov, Dennis Gee, and Dimiter S. Dimitrov. A kinetic model of telomere shortening in infants and adults. Journal of Theoretical Biology, 226(2):169–175, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930300328X.

Smallbone:2007:MCD

[SGG^{+}07] Kieran Smallbone, Robert A. Gatenby, Robert J. Gillies, Philip K. Maini, and David J. Gavaghan. Metabolic changes during carcinogenesis: Potential impact on invasiveness. Journal of Theoretical Biology, 244(4):703–713, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004115.

Smallbone:2005:RAS

[SGGM05] Kieran Smallbone, David J. Gavaghan, Robert A. Gatenby, and Philip K. Maini. The role of acidity in solid tumour growth and invasion. Journal of Theoretical Biology, 235(4):476–484, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000548.

Smallbone:2008:MMT

[SGM08] Kieran Smallbone, Robert A. Gatenby, and Philip K. Maini. Mathematical modelling of tumour acidity. Journal of Theoretical Biology, 255(1):106–112, November 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004128.

Scapoli:2005:SDF

[SGS^{+}05] C. Scapoli, H. Goebl, S. Sobota, E. Mamolini, A. Rodriguez-Larralde, and I. Barrai. Surnames and dialects in France: Population structure and cultural evolution. Journal of Theoretical Biology, 237(1):75–86, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001566.

Saeb:2007:MPA

[SGTF07] Sohrab Saeb, Shahriar Gharibzadeh, Farzad Towhidkhah, and Aydin Farajidavar. Modeling the primary auditory cortex using dynamic synapses: Can synaptic plasticity explain the temporal tuning? Journal of Theoretical Biology, 248(1):1–9, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001415.

Sanchez:2003:UVT

[SH03] María S. Sánchez and Alan Hastings. Uniform vertical transmission and selection in a host-symbiont system. Non-random symbiont distribution generates apparent differential selection. Journal of Theoretical Biology, 225(4):517–530, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003035.

Sakai:2004:SNT

[SH04] Satoki Sakai and Yasushi Harada. Size-number trade-off and optimal offspring size for offspring produced sequentially using a fixed amount of reserves. Journal of Theoretical Biology, 226(3):253–264, February 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002819.

Sakai:2005:POU

[SH05] Satoki Sakai and Yasushi Harada. Production of offspring using current income and reserves: size-number trade-off and optimal offspring size. Journal of Theoretical Biology, 233(1):65– 73, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004527.

Steel:2006:RPC

[SH06] Mike Steel and Jotun Hein. Reconstructing pedigrees: a combinatorial perspective. Journal of Theoretical Biology, 240(3):360–367, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519305004224.

Stone:2007:COC

[SH07] Lewi Stone and Daihai He. Chaotic oscillations and cycles in multi-trophic ecological systems. Journal of Theoretical Biology, 248(2):382–390, September 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002652.

Shackelton:2008:RAG

[SH08a] Laura A. Shackelton and Edward C. Holmes. The role of alternative genetic codes in viral evolution and emergence. Journal of Theoretical Biology, 254(1):128–134, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002750.

Srinivasan:2008:HWC

[SH08b] Manoj Srinivasan and Philip Holmes. How well can springmass-like telescoping leg models fit multi-pedal sagittal-plane locomotion data? Journal of Theoretical Biology, 255(1):1–7, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800341X.

Sherratt:2006:SMF

[She06] Thomas N. Sherratt. Spatial mosaic formation through frequency-dependent selection in Müllerian mimicry complexes. Journal of Theoretical Biology, 240(2):165–174, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004200.

Silver:2002:MID

[SHF02] Frederick H. Silver, Istvan Horvath, and David J. Foran. Mechanical implications of the domain structure of fiberforming collagens: Comparison of the molecular and fibrillar flexibilities of the α1-chains found in Types I–III collagen. Journal of Theoretical Biology, 216(2):243–254, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519302925427.

Sidorov:2002:CCD

[SHHD02] Igor A. Sidorov, Ken S. Hirsch, Calvin B. Harley, and Dimiter S. Dimitrov. Cancer cell dynamics in presence of telomerase inhibitors: Analysis of in vitro data. Journal of Theoretical Biology, 219(2):225–233, November 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931176.

Shudo:2003:OCB

[SHI03] Emi Shudo, Patsy Haccou, and Yoh Iwasa. Optimal choice between feedforward and feedback control in gene expression to cope with unpredictable danger. Journal of Theoretical Biology, 223(2):149–160, July 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300081X.

Shiraishi:2005:EMC

[SHI05] Fumihide Shiraishi, Yuji Hatoh, and Toshinori Irie. An efficient method for calculation of dynamic logarithmic gains in biochemical systems theory. Journal of Theoretical Biology, 234(1):79–85, May 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519304005417. See corrigendum [SHI06b].

Shinbrot:2006:SMD

[Shi06a] Troy Shinbrot. Simulated morphogenesis of developmental folds due to proliferative pressure. Journal of Theoretical Biology, 242(3):764–773, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001718.

Shiraishi:2006:CEM

[SHI06b] Fumihide Shiraishi, Yuji Hatoh, and Toshinori Irie. Corrigendum to “An efficient method for calculation of dynamic logarithmic gains in biochemical systems theory”: [J. Theor. Biol. 234 (2005) 79–85]. Journal of Theoretical Biology, 239(3):398–399, April 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003620. See [SHI05].

Selgrade:2009:MHC

[SHP09] J. F. Selgrade, L. A. Harris, and R. D. Pasteur. A model for hormonal control of the menstrual cycle: Structural consistency but sensitivity with regard to data. Journal of Theoretical Biology, 260(4):572–580, October 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002835.

Santillan:2003:NSN

[SHPDL03] Moisés Santillán, Ricardo Hernández-Pérez, and Rodolfo Delgado-Lezama. A numeric study of the noise-induced tremor in a mathematical model of the stretch reflex. Journal of Theoretical Biology, 222(1):99–115, May 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300016X.

Sherer:2006:ARC

[SHRR06] E. Sherer, R. E. Hannemann, A. Rundell, and D. Ramkrishna. Analysis of resonance chemotherapy in leukemia treatment via multi-staged population balance models. Journal of Theoretical Biology, 240(4):648–661, June 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005035.

Satake:2000:PCF

[SI00] Akiko Satake and Yoh Iwasa. Pollen coupling of forest trees: Forming synchronized and periodic reproduction out of chaos. Journal of Theoretical Biology, 203(2):63–84, March 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519399910664.

Shudo:2001:IDA

[SI01] Emi Shudo and Yoh Iwasa. Inducible defense against pathogens and parasites: Optimal choice among multiple options. Journal of Theoretical Biology, 209(2):233–247, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922598. See erra-

tum [SI04c].

Shudo:2002:ODS

[SI02] Emi Shudo and Yoh Iwasa. Optimal defense strategy: Storage vs. new production. Journal of Theoretical Biology, 219(3):309–323, December 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302931267.

Schlicht:2004:FGD

[SI04a] Robert Schlicht and Yoh Iwasa. Forest gap dynamics and the Ising model. Journal of Theoretical Biology, 230(1):65–75, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001857.

Shudo:2004:DOH

[SI04b] Emi Shudo and Yoh Iwasa. Dynamic optimization of host defense, immune memory, and post-infection pathogen levels in mammals. Journal of Theoretical Biology, 228(1):17–29, May 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303004545.

Shudo:2004:EID

[SI04c] Emi Shudo and Yoh Iwasa. Erratum to “Inducible defense against pathogens and parasites: optimal choice among multiple options” [J. Theor. Biol. (2001) 209, 233–247]. Journal of Theoretical Biology, 231(1):1, November 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002760. See [SI01].

Shoji:2005:LVS

[SI05] Hiroto Shoji and Yoh Iwasa. Labyrinthine versus straightstriped patterns generated by two-dimensional Turing systems. Journal of Theoretical Biology, 237(1):104–116, November 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500158X.

Schlicht:2006:DRW

[SI06] Robert Schlicht and Yoh Iwasa. Direction of regeneration waves in grid-based models for forest dynamics. Journal of Theoretical Biology, 242(2):363–371, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600110X.

Saeki:2009:AHR

[SI09] Koichi Saeki and Yoh Iwasa. Advantage of having regulatory T cells requires localized suppression of immune reactions. Journal of Theoretical Biology, 260(3):392–401, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002975.

Satake:2004:ELI

[SIHH04] Akiko Satake, Yoh Iwasa, Hiroshi Hakoyama, and Stephen P. Hubbell. Estimating local interaction from spatiotemporal forest data, and Monte Carlo bias correction. Journal of Theoretical Biology, 226(2):225–235, January 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003357.

Shoji:2003:SSR

[SIK03] Hiroto Shoji, Yoh Iwasa, and Shigeru Kondo. Stripes, spots, or reversed spots in two-dimensional Turing systems. Journal of Theoretical Biology, 224(3):339–350, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300170X.

Silverman:2002:TCN

[Sil02] B. David Silverman. A two-component nucleation model of protein hydrophobicity. Journal of Theoretical Biology, 216(2):139–146, May 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519302925531.

Silberstein:2009:CIV

[Sil09] Morry Silberstein. The cutaneous intrinsic visceral afferent nervous system: a new model for acupuncture analgesia. Journal of Theoretical Biology, 261(4):637–642, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004287.

Simon:2008:SME

[Sim08] Burton Simon. A stochastic model of evolutionary dynamics with deterministic large-population asymptotics. Journal of Theoretical Biology, 254(4):719–730, October 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003287.

Shoji:2002:DSF

[SIMK02] Hiroto Shoji, Yoh Iwasa, Atushi Mochizuki, and Shigeru Kondo. Directionality of stripes formed by anisotropic reaction–diffusion models. Journal of Theoretical Biology, 214(4):549–561, February 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924804.

Sinkala:2006:SET

[Sin06] Zachariah Sinkala. Soliton/exciton transport in proteins. Journal of Theoretical Biology, 241(4):919–927, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000403.

Sinkkonen:2007:MEA

[Sin07] A. Sinkkonen. Modelling the effect of autotoxicity on density-dependent phytotoxicity. Journal of Theoretical Biology, 244(2):218–227, January 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003456.

Stollenwerk:2003:MPN

[SJ03] Nico Stollenwerk and Vincent A. A. Jansen. Meningitis, pathogenicity near criticality: the epidemiology of meningococcal disease as a model for accidental pathogens. Journal of Theoretical Biology, 222(3):347–359, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000419.

Sumner:2009:MII

[SJ09] J. G. Sumner and P. D. Jarvis. Markov invariants and the isotropy subgroup of a quartet tree. Journal of Theoretical Biology, 258(2):302–310, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000356.

Shipley:2009:DCP

[SJD^{+}09] R. J. Shipley, G. W. Jones, R. J. Dyson, B. G. Sengers, and … J. Malda. Design criteria for a printed tissue engineering construct: a mathematical homogenization approach. Journal of Theoretical Biology, 259(3):489–502, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930900143X.

Segovia-Juarez:2004:ICM

[SJGK04] Jose L. Segovia-Juarez, Suman Ganguli, and Denise Kirschner. Identifying control mechanisms of granuloma formation during M. tuberculosis infection using an agent-based model. Journal of Theoretical Biology, 231(3):357–376, December 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003212.

Sozou:2001:SMC

[SK01] Peter D. Sozou and Thomas B. L. Kirkwood. A stochastic model of cell replicative senescence based on telomere shortening, oxidative stress, and somatic mutations in nuclear and mitochondrial DNA. Journal of Theoretical Biology, 213(4):573–586, December 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924324.

Saramaki:2005:MDE

[SK05] Jari Saramäki and Kimmo Kaski. Modelling development of epidemics with dynamic small-world networks. Journal of Theoretical Biology, 234(3):413–421, June 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006010.

Sakai:2009:OSA

[SK09a] Satoki Sakai and Tomomi Kojima. Overproduction and selective abortion of ovules based on the order of fertilization revisited. Journal of Theoretical Biology, 260(3):430–437, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002744.

Svennungsen:2009:EBV

[SK09b] Thomas O. Svennungsen and Éva Kisdi. Evolutionary branching of virulence in a single-infection model. Journal of Theoretical Biology, 257(3):408–418, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006097.

Smith:2006:EMT

[SKI^{+}06] Noel H. Smith, Kristin Kremer, Jacqueline Inwald, James Dale, Jeffrey R. Driscoll, Stephen V. Gordon, Dick van Soolingen, R. Glyn Hewinson, and John Maynard Smith. Ecotypes of the mycobacterium tuberculosis complex. Journal of Theoretical Biology, 239(2):220–225, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003838.

StHilaire:2007:ANP

[SKK^{+}07] Melissa A. St. Hilaire, Elizabeth B. Klerman, Sat Bir S. Khalsa, Kenneth P. Wright, Charles A. Czeisler, and Richard E. Kronauer. Addition of a non-photic component to a light-based mathematical model of the human circadian pacemaker. Journal of Theoretical Biology, 247(4):583–599, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001671.

Schreiner:2003:HPC

[SKN^{+}03] Wolfgang Schreiner, Rudolf Karch, Martin Neumann, Friederike Neumann, Susanne M. Roedler, and Georg Heinze. Heterogeneous perfusion is a consequence of uniform shear stress in optimized arterial tree models. Journal of Theoretical Biology, 220(3):285–301, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931365.

Seligmann:2006:MTS

[SKR06a] Hervé Seligmann, Neeraja M. Krishnan, and Basuthkar J. Rao. Mitochondrial tRNA sequences as unusual replication origins: Pathogenic implications for homo sapiens. Journal of Theoretical Biology, 243(3):375–385, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002761.

Seligmann:2006:PMO

[SKR06b] Hervé Seligmann, Neeraja M. Krishnan, and Basuthkar J. Rao. Possible multiple origins of replication in primate mitochondria: Alternative role of tRNA sequences. Journal of Theoretical Biology, 241(2):321–332, July 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005205.

Szczerba:2009:CMI

[SKS09] Dominik Szczerba, Haymo Kurz, and Gabor Szekely. A computational model of intussusceptive microvascular growth and remodeling. Journal of Theoretical Biology, 261(4):570–583, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900438X.

Saunders:2000:IRC

[SKW00] Peter T. Saunders, Johan H. Koeslag, and Jabus A. Wessels. Integral rein control in physiology II: a general model. Journal of Theoretical Biology, 206(2):211–220, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921180.

Shiino:2009:CFD

[SKY09] Yuta Shiino, Osamu Kuwazuru, and Nobuhiro Yoshikawa. Computational fluid dynamics simulations on a Devonian spiriferid Paraspirifer bownockeri (Brachiopoda): Generating mechanism of passive feeding flows. Journal of Theoretical Biology, 259(1):132–141, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000812.

Sella:2000:DPC

[SL00] Guy Sella and Michael Lachmann. On the dynamic persistence of cooperation: How lower individual fitness induces higher survivability. Journal of Theoretical Biology, 206(4):465–485, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921301.

Simpson:2007:CIC

[SLB07] Matthew J. Simpson, Kerry A. Landman, and Kaushik Bhaganagarapu. Coalescence of interacting cell populations. Journal of Theoretical Biology, 247(3):525–543, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000926.

Schulte:2005:NFF

[SLH05] Rogier P. O. Schulte, Egbert A. Lantinga, and Michael J. Hawkins. A new family of Fisher-curves estimates Fisher’s alpha more accurately. Journal of Theoretical Biology, 232 (3):305–313, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930400387X.

Slater:2009:CLT

[SLH^{+}09] Jennifer L. Slater, Kerry A. Landman, Barry D. Hughes, Qin Shen, and Sally Temple. Cell lineage tree models of neurogenesis. Journal of Theoretical Biology, 256(2):164–179, January 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005018.

Simpson:2006:LII

[SLHN06] Matthew J. Simpson, Kerry A. Landman, Barry D. Hughes, and Donald F. Newgreen. Looking inside an invasion wave of cells using continuum models: Proliferation is the key. Journal of Theoretical Biology, 243(3):343–360, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002554.

Satake:2007:CES

[SLIL07] Akiko Satake, Heather M. Leslie, Yoh Iwasa, and Simon A. Levin. Coupled ecological-social dynamics in a forested landscape: Spatial interactions and information flow. Journal of Theoretical Biology, 246(4):695–707, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930700029X.

Sontag:2006:DSI

[SLL06] Laura B. Sontag, Matthew C. Lorincz, and E. Georg Luebeck. Dynamics, stability and inheritance of somatic DNA methylation imprints. Journal of Theoretical Biology, 242(4):890–899, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001895.

Schilling:2000:TSD

[SLP00] Christophe H. Schilling, David Letscher, and Bernhard Ø. Palsson. Theory for the systemic definition of metabolic pathways and their use in interpreting metabolic function from a pathway-oriented perspective. Journal of Theoretical Biology, 203(3):229–248, April 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300910737.

Sanyal:2001:TTS

[SM01] D. C. Sanyal and N. K. Maji. Thermoregulation through skin under variable atmospheric and physiological conditions. Journal of Theoretical Biology, 208(4):451–456, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922318.

Sielewiesiuk:2002:CER

[SM02] Jan Sielewiesiuk and Elzbieta Malarczyk. A cycle of enzymatic reactions that behaves like a neuronal circuit. Journal of Theoretical Biology, 214(2):255–262, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924518.

Simons:2003:GOS

[SM03] Gordon Simons and Glenn Morton. The gene-orientation structure of eukaryotes. Journal of Theoretical Biology, 222 (4):471–475, June 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303000614.

Sachs:2006:ESW

[SM06a] Gottfried Sachs and Mochammad Agoes Moelyadi. Effect of slotted wing tips on yawing moment characteristics. Journal of Theoretical Biology, 239(1):93–100, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003188.

Stamatakis:2006:MAM

[SM06b] Michail Stamatakis and Nikos V. Mantzaris. Modeling of ATP-mediated signal transduction and wave propagation in astrocytic cellular networks. Journal of Theoretical Biology, 241(3):649–668, August 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000038.

Szalisznyo:2009:DIS

[SM09a] Krisztina Szalisznyó and László Müller. Dopamine induced switch in the subthreshold dynamics of the striatal cholinergic interneurons: a numerical study. Journal of Theoretical Biology, 256(4):547–560, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004852.

Szilagyi:2009:LSN

[SM09b] András Szilágyi and Géza Meszéna. Limiting similarity and niche theory for structured populations. Journal of Theoretical Biology, 258(1):27–37, May 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006309.

Santillan:2000:RPP

[SMBM00] Moisés Santillán, Joseph M. Mahaffy, Jacques Bélair, and Michael C. Mackey. Regulation of platelet production: The normal response to perturbation and cyclical platelet disease. Journal of Theoretical Biology, 206(4):585–603, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921490.

Stephanou:2008:CMC

[SMCT08] Angélique Stéphanou, Eleni Mylona, Mark Chaplain, and Philippe Tracqui. A computational model of cell migration coupling the growth of focal adhesions with oscillatory cell protrusions. Journal of Theoretical Biology, 253(4):701–716, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002130.

Saffre:2000:ERP

[SMD00] F. Saffre, A.-C. Mailleux, and J. L. Deneubourg. Exploratory recruitment plasticity in a social spider (Anelosimus eximius). Journal of Theoretical Biology, 205(1):37–46, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920423.

Schreiber:2001:HLD

[SMG01] Sebastian J. Schreiber, Nicholas J. Mills, and Andrew P. Gutierrez. Host-limited dynamics of autoparasitoids. Journal of Theoretical Biology, 212(2):141–153, September 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930192363X.

Sismilich:2003:DMM

[SMG^{+}03] M. Sismilich, M. I. Menzies, P. W. Gandar, P. E. Jameson, and J. Clemens. Development of a mathematical method for classifying and comparing tree architecture using parameters from a topological model of a trifurcating botanical tree. Journal of Theoretical Biology, 220(3):371–391, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931778.

Schaller:2007:MAT

[SMH07] Gernot Schaller and Michael Meyer-Hermann. A modelling approach towards epidermal homoeostasis control. Journal of Theoretical Biology, 247(3):554–573, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001397.

Smith:2004:EV

[Smi04] Kenny Smith. The evolution of vocabulary. Journal of Theoretical Biology, 228(1):127–142, May 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303004636.

Smith:2008:TNSa

[Smi08a] Eric Smith. Thermodynamics of natural selection i: Energy flow and the limits on organization. Journal of Theoretical Biology, 252(2):185–197, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000659.

Smith:2008:TNSb

[Smi08b] Eric Smith. Thermodynamics of natural selection II: Chemical Carnot cycles. Journal of Theoretical Biology, 252(2):198–212, May 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000593.

Smith:2008:TNSc

[Smi08c] Eric Smith. Thermodynamics of natural selection III: Landauer’s principle in computation and chemistry. Journal of Theoretical Biology, 252(2):213–220, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000623.

Smith:2008:EAA

[Smi08d] Robert J. Smith. Explicitly accounting for antiretroviral drug uptake in theoretical HIV models predicts longterm failure of protease-only therapy. Journal of Theoretical Biology, 251(2):227–237, March 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006042.

Smith:2009:PMP

[Smi09] Reginald D. Smith. Plant-mycorrhiza percent infection as evidence of coupled metabolism. Journal of Theoretical Biology, 259(1):172–175, July 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000964.

Sewell:2002:APH

[SML02] Christopher Sewell, Jeffrey J. Morgan, and Paul A. Lindahl. Analysis of protein homeostatic regulatory mechanisms in perturbed environments at steady state. Journal of Theoretical Biology, 215(2):151–167, March 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301925366.

Surovstev:2007:WCM

[SML07] Ivan V. Surovstev, Jeffrey J. Morgan, and Paul A. Lindahl. Whole-cell modeling framework in which biochemical dynamics impact aspects of cellular geometry. Journal of Theoretical Biology, 244(1):154–166, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519306003237. See corrigendum [SML08].

Surovtsev:2008:CWC

[SML08] Ivan V. Surovtsev, Jeffrey J. Morgan, and Paul A. Lindahl. Corrigendum to “Whole-cell modeling framework in which biochemical dynamics impact aspects of cellular geometry” [J. Theor. Biol. 244 (2007) 154–166]. Journal of Theoretical Biology, 251(3):552, April 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006339. See [SML07].

Skupsky:2007:BGS

[SMNL07] Ron Skupsky, Colin McCann, Ralph Nossal, and Wolfgang Losert. Bias in the gradient-sensing response of chemotactic cells. Journal of Theoretical Biology, 247(2):242–258, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307000914.

Scianna:2009:ICB

[SMPM09] Marco Scianna, Roeland M. H. Merks, Luigi Preziosi, and Enzo Medico. Individual cell-based models of cell scatter of ARO and MLP-29 cells in response to hepatocyte growth factor. Journal of Theoretical Biology, 260(1):151–160, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002331.

Stirk:2008:SNS

[SMPvdB08] Emily R. Stirk, Carmen Molina-París, and Hugo A. van den Berg. Stochastic niche structure and diversity maintenance in the T cell repertoire. Journal of Theoretical Biology, 255(2):237–249, November 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308003676.

Sasaki:2003:ML

[SN03] Akira Sasaki and Martin A. Nowak. Mutation landscapes. Journal of Theoretical Biology, 224(2):241–247, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001619.

Subramanian:2004:MME

[SN04] K. K. Subramanian and Atul Narang. A mechanistic model for eukaryotic gradient sensing: Spontaneous and induced phosphoinositide polarization. Journal of Theoretical Biology, 231(1):49–67, November 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002565.

Shipman:2005:PPP

[SN05a] P. D. Shipman and A. C. Newell. Polygonal planforms and phyllotaxis on plants. Journal of Theoretical Biology, 236(2):154–197, September 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000950.

Sivozhelezov:2005:HMC

[SN05b] Victor Sivozhelezov and Claudio Nicolini. Homology modeling of cytochrome p450scc and the mutations for optimal amperometric sensor. Journal of Theoretical Biology, 234(4):479–485, June 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304006046.

Sivozhelezov:2006:TFO

[SN06] Victor Sivozhelezov and Claudio Nicolini. Theoretical framework for octopus rhodopsin crystallization. Journal of Theoretical Biology, 240(2):260–269, May 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004169.

Singh:2007:SDH

[SN07] Abhyudai Singh and Roger M. Nisbet. Semi-discrete hostparasitoid models. Journal of Theoretical Biology, 247(4): 733–742, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519307001865.

Sekiguchi:2009:EPE

[SN09] Takuya Sekiguchi and Mayuko Nakamaru. Effect of the presence of empty sites on the evolution of cooperation by costly punishment in spatial games. Journal of Theoretical Biology, 256(2):297–304, January 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005067.

Small:2008:SDV

[SNA^{+}08] Alexander R. Small, Adrian Neagu, Franck Amyot, Dan Sackett, Victor Chernomordik, and Amir Gandjbakhche. Spatial distribution of VEGF isoforms and chemotactic signals in the vicinity of a tumor. Journal of Theoretical Biology, 252(4):593– 607, June 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800060X.

Santhanakrishnan:2009:FWM

[SNCM09] Arvind Santhanakrishnan, Nhi Nguyen, Jennifer G. Cox, and Laura A. Miller. Flow within models of the vertebrate embryonic heart. Journal of Theoretical Biology, 259(3):449–461, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001830.

Snedecor:2003:CTK

[Sne03] Sonya J. Snedecor. Comparison of three kinetic models of HIV-1 infection: Implications for optimization of treatment. Journal of Theoretical Biology, 221(4):519–541, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303932024.

Steck:2003:FEA

[SNT03] R. Steck, P. Niederer, and M. L. Knothe Tate. A finite element analysis for the prediction of load-induced fluid flow and mechanochemical transduction in bone. Journal of Theoretical Biology, 220(2):249–259, January 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931638.

Solow:2001:CTM

[Sol01] Andrew R. Solow. A comment on two models of the remigration interval in marine turtles. Journal of Theoretical Biology, 209(3):371–372, April 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930192274X.

Schilling:2000:AMC

[SP00] Christophe H. Schilling and Bernhard Ø. Palsson. Assessment of the metabolic capabilities of Haemophilus influenzae Rd through a genome-scale pathway analysis. Journal of Theoretical Biology, 203(3):249–283, April 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910889.

Spedding:2001:UCT

[SP01] G. R. Spedding and C. J. Pennycuick. Uncertainty calculations for theoretical flight power curves. Journal of Theoretical Biology, 208(2):127–139, January 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922082.

Sharan:2002:CMO

[SP02] Maithili Sharan and Aleksander S. Popel. A compartmental model for oxygen transport in brain microcirculation in the presence of blood substitutes. Journal of Theoretical Biology, 216(4):479–500, June 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930018.

Saberi:2005:NCC

[SP05a] Kourosh Saberi and Agavni Petrosyan. Neural crosscorrelation and signal decorrelation: insights into coding of auditory space. Journal of Theoretical Biology, 235(1):45–56, July 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304006174.

Stenoien:2005:MEL

[SP05b] Hans K. Stenøien and Bård Pedersen. Mutation and epimutation load in haploid and diploid life forms. Journal of Theoretical Biology, 233(1):119–126, March 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004679.

Seymour:2006:EGE

[SP06] Robert M. Seymour and Andrew Pomiankowski. ESS gene expression of X-linked imprinted genes subject to sexual selection. Journal of Theoretical Biology, 241(1):81–93, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004820.

Saint-Paul:2007:MFH

[SP07a] Gilles Saint-Paul. On market forces and human evolution. Journal of Theoretical Biology, 247(3):397–412, August 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001373.

Szymanowska-Pulka:2007:ACF

[SP07b] Joanna Szymanowska-Pulka. Application of a changing field of growth rates to a description of root apex formation. Journal of Theoretical Biology, 247(4):650–656, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001919.

Sariyar:2006:MCS

[SPAH06] Berna Sariyar, Sinem Perk, Ugur Akman, and Amable Hortaçsu. Monte Carlo sampling and principal component analysis of flux distributions yield topological and modular information on metabolic networks. Journal of Theoretical Biology, 242(2):389–400, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001135.

Soyer:2006:SES

[SPB06] Orkun S. Soyer, Thomas Pfeiffer, and Sebastian Bonhoeffer. Simulating the evolution of signal transduction pathways. Journal of Theoretical Biology, 241(2):223–232, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305005072.

Slot:2002:SRD

[SPC02] L. A. Bruins Slot, P. J. Pauwels, and F. C. Colpaert. Signreversal during persistent activation in μ-opioid signal transduction. Journal of Theoretical Biology, 215(2):169–182, March 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925093.

Srivastava:2008:NSM

[SPE08] Vipin Srivastava, D. J. Parker, and S. F. Edwards. The nervous system might ‘orthogonalize’ to discriminate. Journal of Theoretical Biology, 253(3):514–517, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001707.

Schuster:2008:MMY

[SPF08] Stefan Schuster, Thomas Pfeiffer, and David A. Fell. Is maximization of molar yield in metabolic networks favoured by evolution? Journal of Theoretical Biology, 252(3):497–504, June 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006340.

Schargott:2006:SMB

[SPG06] M. Schargott, V. L. Popov, and S. Gorb. Spring model of biological attachment pads. Journal of Theoretical Biology, 243(1):48–53, November 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001986.

Seiler:2003:SIO

[SPH03] P. Seiler, A. Pant, and J. K. Hedrick. A systems interpretation for observations of bird v-formations. Journal of Theoretical Biology, 221(2):279–287, March 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931912.

Sivozhelezov:2006:MEP

[SPN06] Victor Sivozhelezov, Eugenia Pechkova, and Claudio Nicolini. Mapping electrostatic potential of a protein on its hydrophobic surface: Implications for crystallization of cytochrome p450scc. Journal of Theoretical Biology, 241(1):73–80, July 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004935.

Souza:2009:ECU

[SPS09] Max O. Souza, Jorge M. Pacheco, and Francisco C. Santos. Evolution of cooperation under N-person snowdrift games. Journal of Theoretical Biology, 260(4):581–588, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003166.

Sundaram:2006:ADZ

[SPVN06] Bala Sundaram, Andrew C. Poje, Richard R. Veit, and Herve Nganguia. Acoustical dead zones and the spatial aggregation of whale strandings. Journal of Theoretical Biology, 238(4):764–770, February 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002808.

Sherratt:2002:SCI

[SR02] Thomas N. Sherratt and Gilbert Roberts. The stability of cooperation involving variable investment. Journal of Theoretical Biology, 215(1):47–56, March 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924956.

Sevim:2008:CGR

[SR08] Volkan Sevim and Per Arne Rikvold. Chaotic gene regulatory networks can be robust against mutations and noise. Journal of Theoretical Biology, 253(2):323–332, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001161.

Sumen:2012:CAB

[SRAL12] Cenk Sumen, Timothy E. Reddy, Mark S. Alber, and Peter P. Lee. Corrigendum to: “Agent-based modeling of the context dependency in T cell recognition” [J. Theor. Biol. 236 (2005) 376–391]. Journal of Theoretical Biology, 303(??):152–153, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001427. See [CSR^{+}05].

Sole:2008:CSC

[SRCDS08] Ricard V. Solé, Carlos Rodríguez-Caso, Thomas S. Deisboeck, and Joan Saldaña. Cancer stem cells as the engine of unstable tumor progression. Journal of Theoretical Biology, 253(4):629–637, August 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308001719.

Shoemaker:2003:DSS

[SRG^{+}03] Jason Shoemaker, Gregory T. Reeves, Shakti Gupta, Sergei S. Pilyugin, Thomas Egli, and Atul Narang. The dynamics of single-substrate continuous cultures: the role of transport enzymes. Journal of Theoretical Biology, 222(3):307–322, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000365.

Shorten:2000:CIE

[SRMW00] Paul R. Shorten, A. Bruce Robson, Alan E. Mckinnon, and David J. N. Wall. CRH-induced electrical activity and calcium signalling in pituitary corticotrophs. Journal of Theoretical Biology, 206(3):395–405, October 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921350.

Surovtsev:2000:MMK

[SRN^{+}00] Ivan V. Surovtsev, Ivan A. Razumov, Vyacheslav M. Nekrasov, Alexander N. Shvalov, Juhani T. Soini, Valeri P. Maltsev, Alexander K. Petrov, Valeri B. Loktev, and Andrei V. Chernyshev. Mathematical modeling the kinetics of cell distribution in the process of ligand–receptor binding. Journal of Theoretical Biology, 206(3):407–417, October 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921362.

Sepulchre:2007:MOV

[SRN07] Jacques-A. Sepulchre, Sylvie Reverchon, and William Nasser. Modeling the onset of virulence in a pectinolytic bacterium. Journal of Theoretical Biology, 244(2):239–257, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306003468.

Sen:2008:TDT

[SRR08] Shrabani Sen, Syed Shahed Riaz, and Deb Shankar Ray. Temperature dependence and temperature compensation of kinetics of chemical oscillations; Belousov–Zhabotinskii reaction, glycolysis and circadian rhythms. Journal of Theoretical Biology, 250(1):103–112, January 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004079.

Santoni:2009:CGA

[SRS09] Daniele Santoni and Vincenzo Romano-Spica. Comparative genomic analysis by microbial COGs self-attraction rate. Journal of Theoretical Biology, 258(4):513–520, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000472.

Strain:2002:SDH

[SRWL02] M. C. Strain, D. D. Richman, J. K. Wong, and H. Levine. Spatiotemporal dynamics of HIV propagation. Journal of Theoretical Biology, 218(1):85–96, September 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930559.

Sitaramam:2000:WDC

[SS00] V. Sitaramam and Zuben E. Sauna. What does a common channel for electrolytes and non-electrolytes in the sperm mean? Journal of Theoretical Biology, 206(3):419–428, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921349.

Scheuring:2001:SRP

[SS01] István Scheuring and Eörs Szathmáry. Survival of replicators with parabolic growth tendency and exponential decay. Journal of Theoretical Biology, 212(1):99–105, September 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923604.

Sober:2002:THC

[SS02] Elliott Sober and Michael Steel. Testing the hypothesis of common ancestry. Journal of Theoretical Biology, 218(4):395–408, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930869.

Sakai:2003:SDE

[SS03] Akiko Sakai and Satoki Sakai. Size-dependent ESS sex allocation in wind-pollinated cosexual plants: fecundity vs. stature effects. Journal of Theoretical Biology, 222(3):283–295, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000249.

Stucki:2005:MMR

[SS05a] Jörg W. Stucki and Hans-Uwe Simon. Mathematical modeling of the regulation of caspase-3 activation and degradation. Journal of Theoretical Biology, 234(1):123–131, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005454.

Szczerba:2005:CMF

[SS05b] Dominik Szczerba and Gábor Székely. Computational model of flow-tissue interactions in intussusceptive angiogenesis. Journal of Theoretical Biology, 234(1):87–97, May 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005429.

Sardanyes:2006:BPT

[SS06a] Josep Sardanyés and Ricard V. Solé. Bifurcations and phase transitions in spatially extended two-member hypercycles. Journal of Theoretical Biology, 243(4):468–482, December 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306003055.

Sato:2006:MCO

[SS06b] Yoko Sato and Hiromi Seno. A mathematical consideration for the optimal shell change of hermit crab. Journal of Theoretical Biology, 240(1):14–23, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003607.

Schmidt:2006:ICA

[SS06c] Otto Schmidt and Andreas Schreiber. Integration of cell adhesion reactions — a balance of forces? Journal of Theoretical Biology, 238(3):608–615, February 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002833.

Szathmary:2006:WDL

[SS06d] Eörs Szathmáry and Mauro Santos. He would have deserved to live forever: John Maynard Smith (1920–2004). Journal of Theoretical Biology, 239(2):129, March 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305005229.

Seymour:2009:DCE

[SS09a] Robert M. Seymour and Peter D. Sozou. Duration of courtship effort as a costly signal. Journal of Theoretical Biology, 256(1):1–13, January 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800502X.

Smallbone:2009:FBA

[SS09b] Kieran Smallbone and Evangelos Simeonidis. Flux balance analysis: a geometric perspective. Journal of Theoretical Biology, 258(2):311–315, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000204.

Savage:2000:CDW

[SSA00] Melissa Savage, Bruce Sawhill, and Manor Askenazi. Community dynamics: What happens when we rerun the tape?

Journal of Theoretical Biology, 205(4):515–526, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920551.

Somsen:2002:SOM

[SSB^{+}02] Oscar J. G. Somsen, Marco Siderius, Florian F. Bauer, Jacky L. Snoep, and Hans V. Westerhoff. Selectivity in overlapping MAP kinase cascades. Journal of Theoretical Biology, 218(3):343–354, October 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930821.

Soyer:2006:STN

[SSB06] Orkun S. Soyer, Marcel Salathé, and Sebastian Bonhoeffer. Signal transduction networks: Topology, response and biochemical processes. Journal of Theoretical Biology, 238(2):416–425, January 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002481.

Sachs:2007:SCA

[SSB^{+}07] Rainer K. Sachs, Igor Shuryak, David Brenner, Hatim Fakir, Lynn Hlatky, and Philip Hahnfeldt. Second cancers after fractionated radiotherapy: Stochastic population dynamics effects. Journal of Theoretical Biology, 249(3):518–531, December 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003578.

Strathe:2009:NMM

[SSD09] A. B. Strathe, H. Sørensen, and A. Danfær. A new mathematical model for combining growth and energy intake in animals: The case of the growing pig. Journal of Theoretical Biology, 261(2):165–175, November 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003518.

Sole:2006:ICR

[SSDM06] Ricard V. Solé, Josep Sardanyés, Juana Díez, and Antonio Mas. Information catastrophe in RNA viruses through replication thresholds. Journal of Theoretical Biology, 240(3):353– 359, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004285.

Sriram:2009:DIP

[SSF09] K. Sriram, Sylvain Soliman, and François Fages. Dynamics of the interlocked positive feedback loops explaining the robust epigenetic switching in Candida albicans. Journal of Theoretical Biology, 258(1):71–88, May 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000216.

Sahlin:2009:RTM

[SSJ09] Patrik Sahlin, Bo Söderberg, and Henrik Jönsson. Regulated transport as a mechanism for pattern generation: Capabilities for phyllotaxis and beyond. Journal of Theoretical Biology, 258(1):60–70, May 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000332.

Smidtas:2006:AFY

[SSK06] Serge Smidtas, Vincent Schächter, and François Képès. The adaptive filter of the yeast galactose pathway. Journal of Theoretical Biology, 242(2):372–381, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001111.

Segre:2001:MRC

[SSKL01] Daniel Segré, Barak Shenhav, Ron Kafri, and Doron Lancet. The molecular roots of compositional inheritance. Journal of Theoretical Biology, 213(3):481–491, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924403.

Smith:2008:EDD

[SSL08] Matthew J. Smith, Jonathan A. Sherratt, and Xavier Lambin. The effects of density-dependent dispersal on the spatiotemporal dynamics of cyclic populations. Journal of Theoretical Biology, 254(2):264–274, September 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002920.

Schmitt:2007:PEW

[SSLB07] Armin O. Schmitt, Johannes Schuchhardt, Arne Ludwig, and Gudrun A. Brockmann. Protein evolution within and between species. Journal of Theoretical Biology, 249(2):376–383, November 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003657.

Steel:2009:PIC

[SSM09] Mike Steel, Laszlo Székely, and Elchanan Mossel. Phylogenetic information complexity: Is testing a tree easier than finding it? Journal of Theoretical Biology, 258(1):95–102, May 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000113.

Sherratt:2004:NSU

[SSR04] Thomas N. Sherratt, Michael P. Speed, and Graeme D. Ruxton. Natural selection on unpalatable species imposed by state-dependent foraging behaviour. Journal of Theoretical Biology, 228(2):217–226, May 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000049.

Szamado:2008:EDN

[SSS08] Szabolcs Számadó, Ferenc Szalai, and István Scheuring. The effect of dispersal and neighbourhood in games of cooperation. Journal of Theoretical Biology, 253(2):221–227, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001112.

Stadler:2001:TPF

[SSWF01] Bärbel M. R. Stadler, Peter F. Stadler, Günter P. Wagner, and Walter Fontana. The topology of the possible: Formal spaces underlying patterns of evolutionary change. Journal of Theoretical Biology, 213(2):241–274, November 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924233.

Sanchez:2001:LAD

[ST01] Lucas Sánchez and Denis Thieffry. A logical analysis of the drosophila gap-gene system. Journal of Theoretical Biology, 211(2):115–141, July 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923355.

Sanchez:2003:SFE

[ST03] Lucas Sánchez and Denis Thieffry. Segmenting the fly embryo: a logical analysis of the pair-rule cross-regulatory module. Journal of Theoretical Biology, 224(4):517–537, October 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303002017.

Suzuki:2005:EIG

[ST05] Yukari Suzuki and Yukihiko Toquenaga. Effects of information and group structure on evolution of altruism: analysis of two-score model by covariance and contextual analyses. Journal of Theoretical Biology, 232(2):191–201, January 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003777.

Stanbro:2000:MIPa

[Sta00a] William D. Stanbro. Modeling the interaction of peroxynitrite with low-density lipoproteins. I. Plasma levels of peroxynitrite. Journal of Theoretical Biology, 205(3):457–464, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920794.

Stanbro:2000:MIPb

[Sta00b] William D. Stanbro. Modeling the interaction of peroxynitrite with low-density lipoproteins. II: Reaction/diffusion model of peroxynitrite in low-density lipoprotein particles. Journal of Theoretical Biology, 205(3):465–471, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920800.

Stanbro:2000:MIPc

[Sta00c] William D. Stanbro. Modeling the interaction of peroxynitrite with low-density lipoproteins. III: The role of antioxidants. Journal of Theoretical Biology, 205(3):473–482, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920812.

Staley:2002:DSL

[Sta02] Mark Staley. Darwinian selection leads to Gaia. Journal of Theoretical Biology, 218(1):35–46, September 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930596.

Stayton:2008:CSE

[Sta08] C. Tristan Stayton. Is convergence surprising? An examination of the frequency of convergence in simulated datasets. Journal of Theoretical Biology, 252(1):1–14, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000209.

Stadler:2009:ISU

[Sta09] Tanja Stadler. On incomplete sampling under birth-death models and connections to the sampling-based coalescent. Journal of Theoretical Biology, 261(1):58–66, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003300.

Stephens:2000:CBG

[Ste00] David W. Stephens. Cumulative benefit games: Achieving cooperation when players discount the future. Journal of Theoretical Biology, 205(1):1–16, July 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920411.

Sterrer:2002:OSV

[Ste02a] Wolfgang Sterrer. On the origin of sex as vaccination. Journal of Theoretical Biology, 216(4):387–396, June 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930080.

Stevenson:2002:CEG

[Ste02b] David S. Stevenson. Co-evolution of the genetic code and ribozyme replication. Journal of Theoretical Biology, 217 (2):235–253, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519302930134.

Steuer:2004:ESM

[Ste04] Ralf Steuer. Effects of stochasticity in models of the cell cycle: from quantized cycle times to noise-induced oscillations. Journal of Theoretical Biology, 228(3):293–301, June 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000323.

Steel:2009:BLI

[Ste09] Mike Steel. A basic limitation on inferring phylogenies by pairwise sequence comparisons. Journal of Theoretical Biology, 256(3):467–472, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308005390.

Shi:2008:VLB

[STK08] Veronica Shi, Abdessamad Tridane, and Yang Kuang. A viral load-based cellular automata approach to modeling HIV dynamics and drug treatment. Journal of Theoretical Biology, 253(1):24–35, July 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005619.

Skubic:2004:WPD

[STMH04] Eva Skubic, Michael Taborsky, John M. McNamara, and Alasdair I. Houston. When to parasitize? A dynamic optimization model of reproductive strategies in a cooperative breeder. Journal of Theoretical Biology, 227(4):487–501, April 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004430.

Stoecker:2009:MCB

[STSD09] Urs Stoecker, Ivo A. Telley, Edgar Stüssi, and Jachen Denoth. A multisegmental cross-bridge kinetics model of the myofibril. Journal of Theoretical Biology, 259(4):714–726, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001350.

Shao:2009:PDR

[STW^{+}09] Xiaojian Shao, Yingjie Tian, Lingyun Wu, Yong Wang, Ling Jing, and Naiyang Deng. Predicting DNA- and RNA-binding proteins from sequences with kernel methods. Journal of Theoretical Biology, 258(2):289–293, May 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000289.

Sugimoto:2002:DED

[Sug02] Takeshi Sugimoto. Darwinian evolution does not rule out the Gaia hypothesis. Journal of Theoretical Biology, 218(4):447–455, October 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930912.

Suresh:2003:CTS

[SV03] M. Xavier Suresh and K. Veluraja. Conformations of terminal sialyloligosaccharide fragments — a molecular dynamics study. Journal of Theoretical Biology, 222(3):389–402, June 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930300050X.

Schwacke:2005:CAT

[SV05] John H. Schwacke and Eberhard O. Voit. Computation and analysis of time-dependent sensitivities in generalized mass action systems. Journal of Theoretical Biology, 236(1):21–38, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000731.

Schwacke:2007:PSI

[SV07] John H. Schwacke and Eberhard O. Voit. The potential for signal integration and processing in interacting MAP kinase cascades. Journal of Theoretical Biology, 246(4):604–620, June 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306006035.

Sudar:2008:HMS

[SVG^{+}08] Emina Sudar, Jelena Velebit, Zoran Gluvic, Zorica Zakula, Emilija Lazic, Ljiljana Vuksanovic-Topic, Biljana Putnikovic, Aleksandar Neskovic, and Esma R. Isenovic. Hypothetical mechanism of sodium pump regulation by estradiol under primary hypertension. Journal of Theoretical Biology, 251(4):584–592, April 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308000040.

Serra:2007:WSM

[SVGK07] R. Serra, M. Villani, A. Graudenzi, and S. A. Kauffman. Why a simple model of genetic regulatory networks describes the distribution of avalanches in gene expression data. Journal of Theoretical Biology, 246(3):449–460, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000288.

Soboleva:2005:MMP

[SVN^{+}05] T. K. Soboleva, I. Vetharaniam, A. J. Nixon, R. Montenegro, A. J. Pearson, and J. Sneyd. Mathematical modelling of prolactin-receptor interaction and the corollary for prolactin receptor gene expression in skin. Journal of Theoretical Biology, 234(2):289–298, May 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400579X.

Shorten:2002:IMF

[SVS^{+}02] P. R. Shorten, I. Vetharaniam, T. K. Soboleva, G. C. Wake, and S. R. Davis. Influence of milking frequency on mammary gland dynamics. Journal of Theoretical Biology, 218(4):521–530, October 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930973.

Serra:2004:GNM

[SVS04] R. Serra, M. Villani, and A. Semeria. Genetic network models and statistical properties of gene expression data in knock-out experiments. Journal of Theoretical Biology, 227(1):149–157, March 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003965.

Samuel:2003:CBP

[SW03] Stuart A. Samuel and Gezhi Weng. Characterization of a branch of the phylogenetic tree. Journal of Theoretical Biology, 220(4):457–468, February 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931328.

Schulgasser:2004:HC

[SW04a] Kalman Schulgasser and Allan Witztum. The hierarchy of chirality. Journal of Theoretical Biology, 230(2):281–288, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002589.

Schulgasser:2004:SU

[SW04b] Kalman Schulgasser and Allan Witztum. Spiralling upward. Journal of Theoretical Biology, 230(2):275–280, September 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002577.

Sato:2008:ASE

[SW08a] M. Sato and D. Waxman. Adaptation to slow environmental change, with apparent anticipation of selection. Journal of Theoretical Biology, 252(1):166–172, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307006315.

Schlosser:2008:SMC

[SW08b] Gerhard Schlosser and Günter P. Wagner. A simple model of co-evolutionary dynamics caused by epistatic selection. Journal of Theoretical Biology, 250(1):48–65, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004146.

Slevin:2008:TAI

[SW08c] Keith Slevin and John Waterlow. Theoretical analysis of the importance of recycling in measurements of protein turnover by constant infusion of a labelled amino acid. Journal of Theoretical Biology, 253(1):215–219, July 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800091X.

Steinmeyer:2009:LMS

[SW09] Shelby H. Steinmeyer and Claus O. Wilke. Lethal mutagenesis in a structured environment. Journal of Theoretical Biology, 261(1):67–73, November 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003270.

Song:2006:ESC

[SWB06] Jiangning Song, Minglei Wang, and Kevin Burrage. Exploring synonymous codon usage preferences of disulfide-bonded and non-disulfide bonded cysteines in the E. coli genome. Journal of Theoretical Biology, 241(2):390–401, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500528X.

Sedaghat:2008:CTP

[SWC^{+}08] H. Sedaghat, M. A. Wood, J. W. Cain, C. K. Cheng, C. M. Baumgarten, and D. M. Chan. Complex temporal patterns of spontaneous initiation and termination of reentry in a loop of cardiac tissue. Journal of Theoretical Biology, 254(1):14–26, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002397.

Seki:2007:TSE

[SWI07] Motohide Seki, Joe Yuichiro Wakano, and Yasuo Ihara. A theoretical study on the evolution of male parental care and female multiple mating: Effects of female mate choice and male care bias. Journal of Theoretical Biology, 247(2):281– 296, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001154.

Schwarz:2009:PMC

[SWM09] Roland Schwarz, Matthias Wolf, and Tobias Müller. A probabilistic model of cell size reduction in pseudo-nitzschia delicatissima (bacillariophyta). Journal of Theoretical Biology, 258(2):316–322, May 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900054X.

Savage:2007:GMT

[SWN07] Van M. Savage, Colleen T. Webb, and Jon Norberg. A general multi-trait-based framework for studying the effects of biodiversity on ecosystem functioning. Journal of Theoretical Biology, 247(2):213–229, July 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001129.

Shen:2003:CPA

[SWRH03] Fanyi Shen, Liu Wenji, Gao Rongfu, and Huo Hu. A careful physical analysis of gas bubble dynamics in xylem. Journal of Theoretical Biology, 225(2):229–233, November 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002418.

Savill:2002:MMN

[SWS02] Nicholas J. Savill, Richard Weller, and Jonathan A. Sherratt. Mathematical modelling of nitric oxide regulation of rete peg formation in psoriasis. Journal of Theoretical Biology, 214(1):1–16, January 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924002.

Song:2006:PDC

[SX06] Xinyu Song and Zhongyi Xiang. The prey-dependent consumption two-prey one-predator models with stage structure for the predator and impulsive effects. Journal of Theoretical Biology, 242(3):683–698, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001913.

Shen:2006:FKP

[SYC06] Hong-Bin Shen, Jie Yang, and Kuo-Chen Chou. Fuzzy KNN for predicting membrane protein types from pseudo-amino acid composition. Journal of Theoretical Biology, 240(1):9– 13, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003565.

Sylvester:2006:LDO

[Syl06] Adam D. Sylvester. Locomotor decoupling and the origin of hominin bipedalism. Journal of Theoretical Biology, 242(3):581–590, October 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001779.

Srivastava:2002:SVD

[SYSY02] R. Srivastava, L. You, J. Summers, and J. Yin. Stochastic vs. deterministic modeling of intracellular viral kinetics. Journal of Theoretical Biology, 218(3):309–321, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930293078X.

Shudo:2007:RST

[SYYI07] Emi Shudo, Jin Yang, Akihiko Yoshimura, and Yoh Iwasa. Robustness of the signal transduction system of the mammalian JAK/STAT pathway and dimerization steps. Journal of Theoretical Biology, 246(1):1–9, May 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005595.

Scheuring:2001:DEC

[SZ01] István Scheuring and Orsolya E. Zeöld. Data estimation and the colour of time series. Journal of Theoretical Biology, 213(3):427–434, December 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924282.

Santillan:2004:DIF

[SZ04] Moisés Santillán and Eduardo S. Zeron. Dynamic influence of feedback enzyme inhibition and transcription attenuation on the tryptophan operon response to nutritional shifts. Journal of Theoretical Biology, 231(2):287–298, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003145.

Szymanska:2009:MMH

[SZ09] Zuzanna Szymańska and Maciej Zylicz. Mathematical modeling of heat shock protein synthesis in response to temperature change. Journal of Theoretical Biology, 259(3):562–569, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001180.

Szamado:2003:TDH

[Szá03] Szabolcs Számadó. Threat displays are not handicaps. Journal of Theoretical Biology, 221(3):327–348, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931766.

Sanchez:2003:HSB

[SZC^{+}03] Juan Armando Sánchez, Wei Zeng, Vitor R. Coluci, Carl Simpson, and Howard R. Lasker. How similar are branching networks in nature? A view from the ocean: Caribbean gorgonian corals. Journal of Theoretical Biology, 222(1):135–138, May 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000171.

Schulin-Zeuthen:2005:EPI

[SZLK^{+}05] M. Schulin-Zeuthen, J. B. Lopes, E. Kebreab, D. M. S. S. Vitti, A. L. Abdalla, M. DeL. Haddad, L. A. Crompton, and J. France. Effects of phosphorus intake on phosphorus flow in growing pigs: Application and comparison of two models. Journal of Theoretical Biology, 236(2):115–125, September 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000810.

Surovtsev:2009:MMM

[SZLM09] Ivan V. Surovtsev, Zhigang Zhang, Paul A. Lindahl, and Jeffrey J. Morgan. Mathematical modeling of a minimal protocell with coordinated growth and division. Journal of Theoretical Biology, 260(3):422–429, October 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002616.

Takai:2006:CPP

[Tak06] Kazuyuki Takai. Classification of the possible pairs between the first anticodon and the third codon positions based on a simple model assuming two geometries with which the pairing effectively potentiates the decoding complex. Journal of Theoretical Biology, 242(3):564–580, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001664.

Tannenbaum:2001:SC

[Tan01] Arnold S. Tannenbaum. The sense of consciousness. Journal of Theoretical Biology, 211(4):377–391, August 21, 2001.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923550.

Tanimoto:2007:DTS

[Tan07a] Jun Tanimoto. Does a tag system effectively support emerging cooperation? Journal of Theoretical Biology, 247(4):756–764, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001580.

Tannenbaum:2007:EVL

[Tan07b] Emmanuel Tannenbaum. Extracting viability landscapes from mutagen-response experiments. Journal of Theoretical Biology, 245(1):37–43, March 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004012.

Tannenbaum:2007:WDD

[Tan07c] Emmanuel Tannenbaum. When does division of labor lead to increased system output? Journal of Theoretical Biology, 247(3):413–425, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001282.

Tang:2008:MNS

[Tan08] Moxun Tang. The mean and noise of stochastic gene transcription. Journal of Theoretical Biology, 253(2):271–280, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001471.

Tarnita:2009:MSE

[TAN09] Corina E. Tarnita, Tibor Antal, and Martin A. Nowak. Mutation-selection equilibrium in games with mixed strategies. Journal of Theoretical Biology, 261(1):50–57, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003415.

Tao:2004:IEN

[Tao04a] Yi Tao. Intrinsic and external noise in an auto-regulatory genetic network. Journal of Theoretical Biology, 229(2):147–156, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001122.

Tao:2004:ING

[Tao04b] Yi Tao. Intrinsic noise, gene regulation and steady-state statistics in a two-gene network. Journal of Theoretical Biology, 231(4):563–568, December 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003339.

Tasaki:2002:SDS

[Tas02] Ichiji Tasaki. Spread of discrete structural changes in synthetic polyanionic gel: a model of propagation of a nerve impulse. Journal of Theoretical Biology, 218(4):497–505, October 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930293095X.

Tasaki:2005:RAS

[Tas05] Ichiji Tasaki. Repetitive abrupt structural changes in polyanionic gels: a comparison with analogous processes in nerve fibers. Journal of Theoretical Biology, 236(1):2–11, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000718.

Taylor:2000:SEF

[Tay00] David Taylor. Scaling effects in the fatigue strength of bones from different animals. Journal of Theoretical Biology, 206(2):299–306, September 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921258.

Taylor:2006:MMO

[Tay06] William R. Taylor. A molecular model for the origin of protein translation in an RNA world. Journal of Theoretical Biology, 243(3):393–406, December 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002517.

Tian:2004:BSL

[TB04] Tianhai Tian and Kevin Burrage. Bistability and switching in the lysis/lysogeny genetic regulatory network of bacteriophage λ. Journal of Theoretical Biology, 227(2):229–237, March 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930300420X.

Turner:2006:SSM

[TBB^{+}06] Joanne Turner, Roger G. Bowers, Michael Begon, Susan E. Robinson, and Nigel P. French. A semi-stochastic model of the transmission of Escherichia coli o157 in a typical UK dairy herd: Dynamics, sensitivity analysis and intervention/prevention strategies. Journal of Theoretical Biology, 241(4):806–822, August 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000336.

Turner:2008:NMC

[TBC^{+}08] J. Turner, R. G. Bowers, D. Clancy, M. C. Behnke, and R. M. Christley. A network model of E. coli o157 transmission within a typical UK dairy herd: The effect of heterogeneity and clustering on the prevalence of infection. Journal of Theoretical Biology, 254(1):45–54, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308002452.

Travis:2006:DPN

[TBCD06] J. M. J. Travis, R. W. Brooker, E. J. Clark, and C. Dytham. The distribution of positive and negative species interactions across environmental gradients on a dual-lattice model. Journal of Theoretical Biology, 241(4):896–902, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000385.

Thornley:2008:HBH

[TBR08] Simon Thornley, Chris Bullen, and Mick Roberts. Hepatitis B in a high prevalence New Zealand population: a mathematical model applied to infection control policy. Journal of Theoretical Biology, 254(3):599–603, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003342.

Theraulaz:2002:TPP

[TBS^{+}02] Guy Theraulaz, Eric Bonabeau, Ricard V. Solé, Bertrand Schatz, and Jean-Louis Deneubourg. Task partitioning in a ponerine ant. Journal of Theoretical Biology, 215(4):481–489, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925184.

Tournier:2009:UOI

[TC09] L. Tournier and M. Chaves. Uncovering operational interactions in genetic networks using asynchronous Boolean dynamics. Journal of Theoretical Biology, 260(2):196–209, September 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002653.

Tozeren:2005:OEC

[TCP05] Aydin Tözeren, Charles W. Coward, and Sokol P. Petushi. Origins and evolution of cell phenotypes in breast tumors. Journal of Theoretical Biology, 233(1):43–54, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004497.

Taylor:2007:IFF

[TDW07] Peter D. Taylor, Troy Day, and Geoff Wild. From inclusive fitness to fixation probability in homogeneous structured populations. Journal of Theoretical Biology, 249(1):101–110, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003323.

Tzafriri:2004:TQS

[TE04] A. R. Tzafriri and E. R. Edelman. The total quasi-steady-state approximation is valid for reversible enzyme kinetics. Journal of Theoretical Biology, 226(3):303–313, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303003382.

Tzafriri:2005:VQS

[TE05] A. Rami Tzafriri and Elazer R. Edelman. On the validity of the quasi-steady state approximation of bimolecular reactions in solution. Journal of Theoretical Biology, 233(3):343–350, April 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004850.

Tzafriri:2007:QSS

[TE07] A. Rami Tzafriri and Elazer R. Edelman. Quasi-steady-state kinetics at enzyme and substrate concentrations in excess of the Michaelis–Menten constant. Journal of Theoretical Biology, 245(4):737–748, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005650.

Tennenbaum:2008:SCF

[Ten08] Stephen Tennenbaum. Simple criteria for finding (nearly) optimal vaccination strategies. Journal of Theoretical Biology, 250(4):673–683, February 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307005139.

Takeuchi:2003:STC

[TFYY03] Fumihiko Takeuchi, Yasuhiro Futamura, Hiroshi Yoshikura, and Kenji Yamamoto. Statistics of trinucleotides in coding sequences and evolution. Journal of Theoretical Biology, 222(2):139–149, May 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000213.

Tyson:2004:AW

[TG04] John J. Tyson and Leon Glass. Arthur T. Winfree (1942– 2002). Journal of Theoretical Biology, 230(4):433–439, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304002346.

Tchuraev:2009:GEN

[TG09a] R. N. Tchuraev and A. V. Galimzyanov. Gene and epigene networks: Two levels in organizing the hereditary system. Journal of Theoretical Biology, 259(4):659–669, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001404.

Thomson:2009:RPT

[TG09b] Matthew Thomson and Jeremy Gunawardena. The rational parameterisation theorem for multisite post-translational modification systems. Journal of Theoretical Biology, 261 (4):626–636, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519309004238.

Torabi:2007:CEM

[TGN07] Noorossadat Torabi, Hani Goodarzi, and Hamed Shateri Najafabadi. The case for an error minimizing set of coding amino acids. Journal of Theoretical Biology, 244(4):737–744, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004310.

Tralau:2000:MTT

[TGP^{+}00] Carola Tralau, Gerhard Greller, Markus Pajatsch, Winfried Boos, and Erich Bohl. Mathematical treatment of transport data of bacterial transport systems to estimate limitation in diffusion through the outer membrane. Journal of Theoretical Biology, 207(1):1–14, November 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921404.

Topham:2000:KBG

[TGR^{+}00] Christopher M. Topham, Sheraz Gul, Marina Resmini, Sanjiv Sonkaria, Gerrard Gallacher, and Keith Brocklehurst. The kinetic basis of a general method for the investigation of active site content of enzymes and catalytic antibodies: First-order behaviour under single-turnover and cycling conditions. Journal of Theoretical Biology, 204(2):239–256, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920113.

Tongen:2006:BMA

[TGT06] Anthony Tongen, Alain Goriely, and Michael Tabor. Biomechanical model for appressorial design in magnaporthe grisea. Journal of Theoretical Biology, 240(1):1–8, May 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003541.

Tsoukias:2007:CMO

[TGV^{+}07] Nikolaos M. Tsoukias, Daniel Goldman, Arjun Vadapalli, Roland N. Pittman, and Aleksander S. Popel. A computational model of oxygen delivery by hemoglobin-based oxygen carriers in three-dimensional microvascular networks. Journal of Theoretical Biology, 248(4):657–674, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002913.

Taylor:2000:PGM

[TH00] Christopher F. Taylor and Paul G. Higgs. A population genetics model for multiple quantitative traits exhibiting pleiotropy and epistasis. Journal of Theoretical Biology, 203(4):419–437, April 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910944.

Thompson:2003:ASS

[TH03a] Matthew V. Thompson and N. Michele Holbrook. Application of a single-solute non-steady-state phloem model to the study of long-distance assimilate transport. Journal of Theoretical Biology, 220(4):419–455, February 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931158.

Tufto:2003:ESM

[TH03b] Jarle Tufto and Kjetil Hindar. Effective size in management and conservation of subdivided populations. Journal of Theoretical Biology, 222(3):273–281, June 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000183.

Takahashi:2005:CCS

[TH05] Satoshi Takahashi and Michio Hori. Coexistence of competing species by the oscillation of polymorphisms. Journal of Theoretical Biology, 235(4):591–596, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000743.

Twarock:2006:CVC

[TH06] R. Twarock and R. W. Hendrix. Crosslinking in viral capsids via tiling theory. Journal of Theoretical Biology, 240(3):419– 424, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930500439X.

Theodoridis:2000:SMD

[The00] George C. Theodoridis. On selection mechanisms during initial evolution and on the possible role of ionic balance. Journal of Theoretical Biology, 207(1):15–20, November 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300921520.

Thierie:2004:MTP

[Thi04] Jacques Thierie. Modeling threshold phenomena, metabolic pathways switches and signals in chemostat-cultivated cells: the crabtree effect in saccharomyces cerevisiae. Journal of Theoretical Biology, 226(4):483–501, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003837.

Takeuchi:2007:TCM

[THKU07] Tsutomu Takeuchi, Takamichi Hinohara, Gen Kurosawa, and Kenko Uchida. A temperature-compensated model for circadian rhythms that can be entrained by temperature cycles. Journal of Theoretical Biology, 246(1):195–204, May 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005960.

Taylor:2003:CTD

[THL03] D. Taylor, J. G. Hazenberg, and T. C. Lee. The cellular transducer in damage-stimulated bone remodelling: a theoretical investigation using fracture mechanics. Journal of Theoretical Biology, 225(1):65–75, November 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002224.

Thompson:2000:USC

[Tho00] James R. Thompson. Is the United States country zero for the first-world AIDS epidemic? Journal of Theoretical Biology, 204(4):621–628, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920290.

Thoms:2002:HBC

[Tho02] Silke Thoms. Hydrogen bonds and the catalytic mechanism of human carbonic anhydrase II. Journal of Theoretical Biology, 215(4):399–404, April 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302925282.

Thompson:2005:SPT

[Tho05] Matthew V. Thompson. Scaling phloem transport: Elasticity and pressure-concentration waves. Journal of Theoretical Biology, 236(3):229–241, October 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001189.

Takigawa-Imamura:2006:TAP

[TIM06] Hisako Takigawa-Imamura and Atsushi Mochizuki. Transcriptional autoregulation by phosphorylated and nonphosphorylated KaiC in cyanobacterial circadian rhythms. Journal of Theoretical Biology, 241(2):178–192, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004947.

Taylor:2006:SFT

[TIN06] Christine Taylor, Yoh Iwasa, and Martin A. Nowak. A symmetry of fixation times in evoultionary dynamics. Journal of Theoretical Biology, 243(2):245–251, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002529.

Traulsen:2007:FET

[TIN07] Arne Traulsen, Yoh Iwasa, and Martin A. Nowak. The fastest evolutionary trajectory. Journal of Theoretical Biology, 249(3):617–623, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003955.

Thar:2004:PER

[TK04] Roland Thar and Michael Kühl. Propagation of electromagnetic radiation in mitochondria? Journal of Theoretical Biology, 230(2):261–270, September 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002498.

Takagi:2005:DSB

[TK05] Hiroaki Takagi and Kunihiko Kaneko. Dynamical systems basis of metamorphosis: diversity and plasticity of cellular states in reaction diffusion network. Journal of Theoretical Biology, 234(2):173–186, May 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005685.

Tanaka:2008:MCR

[TK08] Reiko J. Tanaka and Hidenori Kimura. Mathematical classification of regulatory logics for compound environmental changes. Journal of Theoretical Biology, 251(2):363–379, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307005942.

Tildesley:2009:CGP

[TK09a] Michael J. Tildesley and Matt J. Keeling. Corrigendum to “Is R_{0} a good predictor of final epidemic size: Foot-and-mouth disease in the UK”: [J. Theor. Biol. 258 (2009) 623– 629]. Journal of Theoretical Biology, 259(4):863, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002458. See [TK09b].

Tildesley:2009:GPF

[TK09b] Michael J. Tildesley and Matt J. Keeling. Is R_{0} a good predictor of final epidemic size: Foot-and-mouth disease in the UK. Journal of Theoretical Biology, 258(4):623–629, June 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000824. See corri-

gendum [TK09a].

Tomioka:2004:MAN

[TKKA04] Ryota Tomioka, Hidenori Kimura, Tetsuya J. Kobayashi, and Kazuyuki Aihara. Multivariate analysis of noise in genetic regulatory networks. Journal of Theoretical Biology, 229(4):501–521, August 21, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001742.

Tindemans:2006:DVS

[TKM06] Simon H. Tindemans, Norbert Kern, and Bela M. Mulder. The diffusive vesicle supply center model for tip growth in fungal hyphae. Journal of Theoretical Biology, 238(4):937–948, February 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003061.

Tero:2007:MMA

[TKN07] Atsushi Tero, Ryo Kobayashi, and Toshiyuki Nakagaki. A mathematical model for adaptive transport network in path finding by true slime mold. Journal of Theoretical Biology, 244(4):553–564, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600289X.

Tolla:2007:KIM

[TKP07] Caroline Tolla, Sebastiaan A. L. M. Kooijman, and Jean-Christophe Poggiale. A kinetic inhibition mechanism for maintenance. Journal of Theoretical Biology, 244(4):576–587, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004024.

Tsai:2000:SMM

[TL00] Lu Tsai and Liaofu Luo. A statistical mechanical model for predicting b-DNA curvature and flexibility. Journal of Theoretical Biology, 207(2):177–194, November 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921623.

Tao:2002:FDS

[TL02] Yi Tao and Sabin Lessard. Frequency-dependent selection in sexual family-structured populations. Journal of Theoretical Biology, 217(4):525–534, August 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930419.

Thumfort:2000:SAM

[TLA00] Peter P. Thumfort, David B. Layzell, and Craig A. Atkins. A simplified approach for modeling diffusion into cells. Journal of Theoretical Biology, 204(1):47–65, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910713.

Tzeng:2007:MPS

[TLC07] Yun-Huei Tzeng, Wen-Hsiung Li, and Trees-Juen Chuang. Mathematical properties of some measures of evolutionary distance. Journal of Theoretical Biology, 245(4):790–792, April 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306005637.

Tlusty:2007:MEG

[Tlu07] Tsvi Tlusty. A model for the emergence of the genetic code as a transition in a noisy information channel. Journal of Theoretical Biology, 249(2):331–342, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307003669.

Tien:2005:PRM

[TLZ05] Joseph H. Tien, Danielle Lyles, and Mary Lou Zeeman. A potential role of modulating inositol 1,4,5-trisphosphate receptor desensitization and recovery rates in regulating ovulation. Journal of Theoretical Biology, 232(1):105–117, January 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003741.

Tong:2008:NDA

[TLZ^{+}08] Jianbo Tong, Shuling Liu, Peng Zhou, Bulan Wu, and Zhiliang Li. A novel descriptor of amino acids and its application in peptide QSAR. Journal of Theoretical Biology, 253(1):90– 97, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000969.

Thoumine:2000:PML

[TM00] Olivier Thoumine and Jean-Jacques Meister. A probabilistic model for ligand-cytoskeleton transmembrane adhesion: Predicting the behavior of microspheres on the surface of migrating cells. Journal of Theoretical Biology, 204(3):381–392, June 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920241.

Takahashi:2006:ISP

[TM06] Nobuyuki Takahashi and Rie Mashima. The importance of subjectivity in perceptual errors on the emergence of indirect reciprocity. Journal of Theoretical Biology, 243(3):418–436, December 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001974.

Travis:2009:AIR

[TMBD09] Justin M. J. Travis, Karen Mustin, Tim G. Benton, and Calvin Dytham. Accelerating invasion rates result from the evolution of density-dependent dispersal. Journal of Theoretical Biology, 259(1):151–158, July 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001155.

Torres-Montaner:2004:HAN

[TMH04] Antonio Torres-Montaner and David Hughes. A hypothetical anti-neoplastic mechanism associated to reserve cells. Journal of Theoretical Biology, 231(2):239–248, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002917.

Takano:2003:PTS

[TMI03a] Ryohji Takano, Atsushi Mochizuki, and Yoh Iwasa. Possibility of tissue separation caused by cell adhesion. Journal of Theoretical Biology, 221(4):459–474, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931936.

Tohya:2003:DRC

[TMI03b] Shusaku Tohya, Atsushi Mochizuki, and Yoh Iwasa. Difference in the retinal cone mosaic pattern between zebrafish and medaka: Cell-rearrangement model. Journal of Theoretical Biology, 221(2):289–300, March 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931924.

Taghva:2002:AKE

[TML02] Alexander Taghva, Yunmei Ma, and Michael R. Lieber. Analysis of the kinetic and equilibrium binding of ku protein to DNA. Journal of Theoretical Biology, 214(1):85–97, January 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519301924464.

Tolic:2000:MIG

[TMS00] Iva Marija Tolić, Erik Mosekilde, and Jeppe Sturis. Modeling the insulin–glucose feedback system: The significance of pulsatile insulin secretion. Journal of Theoretical Biology, 207(3):361–375, December 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300921805.

Tyson:2001:REC

[TN01] John J. Tyson and Bela Novak. Regulation of the eukaryotic cell cycle: Molecular antagonism, hysteresis, and irreversible transitions. Journal of Theoretical Biology, 210(2):249–263, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922933.

Tao:2004:MGH

[TN04] L. Tao and C. Nicholson. Maximum geometrical hindrance to diffusion in brain extracellular space surrounding uniformly spaced convex cells. Journal of Theoretical Biology, 229(1):59– 68, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001043.

Traulsen:2007:SPE

[TNP07] Arne Traulsen, Martin A. Nowak, and Jorge M. Pacheco. Stochastic payoff evaluation increases the temperature of selection. Journal of Theoretical Biology, 244(2):349–356, January 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003511.

Tyson:2008:MRB

[TNTJ08] R. Tyson, K. D. Newton, H. Thistlewood, and G. Judd. Mating rates between sterile and wild codling moths (Cydia pomonella) in springtime: a simulation study. Journal of Theoretical Biology, 254(2):319–330, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308002439.

Tan:2006:DNS

[TO06] Ning Tan and Qi Ouyang. Design of a network with state stability. Journal of Theoretical Biology, 240(4):592–598, June 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305004716.

Tarnita:2009:SSS

[TOA^{+}09] Corina E. Tarnita, Hisashi Ohtsuki, Tibor Antal, Feng Fu, and Martin A. Nowak. Strategy selection in structured populations. Journal of Theoretical Biology, 259(3):570–581, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001374.

Tracqui:2008:TAA

[TOB08] P. Tracqui, J. Ohayon, and T. Boudou. Theoretical analysis of the adaptive contractile behaviour of a single cardiomyocyte cultured on elastic substrates with varying stiffness. Journal of Theoretical Biology, 255(1):92–105, November 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003949.

Tofilski:2006:ICP

[Tof06] Adam Tofilski. Influence of caste polyethism on longevity of workers in social insect colonies. Journal of Theoretical Biology, 238(3):527–531, February 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002535.

Toro:2000:CSR

[Tor00] Miguel Toro. A comment on sex ratio with intrasex sibling cooperation. Journal of Theoretical Biology, 205(1):161–164, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920484.

Torres:2001:BPL

[Tor01] Jose-Leonel Torres. Biological power laws and Darwin’s principle. Journal of Theoretical Biology, 209(2):223–232, March 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300922586.

Toyoizumi:2009:SPA

[Toy09] Hiroshi Toyoizumi. Sample path analysis of contribution and reward in cooperative groups. Journal of Theoretical Biology, 256(3):311–314, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005134.

Trottier:2005:SPC

[TP05] Helen Trottier and Pierre Philippe. Scaling properties of childhood infectious diseases epidemics before and after mass vaccination in Canada. Journal of Theoretical Biology, 235(3):326–337, August 7, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000391.

Topp:2000:MCM

[TPD^{+}00] Brian Topp, Keith Promislow, Gerda Devries, Robert M. Miura, and Diane T. Finegood. A model of β-cell mass, insulin, and glucose kinetics: Pathways to diabetes. Journal of Theoretical Biology, 206(4):605–619, October 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921507.

Traulsen:2007:PCS

[TPN07] Arne Traulsen, Jorge M. Pacheco, and Martin A. Nowak. Pairwise comparison and selection temperature in evolutionary game dynamics. Journal of Theoretical Biology, 246(3):522– 529, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000069.

Takagi:2004:PAR

[TPP^{+}04] S. Takagi, A. Pumir, D. Pazó, I. Efimov, V. Nikolski, and V. Krinsky. A physical approach to remove anatomical reentries: a bidomain study. Journal of Theoretical Biology, 230(4):489–497, October 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304002395.

Thoms:2001:MCC

[TPWG01] Silke Thoms, Markus Pahlow, and Dieter A. Wolf-Gladrow. Model of the carbon concentrating mechanism in chloroplasts of eukaryotic algae. Journal of Theoretical Biology, 208(3):295–313, February 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922197.

Toth:2007:MET

[TQUN07] Attila Tóth, Ethel Queralt, Frank Uhlmann, and Béla Novák. Mitotic exit in two dimensions. Journal of Theoretical Biology, 248(3):560–573, October 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307002949. See erratum [TQUN08].

Toth:2008:EME

[TQUN08] Attila Tóth, Ethel Queralt, Frank Uhlmann, and Béla Novák. Erratum to “Mitotic exit in two dimensions”: [J. Theor. Biol. 248 (2007) 560–573]. Journal of Theoretical Biology, 250(1):219, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004456. See [TQUN07].

Tsibidis:2008:IBM

[TR08] George D. Tsibidis and Jorge Ripoll. Investigation of binding mechanisms of nuclear proteins using confocal scanning laser microscopy and FRAP. Journal of Theoretical Biology, 253(4):755–768, August 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001872.

Torralba:2003:CPD

[TRM03a] Antonio S. Torralba, Yoel Rodríguez, and Francisco Montero. Corrigendum to “Phenomenological definition of response times with application to metabolic reactions”: [J. theor. Biol. 221, 475–489]. Journal of Theoretical Biology, 223(4):533, August 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930300136X. See [TRM03c, TRM03b].

Torralba:2003:EPD

[TRM03b] Antonio S. Torralba, Yoel Rodríguez, and Francisco Montero. Erratum to “Phenomenological definition of response times with application to metabolic reactions”: [J. theor. Biol. 221, 475–489]. Journal of Theoretical Biology, 223(4):535, August 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303001371. See [TRM03c,

TRM03a].

Torralba:2003:PDR

[TRM03c] Antonio S. Torralba, Yoel Rodriguez, and Francisco Montero. Phenomenological definition of response times with application to metabolic reactions. Journal of Theoretical Biology, 221(4):475–489, April 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519303931961. See corrigendum [TRM03a] and erratum [TRM03b].

Turner:2002:IAC

[TS02] Stephen Turner and Jonathan A. Sherratt. Intercellular adhesion and cancer invasion: a discrete simulation using the extended Potts model. Journal of Theoretical Biology, 216(1):85– 100, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925226.

Thatte:2008:RPS

[TS08] Bhalchandra D. Thatte and Mike Steel. Reconstructing pedigrees: a stochastic perspective. Journal of Theoretical Biology, 251(3):440–449, April 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307006066.

Tschirhart:2000:GEE

[Tsc00] John Tschirhart. General equilibrium of an ecosystem. Journal of Theoretical Biology, 203(1):13–32, March 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910585.

Turner:2004:TTF

[TSC04] Stephen Turner, Jonathan A. Sherratt, and David Cameron. Tamoxifen treatment failure in cancer and the nonlinear dynamics of TGF β. Journal of Theoretical Biology, 229(1):101– 111, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001079.

Tanaka:2009:HAE

[TSK09] Mark M. Tanaka, Scott A. Sisson, and Garry C. King. High affinity extremes in combinatorial libraries and repertoires. Journal of Theoretical Biology, 261(2):260–265, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309003531.

Traulsen:2005:SED

[TSN05] Arne Traulsen, Anirvan M. Sengupta, and Martin A. Nowak. Stochastic evolutionary dynamics on two levels. Journal of Theoretical Biology, 235(3):393–401, August 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000469.

Till:2008:CIE

[TSRB08] Olaf Till, Tobias Siebert, Christian Rode, and Reinhard Blickhan. Characterization of isovelocity extension of activated muscle: a Hill-type model for eccentric contractions and a method for parameter determination. Journal of Theoretical Biology, 255(2):176–187, November 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004244.

Tannenbaum:2006:SQE

[TSS06] Emmanuel Tannenbaum, James L. Sherley, and Eugene I. Shakhnovich. Semiconservative quasispecies equations for polysomic genomes: The haploid case. Journal of Theoretical Biology, 241(4):791–805, August 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000282.

Tiedemann:2007:CBS

[TSZ^{+}07] Hendrik B. Tiedemann, Elida Schneltzer, Stefan Zeiser, Isabel Rubio-Aliaga, Wolfgang Wurst, Johannes Beckers, Gerhard K. H. Przemeck, and Martin Hrabé de Angelis. Cell-based simulation of dynamic expression patterns in the presomitic mesoderm. Journal of Theoretical Biology, 248(1):120–129, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002457.

Thomas:2001:AFD

[TT01] Adrian L. R. Thomas and Graham K. Taylor. Animal flight dynamics I. Stability in gliding flight. Journal of Theoretical Biology, 212(3):399–424, October 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923872.

Taylor:2002:AFD

[TT02] G. K. Taylor and A. L. R. Thomas. Animal flight dynamics II. Longitudinal stability in flapping flight. Journal of Theoretical Biology, 214(3):351–370, February 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924701.

Taylor:2003:AFD

[TT03] G. K. Taylor and A. L. R. Thomas. “Animal Flight Dynamics II. Longitudinal Stability in Flapping flight” [Journal of Theoretical Biology 214 (2002) 351–370]. Journal of Theoretical Biology, 221(4):671, April 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930393205X.

Taylor:2004:SIV

[TT04] David Taylor and Antoine Tilmans. Stress intensity variations in bone microcracks during the repair process. Journal of Theoretical Biology, 229(2):169–177, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001146.

Trommershauser:2001:VEC

[TTKZ01] J. Trommershäuser, S. Titz, B. U. Keller, and A. Zippelius. Variability of excitatory currents due to single-channel noise, receptor number and morphological heterogeneity. Journal of Theoretical Biology, 208(3):329–343, February 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922227.

Tao:2005:CCI

[TTN05] A. Tao, L. Tao, and C. Nicholson. Cell cavities increase tortuosity in brain extracellular space. Journal of Theoretical Biology, 234(4):525–536, June 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304006095.

Tenhumberg:2000:SVF

[TTR00] Brigitte Tenhumberg, Andrew J. Tyre, and Bernie Roitberg. Stochastic variation in food availability influences weight and age at maturity. Journal of Theoretical Biology, 202(4):257–272, February 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910494.

Takamatsu:2009:EDM

[TTT09] Atsuko Takamatsu, Eri Takaba, and Ginjiro Takizawa. Environment-dependent morphology in plasmodium of true slime mold Physarum polycephalum and a network growth model. Journal of Theoretical Biology, 256(1):29–44, January 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308004736.

Thornhill:2000:CPR

[TU00] Richard H. Thornhill and David W. Ussery. A classification of possible routes of Darwinian evolution. Journal of Theoretical Biology, 203(2):111–116, March 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910701.

Tullberg:2003:RSB

[Tul03] Jan Tullberg. Rationality and social behavior. Journal of Theoretical Biology, 224(4):469–478, October 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001942.

Turin:2002:MCO

[Tur02] L. Turin. A method for the calculation of odor character from molecular structure. Journal of Theoretical Biology, 216(3):367–385, June 7, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301925044.

Thaller:2004:RBH

[TW04] S. Thaller and H. Wagner. The relation between Hill’s equation and individual muscle properties. Journal of Theoretical Biology, 231(3):319–332, December 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003182.

Twarock:2004:TAV

[Twa04] R. Twarock. A tiling approach to virus capsid assembly explaining a structural puzzle in virology. Journal of Theoretical Biology, 226(4):477–482, February 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003928.

Tzafriri:2004:ACM

[TWE04] Abraham R. Tzafriri, David Wu, and Elazer R. Edelman. Analysis of compartmental models of ligand-induced endocytosis. Journal of Theoretical Biology, 229(1):127–138, July 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001109.

Tindall:2009:CRM

[TWO^{+}09] M. J. Tindall, J. A. D. Wattis, B. J. O’Malley, L. Pickersgill, and K. G. Jackson. A continuum receptor model of hepatic lipoprotein metabolism. Journal of Theoretical Biology, 257(3):371–384, April 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930800605X.

Takeuchi:2006:ERV

[TY06] Fumihiko Takeuchi and Kenji Yamamoto. Effectiveness of realistic vaccination strategies for contact networks of various degree distributions. Journal of Theoretical Biology, 243(1):39–47, November 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306002128.

Tanaka:2009:PPR

[TY09] Yoshinari Tanaka and Masafumi Yoshino. Predicting the phenotypic response of resource-competing communities to environmental change. Journal of Theoretical Biology, 257(4):627– 641, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006541.

Tsumoto:2006:BMM

[TYI^{+}06] Kunichika Tsumoto, Tetsuya Yoshinaga, Hitoshi Iida, Hiroshi Kawakami, and Kazuyuki Aihara. Bifurcations in a mathematical model for circadian oscillations of clock genes. Journal of Theoretical Biology, 239(1):101–122, March 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500319X.

Tyrcha:2001:ADC

[Tyr01] Joanna Tyrcha. Age-dependent cell cycle models. Journal of Theoretical Biology, 213(1):89–101, November 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924038.

Telschow:2005:BCI

[TYW05] Arndt Telschow, Norio Yamamura, and John H. Werren. Bidirectional cytoplasmic incompatibility and the stable coexistence of two wolbachia strains in parapatric host populations. Journal of Theoretical Biology, 235(2):265–274, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930500024X.

Tao:2007:EFR

[TZS07] Yi Tao, Xiudeng Zheng, and Yuehua Sun. Effect of feedback regulation on stochastic gene expression. Journal of Theoretical Biology, 247(4):827–836, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001403.

Utzny:2001:LTS

[UB01] C. Utzny and N. J. Burroughs. Long-term stability of diverse immunological memory. Journal of Theoretical Biology, 211(4):393–402, August 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301923525.

Uttieri:2007:RBF

[UCSZ07] Marco Uttieri, Daniela Cianelli, J. Rudi Strickler, and Enrico Zambianchi. On the relationship between fractal dimension and encounters in three-dimensional trajectories. Journal of Theoretical Biology, 247(3):480–491, August 7, 2007.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001427.

Uchida:2007:RBC

[UD07] Satoshi Uchida and Barbara Drossel. Relation between complexity and stability in food webs with adaptive behavior. Journal of Theoretical Biology, 247(4):713–722, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002019.

Unckless:2009:PGS

[UH09] Robert L. Unckless and Jeremy K. Herren. Population genetics of sexually antagonistic mitochondrial mutants under inbreeding. Journal of Theoretical Biology, 260(1):132–136, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900263X.

Ueda:2002:ICM

[UHI02] Hiroki R. Ueda, Kenzo Hirose, and Masamitsu Iino. Intercellular coupling mechanism for synchronized and noiseresistant circadian oscillators. Journal of Theoretical Biology, 216(4):501–512, June 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930006.

Ueda:2001:ROW

[UHK01] Hiroki R. Ueda, Masatoshi Hagiwara, and Hiroaki Kitano. Robust oscillations within the interlocked feedback model of drosophila circadian rhythm. Journal of Theoretical Biology, 210(4):401–406, June 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922264.

Umeda:2002:PRC

[UI02] Tamiki Umeda and Kei Inouye. Possible role of contact following in the generation of coherent motion of dictyostelium cells. Journal of Theoretical Biology, 219(3):301–308, December 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931243.

Umeda:2004:CSD

[UI04a] Tamiki Umeda and Kei Inouye. Cell sorting by differential cell motility: a model for pattern formation in Dictyostelium. Journal of Theoretical Biology, 226(2):215–224, January 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003345. See erra-

tum [UI04b].

Umeda:2004:ECS

[UI04b] Tamiki Umeda and Kei Inouye. Erratum to “Cell sorting by differential motility: a model for pattern formation in Dictyostelium”: [J.Theor. Biol. 226 (2004) 215–224]. Journal of Theoretical Biology, 228(1):145–146, May 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930400058X. See [UI04a].

Uitdehaag:2009:DGM

[Uit09] Joost C. M. Uitdehaag. The dependency game: Multiperson reciprocal sharing leads to stable cooperation which can evolve into group formation. Journal of Theoretical Biology, 260(2):253–260, September 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002719.

Uezu:2009:NFC

[UKY^{+}09] Tatsuya Uezu, Sae Kakutani, Mika Yoshida, Akino Nakajima, Takeshi Asao, and Yoshiomi Takagi. Novel features of computer-simulated clonal life of paramecium caudatum. Journal of Theoretical Biology, 258(2):281–288, May 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309000460.

Ursino:2006:TWE

[UL06] Mauro Ursino and Giuseppe-Emiliano La Cara. Travelling waves and EEG patterns during epileptic seizure: Analysis with an integrate-and-fire neural network. Journal of Theoretical Biology, 242(1):171–187, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000774.

Ueno:2008:CNI

[UM08] Taro Ueno and Naoki Masuda. Controlling nosocomial infection based on structure of hospital social networks. Journal of Theoretical Biology, 254(3):655–666, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308003512.

Uriu:2009:TWF

[UMI09] Koichiro Uriu, Yoshihiro Morishita, and Yoh Iwasa. Traveling wave formation in vertebrate segmentation. Journal of Theoretical Biology, 257(3):385–396, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900006X.

Unnikrishnan:2009:HIF

[Unn09] M. K. Unnikrishnan. How is the individuality of a face recognized? Journal of Theoretical Biology, 261(3):469–474, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900383X.

Usami:2006:TSB

[Usa06] Yoshiyuki Usami. Theoretical study on the body form and swimming pattern of anomalocaris based on hydrodynamic simulation. Journal of Theoretical Biology, 238(1):11–17, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002109.

Urzay:2009:WGP

[USTG09] Javier Urzay, Stefan G. Llewellyn Smith, Elinor Thompson, and Beverley J. Glover. Wind gusts and plant aeroelasticity effects on the aerodynamics of pollen shedding: a hypothetical turbulence-initiated wind-pollination mechanism. Journal of Theoretical Biology, 259(4):785–792, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930900201X.

Ullah:2009:ITM

[UW09] Mukhtar Ullah and Olaf Wolkenhauer. Investigating the two-moment characterisation of subcellular biochemical networks. Journal of Theoretical Biology, 260(3):340–352, October 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002471.

Vinga:2004:RCE

[VA04a] Susana Vinga and Jonas S. Almeida. Rényi continuous entropy of DNA sequences. Journal of Theoretical Biology, 231(3):377–388, December 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003224.

Vinogradov:2004:PRQ

[VA04b] Alexander E. Vinogradov and Olga V. Anatskaya. Phenological resonance and quantum life history. Journal of Theoretical Biology, 228(3):417–420, June 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000529.

Veytsman:2006:SMM

[VA06] Boris Veytsman and Leila Akhmadeyeva. Simple mathematical model of pathologic microsatellite expansions: When self-reparation does not work. Journal of Theoretical Biology, 242(2):401–408, September 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001147.

VanWassenbergh:2005:DMM

[VAAH05] Sam Van Wassenbergh, Peter Aerts, Dominique Adriaens, and Anthony Herrel. A dynamic model of mouth closing movements in clariid catfishes: the role of enlarged jaw adductors. Journal of Theoretical Biology, 234(1):49–65, May 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005284.

vanAlbada:2009:MFMb

[vAGDR09] S. J. van Albada, R. T. Gray, P. M. Drysdale, and P. A. Robinson. Mean-field modeling of the basal ganglia-thalamocortical system. II: Dynamics of Parkinsonian oscillations. Journal of Theoretical Biology, 257(4):664–688, April 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006474.

VanDongen:2006:PSB

[Van06a] Stefan Van Dongen. Prior specification in Bayesian statistics: Three cautionary tales. Journal of Theoretical Biology, 242(1):90–100, September 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306000609.

Vandermeer:2006:OSF

[Van06b] John Vandermeer. Omnivory and the stability of food webs. Journal of Theoretical Biology, 238(3):497–504, February 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305002407.

vanAlbada:2009:MFMa

[vAR09] S. J. van Albada and P. A. Robinson. Mean-field modeling of the basal ganglia-thalamocortical system. I: Firing rates in healthy and Parkinsonian states. Journal of Theoretical Biology, 257(4):642–663, April 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006486.

Vazquez:2007:EOS

[Vaz07] Alexei Vazquez. Epidemic outbreaks on structured populations. Journal of Theoretical Biology, 245(1):125–129, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004279.

Venner:2006:DOI

[VCBV^{+}06] Samuel Venner, Iadine Chadès, Marie-Claude Bel-Venner, Alain Pasquet, François Charpillet, and Raymond Leborgne. Dynamic optimization over infinite-time horizon: Webbuilding strategy in an orb-weaving spider as a case study. Journal of Theoretical Biology, 241(4):725–733, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306000105.

Vautrin:2007:EID

[VCGV07] E. Vautrin, S. Charles, S. Genieys, and F. Vavre. Evolution and invasion dynamics of multiple infections with wolbachia investigated using matrix based models. Journal of Theoretical Biology, 245(2):197–209, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930600470X.

vandenBerg:2003:APM

[vdBR03] Hugo A. van den Berg and David A. Rand. Antigen presentation on MHC molecules as a diversity filter that enhances immune efficacy. Journal of Theoretical Biology, 224(2):249–267, September 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303001620.

vandenBerg:2004:DCA

[vdBR04a] Hugo A. van den Berg and David A. Rand. Dynamics of T cell activation threshold tuning. Journal of Theoretical Biology, 228(3):397–416, June 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000517.

vandenBerg:2004:FMD

[vdBR04b] Hugo A. van den Berg and David A. Rand. Foreignness as a matter of degree: the relative immunogenicity of peptide/ MHC ligands. Journal of Theoretical Biology, 231(4):535–548, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304003297.

vandenBerg:2009:PGP

[vdBvdB09] Femke van den Berg and Frank van den Bosch. The presence of generalist plant pathogens might not explain the long-term coexistence of plant species. Journal of Theoretical Biology, 257(3):446–453, April 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006176.

vandenBerg:2007:CCD

[vdBWLS07] Hugo A. van den Berg, Linda Wooldridge, Bruno Laugel, and Andrew K. Sewell. Coreceptor CD8-driven modulation of T cell antigen receptor specificity. Journal of Theoretical Biology, 249(2):395–408, November 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307003621.

vanderGulik:2009:FPE

[vdGMG^{+}09] Peter van der Gulik, Serge Massar, Dimitri Gilis, Harry Buhrman, and Marianne Rooman. The first peptides: The evolutionary transition between prebiotic amino acids and early proteins. Journal of Theoretical Biology, 261(4):531–539, December 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900424X.

Verdasca:2005:RES

[VdGN^{+}05] J. Verdasca, M. M. Telo da Gama, A. Nunes, N. R. Bernardino, J. M. Pacheco, and M. C. Gomes. Recurrent epidemics in small world networks. Journal of Theoretical Biology, 233(4):553– 561, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005193.

VanDien:2003:DES

[VdL03] Stephen J. Van Dien and Victor de Lorenzo. Deciphering environmental signal integration in σ54-dependent promoters with a simple mathematical model. Journal of Theoretical Biology, 224(4):437–449, October 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001917.

Vereecken:2000:PMM

[VDV00] Karen M. Vereecken, Els J. Dens, and Jan F. Van Impe. Predictive modeling of mixed microbial populations in food products: Evaluation of two-species models. Journal of Theoretical Biology, 205(1):53–72, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300920460.

Veitia:2003:NEM

[Vei03] Reiner A. Veitia. Nonlinear effects in macromolecular assembly and dosage sensitivity. Journal of Theoretical Biology, 220(1):19–25, January 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931055.

Verhagen:2004:HBS

[Ver04] Jan H. G. Verhagen. Hydrodynamics of burst swimming fish larvae; a conceptual model approach. Journal of Theoretical Biology, 229(2):235–248, July 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001377.

Vahedi:2009:SRD

[VFC^{+}09] Golnaz Vahedi, Babak Faryabi, Jean-Francois Chamberland, Aniruddha Datta, and Edward R. Dougherty. Sampling-rate-dependent probabilistic Boolean networks. Journal of Theoretical Biology, 261(4):540–547, December 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003889.

Varfolomeyev:2001:HGH

[VG01] S. D. Varfolomeyev and K. G. Gurevich. The hyperexponential growth of the human population on a macrohistorical scale. Journal of Theoretical Biology, 212(3):367–372, October 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923847.

Vilas:2006:SIP

[VGBA06] Carlos Vilas, Míriam R. García, Julio R. Banga, and Antonio A. Alonso. Stabilization of inhomogeneous patterns in a diffusion-reaction system under structural and parametric uncertainties. Journal of Theoretical Biology, 241(2):295–306, July 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305005175.

Varon:2002:KAG

[VGCGM^{+}02] R. Varón, F. García-Cánovas, M. García-Moreno, E. Valero, M. Molina-Alarcón, M. J. García-Meseguer, J. A. Vidal De Labra, and C. Garrido-Del Solo. Kinetic analysis of the general modifier mechanism of botts and Morales involving a suicide substrate. Journal of Theoretical Biology, 218(3):355–374, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930833.

Vilar:2009:NQM

[VGDSU09] Santiago Vilar, Humberto González-Díaz, Lourdes Santana, and Eugenio Uriarte. A network-QSAR model for prediction of genetic-component biomarkers in human colorectal cancer. Journal of Theoretical Biology, 261(3):449–458, December 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309003452.

Valero:2007:CKB

[VGMM^{+}07a] Edelmira Valero, Manuela García-Moreno, Jesualdo Masiá, María-José García-Meseguer, and Ramón Varón. Corrigendum to “Kinetic behaviour of proenzymes activation in the presence of different inhibitors for both activating and activated enzymes”: [J. Theor. Biol. 245 (2007) 175–192]. Journal of Theoretical Biology, 247(1):212, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519307001440. See [VGMM^{+}07b, VGMM^{+}08].

Valero:2007:KBP

[VGMM^{+}07b] Edelmira Valero, Manuela García-Moreno, Jesualdo Masiá, María-José García-Meseguer, and Ramón Varón. Kinetic behaviour of proenzymes activation in the presence of different inhibitors for both activating and activated enzymes. Journal of Theoretical Biology, 245(1):175–192, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306004309. See corrigenda

[VGMM^{+}07a, VGMM^{+}08].

Valero:2008:CKB

[VGMM^{+}08] Edelmira Valero, Manuela García-Moreno, Jesualdo Masiá, María-José García-Meseguer, and Ramón Varón. Corrigendum to “Kinetic behaviour of proenzymes activation in the presence of different inhibitors for both activating and activated enzymes”: [J. Theor. Biol. 245 (2007) 175–192]. Journal of Theoretical Biology, 250(3):580, February 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519307005279. See [VGMM^{+}07b, VGMM^{+}07a].

Varon:2006:KAG

[VGMVR^{+}06] R. Varón, M. García-Moreno, D. Valera-Ruipérez, F. García-Molina, F. García-Cánovas, R. G. Ladrón de Guevara, J. Masiá-Pérez, and B. H. Havsteen. Kinetic analysis of a general model of activation of aspartic proteinase zymogens. Journal of Theoretical Biology, 242(3):743–754, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001640.

Vainstein:2005:CEG

[VGS^{+}05] V. Vainstein, Y. Ginosar, M. Shoham, D. O. Ranmar, A. Ianovski, and Z. Agur. The complex effect of granulocyte colony-stimulating factor on human granulopoiesis analyzed by a new physiologically-based mathematical model. Journal of Theoretical Biology, 234(3):311–327, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005806.

Venkatasubramanian:2006:IEM

[VHF06] Raja Venkatasubramanian, Michael A. Henson, and Neil S. Forbes. Incorporating energy metabolism into a growth model of multicellular tumor spheroids. Journal of Theoretical Biology, 242(2):440–453, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001184.

Venkatasubramanian:2008:ICC

[VHF08] Raja Venkatasubramanian, Michael A. Henson, and Neil S. Forbes. Integrating cell-cycle progression, drug penetration and energy metabolism to identify improved cancer therapeutic strategies. Journal of Theoretical Biology, 253(1):98–117, July 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308000799.

Vincent:2005:DDB

[Vin05] Julian F. V. Vincent. Deconstructing the design of a biological material. Journal of Theoretical Biology, 236(1):73–78, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000779.

vanKooten:2005:BAE

[vKdRP05] Tobias van Kooten, André M. de Roos, and Lennart Persson. Bistability and an Allee effect as emergent consequences of stage-specific predation. Journal of Theoretical Biology, 237(1):67–74, November 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001530.

vanKooten:2007:PDC

[vKPdR07] Tobias van Kooten, Lennart Persson, and André M. de Roos. Population dynamical consequences of gregariousness in a size-structured consumer-resource interaction. Journal of Theoretical Biology, 245(4):763–774, April 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306005674.

Vilela:2009:IPW

[VL09] Rafael D. Vilela and Benjamin Lindner. Are the input parameters of white noise driven integrate and fire neurons uniquely determined by rate and CV? Journal of Theoretical Biology, 257(1):90–99, March 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005833.

vanLeeuwen:2007:EIB

[vLBJK07] Ingeborg M. M. van Leeuwen, Helen M. Byrne, Oliver E. Jensen, and John R. King. Elucidating the interactions between the adhesive and transcriptional functions of βcatenin in normal and cancerous cells. Journal of Theoretical Biology, 247(1):77–102, July 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000495.

Varanda:2000:HAA

[VLFN00] Wamberto A. Varanda, Larry S. Liebovitch, José N. Figueiroa, and Romildo A. Nogueira. Hurst analysis applied to the study of single calcium-activated potassium channel kinetics. Journal of Theoretical Biology, 206(3):343–353, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921313.

vonLaer:2006:IGM

[vLHH06] Dorothee von Laer, Susanne Hasselmann, and Klaus Hasselmann. Impact of gene-modified T cells on HIV infection dynamics. Journal of Theoretical Biology, 238(1):60–77, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305002079.

Viscido:2001:RSH

[VMW01] Steven V. Viscido, Matthew Miller, and David S. Wethey. The response of a selfish herd to an attack from outside the group perimeter. Journal of Theoretical Biology, 208(3):315–328, February 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922215.

Viscido:2002:DSH

[VMW02] Steven V. Viscido, Matthew Miller, and David S. Wethey. The dilemma of the selfish herd: The search for a realistic movement rule. Journal of Theoretical Biology, 217(2):183–194, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930250.

Vance:2007:PPG

[VN07] Richard R. Vance and Andrew L. Nevai. Plant population growth and competition in a light gradient: a mathematical model of canopy partitioning. Journal of Theoretical Biology, 245(2):210–219, March 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004917.

Voelkl:2008:ISS

[VN08] Bernhard Voelkl and Ronald Noë. The influence of social structure on the propagation of social information in artificial primate groups: a graph-based simulation approach. Journal of Theoretical Biology, 252(1):77–86, May 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000532.

Veitia:2000:PPS

[VO00] Reiner Veitia and Chris Ottolenghi. Placing parallel stranded DNA in an evolutionary context. Journal of Theoretical Biology, 206(2):317–322, September 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921192.

Voit:2003:BGR

[Voi03] Eberhard O. Voit. Biochemical and genomic regulation of the trehalose cycle in yeast: review of observations and canonical model analysis. Journal of Theoretical Biology, 223(1):55– 78, July 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303000729.

VonWangenheim:2001:MIT

[VP01] Karl-Hartmut Von Wangenheim and Hans-Peter Peterson. A mechanism of intracellular timing and its cooperation with extracellular signals in controlling cell proliferation and differentiation, an amended hypothesis. Journal of Theoretical Biology, 211(3):239–251, August 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923422.

Virkki:2007:CCB

[VPGA07] Arho Virkki, Olli Polo, Mats Gyllenberg, and Tero Aittokallio. Can carotid body perfusion act as a respiratory controller? Journal of Theoretical Biology, 249(4):737–748, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307004274.

VanderLinden:2006:PPT

[VRAF06] M. G. Van der Linden, T. G. Rêgo, D. A. M. Araújo, and S. T. Farias. Prediction of potential thermostable proteins in Xylella fastidiosa. Journal of Theoretical Biology, 242(2):421–425, September 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001160.

VanDenBerg:2001:RSC

[VRB01] H. A. Van Den Berg, D. A. Rand, and N. J. Burroughs. A reliable and safe T cell repertoire based on low-affinity T cell receptors. Journal of Theoretical Biology, 209(4):465–486, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922817.

Veluraja:2008:DSL

[VS08] K. Veluraja and Arunachalam N. Seethalakshmi. Dynamics of sialyl Lewis a in aqueous solution and prediction of the structure of the sialyl Lewis a -SelectinE complex. Journal of Theoretical Biology, 252(1):15–23, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000039.

Vainstein:2007:DMD

[VSA07] Mendeli H. Vainstein, Ana T. C. Silva, and Jeferson J. Arenzon. Does mobility decrease cooperation? Journal of Theoretical Biology, 244(4):722–728, February 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004218.

Valandro:2000:IQP

[VSCG00] L. Valandro, B. Salvato, R. Caimmi, and L. Galzigna. Isomorphism of quasispecies and percolation models. Journal of Theoretical Biology, 202(3):187–194, February 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910366.

Voronin:2006:PMP

[VSP06] Yegor A. Voronin, Igor A. Sidorov, and Vinay K. Pathak. A probability model predicting initiation efficiency of retroviral vectors with two primer-binding sites. Journal of Theoretical Biology, 242(2):347–355, September 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001081.

Vieira:2008:GCMa

[VTC08a] Ricardo Augusto Mendonça Vieira, Luis Orlindo Tedeschi, and Antonello Cannas. A generalized compartmental model to estimate the fibre mass in the ruminoreticulum: 1. Estimating parameters of digestion. Journal of Theoretical Biology, 255(4):345–356, December 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004256.

Vieira:2008:GCMb

[VTC08b] Ricardo Augusto Mendonça Vieira, Luis Orlindo Tedeschi, and Antonello Cannas. A generalized compartmental model to estimate the fibre mass in the ruminoreticulum: 2. Integrating digestion and passage. Journal of Theoretical Biology, 255(4):357–368, December 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930800427X.

Vittori:2006:PEA

[VTG^{+}06] Karla Vittori, Grégoire Talbot, Jacques Gautrais, Vincent Fourcassié, Aluizio F. R. Araújo, and Guy Theraulaz. Path efficiency of ant foraging trails in an artificial network. Journal of Theoretical Biology, 239(4):507–515, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003577.

VanDongen:2005:HVD

[VTL05] Stefan Van Dongen, Willem Talloen, and Luc Lens. High variation in developmental instability under non-normal developmental error: a Bayesian perspective. Journal of Theoretical Biology, 236(3):263–275, October 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001190.

vanVeelen:2005:UPE

[vV05] Matthijs van Veelen. On the use of the price equation. Journal of Theoretical Biology, 237(4):412–426, December 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305001906.

vanVeelen:2006:WKG

[vV06] Matthijs van Veelen. Why kin and group selection models may not be enough to explain human other-regarding behaviour. Journal of Theoretical Biology, 242(3):790–797, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001743.

vanVeelen:2007:HML

[vV07] Matthijs van Veelen. Hamilton’s missing link. Journal of Theoretical Biology, 246(3):551–554, June 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000057.

vanVeelen:2009:GSK

[vV09] Matthijs van Veelen. Group selection, kin selection, altruism and cooperation: When inclusive fitness is right and when it can be wrong. Journal of Theoretical Biology, 259(3):589–600, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001891.

vanVeelen:2007:LWA

[vVH07] Matthijs van Veelen and Astrid Hopfensitz. In love and war: Altruism, norm formation, and two different types of group selection. Journal of Theoretical Biology, 249(4):667–680, December 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004195.

VanOoyen:2000:DNC

[VW00] Arjen Van Ooyen and David J. Willshaw. Development of nerve connections under the control of neurotrophic factors: Parallels with consumer-resource systems in population biology. Journal of Theoretical Biology, 206(2):195–210, September 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300921143.

Vasiev:2003:MDD

[VW03] Bakhtier Vasiev and Cornelis J. Weijer. Modelling of Dictyostelium discoideum slug migration. Journal of Theoretical Biology, 223(3):347–359, August 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001036.

Vasco:2007:TDP

[VWR07] Daniel A. Vasco, Helen J. Wearing, and Pejman Rohani. Tracking the dynamics of pathogen interactions: Modeling ecological and immune-mediated processes in a two-pathogen single-host system. Journal of Theoretical Biology, 245(1):9– 25, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306003638.

Wiegmann:2007:MCU

[WA07] Daniel D. Wiegmann and Lisa M. Angeloni. Mate choice and uncertainty in the decision process. Journal of Theoretical Biology, 249(4):654–666, December 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004055.

Wagner:2003:RMB

[Wag03] Andreas Wagner. Risk management in biological evolution. Journal of Theoretical Biology, 225(1):45–57, November 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002194.

Wahl:2002:EDL

[Wah02] L. M. Wahl. Evolving the division of labour: Generalists, specialists and task allocation. Journal of Theoretical Biology, 219(3):371–388, December 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931334.

Wakano:2004:DGE

[Wak04] Joe Yuichiro Wakano. Drastic growth effect may explain sympatric cannibalistic polymorphism. Journal of Theoretical Biology, 226(1):69–77, January 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003217.

Wakano:2005:EEF

[Wak05] Joe Yuichiro Wakano. Evolution of extraordinary femalebiased sex ratios: The optimal schedule of sex ratio in local mate competition. Journal of Theoretical Biology, 237(2):193–202, November 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930500161X.

Wakano:2007:ECS

[Wak07] Joe Yuichiro Wakano. Evolution of cooperation in spatial public goods games with common resource dynamics. Journal of Theoretical Biology, 247(4):616–622, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001907.

Wallraff:2000:SNB

[Wal00] Hans G. Wallraff. Simulated navigation based on observed gradients of atmospheric trace gases (models on pigeon homing, part 3). Journal of Theoretical Biology, 205(1):133–145, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920526.

Wood:2006:MAG

[WAL06] A. J. Wood, G. J. Ackland, and T. M. Lenton. Mutation of albedo and growth response produces oscillations in a spatial daisyworld. Journal of Theoretical Biology, 242(1):188–198, September 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000786.

Wallace:2007:CIB

[Wal07a] Rodrick Wallace. Culture and inattentional blindness: a global workspace perspective. Journal of Theoretical Biology, 245(2):378–390, March 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004735.

Walters:2007:MPD

[Wal07b] K. Walters. Modelling the probability distribution of the number of DNA double-strand breaks due to sporadic alkylation of nucleotide bases. Journal of Theoretical Biology, 245(1):161– 168, March 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004528.

Walker:2008:MEM

[Wal08] Michael M. Walker. A model for encoding of magnetic field intensity by magnetite-based magnetoreceptor cells. Journal of Theoretical Biology, 250(1):85–91, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004596.

Wang:2000:MPR

[WAMO00] Jin-Sheng Wang, Tetsuya Araki, Masayoshi Matsuoka, and Takahira Ogawa. A model of photoinhibition related to mRNA instability in ethylene production by a recombinant cyanobacterium. Journal of Theoretical Biology, 202(3):205–211, February 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519399910536.

Wang:2000:SDD

[Wan00] James H.-C. Wang. Substrate deformation determines actin cytoskeleton reorganization: a mathematical modeling and experimental study. Journal of Theoretical Biology, 202(1):33–41, January 7, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519399910354.

Wang:2006:MPP

[Wan06] Hongyun Wang. Motor potential profile and a robust method for extracting it from time series of motor positions. Journal of Theoretical Biology, 242(4):908–921, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001597.

Waxman:2006:FGM

[Wax06] D. Waxman. Fisher’s geometrical model of evolutionary adaptation-beyond spherical geometry. Journal of Theoretical Biology, 241(4):887–895, August 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000361.

Waxman:2009:FLM

[Wax09] D. Waxman. Fixation at a locus with multiple alleles: Structure and solution of the Wright Fisher model. Journal of Theoretical Biology, 257(2):245–251, March 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006139.

Walters:2002:GFG

[WB02] K. Walters and P. G. Blackwell. Group formation games with reappraisal. Journal of Theoretical Biology, 217(2):149–157, July 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930286.

Wierschem:2004:CBP

[WB04] Keola Wierschem and Richard Bertram. Complex bursting in pancreatic islets: a potential glycolytic mechanism. Journal of Theoretical Biology, 228(4):513–521, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000724.

Woods:2006:IET

[WB06] H. Arthur Woods and Roger T. Bonnecaze. Insect eggs at a transition between diffusion and reaction limitation: Temperature, oxygen, and water. Journal of Theoretical Biology, 243(4):483–492, December 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003006.

Waltman:2002:MMB

[WBB02] Paul Waltman, James Braselton, and Lorraine Braselton. A mathematical model of a biological arms race with a dangerous prey. Journal of Theoretical Biology, 218(1):55–70, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930572.

Whittaker:2009:MMF

[WBD^{+}09] Robert J. Whittaker, Richard Booth, Rosemary Dyson, Clare Bailey, Louise Parsons Chini, Shailesh Naire, Sevil Payvandi, Zimei Rong, Hannah Woollard, Linda J. Cummings, Sarah L. Waters, Lina Mawasse, Julian B. Chaudhuri, Marianne J. Ellis, Vipin Michael, Nicola J. Kuiper, and Sarah Cartmell. Mathematical modelling of fibre-enhanced perfusion inside a tissue-engineering bioreactor. Journal of Theoretical Biology, 256(4):533–546, February 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005353.

Weld:2004:MPG

[WBH04] Richard J. Weld, Christine Butts, and Jack A. Heinemann. Models of phage growth and their applicability to phage therapy. Journal of Theoretical Biology, 227(1):1–11, March 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303002625.

Welter:2008:EVN

[WBR08] M. Welter, K. Bartha, and H. Rieger. Emergent vascular network inhomogeneities and resulting blood flow patterns in a growing tumor. Journal of Theoretical Biology, 250(2):257–280, January 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004584.

Welter:2009:VRA

[WBR09] M. Welter, K. Bartha, and H. Rieger. Vascular remodelling of an arterio-venous blood vessel network during solid tumour growth. Journal of Theoretical Biology, 259(3):405–422, August 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001635.

Wang:2006:ESA

[WBSY06] Yue-Lan Wang, Ju Bao, Yan Sun, and Jie Yang. Energy and structural analysis of double nucleic acid triplets. Journal of Theoretical Biology, 238(1):85–103, January 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002201.

West:2002:GMH

[WC02] Stuart A. West and J. Paul Cunningham. A general model for host plant selection in phytophagous insects. Journal of Theoretical Biology, 214(3):499–513, February 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924750.

Wan:2007:SEM

[WC07a] Hui Wan and Jing’an Cui. An SEIS epidemic model with transport-related infection. Journal of Theoretical Biology, 247(3):507–524, August 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307001555.

Wood:2007:FBA

[WC07b] A. J. Wood and J. B. Coe. A fitness based analysis of daisyworld. Journal of Theoretical Biology, 249(2):190–197, November 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307003645.

Wang:2006:CGO

[WCA06] Ruiqi Wang, Luonan Chen, and Kazuyuki Aihara. Construction of genetic oscillators with interlocked feedback networks. Journal of Theoretical Biology, 242(2):454–463, September 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306001202.

Wu:2009:ASR

[WCH^{+}09] Sheng-Nan Wu, Bing-Shuo Chen, Tai-I Hsu, Hsung Peng, Yung-Han Wu, and Yi-Ching Lo. Analytical studies of rapidly inactivating and noninactivating sodium currents in differentiated NG108-15 neuronal cells. Journal of Theoretical Biology, 259(4):828–836, August 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002185.

Wu:2008:CSI

[WCLL08] Sheng-Nan Wu, Bing-Shuo Chen, Ming-Wei Lin, and Yen-Chin Liu. Contribution of slowly inactivating potassium current to delayed firing of action potentials in NG108-15 neuronal cells: Experimental and theoretical studies. Journal of Theoretical Biology, 252(4):711–721, June 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000520.

Williams:2003:EWC

[WD03] R. J. P. Williams and J. J. R. Fraústo Da Silva. Evolution was chemically constrained. Journal of Theoretical Biology, 220(3):323–343, February 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931523.

Weindl:2009:MPS

[WDH^{+}09] Johanna Weindl, Zaher Dawy, Pavol Hanus, Juergen Zech, and Jakob C. Mueller. Modeling promoter search by E. coli RNA polymerase: One-dimensional diffusion in a sequence-dependent energy landscape. Journal of Theoretical Biology, 259(3):628–634, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900232X.

Wackermann:2006:DKM

[WE06] Jirí Wackermann and Werner Ehm. The dual klepsydra model of internal time representation and time reproduction. Journal of Theoretical Biology, 239(4):482–493, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305003590.

Wells:2000:ETH

[Wel00a] J. C. K. Wells. Environmental temperature and human growth in early life. Journal of Theoretical Biology, 204(2):299–305, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920150.

Wells:2000:NSS

[Wel00b] Jonathan C. K. Wells. Natural selection and sex differences in morbidity and mortality in early life. Journal of Theoretical Biology, 202(1):65–76, January 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519399910445.

Wells:2002:TEH

[Wel02] Jonathan C. K. Wells. Thermal environment and human birth weight. Journal of Theoretical Biology, 214(3):413–425, February 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924658.

Wells:2003:TPH

[Wel03] Jonathan C. K. Wells. The thrifty phenotype hypothesis: Thrifty offspring or thrifty mother? Journal of Theoretical Biology, 221(1):143–161, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931833.

Westneat:2003:BMA

[Wes03] Mark W. Westneat. A biomechanical model for analysis of muscle force, power output and lower jaw motion in fishes. Journal of Theoretical Biology, 223(3):269–281, August 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000584.

Westerhoff:2008:SCS

[Wes08] Hans V. Westerhoff. Signalling control strength. Journal of Theoretical Biology, 252(3):555–567, June 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005772.

Wiback:2004:MCS

[WFGP04] Sharon J. Wiback, Iman Famili, Harvey J. Greenberg, and Bernhard Ø. Palsson. Monte Carlo sampling can be used to determine the size and shape of the steady-state flux space. Journal of Theoretical Biology, 228(4):437–447, June 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304000554.

Wlodek:2003:DEL

[WG03] Danuta Wlodek and Michael Gonzales. Decreased energy levels can cause and sustain obesity. Journal of Theoretical Biology, 225(1):33–44, November 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002182.

White:2006:RII

[WG06] K. A. Jane White and C. A. Gilligan. The role of initial inoculum on epidemic dynamics. Journal of Theoretical Biology, 242(3):670–682, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001585.

Whiteley:2000:EIO

[WGH00] J. P. Whiteley, D. J. Gavaghan, and C. E. W. Hahn. The effect of inspired oxygen concentration on the ventilation– perfusion distribution in inhomogeneous lungs. Journal of Theoretical Biology, 204(4):575–585, June 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920368.

Whiteley:2001:MIG

[WGH01] J. P. Whiteley, D. J. Gavaghan, and C. E. W. Hahn. Modelling inert gas exchange in tissue and mixed-venous blood return to the lungs. Journal of Theoretical Biology, 209(4):431–443, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301922787.

Weber:2000:OSD

[WH00] Thomas P. Weber and Anders Hedenström. Optimal stopover decisions under wind influence: the effects of correlated winds. Journal of Theoretical Biology, 205(1):95–104, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920472.

Welton:2001:TID

[WH01] Nicky J. Welton and Alasdair I. Houston. A theoretical investigation into the direct and indirect effects of state on the risk of predation. Journal of Theoretical Biology, 213(2):275–297, November 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301924191.

Wilkins:2003:IMC

[WH03] Jon F. Wilkins and David Haig. Inbreeding, maternal care and genomic imprinting. Journal of Theoretical Biology, 221(4):559–564, April 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303932061.

Wolkenhauer:2007:ACM

[WH07] Olaf Wolkenhauer and Jan-Hendrik S. Hofmeyr. An abstract cell model that describes the self-organization of cell function in living systems. Journal of Theoretical Biology, 246(3):461– 476, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307000203.

Wang:2007:QKM

[WHH07] Jinhui Wang, Xudong Huang, and Weidong Huang. A quantitative kinetic model for ATP-induced intracellular Ca^{2+} oscillations. Journal of Theoretical Biology, 245(3):510–519, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306005492.

White:2005:AML

[Whi05] D. A. White. Architectural mutation and leaf form, for the palmate series. Journal of Theoretical Biology, 235(2):289–301, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000263.

Whitehead:2007:LCE

[Whi07] Hal Whitehead. Learning, climate and the evolution of cultural capacity. Journal of Theoretical Biology, 245(2):341–350, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004681.

Wilkinson:2001:PPD

[Wil01] Michael H. F. Wilkinson. Predation in the presence of decoys: An inhibitory factor on pathogen control by bacteriophages or bdellovibrios in dense and diverse ecosystems. Journal of Theoretical Biology, 208(1):27–36, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921970.

Williams:2002:PPT

[Wil02] R. J. P. Williams. The problem of proton transfer in membranes. Journal of Theoretical Biology, 219(3):389–396, December 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302931346.

Wilkins:2006:CSD

[Wil06a] Jon F. Wilkins. Competitive signal discrimination, methylation reprogramming and genomic imprinting. Journal of Theoretical Biology, 242(3):643–651, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001767.

Wilkins:2006:TSR

[Wil06b] Jon F. Wilkins. Tissue-specific reactivation of gene expression at an imprinted locus. Journal of Theoretical Biology, 240(2):277–287, May 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004108.

Williams:2008:UHS

[Wil08a] Paul D. Williams. Unhealthy herds: Some epidemiological consequences of host heterogeneity in predator-host-parasite systems. Journal of Theoretical Biology, 253(3):500–507, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930800146X.

Willson:2008:RCP

[Wil08b] Stephen J. Willson. Reconstruction of certain phylogenetic networks from the genomes at their leaves. Journal of Theoretical Biology, 252(2):338–349, May 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000787.

Wills:2009:IGP

[Wil09] Peter R. Wills. Informed generation: Physical origin and biological evolution of genetic codescript interpreters. Journal of Theoretical Biology, 257(3):345–358, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006693.

Winfree:2004:AWA

[Win04] Rachael Winfree. Art Winfree, artist of science. Journal of Theoretical Biology, 230(4):441–443, October 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304002358.

Winther:2006:FWP

[Win06] Rasmus Grønfeldt Winther. Fisherian and Wrightian perspectives in evolutionary genetics and model-mediated imposition of theoretical assumptions. Journal of Theoretical Biology, 240(2):218–232, May 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004133.

Witting:2003:MLH

[Wit03] Lars Witting. Major life-history transitions by deterministic directional natural selection. Journal of Theoretical Biology, 225(3):389–406, December 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303002741.

Wodarz:2001:RCH

[WJ01] Dominik Wodarz and Vincent A. A. Jansen. The role of T cell help for anti-viral CTL responses. Journal of Theoretical Biology, 211(4):419–432, August 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923586.

Weiss:2000:ICP

[WJMH00] Olaf Weiss, Miguel A. Jiménez-Montaño, and Hanspeter Herzel. Information content of protein sequences. Journal of Theoretical Biology, 206(3):379–386, October 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921386.

Wei:2005:SIM

[WK05] Wei Wei and Stephen M. Krone. Spatial invasion by a mutant pathogen. Journal of Theoretical Biology, 236(3):335–348, October 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001256.

Webb:2007:HPI

[WKB07] Steven D. Webb, Matt J. Keeling, and Mike Boots. Hostparasite interactions between the local and the mean-field: How and when does spatial population structure matter? Journal of Theoretical Biology, 249(1):140–152, November 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002925.

Wang:2003:KAM

[WKG03] Steven S.-S. Wang, Vasiliki Kazantzi, and Theresa A. Good. A kinetic analysis of the mechanism of β-amyloid induced G protein activation. Journal of Theoretical Biology, 221(2):269–278, March 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303931894.

Woolf:2001:UBH

[WKL01] Peter J. Woolf, Terry P. Kenakin, and Jennifer J. Linderman. Uncovering biases in high throughput screens of G-protein coupled receptors. Journal of Theoretical Biology, 208(4):403–418, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922276.

Wu:2009:FPB

[WKRD09] H. Wu, J. W. Kim, P. A. Robinson, and P. M. Drysdale. Firing pattern of bursting neurons under sinusoidal drive in meanfield modeling. Journal of Theoretical Biology, 259(1):101– 108, July 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001313.

Willett:2002:SCO

[WL02] L. J. Willett and S. K. Loyalka. Sticking coefficientorchestrated selection in a kinetic model for the first origin. Journal of Theoretical Biology, 218(1):13–33, September 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930602.

Wei:2003:LSN

[WL03] Chunfang Wei and Philip M. Lintilhac. Loss of stability — a new model for stress relaxation in plant cell walls. Journal of Theoretical Biology, 224(3):305–312, October 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930300167X.

Woolf:2004:ADI

[WL04] Peter J. Woolf and Jennifer J. Linderman. An algebra of dimerization and its implications for G-protein coupled receptor signaling. Journal of Theoretical Biology, 229(2):157–168, July 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001134.

Wang:2007:CVI

[WL07a] Zhiping Wang and Xianning Liu. A chronic viral infection model with immune impairment. Journal of Theoretical Biology, 249(3):532–542, December 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004018.

Worden:2007:EEP

[WL07b] Lee Worden and Simon A. Levin. Evolutionary escape from the prisoner’s dilemma. Journal of Theoretical Biology, 245(3):411–422, April 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004863.

Wang:2009:IMT

[WL09] Hsiuying Wang and Wen-Hsiung Li. Increasing MicroRNA target prediction confidence by the relative R^{2} method. Journal of Theoretical Biology, 259(4):793–798, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900229X.

Wise:2008:TDM

[WLFC08] S. M. Wise, J. S. Lowengrub, H. B. Frieboes, and V. Cristini. Three-dimensional multispecies nonlinear tumor growth — I: Model and numerical method. Journal of Theoretical Biology, 253(3):524–543, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001525.

Wu:2007:MPD

[WLHB07] Hsin-Hue Wu, How-Jing Lee, Shwu-Bin Horng, and Ludek Berec. Modeling population dynamics of two cockroach species: Effects of the circadian clock, interspecific competition and pest control. Journal of Theoretical Biology, 249(3):473–486, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307003980.

Wang:2006:RAG

[WLZ^{+}06] Yuhong Wang, Wei Li, Tongtong Zhang, Chuanfang Ding, Zhibin Lu, Nanye Long, John P. Rose, Bi-Cheng Wang, and Dawei Lin. Reconstruction of ancient genome and gene order from complete microbial genome sequences. Journal of Theoretical Biology, 239(4):494–498, April 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930500353X.

Wagner:2000:MEG

[WM00] Günter P. Wagner and Jason Mezey. Modeling the evolution of genetic architecture: a continuum of alleles model with pairwise A×A epistasis. Journal of Theoretical Biology, 203(2):163–175, March 21, 2000. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300910749.

Wilson:2004:MNC

[WM04] D. P. Wilson and D. L. S. McElwain. A model of neutralization of chlamydia trachomatis based on antibody and host cell aggregation on the elementary body surface. Journal of Theoretical Biology, 226(3):321–330, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003527.

Wu:2002:SMG

[WMLC02] Rongling Wu, Chang-Xing Ma, Ramon C. Littell, and George Casella. A statistical model for the genetic origin of allometric scaling laws in biology. Journal of Theoretical Biology, 219(1):121–135, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931140.

Wiback:2003:RMF

[WMP03] Sharon J. Wiback, Radhakrishnan Mahadevan, and Bernhard Ø. Palsson. Reconstructing metabolic flux vectors from extreme pathways: defining the α-spectrum. Journal of Theoretical Biology, 224(3):313–324, October 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303001681.

Wallhead:2008:SIP

[WMS08a] Philip J. Wallhead, Adrian P. Martin, and Meric A. Srokosz. Spatially implicit plankton population models: Transient spatial variability. Journal of Theoretical Biology, 253(3):405–423, August 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308001446.

Williams:2008:MLG

[WMS08b] George S. B. Williams, Evan J. Molinelli, and Gregory D. Smith. Modeling local and global intracellular calcium responses mediated by diffusely distributed inositol 1,4,5-trisphosphate receptors. Journal of Theoretical Biology, 253(1):170–188, July 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308001136.

Willensdorfer:2005:MEG

[WN05] Martin Willensdorfer and Martin A. Nowak. Mutation in evolutionary games can increase average fitness at equilibrium. Journal of Theoretical Biology, 237(4):355–362, December 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305001840.

Wang:2008:ICA

[WNT08] Wendi Wang, Shinji Nakaoka, and Yasuhiro Takeuchi. Invest conflicts of adult predators. Journal of Theoretical Biology, 253(1):12–23, July 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004468.

Webb:2004:IML

[WO04] Steven D. Webb and Markus R. Owen. Intra-membrane ligand diffusion and cell shape modulate juxtacrine patterning. Journal of Theoretical Biology, 230(1):99–117, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304001882.

Wodarz:2001:HDV

[Wod01] Dominik Wodarz. Helper-dependent vs. helper-independent CTL responses in HIV infection: Implications for drug therapy and resistance. Journal of Theoretical Biology, 213(3):447–459, December 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301924269.

Wodarz:2007:ESC

[Wod07] Dominik Wodarz. Effect of stem cell turnover rates on protection against cancer and aging. Journal of Theoretical Biology, 245(3):449–458, April 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306004899.

Wagner:2007:SME

[WOLS07] Günter P. Wagner, Wolfgang Otto, Vincent Lynch, and Peter F. Stadler. A stochastic model for the evolution of transcription factor binding site abundance. Journal of Theoretical Biology, 247(3):544–553, August 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000902.

Wong:2005:HDS

[Won05] Pierre Wong. A hypothesis of the disc-sphere transformation of the erythrocytes between glass surfaces and of related observations. Journal of Theoretical Biology, 233(1):127–135, March 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304004680.

Wong:2006:BHE

[Won06] Pierre Wong. The behavior of the human erythrocyte as an imperfect osmometer: a hypothesis. Journal of Theoretical Biology, 238(1):167–171, January 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002493.

Watts:2001:TAT

[WP01] Aaron B. Watts and Hasu Patel. The theory and application of transition state pK_{a} values: The reaction of papain with a series of trimethylene disulphide reactivity probes. Journal of Theoretical Biology, 209(4):417–429, April 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301922763.

Wang:2003:RNA

[WPE03] Hongyun Wang, Charles S. Peskin, and Timothy C. Elston. A robust numerical algorithm for studying biomolecular transport processes. Journal of Theoretical Biology, 221(4):491–511, April 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303932000.

Wanant:2000:IRB

[WQ00] Sumanas Wanant and Michael J. Quon. Insulin receptor binding kinetics: Modeling and simulation studies. Journal of Theoretical Biology, 205(3):355–364, August 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920691.

Weitz:2003:SDR

[WR03] Joshua S. Weitz and Daniel H. Rothman. Scale-dependence of resource-biodiversity relationships. Journal of Theoretical Biology, 225(2):205–214, November 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930300239X.

Wang:2004:SSO

[WR04] Wendi Wang and Shigui Ruan. Simulating the SARS outbreak in Beijing with limited data. Journal of Theoretical Biology, 227(3):369–379, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004338.

Wexler:2007:PDP

[WR07a] Ydo Wexler and Oleg Rokhlenko. Prisoner’s dilemma posed by fitness-associated recombination strategies. Journal of Theoretical Biology, 247(1):1–10, July 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000537.

Wu:2007:MIN

[WR07b] Huiying Wu and P. A. Robinson. Modeling and investigation of neural activity in the thalamus. Journal of Theoretical Biology, 244(1):1–14, January 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003080.

Wenzler:2001:TIT

[WRF01] D. Wenzler, M. Reinhardt, and L. Fukshansky. Topology of an intracellular transduction chain (phototropism of phycomyces ): 1. Joint review of functional, temporal, and spatial aspects. Journal of Theoretical Biology, 211(4):313–332, August 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923513.

Walther:2004:MMR

[WRG^{+}04] Th. Walther, H. Reinsch, A. Große, K. Ostermann, A. Deutsch, and Th. Bley. Mathematical modeling of regulatory mechanisms in yeast colony development. Journal of Theoretical Biology, 229(3):327–338, August 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304001432.

Wedge:2009:SMD

[WRKK09] David C. Wedge, William Rowe, Douglas B. Kell, and Joshua Knowles. In silico modelling of directed evolution: Implications for experimental design and stepwise evolution. Journal of Theoretical Biology, 257(1):131–141, March 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005845.

Wiktorsson:2004:MMS

[WRNB04] Magnus Wiktorsson, Tobias Rydén, Elna Nilsson, and Göran Bengtsson. Modelling the movement of a soil insect. Journal of Theoretical Biology, 231(4):497–513, December 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304003157.

Wagner:2002:CAM

[WS02a] Clemens Wagner and Jörg W. Stucki. Construction of an associative memory using unstable periodic orbits of a chaotic attractor. Journal of Theoretical Biology, 215(3):375–384, April 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301925032.

Wick:2002:WMH

[WS02b] David Wick and Steven G. Self. What’s the matter with HIV-directed killer T cells? Journal of Theoretical Biology, 219(1):19–31, November 7, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302931036.

Wagner:2003:QIH

[WS03a] Günter P. Wagner and Peter F. Stadler. Quasi-independence, homology and the unity of type: a topological theory of characters. Journal of Theoretical Biology, 220(4):505–527, February 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930393150X.

Wuchty:2003:CCN

[WS03b] Stefan Wuchty and Peter F. Stadler. Centers of complex networks. Journal of Theoretical Biology, 223(1):45–53, July 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303000717.

Wodarz:2004:ACR

[WS04] Dominik Wodarz and Akira Sasaki. Apparent competition and recovery from infection. Journal of Theoretical Biology, 227(3):403–412, April 7, 2004. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303004387.

Weibull:2006:NSS

[WS06] Jörgen W. Weibull and Marcus Salomonsson. Natural selection and social preferences. Journal of Theoretical Biology, 239(1):79–92, March 7, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305003164.

Wickramasuriya:2009:IHP

[WS09] A. Wickramasuriya and J. Schmitt. Improving horizontal plane locomotion via leg angle control. Journal of Theoretical Biology, 256(3):414–427, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004694.

Wick:2002:DST

[WSC02] David Wick, Steven G. Self, and Lawrence Corey. Do scarce targets or T killers control primary HIV infection? Journal of Theoretical Biology, 214(2):209–214, January 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924579.

Walter:2005:PSA

[WSM05] Nicole A. R. Walter, Justin Stebbing, and Walter Messier. The potential significance of adaptive evolution and dimerization in chimpanzee intercellular cell adhesion molecules (ICAMs). Journal of Theoretical Biology, 232(3):339–346, February 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930400390X.

Widder:2007:DPG

[WSS07] Stefanie Widder, Josef Schicho, and Peter Schuster. Dynamic patterns of gene regulation i: Simple two-gene systems. Journal of Theoretical Biology, 246(3):395–419, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307000197.

Wagner:2001:MSN

[WT01] Mirko Wagner and Jens Timmer. Model selection in nonnested hidden Markov models for ion channel gating. Journal of Theoretical Biology, 208(4):439–450, February 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300922306.

Wild:2005:KSA

[WT05] Geoff Wild and Peter D. Taylor. A kin-selection approach to the resolution of sex-ratio conflict between mates. Journal of Theoretical Biology, 236(2):126–136, September 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305000925.

Wild:2007:DLW

[WT07] Geoff Wild and Arne Traulsen. The different limits of weak selection and the evolutionary dynamics of finite populations. Journal of Theoretical Biology, 247(2):382–390, July 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307001336.

Wang:2009:APP

[WT09a] Wendi Wang and Yasuhiro Takeuchi. Adaptation of prey and predators between patches. Journal of Theoretical Biology, 258(4):603–613, June 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000770.

Woelfing:2009:SSI

[WT09b] Benno Woelfing and Arne Traulsen. Stochastic sampling of interaction partners versus deterministic payoff assignment. Journal of Theoretical Biology, 257(4):689–695, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308006711.

Tam:2009:CRS

[wTA09] Tze wai Tam and Put O. Ang. Catastrophic regime shifts in coral communities exposed to physical disturbances: Simulation results from object-oriented 3-dimensional coral reef model. Journal of Theoretical Biology, 259(2):193–208, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309001234.

Warren:2009:MMC

[WTC09] N. J. Warren, M. H. Tawhai, and E. J. Crampin. A mathematical model of calcium-induced fluid secretion in airway epithelium. Journal of Theoretical Biology, 259(4):837–849, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309002008.

Wang:2008:IVE

[WTL08] Hsiuying Wang, Yun-Huei Tzeng, and Wen-Hsiung Li. Improved variance estimators for one- and two-parameter models of nucleotide substitution. Journal of Theoretical Biology, 254(1):164–167, September 7, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002178.

Wang:2006:PPS

[WTSN06] Wendi Wang, Yasuhiro Takeuchi, Yasuhisa Saito, and Shinji Nakaoka. Prey-predator system with parental care for predators. Journal of Theoretical Biology, 241(3):451–458, August 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305005400.

Wolf:2005:DTA

[WVA05a] Denise M. Wolf, Vijay V. Vazirani, and Adam P. Arkin. Diversity in times of adversity: probabilistic strategies in microbial survival games. Journal of Theoretical Biology, 234(2):227– 253, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005740.

Wolf:2005:MMP

[WVA05b] Denise M. Wolf, Vijay V. Vazirani, and Adam P. Arkin. A microbial modified prisoner’s dilemma game: how frequencydependent selection can lead to random phase variation. Journal of Theoretical Biology, 234(2):255–262, May 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304005752.

Williams:2000:CSM

[WW00] E. D. Williams and G. T. Williams. Cancer, somatic mutation and stem cells. Journal of Theoretical Biology, 205(1):165– 166, July 7, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300920502.

Welch:2005:NMP

[WW05a] John J. Welch and David Waxman. The nk model and population genetics. Journal of Theoretical Biology, 234(3):329–340, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005818.

White:2005:RCM

[WW05b] Steven M. White and K. A. Jane White. Relating coupled map lattices to integro-difference equations: dispersaldriven instabilities in coupled map lattices. Journal of Theoretical Biology, 235(4):463–475, August 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000536.

Willadsen:2007:RSS

[WW07] Kai Willadsen and Janet Wiles. Robustness and state-space structure of Boolean gene regulatory models. Journal of Theoretical Biology, 249(4):749–765, December 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004237.

Welch:2008:CIC

[WW08] John J. Welch and David Waxman. Calculating independent contrasts for the comparative study of substitution rates. Journal of Theoretical Biology, 251(4):667–678, April 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006522.

Walker:2006:IAM

[WWS^{+}06] Dawn Walker, Steven Wood, Jennifer Southgate, Mike Holcombe, and Rodney Smallwood. An integrated agentmathematical model of the effect of intercellular signalling via the epidermal growth factor receptor on cell proliferation. Journal of Theoretical Biology, 242(3):774–789, October 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306001706.

Wang:2008:DHM

[WWS08] Kaifa Wang, Wendi Wang, and Shiping Song. Dynamics of an HBV model with diffusion and delay. Journal of Theoretical Biology, 253(1):36–44, July 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005668.

Wang:2006:USG

[WYC06] Shuang-Quan Wang, Jie Yang, and Kuo-Chen Chou. Using stacked generalization to predict membrane protein types based on pseudo-amino acid composition. Journal of Theoretical Biology, 242(4):941–946, October 21, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001901.

Wang:2005:SPM

[WYXC05] Meng Wang, Jie Yang, Zhi-Jie Xu, and Kuo-Chen Chou. SLLE for predicting membrane protein types. Journal of Theoretical Biology, 232(1):7–15, January 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003455.

Wang:2008:CPS

[WZ08] Jun Wang and Xiaoqi Zheng. Comparison of protein secondary structures based on backbone dihedral angles. Journal of Theoretical Biology, 250(2):382–387, January 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307005103.

Wang:2000:CCS

[WZL00] Zong-Ling Wang, Jing-Guang Zhang, and Xin-Min Liu. Competition and coexistence in spatially subdivided habitats. Journal of Theoretical Biology, 205(4):631–639, August 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519300920964.

Wang:2008:EGD

[WZL^{+}08] Shichang Wang, Boyu Zhang, Zhenqing Li, Ross Cressman, and Yi Tao. Evolutionary game dynamics with impulsive effects. Journal of Theoretical Biology, 254(2):384–389, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002518.

Wei:2007:MIS

[WZW^{+}07] Huachun Wei, Rui Zhang, Chunfang Wang, Huiqin Zheng, Aixiu Li, Kuo-Chen Chou, and Dong-Qing Wei. Molecular insights of SAH enzyme catalysis and implication for inhibitor design. Journal of Theoretical Biology, 244(4):692–702, February 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004127.

Xu:2007:SMI

[XAP07] Yaji Xu, Linda J. S. Allen, and Alan S. Perelson. Stochastic model of an influenza epidemic with drug resistance. Journal of Theoretical Biology, 248(1):179–193, September 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002299.

Xiao:2005:UDS

[XBCF05] Yanni Xiao, Roger G. Bowers, Damian Clancy, and Nigel P. French. Understanding the dynamics of salmonella infections in dairy herds: a modelling approach. Journal of Theoretical Biology, 233(2):159–175, March 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004710.

Xu:2005:FSS

[XBGN05] Cailin Xu, Mark S. Boyce, Madhav Gadgil, and Vidyanand Nanjundiah. Forecasting spatially structured populations: the role of dispersal and scale. Journal of Theoretical Biology, 233(2):177–189, March 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304004722.

Xiao:2009:EDH

[XBT09] Yanni Xiao, Roger G. Bowers, and Sanyi Tang. The effect of delayed host self-regulation on host-pathogen population cycles in forest insects. Journal of Theoretical Biology, 258(2):240–249, May 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309000526.

Xu:2002:SRN

[XD02] Tian Hong Xu and Ke Jing Deng. Sex and retrotransposons: a new approach to the problem. Journal of Theoretical Biology, 218(2):259–260, September 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930742.

Xu:2006:SSM

[XFAS06] Dashun Xu, Zhilan Feng, Linda J. S. Allen, and Robert K. Swihart. A spatially structured metapopulation model with patch dynamics. Journal of Theoretical Biology, 239(4):469– 481, April 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003528.

Xiao:2007:PAI

[XFBC07] Yanni Xiao, Nigel P. French, Roger G. Bowers, and Damian Clancy. Pair approximations and the inclusion of indirect transmission: Theory and application to between farm transmission of salmonella. Journal of Theoretical Biology, 244(3):532–540, February 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306003912.

Xie:2007:MRM

[Xie07] Ping Xie. Model for RuvAB-mediated branch migration of Holliday junctions. Journal of Theoretical Biology, 249(3):566–573, December 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307004134.

Xie:2009:PMD

[Xie09] Ping Xie. A possible mechanism for the dynamics of transition between polymerase and exonuclease sites in a high-fidelity DNA polymerase. Journal of Theoretical Biology, 259(3):434–439, August 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001659.

Xie:2007:MCR

[XK07] Z. Xie and D. Kulasiri. Modelling of circadian rhythms in drosophila incorporating the interlocked PER/TIM and VRI/PDP1 feedback loops. Journal of Theoretical Biology, 245(2):290–304, March 21, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306005182.

Xu:2009:HMD

[XM09] Rui Xu and Zhien Ma. An HBV model with diffusion and time delay. Journal of Theoretical Biology, 257(3):499–509, April 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309000058.

Xiao:2005:AGC

[XSD^{+}05] Xuan Xiao, Shihuang Shao, Yongsheng Ding, Zhengde Huang, Xiaojing Chen, and Kuo-Chen Chou. An application of gene comparative image for predicting the effect on replication ratio by HBV virus gene missense mutation. Journal of Theoretical Biology, 235(4):555–565, August 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000688.

Xu:2006:ENF

[XT06] Bing-Liang Xu and Yi Tao. External noise and feedback regulation: Steady-state statistics of auto-regulatory genetic network. Journal of Theoretical Biology, 243(2):214–221, November 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306002293.

Xiao:2008:PPS

[XWC08] Xuan Xiao, Pu Wang, and Kuo-Chen Chou. Predicting protein structural classes with pseudo amino acid composition: An approach using geometric moments of cellular automaton image. Journal of Theoretical Biology, 254(3):691–696, October 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308003263.

Xing:2001:ADS

[XWD^{+}01] Eric Poe Xing, Denise M. Wolf, Inna Dubchak, Sylvia Spengler, Manfred Zorn, Ilya Muchnik, and Casimir Kulikowski. Automatic discovery of sub-molecular sequence domains in multi-aligned sequences: a dynamic programming algorithm for multiple alignment segmentation. Journal of Theoretical Biology, 212(2):129–139, September 21, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301923197.

Yagil:2009:SCD

[Yag09] G. Yagil. The structural complexity of DNA templates — implications on cellular complexity. Journal of Theoretical Biology, 259(3):621–627, August 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309001611.

Yu:2004:CGR

[YAL04] Zu-Guo Yu, Vo Anh, and Ka-Sing Lau. Chaos game representation of protein sequences based on the detailed HP model and their multifractal and correlation analyses. Journal of Theoretical Biology, 226(3):341–348, February 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303003564.

Ylikarjula:2000:EPN

[YALT00] Janica Ylikarjula, Susanna Alaja, Jouni Laakso, and David Tesar. Effects of patch number and dispersal patterns on population dynamics and synchrony. Journal of Theoretical Biology, 207(3):377–387, December 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921817.

Yamauchi:2003:FAB

[Yam03] Atsushi Yamauchi. Factors affecting binary sex evolution with respect to avoidance of vertical transmission of deleterious intracellular parasites. Journal of Theoretical Biology, 222(4):505–515, June 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000651.

Yaniv:2005:API

[Yan05] Osnat Yaniv. An asymmetric parental investment conflict with continuous strategy sets. Journal of Theoretical Biology, 237(3):245–256, December 7, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001761.

Yaniv:2006:PME

[Yan06] Osnat Yaniv. Properties of a mixed ESS candidate in continuous strategy sets. Journal of Theoretical Biology, 238(4):795–804, February 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305002894.

Yan:2008:SRL

[Yan08] Ping Yan. Separate roles of the latent and infectious periods in shaping the relation between the basic reproduction number and the intrinsic growth rate of infectious disease outbreaks. Journal of Theoretical Biology, 251(2):238–252, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930700598X.

Yang:2009:MMH

[Yan09] Jie Yang. Molecular modeling of human BAD and its interaction with PKAc or PP1c. Journal of Theoretical Biology, 257(1):159–169, March 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006012.

Yates:2005:SMD

[YB05] Gillian E. Yates and Mark Broom. A stochastic model of the distribution of unequal competitors between resource patches. Journal of Theoretical Biology, 237(3):227–237, December 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305001748.

Yates:2007:SMK

[YB07] G. E. Yates and M. Broom. Stochastic models of kleptoparasitism. Journal of Theoretical Biology, 248(3):480–489, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002305.

Yates:2000:CMR

[YBV^{+}00] Andrew Yates, Claudia Bergmann, J. Leo Van Hemmen, Jaroslav Stark, and Robin Callard. Cytokine-modulated regulation of helper T cell populations. Journal of Theoretical Biology, 206(4):539–560, October 21, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300921477.

Yakubu:2002:IBL

[YCC02] Abdul-Aziz Yakubu and Carlos Castillo-Chavez. Interplay between local dynamics and dispersal in discrete-time metapopulation models. Journal of Theoretical Biology, 218(3):273–288, October 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302930754.

Yu:2006:PRR

[YCC^{+}06] Xiaojing Yu, Jianping Cao, Yudong Cai, Tieliu Shi, and Yixue Li. Predicting rRNA-, RNA-, and DNA-binding proteins from primary structure with support vector machines. Journal of Theoretical Biology, 240(2):175–184, May 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004212.

Yates:2004:CCS

[YCS04] Andrew Yates, Robin Callard, and Jaroslav Stark. Combining cytokine signalling with T-bet and GATA-3 regulation in Th1 and Th2 differentiation: a model for cellular decision-making. Journal of Theoretical Biology, 231(2):181–196, November 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304002863.

Yee:2009:MAS

[YD09] Thomas W. Yee and Thomas Dirnböck. Models for analysing species’ presence/absence data at two time points. Journal of Theoretical Biology, 259(4):684–694, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309002276.

Yan:2005:DDE

[YDFQ05] Xiting Yan, Minghua Deng, Wing K. Fung, and Minping Qian. Detecting differentially expressed genes by relative entropy. Journal of Theoretical Biology, 234(3):395–402, June 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304005995.

Yang:2006:AFA

[YDL06a] Jie Yang, Xian-Chi Dong, and Yao Leng. Application of FTTP to alpha-helix or beta-strand motifs. Journal of Theoretical Biology, 242(1):199–219, September 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306000798.

Yang:2006:CBA

[YDL06b] Jie Yang, Xian-Chi Dong, and Yao Leng. Conformation biases of amino acids based on tripeptide microenvironment from PDB database. Journal of Theoretical Biology, 240(3):374– 384, June 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305004297.

Yamamura:2007:OPA

[YFH^{+}07] Norio Yamamura, Noboru Fujita, Motoyuki Hayashi, Yusuke Nakamura, and Atsushi Yamauchi. Optimal phenology of annual plants under grazing pressure. Journal of Theoretical Biology, 246(3):530–537, June 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307000252.

Yoshida:2005:SIC

[YFK05] Hiroshi Yoshida, Chikara Furusawa, and Kunihiko Kaneko. Selection of initial conditions for recursive production of multicellular organisms. Journal of Theoretical Biology, 233(4):501– 514, April 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519304005065.

Yi:2003:MSV

[YFKP03] Chung-Seon Yi, Aaron L. Fogelson, James P. Keener, and Charles S. Peskin. A mathematical study of volume shifts and ionic concentration changes during ischemia and hypoxia. Journal of Theoretical Biology, 220(1):83–106, January 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931547.

Yamamura:2004:EMT

[YHBW04] Norio Yamamura, Masahiko Higashi, Narayan Behera, and Joe Yuichiro Wakano. Evolution of mutualism through spatial effects. Journal of Theoretical Biology, 226(4):421–428, February 21, 2004. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303003618.

Yearsley:2002:LPF

[YHG^{+}02] Jonathan Yearsley, Ian M. Hastings, Iain J. Gordon, Ilias Kyriazakis, and Andrew W. Illius. A lifetime perspective on foraging and mortality. Journal of Theoretical Biology, 215(4):385– 397, April 21, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925294.

Yokomizo:2003:CEA

[YHI03] Hiroyuki Yokomizo, Patsy Haccou, and Yoh Iwasa. Conservation effort and assessment of population size in fluctuating environments. Journal of Theoretical Biology, 224 (2):167–182, September 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303001449.

Yokomizo:2004:MYO

[YHI04] Hiroyuki Yokomizo, Patsy Haccou, and Yoh Iwasa. Multipleyear optimization of conservation effort and monitoring effort for a fluctuating population. Journal of Theoretical Biology, 230(2):157–171, September 21, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304001912.

Yokomizo:2007:OCS

[YHI07] Hiroyuki Yokomizo, Patsy Haccou, and Yoh Iwasa. Optimal conservation strategy in fluctuating environments with species interactions: Resource-enhancement of the native species versus extermination of the alien species. Journal of Theoretical Biology, 244(1):46–58, January 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306002505.

Yang:2006:LCD

[YHM^{+}06] Ling Yang, Zhangang Han, W. Robb MacLellan, James N. Weiss, and Zhilin Qu. Linking cell division to cell growth in a spatiotemporal model of the cell cycle. Journal of Theoretical Biology, 241(1):120–133, July 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305004923.

Yamauchi:2003:FSR

[YK03] Atsushi Yamauchi and Yukiko Kamite. Facultative sexual reproduction under frequency-dependent selection on a single locus. Journal of Theoretical Biology, 221(3):411–424, April 7, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930393195X.

Yearsley:2005:LHM

[YKG^{+}05] Jonathan M. Yearsley, Ilias Kyriazakis, Iain J. Gordon, Sarah L. Johnston, John R. Speakman, Bert J. Tolkamp, and Andrew W. Illius. A life history model of somatic damage associated with resource acquisition: damage protection or prevention? Journal of Theoretical Biology, 235(3):305–317, August 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305000275.

You:2009:CBT

[YKiP^{+}09] Sang Koo You, Dae Hyuk Kwon, Yong ik Park, Sun Myong Kim, Myung-Hoon Chung, and Chul Koo Kim. Collective behaviors of two-component swarms. Journal of Theoretical Biology, 261(3):494–500, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004056.

Yi:2000:FTN

[YL00] Tao Yi and Sabin Lessard. Fundamental theorem of natural selection and frequency-dependent selection: Analysis of the matrix game diploid model. Journal of Theoretical Biology, 206(1):17–25, September 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300920976.

Yang:2006:DSD

[YLCW06] Zhaochun Yang, Jeen-Shang Lin, Jianxin Chen, and James H-C. Wang. Determining substrate displacement and cell traction fields — a new approach. Journal of Theoretical Biology, 242(3):607–616, October 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306001883.

Yi:2000:ESS

[YLL00] Tao Yi, Sabin Lessard, and Mathieu Lemire. Evolutionarily stable strategy in a sex- and frequency-dependent selection model. Journal of Theoretical Biology, 204(2):191–200, May 21, 2000. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300910981.

Yang:2003:DRP

[YLM03] Qingwu Yang, Paul A. Lindahl, and Jeffrey J. Morgan. Dynamic responses of protein homeostatic regulatory mechanisms to perturbations from steady state. Journal of Theoretical Biology, 222(4):407–423, June 21, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303000523.

Yaniv:2004:PIC

[YM04] Osnat Yaniv and Uzi Motro. The parental investment conflict in continuous time: St. Peter’s fish as an example. Journal of Theoretical Biology, 228(3):377–388, June 7, 2004. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304000499.

Yaniv:2005:TDAa

[YM05a] Osnat Yaniv and Uzi Motro. Time-dependent animal conflicts: 1. The symmetric case. Journal of Theoretical Biology, 232(2):261–275, January 21, 2005. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519304003844.

Yaniv:2005:TDAb

[YM05b] Osnat Yaniv and Uzi Motro. Time-dependent animal conflicts: 2. The asymmetric case. Journal of Theoretical Biology, 232(2):277–284, January 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304003856.

Yan:2001:DRF

[YMC01] Bo Yan, Robert M. Miura, and Yi-Der Chen. Direction reversal of fluctuation-induced biased Brownian motion on distorted ratchets. Journal of Theoretical Biology, 210(2):141– 150, May 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519300922884.

Yan:2004:CPC

[YMLK04] Bin Yan, Barbara A. Methé, Derek R. Lovley, and Julia Krushkal. Computational prediction of conserved operons and phylogenetic footprinting of transcription regulatory elements in the metal-reducing bacterial family geobacteraceae. Journal of Theoretical Biology, 230(1):133–144, September 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519304001900.

Yoshiyama:2002:CTV

[YN02] Kohei Yoshiyama and Hisao Nakajima. Catastrophic transition in vertical distributions of phytoplankton: Alternative equilibria in a water column. Journal of Theoretical Biology, 216(4):397–408, June 21, 2002. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519302930079.

Yusim:2001:MPT

[YNL01] Karina Yusim, Shlomo Nir, and Noam Lahav. A model for proto-tRNA loading. Journal of Theoretical Biology, 208(1):109–116, January 7, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519300922045.

Yamauchi:2009:STC

[YNO09] Atsushi Yamauchi, Takaaki Nishida, and Takayuki Ohgushi. Stochastic tunneling in the colonization of mutualistic organisms: Primary succession by mycorrhizal plants. Journal of Theoretical Biology, 261(1):74–82, November 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003245.

Yoshigaki:2003:WDC

[Yos03] Tomoyoshi Yoshigaki. Why does a cleavage plane develop parallel to the spindle axis in conical sand dollar eggs? A key question for clarifying the mechanism of contractile ring positioning. Journal of Theoretical Biology, 221(2):229–244, March 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931869.

Yamaguchi:2007:DTM

[YOYT07] Sachi Yamaguchi, Yuki Ozaki, Yoichi Yusa, and Satoshi Takahashi. Do tiny males grow up? Sperm competition and optimal resource allocation schedule of dwarf males of barnacles. Journal of Theoretical Biology, 245(2):319–328, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930600484X.

Yang:2009:PPS

[YPY^{+}09] Jian-Yi Yang, Zhen-Ling Peng, Zu-Guo Yu, Rui-Jie Zhang, Vo Anh, and Desheng Wang. Prediction of protein structural classes by recurrence quantification analysis based on chaos game representation. Journal of Theoretical Biology, 257(4):618–626, April 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006723.

Yates:2007:LPI

[YS07] George T. Yates and Thomas Smotzer. On the lag phase and initial decline of microbial growth curves. Journal of Theoretical Biology, 244(3):511–517, February 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306003894.

Yu:2009:TCN

[YSW09] Jia-Feng Yu, Xiao Sun, and Ji-Hua Wang. TN curve: a novel 3d graphical representation of DNA sequence based on trinucleotides and its applications. Journal of Theoretical Biology, 261(3):459–468, December 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900366X.

Yeates:2001:UMD

[YTAK01] M. P. Yeates, B. J. Tolkamp, D. J. Allcroft, and I. Kyriazakis. The use of mixed distribution models to determine bout criteria for analysis of animal behaviour. Journal of Theoretical Biology, 213(3):413–425, December 7, 2001. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519301924257.

Yang:2008:RPH

[YTL08] Yuchen Yang, Erwin Tantoso, and Kuo-Bin Li. Remote protein homology detection using recurrence quantification analysis and amino acid physicochemical properties. Journal of Theoretical Biology, 252(1):145–154, May 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000453.

Yunes:2005:EHM

[Yun05] Rosendo A. Yunes. The evolution of the human mind and logic-mathematics structures. Journal of Theoretical Biology, 236(1):95–110, September 7, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305000809.

Yamauchi:2006:PCS

[YY06] Atsushi Yamauchi and Norio Yamamura. Persistence conditions of symmetric social hybridogenesis in haplo-diploid hymenoptera. Journal of Theoretical Biology, 240(1):24–31, May 7, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519305003619.

Yin:2007:PPC

[YY07] Changchuan Yin and Stephen S.-T. Yau. Prediction of protein coding regions by the 3-base periodicity analysis of a DNA sequence. Journal of Theoretical Biology, 247(4):687–694, August 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307001543.

Yang:2009:TFN

[YYA09] Eric Yang, Martin L. Yarmush, and Ioannis P. Androulakis. Transcription factor network reconstruction using the living cell array. Journal of Theoretical Biology, 256(3):393–407, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308005109.

Yokomizo:2003:OCE

[YYI03] Hiroyuki Yokomizo, Jin Yamashita, and Yoh Iwasa. Optimal conservation effort for a population in a stochastic environment. Journal of Theoretical Biology, 220(2):215–231, January 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519303931572.

Yamaguchi:2008:MGS

[YYY^{+}08] Sachi Yamaguchi, Yoichi Yusa, Shigeyuki Yamato, Satoru Urano, and Satoshi Takahashi. Mating group size and evolutionarily stable pattern of sexuality in barnacles. Journal of Theoretical Biology, 253(1):61–73, July 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308000441.

Yang:2006:NMA

[YZGW06] Huijie Yang, Fangcui Zhao, Jianzhong Gu, and Binghong Wang. Nonlinear modelling approach to human promoter sequences. Journal of Theoretical Biology, 241(4):765–773, August 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306000294.

Yu:2009:CVI

[YZHZ09] Tao Yu, Xiaoqin Zou, Sheng-You Huang, and Xian-Wu Zou. Cutoff variation induces different topological properties: a new discovery of amino acid network within protein. Journal of Theoretical Biology, 256(3):408–413, February 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005079.

Zwolak:2009:QME

[ZAB^{+}09] Jason Zwolak, Nassiba Adjerid, Elife Z. Bagci, John J. Tyson, and Jill C. Sible. A quantitative model of the effect of unreplicated DNA on cell cycle progression in frog egg extracts. Journal of Theoretical Biology, 260(1):110–120, September 7, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002434.

Zachos:2009:NCG

[Zac09] Louis G. Zachos. A new computational growth model for sea urchin skeletons. Journal of Theoretical Biology, 259(3):646–657, August 7, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001672.

Zhang:2007:DTD

[ZAD07] Le Zhang, Chaitanya A. Athale, and Thomas S. Deisboeck. Development of a three-dimensional multiscale agent-based tumor model: Simulating gene-protein interaction profiles, cell phenotypes and multicellular patterns in brain cancer. Journal of Theoretical Biology, 244(1):96–107, January 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002785.

Zavala:2007:AMS

[ZAdlPLM07] Miguel A. Zavala, Óscar Angulo, Rafael Bravo de la Parra, and Juan C. López-Marcos. An analytical model of stand dynamics as a function of tree growth, mortality and recruitment: The shade tolerance-stand structure hypothesis revisited. Journal of Theoretical Biology, 244(3):440–450, February 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004000.

Zahalak:2000:TSC

[Zah00] George I. Zahalak. The two-state cross-bridge model of muscle is an asymptotic limit of multi-state models. Journal of Theoretical Biology, 204(1):67–82, May 7, 2000. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519300910841.

Zamir:2001:FDM

[Zam01] M. Zamir. Fractal dimensions and multifractility in vascular branching. Journal of Theoretical Biology, 212(2):183– 190, September 21, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519301923677.

Zamir:2003:CTM

[Zam03] M. Zamir. Critique of the test of multifractality as applied to biological data. Journal of Theoretical Biology, 225(3):407–412, December 7, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303002613.

Zhang:2006:IDS

[ZC06] Yusen Zhang and Wei Chen. Invariants of DNA sequences based on 2DD-curves. Journal of Theoretical Biology, 242(2):382–388, September 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519306001123.

Zhao:2009:DAL

[ZC09] Zhong Zhao and Lansun Chen. Dynamic analysis of lactic acid fermentation in membrane bioreactor. Journal of Theoretical Biology, 257(2):270–278, March 21, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308006024.

Zioupos:2001:TFB

[ZCC01] Peter Zioupos, John D. Currey, and Adrià Casinos. Tensile fatigue in bone: Are cycles-, or time to failure, or both, important? Journal of Theoretical Biology, 210(3):389–399, June 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519301923161.

Zhou:2005:EDH

[ZCD05] Tong Zhou, J. M. Carlson, and John Doyle. Evolutionary dynamics and highly optimized tolerance. Journal of Theoretical Biology, 236(4):438–447, October 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305001323.

Zhou:2007:UCA

[ZCLZ07] Xi-Bin Zhou, Chao Chen, Zhan-Chao Li, and Xiao-Yong Zou. Using Chou’s amphiphilic pseudo-amino acid composition and support vector machine for prediction of enzyme subfamily classes. Journal of Theoretical Biology, 248(3):546–551, October 7, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307002809.

Zhang:2008:PPS

[ZDC08] Tong-Liang Zhang, Yong-Sheng Ding, and Kuo-Chen Chou. Prediction protein structural classes with pseudo-amino acid composition: Approximate entropy and hydrophobicity pattern. Journal of Theoretical Biology, 250(1):186–193, January 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519307004420.

Zhu:2006:SGC

[ZF06] Wen Zhu and Stephen Freeland. The standard genetic code enhances adaptive evolution of proteins. Journal of Theoretical Biology, 239(1):63–70, March 7, 2006. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305003140.

Zhang:2008:PCO

[ZF08] Guang-Ya Zhang and Bai-Shan Fang. Predicting the cofactors of oxidoreductases based on amino acid composition distribution and Chou’s amphiphilic pseudo-amino acid composition. Journal of Theoretical Biology, 253(2):310–315, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001288.

Zacarias:2005:PTR

[ZFVH05] G. D. Zacarias, C. P. Ferreira, and J. X. Velasco-Hernández. Porosity and tortuosity relations as revealed by a mathematical model of biofilm structure. Journal of Theoretical Biology, 233(2):245–251, March 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004771.

Zohdi:2004:PMA

[ZHB04] T. I. Zohdi, G. A. Holzapfel, and S. A. Berger. A phenomenological model for atherosclerotic plaque growth and rupture. Journal of Theoretical Biology, 227(3):437–443, April 7, 2004. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519303004478.

Zhdanov:2003:PIP

[Zhd03] Vladimir P. Zhdanov. Propagation of infection and the prey-predator interplay. Journal of Theoretical Biology, 225 (4):489–492, December 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519303002911.

Zhivotovsky:2002:MEE

[Zhi02] Lev A. Zhivotovsky. A model of the early evolution of soma-to-germline feedback. Journal of Theoretical Biology, 216(1):51– 57, May 7, 2002. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519302925336.

Zhu:2009:APP

[Zhu09] Hao Zhu. Is anisotropic propagation of polarized molecular distribution the common mechanism of swirling patterns of planar cell polarization? Journal of Theoretical Biology, 256(3):315–325, February 7, 2009. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308004621.

Zeineddine:2005:ESC

[ZJ05] Mohammed Zeineddine and Vincent A. A. Jansen. The evolution of stability in a competitive system. Journal of Theoretical Biology, 236(2):208–215, September 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305000937.

Zhang:2006:IEH

[ZJ06] Da-Yong Zhang and Xin-Hua Jiang. Interactive effects of habitat productivity and herbivore pressure on the evolution of anti-herbivore defense in invasive plant populations. Journal of Theoretical Biology, 242(4):935–940, October 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519306002141.

Zajac:2003:SCE

[ZJG03] Mark Zajac, Gerald L. Jones, and James A. Glazier. Simulating convergent extension by way of anisotropic differential adhesion. Journal of Theoretical Biology, 222(2):247–259, May 21, 2003. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930300033X.

Zhou:2008:PTI

[ZKFW08] Yicang Zhou, Kamran Khan, Zhilan Feng, and Jianhong Wu. Projection of tuberculosis incidence with increasing immigration trends. Journal of Theoretical Biology, 254(2):215–228, September 21, 2008. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308002713.

Zheng:2005:BPI

[ZKH^{+}05] M. Zheng, Y. Kashimori, O. Hoshino, K. Fujita, and T. Kambara. Behavior pattern (innate action) of individuals in fish schools generating efficient collective evasion from predation. Journal of Theoretical Biology, 235(2):153–167, July 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519305000056.

Zhang:2009:NRA

[ZLLZ09] Li Zhang, Bo Liao, Dachao Li, and Wen Zhu. A novel representation for apoptosis protein subcellular localization prediction using support vector machine. Journal of Theoretical Biology, 259(2):361–365, July 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309001349.

Zhang:2007:CAB

[ZLN07] Yong-Tao Zhang, Arthur D. Lander, and Qing Nie. Computational analysis of BMP gradients in dorsal-ventral patterning of the zebrafish embryo. Journal of Theoretical Biology, 248(4):579–589, October 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519307002688.

Zhao:2008:NGB

[ZLXY08] Hongya Zhao, Alan Wee-Chung Liew, Xudong Xie, and Hong Yan. A new geometric biclustering algorithm based on the Hough transform for analysis of large-scale microarray data. Journal of Theoretical Biology, 251(2):264–274, March 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307006017.

Zhu:2003:MMC

[ZM03] H. X. Zhu and J. R. Melrose. A mechanics model for the compression of plant and vegetative tissues. Journal of Theoretical Biology, 221(1):89–101, March 7, 2003. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519303931730.

Zhi:2001:MBS

[ZMQX01] Wang Zhi, Zhao Ming, and Yu Qi-Xing. Modeling of branching structures of plants. Journal of Theoretical Biology, 209(4):383–394, April 21, 2001. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519301922520.

Zou:2008:MSY

[ZPP08] Xiufen Zou, Tao Peng, and Zishu Pan. Modeling specificity in the yeast MAPK signaling networks. Journal of Theoretical Biology, 250(1):139–155, January 7, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004535.

Zhu:2007:SGR

[ZRSK07] Rui Zhu, Andre S. Ribeiro, Dennis Salahub, and Stuart A. Kauffman. Studying genetic regulatory networks at the molecular level: Delayed reaction stochastic models. Journal of Theoretical Biology, 246(4):725–745, June 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307000513.

Zhang:2007:EII

[ZSF^{+}07] Pei Zhang, Gregory J. Sandland, Zhilan Feng, Dashun Xu, and Dennis J. Minchella. Evolutionary implications for interactions between multiple strains of host and parasite. Journal of Theoretical Biology, 248(2):225–240, September 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519307002238.

Zammataro:2007:ECM

[ZSRB07] Luca Zammataro, Guido Serini, Todd Rowland, and Federico Bussolino. Embryonic cleavage modeling as a computational approach to sphere packing problem. Journal of Theoretical Biology, 245(1):77–82, March 7, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519306004565.

Zintzaras:2002:LUC

[ZSS02] Elias Zintzaras, Mauro Santos, and Eörs Szathmáry. “living” under the challenge of information decay: The stochastic corrector model vs. hypercycles. Journal of Theoretical Biology, 217(2):167–181, July 21, 2002. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519302930262.

Zhang:2001:NAP

[ZSZ01] Ziding Zhang, Zhi-Rong Sun, and Chun-Ting Zhang. A new approach to predict the helix/strand content of globular proteins. Journal of Theoretical Biology, 208(1):65–78, January 7, 2001. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930092201X.

Zhang:2006:GWA

[ZSZ^{+}06] Zhihua Zhang, Hong Sun, Yong Zhang, Yi Zhao, Baochen Shi, Shiwei Sun, Hongchao Lu, Dongbo Bu, Lunjiang Ling, and Runsheng Chen. Genome-wide analysis of mammalian DNA segment fusion/fission. Journal of Theoretical Biology, 240(2):200–208, May 21, 2006. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519305004194.

Zurakowski:2006:MPC

[ZT06] Ryan Zurakowski and Andrew R. Teel. A model predictive control based scheduling method for HIV therapy. Journal of Theoretical Biology, 238(2):368–382, January 21, 2006. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519305002067.

Zhang:2008:SEM

[ZT08] Tailei Zhang and Zhidong Teng. An SIRVS epidemic model with pulse vaccination strategy. Journal of Theoretical Biology, 250(2):375–381, January 21, 2008. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307004614.

Zhang:2009:MCT

[ZTKH09] Henggui Zhang, Tao Tao, Sanjay Kharche, and Simon M. Harrison. Modelling changes in transmural propagation and susceptibility to arrhythmia induced by volatile anaesthetics in ventricular tissue. Journal of Theoretical Biology, 257(2):279–291, March 21, 2009. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519308006358.

Zheng:2003:PIL

[ZW03] Wei-Mou Zheng and Fang Wu. In-phase implies large likelihood for independent codon model: distinguishing coding from non-coding sequences. Journal of Theoretical Biology, 223(2):199–203, July 21, 2003. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519303000869.

Zurakowski:2007:MDA

[ZW07] Ryan Zurakowski and Dominik Wodarz. Model-driven approaches for in vitro combination therapy using ONYX-015 replicating oncolytic adenovirus. Journal of Theoretical Biology, 245(1):1–8, March 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930600453X.

Zu:2008:EUS

[ZWT^{+}08] Jian Zu, Wendi Wang, Yasuhiro Takeuchi, Bo Zu, and Kaifa Wang. On evolution under symmetric and asymmetric competitions. Journal of Theoretical Biology, 254(2): 239–251, September 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251930800297X.

Zhang:2008:MVC

[ZXWF08] Chiyu Zhang, Shungao Xu, Ji-Fu Wei, and Donald R. Forsdyke. Microsatellites that violate Chargaff’s second parity rule have base order-dependent asymmetries in the folding energies of complementary DNA strands and may not drive speciation. Journal of Theoretical Biology, 254(1):168–177, September 7, 2008. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308002646.

Zhou:2005:FMD

[ZYD^{+}05] Li-Qian Zhou, Zu-Guo Yu, Ji-Qing Deng, Vo Anh, and Shun-Chao Long. A fractal method to distinguish coding and non-coding sequences in a complete genome based on a number sequence representation. Journal of Theoretical Biology, 232(4):559–567, February 21, 2005. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519304004230.

Zhao:2007:CHP

[ZYW07] Fangcui Zhao, Huijie Yang, and Binghong Wang. Complexities of human promoter sequences. Journal of Theoretical Biology, 247(4):645–649, August 21, 2007. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519307001567.

Zhang:2007:SAT

[ZZLT07] Boyu Zhang, Zhibin Zhang, Zhenqing Li, and Yi Tao. Stability analysis of a two-species model with transitions between population interactions. Journal of Theoretical Biology, 248(1):145–153, September 7, 2007. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930700224X.

Zhang:2007:MDS

[ZZW07] Yong Zhang, Xin Zhao, and Peng-Ye Wang. Molecular dynamics study of the fibril elongation of the prion protein fragment PrP106-126. Journal of Theoretical Biology, 245(2):238–242, March 21, 2007. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519306004942.